A Complete Bibliography of Publications in Algorithms

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

04 November 2025 Version 1.61

Title word cross-reference

(1 + 1) [1378]. (8/3)pn^{3} [277]. (α, β) [1471]. (H, S) [1083]. (n^{2} − 1) [277]. (s, t) [1665]. (σ, Z) [1745]. 0 [3183, 3549]. 0/1 [1944]. 0 < p < 1 [145]. 1 [892, 3183, 3549]. 2 [66, 1268, 1204, 1239, 892, 101, 259, 1165, 1191, 1358, 536, 731]. 3 [152, 683, 1269, 635, 1563, 1251, 1211, 1008, 852, 1361, 91, 606, 819]. 3/2 [361]. 4 [683]. ^{Æ} [3643]. ^{α} [2520]. ^{Cross−} [469]. _{2} [3081, 2661, 2339]. _{2.5} [3419, 2135, 2475]. _{3} [3211]. _{4} [3211]. _{6} [1642]. A [1909]. A∗ [3561, 575]. alpha [3831]. β [3636, 1361]. C [3438, 1100, 1506]. O [2415]. D [2850, 1709]. D∗ [3155]. ℓ^{1} [160]. ℓ^{p} [145]. ℓ_{1} [229]. ϵ [1576]. F^{1} [2211]. γ [28, 897]. h [73]. H_{∞} [343, 2131, 2203, 3120]. K [1382, 3192, 921, 1266, 687, 238, 1304, 1721, 2641, 3784, 2714, 890, 1380, 2038, 1034, 1744, 1313, 195, 1353, 1145, 727, 1124, 4, 541, 2125, 1656, 767, 2029, 2982]. k < 2 [1353]. kd [1844]. L^{p} [145]. L_{0} [751, 3473, 1918]. L_{1} [102, 104]. l_{2,1} [1555]. l_{p} [1592]. Λ [3453, 1423, 1622]. m [2360, 2968, 3244]. m_{λ} [443]. μ [2757, 2730]. N [46, 3464, 2648]. O(log_{2} N) [1599]. O(n) [98]. ω [2411]. p [1052, 1378]. Q [690, 2693, 3079]. r [2616]. T [3367]. t/k [2057]. U [2636]. X [974].

-alanine [1361]. -Analysis [2730]. -Anonymity [195]. -Approximation [361]. -Ary [2968]. -Back [2648]. -Based [2074]. -Bézier [3079]. -Center [1656]. -Circle [1313]. -Circulant [2411]. -Colored [1358]. -Component [1361]. -Connected [1269]. -Consensus [4]. -Contractive-Type [1745]. -Coverage [66]. -Cut [541]. -D [91]. -Diagnosability [2057]. -Distance-Hereditary [1353]. -Electrolyte [1361]. -epf [1313]. -Evolutionary [1378]. -Flip [2616]. -Grams [46]. -Graphs [897]. -Immediate [2125]. -Learning [690]. -Limited [594]. -Means [1034, 1744, 1382, 3192, 3438, 1266, 687, 1304, 1100, 1506, 1721, 2641, 3784, 2714, 890, 2038]. -Median [1378]. -Medoids [921]. -mer [1380]. -Model [1622]. -MSV [3636]. -Multiple-Knapsack [1944]. -Net [2636]. -NN [1124, 727]. -Nonexpansive [1471]. -Norm [1555, 1918]. -Operations [1423]. -OPT [1211]. -Optimal [2850]. -Order [1165]. -Pareto [2029]. -Partite [767]. -Paths [1665]. -Permutations [1423]. -Player [1268]. -Proper [1576]. -Puzzle [277]. -QAM [3244]. -Refined [1052]. -Refinement [73]. -Round [98]. -SAT [1251]. -Separation [1709]. -Stable [1909]. -Subgraph [541]. -Synthesis [2730]. -Tree [1844]. -Tuple [536]. -Value [2693]. -Width [974].

/Sliding [3120].

0.35 [2624, 2593].

1 [353, 499, 74]. 1-Safe [3689]. 1.5-Dimensional [3255]. 10 [455, 2801]. 11 [1810]. 12 [869]. 12-Lead [3021]. 13 [1334]. 15 [869]. 15.3 [3437]. 16-Criterion [3848]. 17 [455]. 185 [213]. 18F [3338]. 19 [1289, 2401, 1431, 1175, 2304, 2700, 1869, 1570, 3161, 2557, 2223, 1449, 1342]. 1D [3798]. 1D/2D [3798]. 1st [2036].

2 [3665, 390, 3444, 2632, 1598]. 2-5 [2871]. 2-Dimensional [3207]. 2-Phase [521]. 2.0 [1308]. 2.5D [1392]. 2015 [387]. 219 [1334]. 256 [1867]. 2D [3596, 3134, 2107, 3446, 2094, 3077, 3837, 2682, 3521, 3798, 2772, 2523]. 2D/3D [3837]. 2nd [3268].

3 [1810, 538]. 3-Dimensional [2287]. 3-RRR [538]. 3.5 [3116]. 305 [3437]. 309.- [74]. 3D [2791, 3761, 2815, 2007, 2254, 2429, 1337, 1847, 2920, 3355, 2284, 3837, 2492, 3424, 2905, 3064, 1985, 3635, 2657, 1347, 1693, 1862]. 3DCNN [2452]. 3PDP [3007]. 3SAT [136].

4 [3020]. 4.0 [2755]. 479.50.- [74]. 4WIDEV [3661].

5-Point [102]. 5G [1611, 3701, 2030]. 5G-Band [1611].

6-Approximation [3539]. 60th [1210]. 6D [3457]. 6G [2536, 1814]. 6G/B6G [2536].

7 [213].

87 [460].

9 [353, 460]. 978 [74]. 978-0-387-30770-1 [74].

A* [1775]. A2ES [2213]. AADE [3591]. AASA [3053]. Abandoned [3385]. ABC [816]. Abdominal [3338]. ABI [174]. Abilities [3122]. Ability [2724, 3675]. Abnormal [2040, 2518, 3063, 732, 1602]. Abnormalities [808, 3563, 3738]. Absolute [3829]. Absorption [1975]. Abstraction [2867]. Abuse [1720]. AC [3366, 3783]. AC-DC [3366]. Academic [236, 2742, 3470, 1763, 2708]. Accelerate [373, 3692, 2103, 1668]. Accelerated [2341, 2695, 1794, 3097, 686, 2291]. Accelerating [1910, 3793, 1610, 990, 1402, 1474]. Acceleration [3379, 1843, 2774, 314, 317]. Accelerator [1147, 1754, 681, 2193, 856]. Accelerators [1989]. Acceptability [703]. Acceptance [1080, 2024]. Accepting [3183]. Access [916, 139, 914, 1328, 3750, 121, 2951, 2882, 461, 3376, 3248]. Accessibility [280]. Accident [3447]. Accident-Induced [3447]. Accidents [2340, 1569]. Accidents-Artificial [2340]. According [2272]. Account [2658, 2325]. Accounting [1923, 1222, 1332]. Accounts [3083, 3920]. Accuracy [2375, 1935, 2583, 1435, 3076, 3722, 2759, 2641, 960, 2983, 3669]. Accurate [2864, 3599, 2904, 2883, 2785, 3216, 3253, 2014, 2353, 3667, 1098]. Accurately [3066]. ACE [742]. Achievable [2488]. Achievement [849]. Achieves [1773]. Achieving [2913]. Acknowledgement [972, 208, 246, 334, 417, 542, 765]. Acknowledgment [1648, 2146]. ACO [1613, 2120, 3903, 1195]. ACO-BmTSP [2120, 3903]. Acoustic [3100, 2266, 2220, 1984, 2523, 2300, 1013, 1200]. Acoustical [29]. Acquired [2785]. Acquiring [2589]. Acquisition [127, 2048, 306]. Across [3620, 1253]. Act [823, 3190]. ACT-FRCNN [3190]. Actinidia [3815]. Action [2628, 1329, 1652, 2694, 1214, 3246, 2452, 2717, 1111, 1286]. Action-Translator [2694]. Actionable [3558, 3595]. Actions [2934, 3645]. Activation [2705, 2644, 3350, 2523]. Activation-Based [2705]. Active [2973, 3846, 29, 1504, 1623, 1855, 543, 6, 2775, 3352, 953, 1762, 3818, 729, 3060]. Activities [1939, 26, 3362, 1540]. Activity [1493, 24, 2295, 141, 2761, 3462, 2113, 2887, 3110, 2160, 3239, 383, 2073, 3383]. Actor [1991, 1465, 453]. Actor-Based [1991]. Actor-Critic [453]. Actual [27]. Actuarial [2182]. Actuation [1785]. Actuator [1504, 3215, 543, 2853]. Acute [1939, 3533]. Acyclic [1585]. Ad [647, 2702]. Ad-Hoc [2702]. AdaBoost [3010, 3487]. AdaBoost-Mixed [3487]. Adam [3263, 3510, 2929, 2842]. Adaptation [1031, 840, 424, 3829, 478, 761, 2924, 2832, 3510, 800, 2098, 2362, 392, 383, 1226, 1726, 877, 3246]. Adapted [1205, 3587, 3388, 2032]. Adapters [1667]. Adaptive [3191, 1276, 3650, 1143, 1850, 1160, 2360, 1940, 3643, 783, 3072, 3804, 1138, 632, 1843, 1142, 127, 3781, 491, 3639, 1654, 746, 2347, 885, 976, 3260, 1161, 631, 922, 1557, 853, 3736, 1309, 409, 202, 366, 2745, 3750, 258, 687, 978, 2666, 602, 747, 2590, 3505, 1304, 1623, 1649, 3150, 1444, 1488, 1919, 3023, 3253, 2097, 1434, 3694, 448, 2371, 3726, 3890, 2534, 3349, 3103, 748, 2393, 3384, 637, 3708, 3112, 3406, 3170, 73, 451, 1419, 3257, 3577, 256, 1322, 3476, 845, 1487, 1758, 2216, 2714, 2976, 3053, 3586, 2075, 47, 2838, 2392, 2933, 3308, 2738, 3915, 1012, 3724, 1101]. Adaptive [988, 1512, 1742, 782, 239, 3487, 1712, 377, 1131, 1602, 1789, 3546, 3270, 718, 3921, 1604, 3570, 1424]. Adaptive-Size [922]. Adaptive-Step [3191]. Adaptivity [1055]. ADC [3908]. ADDAEIL [3611]. Adder [1446]. Adder/Subtractor [1446]. Adding [1892, 1011, 2731, 2372, 1379]. Adding-Doubling [2731]. Addition [3122]. Additional [1741]. Additive [3883, 2807, 2066]. Address [1739, 713, 3370]. Addressing [2395, 3289, 3895, 3520, 2265]. ADHD [3705]. Adil [1210]. Adjacencies [2646]. Adjacency [2367, 3803, 1671]. Adjoint [2063]. Adjustable [2334, 670, 1367]. Adjusting [3381]. Adjustment [3507, 1173, 3182, 3541]. Administrative [3162]. Admission [1289, 3237]. Admitted [1110]. ADMM [1639, 2347]. ADMs [1399]. Adopting [3231]. Adoption [2238]. ADRC [876, 3811]. Adsorption [3794, 3472]. Adults [3040, 356]. Advanced [3054, 2864, 3063, 3793, 633, 3495, 956, 3496, 2911, 3442, 3211, 3343, 3216, 2013, 1085, 1285, 3232, 1899, 1710]. Advancements [3435, 2840, 3497, 1128, 3073]. Advances [1646, 342, 3006, 3255, 59, 1291, 2735, 41, 3522]. Advancing [2486, 3357, 3396, 2818, 3345]. Advantages [1846, 3322, 3606]. Advection [1557]. Advection-Diffusion [1557]. Adversarial [1335, 1892, 2280, 3816, 2819, 3307, 1748, 3219, 2054, 2637, 2011, 3226, 2749, 2558, 1966, 1534, 1851, 3173, 2258, 2859, 674, 2078, 2079, 1882, 3300]. Adversaries [894]. Adverse [2760, 3185]. Advertising [2684, 1676, 768]. Advising [2742]. Aegean [3236]. Aerial [2356, 2127, 983, 2741, 2857, 3238, 1821, 3476, 3534, 2933, 2613, 1235]. Aerobatic [2467]. Aeroelastic [1735]. Aerosols [1608]. AetherGeo [3630]. AFE [2933]. AFE-YOLOv8 [2933]. Affecting [1899]. Affective [2213, 3765, 2499]. Affects [2086]. Affine [746, 122, 2923]. Affinity [369, 1300]. Affordable [3642]. Africa [3403]. African [803, 2999, 3726]. After [3810, 3040, 2819, 1821]. Aftersales [3442]. Again [1248]. Against [1018, 51, 1892, 2280, 1636, 1650, 3071]. Age [3508, 3390, 546, 3235, 1827]. Age-of-Information [1827]. Ageism [3307]. Agent [803, 714, 139, 1582, 3292, 3144, 823, 1457, 1869, 2678, 690, 3886, 3522, 2421, 3560, 1280, 3242, 3908, 3121, 2710, 1826, 814]. Agent-Based [139, 1869, 2678, 2421, 3560, 1280, 2710]. Agent-Deed-Consequence [3908]. Agentic [3797, 3751]. Agents [820, 3205, 595, 182, 3396, 3280].

Agglomerative [2342, 1378, 271, 1774, 1795]. Aggregate [1207, 1584, 1786, 2823, 2901]. Aggregated [2217, 949]. Aggregates [881, 2269]. Aggregation [3614, 1096, 2289, 2808, 640, 276, 2450, 1332, 560, 3529, 2192, 651, 3542]. Aggressiveness [888]. Aging [3761, 1782, 3543, 2417]. Agnostic [1004, 1573]. AGR [2515]. Agreement [2583]. Agreements [1029]. Agricultural [2394, 3840, 1241, 1692, 3223, 1531, 501]. Agriculture [2001, 3474]. AGV [784]. Ahead [2260, 1425, 2599, 501]. Ahmad [455]. AHP [594, 3572]. AHRS [378]. AI [1982, 1989, 3651, 2406, 3433, 2410, 3656, 3409, 3751, 2793, 3089, 3703, 3558, 3569, 3295, 3806, 3805, 2550, 3101, 3664, 3615, 3299, 1881, 2318, 3892, 3898, 3396, 3762, 1534, 3172, 3666, 3033, 3907, 3478, 2884, 2979, 3534, 2716, 3585, 3669, 3919, 3545]. AI-Based [3664, 1534, 3534]. AI-Driven [3805, 2716, 3545]. AI-Enhanced [3299, 3892]. AI-Enriched [3478]. AI-Powered [3907]. Aid [2327]. Aided [57, 1357, 3089, 3557, 965, 1464, 85, 316, 904, 53, 3126]. Aiding [2801, 875]. Aiming [1001]. Air [2013, 3515, 1324, 3895, 530, 2383, 2847]. Aircraft [3663, 3325, 3159, 3056, 187, 2853, 3584, 1762, 1849]. Airfoil [675]. Airline [1746]. Airport [2585, 187]. Airspace [2548, 1727]. Airstrips [3354]. AIS [730, 2719]. AKT [3797]. al [455, 1334, 3049, 869]. alanine [1361]. Alarm [235, 2555, 2874]. Alcohol [3699]. ALE [47]. ALE-PSO [47]. Alerting [3150]. Algebra [1694, 3779]. Algebra-Based [3779]. Algebraic [526, 1503, 3149, 1024, 2269, 1844]. Algorethics [3101]. Algorithm [3191, 3254, 1406, 151, 1017, 2614, 1933, 1990, 1126, 3379, 224, 166, 654, 3038, 3412, 3650, 1732, 692, 3493, 1360, 2618, 2395, 2695, 3850, 29, 1642, 1994, 2583, 1818, 1910, 363, 1995, 3705, 169, 3131, 201, 1552, 476, 3643, 1576, 424, 426, 512, 1463, 777, 610, 702, 788, 1429, 3596, 1232, 1264, 1159, 1201, 921, 2172, 1178, 3857, 3438, 879, 510, 799, 3356, 2211, 3612, 2127, 1554, 600, 2042, 514, 266, 522, 816, 787, 791, 2215, 3551, 1353, 3564, 1633, 746, 17, 374, 282, 506, 2774, 3087, 330, 956, 3689, 554, 2542, 1582, 2281, 3359, 3673]. Algorithm [2970, 489, 537, 1817, 723, 1509, 3693, 1161, 2457, 631, 521, 473, 2916, 1077, 3709, 695, 64, 3206, 2433, 2685, 2370, 2462, 2489, 1874, 529, 420, 567, 2628, 3792, 3499, 1237, 2850, 3208, 3561, 1806, 2799, 2999, 3128, 3010, 3795, 2545, 1669, 190, 1779, 1096, 2355, 2630, 366, 2326, 1051, 1619, 662, 48, 38, 3593, 1413, 2745, 3067, 3619, 778, 1314, 3740, 1185, 3210, 3410, 3446, 3501, 2167, 234, 258, 3093, 2412, 30, 1528, 67, 3871, 2313, 2435, 1692, 2666, 602, 747, 3503, 2290, 3505, 3787, 2128, 3873, 1209, 3448, 3301, 3783, 2377, 3098, 36, 2407, 2271, 3450, 3711, 321, 3875]. Algorithm [1304, 1649, 2709, 2806, 3150, 887, 2852, 3151, 2008, 3507, 1157, 2094, 1287, 1378, 1615, 269, 2690, 1100, 1506, 2942, 2153, 185, 2800, 2447, 2143, 35, 566, 3510, 1250, 360, 259, 3077, 3812, 2054, 2011, 2012, 1658, 1682, 222, 3365, 458, 1836, 913, 2491, 3156, 2784, 304, 257, 285, 268, 352, 401, 432, 448, 456, 743, 557, 655, 1173, 1278, 1516, 2014, 1961, 1775, 2205, 2726, 3617, 3332, 3884, 3256, 3514, 1568, 3515, 219, 422, 650, 3690, 45, 296, 397, 462, 736, 561, 547, 558, 828, 2923, 2868, 3416, 670, 808, 3889, 3002, 2357, 3516, 248, 2056, 2057, 3160, 2325, 327, 1802, 3622, 582, 226, 1188, 3517, 3624, 2133].

Algorithm [2995, 2382, 1041, 1116, 2556, 1215, 2794, 2240, 121, 1171, 1085, 3331, 134, 23, 3893, 2098, 193, 3228, 276, 331, 470, 265, 979, 419, 3164, 1367, 1573, 1030, 2198, 2731, 2131, 728, 2203, 2913, 3288, 3265, 3823, 1760, 2217, 2917, 1876, 2121, 2018, 1348, 722, 1279, 54, 1967, 3039, 2737, 3232, 3388, 2442, 902, 2703, 1683, 1897, 2598, 3464, 3654, 2483, 2450, 2560, 763, 2384, 398, 361, 3234, 3284, 3633, 609, 3784, 1900, 1761, 3354, 2061, 2062, 1535, 2768, 392, 475, 612, 2841, 855, 3406, 2816, 1605, 73, 2565, 2974, 3470, 1059, 1807, 1865, 647, 3526]. Algorithm [1078, 1152, 1507, 1284, 603, 1821, 3659, 2126, 2952, 1845, 3677, 1864, 2646, 891, 1715, 1448, 3772, 1459, 1322, 2567, 3475, 3240, 2772, 3476, 3477, 2308, 2316, 2287, 2983, 154, 3119, 1445, 5, 2220, 1441, 3317, 660, 538, 3242, 378, 2067, 3424, 3036, 2945, 1795, 938, 910, 3370, 1193, 2376, 827, 1038, 833, 1608, 1037, 644, 837, 1930, 2710, 346, 2929, 1478, 1597, 349, 3079, 187, 1094, 2374, 1470, 925, 1391, 2169, 1501, 2233, 731, 1320, 3053, 2480, 1647, 2982, 1405, 2402, 2954, 2075, 47, 1599, 2838, 1521, 642, 2475, 273, 262, 326, 515, 425, 878, 824]. Algorithm [805, 870, 1495, 1543, 1660, 2653, 2525, 2654, 2736, 3482, 3661, 3655, 3811, 3539, 613, 1130, 1650, 2163, 2738, 545, 838, 1544, 1985, 3183, 3841, 1197, 408, 477, 578, 217, 244, 807, 1393, 3724, 3818, 2802, 656, 1447, 307, 502, 1056, 988, 1512, 3541, 573, 348, 806, 2037, 691, 2487, 2822, 402, 1198, 1878, 104, 239, 2358, 3920, 3488, 3834, 3125, 3188, 658, 786, 3562, 309, 300, 496, 518, 697, 549, 575, 780, 873, 815, 752, 1131, 1492, 1547, 1789, 1916, 1793, 1728, 1627, 2082, 1713, 2307, 2847, 2991, 2877, 3046, 3395, 3270, 3652, 453, 415, 679, 769, 1422, 1421, 1737, 2829, 3921, 3922, 958, 2831, 1510, 1548, 3549, 255, 384]. Algorithm [729, 1988, 3251, 3550, 186, 703, 3399, 3285]. Algorithm-Based [1990, 3493, 3783, 2847]. Algorithmic [3697, 1902, 3637, 170, 2089, 885, 1400, 2002, 3255, 2934, 1588, 1282, 3016, 700, 3758, 3311, 909, 3895, 3580, 2122, 2151, 96, 943, 1979, 367, 100, 3770, 3835, 1229]. Algorithmics [871, 344]. Algorithms [1675, 1231, 3271, 353, 455, 1031, 2302, 3347, 1922, 1431, 1517, 2228, 734, 1455, 2811, 373, 3083, 1709, 2894, 2118, 390, 90, 971, 3198, 3411, 2104, 3853, 2968, 28, 1580, 2910, 556, 1945, 115, 1481, 2508, 3061, 753, 761, 236, 3599, 1262, 2512, 813, 624, 2762, 1398, 1386, 3336, 2229, 2879, 3495, 1081, 336, 1133, 945, 3497, 1582, 669, 2756, 2792, 400, 1027, 1928, 3089, 3498, 372, 3267, 3312, 457, 3204, 31, 3371, 37, 1504, 1583, 3757, 1088, 2977, 3283, 1343, 1810, 207, 460, 3078, 2453, 739, 3315, 854, 202, 2805, 868, 1953, 740, 3020, 3712]. Algorithms [323, 27, 3636, 3696, 1955, 3750, 2688, 2283, 2548, 570, 817, 3211, 418, 1635, 3289, 1559, 1652, 3449, 3649, 78, 213, 3019, 183, 2739, 446, 2830, 19, 1247, 951, 3043, 2591, 3044, 720, 1144, 3295, 2415, 2451, 3268, 1891, 576, 3153, 3219, 1663, 1490, 3030, 2052, 1861, 3058, 2272, 3794, 2778, 2595, 2055, 581, 3225, 1213, 279, 1252, 3615, 146, 794, 509, 736, 1786, 3158, 3426, 3726, 3282, 3845, 1838, 167, 1365, 272, 212, 297, 3104, 977, 121, 1285, 2329, 3011, 2284, 2640, 3106, 1189, 3302, 1423, 1170, 3029, 1591, 2893, 6, 1592, 1689, 2866, 3463, 2459, 3230].

Algorithms [1277, 1721, 2432, 3233, 2642, 18, 959, 2208, 3017, 2600, 2912, 494, 3606, 1008, 2963, 405, 684, 444, 2269, 1716, 2065, 3589, 153, 2734, 3056, 2027, 2918, 3073, 1898, 1875, 1334, 1089, 3580, 3049, 1538, 3048, 2494, 2028, 1539, 1749, 1616, 634, 2677, 1972, 869, 2366, 1472, 2349, 1668, 254, 3176, 81, 1734, 1914, 1132, 2068, 588, 1978, 3745, 16, 2191, 3245, 1257, 2354, 119, 2728, 798, 2848, 1005, 1854, 3534, 653, 2701, 2662, 589, 2300, 2608, 3124, 1370, 1671, 2610, 2650, 1449, 2707, 286, 365, 2530, 618, 952, 2036, 2351, 2828, 216, 1197, 3185, 555]. Algorithms [1870, 3544, 2499, 218, 377, 895, 1350, 856, 2758, 553, 44, 1203, 2596, 2041, 1601, 2722, 2619, 74, 972, 1648, 2146]. Algorithms-Based [2866]. Alignment [930, 46, 3859, 313, 2773, 3763, 3029, 2862, 3677, 2317]. All-Optical [1275]. All-Out [3179]. All-Pairs [3812]. Allen [1090]. Alleviate [3447]. Allocating [1664, 1467]. Allocation [1575, 3427, 3351, 1328, 234, 3027, 1586, 2809, 3880, 352, 3455, 3760, 558, 2024, 998, 1403, 3577, 2122, 1691, 837, 2848, 545, 496, 2702, 769]. Allocations [2058]. Allows [1598]. Alloys [967]. Alone [2495]. along [1887]. Alpha [151, 2467, 2022]. Alpha-Divergences [2022]. Alphabets [126, 868]. Alterations [2405]. Alternating [2503, 778, 2712, 1960, 351, 1033, 2083]. Alternative [3421, 1468, 2389, 1970, 2475, 3572]. Alternatives [113, 1801, 2809, 1310]. Althöfer [151]. Alzheimer [3415, 3686, 2302, 2864, 3309]. AMARA [2313]. AMARA-fMRI [2313]. Amateur [2113]. Amazon [3354, 1129, 2716]. Amazons [2991]. Ambiguities [1651]. American [1263, 3057]. Amidst [3157]. Among [3903, 3403, 2830]. AMOSA [959]. Amplifier [275]. Amplitude [928]. Amplitude-Aware [928]. AMR [2026]. AMR-Assisted [2026]. Anaerobic [3179]. Analog [2584, 1621, 557]. Analogical [2229]. Analyses [1630]. Analysis [2401, 1231, 2757, 2321, 2184, 2039, 3054, 3129, 3493, 490, 3433, 871, 2619, 3851, 388, 3797, 3197, 1938, 440, 531, 3718, 2990, 888, 3729, 2894, 3601, 2501, 1062, 905, 3, 1860, 2418, 1519, 2584, 3198, 3009, 3409, 2840, 2497, 3857, 2539, 3045, 799, 1842, 514, 3439, 191, 3564, 176, 2043, 3773, 1948, 3075, 1004, 1383, 229, 1499, 3498, 3024, 3382, 3394, 3139, 1803, 1103, 2898, 3092, 1558, 1105, 3866, 1895, 1343, 2373, 32, 1829, 3369, 2420, 854, 3441, 3735, 2814, 2546, 1873, 2234, 3840, 1104, 3761, 2436, 3052, 3006, 1606, 3263, 3636, 3296, 1955, 67, 2865]. Analysis [2741, 3095, 2769, 1209, 2378, 138, 2271, 3450, 1267, 2050, 2830, 967, 3744, 3830, 1903, 3381, 2592, 1490, 2783, 3794, 1483, 438, 1326, 429, 913, 3664, 403, 432, 1173, 1477, 1278, 1377, 3660, 3886, 1024, 1962, 1812, 462, 3282, 3746, 3355, 2164, 2218, 1931, 1571, 1838, 3573, 2259, 2639, 212, 1291, 771, 3103, 2794, 21, 3459, 2329, 3672, 3011, 276, 709, 2731, 1217, 2187, 1254, 3029, 1907, 601, 2721, 1348, 3676, 2730, 3298, 3898, 546, 2125, 2516, 1968, 801, 2483, 12, 2560, 614, 1899, 1808, 3047, 1009, 2746, 2186, 2674, 3523, 1716, 520, 2565, 1695, 3589, 2734, 1465, 1122]. Analysis [2939, 3073, 469, 1284, 79, 1222, 1121, 2126, 1971, 2122, 3116, 3772, 1972, 3630, 2413, 953, 3241, 378, 3403, 2068, 588, 2288, 1099, 3631, 591, 1711, 3532, 1856, 119, 368, 1163, 160, 2824, 3591, 2484, 3535, 2716, 1315, 1449, 3775, 2472, 1782, 1402, 326, 501, 1427, 1698, 2611, 3770, 2163, 2874, 3071, 2572, 2573, 260, 307, 332, 3309, 435, 104, 364, 3488, 3648, 2038, 3305, 525, 1763, 2176, 3126, 3570, 565, 1461, 44, 3572, 2409]. Analysis-Based [1326, 104]. Analytic [534, 1810, 2809, 2658, 321, 397, 2061, 2902, 1574, 594, 155]. Analytical [2401, 2090, 3502, 2247, 102, 3223, 3331, 1537, 1705, 2267, 3118]. Analytical-Numerical [3223]. Analytics [1002, 556, 1469, 3519, 1710, 3326, 2169]. Analytics-Driven [1002]. Analyzing [1709, 3496, 3301, 1144, 2027, 1896, 2484, 3250]. AnalyzIR [604]. Ancestral [181, 1408]. Anchor [3213, 2317]. Anchor-Based [3213]. and/or [3364, 2132, 3156, 2349]. Anderson [1617]. Anesthesia [637]. Aneurysm [3338]. Angiography [2635]. Angle [531, 2583, 1633, 828, 1802, 3488, 658]. Angle-Propagation [1802]. Angles [2791, 2974]. Angular [1633]. Animals [1647]. Animation [2987, 88]. Anisotropic [157]. ANN [1790, 3795, 3722, 3890, 2139, 2333]. Annealed [233]. Annealer [821]. Annealing [1164, 1435, 3228, 1152, 1507, 827, 3481]. Annotated [2924, 7]. Annotation [1102, 3485]. ANNs [1611]. Anomalies [1308, 3866, 2723, 2168, 3251]. Anomaly [1934, 2235, 2410, 3856, 2858, 1948, 1557, 3369, 2547, 3263, 2943, 3321, 3100, 3578, 3611, 148, 1605, 3049, 2771, 2278, 1856, 2102, 3314, 859, 2514, 2244]. Anonymity [195, 408]. Anonymization [1739, 2519, 935]. Anonymous [1718]. Ansatz [778]. Ansätzes [1960]. Answer [1805, 2262, 1438, 2299]. Answering [2935]. Ant [1523, 2720, 242, 48, 1242, 1692, 1623, 182, 3887, 547, 3517, 2098, 2640, 3823, 2737, 1533, 1095, 2120, 1467, 106, 1197, 697, 762, 3922]. Ant-Balanced [2120]. Ant-Like [182]. Ant-Lion [1623]. Antenna [2995, 2913, 365, 2738]. Antennas [3423, 1611, 2283, 365]. Anti [1262, 2398, 3530, 1762, 2400, 3547]. Anti-Interference [2398]. Anti-Lock [1262]. Anti-Ship [3530]. Anti-Skid [1762]. Anti-Spoofing [2400]. Anti-Sway [3547]. Antibiotics [10]. Antifungal [2904]. Antioxidants [3069]. Antiretroviral [3598]. Anxiety [2622]. Any [131]. Aortic [3338, 3533]. Apache [1014, 3035, 3326]. Aperture [115, 1820]. APF [3655]. APF-Dijkstra [3655]. API [1784]. APIT [522]. Apnea [3172]. Apollonian [2212]. AppArmor [3431]. Appearance [1022, 260]. Appearance-based [260]. Apple [1699]. Appliance [3737]. Applicability [291, 464, 454, 2649, 2610]. Application [3492, 3734, 1431, 1818, 915, 3063, 2584, 1201, 1178, 3856, 3061, 3612, 2127, 3200, 127, 816, 850, 1202, 1555, 1169, 176, 540, 1090, 1811, 31, 1784, 2140, 2911, 3304, 1558, 3795, 3264, 2047, 2048, 342, 717, 48, 1046, 1241, 3870, 2548, 3055, 747, 59, 2175, 138, 2194, 563, 3318, 3149, 3150, 1032, 2800, 2764, 3882, 2272, 1682, 3225, 913, 1173, 1962, 621, 3299, 2164, 1390, 3456, 1879, 1085, 1804, 1572, 288, 1367, 1298, 1964, 3676, 3555, 3168, 358, 1900, 1761, 3901, 1899, 684, 539, 2816, 2565, 468, 1292, 2027, 1336, 1507, 1222, 2522, 2646, 3666, 2029, 1972, 3477, 1701]. Application [2502, 1361, 3120, 1977, 1979, 1225, 3121, 1369, 1981, 1037, 676, 368, 33, 2300, 2982, 2385, 2074, 2838, 1130, 651, 2766, 217, 524, 3618, 3920, 1372, 658, 2860, 453, 2758, 3291, 2083]. Application-Guided [2584]. Applications [1675, 3642, 1498, 3839, 2937, 1922, 302, 2252, 3132, 971, 1678, 1052, 1138, 3409, 3751, 2740, 2242, 3858, 2508, 624, 3749, 2870, 1398, 1386, 3336, 2229, 1148, 3201, 3075, 3498, 2001, 1119, 1205, 3207, 2977, 3283, 2411, 1813, 3593, 3443, 1273, 2006, 1661, 3505, 2128, 2678, 2227, 1480, 832, 822, 3511, 2053, 2380, 2761, 2055, 2909, 429, 3706, 2553, 2995, 2516, 3654, 2561, 398, 3236, 2563, 1932, 2186, 2274, 3524, 1841, 2857, 3406, 1082, 3904, 3589, 1865, 2918, 3238, 1766, 3474, 2772, 3626, 2884, 943, 3403, 2191, 362, 367, 3814, 2976, 1370, 2206, 391, 870, 2351, 2833]. Applications [3834, 2157, 1793, 2397, 3743]. Applied [2406, 424, 1578, 278, 2476, 917, 3160, 1705, 1038, 2070, 1260, 2358]. Apply [2614]. Applying [1992, 2455, 24, 2921, 2112, 3235, 1644, 1280, 3908, 3070, 1449, 232, 118, 1101, 841, 1235]. Appointments [1997]. Apprenticeship [3768]. Approach [3192, 3028, 2973, 1934, 590, 1575, 3194, 2732, 3430, 3493, 3196, 1388, 2040, 3427, 1404, 1442, 3346, 820, 1744, 1261, 2622, 1860, 1699, 3804, 2422, 3037, 3363, 1061, 1581, 3859, 3860, 139, 442, 1202, 1180, 1639, 1624, 191, 225, 1998, 3773, 1072, 1167, 2443, 2984, 1383, 1357, 2000, 3498, 1628, 3816, 3205, 954, 443, 2851, 1271, 1768, 2340, 1503, 190, 586, 1873, 1458, 2099, 1813, 2496, 1392, 2167, 466, 1877, 3094, 2845, 892, 3147, 3097, 164, 2941, 1480, 1084, 882, 3557, 2094, 3214, 1959, 1091, 132, 3594, 738, 1663, 3454, 2533, 2013, 2909]. Approach [271, 308, 620, 701, 1863, 3660, 3157, 2095, 3581, 2311, 3161, 52, 3459, 2813, 314, 3164, 3895, 3330, 530, 2125, 2490, 2887, 2498, 2561, 2362, 2023, 614, 2421, 2160, 2768, 1401, 3007, 2353, 745, 2066, 1023, 1292, 2985, 2602, 3306, 2494, 3659, 1970, 3474, 2222, 665, 2469, 1691, 3117, 2897, 2680, 2928, 3906, 1733, 811, 1300, 1149, 2859, 1731, 173, 441, 2223, 3121, 918, 2513, 2478, 3608, 1904, 2747, 2609, 2076, 3181, 1676, 3345, 579, 606, 3770, 3353, 3293, 3538, 357, 992, 810, 1198, 733, 2320, 2655, 3767, 3737, 789, 2531, 2409, 1229]. Approaches [3579, 13, 3009, 2416, 2526, 192, 2005, 1825, 9, 1958, 437, 3016, 3746, 3311, 2558, 1772, 749, 2250, 2645, 1283, 2464, 1508, 536, 3534, 1729, 2168, 2760, 53, 3126, 1983]. Appropriate [399]. Approximability [159, 2953]. Approximate [1178, 1806, 2355, 778, 1114, 1682, 1452, 2995, 2492, 2837, 65, 402]. Approximated [2757, 2411]. Approximately [1429]. Approximating [323, 2915, 360, 3587, 111, 117]. Approximation [3340, 1021, 1150, 1148, 3423, 3647, 615, 1205, 1299, 1088, 1399, 202, 1242, 1207, 1584, 1316, 1794, 185, 576, 3831, 1682, 1252, 994, 979, 1423, 3108, 2442, 361, 214, 3902, 1673, 3798, 634, 1974, 3539, 1725, 3559, 1422, 2083]. Approximation-Based [3559]. Approximations [2572, 1849]. APT’s [3656]. Aquatic [35]. Aquifer [3081]. Arabic [2401, 3129, 350, 926, 3296, 1464, 3042]. Arabidopsis [2490]. Arbitrarily [2492]. Arbitrary [3689, 3077, 923, 994, 3118]. Arboreal [3774]. Arborescence [2669]. Arborescences [2627]. Arc [3749, 506, 1358, 307].

Arc-Completion [1358]. Arc-Weighted [506]. ArCAR [1464]. Archimedean [2668]. Architectural [3214, 429]. Architecture [3415, 1908, 2302, 1744, 2927, 3013, 3781, 2878, 893, 3868, 2631, 2942, 1302, 1385, 1003, 1109, 3894, 1967, 990, 3769, 1927, 3789, 1407, 2206, 3765, 3291]. Architecture-Based [2942]. Architectures [3849, 203, 1069, 3840, 196, 1288, 2559, 1599]. Archive [2037]. Arctic [3206]. Area [2834, 3419, 3297, 1635, 1186, 743, 1572, 884, 2660, 1465, 256, 2566, 665, 2890, 545]. Areas [957, 1362, 1879, 441, 1809, 3752]. Areas-of-Interest [441]. Arguments [1669]. ARIMA [3722, 3769, 399, 766, 792]. ARIMA-ANN [3722]. Arising [2109, 1620]. Arithmetic [2970, 97, 2022]. Arithmetics [3374]. Arkouda [1469]. Arm [3191, 3164, 2255, 852]. Arm-Based [2255]. Armed [544, 2316, 3308]. Armor [3431]. Arnold [3806, 3345]. Arrangements [3567]. Array [2774, 2995, 2738, 1230]. Array-Based [2774, 2995]. Arrays [1614, 1633, 2468, 2913, 129]. Arrest [3380]. Arrhythmia [1830, 3021, 776]. Arrhythmias [3342]. Arrival [3139, 3690, 384]. Art [2439, 193, 2484]. Artery [2776]. Article [3660]. Articulated [799, 553]. Artifact [2313, 899]. Artifacts [314, 3017]. Artificial [2614, 2341, 3839, 3430, 2189, 1992, 3850, 2518, 1995, 3198, 3286, 370, 3588, 266, 633, 2762, 3861, 3075, 3359, 1462, 2911, 2922, 1952, 2340, 3636, 1692, 59, 379, 2678, 1304, 3808, 3506, 832, 1891, 1855, 268, 1035, 1368, 736, 1912, 3311, 1890, 2198, 3337, 2248, 3461, 2733, 1822, 2972, 609, 3662, 463, 2918, 1898, 2675, 3117, 3906, 2405, 2945, 3807, 3403, 3745, 2440, 2480, 1785, 3591, 1521, 3368, 691, 939, 2381, 1066, 2515]. Artificial-Intelligence-Based [3198]. Artwork [1904]. Ary [2968, 1131]. Ascertaining [2458]. ASD [2398]. Asia [3538]. Asian [2213, 3117]. ASP [2268, 2242]. ASP-Based [2242]. Aspect [2620, 2561, 674]. Aspect-Based [2620]. Aspect-Oriented [2561]. Aspects [1282, 3227, 1106, 2529, 3473]. Assemblies [24]. Assembly [1868, 297, 2183]. Assess [2781, 3662]. Assessing [2732, 2724, 3502, 3298, 2642, 2208, 1781]. Assessment [3642, 3196, 3605, 1039, 410, 460, 3264, 3870, 1955, 356, 3876, 3152, 566, 621, 3573, 52, 2880, 174, 1254, 730, 2190, 652, 684, 2974, 1465, 2399, 2088, 591, 3070, 2179, 3483, 2192, 3841, 3487, 3395, 1154]. Assessments [2172]. Asset [2962]. Assets [1642]. Assigning [3507]. Assignment [2975, 970, 313, 1238, 3871, 1272, 1363, 2869, 3228, 58, 1973, 487, 3571, 814, 1988]. Assignment-Enhanced [3571]. Assimilated [3236]. Assimilation [2296]. Assistant [2153, 3395]. Assisted [1232, 1328, 1586, 3449, 2493, 3256, 3349, 2026, 325, 1094, 2571, 2861, 2702]. Associated [1611, 3324, 2806, 341, 1799]. Associative [1227]. Astronomical [3437, 2376]. ASVM [879]. Asymmetric [1268, 695]. Asymptotic [2442]. Asymptotically [2415]. Asynchronous [1313, 3135, 914, 1169, 2305, 3565, 1307, 117, 1545]. At-Least- [541]. Athletes [2113]. Atmospheric [3236]. Atom [2695, 1577, 1893]. Atomic [2695, 3233]. Atrial [1830, 2081]. ATT&CK [1913]. Attachment [2109, 878]. Attachments [1389]. Attack [3412, 2947, 2896, 3409, 3861, 1636, 1161, 1837, 2839, 3156, 3888, 2117, 1522, 3585]. Attack-Response [3412]. Attacker [1853]. Attacks [2812, 2537, 51, 2623, 3085, 1892, 2280, 2756, 2819, 2449, 3321, 1748, 3701, 2558, 909, 1966, 2890, 1638, 3546]. Attempts [3911]. Attendance [1332]. Attention [3846, 1524, 2821, 2279, 2957, 2919, 1951, 2943, 3790, 1754, 2729, 3390, 3706, 3885, 3821, 1003, 2978, 3828, 3159, 3457, 3386, 3231, 1851, 1927, 2323, 2513, 3341, 2452, 2102, 2525, 2654, 2766, 1747, 3725, 1456, 3621, 2657, 1604, 3842]. Attention-Based [2729, 3885, 2323, 2513, 2102]. Attention-Boosted [1951]. Attention-Driven [2452]. Attention-Survival [2279]. Attentional [3777]. Attentions [3341]. Attentive [2254, 1542]. Attentively [1951]. Attitude [1633]. Attraction [1212]. Attractiveness [977]. Attribute [1096, 1801, 3166, 1724, 1332, 560, 536, 533, 651, 666]. Attribution [3133, 2096]. ATTRIVAR [2873]. AUC [3184]. Auction [2423, 516, 319]. Audio [1759, 3856, 3360, 2959, 2400]. Audiovisual [2149]. Audit [1595, 1369]. Auditability [2087]. Auditing [3569]. Auditory [1524]. Augment [1769]. Augmentation [2615, 2111, 2001, 3138, 1886, 2667, 3226, 2201, 2856, 2917, 2559, 1769, 2665, 3666, 2142, 2748, 2514]. Augmented [2618, 3024, 519, 884, 1376, 3715]. Austria [2502]. Authentication [2087, 2410, 1654, 537, 1161, 433]. Authorship [2899, 2096]. Autism [1770, 3286, 1799]. Auto [3650, 1177, 540, 2014, 670, 626, 643]. Auto-Adjustable [670]. Auto-Diagnosis [1177]. Auto-Encoder [540, 2014, 643]. Auto-Tuning [3650, 626]. Autocontouring [3446, 3501]. Autoencoder [866, 1636, 1951, 3881, 3174, 2204, 2514, 2226, 2261]. Autoencoder-Based [3174, 2261]. Autoencoders [930, 3208, 2564]. Autoencoders-Based [3208]. Autoencoding [1743]. Automata [2996, 803, 430, 3374, 2435, 2675, 3119]. Automata-Based [2675]. Automated [3629, 639, 3493, 3434, 2938, 1880, 64, 3761, 3020, 2661, 3145, 598, 3147, 78, 619, 3062, 2885, 2703, 60, 2528, 1655, 3469, 1153, 1334, 2901, 3391, 2106, 2237, 1408, 3568, 3543]. Automatic [3192, 1549, 1798, 303, 3729, 22, 2129, 2965, 3551, 2625, 1835, 2049, 1830, 1955, 2313, 2593, 3224, 1847, 849, 1365, 3597, 3518, 1340, 3162, 314, 3384, 2788, 2225, 2559, 770, 1808, 1883, 463, 2030, 154, 2763, 2402, 485, 1512, 3189]. Automatically [2589, 144, 2931]. Automating [2594, 2288]. Automation [3358, 2522, 3376, 2824, 2989]. Automaton [1486]. Automobile [1056]. Automotive [2195, 3805, 3594, 3100, 2471, 3042, 855, 3189, 3250]. Autonomic [3822]. Autonomous [995, 3271, 991, 803, 702, 3610, 1097, 3078, 3144, 9, 1925, 3782, 2488, 2796, 3512, 1109, 20, 2345, 2494, 937, 2760, 1341, 3185, 2574]. Autoregressive [2506, 3158, 1070, 307]. Auxiliary [256, 474, 477, 705]. Availability [2950, 3562]. Available [1354]. Avalanche [3757]. Average [3829, 2249, 2940, 101, 294]. Averages [1032, 3691]. Averaging [145, 180, 2523]. Aviation [3148]. Avoidance [995, 3297, 730, 784, 2303]. Aware [2485, 2423, 3859, 928, 1068, 3359, 678, 987, 3611, 3763, 2091, 3294, 2198, 1277, 3687, 2763, 3317, 2, 3377, 1447, 332, 3771, 724, 2725]. Awareness [1724, 966]. Axial [2203]. Axis [2667, 3798]. Azerbaijani [2899].

B [3587]. B-Spline [3587]. B6G [2536]. Back [1241, 1248, 2648, 2877]. Backbone [1751]. Backbones [1901]. Backdoor [2839]. Backdoors [1901]. Background [2219, 3453, 2790, 859, 872]. Background-Subtraction [872]. Backlash [714]. Backpropagation [1068]. Backscattered [2453]. Backstepping [2205]. Backtracking [413, 2358]. Backtracking-Based [413]. Backward [3874, 3405]. Bacteria [2650]. Bacterial [548, 2939, 2650, 805, 3188, 3361]. Bacterial-Foraging [548]. Badger [3210]. BAG [1801]. BAG-DSM [1801]. Baggage [2830]. Bagirov [1210]. Baidu [3752]. Balance [447, 5]. Balanced [1936, 906, 848, 2549, 161, 2120, 774, 1094, 1599, 1773]. Balancer [1700]. Balancing [3080, 2950, 3443, 3224, 3101, 2198, 361, 310, 2183, 2478, 3329, 3800]. Balatonboglár [1251]. Ball [552]. Ballast [3024, 3029]. Ballistic [892]. Balls [1885]. Bamboo [818]. Banach [302, 2085, 398, 968, 1471]. Band [1611, 3143, 2839, 39, 128]. Bandit [544, 2316]. Bandits [3308]. Bands [889]. Bandwidth [1308]. Bang [692]. Bank [243, 1410, 645, 2790]. Banking [3130]. Banks [608]. Bar [2126]. Barcodes [1248]. Bare [1565]. Bargaining [2058]. Bark [3669]. Barrier [1121]. Barriers [3662, 1222]. Bars [23, 1616]. Base [668, 89, 2665, 3478, 827]. Based [3191, 1382, 3491, 2780, 2614, 2404, 3415, 1990, 3271, 1991, 3193, 224, 1335, 3018, 3713, 1031, 1934, 590, 3012, 3847, 1901, 3038, 2812, 2235, 2144, 3650, 3347, 2947, 3129, 2423, 3493, 2189, 1360, 1388, 3427, 2108, 1609, 2470, 2583, 3346, 2990, 2620, 2256, 1818, 2896, 3705, 3729, 714, 1418, 820, 3435, 1744, 476, 2260, 1261, 3579, 1911, 2427, 2466, 3132, 610, 2821, 1678, 1519, 1699, 50, 3198, 3286, 3009, 2422, 2993, 2268, 3857, 3438, 879, 2242, 799, 1179, 103, 3037, 3061, 2540, 3356, 2211, 25, 2767, 2624, 139, 393, 484, 596, 787, 850, 791, 1202, 1069]. Based [1180, 1514, 1639, 1624, 2965, 2907, 1654, 3749, 191, 170, 1670, 2043, 3773, 282, 540, 2774, 330, 3497, 976, 2542, 617, 1582, 3423, 3838, 3864, 1027, 3673, 1357, 123, 489, 235, 372, 1817, 3024, 3260, 2001, 2787, 1181, 3203, 3693, 1811, 457, 663, 1837, 3138, 874, 75, 1784, 2093, 511, 3647, 3709, 3140, 2145, 1526, 64, 2543, 2433, 2685, 779, 3304, 3680, 3207, 2526, 1846, 312, 443, 567, 630, 1183, 1308, 2628, 207, 1686, 3208, 2629, 1018, 2224, 410, 460, 1305, 1045, 3369, 1625, 3500, 3315, 1238, 2705, 2671, 3720, 854, 1669, 1607, 190, 1096, 366, 1913, 2476, 2827, 711, 3263, 3869, 2438]. Based [3870, 3445, 2454, 1825, 2519, 498, 483, 3296, 1392, 1328, 1868, 2631, 3446, 2167, 2283, 863, 466, 258, 790, 1207, 2430, 3678, 2632, 3094, 3297, 2315, 2330, 30, 2313, 1631, 2704, 371, 1692, 1491, 2666, 648, 1687, 3398, 238, 1244, 1410, 1489, 2175, 2590, 2290, 3095, 3504, 3505, 3787, 1869, 830, 3873, 3147, 2678, 164, 1886, 3783, 1245, 2377, 2194, 199, 2407, 2832, 2286, 3324, 619, 1022, 3318, 446, 2969, 369, 3874, 3373, 19, 1304, 696, 3148, 3557, 2289, 3321, 182, 2673, 3213, 1444, 2239, 969, 2527, 1293, 1157, 2094, 3214, 3215, 1959, 1615, 967, 2635, 2636, 2942, 1891, 1134, 3453, 2800]. Based [437, 149, 1855, 1903, 839, 1780, 3880, 3022, 1663, 2594, 1530, 566, 3155, 2297, 3794, 2992, 3766, 1565, 1326, 2054, 1722, 1588, 581, 3225, 2729, 3365, 3646, 429, 1791, 2796, 2667, 913, 2784, 3664, 304, 257, 308, 352, 403, 507, 706, 655, 701, 1173, 1035, 1172, 1485, 1337, 1330, 1411, 1516, 1961, 1629, 1905, 1775, 2371, 2505, 2776, 2920, 3611, 3455, 3332, 3885, 3760, 3556, 3821, 3887, 3888, 1568, 2181, 650, 3420, 3690, 1812, 397, 413, 462, 675, 736, 558, 828, 775, 1214, 1569, 2978, 2867, 3726, 3700, 3327, 3836, 3625, 1570, 3890, 2357, 1931, 755, 2095, 2554, 2749, 3159, 929, 1802, 3622, 548, 1838, 226, 1188, 1531, 3731, 3701, 3624, 2995]. Based [773, 3068, 2639, 1102, 3062, 1174, 1041, 3349, 2555, 3269, 3819, 1752, 2880, 134, 3322, 2557, 3748, 331, 1532, 935, 265, 3164, 1367, 1253, 1857, 2198, 1217, 3894, 3288, 3265, 1254, 3310, 3029, 3339, 3021, 2121, 2019, 3676, 1924, 1426, 1613, 2866, 54, 530, 546, 2021, 2445, 2986, 2314, 3583, 3232, 2125, 3691, 2706, 2113, 3320, 275, 1626, 2255, 2641, 2432, 3396, 2463, 2450, 2560, 2023, 3762, 1816, 2190, 3555, 2972, 1651, 763, 1618, 948, 2562, 2319, 1796, 652, 2908, 3284, 2682, 3467, 609, 2139, 2265, 2755, 3902, 1866, 1534, 1219, 2421, 1883, 2160, 2274, 3111]. Based [444, 3523, 1166, 3468, 2841, 2210, 2269, 2063, 1655, 3406, 990, 3469, 2343, 2115, 3904, 2196, 2665, 2261, 876, 1465, 1851, 3257, 233, 564, 851, 3113, 3418, 742, 1078, 1432, 2985, 3577, 2027, 2602, 2918, 3684, 3472, 399, 3563, 1190, 3073, 1001, 1153, 1334, 2389, 1595, 757, 2675, 3659, 2122, 1057, 1006, 2646, 1416, 1412, 2029, 1715, 3779, 986, 3802, 3475, 3174, 2535, 2897, 2316, 3175, 1733, 3527, 2405, 3560, 2399, 776, 2606, 2278, 1192, 1280, 1710, 2349, 953, 2763, 1441, 594, 3004, 1223, 3242, 378, 281, 254, 2945, 938, 198, 2568, 1508, 1460, 3243, 3244, 2033]. Based [1659, 2323, 3480, 1028, 1376, 857, 1369, 2070, 1820, 644, 2696, 3733, 2710, 346, 591, 784, 918, 1771, 2842, 3530, 2714, 2513, 3531, 2072, 3719, 3786, 3552, 516, 676, 536, 1094, 3609, 119, 2374, 3341, 1984, 1735, 247, 580, 1542, 2890, 3246, 1785, 890, 2649, 3534, 1647, 653, 2243, 1894, 3481, 3723, 1352, 2609, 2074, 2102, 3537, 70, 2338, 642, 2905, 2472, 1402, 227, 273, 262, 326, 515, 425, 710, 646, 719, 606, 847, 805, 870, 1427, 1543, 1819, 1688, 1826, 2392, 2653, 2525, 2654, 3247, 3064, 2748, 2736, 2981, 3319, 3482, 3715, 3584, 3353, 3661, 3655, 3811, 3484, 474, 613, 625, 2275, 3182, 3738]. Based [1650, 3914, 838, 1985, 1764, 2874, 3183, 3841, 1260, 467, 1197, 999, 3248, 477, 593, 578, 3540, 244, 357, 766, 807, 1355, 1393, 1545, 2078, 3186, 3724, 3818, 992, 1286, 2802, 1473, 1101, 656, 964, 260, 307, 332, 559, 1056, 988, 1512, 2079, 1839, 3725, 3710, 573, 348, 493, 1926, 2037, 555, 691, 2822, 2971, 705, 1198, 1878, 1380, 2180, 239, 231, 733, 872, 2358, 2379, 1331, 3618, 3920, 3765, 2514, 786, 1638, 1690, 1712, 3752, 2320, 3562, 319, 309, 377, 496, 518, 697, 552, 575, 826, 864, 899, 752, 963, 1026, 1547, 1456, 1350, 1916, 1627, 2082, 1763, 1643, 2400, 2346, 2387, 2273, 2655, 2656, 3015, 2702, 2847, 3046, 2860, 3767]. Based [3621, 3674, 3574, 3270, 3559, 497, 415, 1421, 1737, 2829, 3921, 3922, 3679, 3490, 958, 2831, 53, 1510, 1548, 1604, 2699, 294, 384, 553, 729, 3189, 3127, 3250, 3251, 3842, 1066, 1424, 1672, 2725, 3300, 703, 3001, 1461, 2797, 2531, 44, 1230, 3190, 3399, 3285, 1983, 14, 1000, 730, 104, 255]. Based-Intuitionistic [1096]. Bases [1670, 3149, 97, 2038]. Basic [1553, 3061, 1556, 2565]. Basins [1212]. Basis [1261, 2418, 2544, 3694, 62, 3047, 2151, 3481, 3791, 705]. Bat [1582, 222, 828, 494, 56, 1548, 3003]. Batch [2048, 2148, 3104, 3339]. Batch-to-Batch [3339]. Batching [1442, 2255, 3602]. Batteries [2156, 3555]. Battery [3366, 3649, 2021, 2072, 2976, 1005, 3800, 1987]. Battery-Connected [1855]. Battlefield [3893]. Bayes [1573, 2103, 3767]. Bayesian [1386, 3809, 471, 1178, 2512, 1097, 2803, 2200, 1271, 1801, 2048, 1087, 3263, 2881, 2328, 1007, 45, 3828, 3573, 54, 2384, 3467, 1222, 1121, 1256, 2316, 2681, 587, 3704, 918, 3532, 465, 3728, 2379, 1079, 697]. Bayesian-MAP [918]. BCCC [2056, 2057]. Be [105, 1645, 2282, 1359, 3116]. Beam [2420, 754, 1195]. Beam-ACO [1195]. Beamformer [2468]. Beamforming [777, 365]. Beamwidth [2468]. Bearing [928, 816, 1526, 2550, 701, 3828, 552]. Bearings [3310, 1969, 3487]. Beauty [1811]. Bed [1365, 2223]. Bed-Leaving [1365]. Bee [3850, 1995, 753, 1619, 1692, 1304, 1368, 736, 494, 609, 2480, 1521, 3728]. Bee-Ant [1692]. Beekeeping [2450]. Before [309]. Behaved [747, 882]. Behavior [2888, 1027, 3638, 30, 3301, 138, 1959, 403, 3512, 3758, 546, 2498, 3662, 3073, 3176, 1478, 1647, 2077, 2130]. Behavior-Based [3073]. Behavioral [1741, 3888, 3102]. Behaviors [3145, 2677, 964]. Behaviour [1322, 685]. behind [1655]. Being [3671, 3002]. Belief [1850, 663, 3557, 1461]. BELMKN [587]. Belt [702, 385, 2483, 1332]. Belts [2255]. Benchmark [2003, 1954, 1326, 550, 3886, 3824, 3337, 1705, 1704]. Benchmarking [3651, 276, 3687, 3543]. Bender [2788]. Bending [828, 1755, 641]. Beneath [2550]. Benefit [2940, 1222, 1121]. BENK [2697]. Bennett [51]. Beran [2697]. Bernoulli [1755]. BERT [1913, 2091, 3672, 2716, 3765]. BERT-Based [1913, 3765]. Bertalanffy [3817]. Berth [631]. BES [3203]. Best [1211, 744, 869, 759, 1358, 1711, 907]. Beta [151]. Better [2279, 3703, 940, 2415, 1570, 144, 150]. Between [723, 338, 3157, 3600, 3383, 3025, 1110, 2256, 2466, 629, 1946, 3040, 2281, 635, 942, 2214, 2551, 650, 601, 3069, 1089, 2323, 1320, 2038]. Beyond [1020, 3592, 551, 3360, 2030]. Bézier [3079]. BFS [61, 330]. BFS-Based [330]. Bi [1017, 3028, 2792, 887, 2095, 3159, 985, 599, 724]. Bi-Directional [985]. Bi-Level [887, 3159, 724]. Bi-LSTM-Based [2095]. Bi-Objective [1017, 3028, 2792, 599]. Bias [2938, 1902, 2693, 2798, 1226, 710]. Bias-Field [2938]. Biased [1888, 951, 1658, 977, 2703]. Biased-Randomized [1888, 951, 1658, 977, 2703]. Biases [2257]. bIAVOA [3701]. Bibliography [7]. Bibliometric [3855, 3564, 3075, 2678, 1490, 3523, 2727]. Biclique [615, 541]. Bicliques [504]. Bicolouring [3034]. Bicriteria [896]. Bidding [1517]. Bidirectional [491, 713]. Bifurcations [1228]. Big [692, 1034, 1744, 336, 956, 934, 2006, 1956, 321, 2328, 1044, 804, 3326, 1405, 3126, 692, 1879]. BiGridLSTM [713]. Bike [2434, 3041, 184, 1298]. Bike-Sharing [2434, 3041, 1298]. Bilateral [1937, 3197, 2635, 453]. Bilayer [562]. Bilevel [1188, 2803]. Bilinear [291, 723, 1571]. BiLSTM [1115, 3465, 3769, 1111, 1456]. BiLSTM-ARIMA [3769]. Bin [2950, 3367, 3619]. Binarization [1950]. Binarized [3596]. Binary [2616, 2896, 3768, 3867, 1087, 2712, 1025, 3321, 969, 1573, 1760, 382, 990, 1358, 3704, 3572]. Binary-Explainable [1358]. Binding [3066, 22, 34, 41]. Binomial [1796, 1222, 2873]. Bio [3590, 1605, 1257, 798, 1370, 206]. Bio-Inspired [3590, 1605, 1257, 798, 1370, 206]. Bioacoustic [79]. Biochemical [3497]. Biodegradability [89]. Biodiesel [2838]. Bioequivalence [3520]. Biofeedback [429]. Biogeography [493]. Biogeography-Based [493]. Bioinformatics [2104, 2413]. Bioinspired [3719]. Biological [414, 1046, 1772, 1602]. Biologically [1109, 1107, 3833]. Biology [1031, 1928, 7]. Biology-Inspired [1031, 1928]. Biomarker [3761]. Biomarkers [2670]. Biomass [3675]. Biomedical [2619, 1068, 1952, 1134, 3589, 1701, 3241, 3834]. Biometric [2149, 461, 3182, 3189, 3250]. Biomimetic [2813]. Biomolecules [46]. Bioreactor [3104]. Bioreactors [2208]. Biorthogonal [72]. Biparametric [2723, 317]. Bipartite [3026, 2093, 2403, 3034, 571, 886, 2876]. Bipolar [1974]. Bird [3299, 728]. Birds [236, 987]. Birthday [1210]. Bisimulation [662]. Bison [3160]. Bispectrum [463, 307]. Bit [1798, 240, 287, 1932, 2985, 2114]. Bit-Width [2114]. Bitemporal [1950]. Bitwise [961]. Black [2129, 1308, 1748, 1251, 2011, 3011, 1279, 960, 2308, 2264]. Black-and-White [1251]. Black-Box [2129, 1748, 2011, 1279, 960]. Blast [3795, 863]. Blast-Induced [3795]. Blast-Resistant [863]. Blending [1659]. Blind [1181, 1413, 3907, 918, 3185, 3487, 899]. Blister [3127]. Block [14, 844, 1452, 1041, 3349, 2019, 3231, 109, 394, 1141]. Block-based [14]. Blockchain [2423, 1797, 2623, 2858, 2548, 3653, 1632, 2449, 3782, 2117, 1595, 1369, 907, 3545]. Blockchain-Based [2423, 1595, 1369]. Blockchain-Enabled [2623, 3782]. Blockchains [1817, 2816]. Blocking [1949, 394]. Blocklength [793]. Blocks [1587, 2026]. Blood [3629, 3347, 2615, 2581, 1858, 472, 1277]. Blow [2263]. Blow-Up [2263]. Blowfish [1933]. BLSTM [2081]. Bluetooth [3823]. BmTSP [2120, 3903]. Board [2197, 1601, 2722]. Boards [132]. Bodies [29, 1297]. Body [2723, 545, 245, 2244]. Bohai [3299]. Boilers [3668]. Boiling [3882]. Boltzmann [573]. Bonds [2281]. Bone [230, 2236, 2831]. Bone-Conducted [2236]. Bones [1565]. Bonferroni [536, 533]. Book [74]. Boolean [2990, 437, 1637, 2605, 712, 2157]. Boost [623]. Boosted [1578, 1951]. Boosters [2948]. Boosting [1901, 440, 2301, 3494, 2009, 1722, 3459, 1543, 3064, 1753]. Bootstrap [1271, 3870]. Bootstrap-Based [3870]. Both [2267]. Botnet [3496]. Bots [2406]. Bottleneck [449, 1209, 936]. Bottlenecks [2592]. Bottlenose [3644]. Bottom [1023, 3801]. Bottom-Hole [3801]. Bottom-Up [1023]. Bound [1576, 2903, 1412]. Boundaries [101, 3521, 1844]. Boundary [2116, 1709, 2511, 3200, 2355, 1392, 1207, 2875, 1377, 1024, 1165, 2345, 2364, 1594, 2119, 1975, 2914, 3913, 1492]. Bounded [1991, 166, 974, 3612, 975, 1783, 422, 585, 1333, 1336]. Bounded-Error [1783, 1336]. Bounding [1788, 3481]. Bounds [3613, 3411, 1127, 2794, 2335, 985, 3481, 155]. Boussinesq [1877]. Box [2129, 3612, 1308, 2550, 1748, 2011, 1279, 960, 1507, 2264]. Box-Bounded [3612]. Boxes [2240, 811]. BP [792]. BPPV [3851]. Bradykinesia [3779]. Bragg [2184]. Brain [57, 24, 220, 2938, 2512, 1167, 2878, 2787, 2433, 3761, 77, 2594, 2259, 2639, 2813, 470, 2291, 1520, 1799, 604, 869, 759, 3909, 1487, 3177, 3341, 3912, 3181, 2872, 2829]. Brain-Computer [2512]. Brain-Inspired [1167, 470]. Brain-Storm [1487]. Brains [2625]. Brainwaves [2016]. Braking [1262, 3224, 2345, 520, 1762, 2847]. Branch [1576, 1239, 2709, 2008, 694, 2307, 998]. Branch-and-Bound [1576]. Branch-and-Price [2307]. Branching [629, 356, 1374]. Brassard [51]. Brazilian [1388]. Breakdown [2381]. Breakers [1868, 1897]. Breakthroughs [3426]. Breast [2670, 879, 3089, 2806, 1891, 3819, 3388, 2604, 2798, 2914, 83, 86]. Breathing [2880, 2332]. Bregman [360, 398, 3314]. Bridge [397, 1495]. Bridges [1428, 462, 736]. Bridging [3838, 2335]. Brief [1139, 355, 1536]. Brillouin [3022]. Broadband [777]. Broadcast [3689]. Broadcasting [2967, 3328]. Broadleaf [2741]. Brokerage [1674]. Bromodomain [22]. Bromodomain-Binding [22]. Bros [2964]. Brownian [1889, 737, 2946]. Broyden [298]. Bruising [2196]. Brushless [1441]. Brushstroke [2785]. BTS [3261]. Bubbles [1714]. Buck [2138, 3807]. Budget [576]. Buffalo [803]. Buffer [3139]. Buffers [599]. Build [3059]. Building [2834, 235, 2167, 1172, 3355, 1871, 1532, 3229, 2190, 3047, 1137]. Buildings [863, 3557, 1959, 3152, 2568]. Built [3895]. Bump [3388]. Bump-Hunting [3388]. Bundle [2477, 1033]. Bundling [1224]. Buoyant [1877]. Burgers [261]. Bursting [3045]. Bus [2597, 2598, 3907, 1303, 3182]. Business [3014, 3459, 1194, 1856, 1778]. Butterfly [500, 602]. Buy [1593]. Buy-Online-Pick-Up-in-Store [1593]. BWO [3548]. BWO-ICEEMDAN-iTransformer [3548]. Bypasses [2547]. Byzantine [3278, 672]. Byzantine-Tolerant [3278].

C [1665, 3467]. C-Shaped [1665]. C-Spine [3467]. CaAIS [2675]. Cable [2356, 3897, 233]. Cable-Connected [2356]. Cache [2958, 1466, 1424]. CAD [1917, 3805]. Cadastral [937]. CADrx [77]. Caffeine [3699]. Cagniard [1820]. Cahn [1112, 1090, 3617]. Calcification [2567]. Calculating [3312, 3879, 3677, 1980]. Calculation [2695, 1123, 528, 2210, 2338, 3399]. Calculator [1588]. Calculus [302, 2167, 892]. Calculus-Based [2167]. Calibrated [2906]. Calibration [2711, 1649, 1365, 354]. Calls [56]. Calorimeter [1719]. Camera [1463, 2555, 1135, 354]. Cameras [3143, 1904]. Camstyle [1604]. Can [3142, 2257, 1645, 2639, 2282, 3116, 2169]. Canada [3508]. Cancellable [1263]. Cancellation [279, 2220, 2379]. Cancer [2966, 2670, 1395, 1867, 1167, 3089, 2999, 3010, 2005, 3501, 1653, 356, 2806, 3508, 3456, 3819, 2893, 3231, 3388, 3902, 2604, 2413, 2798, 188, 316, 83, 86]. Candidate [3507]. Candidates [2466]. Canoeing [3493]. Canonical [3382, 571, 3374, 1163, 710]. Capabilities [2252, 1805]. Capability [3848]. Capacitated [562, 600, 2630, 3417, 3302, 3137]. Capacity [3366, 1526, 1267, 2550, 3364, 3455, 1634, 3731, 3479, 3179, 2076, 1839]. Capillary [2166]. Capsule [1636, 2419, 1965, 1347]. Capture [2332]. Capturing [1288]. Caputo [2355, 1673]. Car [1549, 639, 103, 1000, 2483, 3352, 2725]. Car-Following [2725]. CARA [1447]. Carbon [3755, 1863, 3344, 3368, 873]. Carbon-Energy [1863]. Card [546]. Cardiac [3346, 1460, 3380, 3585]. Cardinal [1388]. Cardinality [311]. Cardiometabolic [2038]. Cardiovascular [1686, 3440, 2735, 2835]. Care [2663, 2004, 896, 2177, 3563, 3906]. Career [3892]. Careful [2641]. Cargo [2456]. Carlo [373, 22, 1052, 2086, 3498, 2420, 2881, 568, 3150, 3878, 3659, 964, 573, 2991]. Carlo-Based [3659]. Carousel [2530]. Carrier [2659, 3159]. Carrier-Based [3159]. Carsharing [1776, 986, 1427]. Cartesian [1788, 2093]. CartPole3D [1943]. Cas9 [3813]. Cascade [1406, 62, 2392]. Cascade-Correlation [1406]. Cascaded [213, 211, 623]. Cascading [3448, 3071]. Case [1289, 2757, 654, 3850, 871, 440, 1049, 2456, 3438, 2086, 1670, 1582, 2394, 3371, 313, 2785, 2851, 3209, 434, 1392, 2700, 1631, 2892, 333, 3222, 1589, 3299, 1044, 3068, 1752, 2951, 3461, 3595, 1708, 1808, 2177, 3047, 3236, 855, 2645, 1333, 3801, 1507, 3115, 1971, 1307, 2979, 2223, 1478, 3537, 1315, 3188, 3752, 2577]. Case-Based [1670, 1631]. Case-Study [2645]. Cases [2302, 892, 1189, 2202]. CASF [3763]. Cash [243]. Casting [3677]. Cat [1157]. Catalan [3401]. Catalan-like [3401]. Catalog [122]. Categorical [708, 379, 2932, 1432]. Categories [850]. Categorization [350]. Caterpillar [2403]. Catheter [1757]. Cathode [2150]. Cathodic [827]. Cauchy [1834]. Causal [3133, 2975, 2750, 3369, 2795, 3398]. Causality [1037, 1782]. Cause [3133, 3263, 2759]. Causes [2249]. Cautionary [3816]. Cayley [1622]. CBM [2427, 2466, 3453]. CCA [3443]. CCFNet [2825]. CCTV [2237]. CDCL [1901]. CDCL-Based [1901]. Cech [954]. CEE [1241]. CEEMDAN [3860, 3400]. CEEMDAN-IHO-SVM [3400]. Cell [2615, 2581, 2424, 1018, 3447, 1962, 1912, 1655, 1605, 1972, 2288, 3391, 2072, 2954, 3485, 3800]. Cells [2966, 531, 1929, 1277, 1001, 188, 2385]. Cellular [803, 1328, 2435, 706, 2675, 2030, 3119, 963]. Censoring [2510, 3598]. Center [2792, 2057, 1808, 1656, 2952]. Centers [693, 2569]. Central [2949, 1757]. CentralBark [2836]. Centrality [905, 1312, 495, 1971, 3608]. Centralized [3242, 1422]. Centric [2580, 2180]. Centroid [691]. Cepstral [1672]. Cereal [1679]. Certain [1807, 677]. Certification [938]. Cervical [2966]. Cesàro [3480]. CFD [3469]. CFD-Based [3469]. CFGs [3002]. Chain [373, 2623, 3090, 2330, 1488, 566, 2437, 2768, 3659, 1596, 1864, 1459, 1482, 2263, 3124, 964]. Chain-Oriented [2330]. Chains [2165, 1928, 2369, 775, 2713, 2928, 589, 3537]. Challenge [976]. Challenges [2536, 1759, 3433, 3435, 3751, 1846, 941, 2633, 3808, 3157, 3426, 3272, 3895, 3604, 2735, 3306, 2884, 3370, 2798, 2440, 1711, 3376, 1005, 1738, 2245]. Chamber [3632]. Chameleon [645]. Change [2583, 714, 2304, 1950, 177, 3295, 347, 1070, 3466, 3598]. Change-Point [1070]. Changes [3867, 77, 3017]. Changing [1528]. Channel [3003, 2360, 3613, 582, 837, 2848, 1688, 2766, 988, 377, 228]. Channels [1780, 279, 2017, 3244]. Chaos [632, 1349, 3696, 1032, 2778, 1705, 3641, 1098]. Chaos-Enhanced [3641]. Chaotic [3658, 2820, 431, 1413, 1246, 1304, 1032, 2553, 2291, 233, 1487, 2338, 3484, 3046]. Character [121, 1464, 1733]. Character-Level [1464]. Characteristic [2336, 3172, 1001, 433]. Characteristics [2732, 3301, 3099, 3022, 462, 2602, 2772, 3391, 3483, 332, 245]. Characterization [1858, 3320]. Characterizing [3854, 3283]. Charging [2767, 2921, 2158, 3731]. Chart [2873]. Chatbot [1151, 3664]. Chatbots [2422]. ChatGPT [2515, 3604, 3116]. ChatGPT-3.5 [3116]. Cheap [2009]. Cheating [3920]. Chebyshev [3254, 2053, 3046]. Checked [2830]. Checking [2910, 2685, 1438, 3375, 223]. Chellappa [3740]. Chemical [25, 2992, 2425, 1085, 2198, 310, 1066]. Chemicalization [3855]. Chemistry [3734]. Chemotherapy [1997]. Chess [1756]. Chest [1431, 2629, 2322, 2557, 3334]. CHF [74]. Child [3759]. Children [3272]. Chiller [3794]. Chimp [1544, 1793]. China [3419, 3299, 1122, 3752]. Chinampas [2276]. Chinese [535, 3420, 2789, 3920, 518, 2176]. Chip [1754, 3700, 2772]. Choice [120, 207, 887, 2498, 1332, 2029, 2130, 3791]. Choose [2279]. Choosing [907]. Choreography [2845]. Chornobyl [2287]. Chosen [1645]. Chromatic [1415]. Chromatin [1929, 2484]. Chromosomal [181]. Chromosome [1408]. Chronic [3082, 1907, 1771]. Chronotype [685]. CHTKC [1711]. Church [903]. Churn [2888, 3478]. ChurnKB [3478]. Cicada [1984]. Cine [3446, 2667]. Cine-MRI [2667]. Circadian [3605]. Circle [1313, 3062, 1407]. Circle-U-Net [1407]. Circles [2212]. Circuit [1860, 1621, 1868, 1527, 3449, 2407, 1960, 2324, 3039, 1897, 367, 332]. Circuit-Width [2324]. Circuits [3428, 2377, 2448, 2871]. Circulant [2093, 2411, 3516, 555]. Circular [2913]. Circulatory [676]. Cities [2536, 1182, 1490, 2979]. Citizen [1651]. Citric [937]. City [1951, 2425, 3282, 884, 2445, 2335, 2561, 2823, 869, 759, 3908]. Citywide [963]. Civil [1052, 2295, 1535, 1777, 643]. CKD [3063]. Cladding [1035]. Clandestine [3354]. Class [1382, 2622, 1514, 2246, 3367, 853, 717, 751, 3869, 3289, 2257, 3321, 2380, 406, 1165, 716, 3375, 3803, 2563, 843, 2119, 1332, 2232, 286, 555, 3486, 3674]. Class-Incremental [3674]. Classes [3820, 2778, 2357, 2372, 1807, 91, 367, 677]. Classical [471, 1019, 1027, 2096, 297, 2208, 3525, 1016, 3745]. Classification [3379, 3629, 3847, 2580, 2937, 3651, 1014, 2581, 2617, 303, 2228, 2582, 3605, 1939, 84, 3346, 2620, 2129, 2211, 393, 732, 1180, 1147, 1555, 2512, 2676, 540, 3342, 1004, 1928, 2001, 3709, 312, 529, 420, 1835, 1952, 1305, 2999, 3010, 2900, 2546, 3840, 2005, 2322, 3554, 1087, 2815, 1025, 1956, 1584, 2430, 2704, 2865, 371, 1720, 2590, 2892, 199, 2941, 1529, 3321, 439, 969, 3806, 1959, 1674, 508, 1134, 3218, 1748, 1791, 1847, 2209, 640, 1743, 3160, 2398, 680, 3348, 470, 1573, 2187, 728, 2846, 3357, 3232, 3388, 3465, 2060, 1440, 2319, 3606, 3902, 1866, 3360, 3407]. Classification [520, 2959, 2665, 3238, 3306, 3172, 56, 3173, 3779, 3475, 2897, 2464, 3906, 150, 776, 1472, 3004, 2779, 2033, 2376, 2798, 1659, 2383, 2929, 83, 2142, 1227, 10, 2662, 2182, 70, 2338, 2152, 2789, 2525, 2521, 2981, 2836, 3669, 474, 3738, 3776, 3186, 638, 1380, 3187, 1155, 3562, 3621, 1603, 1347, 1718, 2244]. Classifications [1983]. Classifier [1382, 1498, 3259, 862, 472, 1723, 1743, 1931, 3824, 1966, 1375, 2232, 2264]. Classifiers [2257, 2371, 2638, 2558, 520, 2196, 2604, 82]. Classify [3440, 1754, 2886, 3168]. Classifying [1092, 1167, 1000, 3672, 2932]. Classroom [3187]. Clause [1272]. Cleaning [2495]. Clear [63]. Cleft [3125]. Click [1244]. Click-Through [1244]. Clicks [3644, 70]. Clickstream [2684]. Client [3370]. Climate [3491, 2304, 3698, 1241, 1172]. Climbing [1080, 3531]. Clingo [2252, 2268]. Clinical [2732, 84, 1757, 1395, 2359, 3684, 2801, 3533, 3065, 2828, 2192, 3543]. Clinically [3595]. Clinics [1689]. Clipping [2284]. Clique [95, 426, 3575, 1282, 3016, 193, 153, 1193, 712]. Clique-Width [3575]. Cliques [3328, 1136, 483, 2769, 767]. Clock [3513, 286]. Clogging [3320]. Cloning [687]. CLOPE [271]. Close [3437, 2127, 1371]. Closed [2616, 1524, 37, 2800, 2016, 637, 1807, 1755]. Closed-Form [2616, 2800, 1807, 1755]. Closed-Loop [1524, 37, 637]. Closeness [1145]. Closest [3259, 2752, 1592]. Closure [3812]. Cloud [1732, 3351, 2540, 1180, 336, 3359, 3774, 1784, 678, 1450, 688, 2815, 2438, 3732, 3027, 447, 138, 321, 3782, 3880, 3155, 693, 257, 401, 2554, 2807, 3274, 452, 2952, 1448, 2399, 1727, 273, 407, 1728]. Cloud-Based [1784, 2438, 2399]. Clouds [3869, 3404, 1985, 1693]. Clues [1113]. Cluster [3012, 3864, 2916, 3680, 1565, 2088, 319]. Cluster-Based [3864, 1565]. Clustered [3284]. Clustering [1382, 3192, 1231, 2757, 1934, 1680, 142, 1744, 3579, 1306, 340, 2342, 3438, 514, 866, 2111, 867, 2281, 727, 1124, 2685, 420, 1343, 2718, 708, 3627, 1051, 3445, 3076, 498, 258, 687, 2549, 3505, 3873, 1245, 369, 1304, 2328, 1100, 1506, 2800, 149, 2297, 1565, 550, 308, 456, 1000, 1590, 1774, 3327, 773, 1752, 2641, 2432, 161, 2601, 3784, 2564, 1656, 804, 647, 1432, 1300, 1445, 1795, 3244, 441, 2714, 1194, 1598, 465, 10, 890, 1647, 3535, 1370, 227, 262, 2241, 2479, 1197, 260, 249, 3618, 2038, 3015, 411, 255]. Clustering-Based [227, 3618]. Clustering/Distribution [2757]. Clusters [1317, 3576, 607, 150]. Clutter [1224]. CMOS [2320]. CNC [3556, 3821]. CNF [2545]. CNN [2581, 3705, 1069, 1357, 1886, 1615, 1516, 3897, 2196, 742, 3273, 3329, 2193, 474, 1456]. CNN-Based [1069, 2196]. CNN-BiLSTM [1456]. CNN-GRU-LSTM [3705]. CNN-Hopfield [1615]. CNN-LSTM [3897]. CNN-LSTM-GRU [3329]. CNNs [3761, 3231]. Co [340, 969, 1640, 2726, 3894, 2887, 158, 1402, 3081, 2661, 3901, 2339, 2383]. Co-Activity [2887]. Co-Clustering [340]. Co-Design [3894]. Co-Graph [2726]. Co-Lexicographic [1640]. Co-Occurrence [158]. Co-Processing [1402]. Co-Training [969]. Coal [3483]. Coal-Fired [3483]. Coalescence [1384]. Coalescing [2713]. Coalition [595]. Coarsely [936]. Cobb [2791]. Cockroach [382]. Cocoa [1478]. CoCoSo [3196]. Code [914, 2046, 2719, 450, 304, 1629, 3597, 2170, 2558, 58, 2136, 3306, 1540, 2191, 2243]. Code-Based [2243]. Codes [92, 955, 2968, 912, 3787, 822, 2232, 936, 2652, 895]. Coding [1344, 3856, 3857, 1123, 874, 3349, 3310, 1852, 2897, 1101]. Coding-Assisted [3349]. Codon [1518]. Coefficient [2757, 788, 522, 3044, 2208, 2267, 1672]. Coefficient-Triangle [522]. Coefficients [1201, 1100, 1506, 3896]. Coevolution [1394, 2075, 697]. COG [3877]. Cognition [1346]. Cognitive [1524, 1118, 3861, 2460, 1956, 1502, 3898, 1300, 2238, 3914, 348, 496, 228]. Cognitive-Driven [1524]. Cognitively [2460]. Coherence [619, 795, 1003, 1102, 2528]. Coherent [3380]. Cohesionless [2550]. Cohesive [344]. Cohort [3229, 1771, 676]. Cohorts [3727]. Coil [1504, 3609]. Coiling [3844]. Coins [2233]. Cold [2265, 3124, 1676]. Cold-Start [2265, 1676]. coli [1550]. Collaboration [1501]. Collaborative [370, 139, 147, 522, 2825, 3087, 1116, 3247]. Collecting [3037]. Collection [1308, 429, 569, 1472]. Collective [51, 1027, 3144]. Collision [995, 3297, 730, 1095, 784, 2509]. Collision-Free [1095]. Collocation [900]. Colloidal [1320]. Colonies [2720]. Colonoscopy [84, 3407]. Colony [1523, 3850, 1995, 753, 2720, 242, 1619, 1692, 1304, 3887, 1368, 736, 547, 3517, 2098, 2640, 3823, 2737, 1533, 494, 1095, 609, 1655, 1467, 2480, 1521, 1197, 697, 762, 3922]. Color [1344, 508, 1780, 2992, 2384, 262, 646]. Color-Based [646]. Colorectal [85, 2033, 3542]. Colored [159, 2457, 695, 265, 1358, 710]. Coloring [831, 2646, 1494, 677]. Colour [1476]. Colouring [1669, 571, 2956]. Column [1556, 1616, 2031]. Columnar [1310]. Combination [3494, 3503, 3744, 1485, 1812, 1927, 815]. Combinations [1306, 1345, 1640]. Combinatorial [1990, 1523, 636, 395, 581, 382, 569, 3580, 2349, 2744, 1501, 1809, 1725, 319]. Combinatorics [534, 1810]. Combined [431, 746, 2970, 1891, 809, 385, 614, 1453, 2571, 999, 1286]. Combining [3051, 3860, 3561, 3362, 1434, 1216, 3465, 1296, 937, 3919, 859]. Comfort [1995, 3698, 3224]. Command [3042]. Commensurate [2418]. Comment [353]. Commenting [2136]. Commerce [2543, 2684, 3280, 768]. Commercial [2044]. Commit [1540]. Commitment [820, 1169]. Commitment-Based [820]. Commodities [699]. Commodity [501]. Common [166, 360, 550, 1332, 1301, 2874, 186, 3001]. Common-Metric [1332]. Communication [2817, 2536, 823, 3753, 3688, 2709, 3043, 1362, 1872, 304, 3276, 3892, 1364, 2450, 1541, 3174, 1827, 3644, 3528, 1195, 3534, 2651, 1832, 496]. Communication-Efficient [2817, 1872]. Communications [837, 391]. Communities [1911, 2883, 2507]. Community [1624, 919, 483, 1410, 1485, 1102, 331, 495, 2105, 292, 3775, 396, 1643]. Commuting [3786]. Comorbidities [1771]. Compact [785, 2669, 659, 2551, 1191, 3616]. Compaction [903]. Compactly [1420]. Compactness [255]. Companies [1359, 1710]. Companion [1788]. Company [461]. Comparative [3316, 2184, 3847, 2899, 3009, 3061, 753, 1499, 511, 3205, 1846, 3142, 3840, 2476, 3590, 3636, 3296, 2865, 2449, 295, 3218, 3883, 3646, 2218, 1452, 3459, 3672, 1348, 3898, 3396, 959, 494, 293, 3784, 3047, 520, 3239, 953, 1449, 3485]. Comparing [1805, 2661, 3029, 3461, 2644, 1540, 2366, 1983]. Comparison [3025, 3759, 2260, 390, 13, 2524, 236, 2416, 2229, 243, 2544, 1504, 2977, 1018, 3067, 498, 2247, 213, 211, 3845, 3837, 2459, 2177, 1166, 2645, 1089, 2604, 1508, 3745, 1470, 2073, 2475, 2521, 216, 992, 3800, 1079]. Comparisons [1262, 990]. Compatibility [797, 3069]. Compatible [75, 2573]. Compendium [932]. Compensating [1600]. Compensation [1860, 2112, 1964, 3112, 3048, 247, 710, 766]. Competency [3084]. Competing [2085, 1524]. Competition [2466]. Competitive [3493, 2894, 3, 424, 834, 146, 58, 233, 217]. Competitiveness [3726]. Complementarity [322, 406, 1132]. Complete [2996, 2589, 2403, 571, 105, 1307, 3591]. Completion [957, 1358, 597, 2955, 625]. Complex [2757, 3194, 3045, 1317, 2919, 3610, 511, 954, 1895, 2883, 1834, 2453, 1051, 38, 2906, 2942, 2729, 1007, 1330, 3884, 45, 3889, 327, 3521, 672, 1223, 2696, 1522, 357, 2573, 575]. Complex-Valued [2757, 327]. Complexes [76, 962]. Complexity [1933, 3026, 3196, 1344, 1090, 2140, 1088, 630, 1399, 3696, 251, 2271, 3450, 566, 1483, 2639, 1170, 1364, 1065, 2681, 671, 2789, 1690, 679]. Compliance [2192]. Complications [2341, 3810]. Comply [3644]. Component [1575, 1519, 1948, 229, 3866, 3443, 3148, 3381, 1571, 1963, 1121, 2122, 1361, 1195, 160, 2484, 2163, 435, 3285]. Components [3596, 2436, 2330, 913, 2417]. Composite [3783, 1981, 527]. Composites [967, 52]. Composition [3359, 1874, 3594, 1174, 2063, 3536, 1230]. Compound [72]. Compounds [3069, 1066]. Comprehensive [1231, 2740, 3637, 3408, 3336, 2843, 2787, 3055, 2892, 1812, 3573, 3824, 1907, 1015, 2483, 3167, 2117, 3687, 3306, 3909, 3185, 3648]. Compressed [175, 240, 427, 110, 1364, 2065, 1274, 65, 870]. Compressible [1877]. Compressible/Incompressible [1877]. Compressing [733]. Compression [61, 2481, 971, 87, 1093, 3387, 3793, 116, 123, 1045, 941, 125, 3032, 1295, 2554, 1041, 134, 174, 80, 128, 339, 451, 860, 465, 1101, 2114, 3285]. Compression-Based [123]. Compressive [204, 306, 413, 2344, 485, 348, 435]. Compressively [1071]. Compressors [3835]. Computability [3613]. Computation [1934, 1576, 1178, 3135, 785, 3393, 1373, 1351, 1209, 101, 3278, 2551, 2382, 2715, 3281, 23, 1852, 2065, 2838]. Computational [1933, 2306, 3287, 3855, 24, 556, 3858, 813, 3736, 3761, 2099, 3712, 1319, 472, 1903, 7, 1483, 2595, 3016, 130, 3617, 162, 2788, 41, 2190, 153, 1106, 1970, 3473, 1416, 2031, 2032, 2681, 1192, 1975, 3241, 2980, 1729, 2707, 3489, 679]. Computationally [3829, 3310]. Compute [1552, 3812, 1694, 121].

Computed [1937, 1392, 2143, 3386, 3684]. Computer [57, 2539, 2512, 17, 3089, 3020, 3696, 3506, 2051, 2594, 3348, 1464, 2208, 2494, 85, 316, 904, 53, 2409]. Computer-Aided [57, 3089, 1464, 85, 53]. Computerized [272]. Computing [166, 1517, 1021, 1442, 2256, 2213, 426, 1577, 3013, 2672, 2524, 3035, 3408, 2774, 1167, 330, 2977, 3283, 534, 1810, 1206, 3389, 3750, 3027, 3502, 1652, 2517, 138, 172, 3324, 720, 2214, 3880, 3155, 3454, 1861, 338, 1252, 1337, 3516, 962, 1752, 1852, 3463, 1617, 1866, 3525, 1539, 1006, 3475, 1129, 776, 504, 310, 1727, 165, 589, 898, 273, 244, 792, 1988]. Computing-Based [1752]. Computing-Overview [1021]. Concatenated [2638]. Concave [3378]. Concealment [826]. Concentration [1062, 1361, 1456]. Concept [1103, 2141, 1663, 1966, 3695, 1153, 1334, 937, 1778, 2361]. Concepts [3061, 2046, 211, 2329, 1704, 213]. Concerning [1569]. Concrete [3693, 1886, 2636, 1616]. Concrete-Crack [2636]. Concurrent [427, 2313]. CONDA [1103]. CONDA-PM [1103]. Condensation [3882]. Condensed [3837]. Condition [280, 3076, 3068, 3897, 2092, 1303, 2577]. Condition-Based [3068]. Conditional [3718, 278, 2289, 1800, 2980, 1260, 3300]. Conditioning [3494]. Conditions [639, 289, 454, 1455, 1709, 1860, 1942, 2511, 2919, 2630, 1825, 1444, 1362, 1165, 2471, 1816, 2119, 592, 3185]. Conducted [2236]. Conduction [2116]. Conductive [772]. Conductivity [2064]. Cones [3804]. Conference [2036]. Confidence [1941, 2510, 3711, 2642, 1028]. Confidence-Based [1028]. Configurable [1917, 1466]. Configuration [2042, 907]. Configurations [3854, 17, 1735, 106]. Configuring [1378]. Confined [2467]. Conflict [791, 1202, 1884, 3672, 2876]. Conflicting [3778]. Conflicts [1836, 2887]. Conformance [1856]. Conformations [2556]. Conformity [3762]. Congested [1906]. Congestion [1176, 1514, 2903, 3447, 2237, 1447, 1026]. Congestion-Aware [1447]. Conglomerated [2959]. Congruential [1366]. Conical [3151, 833]. Conjecture [353, 324, 677]. Conjugate [322, 3633, 1475, 1620, 780, 664]. Connected [995, 166, 3328, 2582, 3434, 3596, 2356, 1269, 3078, 978, 1855, 1251, 3384, 3484, 2163, 2509]. Connected-Component [2163]. Connected-Components [3596]. Connection [723, 2335]. Connectionist [191]. Connections [3194, 3712, 3578]. Connectivity [974, 1609, 1551, 1387, 797, 571, 3364, 2056, 1323, 601, 1162, 3177]. Connector [786]. ConNetClus [1763]. ConNetClus-Based [1763]. Consensus [3729, 2548, 3297, 1632, 2328, 4, 929, 2121, 2117, 565]. Consensus-Based [3729, 3297]. Consequence [3908]. Consequences [598]. Conserving [358]. Consideration [1703, 2612]. Considerations [3363]. Considering [3054, 3394, 3698, 3301, 2050, 2637, 3344, 2345, 1966, 1920, 1507, 1974, 1906, 3483, 3918, 786, 453, 769, 3550, 565]. Consistency [2020, 592, 179]. Consistent [874]. Consortium [2650]. Constant [2468, 1333, 985, 1307]. Constant-Beamwidth [2468]. Constant-Time [1307]. Constants [3100]. Constellation [630, 2574]. Constellations [2197].

Constituents [120, 35]. Constrained [974, 3132, 386, 2669, 3592, 1825, 3750, 1527, 2549, 172, 1909, 3008, 829, 167, 3107, 1921, 275, 2561, 3721, 1470, 1196, 1050, 2610, 3913, 1500, 992, 1512]. Constraint [1806, 1917, 2123, 2837, 2853, 186]. Constraints [600, 787, 1236, 3727, 630, 673, 1080, 896, 3297, 882, 3364, 1658, 143, 3375, 1311, 2876, 2299, 603, 986, 3479, 3180, 821, 3655, 1742, 493]. Constriction [171]. Constructed [26, 3122]. Constructing [1269, 2724, 2642, 1796, 341, 1129, 3737]. Construction [3809, 1824, 1639, 1462, 1105, 2431, 3149, 2634, 1502, 3008, 2386, 3282, 3002, 2056, 1285, 2442, 2887, 1486, 1920, 1059, 1122, 1777, 1057, 936, 1771, 368, 1894]. Constructive [1936, 1515, 2706, 948]. Consumer [3129, 2044]. Consumption [1995, 1463, 2326, 3263, 3301, 3557, 3509, 3845, 3758, 3460, 2601, 2979, 3121]. Contact [303, 15, 2805, 3528]. Container [2990, 1293, 811, 3571]. Containers [741]. Containership [741]. Containment [1457, 589]. Contemporary [2954]. Content [1181, 1025, 149, 3420, 135, 1215, 1968, 3591, 2134, 53]. Content-Based [53]. Context [3741, 355, 1803, 713, 1046, 2164, 3758, 2882, 3708, 1651, 2755, 118, 752, 1713]. Context-Free [118]. Context-Sensitive [355]. Context-Sensitivity [713]. Contexts [577]. Contextual [3877, 148, 3308, 3635]. Contingency [1105]. Contingent [3490]. Continua [42]. Continued [2465]. Continuity [2563]. Continuous [107, 3131, 2360, 1264, 510, 2957, 3292, 323, 1635, 1378, 581, 3694, 2270, 1279, 3466, 2160, 2343, 1695, 851, 577, 3677, 2075, 2428, 1500, 488, 2509, 555, 2499, 1397, 453]. Continuous-Time [3292, 581, 1500]. Continuous-Wave [2343]. Continuously [1106]. Contour [1930]. Contouring [404]. Contourlet [879, 2787]. Contours [3501]. Contract [814, 497]. Contract-Based [497]. Contraction [940, 1412]. Contractive [1745]. Contracts [2443]. Contradiction [459]. Contradiction-Specific [459]. Contrail [1194]. Contrast [230, 239]. Contrastive [2856, 2935, 2748, 2514, 2273, 1693]. Contributing [1559]. Contribution [1097]. Control [224, 1550, 3194, 1143, 2618, 1993, 1524, 3197, 2470, 531, 1339, 2296, 1429, 661, 449, 3275, 2356, 478, 3199, 127, 596, 791, 1262, 1633, 1039, 3698, 631, 37, 1504, 1874, 1496, 2671, 2326, 717, 1741, 2476, 1046, 2751, 3750, 3297, 2412, 3211, 3343, 1457, 830, 343, 1612, 696, 1444, 437, 690, 1250, 3511, 3224, 764, 1909, 2055, 3303, 3694, 2205, 2961, 3332, 3760, 3515, 3420, 2386, 621, 793, 3889, 1390, 375, 929, 686, 1189, 2131, 2203, 3339, 6, 2885, 2730, 637, 1270, 2059, 2312, 404, 2908, 1900, 1761, 3169, 1535, 876, 564, 3414, 3600, 1453, 1875, 2138, 993, 3634, 2930]. Control [2532, 3476, 538, 3807, 588, 461, 837, 487, 1597, 2853, 3814, 798, 247, 2890, 1785, 3534, 1521, 2873, 2472, 1379, 425, 719, 1341, 3484, 2738, 807, 1762, 3818, 992, 623, 810, 1512, 1742, 3541, 3800, 782, 3835, 2487, 2971, 1725, 1638, 3546, 3632, 3270, 3547, 2770, 841, 729, 3060]. Control-Oriented [531]. Controllability [1236, 3522]. Controllable [1538]. Controlled [1031, 1236, 3165, 2842]. Controller [639, 991, 2947, 692, 2618, 888, 3643, 1232, 753, 3639, 2520, 983, 3783, 1623, 690, 1855, 2715, 1426, 626, 1921, 801, 293, 1120, 468, 3472, 911, 3120, 1441, 538, 362, 2976, 1521, 3050, 425, 2797, 617]. Controllers [2118, 1018, 2412, 3215, 834, 494, 2963, 475, 612, 953, 2945, 579]. Controls [981]. Convection [1994, 3044]. Convective [1112, 3618]. Convenience [2369]. Convergecast [210]. Convergence [315, 289, 388, 2085, 1938, 2086, 2246, 3128, 3735, 318, 3692, 3458, 3163, 416, 2020, 2186, 3171, 1078, 1089, 968, 2611]. Convergent [274]. Converging [748]. Conversation [1754, 2091]. Conversion [985, 1672]. Convert [1251]. Converter [2138, 3807, 2976]. Converters [1678]. Converting [1637]. Convex [3439, 3378, 739, 2477, 2503, 2783, 2551, 2961, 3721, 1033, 2653, 1638]. Convex-Concave [3378]. Convex-Hull [739]. Convexity [102]. Conveyor [828, 385, 2255]. Convolution [1147, 760, 69, 3821, 3468, 1807, 2119, 2654, 856]. Convolution-Type [2119]. Convolutional [3018, 2581, 1010, 2360, 2540, 3041, 540, 2444, 1428, 2003, 2900, 2547, 2162, 1087, 1413, 1392, 1621, 1687, 2254, 795, 3219, 2593, 2011, 1847, 640, 2978, 1570, 3355, 1140, 2886, 2986, 1440, 2564, 3468, 3175, 1733, 681, 3004, 2859, 3716, 2071, 2243, 898, 1819, 3571, 559, 364, 859, 872, 679]. Convolutional-Neural-Network-Based [1733]. Convolutions [3149, 3578, 3350, 3685]. Convulsive [3150]. COOBBO [244]. Cooking [1318]. Cooling [2571]. Cooperation [1280, 1346, 411, 768]. Cooperative [1850, 1940, 1275, 489, 2809, 1480, 3222, 3886, 582, 2058, 1333, 2323, 1394, 2075, 1742, 1643]. Coordinate [2215, 543, 1963, 2766]. Coordination [3394, 1457, 3760, 3514, 515, 1718]. Copositive [1663]. Copper [849]. Copula [1096]. Copulas [112]. Cord [385]. Core [203, 1317, 458, 3701, 198, 2414, 1320, 1660]. Core-Clusters [1317]. Cores [1249]. Coresets [1034]. Corneal [3062]. Corner [1971]. Coronary [2635, 2776]. Coronavirus [1441]. Corpora [725]. Corporate [3865]. Corpus [474]. Corrected [11, 3616]. Correcting [2232, 2652]. Correction [2938, 2370, 1810, 2629, 460, 3720, 213, 3826, 2053, 1334, 3049, 869, 2872]. Corrections [2380]. Correlated [3139, 2851, 722, 2932]. Correlation [1406, 2173, 3382, 67, 1410, 3022, 304, 3763, 1312, 62, 3029, 1899, 1163, 2038, 3285]. Correlation-Alignment [3763]. Correlations [3040, 3727, 3624, 3317]. Correntropy [2706]. Correntropy-Based [2706]. Correspondence [942]. Corrosion [3115, 2428]. Corrosions [3004]. Cos [3701]. Cosine [1232, 3869, 3023, 650, 2801, 608]. Cosine-Modulated [608]. Cost [3429, 654, 1936, 1430, 147, 2443, 2669, 1351, 1825, 1564, 693, 764, 1743, 3731, 23, 2103, 1816, 3047, 3815, 1507, 1222, 1121, 1416, 1459, 2478, 3591, 658]. Cost-Balanced [1936]. Cost-Benefit [1121]. Cost-Effective [3815, 3591]. Cost-Efficient [1816]. Cost-Quality-Risk [1825]. Cost-Sensitive [1743]. Cost-Sharing [1351]. Costs [1644, 1710]. Couchbase [2161]. CouchDB [2161]. Cough [2959]. Count [121, 1121]. Counter [854].

Counter-Terrorism [854]. Counting [3804, 3623, 1739, 3680, 2099, 2392]. Countries [1241, 3726, 2889]. Coupled [3582, 209, 711, 528, 418, 532, 404, 2389, 3048, 1414, 1887]. Coupling [1657, 2648, 805, 384]. Cournot [496]. Course [370, 171, 3410, 375, 3892, 3519]. Course-Keeping [375]. Courses [3470]. CoV [3665, 2632, 1598]. Cover [615, 689, 2729, 3365, 74, 551, 111]. Coverage [66, 491, 2660]. Covering [2404]. Covers [2403, 3364]. Covert [2985]. COVID [2401, 1431, 1175, 2304, 2700, 1869, 1570, 3161, 2557, 2223, 1449, 1342, 1289]. COVID-19 [2401, 1431, 1175, 2304, 2700, 1869, 1570, 3161, 2557, 2223, 1449, 1342, 1289]. CPM [2784]. CPU [422]. CPU/GPU [422]. CR [234]. Crack [3714, 1880, 1886, 2636, 3831]. Cracked [342, 42]. Cracks [1023]. Crane [2990, 3279, 3145, 3547]. Crash [2941, 1121]. Crashes [3569]. Crawling [1871]. Creating [2281, 1249, 3125]. Creation [2494]. Creative [3659]. Credibility [1825]. Credit [546, 1508, 883, 1753]. Creep [2734]. Crested [3891]. CRF [2256]. Crime [3722, 3355]. Criminal [3775]. CRIS [1049]. Crisis [1410, 3672]. CRISPR [3813]. CRISPR-Cas9 [3813]. Criteria [1388, 1481, 922, 3208, 3264, 3052, 1489, 2658, 172, 1502, 2863, 1359, 2746, 3904, 1465, 1315, 407, 2931]. Criterion [3848, 1709, 844, 1166, 2810, 1054, 605, 394]. Critic [453]. Critical [169, 2229, 3279, 3043, 2592, 2465, 1708, 3073, 3179, 2134]. Critical-Line [169]. Crop [2676, 2001, 3840, 2326]. Crop/Weed [2001]. Crops [1679]. Cross [3759, 2825, 812, 3678, 2832, 3043, 3885, 1070, 404, 1221, 399, 757, 1864, 2100, 1225, 2317, 2736, 474, 3776, 1135, 3921, 1200]. Cross-Camera [1135]. Cross-Coupled [404]. Cross-Cultural [3759]. Cross-Docking [1864]. Cross-Domain [3043, 474]. Cross-Entropy [1070, 1225, 3921]. Cross-Fusion [2825]. Cross-Language [1200]. Cross-Layer [812]. Cross-Modal [3678, 3885, 3776]. Cross-Network [2317]. Cross-Origin [1221]. Cross-Project [2832, 757]. Cross-Scale [2736]. Cross-Validation [399]. CrossFit [1681]. Crossover [2362, 2216, 3123]. Crosspost [2507]. CrosspostNet [2507]. Crosswords [1626]. Crow [2970]. Crowd [1739, 1104, 582, 2237, 1286]. Crowdfunding [2901, 2134]. Crowdsourced [3014]. Crowdsourcing [1325, 497]. Crown [2392]. Crowned [3891]. Crowns [2741]. Crucial [8]. Crude [3025, 2437, 3609]. Cruise [2445]. Crunch [692]. Cryptanalysis [1637, 3471, 3802]. Cryptocurrencies [2121]. Cryptocurrency [1992, 1744, 2621, 1829, 1063]. Cryptographic [1933]. Cryptography [563, 1372]. Crystal [2954]. CS [266, 296]. CSI [1586]. CSpredR [3319]. CSR [2002]. CT [3665, 242, 3086, 683, 3338, 1071, 3467, 2353, 2818, 918, 91, 642, 819, 260, 899]. CT/PET [1071]. CTC [85]. CTprintNet [2353]. Cu [1361]. Cu- [1361]. Cube [2855, 1523, 1374]. Cube-Free [1374]. Cubes [137]. Cubic [102, 1673, 1541, 560, 3079, 666, 104]. Cuboids [1320]. Cuckoo [236, 567, 561, 3526]. CUDA [3271, 2163, 2157, 3399]. CUDA-Based [3399].

Cuenca [3494]. Cultivars [2917]. Cultivations [1550]. Cultural [3759, 2422, 1465]. Culture [446]. Cumulative [858, 3173, 3788]. Currency [1128]. Current [1002, 3054, 1086, 3423, 2503, 2866, 2151, 827, 644, 1005, 3124, 3248, 623, 53]. Currents [3649]. Curriculum [3299]. Cursory [1605]. Curvature [3140, 1812, 2842]. Curvature-Based [3140, 1812]. Curvature-Controlled [2842]. Curvatures [1999]. Curve [2239, 338, 2804]. Curves [1999, 1024, 2336]. Custom [3355]. Customer [2888, 2165, 1808, 2177, 3478, 3176, 2479]. Customers [3139, 2684]. Customizable [940]. Customized [1250, 1303]. Customs [3189]. Cut [1417, 3872, 1562, 541]. Cutting [1430, 1911, 3498, 3255, 1562, 584, 2031]. Cutting-Edge [3498]. CVE2ATT&CK [1913]. CVEs [1913]. Cyber [1730, 3156, 1967, 2890, 2168, 1738]. Cyber-Attack [3156]. Cyber-Manufacturing [1967]. Cyber-Physical [1730, 2168]. Cyberattack [1838]. Cyberattacks [2168]. Cybersecurity [2410, 3409, 1815]. Cycle [3514, 1771, 671, 2612, 1066]. Cycles [506, 2053, 716]. Cyclic [2140, 1949, 3619, 81, 2288, 485, 2847]. Cyclical [838]. Cycling [241]. Cyclostationary [3487]. Cyclotomic [912]. CYK [1199]. Cylinder [425, 807]. Cylindrical [3211, 3343, 842, 1749].

D [819, 2037, 3643, 2221, 152, 683, 1204, 235, 2112, 1239, 635, 3211, 3343, 892, 101, 1563, 259, 1565, 3646, 1964, 1008, 852, 1191, 2126, 2945, 91, 731, 606, 479]. D-Bar [2126]. D-Domination [2037]. D-Grid [1204]. D-S [235]. D-STATCOMs [2221, 2112, 1964]. D2D [1328, 837]. DAE [3138]. DAE-GAN [3138]. DAG [2121]. DAG-Based [2121]. Daily [3599, 3383]. Dairy [3461, 3916]. Damage [1428, 347, 927, 736, 73, 2535, 1137, 3621]. Damped [1903]. Danger [1588]. Dangerous [2305, 1569, 1977]. Dark [1865, 988]. Data [3192, 2401, 3491, 107, 1933, 3379, 1335, 2375, 2253, 654, 3849, 2302, 3809, 2537, 3432, 3130, 2583, 1049, 2296, 1034, 1744, 1062, 905, 2621, 414, 89, 2622, 971, 1086, 13, 2456, 2840, 921, 278, 2623, 2765, 2539, 556, 2242, 3858, 87, 1093, 3387, 1902, 3199, 3041, 867, 3861, 1387, 1929, 2089, 1167, 336, 956, 862, 116, 1383, 372, 3024, 2001, 934, 457, 3138, 3204, 3371, 1784, 3425, 1182, 1308, 2785, 2999, 708, 3142, 2924, 3840, 1051, 3832, 2906, 3445, 1830, 1185, 3014, 1955, 2006, 1956, 2486, 1207, 1584, 1717, 2430, 2845, 598, 1792, 447]. Data [1661, 781, 2769, 1209, 1794, 1886, 1245, 2500, 2194, 880, 3569, 356, 321, 3373, 2328, 3707, 306, 1134, 3217, 3879, 1029, 3881, 1249, 1490, 2331, 822, 693, 1565, 835, 1483, 3225, 429, 2667, 886, 965, 1278, 2014, 2505, 3565, 650, 1962, 3226, 1312, 1743, 2201, 3282, 3625, 829, 1044, 2164, 2218, 2057, 2095, 3160, 3456, 680, 148, 1323, 143, 771, 1752, 3748, 174, 276, 935, 748, 419, 314, 645, 288, 1253, 909, 2187, 2917, 2721, 1965, 2559, 770, 722, 1967, 3899, 2135, 1769, 2599, 1708, 2450, 2560, 3274, 2562, 852, 3236, 3169, 2775, 2746, 3902, 2421, 1220, 3695]. Data [725, 1221, 2269, 1848, 804, 2826, 523, 1716, 452, 2565, 2065, 180, 2665, 2092, 3418, 889, 2126, 2952, 1256, 665, 1715, 860, 1972, 1459, 2684, 3906, 150, 1733, 2413, 973, 2399, 2647, 1896, 1445, 2901, 2691, 2779, 1224, 2202, 3244, 2376, 1844, 2183, 2323, 1037, 3528, 2072, 592, 3079, 676, 1711, 2142, 3376, 3532, 1704, 3533, 368, 3326, 2890, 117, 2824, 3329, 2662, 2523, 2168, 2608, 3180, 2385, 2989, 3620, 1782, 2789, 1402, 2264, 643, 1427, 1660, 2611, 3247, 3314, 3661, 3485, 3585, 3776, 1130, 2479, 2204, 477, 359, 2802, 1101, 656, 638, 1546, 1839, 3603, 3543, 3670, 2361]. Data [3920, 2514, 517, 3562, 1026, 3015, 2702, 3546, 3598, 841, 3126, 2531, 3285, 3202, 1600, 1879, 1405]. Data-Aided [965]. Data-Algorithmic [1902]. Data-Based [3562]. Data-Driven [3809, 3432, 1062, 2456, 2539, 3199, 3041, 3204, 2845, 1792, 3373, 3217, 1323, 3169, 2092, 2890, 1782, 1402, 3661, 3920, 2702, 3546]. Data-Efficient [3832]. Data-Intensive [2242]. Data-Oriented [1224]. Data-Parallel [1546]. Data-Processing [2269]. Data-Scarce [2765]. Data-Selective [748]. Data-Sets [1134]. Data-Structures [2065]. Database [123]. Databases [1049, 2161, 850, 727, 10, 1310]. Dataset [2213, 2765, 2107, 3024, 3218, 3355, 2742, 2844, 2494, 2763, 2747, 2836]. Datasets [1014, 2304, 2434, 2895, 2458, 1305, 1625, 2546, 2943, 2865, 2148, 1116, 2337, 1432, 3238, 3173, 3115, 3175, 1099, 1831]. DataToCare [3563]. Date [1568, 3524]. Dates [2956]. Davidson [1163]. Day [2260, 2599, 501]. Day-Ahead [2260, 2599]. DBN [2280]. DBN-LSTM [2280]. Dbpedia [1151]. DC [3366, 1678, 1108, 3783, 690, 1590, 1876, 1897, 1974, 1441, 3807, 1260, 2797]. DC-AC [3783]. DC-GAN [1260]. DC/DC [3807]. DCA [1108]. DCD [1078]. DCD-Based [1078]. DDL [3273]. DDoS [2280, 3888, 3701]. De-Noising [705]. De-Quantization [641]. DEA [3636, 1166, 3538]. DEA- [3636]. Deadlock [2093]. Deadlock-Free [2093]. Deal [3833]. Dealing [1902, 1651, 2061, 1713]. DeBERTa [3767]. Debiasing [1335]. Deblending [3001]. Deblurring [2653, 1926]. Debt [1110]. Decentralized [2087, 1339, 3378, 2751, 1700, 2197]. Deception [2304]. DeCGAN [3515]. Decimal [1446]. Decision [2578, 590, 1021, 84, 1339, 3729, 2975, 1481, 787, 3408, 1670, 3266, 2365, 3209, 3264, 1096, 1801, 1046, 2175, 1559, 2431, 1502, 2863, 2441, 1249, 3512, 735, 1044, 2471, 3893, 709, 2600, 1359, 2746, 1932, 684, 3358, 1283, 1057, 3764, 560, 943, 1734, 1979, 536, 2570, 227, 533, 1738, 2521, 3064, 2428, 2137, 407, 651, 666, 2499, 2931, 3632, 2829, 2699, 1424, 703, 3137, 2245, 3572]. Decision-Maker [2578]. Decision-Making [590, 3729, 3408, 3264, 3512, 684, 1283, 560, 943, 536, 2570, 533, 1738, 666, 2245]. Decision-Map [2521]. Decision-Time [3632]. Decisions [1528, 1293, 3466, 3716]. Declarative [2242, 1981]. Decline [3040]. Decodable [822]. Decoder [3790]. Decoding [1524, 2919, 3419, 427, 1568, 2813, 936, 895]. Decompensation [2663]. Decomposition [1788, 1360, 1938, 2689, 95, 1169, 1090, 3277, 1558, 1108, 1243, 721, 257, 376, 3891, 2786, 2484, 501, 1887, 2037, 3570]. Decomposition-Based [1360]. Decomposition-Ensemble [501]. Decompositions [916, 2093, 2403, 1329, 942, 2337, 1725]. Deconstructing [3868]. Deconvolution [3437, 3487, 899, 3570]. Decoupled [2900]. Decoupling [354]. Decrypted [2475]. Deduction [3182]. Deductive [3558]. Deduplication [135]. Deduplication-Enabled [135]. Deed [3908]. Deep [1406, 1675, 2614, 2973, 3629, 3193, 3846, 2812, 2580, 1850, 3129, 1431, 2617, 2040, 1404, 1609, 2149, 2990, 930, 3262, 2622, 2292, 3372, 3409, 2342, 3275, 2624, 596, 1233, 1202, 1069, 2495, 862, 669, 2388, 3382, 2787, 3090, 663, 3413, 1880, 3709, 3140, 779, 3283, 2629, 2819, 1239, 2924, 2049, 3143, 2090, 3446, 3750, 1025, 3653, 2741, 707, 3398, 3504, 2941, 619, 795, 1022, 2969, 3321, 1488, 508, 3744, 3452, 3453, 1748, 3219, 2285, 2992, 2637, 2011, 1290, 3365, 3646, 2909, 2667, 1115, 1590, 1905, 2371, 2776, 3512, 3332, 3821, 422, 535, 694, 1063, 3891, 1323, 3102, 2259, 2639, 2426, 3062]. Deep [2555, 3269, 2557, 3748, 3164, 1857, 1907, 1965, 2559, 770, 1924, 1464, 3607, 3462, 3230, 1277, 3899, 2445, 2135, 3583, 2310, 2528, 2060, 3901, 1685, 2564, 1520, 2754, 2210, 2353, 2115, 2066, 2665, 3049, 2771, 3474, 1006, 2753, 2178, 2818, 2567, 1016, 2502, 1719, 2220, 2763, 3176, 2859, 1977, 3909, 1460, 3825, 2033, 2798, 2513, 2034, 3531, 950, 2570, 1542, 2106, 2243, 3070, 3422, 2747, 2760, 2077, 3181, 1067, 1819, 2317, 2652, 2241, 3715, 3353, 2836, 1137, 1355, 1484, 3725, 3919, 2971, 2575, 364, 3648, 552, 826, 2847, 2872, 1882, 1987, 1461, 3291, 2596]. Deep&Cross [1244]. Deep-Fake [2077]. Deep-Learned [2342]. Deep-Learning [3102, 3607]. Deep-Learning-Based [3193, 2629, 3453, 2667]. Deep-Reinforcement-Learning-Based [2445]. Deep-Shallow [3049, 2771]. Deeper [3021]. Deepfake [1759]. Deephive [3147]. DeepLabV3 [3714]. DeepONet [1924]. Deepwater [3541]. Defacement [894]. Defaults [3865]. Defect [3781, 3780, 3140, 2832, 3876, 1047, 1140, 3622, 385, 757, 2178, 3529, 3915, 1505, 3127]. Defects [2590, 3310, 750]. Defending [1892, 2280]. Defense [3656, 108, 1738, 2078]. Deferred [3480]. Deficiencies [1711]. Deficits [2957]. Defined [2947, 90, 2046, 831, 3213, 2826, 3050]. Deformable [843, 91]. Deformation [3913]. Defuzzify [1946]. Degenerate [176]. Degeneration [552]. Degradation [620]. Degree [166, 2519, 2247, 3224, 1024, 3107, 2333, 515, 579, 1511]. Degree-Based [2519]. Degree-Constrained [3107]. Degressive [1131]. Dehazing [988]. DeHoop [1820]. Dekker [3464]. Delaunay [76]. Delay [945, 617, 1018, 2286, 296, 276, 1064, 247, 1405, 525]. Delay-Prone [945]. Delayed [414, 934, 3378, 2004, 3512, 2559, 3721, 3827]. Delayed-Enhancement [2559]. Delays [3197, 2457, 705]. Deletion [2726, 1379]. Deletions [109]. Deliver [2365]. Deliveries [2661]. Delivery [2988, 3133, 1568, 1786, 1171, 902, 2299, 3053, 1729]. Delphi [863]. Delta [2770]. Delving [2795]. Demand [3698, 1491, 2830, 1488, 3344, 3463, 2597, 2908, 1541, 1906, 1427, 513]. Demand-Responsive [2597]. Demand-Supply [1488]. Demands [3041, 1298, 1292]. Dementia [3847]. Democratizing [3815]. Demographic [2602]. Dendritic [1605]. Denial [2537, 2756, 3375, 1638]. Denial-of-Service [2537]. Denoising [1357, 3386, 2606, 2523]. Dense [2636, 1655]. DenseNet [2620, 3738]. DenseNet-Based [3738]. DenseNet201 [3350]. Densest [2111, 901, 541]. DenseZDD [659]. Densities [1374]. Density [3340, 1680, 3505, 775, 1116, 3322, 2393, 1106, 3050]. Density-Based [3505]. Dental [3198]. Deoxy [601]. Deoxy-Haemoglobin [601]. Depalletizing [3020]. DEPANet [3529]. Departing [2830]. Department [2159]. Dependence [620]. Dependencies [1888, 1658, 2282]. Dependency [3888]. Dependent [225, 1240, 1281, 1468, 3617, 143, 3896, 1338, 2076, 2905, 2651, 118]. Depending [2267]. Deployable [1389]. Deployed [2977]. Deployment [2456, 2921, 473, 2164, 1981, 1195, 3482]. Deposition [1361]. Depot [1287, 605]. Depressive [2622, 3065]. Depth [1463, 3009, 799, 629, 3055, 1527, 2555, 2777, 2955, 3842]. Depth-Camera-Based [2555]. Deregulation [3797]. Derek [98]. Derivative [1126, 981, 315, 388, 3197, 3643, 280, 2167, 2175, 2634, 267, 3694, 3896, 2963, 1673, 845, 3480, 337]. Derivative-Free [2167, 2175, 267, 337]. Derivatives [3221, 33]. Derived [2467, 3338, 3371, 1116]. Dermoscopic [2005, 2547]. Descent [1963, 2147, 559, 453]. Descent5 [664]. Description [3597, 1536]. Descriptive [1344]. Descriptor [657, 2649]. Descriptor-Based [2649]. Descriptors [2966, 915, 46]. Design [3412, 981, 1463, 1232, 25, 3363, 914, 1670, 2879, 2089, 374, 2774, 1389, 511, 3206, 2520, 2850, 2141, 3627, 434, 2049, 1611, 2438, 472, 3014, 3210, 2262, 2330, 1703, 2175, 830, 78, 2407, 3214, 3215, 3153, 1855, 1249, 2012, 1588, 1172, 2205, 2997, 397, 675, 3299, 2150, 2995, 2713, 1787, 2889, 728, 3894, 3288, 585, 1689, 1921, 801, 1897, 275, 2597, 2463, 2450, 3754, 1219, 2274, 1746, 3469, 3257, 1453, 1336, 1595, 3114, 3659, 1970, 911, 1416, 3666, 1972, 3802, 2532, 2032, 3119, 1280, 3120, 538, 2945, 1028, 857, 461, 3719, 3591, 3422, 2716, 2954]. Design [47, 1706, 2472, 3064, 608, 1451, 1849, 1870, 2358, 786, 2320, 1916, 8, 228, 553]. Design-Overview [2879]. Designing [984, 557, 1216, 2183, 368, 1854]. Designs [722, 856]. Desired [1066]. Despite [3017]. Destination [2982]. Destruction [3755]. Destructive [3497]. Detect [506, 1485, 3354, 2502, 1228, 3065, 309]. Detecting [3665, 3083, 2304, 3857, 3085, 2280, 3861, 3496, 2756, 3866, 2482, 1635, 3569, 2806, 1362, 1836, 2426, 2603, 3240, 2859, 3911, 3833, 3353, 1156, 527, 1778, 1230]. Detection [2973, 3415, 3379, 3714, 1934, 3686, 2235, 2580, 2144, 2966, 3054, 2864, 2947, 1759, 1431, 3430, 3195, 3822, 2040, 1939, 2583, 2990, 1770, 3435, 2622, 894, 2410, 610, 2292, 3063, 3286, 1943, 3856, 2765, 3438, 879, 3045, 115, 2540, 1317, 1202, 1624, 3086, 3059, 3780, 3610, 2512, 2230, 2858, 3749, 1353, 3773, 1948, 282, 1081, 3497, 976, 229, 3088, 3089, 3382, 3279, 1837, 2158, 2916, 3140, 3092, 1557, 1950, 3680, 1846, 630, 2904, 1309, 3369, 3010, 287, 825, 919, 157, 2547, 2827, 3263, 2943, 3590, 1830, 1314, 27, 483, 3678, 1687, 1720, 3096, 3097, 1886, 3876, 1438, 1047, 3321, 2636]. Detection [3307, 3451, 2285, 3100, 2009, 3578, 1859, 458, 1884, 459, 655, 701, 2014, 2371, 2776, 3611, 3747, 3888, 1723, 2552, 3785, 2856, 3700, 3327, 808, 3355, 3622, 1838, 1070, 3701, 1912, 148, 2259, 1340, 920, 3386, 2557, 331, 1890, 1532, 419, 3164, 1857, 3021, 2721, 3607, 3583, 3231, 495, 2483, 3017, 1708, 1375, 2023, 3762, 3467, 614, 2105, 2062, 1534, 1883, 3406, 1605, 1023, 3418, 742, 3238, 3306, 469, 3049, 2771, 3115, 2604, 2178, 292, 2567, 3475, 1016, 2535, 3033, 2030, 845, 154, 776, 2278, 480, 660, 3424, 3789, 1977, 85, 2979, 441, 857, 2696, 3273, 2071, 950, 316, 3552, 1856, 119, 3341]. Detection [3178, 160, 1320, 3329, 2168, 2760, 3334, 2102, 3912, 1679, 750, 836, 1688, 2392, 2654, 2933, 2736, 3314, 3669, 3914, 1443, 2766, 3915, 819, 3540, 1505, 1545, 1887, 3377, 260, 307, 396, 1056, 2079, 3685, 733, 859, 2514, 3648, 518, 1602, 1492, 1547, 2081, 2613, 2655, 2176, 3767, 3385, 1737, 2657, 2831, 86, 3127, 3550, 3190, 3291, 3202, 2244]. Detector [3597, 1253, 1107]. Detectron2 [2230]. Deteriorating [2480]. Deterioration [2025]. Determinants [1731]. Determination [22, 3612, 1633, 3731, 1573, 3531, 703]. Determine [662, 1559, 846, 276, 2570]. Determined [2868]. Determining [50, 2042, 1297, 1105, 3792, 3023, 2272, 3889, 2332, 1535, 2509]. Deterministic [1034, 1682, 3332, 2121, 1900, 1761, 1700, 1474]. Detonation [1580]. Detours [1579]. Developer [2136]. Developing [2401, 3432, 2281, 2924, 2369, 1904]. Development [3491, 2817, 2849, 3861, 3338, 2141, 59, 3448, 3663, 3101, 1212, 3726, 3518, 2844, 2450, 2139, 2421, 463, 1465, 3659, 2567, 3719, 2649, 1422]. Developments [1989]. Deviation [3394, 2853, 3488]. Device [3013, 706, 2450, 2679, 319]. Device-to-Device [319]. Device-to-Device-Enabled [706]. Devices [2659, 2044, 3216, 3893, 2232, 2655]. Dewetron [288]. DFS [796]. DGA [1375]. DGNCA [3780]. DGNCA-Net [3780]. DGSS [3610]. DGSS-YOLOv8s [3610]. DGT [1040]. Diabetes [3397, 1939, 37, 3142, 3900, 2602, 3600, 1898]. Diabetic [2821, 2230, 758]. Diagnosability [2057]. Diagnose [342]. Diagnosing [2302, 1387, 927, 904]. Diagnosis [2966, 2937, 2617, 57, 3705, 2927, 1177, 3860, 928, 816, 2907, 2043, 3342, 3089, 1686, 472, 1621, 1891, 3325, 3556, 3821, 2181, 3828, 2057, 272, 3819, 1367, 3310, 2972, 3234, 1969, 2261, 1048, 2818, 2030, 2691, 3716, 3403, 188, 316, 3400, 613, 2828, 3309, 2387, 2576, 53]. Diagnostic [3432, 3348, 2986, 1816, 1005]. Diagnostics [976, 1960, 1007, 3906, 2656]. Diagonal [2688, 1696]. Diagonally [754]. Diagrams [3877, 1156, 2699]. Diagrid [1959]. Dial [783]. Dial-a-Ride [783]. Diameter [1150, 3516, 276, 3111, 3798]. Diameter-Aggregation [276]. Dichotomous [1073]. Dictionaries [110]. Dictionary [922, 1183, 751, 875, 465, 2699]. Diesel [3068]. Difference [3829, 2477, 2493, 1736, 1058, 481, 2299, 1376, 731, 1346]. Difference-Based [1376]. Differenced [2659]. Differences [3292, 3550]. Different [3429, 2659, 1994, 390, 13, 3796, 1069, 17, 2589, 1499, 2544, 1018, 323, 1703, 3099, 1562, 1645, 1776, 546, 3017, 2177, 2196, 889, 1821, 953, 254, 1981, 91, 3005, 1380, 768]. Differentiable [2631, 3529]. Differential [3254, 3316, 3492, 1126, 1073, 1940, 1201, 753, 761, 3428, 2950, 2086, 885, 980, 2683, 3074, 420, 3867, 409, 754, 2355, 941, 3094, 2666, 2493, 1783, 1623, 1958, 532, 1919, 2380, 401, 1165, 2240, 2131, 3837, 3654, 2362, 1594, 2119, 1191, 891, 3772, 96, 3243, 1226, 1758, 1726, 2216, 2358, 309, 697, 815, 1627, 3842]. Differential-Evolution-Based [697]. Differentially [3284]. Differentiated [408]. Differentiators [2488]. Diffusion [3686, 1040, 1557, 77, 1649, 3103, 2020, 2442, 2267, 2606, 3733, 2824, 3586, 216, 2782, 2514, 3559]. Diffusion-Based [2514]. Diffusion-Weighted [77]. Diffusivity [481]. Digit [3629]. Digital [2485, 1404, 2470, 2584, 87, 1093, 2692, 431, 772, 2096, 3442, 3653, 3148, 3663, 3149, 3295, 3307, 54, 2986, 2294, 3236, 2066, 1781, 2448, 1854, 2385, 2610, 2275, 3916, 3125, 1372, 86]. Digital-Twin [2692]. Digitization [2067]. Digits [2712]. Digraph [2538, 506, 111]. Digraphs [831, 3575]. Dijkstra [3655]. Dilemma [3214]. Dilemma-Based [3214]. Dimension [2953, 264, 489, 3867, 1209, 1134, 1567, 3406, 1065, 389]. Dimensional [1430, 1525, 220, 980, 229, 3024, 3203, 3207, 2999, 1607, 1051, 3255, 3593, 1243, 2052, 2488, 1884, 1590, 1377, 2014, 3256, 2337, 3748, 3330, 2031, 2287, 3404, 3530, 3609, 1735, 3586, 870, 3314, 3791, 359, 1372, 2861, 1230, 15]. Dimensionality [3195, 1625, 3840, 3445, 1166, 2611, 3618]. Dimensions [2855, 2461, 2390, 2875, 160]. Dioxide [3755]. Dipole [33]. Direct [363, 1086, 825, 2661, 196, 2175, 272, 2866, 3798, 2472, 992, 354]. Direct-Comparison [992]. Directed [797, 1585, 1251, 2726, 376, 962, 3177, 396, 3015]. Direction [2583, 1863, 3633, 351, 3273, 384]. Directional [1609, 714, 3551, 707, 1075, 985]. Directions [1759, 3751, 3604, 2440]. Directivity [2913]. Directly [754]. Dirichlet [2194, 1691]. Dirichlet-to-Neumann [2194]. Disabled [2124]. Disassembly [3247]. Disaster [3276, 516, 2570]. Disasters [1821]. Discharge [3138, 2004]. Discharges [1087]. Discontinuities [60]. Discontinuity [2109]. Discontinuous [152, 1377, 2442]. Discount [2666]. Discoverability [3742]. Discovering [2845, 2605, 1731, 2134]. Discovery [1709, 201, 3133, 2140, 194, 2795, 3147, 3335, 2357, 1323, 2329, 3232, 41, 3368]. Discrete [3131, 2689, 3401, 1264, 1888, 785, 2165, 2489, 3696, 343, 1889, 3023, 3744, 69, 764, 1658, 1589, 2997, 755, 2703, 1900, 1761, 1777, 1336, 911, 1448, 985, 3825, 2137, 2499, 3549, 1474, 2995]. Discrete-Event [1888, 1658, 1777, 3825]. Discrete-Time [2165, 343, 1889, 764, 1336, 911]. Discretization [3013, 3200, 2493]. Discriminant [2187]. Discrimination [3032]. Discriminative [2256, 733]. Discussion [2205]. Discussions [3026]. Disease [3415, 3686, 3194, 2302, 2864, 2582, 1993, 1609, 3082, 2391, 2171, 2043, 1314, 2638, 1907, 2735, 2972, 3779, 2691, 676, 2835, 3334, 1259, 1679, 2525, 3685, 3309]. Diseases [3432, 57, 1686, 3440, 795, 3033, 1771, 2573]. Disentanglement [3917]. Disjoint [1418, 159, 2056, 1396]. Disk [2]. Disorder [1770, 3286, 3451, 548, 3065]. Disordered [790, 1622, 1355, 1690, 1350]. Disorders [2622, 3590, 2844, 2038]. Dispatch [2970, 222, 256]. Dispatching [2724, 2368]. Dispersed [1641]. Dispersion [476, 1714, 2630, 1620]. Displacement [3318, 2203, 1521]. Display [977]. Disruption [3582, 3093]. Dissection [940]. Dissimilarity [757]. Dissolution [2090, 3109]. Dissolved [1550]. Distance [2672, 1353, 1305, 635, 1207, 1209, 720, 1134, 2214, 2551, 3016, 4, 338, 1337, 1312, 3516, 2794, 1364, 1841, 109, 1795, 2088, 349, 3786, 1146, 1473, 1380]. Distance-Based [1305, 1207]. Distance-Hereditary [3016]. Distances [1133, 2686]. Distancing [1182]. Distillation [3025, 2628, 2396, 3503, 2590, 3876, 1629, 1764, 3685]. Distillation-Based [1629]. Distilled [1385]. Distinct [2615]. Distinguishing [2460]. Distortions [2420]. Distribute [3903]. Distributed [1014, 906, 905, 3013, 24, 1843, 1150, 2907, 636, 2756, 116, 931, 874, 287, 202, 919, 3443, 1956, 572, 528, 3878, 2415, 1091, 3022, 616, 429, 3760, 773, 1102, 143, 1814, 18, 2908, 804, 452, 3048, 2138, 1700, 1659, 516, 1785, 2523, 2300, 1599, 2241, 3539, 3248, 1832, 1545, 2574, 1199, 1712, 2702, 3546, 1422]. Distributing [1480]. Distribution [2757, 1641, 3427, 3718, 1818, 2894, 2221, 1525, 51, 3394, 2683, 595, 2664, 2112, 2503, 2741, 3373, 699, 3747, 3514, 3731, 121, 1964, 2335, 1618, 1796, 2768, 634, 2308, 1280, 165, 1258, 1301, 3247, 3841, 3248, 3540, 3724, 1546, 1643]. Distributional [442, 1060, 2408]. Distributions [3192, 785, 3727, 1751, 994, 1253, 1099, 1620]. Disturbance [1143, 888, 1623, 2730, 1762, 3818, 3060]. Disturbances [661]. Dive [2580, 3653]. Diverged [2862]. Divergence [157, 3154, 2681, 3314]. Divergences [2022]. Diverse [3091, 2128, 2556, 2018, 3238, 2323, 2317]. Diversification [428]. Diversity [3688, 1416, 594, 1198]. Dividing [1554]. Divisible [138]. Division [2720, 3349, 2951, 2567]. DIY [3815]. DLMinTC [3365]. DNA [1093, 1167, 2047, 125, 297, 1041, 1543]. DNN [2352, 1843, 1253, 1983]. DNN-Based [1983]. Do [2046, 3324, 2727, 1460]. Docking [32, 1864, 1436, 2409]. Document [1231, 627, 350, 2161, 2000, 3720, 149, 640, 989, 674]. Documents [2219, 2827, 284]. Does [1058]. DOF [404]. Doha [1219]. Dolphin [3644]. Domain [3018, 3713, 1938, 840, 3697, 3856, 2298, 2512, 3749, 3774, 2046, 2924, 434, 648, 2832, 3753, 3826, 3043, 3213, 721, 839, 268, 507, 1288, 2935, 293, 2421, 2343, 997, 2649, 1819, 2392, 474, 467, 2833, 3737, 1604, 3001]. Domain-Adaptive [2392]. Domain-Independent [2421]. Domain-Specific [3697, 434, 2935]. Domains [1670, 1499, 2373, 3118, 486]. Dombi [560]. Domestic [3668]. Dominance [603, 493]. Dominant [1606, 2332, 3786]. Dominated [1619, 1568, 3244]. Dominating [1418, 1327, 978, 2915, 3364, 1252, 162, 910, 3539]. Domination [3411, 1996, 3227, 2037]. Doors [3518]. Doppler [2659]. Dose [3699, 3386]. Dosing [719]. Double [489, 2403, 2373, 2430, 3512, 2868, 155, 807, 3547]. Double-Cylinder [807]. Double-Optimal [2868]. Double-Threshold [489]. Doubling [2731]. Doubly [279]. Doubly-Selective [279]. Down [1530, 2609]. Downwind [774]. DP [1944]. DPD3QN [3512]. DQN [2303, 1474]. DRAM [139]. Drawbacks [3419]. Drawing [376, 2166, 3395]. Drawings [345]. DRFAN [3706]. Drift [1103, 1309, 3611, 1966, 3695, 1778, 2361]. Drift-Aware [3611]. Drill [1798]. Drilling [1883]. Drive [2454, 3760]. Driven [2578, 1575, 991, 1002, 3809, 3432, 1524, 3797, 1646, 1062, 2456, 3409, 2539, 3199, 3041, 2541, 3564, 3204, 874, 2916, 3441, 2845, 1792, 1632, 3373, 3805, 3217, 3831, 737, 2201, 1323, 1515, 3803, 3169, 1695, 2092, 3769, 2469, 3906, 1224, 2890, 2480, 2452, 2716, 2402, 1782, 1402, 3770, 3661, 3827, 1443, 3545, 3920, 3188, 2702, 3546]. Driver [3846, 3250, 3550]. Drivers [3662, 2307]. Drives [1179]. Driving [3059, 1308, 3512, 3518, 3460, 2494, 3185, 3550]. DRL [2614]. Drone [3782, 2796, 2123]. Drones [2456, 2940]. Drop [1507, 1839]. Dropout [3852, 2645, 559]. Dropped [1645]. DRSN [3556]. DRSN-GCE [3556]. Drug [1790, 400]. DrugFinder [2346]. Druggable [2346]. Drugs [3520]. Drum [876]. DS [1094]. DScanNet [3622]. DSM [1801]. DSP [607]. DTI [157]. DTW [1000]. Dual [3262, 2767, 1127, 2629, 3043, 2231, 3512, 509, 1003, 585, 3323, 2103, 1311, 3667, 270, 533, 3811, 3788, 3685, 3771, 3737, 1604]. Dual-Energy [2629]. Dual-Frequency [3262, 509]. Dual-Fuel [3667]. Dual-Loop [3811]. Dual-Loss [3771]. Dual-Peak [3788]. Dual-Population [2767]. Dual-Priority [3043, 3512]. Dual-Rate [270]. Dual-Scale [3685]. Dual-Strategy [3737]. Duality [3180]. Dubins [164, 3007]. Due [1945, 2956]. Due-Windows [1945]. Dueling [3512]. Duplication [1479]. Duplication-Transfer-Loss [1479]. Duration [949, 3827]. Durations [947]. during [683, 1215, 2648, 1839]. DVB [970]. DVB-T2 [970]. DWT [3285]. Dynamic [1017, 3028, 2578, 1993, 2108, 726, 714, 916, 3084, 526, 424, 1860, 1429, 24, 1843, 1481, 478, 1888, 761, 139, 500, 629, 1514, 3041, 2774, 3382, 3693, 2627, 797, 1329, 2799, 335, 209, 2630, 1083, 3410, 1281, 3211, 3343, 1692, 3398, 2175, 1272, 1363, 2368, 1632, 3877, 1958, 1186, 1385, 130, 1007, 352, 448, 917, 1173, 1330, 2015, 3821, 3514, 45, 1684, 3624, 977, 1285, 2640, 2345, 1613, 1689, 796, 2314, 495, 2310, 2597, 2643, 2932, 569, 392, 2837, 2115, 564, 3577, 1777, 2675, 2790, 2222, 1459, 2316, 845, 973, 3280, 1700, 1758, 3273, 3177, 2570, 2523, 1706, 232, 719]. Dynamic [1495, 2736, 1500, 2874, 3915, 3675, 578, 3724, 2303, 3645, 872, 658, 496, 1916, 2702, 294, 2245, 1294]. Dynamical [528, 1694, 1900, 900, 2261, 1865, 1190, 2677, 1228, 1522]. Dynamically [1952, 2713]. Dynamics [3699, 2582, 3287, 414, 2720, 2820, 264, 1947, 931, 2045, 1779, 3458, 846, 2190, 1651, 2343, 865, 2122, 3560, 1192, 2263, 3814, 2707, 810, 2499, 3270]. DynASP2.5 [1329].

E-Commerce [2543, 2684, 3280, 768]. E-Government [3849]. e-Grocery [1280]. E-Learning [114]. E-Scooter [2737]. E-TUNI [3405, 3814]. EAND [2919]. EAND-LPRM [2919]. Earliest [3351]. Earliness [1945]. Early [3686, 2966, 1770, 2622, 3086, 1952, 2159, 1786, 3272, 1907, 3470, 188, 316, 3533, 3743]. Earth [3651, 3528, 3670]. Earthquake [3318].

Earthquake-Induced [3318]. Earthquakes [3657]. Easily [468]. East [3757]. Eastern [2979]. EC2 [1129]. ECC [3766]. ECC-Based [3766]. Eccentricity [165]. ECG [2235, 2213, 3063, 3342, 862, 3021, 3017]. ECG-Rhythm [862]. Echelon [2335, 1152, 1507, 3137]. Echo [2220]. Echocardiogram [3134]. Echolocation [56]. Eco [2849]. Eco-Innovative [2849]. ECOC [2232]. Ecology [35]. Ecology-Relevant [35]. Economic [2260, 2084, 2970, 1661, 222, 256, 3775]. Economics [2122]. Economies [3538]. Economy [1002, 3698, 3344, 829, 3758]. Economy-Related [829]. Ecosystem [3663, 2222]. ECT [772]. Ecuador [3494]. EDAS [863]. Edge [166, 2937, 159, 3013, 2473, 3035, 976, 3498, 2977, 3283, 157, 3750, 3027, 1585, 957, 3782, 3880, 2761, 689, 2726, 3294, 3386, 3778, 3323, 3762, 1827, 1224, 3529, 2169, 1107, 898, 1067, 2206, 3482, 756, 1545, 3917, 792]. Edge-Colored [159]. Edge-Guided [3323, 3529]. Edge-Missing [957]. Edge-Nodes [756]. Edge-Uniform [689]. Edges [1011, 1887]. Edit [720, 1841, 109, 349, 1473]. Edited [74]. Edition [3268]. Editor [2041, 2508, 1, 2055]. Editorial [971, 215, 1386, 3006, 880, 2271, 3450, 1213, 212, 2912, 1716, 1701, 1601, 2722]. Editors [2228, 1210, 1170]. Education [3852, 3588, 3264, 2742, 3604, 3396]. Educational [3851, 3881, 2645, 114, 3137]. Edwards [1617]. EEG [3003, 3847, 1609, 3705, 1519, 1387, 1183, 1567, 2201, 920, 2886, 463, 889, 2648, 3177, 2338, 3738, 3914, 3187, 1155, 1718]. EEG-Based [3847, 3705, 3914]. EEMD [816]. Effect [350, 1069, 2249, 2004, 2673, 2697, 2309, 2667, 1863, 1318, 2414]. Effective [1360, 2171, 3860, 1387, 330, 956, 3382, 3371, 2632, 3451, 306, 2715, 3607, 1822, 1683, 3815, 1120, 2124, 2374, 1598, 3591, 656, 2114, 2081]. Effectively [2455, 173, 3481]. Effectiveness [3601, 2377, 2674, 3911, 2770]. Effects [554, 3875, 149, 1302, 1279, 2025, 745, 2069, 1037, 2480, 565]. Efficacy [2665]. Efficiencies [1798]. Efficiency [2395, 3287, 3668, 3735, 2709, 1891, 1861, 2491, 3664, 1838, 2217, 722, 3553, 2568, 3770, 3538, 1832, 319]. Efficient [2780, 2341, 2817, 1680, 3651, 2616, 2188, 2352, 315, 2620, 1770, 1086, 1944, 3829, 3035, 3200, 3428, 2215, 2919, 3136, 3780, 330, 956, 669, 3864, 3498, 631, 75, 615, 1299, 1308, 3499, 1237, 2718, 2224, 3010, 329, 1607, 688, 740, 3832, 662, 760, 1413, 181, 1185, 978, 3871, 2698, 1751, 1243, 1209, 2334, 880, 619, 720, 1872, 3452, 3253, 1250, 2011, 1836, 1477, 2808, 3617, 2357, 2311, 3457, 1085, 1285, 3106, 3310, 796, 1015, 1769, 744, 18, 161, 2335, 2450, 1816, 3042, 1486, 998, 1403, 1220, 3524, 2816, 523, 436, 647, 1283, 2952, 2753, 1459, 154, 1361, 2681, 3641]. Efficient [968, 3602, 1844, 3644, 1338, 1407, 349, 1598, 1227, 2169, 1414, 653, 2402, 1671, 2777, 337, 445, 3353, 249, 1546, 3685, 3835, 218, 197, 549, 1492, 2991, 3767, 3385]. EfficientNet [2624, 2942]. EfficientNets [1529]. Ego [3271]. Ego-Motion [3271]. EGSDK [3323]. EGSDK-Net [3323]. Egyptian [2162]. eHealth [2164]. Ehécatl [2548]. Eigenfunctions [3171]. Eigenvalue [1470, 925]. Eigenvalues [1333]. Eigenvector [347]. Eight [3143, 239]. Eight-Band [3143]. Eight-Scale [239]. Eighth [289, 388, 267]. Eighth-Order [267]. Eisenstat [2850]. Elastic [754, 1440]. Elasticity [665]. Elastodynamic [1377]. Elastomer [939]. Elderly [3903]. Election [1189].

ELECTRE [1998]. ELECTRE-MOr [1998]. Electric [2485, 3427, 788, 2767, 2084, 2158, 3448, 2345, 3460, 801, 564, 1001, 3479, 2072, 3682, 2860, 789, 1987]. Electrical [393, 2544, 2112, 2503, 832, 1564, 2604, 2126, 3300]. Electricity [3427, 2260, 1061, 3509, 403, 3188, 453]. Electro [543, 1657, 2389]. Electro-Hydraulic [543]. Electro-Magneto-Mechanical [2389]. Electro-Thermal [1657]. Electroacupuncture [3181]. Electrocardiogram [1686, 2081]. Electrocardiography [3465, 3739]. Electrochemical [2150, 2156]. Electrode [3854]. Electroencephalogram [3184]. Electroencephalography [2639]. Electrolarynx [39]. Electrolysers [2614]. Electrolyte [1361]. Electromagnetic [1086, 3511, 385, 1719, 1820]. Electromechanical [3048]. Electromyography [2232]. Electronic [2377, 3189, 3250]. Electronics [568]. Electrooculography [1688]. Element [152, 2292, 1052, 3712, 3095, 675, 52, 11, 2788, 2644, 1616, 2861]. Elemental [3536]. Elements [785, 1377, 406, 68]. Elevating [2506]. Elevation [2485, 2385]. Elevator [2907]. Elias [92]. Elicitation [2486, 3573, 2902]. Elite [2863, 415]. Elitism [308]. Ellipse [3108]. Ellipses [2642, 2212]. Ellipsoid [1336]. Ellipsoidal [3192, 1766]. Elliptic [2723]. Elliptical [112]. ELM [508, 613, 1789]. ELM-RBF [1789]. Elman [766]. Elongated [3377]. Elongation [849]. Elston [1525]. Elucidating [2939]. Email [3911]. Embedded [924, 3132, 1097, 967, 3883, 2170, 3330, 2951, 2232, 2241]. Embedding [906, 924, 1233, 1631, 1075, 459, 1174, 1752, 3803, 821, 2766, 359, 3621]. Embeddings [668, 3091, 3799, 3717]. EMD [385, 530, 3710, 2379]. Emergencies [1790]. Emergency [2341, 787, 1182, 2159, 2272, 520, 3185, 2854, 206]. Emergent [191]. Emerging [2536, 1273, 78, 2016, 1122]. EMG [3894, 3904]. emgr [622]. Emission [1514, 2970, 1507, 2266]. Emissions [2661, 3901]. Emitter [2954]. Emotion [3846, 2895, 1720, 1754, 3885, 2091, 2201, 463, 3033, 1659, 2747, 3184, 2180, 1013]. Emotion-Centric [2180]. Emotional [2213, 3593, 3758, 3102, 470, 3765, 1200]. Emotions [2546, 2942, 2152]. Emphasis [3912]. Empirical [1803, 3277, 622, 3099, 1565, 2794, 2421, 3742, 35]. Employee [2234, 1352]. Employees [1899]. Employing [2852, 1377, 2886, 2602, 1745]. Employment [1920]. Empowered [2506]. Emulation [3576, 1706]. Enabled [2623, 3201, 3782, 706, 135, 1979, 1137]. Enabling [2975, 1879]. Encapsulant [3069]. Enclosures [2377, 1336]. EnCodec [3360]. Encoded [868, 1131]. Encoder [540, 2014, 1927, 643, 1511]. Encoding [3346, 793, 327, 1041, 2801, 1511, 3319, 2699]. Encoding-Based [3346]. Encodings [3756]. Encouraging [2849]. Encrypted [452]. Encryption [3658, 3766, 1828, 2327]. Encyclopedia [74]. End [2949, 2481, 1297, 3712, 1564, 3465, 2494, 2316, 1764]. End-Plate [3712]. End-to-End [2949, 2481, 3465, 2494, 2316, 1764]. Endometriosis [2765].

Endoscopy [1965]. Endurance [3455]. Energetic [3612]. Energy [3651, 1517, 3366, 1995, 3436, 3363, 2089, 3864, 2629, 2224, 2048, 2326, 3144, 3263, 2050, 3557, 3152, 543, 693, 987, 1863, 3332, 3845, 2556, 2198, 3288, 3265, 3823, 2217, 3460, 1015, 2294, 2103, 18, 161, 3553, 2450, 3167, 1651, 2601, 358, 1403, 3606, 647, 564, 2952, 1454, 869, 759, 2568, 2979, 3479, 837, 1094, 3538, 545, 964, 1809, 2612, 3835, 724, 549, 3574, 769, 792, 1987]. Energy-Aware [987, 2198, 724]. Energy-Conserving [358]. Energy-Efficient [3651, 3864, 2224, 1403, 647, 3835, 549]. Energy-Harvesting [769]. Enforcing [170, 3727, 2859]. Engagement [2539, 3201, 1754, 3102, 2169]. Engine [2199, 725, 3667, 1870, 841]. Engineered [2818]. Engineering [3412, 1052, 3074, 2295, 3264, 3593, 3210, 2330, 395, 3153, 1530, 1007, 2742, 3892, 2274, 3666, 3119, 3478, 3241, 3641, 3916, 3834, 1916, 1793]. Engineering-Oriented [3916]. Engines [2978, 3068]. English [2546]. Enhance [2371]. Enhanced [562, 1523, 3714, 2966, 1010, 3131, 3436, 2689, 1942, 3063, 449, 2720, 2919, 3639, 3495, 1928, 3140, 3866, 2359, 3554, 2486, 2148, 3783, 3321, 2852, 3154, 3514, 146, 1003, 3299, 3843, 3892, 2721, 3897, 3407, 1284, 1448, 3641, 2710, 878, 3538, 3484, 3571, 3776, 1544, 496, 2576, 1753]. Enhancement [1641, 230, 1555, 2236, 2512, 2676, 3260, 708, 2546, 2635, 39, 507, 1590, 3515, 1071, 2559, 742, 758, 2696, 1094, 3609, 2933, 3544, 239, 1497]. Enhancing [3839, 3347, 3129, 3195, 2395, 2391, 3287, 2621, 3085, 3363, 3090, 3010, 2496, 2943, 3722, 2709, 2550, 3219, 3881, 3511, 2491, 3823, 2983, 3907, 2354, 2142, 2944, 2300, 2182, 2716, 3742, 3585, 1747, 3717]. eNightTrack [2555]. Enough [2127, 1566, 1371]. Enriched [3478]. Enrichment [2477, 1116]. Ensemble [3415, 1031, 2580, 3195, 2391, 885, 1004, 1357, 3757, 1271, 2546, 1697, 2460, 1025, 2941, 969, 3611, 808, 1063, 2638, 2804, 770, 3357, 2060, 3467, 1166, 960, 1082, 2132, 2691, 2779, 1659, 3910, 501, 638, 3710, 1198, 1013, 1200]. Ensemble-Based [3611]. Ensembles [1578, 2441, 1060, 1931, 2217, 3899, 56, 1028, 2035]. Ensembling [2756, 1529]. Ensuring [1232]. Entanglement [2396]. Enterprise [2242, 2471, 1781]. Enterprises [2394, 3301, 3121]. Enthusiasm [2094]. Enthusiasm-Based [2094]. Entities [2903, 2645]. Entity [442, 3203, 3763]. Entity-Relation [3203]. Entropy [2896, 240, 928, 934, 1813, 3707, 122, 566, 822, 2998, 1070, 2382, 54, 2021, 3234, 1166, 2210, 1432, 2278, 2681, 1225, 2338, 515, 1101, 1350, 2387, 3921]. Entropy-Based [566, 1432, 2278, 2338]. Entropy-Regularized [3707]. Entry [1389, 2247]. Enumerating [3364, 259, 767]. Enumeration [933, 3227, 21]. Envelope [644]. Envelopment [2746, 2565]. Environment [1991, 3351, 443, 138, 2673, 3100, 2796, 1695, 237, 580, 2509, 883]. Environmental [3855, 2084, 3561, 3502, 3722, 2940, 2386, 3459, 2383, 3124, 3669, 3916]. Environmentally [771]. Environments [3759, 1551, 2781, 2692, 2280, 2495, 336, 3774, 2724, 321, 35, 1787, 3462, 2675, 1129, 1758, 2610, 300, 377, 1474]. Envy [1541]. Envy-Free [1541]. epf [1313]. Epidemic [2215, 1236].

Epidemics [3702]. Epidemiological [3194, 2474]. EpiInfer [3702]. Epileptic [3744, 1567]. Epileptogenic [2639]. Epipolar [480]. Epipolar-Geometry [480]. Epistasis [1384]. Epitope [36]. EPPM [2887]. EPSO [3593]. EPSOM [2870]. EPSOM-Hyb [2870]. Equal [1494]. Equality [821]. Equalization [2895]. Equalized [2860]. Equation [2757, 2116, 3817, 2511, 1019, 261, 2984, 1090, 941, 3253, 1903, 3617, 290, 2868, 265, 2020, 2442, 1486, 2529, 2293, 2267, 3586, 3728]. Equation-Error [265]. Equations [3254, 1126, 1020, 328, 455, 423, 454, 1455, 2085, 2811, 1938, 305, 324, 203, 1201, 3428, 3820, 1112, 1349, 980, 482, 409, 754, 274, 2355, 2493, 1783, 1958, 532, 1919, 1165, 2325, 3837, 3654, 1637, 1594, 1537, 1807, 2119, 341, 1191, 1089, 96, 1975, 2110, 3243, 310, 2147, 774, 299, 488, 524, 2743, 3834, 1141, 3559, 317, 1988, 353]. Equatorial [3683]. Equilibria [1576, 1275]. Equilibrium [3135, 1805, 1351, 2012]. Equipartition [3521]. Equipment [3092, 2377, 2759, 1568, 2092, 594, 357]. Equisum [908]. Equity [2962]. Equivalence [2256, 1117, 2357, 265]. Equivalence-Based [265]. Equivalent [2377]. Equivariant [3404]. Era [2375, 3806]. Erased [1135]. Erratum [455]. Error [2990, 2292, 3639, 2044, 418, 1783, 3826, 438, 2998, 929, 1055, 265, 1573, 1336, 2232, 2220, 2652, 155, 766, 826]. Error-Adaptive [3639]. Error-Correcting [2652]. Errors [1092, 2700, 1249, 1923, 2325]. Escherichia [1550]. Estate [3282, 3345]. Estimate [2357, 1862]. Estimates [3429, 3718, 1055]. Estimating [3091, 673, 3143, 2697, 416, 601, 1371, 2648, 1630, 3645, 1986]. Estimation [3340, 3271, 3193, 2659, 3130, 1642, 1463, 471, 788, 1525, 2473, 3668, 2443, 976, 2388, 372, 2851, 3795, 2881, 466, 2632, 379, 3790, 1783, 1889, 1696, 3373, 2800, 2143, 3156, 737, 965, 2693, 509, 3890, 716, 3457, 3349, 2393, 2217, 2156, 801, 1772, 3047, 1396, 1106, 1777, 1875, 1766, 2293, 2810, 480, 3317, 3177, 653, 3179, 2074, 2777, 710, 836, 3661, 216, 2871, 3488, 3752, 377, 718, 384, 3842, 2409]. Estimations [2247, 1906]. Estimator [2870, 3421, 2697, 845, 1545]. Estimators [1433]. Ethical [3306]. Ethics [3908]. Euclidean [380, 4, 948, 2686]. Eukaryotic [1785]. Euler [1755, 3172, 3405, 3814]. Euler-Type [3405, 3814]. EUR [74]. European [3668, 2711, 3413, 3222]. EUSO [2531]. Eutrophying [993]. Evacuation [959, 1286]. Evaluate [2704, 1489, 2893]. Evaluated [3815, 3902]. Evaluating [2616, 3848, 3862, 2249, 2819, 2377, 581, 722, 2964, 3177, 3911, 2824, 3912]. Evaluation [2375, 1388, 3366, 3013, 2161, 1833, 1181, 934, 678, 3315, 1241, 3501, 823, 863, 3145, 2330, 2007, 2759, 1247, 450, 428, 1564, 2761, 1588, 3420, 1812, 3726, 3746, 3893, 3396, 12, 2963, 1057, 2031, 1734, 3602, 2824, 2168, 2073, 3182, 407, 3488, 1397, 3395, 2770]. Evaluations [981, 1757]. Evapotranspiration [5]. Evasion [1636, 1091]. EVCS [2158]. Event [3379, 2395, 1888, 2589, 1784, 1183, 3335, 1658, 1912, 1340, 2703, 1220, 2837, 1777, 3825, 284]. Event-Related [1183]. Events [1939, 2799, 2285]. Every [1655, 2364]. Everything [3444, 105]. Evidence [235, 3678, 3662, 1094]. EvoDevo [3719]. Evolution [1073, 3699, 1940, 753, 761, 2086, 885, 1400, 2683, 3074, 3867, 1051, 2700, 3094, 2666, 3657, 401, 800, 3597, 2362, 3772, 2383, 1226, 1758, 1726, 2216, 2448, 1894, 3123, 2402, 1839, 2358, 309, 697, 815]. Evolution-Driven [2402]. Evolutional [2287]. Evolutionary [1990, 1934, 1680, 3605, 2470, 512, 2127, 1702, 631, 3499, 190, 3636, 2262, 2412, 2050, 1958, 1378, 2011, 557, 3884, 3845, 1188, 3624, 2715, 2098, 2558, 1518, 2641, 444, 2250, 843, 3073, 480, 2070, 3719, 2838, 2650, 502, 3920, 2346]. Evolutionary-Fuzzy [444]. Evolving [639, 1133, 1321, 2368, 689, 60, 2605, 1668, 2448]. EWMA [3257]. Exact [1017, 1306, 3654, 1113, 3171, 153, 81, 1193, 1338, 2076, 952]. Exact-Local [2076]. Exactness [3395]. Exam [3396]. Examination [221, 2095, 2196, 2727]. Examining [3777]. Example [2302, 2899, 2054, 3672, 2073, 3771]. Excel [1439]. Exchange [794, 3708, 3117]. Excitation [1535]. Excited [532]. Exclusion [3233, 3464, 2573, 186]. Exclusive [427]. Executable [2930]. Execution [386, 773, 1787, 2951]. Exemplar [2935]. Exercise [2051]. Exhaust [3901]. Exhaustive [1517, 21]. Existence [3492, 3135, 482]. Existing [2512, 94]. Exogenous [2599]. Exoskeleton [2532]. Expanding [291, 464, 1422]. Expansion [2780, 1261, 3612, 2227, 1485, 3625, 541, 712, 1098]. Expected [1825, 1058, 277]. Expensive [76, 3256, 3584]. Experience [3512]. Experiences [1474]. Experiential [3892]. Experiment [2395, 2427, 2466, 48, 3453]. Experimental [1143, 871, 3197, 2416, 3140, 797, 1496, 2420, 2378, 722, 626, 1766, 2532, 2032, 1734, 2069]. Experimentation [739]. Experiments [2306, 2850, 1018, 2515, 1838, 669]. Experiments-Based [1018]. Expert [472, 3573, 763, 1461, 2708]. Experts [1460]. Expiration [3524]. Explain [3716]. Explainability [3452, 3695, 2132, 3620]. Explainable [3415, 668, 3432, 2040, 2518, 3082, 3066, 3852, 3089, 2365, 3369, 1653, 3808, 3557, 3806, 1890, 3272, 3337, 1695, 2261, 1358, 2884, 3341, 3585, 3916, 1982]. Explaining [2521, 3865, 1983]. Explanation [2867, 1747]. Explanations [3372, 2807, 3231]. Explicit [3356, 3221, 3114, 3800]. Exploitation [3080]. Exploiting [1901, 2344, 960, 792]. Exploration [3080, 3566, 848, 1083, 2892, 2556, 3606, 1685, 1821, 774, 2651, 3308]. Exploratory [1164, 2769, 1326, 3243]. Explore [1653]. Explorer [2564]. Exploring [1798, 3839, 2781, 2538, 1867, 2984, 3774, 2678, 1903, 1791, 3016, 2552, 3272, 2917, 3898, 2665, 3921]. Exponent [3643]. Exponential [2683, 3647, 430, 1652, 2239, 3257, 2110, 3179, 155]. Exponentially [3222]. Exponents [528]. Express [1301]. Expression [1381, 2806, 1312, 755, 1637, 3178, 3725, 1983]. Expressive [3375, 2964]. Extended [1455, 3436, 3817, 1996, 3639, 3031, 2900, 983, 3154, 2778, 1390, 1174, 3654, 1538, 68, 2069]. Extended-State [3639]. Extending [454, 2140]. Extensible [3337]. Extension [3431, 3149, 839, 1970, 2946, 1513]. Extensions [2723, 2876]. Extensive [2046]. Exterior [3752]. External [2443, 3338, 1287, 3353, 2479]. External-Fleet [2443]. Extinct [1479]. Extinction [190]. Extinction-Based [190]. Extra [3155]. Extracted [2818]. Extracting [2794, 150, 158]. Extraction [3847, 1578, 2938, 2211, 3203, 1183, 1834, 2453, 3052, 1047, 149, 3022, 3156, 3664, 655, 3458, 2421, 3468, 463, 3239, 2897, 2185, 188, 750, 3915, 3543, 1155, 2397, 3621, 3652]. Extrapolate [2465]. Extrapolation [3692]. Extremality [1622]. Extreme [3494, 1928, 711, 1278, 1774, 680, 3459, 3339, 587, 638, 1753, 1543]. Extremities [90]. Extrinsic [2200]. Eye [2165, 3059, 2438, 3906, 1688]. Eye-Tracking [2438]. Eyebrows [1812]. Eyes [1812].

F [455]. Fabric [3140, 508, 3915]. Facade [863]. Face [2375, 1811, 3032, 2054, 1516, 1812]. Facebook [3583]. FaceNet [2375]. Faces [3157]. Facial [2895, 755, 3725, 1983]. Facilitating [3426, 3282]. Facilities [1552, 1572, 2335, 1152, 1507, 2124]. Facility [1321, 585, 1354, 1439, 1311, 1292, 2570]. Facing [3754]. Fact [1438]. Fact-Checking [1438]. Factor [34, 41, 2560, 684, 986, 1315, 300, 1737]. Factored [325]. Factoring [1594]. Factorization [868, 1295, 1409, 2097, 1740, 2270, 1542, 3603]. Factors [3494, 3561, 2249, 1559, 462, 3161, 1899, 594, 3669, 2134]. FACTS [2308]. Fading [2360, 2017]. FADIT [989]. Failure [3196, 1092, 49, 3448, 52, 3459, 3071, 565]. Failures [3112]. Fair [1790, 2542, 1954, 3237, 2427, 2466, 3453]. Fairness [1335, 3687, 943, 2798]. Fairness-Aware [3687]. Fake [2580, 3083, 3088, 3678, 1438, 2727, 2077]. FakeBob [1892]. Fakes [2009]. Falkenauer [3367]. Fall [2555]. Falls [1982]. False [2395, 1723, 2552, 1857, 3418, 3546]. Falsification [3861]. Families [2188, 659, 1420, 2723, 1106]. Family [328, 1349, 1947, 980, 318, 267, 2239, 360, 1212, 2723, 2186, 93, 1258, 299, 317]. Fanfiction [2096]. Fano [92]. Fans [654]. Farm [1254]. Farms [1404]. Farthest [2752]. Fast [692, 46, 1148, 659, 1509, 2370, 2207, 3504, 172, 2407, 2635, 835, 2010, 1588, 1884, 3156, 2784, 923, 2998, 2794, 470, 748, 989, 902, 3521, 2210, 660, 85, 1338, 1320, 1985, 2822, 733, 856, 1098, 1910]. Faster [1133, 940, 3212, 1687, 219, 376, 1592, 214, 742, 679]. FASTSET [835]. Father [100]. Fatigue [3059, 1688, 3914, 3550]. Fault [2144, 3054, 2927, 3860, 928, 816, 2907, 3749, 3773, 3382, 3092, 1621, 3096, 2272, 1960, 701, 3325, 3556, 3821, 2181, 3828, 2056, 419, 1367, 1375, 1816, 3234, 1883, 1969, 233, 3475, 2030, 2232, 845, 2072, 2853, 1005, 589, 1782, 389, 1698, 613, 2874, 467, 999, 3540, 307, 332, 2387, 2576, 2656, 3737]. Fault-Containment [589]. Fault-Detection [845]. Fault-Diagnosis [2387]. Fault-Tolerant [2056]. Fault-Tolerant-Embedded [2232]. Faults [1966]. Faulty [3504]. FCM [1940]. FD [1938]. FDG [3086, 3338]. FE [1938, 3776]. FE-MCFN [3776]. FE/FD [1938]. Feasibility [1234, 2286, 1173, 1030, 398]. Feasible [3394]. FEAST [925]. Feature [3012, 1073, 3847, 2580, 2149, 3259, 3063, 3856, 2211, 3860, 1233, 2111, 26, 2724, 1183, 2904, 1309, 1625, 3445, 1956, 790, 1751, 3096, 2289, 3805, 2451, 3268, 2097, 2808, 1723, 1368, 640, 755, 1324, 3456, 3891, 3457, 3824, 3459, 3748, 1217, 3894, 2886, 1760, 1348, 1828, 3902, 3468, 3406, 2250, 3904, 463, 3113, 3239, 3474, 2897, 1149, 2399, 1472, 3478, 2763, 2691, 1508, 2696, 2513, 3609, 1598, 1894, 1259, 262, 2530, 2933, 2736, 3585, 1135, 1053, 1544, 1985, 3915, 593, 1155, 552, 1350, 2400, 3621, 3674, 2657, 255, 3717, 3550]. Feature-Aware-Based [2763]. Feature-Based [2289, 3406]. Feature-Centric [2580]. Feature-Level [3860]. Feature-Weighted [593]. Featured [64]. Featured-Based [64]. Features [3347, 2173, 2899, 1181, 2547, 1022, 3318, 1564, 3882, 655, 3157, 3581, 1838, 3102, 1217, 2818, 3383, 467, 485, 1012, 1380, 872, 86, 1672, 1013]. Featuring [2772]. Febrile [3432]. Fed [3104, 3339, 2071, 1924]. Fed-Batch [3104, 3339]. Fed-DeepONet [1924]. Federated [2817, 2947, 1842, 1846, 3653, 3503, 1586, 2709, 1872, 3707, 2013, 3565, 2017, 1924, 1708, 3284, 3370, 1918]. Feedback [3211, 3215, 279, 935, 539, 197]. Feedback-Based [935]. Feeder [1216, 2597]. FEM [3831, 2139, 1191]. Fence [576]. FenceTalk [2552]. Fermentation [3339]. Ferromagnetic [1047]. Ferroptosis [3717]. Ferroptosis-Related [3717]. Fertilizer [3855]. Fetal [272, 12, 3739]. Few [1751, 3581, 3819, 2935, 2353, 3912, 3771, 3674]. Few-Shot [1751, 3581, 3819, 2935, 3912, 3771, 3674]. Few-View [2353]. FFT [644]. FHE [3732]. Fi [2044]. Fiber [3191, 2184, 3022, 2343, 1616, 660, 2414, 3353]. Fiber-Based [3191]. Fiber-Reinforced [1616]. Fibers [1092, 2166]. Fibrillation [1830, 2081]. Fidelity [3753, 3688, 3706, 2985, 1705, 1192, 1735, 1831, 1849, 1878]. Fidex [3372]. FidexGlo [3372]. Field [1677, 2938, 2774, 157, 1076, 49, 2658, 3299, 412, 2492, 2287, 1820, 33, 3913]. Fields [29, 278, 1589, 33]. Fifth [290]. Fifth-Order [290]. Fight [3796]. Fighter [3056]. figures [74]. Files [1093, 427, 1043]. Filion [1810]. Filling [2505]. Film [3235]. Filter [3639, 431, 746, 1357, 723, 893, 2900, 983, 830, 2635, 1855, 1836, 507, 620, 775, 1390, 52, 3676, 3465, 1001, 2790, 3402, 608, 309]. Filter-Based [1001]. Filtered [954, 463]. Filtering [3713, 3839, 2395, 2695, 1937, 1944, 3793, 370, 176, 1148, 372, 1116, 748, 3056, 2822, 377, 549, 3652]. Filters [1884, 557, 1599, 1107]. Filtration [1962, 1975]. Finance [813]. Financial [2446, 732, 1948, 3413, 1410, 704, 700, 3298, 745, 3769, 3526, 1459, 1315, 685, 2204]. Find [3077, 2556]. Finding [1020, 1168, 2172, 2111, 1145, 2627, 2883, 301, 2664, 2047, 323, 1665, 344, 2592, 1211, 1683, 767, 3392, 133, 3834, 2708]. Findings [2239]. Fine [3790, 1531, 3383, 1056, 3621]. Fine-Grained [3790, 1531, 3383, 1056]. Finger [3059, 1765]. Fingerprinting [3216, 3269, 1750]. Fingerprints [34]. Finish [3351]. Finite [152, 2292, 1052, 3816, 430, 3292, 1768, 3712, 3095, 2493, 1612, 1246, 360, 1736, 793, 52, 11, 481, 2644, 1222, 1616, 68, 731, 2861, 2758]. Finite-Difference [1736]. Finite-Time [1612, 1246]. Finitely [406]. FIR [296]. Fire [1552, 610, 235, 3260, 48]. Fired [3483]. Firefighter [576]. Firefly [1978]. Fireworks [476, 366, 432, 448, 346]. Firing [2016, 3391]. Firm [2155, 1899]. First [2406, 315, 388, 2950, 280, 1457, 1559, 3876, 2019, 80, 1033, 3645]. First-Order [1457, 3876, 2019, 1033]. First-Person [2406, 3645].

First-Year [1559]. Fish [1027, 691]. Fisher [2447, 2611]. Fissure [242, 1180, 819]. Fit [156, 112]. Fitness [2898, 3772]. Fitted [2239]. Fitting [252, 1119, 1834, 1046, 1661, 2804, 2682, 1574]. Fitting-Cauchy [1834]. Fixed [1139, 1020, 2467, 2903, 204, 3872, 944, 132, 360, 1572, 2312, 3469, 1745, 1471]. Fixed-Parameter [1139]. Fixed-Point [1020, 944, 1745]. Fixed-Shape [1572]. Fixed-Time [2312]. Fixed-Wing [3469]. Flag [962]. Flagellum [1785]. Flame [2767, 3551, 3106]. Flash [3502]. Flat [673]. Fleet [2443, 2394, 2940, 2823]. Fleets [3144, 2369]. Flexibility [3566, 347]. Flexible [1945, 2356, 3393, 3499, 2799, 2503, 3389, 2454, 1868, 3093, 2290, 696, 458, 987, 248, 2556, 3167, 1845, 1296, 2480, 2192, 724]. Flexible-Joint [696]. Flexible-Size [1296]. Flexure [275]. Flexure-Based [275]. Flight [1264, 1202, 2205, 1426]. Flip [2616]. Flipping [1582, 839]. Floating [374]. Flock [1900, 2829]. Flora [2676]. Flow [1360, 512, 491, 2107, 225, 243, 861, 2489, 1951, 574, 157, 1458, 3343, 1437, 3649, 1561, 2830, 3044, 844, 1530, 1658, 3106, 1689, 2348, 2139, 3901, 1221, 842, 1538, 1974, 996, 836, 1660, 2079, 599, 524, 3618, 2743]. Flow-Down [1530]. Flow-Level [2079]. Flow-Shop [1360, 2489]. Flower [3680, 2751, 1978]. Flowers [1478]. Flows [189, 503, 2370, 205, 1877, 168, 2133]. Flowshop [1054, 394]. Flowtime [394]. Floyd [3812]. Fluctuation [2050]. Fluid [3287, 2107, 3673, 2190, 1192, 2707]. FluidDynamic [103]. Fluidsim [103]. Fluorescence [67]. Flux [11, 3004]. Flux-Corrected [11]. Fly [224]. Flying [1551]. FMEA [3196]. FMEA-ROC-CoCoSo [3196]. fMRI [2313, 1799]. fNIRS [598, 2313, 601]. Focality [1773]. Focused [777, 1239]. FocusNet [3876]. Fog [2524, 3454, 684, 2027, 776]. Fog-Based [2027]. Fog-Computing-Based [776]. Fog-Haze [684]. Folded [2925]. Folding [21, 2364]. Folkcraft [1885]. Following [995, 3303, 2713, 2725]. Food [3777, 466, 3524, 2263, 2833]. Foot [3059, 1835]. Footing [1526]. FOPID [1232]. Foraging [548, 805, 3188, 3361]. Forbidden [1551]. Force [2695, 3760, 376, 412, 33, 3661]. Force-Directed [376]. Force-Field [412, 33]. Forced [827]. Forces [241]. Forcing [2333]. Forecast [663, 3702, 1411, 405, 1710, 3121, 3682, 741, 1331]. Forecasted [1292]. Forecasting [654, 3436, 2621, 2506, 1061, 1425, 147, 1074, 3868, 3076, 2315, 1491, 3722, 2504, 2830, 141, 2289, 3508, 2331, 2013, 403, 809, 735, 829, 1063, 3891, 3011, 2807, 3748, 3463, 530, 2135, 3555, 982, 745, 3769, 3526, 3910, 505, 2989, 3345, 501, 999, 1331, 3188, 1987, 3548]. Forecasts [3236]. Foremost [1957]. Forensic [3734]. Forensics [1580]. Forest [1552, 2975, 928, 3088, 3369, 1559, 2796, 3906, 2502, 2102, 2376]. Forestry [1531]. Forests [1578, 505, 227, 2361]. Foreward [1170]. Foreword [2228, 2041, 2508, 1, 1210, 2055]. FOREX [1128]. ForGAN [3509]. Forgery [2482]. Forgetful [2361]. Forging [807]. Fork [1585]. Fork-Join [1585]. Form [2616, 1958, 2800, 1807, 1755, 1333, 1574]. Formal [1991, 2329, 114, 1915]. Formation [1339, 1313, 3297, 3755, 2471, 3837, 3169, 3036]. Formed [1297]. Forming [2644]. Forms [723, 1077]. Formula [2616, 1976]. Formulas [3626]. Formulation [490, 311, 1877, 1537, 949]. Formulations [1430, 42, 2335]. Forums [950]. Forward [2679]. Foundation [2545, 754, 1495]. Foundations [2550, 160]. Founding [100]. Four [1425, 3074, 959, 501]. Fourier [1910, 2995, 2107, 2207, 155]. Fourth [1947, 754, 1058]. Fourth-Order [1947, 754, 1058]. Fovea [231]. FPGA [692, 1463, 1147, 772, 2224, 1588, 3894, 1446, 2193, 1547]. FPGA-Based [1588, 3894, 1547]. FPGAs [990]. FPT [2726, 1064]. FPTAS [1294]. Fractal [264, 489, 1567, 389]. Fractional [3492, 2757, 1126, 2039, 1143, 2618, 302, 2501, 2418, 632, 980, 2711, 667, 3647, 2355, 717, 3696, 2751, 616, 737, 621, 929, 2715, 3103, 626, 637, 3654, 2963, 900, 993, 634, 2677, 3641, 2946, 3486, 2877, 3559]. Fractional-in-Space [3103]. Fractional-Order [1143, 2501, 2418, 632, 2751, 2715, 637, 3641]. Fractional-Power [929]. Fractions [2465]. Fracture [3831, 3164, 3467]. Fractures [3040, 2125]. Fragile [433]. Fragment [297]. Frame [2973, 386, 1117, 3627, 358]. Frame-Level [2973]. Frames [1563]. Framework [1154, 1523, 2817, 3846, 2144, 2423, 2617, 2582, 2962, 440, 84, 714, 1062, 2473, 3656, 42, 3045, 3858, 242, 3864, 3382, 3139, 1803, 1103, 2249, 3500, 3761, 740, 3712, 2943, 3296, 3732, 622, 2704, 1720, 3826, 3318, 2431, 1247, 3148, 568, 3877, 2869, 1530, 3155, 3454, 429, 2015, 3888, 3514, 2997, 3785, 2386, 3891, 548, 3068, 1340, 3819, 846, 1852, 1822, 3595, 3900, 2598, 1816, 3274, 3407, 3469, 3525, 758, 3577, 79, 1827, 1979, 1771, 3177, 1194, 368, 1904, 2786, 1405, 1050, 3742, 2152, 643, 3715, 3913, 2509, 3916, 1484, 249, 3645, 3919, 3187, 815, 1643]. Frameworks [2884]. Frankot [3740]. Fraudulent [3240]. FRCNN [3190]. Fredholm [2119, 980, 482, 1537, 1807]. Free [151, 291, 1823, 2811, 955, 1264, 3199, 2093, 2167, 2175, 267, 2527, 2011, 462, 1912, 2555, 1095, 1374, 842, 1541, 1727, 337, 3539, 118, 3541, 3546]. Free-Interface [462]. Freedom [2333, 579]. FreeSurfer [3309]. Freeway [1514, 2381]. Freight [2674, 3111, 3571]. French [3726, 2225]. French-Speaking [3726]. Frequencies [3157, 2648]. Frequency [3018, 3713, 2618, 3262, 115, 858, 3774, 1499, 2903, 3753, 3826, 3744, 3022, 2594, 509, 2639, 3349, 23, 3029, 3896, 2343, 119, 2890, 2976, 2479, 3918, 1638, 3570, 1672, 3060]. Frequency-Based [3029]. Frequency-Domain [3018, 3826]. Frequent [740, 2700, 2265, 117]. Friction [788, 3667, 2509]. Frobenius [1788]. Frog [248, 346]. Frog-Leaping [248]. Front [2863, 3624, 3104]. Frontier [3075, 2367]. Frontiers [1170]. Frost [3693]. Frozen [341, 1676]. Fruit [224]. Fuel [531, 1018, 3325, 1001, 1972, 3667, 613]. Fulfillment [1451]. Full [2443, 1877, 358, 2679, 3913]. Full-Field [3913]. Full-Scale [358]. Fully [2367, 2900, 1392, 3225, 2978, 796, 1828, 1599, 1512]. Fully-Balanced [1599]. Function [3340, 1261, 3613, 3356, 131, 2544, 529, 2373, 207, 2800, 3694, 2920, 62, 1288, 2442, 2492, 3047, 180, 3798, 3350, 1416, 3802, 3392, 2333, 3243, 346, 3481, 2523, 326, 824, 155, 2572, 658, 377, 2872, 415]. Function-Based [207, 2800]. Functional [1154, 3040, 2545, 2313, 3805, 45, 1323, 601, 237, 3532, 527, 104]. Functionally [3343]. Functions [3579, 152, 424, 2418, 1138, 3312, 1205, 853, 1108, 1046, 2167, 1661, 122, 1722, 3587, 1165, 143, 2644, 1106, 2605, 1976, 3179, 3584, 3791]. Fund [1639, 1859]. Fundamental [2584, 1337]. Fundus [758]. Fungus [2426]. Further [68, 3125]. Fuse [2414]. Fused [1146]. Fusing [2126]. Fusion [2149, 2927, 3009, 2825, 235, 1811, 3678, 3744, 1859, 3763, 1723, 3157, 3282, 3457, 1531, 288, 1724, 3113, 3424, 2691, 3609, 1765, 1542, 262, 2736, 3319, 3655, 3776, 1130, 2237, 3725, 552, 752, 2400, 3249, 3767, 245, 3550]. Fusion-Based [1542, 3767]. Future [3566, 2536, 1759, 1021, 3751, 2587, 3201, 3653, 3806, 3604, 2735, 2440, 1005, 3124, 53]. Futures [3295, 3609, 501]. Fuzzification [1100, 1506]. Fuzzy [3492, 1154, 2321, 1031, 1934, 1550, 1732, 499, 476, 3579, 424, 632, 3438, 1945, 2356, 478, 2211, 753, 761, 139, 147, 484, 1670, 956, 2626, 1105, 443, 3792, 2396, 1096, 2234, 3389, 3144, 472, 1825, 3740, 1956, 863, 3449, 199, 696, 2633, 887, 3151, 1502, 1100, 1506, 690, 834, 3843, 773, 3573, 2715, 3459, 1085, 1285, 728, 1254, 1433, 2362, 494, 3234, 1883, 444, 392, 475, 1848, 468, 1284, 3048, 2122, 1057, 2646, 3402, 1482, 1745, 492, 254, 198, 833, 1028, 3245, 362, 1296, 2946, 1370, 3484, 3776, 407, 2192, 651, 357, 807, 883, 1350, 255, 703, 2797]. Fuzzy-Based [1254, 1057]. Fuzzy-Controlled [1031]. Fuzzy-Enhanced [3776]. Fuzzy-Logic-Assisted [3449]. Fuzzy-Logic-Based [254]. Fuzzy-Set [3459].

G [1810]. GA [976, 293, 1987]. GA-Adaptive [976]. GA-Reinforced [1987]. Gabor [3140, 2790, 1107, 86]. Gadolinium [1497]. GAF [3009]. GAF-Net [3009]. GAGAN [3219]. Gain [1524, 2883, 719]. Gaining [3650]. Gaining-Sharing [3650]. Gait [2043]. Galaxies [3651]. Galaxy [3651]. Galerkin [1052]. Gallup [3527]. Game [1677, 1268, 3656, 3061, 1741, 2050, 3615, 1881, 2131, 98, 502, 496, 1602, 768]. Game-Based [3061]. Games [1576, 1176, 3135, 595, 1351, 1319, 3222, 2058, 2563, 3358, 3679]. Gamma [1178, 2287, 2648]. Gamma-Field [2287]. Gamma-Ray [1178]. GAMS [2221]. GAN [3138, 2201, 1260, 2936]. Gangue [2736]. GANs [3686]. Gantry [2715]. Gap [2335, 2599, 2071]. Gapped [3212]. Gappy [1090]. Garbage [569]. Garden [818]. Gas [3668, 3599, 1087, 1955, 2425, 2997, 3121, 2428, 2137]. Gas-Gathering [2997]. Gasket [786]. Gaslike [3477]. Gate [2774, 1916]. Gated [3885]. Gathering [2305, 2997, 276]. GATransformer [3341]. Gauss [1132]. Gaussian [252, 3262, 1425, 1202, 1234, 3421, 2718, 1030, 2278, 2647, 70, 1706, 2572, 300, 377, 3270]. Gaussian-Kernel-Based [70]. Gaze [1069, 2165, 3790]. GBM [77]. GCE [3556]. GDUI [2782]. Gegenbauer [3647]. Gegenbauer-Based [3647]. Gel [3794]. Gel-Water [3794]. Gels [2806]. Gemini [3848]. Gen [3666]. Gen-AI [3666]. GenAI [3806]. Gender [1902, 2257]. Gene [1276, 3551, 867, 529, 2999, 1312, 1408]. General [2579, 3031, 2870, 3567, 1080, 2099, 2128, 1480, 2633, 2869, 2515, 1253, 1255, 569, 492, 3403, 1484, 2114]. General-Purpose [2128]. Generalised [250]. Generalist [3703]. Generalization [2148, 3122, 1819]. Generalized [3254, 302, 454, 2473, 1842, 1555, 3312, 954, 2112, 2355, 2377, 136, 568, 1032, 3731, 3228, 3654, 672, 351, 310, 1980, 1037, 1033, 949, 106, 1471, 533, 578, 402, 3126, 186, 1098]. Generalizing [2022]. Generate [3848, 2031, 3175, 3341, 2204]. Generate-and-Solve [2031]. Generated [1025, 1564, 2859]. Generates [2412]. Generating [2968, 320, 3727, 1801, 2435, 259, 2778, 2011, 2240, 2018, 1626, 1113, 3116, 3527, 1831, 3775]. Generation [1641, 3850, 3494, 3061, 1093, 2541, 17, 3423, 1161, 1556, 567, 1625, 3076, 2719, 2050, 3152, 3219, 3881, 2488, 1250, 2054, 3225, 2808, 3420, 2867, 3597, 1085, 2225, 3396, 1218, 2522, 2031, 2349, 3177, 1227, 3586, 2571, 3537, 904, 3568, 623, 2931, 3758]. Generative [1335, 3686, 2793, 3816, 2719, 3214, 3219, 2694, 3226, 2749, 3520, 3662, 2258, 3478, 2859, 3719, 674, 2777, 1260, 3300]. Generator [2626, 858, 1366, 1606, 3414]. Generators [2496, 2964]. Generic [173, 3591]. Genes [2490, 1016]. Genetic [3271, 654, 2732, 1732, 2172, 1945, 799, 127, 2767, 514, 787, 3599, 3495, 956, 31, 2724, 926, 3712, 3410, 2435, 2290, 3417, 3449, 2368, 1649, 1615, 3219, 913, 2014, 1568, 1365, 1689, 2348, 1897, 3606, 2768, 3406, 1845, 2605, 2029, 1322, 2772, 538, 3245, 2944, 3123, 2838, 1155, 3834, 2038, 1131, 1789, 1627, 2758, 553, 729, 44]. Genetic-Fuzzy [956]. Genetic-Whale [926]. Genome [3554, 3879, 3456, 3902, 2448]. Genomes [181]. Genomic [2917]. Genomics [1962]. Gentle [367]. Genus [3357]. Genz [1525]. Geo [441]. Geo-Clustering [441]. GeoAI [1004]. Geographic [3249]. Geography [3299]. Geohash [3136]. Geolocation [1002]. Geologic [3630]. Geomechanical [1098]. Geometric [2089, 576, 1589, 338, 3779, 3631, 163, 3489, 1492]. Geometrical [2127, 1611, 2723]. Geometries [1290, 3322]. Geometry [31, 2595, 2554, 3320, 2522, 480, 847, 3913]. Geopolitical [745]. GeoSpark [1124]. Geriatric [356]. Gesellschaft [387]. Gesture [2444, 852, 2232, 3715]. Gibbs [628, 2179]. GIN [3884]. GIN-Guided [3884]. Ginzburg [3728]. Girding [948]. GIS [3438]. GIS-Based [3438]. GISs [3280]. Given [259]. Glacier [2566]. Glaciers [2566]. Glands [2040, 2518]. Glaucoma [3716]. GLMI [3136]. Global [3080, 1164, 363, 915, 3372, 510, 1179, 1425, 127, 442, 3041, 3260, 313, 511, 853, 2047, 1953, 3872, 269, 2534, 2556, 3105, 2459, 3899, 869, 759, 3119, 3536, 1394, 2075, 3682, 824, 2611, 502, 1916, 415]. Global-Best [869, 759]. Globally [75, 1527]. Glove [3779]. Glowworm [805]. Glucose [107, 3347, 3900]. Glue [719]. Gluon [2427]. Glycemic [37]. GmbH [74]. GMHCA [3885]. GMHCA-MCBILSTM [3885]. GMRES [2923]. GNDO [1876]. GNN [3418, 3917]. GNN-Based [3418]. GNSS [2215]. Go [3313]. Goal [1941, 887, 1055, 1695, 1482]. Goal-Driven [1695]. Godunov [606]. Godunov-Based [606]. GOMORS [1650]. Good [3057].

Goodness [156, 112]. Goodness-of-Fit [112]. Goods [1362, 1569, 1977]. Google [3848, 3333]. Gout [2779]. Governance [1352]. Government [3849, 3591]. GPC [1120, 247]. GPGPU [3035]. GPN [2806]. GPS [13]. GPT [3672]. GPU [2695, 3637, 2688, 422, 2731]. GPU-Accelerated [2695]. GPUs [221, 1452]. Grade [3792]. Graded [3343]. Grades [2659]. Gradient [440, 1578, 3494, 3356, 322, 713, 567, 2664, 635, 1610, 2783, 2011, 3332, 1923, 3459, 3322, 1924, 3173, 1475, 1192, 3242, 2147, 2842, 1735, 2905, 1543, 3064, 477, 559, 780, 2877, 453, 664, 294, 1753]. Gradient-Based [567, 3322, 1924, 1192, 1735, 294]. Gradient-Boosted [1578]. Gradient-Free [2011]. Gradients [2399, 2936]. Grading [3062, 3595, 2754]. Gradual [3410, 198]. Grained [3790, 1531, 3383, 1056]. Gram [3402]. Gramian [622]. Grammar [120, 1045, 134, 2944]. Grammar-Based [134]. Grammar-Guided [2944]. Grammars [118]. Grammatical [3657, 3597, 2383, 2448, 1894, 3123, 2402]. Grams [46]. Granular [1021, 1305, 1752, 1006]. Granularity [2295, 3581]. Granulation [1177, 3866, 1752]. Granulation-Enhanced [3866]. Graph [3665, 1136, 61, 2226, 1646, 1911, 2538, 90, 1577, 3697, 3857, 1150, 3041, 3730, 1398, 1148, 156, 3496, 2750, 2367, 931, 1811, 1469, 2093, 1205, 312, 2628, 2883, 797, 1625, 2546, 1813, 3638, 1185, 2712, 942, 1631, 164, 183, 2052, 3278, 1251, 3812, 3390, 886, 1252, 2209, 2726, 376, 297, 1752, 2956, 1217, 2463, 361, 1095, 2062, 3240, 2823, 2507, 1162, 1494, 3528, 3719, 2607, 3341, 2786, 1671, 1511, 3917, 1473, 3249, 3652, 1603]. Graph-Based [3857, 164, 3719]. Graph-Retraining [2209]. Graphic [2073]. Graphical [2870, 935]. Graphs [166, 1824, 159, 1996, 624, 1353, 1269, 1148, 1145, 330, 1133, 320, 2403, 1583, 897, 3034, 2805, 2099, 571, 1585, 2769, 1415, 1957, 3098, 1665, 2773, 689, 1290, 3365, 1282, 3016, 3156, 1213, 3516, 135, 250, 1217, 3107, 2372, 796, 946, 495, 3522, 767, 1274, 2876, 1970, 2646, 1358, 1668, 1224, 1893, 165, 163, 677, 1660, 3539, 2, 3258, 3737]. Grasp [2897]. Grasshopper [885, 2482, 1038]. Grassmann [1114]. Gravitation [326]. Gravitational [558, 444]. Gravity [3608]. Gray [603, 2968]. Grayscale [2125]. Greece [3047, 3236]. Greedy [2404, 1418, 1327, 3087, 2457, 2664, 331, 1795, 2100, 2530, 1393, 549, 1316]. Greedy-Gradient [2664]. Greek [440, 3591]. Green [3651, 3261, 3441, 3384, 3243]. Greenhouse [2326, 43]. Greiner [3399]. Grey [2550, 3794, 327, 3288, 960, 468, 2747, 578]. Grey-Box [960]. Greywolf [2571]. Grid [2967, 3366, 63, 3423, 1204, 713, 2283, 234, 1792, 238, 3447, 3281, 2019, 3391, 3586, 2651, 232, 3484, 3652, 3300]. Grid-Based [238, 2019]. Grid-Connected [3484]. Grids [3292, 2503, 2371, 3573, 412, 1626, 804, 3418, 774]. GRM [2047]. GRNN [2209]. Grocery [1280, 1729]. Gromov [1579, 2214, 1146]. Grooming [1399]. Groove [552]. Ground [1834, 2505]. Grounding [2252, 3540, 3300]. Groundwater [1208]. Group [1993, 2108, 3729, 220, 2720, 1555, 225, 1780, 2293, 997, 1597, 1495, 407, 1926, 703]. Grouped [3813]. Grouping [1732, 3877, 346]. Groups [1889, 2050, 546]. Grover [3039]. Groves [937]. Growing [1406, 1180, 3222, 1655]. Growth [3817, 3338, 3014, 3538]. GRU [3705, 3828, 652, 3329]. Gu [2850]. Guangxi [3752]. Guarantee [2666]. Guaranteed [1496, 764, 1416]. Guaranteed-Cost [764]. Guarantees [1831]. Guessing [2172, 2210]. Guest [2912, 2228, 2041, 2508, 1210, 2055, 1170]. GUI [1804]. Guidance [2608, 1286, 1235]. Guide [2145, 1917, 598, 3677]. Guided [2584, 195, 683, 1127, 893, 2934, 3454, 1740, 3884, 1684, 3456, 3323, 2679, 3407, 1691, 3529, 2944, 3813, 2782, 2936, 3618, 3549]. Guidelines [871, 1854]. Gumbel [3920]. Gumbel-Based [3920]. GVNS [569]. GW [1590]. GW-DC [1590]. GWP [2064]. Gyro [766].

Haar [419]. Habitat [3299]. Habituation [3777]. Hadoop [725]. Haemoglobin [3327, 601]. Hair [3395]. Half [1243]. Halting [1117]. Halving [2100]. Hamiltonian [2881, 3878, 1665, 671]. Hamiltonicity [571]. Hammer [725]. Hammerstein [2119, 270, 477, 3486]. Hamming [450, 1364]. Hand [2957, 2444, 3779, 2897, 3645]. Handcrafted [3499, 2547]. Handling [3222]. Handoff [228]. Hands [2126]. Hands-On [2126]. Handwritten [1733]. Hankel [2083]. Happiness [3699]. Hard [512, 3367, 1299, 2633, 74, 1256, 1320, 106, 3771, 780]. Hard-Core [1320]. Hardness [1268, 2953, 1088, 1319, 1242, 1189]. Hardware [2341, 1463, 2910, 2453, 3597, 3894, 1101]. Hardware-Accelerated [2341]. Hardware-Based [1101]. Hardware-Software [3894]. Harmonic [1818, 305, 2934]. Harmonics [1792]. Harmony [753, 761, 1997, 374, 420, 392, 475, 1627, 3549]. HARP [893]. Hartmann [2420]. Harvest [1699]. Harvesting [1403, 837, 545, 769]. Hash [46, 537, 854, 3802, 1704]. Hash-Based [854]. Hashcash [2537]. Hashing [1817, 1161, 3058]. Hastings [1591, 3728]. Hate [3430, 1890, 3583]. Haul [2847]. Hauling [3005]. Hausdorff [2551]. Having [328, 455, 423]. HAVOK [3045]. Haystack [2117]. Hazard [3757]. Hazards [1249]. Haze [684]. HCV [1016]. HD [1243]. HD-Tree [1243]. Head [2924, 3885, 2978]. Headlines [3298, 2789]. Headway [2597]. Health [3346, 2249, 2795, 896, 3671, 1029, 3883, 1044, 3311, 2023, 1710, 2801, 997, 643, 3603, 3570]. Health-Related [2795]. Healthcare [1675, 1908, 1732, 3433, 3409, 3858, 2762, 3838, 2140, 3425, 3653, 3101, 3225, 1800, 3595, 3577, 2918, 3033, 1701]. Healthcare-Cloud [1732]. Heap [802]. Heart [3347, 2391, 2171, 272, 2638, 728, 2516, 12, 614, 2691, 3585]. Heartbeat [776]. Heat [2116, 1994, 2511, 3438, 2984, 1568]. Heating [3668, 2571]. Heavy [2295, 3263, 1362, 1777, 2847]. Heavy-Haul [2847]. Hedge [2996, 1859]. Helicopter [3087]. Helicopters [1388]. Helmet [2859]. Helmholtz [2757]. Help [419]. Hemodynamic [3380]. HER2 [2754, 2806]. Herd [1582, 308, 456]. Hereditary [1353, 3016]. Heritage [3895]. Hesitant [684, 651, 666]. Hessian [635, 359]. Hestenes [3633]. Heston [774]. Heterogeneity [3886, 1222, 1332, 2725]. Heterogeneity-Aware [2725]. Heterogeneous [668, 3351, 3730, 3576, 1509, 861, 1526, 301, 3292, 172, 306, 2697, 2309, 929, 11, 902, 804, 1539, 811, 3479, 3788, 1763]. Heteroscedastic [2322]. Heterostructures [2701]. Heuristic [2404, 1990, 1418, 783, 1327, 236, 1582, 2664, 1240, 586, 2666, 2334, 844, 226, 2534, 1515, 1613, 2866, 2459, 2887, 3167, 3754, 1618, 2963, 855, 2027, 1054, 1795, 2691, 81, 394, 3482, 3788, 952, 549, 873, 789, 3137]. Heuristic-Metaheuristic [2691]. Heuristics [1936, 2524, 478, 2127, 25, 1888, 1127, 574, 295, 2098, 948, 3523, 3170, 179, 1296, 206, 3549, 3252]. Heuristics-Guided [3549]. Hexadecimal [1584]. HFGAD [3790]. HFS [1462]. HFS-RF-PSO [1462]. HHL [3743]. HHT [3831]. HHT- [3831]. Hidden [490, 26, 2706, 79, 877]. Hierarchical [2457, 529, 1801, 570, 3790, 1022, 122, 450, 3760, 2997, 2437, 1222, 1121, 2339, 1700, 918, 1194, 10, 1346, 3542, 752, 3490, 411]. Hierarchical-Matching-Based [1022]. Hierarchies [940, 3107, 1412, 150]. Hierarchy [2809, 2658, 397, 2061, 2902, 594, 932]. Hieroglyphs [2162]. HIFU [2012]. HIFUN [1397]. High [3038, 464, 315, 1455, 240, 1525, 2467, 1112, 1234, 1308, 2999, 1607, 3735, 1051, 2741, 1243, 199, 3753, 3688, 1564, 822, 2784, 2014, 3706, 3256, 3690, 3785, 809, 3299, 3700, 3163, 23, 2337, 3748, 1532, 3463, 2190, 2139, 2674, 1851, 2985, 3414, 1466, 2684, 1441, 3789, 1914, 3370, 1711, 1735, 2976, 2835, 2233, 10, 719, 606, 3314, 359, 1545, 1484, 3818, 1839, 3616, 1926, 1878, 1141, 3546, 3490, 3001]. High-Alpha [2467]. High-Bandwidth [1308]. High-Dimensional [1525, 2999, 1607, 1051, 1243, 3256, 2337, 3748, 1735, 3314, 359]. High-Efficiency [3735]. High-End [1564]. High-Entropy [822]. High-Fidelity [3753, 3688, 3706, 2985, 1878]. High-Frequency [2976]. High-Gain [719]. High-Order [240, 606, 3616, 1926]. High-Performance [3038, 3463, 1466]. High-Power [3414]. High-Precision [3690, 809, 3700]. High-Quality [3490]. High-Resolution [3785, 1532, 3789, 1839, 3001]. High-Risk [2835]. High-Speed [3818, 3546]. High-Throughput [2784, 1711]. High-Value [2741, 2684]. Higher [3038, 3852, 3588, 1999, 229, 503, 2751, 3221, 2913, 3396, 865, 1831, 2743, 932]. Higher-Fidelity [1831]. Higher-Order [1999, 503, 2751, 3221, 865, 2743]. Highest [290]. Highlighted [2385]. Highlighting [3570]. Highly [3026, 3612, 1071, 3520, 2862, 3056, 2779]. Highway [3099, 3385]. Hilbert [3809, 1976]. Hill [1080, 3531]. Hill-Climbing [1080]. Hilliard [1112, 3617]. Hindmarsh [1783]. Hinge [799, 736]. Hinged [736]. HIPACO [3922]. Hippocampal [2909]. Hippocampus [3761]. Hires [2153]. Hiring [2153]. Histogram [2950, 2895, 673, 2992, 1476, 642]. Histogram-Based [642]. Histological [2625]. Histone [2484]. Histopathology [1529]. Historical [1880, 2219, 851]. History [1031, 3178]. History-Based [1031]. HITS [3841]. HIV [1779, 3019]. HMM [2256, 383]. HMM-Based [2256]. Hobbing [2266]. Hoc [647, 2702]. Hölder [280]. Hole [1353, 3801]. Holistic [738]. Hollow [2166]. Home [896, 1729, 3737]. Homes [3383]. Homogeneous [1889, 1660]. Homological [2367]. Homology [962, 1228, 2244]. Homomorphic [1828, 2327]. Homotopy [341]. Honey [3210]. Honeycomb [2861]. Honor [1210]. Hop [306]. Hopf [1228].

Hopfield [1615]. Hopscotch [1994]. Horizon [992]. Horizons [3205]. Horizontal [1425, 828, 1280]. Hormann [3399]. Horse [1760]. Hospital [1289, 1236]. Hospitalization [3810]. Hospitalized [1236]. Hosting [2618, 3366]. Hot [2759, 3553, 3609]. Hot-Rolled [3609]. Hotels [3553]. Hough [3062]. Hour [1425]. Hourly [1115]. Householder [464, 2233]. Householder-Like [464]. HPANet [3542]. HRIDM [3021]. HRRS [1817]. HSS [454, 3747]. HSS-YOLO [3747]. HT [3203]. HT-BES [3203]. Hub [2462, 1412]. Hub-and-Spoke [2462]. Hubs [3608]. Hull [739]. Human [1493, 3084, 220, 26, 1998, 282, 956, 3201, 2046, 1314, 3558, 823, 2486, 3313, 2892, 3506, 2942, 2761, 3785, 1214, 1188, 2880, 3462, 2528, 2023, 3110, 2332, 2160, 1746, 3239, 383, 3246, 1501, 2452, 3827, 1260, 411, 245, 2244]. Human-Algorithm [1501]. Human-Computer [3506]. Human-Defined [2046]. Human-Driven [3827]. Human-Engagement-Enabled [3201]. Human-Inspired [3084]. Human-Machine [956]. Human-Readable [3558, 2486]. Humanitarian [1596]. Humans [3759]. Hunting [3388]. Hurricane [2315]. HVAC [692, 2339]. Hyb [2870]. Hybrid [3028, 1934, 3412, 3196, 1388, 3850, 2811, 3366, 3083, 3705, 783, 512, 1944, 2268, 2298, 370, 266, 866, 1997, 225, 540, 2443, 1582, 1027, 3088, 2970, 2878, 1462, 2683, 473, 3277, 3206, 3074, 2370, 2489, 1237, 1428, 2799, 2420, 3795, 919, 926, 2496, 3445, 2454, 3094, 2175, 3417, 3722, 3873, 3321, 269, 2550, 3219, 1658, 2491, 2015, 3455, 3706, 3332, 1723, 422, 1776, 3581, 3891, 2638, 3104, 3106, 288, 2913, 3021, 2893, 2019, 3462, 2348, 3357, 2135, 3900, 1897, 2023, 358, 3606, 2250, 3904, 3525, 1059, 564, 3769, 758, 3801, 3474, 2772, 492, 2983, 3120, 3036, 1038, 1979]. Hybrid [2710, 3530, 628, 3329, 890, 2154, 3723, 2076, 425, 3319, 3811, 766, 3568, 599, 3919, 3834, 3648, 309, 762, 963, 2872, 3574, 1510, 3060, 1013]. HybridAlgorithm [1837]. Hybridization [1582, 3088, 2291, 1006]. Hybrids [2680]. Hydraulic [3287, 543, 3831, 425, 3484, 807]. Hydro [2048]. Hydro-Energy [2048]. Hydrodynamics [268]. Hydrogen [3441, 3455, 3332, 2908, 3905]. Hydrogen-Based [2908]. Hydrogen-Powered [3455]. Hydrothermal [1169]. Hyena [2348]. Hyper [1990, 1523, 1411, 2534, 1613, 3754, 3523, 3788, 873]. Hyper-Cube [1523]. Hyper-Heuristic [1990, 2534, 1613, 3754, 3788]. Hyper-Heuristics [3523]. Hyper-Parameter [1411]. Hyperbaric [3632]. Hyperbolic [1974, 239]. HyperDE [2534]. Hyperdimensional [1167]. Hypergraph [1577]. Hypergroup [3792]. Hyperloop [2674]. Hyperparameter [2391, 2129, 3446, 2527, 3828, 2350, 1914, 2100, 2402]. Hyperparameter-Free [2527]. Hyperparameters [1010]. Hyperparathyroidism [2518]. Hyperspectral [2994, 1004, 1950, 3840, 1245, 1215, 2562, 128, 339, 523, 1716, 2196, 3776, 3202]. Hypothenemus [3889]. Hypothesis [315, 388, 430, 1859, 775, 541, 1433, 712]. Hypothetical [2053]. HypPy [2994].

I-PD [2945]. I3D [1214]. I3D-Shufflenet [1214]. I4.0 [3442]. IaaS [3351].

IAS [976]. iCC [1050]. ICD [2801]. ICD-10 [2801]. Ice [3698]. ICEEMDAN [3548]. ICNN [1570]. iCOR [3465]. ICU [2223]. ID3 [515]. IDC [2806]. IDE [2039]. Idea [907]. Ideal [1709, 582]. Ideality [2458]. Ideation [950]. Identical [3582]. Identifiability [1503]. Identification [3025, 2341, 2670, 2149, 2896, 2360, 3009, 28, 1867, 431, 3091, 1077, 2140, 3304, 2526, 1428, 2436, 1619, 2476, 2454, 570, 418, 619, 795, 696, 3044, 1958, 3451, 1385, 296, 736, 265, 1253, 2866, 1255, 2986, 1724, 3113, 399, 1001, 1875, 681, 1915, 660, 16, 644, 2072, 2835, 2243, 2385, 270, 1543, 3661, 3585, 1135, 1053, 477, 1355, 1750, 3486, 705, 1870, 2346, 1604, 245, 294, 3189, 3250, 3251, 3739]. Identification-Oriented [2896]. Identify [2685, 1799, 1037, 486]. Identifying [2456, 1669, 3448, 1330, 3161, 1822, 2125, 3166, 2680, 1223, 3911, 3608, 3669]. Identity [2087]. Idle [3389]. IDS [1850]. IGA [2772]. IGLOO [2556]. Ih [1236]. IHO [3400]. II [3261, 3862, 521, 3450, 959, 1650, 786]. Iicu [1236]. IID [3707, 3370]. III [1395, 2701, 1379]. IIoT [2461, 1174]. ILC [2806, 616]. Ill [754]. Ill-Posed [754]. Illnesses [2986]. Illumination [3720]. Illusions [2687]. Illustration [356]. ILP [1483]. Image [3713, 3658, 1160, 2619, 1609, 2689, 3437, 3198, 3134, 866, 1233, 2825, 683, 1004, 1357, 123, 537, 2001, 2787, 1811, 607, 1161, 893, 1952, 410, 460, 3720, 2900, 2482, 2322, 48, 38, 2827, 3444, 3732, 466, 2741, 648, 3787, 3032, 199, 3753, 3826, 957, 2673, 944, 1748, 2143, 3154, 1780, 3022, 3219, 2009, 3766, 507, 1590, 1847, 3706, 413, 694, 2218, 2749, 3581, 1071, 3386, 1476, 1804, 3348, 2170, 989, 2846, 1966, 2125, 2564, 3407, 843, 451, 3589, 758, 3684, 3173, 1781, 2567, 3477, 2067, 2929, 1257, 83, 1107, 641, 262, 646, 961, 2653, 2179, 2836, 1012, 1484, 1887, 249, 1926, 2822]. Image [239, 1372, 3562, 433, 826, 53]. Image-Based [1609, 1811, 466, 2673, 3684]. Image-Guided [683]. Image-Matching [2822]. Image-to-Image [2001]. Imagery [2937, 390, 64, 2839, 2426, 1532, 2680, 2288, 2613]. Images [1662, 2973, 3629, 1431, 2695, 1937, 252, 57, 1911, 3437, 3596, 3804, 230, 242, 683, 2625, 1817, 1181, 1950, 157, 2005, 2547, 3446, 125, 1245, 619, 795, 1529, 3452, 2593, 2992, 3646, 1003, 1570, 2259, 2639, 3062, 2557, 2731, 2528, 3467, 3784, 2062, 2754, 128, 339, 2258, 2859, 2033, 2071, 918, 2914, 83, 3178, 937, 642, 1137, 1260, 819, 2782, 260, 3752]. Imaginary [137]. Imaging [3686, 220, 3749, 1929, 3497, 772, 2911, 2359, 2313, 2635, 1905, 1847, 167, 1912, 1215, 2196, 1405, 2574, 3670, 3300]. Imaging-Based [3497]. Imbalance [3289, 2647]. Imbalanced [2434, 862, 1305, 2148, 1743, 2779]. Imitating [3313]. IMM [3056]. Immediate [2125]. Immersed [1024]. Immigration [2290]. Immobile [1663]. Immune [370, 3796, 2965, 2675, 1602]. Immunoprecipitation [2484]. Impact [2732, 1937, 3851, 2446, 1552, 2623, 2084, 2295, 1653, 704, 3100, 52, 3459, 3898, 2294, 725, 3239, 3536, 2789, 2, 3914, 3916]. Impacting [2812]. Impairment [2460]. Impedance [2544, 832, 2604, 2126, 1495, 3300]. Imperceptibility [2689]. Imperfect [1586, 577]. Imperfections [1018]. Imperfectly [1138]. Imperialist [424, 146, 217]. Implantation [1967]. Implement [1120]. Implementable [709]. Implementation [3433, 2896, 169, 1552, 2585, 1147, 772, 3500, 739, 2545, 1917, 2934, 2139, 3069, 3472, 1191, 1595, 2529, 2287, 2681, 461, 1037, 1854, 3591, 3422, 273, 155, 1101, 2157, 1203]. Implementations [1994, 439, 1079]. Implementing [3141, 2960, 3355, 2243]. Implications [2626, 3151, 601]. Implicit [1874, 1373, 754, 1909, 247]. Importance [258, 3876, 3166, 2130, 3841]. Impossibility [3689, 3278]. Impossible [1669]. Impractical [2415]. Imprecise [1946]. Impregnated [1375]. Improve [2129, 2279, 2500, 1134, 1044, 2917, 1689, 2703, 83, 368]. Improved [3054, 1143, 2618, 3196, 305, 3259, 1327, 3037, 1581, 2767, 442, 816, 787, 2965, 3551, 3087, 2542, 3359, 2878, 2999, 3735, 2322, 2751, 3210, 1687, 3505, 446, 1246, 352, 456, 3565, 3887, 1368, 558, 828, 248, 2095, 1802, 3891, 548, 385, 3517, 3269, 470, 3164, 3288, 3265, 3234, 3354, 2062, 876, 3526, 1078, 3350, 3471, 1970, 3116, 2772, 2535, 2801, 3370, 2147, 644, 346, 2929, 3552, 247, 2102, 1394, 847, 824, 1660, 2654, 1479, 3183, 3841, 1197, 3818, 1056, 2303, 3188, 525, 575, 814, 1789, 1916, 2082, 2613, 3046, 354, 1421, 2829, 3922, 958, 2831, 479, 3127, 3251, 3300, 3291]. Improvement [3076, 1502, 432, 2098, 3396, 948, 515]. Improvements [283, 1054]. Improves [2895, 2862]. Improving [2375, 2195, 2001, 3091, 1266, 2666, 602, 177, 1267, 1891, 3022, 271, 2742, 58, 2641, 3467, 3784, 1534, 383, 1758, 2786, 3122, 3669, 2871]. Imputation [3371, 2505, 2095, 1851, 3603]. In-Context [3741]. In-Depth [3009, 3055]. In-Line [1798]. In-Process [2266]. In-Store [1593]. In-Transit [1402]. Inaccurate [2183]. Inapproximability [541, 712]. Incentive [75, 497]. Inception [2753]. Inception-Based [3021]. InceptionV3 [2023]. InceptionV3-SVM-Based [2023]. Incidence [3508]. Incidents [1156]. Incipient [1375]. Including [3366, 3104]. Inclusion [2446]. Inclusive [2889]. Income [1553, 2889]. Incomplete [797, 1452, 1256, 1470]. Incompressible [2370, 1877]. Inconsistencies [2061]. Incorporated [1278, 3603]. Incorporating [2496, 3875, 2469]. Increment [1935]. Incremental [1824, 84, 142, 2252, 3382, 3378, 3869, 3611, 680, 1064, 1668, 1397, 3674]. Incubated [1359]. Independence [207, 2980]. Independent [3411, 1996, 2236, 1282, 770, 2421, 112]. Indetermsoft [3155]. Indetermsoft-Set-Based [3155]. Index [3136, 659, 1717, 1243, 2013, 3844, 3843, 436, 349, 3770, 1548, 1066, 1310, 3117]. Indexes [708, 65]. Indexing [1381, 181, 572, 3212]. Indian [3683]. Indications [2113]. Indicator [20, 2568]. Indicators [3722, 1315]. Indices [3735, 498, 1663, 2479]. Individual [1630, 3536, 1137, 3550]. Individualized [2350]. Indonesia [1507]. Indoor [1995, 522, 3698, 3216, 3690, 3269, 2983, 2513, 2609, 3922]. Induced [3795, 3447, 3318, 3169, 504, 33, 712]. Induced-Dipole [33]. Inductance [2918]. Induction [55, 3414]. Inductive [839]. Inductor [3262]. Industrial [1908, 2461, 3204, 3092, 3144, 830, 3301, 1270, 2882, 3762, 2092, 2299, 3634, 2178, 2945, 1369, 2275, 3270]. Industrial-Scale [2299]. Industrialized [1122]. Industries [850]. Industry [2888, 3594, 3461, 2703, 1920, 855, 1781, 3666, 3121, 1467, 1301, 2755].

Inefficiency [1351]. Inequalities [2439, 1637, 2563, 1453, 525]. Inequality [3215]. Inequality-Based [3215]. Inertia [1995, 838]. Inertia-Weighted [838]. Inertial [2583, 3451, 2880, 378]. Inexact [2110, 1033]. Infarction [1497]. Infection [2108, 2473, 1779, 3702]. Infections [3403]. Infectious [3194]. Infer [226, 3858]. Inference [1843, 2416, 1386, 2229, 3263, 3390, 1312, 1853, 1162, 1137, 3728, 883, 3919, 1066]. Inferential [3699, 3000, 3258]. Inferring [1002, 2577, 3598]. Infinite [3139, 3444, 992]. Inflated [3598]. Influence [2501, 1327, 2720, 1861, 462, 1189, 1125, 2647, 3775]. Influences [3502]. Influential [1330, 1223, 3608]. Infomap [503]. Informatik [387]. Information [2579, 590, 2896, 2584, 50, 1177, 2474, 2434, 2279, 1624, 3059, 2589, 235, 3203, 922, 3866, 3052, 2334, 2447, 428, 1861, 2533, 2637, 3664, 1723, 3836, 1752, 3708, 2775, 2810, 1827, 1162, 936, 2333, 2440, 1542, 937, 3833, 100, 515, 847, 1543, 2237, 651, 3919, 2346, 2708]. Information-Based [50, 1624]. Informed [3072, 2511, 2388, 2277, 1657, 2643, 3586, 2074]. Informer [1851]. Informer-WGAN [1851]. InfoVis [88]. Infrared [314, 3676, 281, 1926]. Infrastructure [2536, 1552, 3897, 2535, 3591, 3351]. Infrastructures [734, 1708, 1700, 3479]. Ingestion [3274]. Inhomogeneous [429]. Initial [2780, 1272, 1363, 616, 2019]. Initialization [3689, 2133, 1414]. Initialized [2679]. Initiation [2414]. Injection [3773, 2547, 3565, 2601, 3546]. Injuries [1681]. Injury [2938, 2608]. Inner [720]. Innovation [2155, 2084, 3838, 3101, 2122, 3770, 477]. Innovation-Driven [3770]. Innovations [2536, 3653]. Innovative [2849, 3131, 2506, 2791, 3511, 2884]. Inpaint [1484]. Inpainting [1484]. Input [1234, 372, 2457, 3093, 1847, 3349, 1254, 254, 487]. Inputs [1388, 726]. INS [784]. INS-UWB [784]. Inscriptions [2831]. Insect [1046]. Inserting [682]. Insertion [844, 948, 394, 1379, 549]. Insertions [3623]. Insights [3855, 3055, 2013, 3016, 2846, 3899, 2202, 3245]. Insolation [1942]. Inspect [1049]. Inspection [2834, 3706, 3747, 422, 3762, 1547, 2860, 3189]. Inspired [1231, 1031, 3084, 556, 510, 236, 1167, 1928, 2496, 3590, 747, 2690, 2687, 1109, 470, 2737, 1008, 1605, 1257, 798, 1647, 2954, 1370, 2838, 1107, 502, 206]. Inspiring [3651]. Instability [20, 2648]. Instagram [3083]. Installation [1757]. Instance [1646, 1581, 1343, 2441, 1412, 1765]. Instance-Based [1412]. Instances [1823, 1306, 3367, 2011, 1129, 133]. Institution [2969]. Instruments [2275]. Insulator [3780, 3876]. Insulators [1687]. Insurance [1058]. Intangible [1635]. Integer [3013, 3853, 311, 1169, 1768, 2503, 3711, 132, 2956, 3392, 1467, 1707, 1725, 2877, 1066]. Integer-Fractional [2877]. Integers [908, 835]. Integrability [1865]. Integral [981, 3197, 3647, 482, 3694, 2205, 2963, 1537, 1807, 1755, 1975, 1725, 525]. Integrated [1908, 788, 78, 2431, 1293, 3514, 3895, 2598, 2032, 2399, 3532, 2612, 3670, 3632, 1461, 1342]. Integrates [803]. Integrating [3699, 2583, 2621, 3286, 370, 3059, 3861, 617, 335, 3209, 3445, 3746, 3231, 3662, 1777, 3474, 2818, 3033, 94, 5, 3280, 3910, 3682, 1832, 3685, 3765]. Integration [1641, 2864, 2221, 3860, 2373, 2503, 3740, 184, 2517, 3883, 935, 1876, 2887, 358, 2063, 3905, 938, 3405, 3329, 3070, 573, 724, 864, 3717].

Integrative [24, 655]. Integrator [3405, 3814]. Integrity [184, 3101, 3375]. Integro [980, 1783, 532, 438, 1594, 2119, 1191]. Integro-Differential [980, 1783, 532, 1594, 2119, 1191]. Intellectual [3430, 2006]. Intelligence [2614, 2341, 2988, 3430, 2189, 1992, 2518, 1010, 3198, 3286, 556, 3588, 139, 2155, 2587, 2762, 3861, 3075, 1462, 2911, 2922, 2476, 55, 2678, 1084, 3808, 3506, 269, 832, 1891, 1855, 2761, 308, 2995, 1912, 3311, 1890, 3337, 2248, 3461, 1822, 2972, 3662, 2918, 1898, 3117, 3906, 3280, 2945, 3403, 3745, 2440, 3591, 216]. Intelligence-Based [2614, 2972, 2918]. Intelligence-Driven [3906]. Intelligent [3713, 1798, 1850, 3432, 3822, 714, 788, 3059, 3204, 3074, 2671, 2436, 1586, 78, 3782, 927, 3760, 3747, 3325, 3887, 2181, 884, 3165, 2912, 3234, 475, 3472, 3801, 2030, 1472, 3535, 2608, 1521, 3775, 3635, 2854, 3487, 939, 3251, 3291]. Intensity [2938, 1863, 2343, 2202, 1773, 1331, 1737]. Intensive [2663, 2242, 3563]. Intent [3880]. Intent-Based [3880]. Intentions [1239]. Inter [13, 3501, 3213]. Inter-Domain [3213]. Inter-Observer [3501]. Inter-tropical [13]. Interacting [1730, 846, 33]. Interaction [32, 379, 3506, 2875, 146, 1684, 1570, 3600, 88, 527, 3765, 2273]. Interaction-Based [1570]. Interactions [1081]. Interactive [87, 2692, 3394, 1469, 2010, 3836, 3843, 1932, 3064, 2358]. Interception [892]. Interceptive [2934]. Interconnected [2963, 2651]. Interconnections [3302]. Interdependence [565]. Interdependent [2658]. Interest [3730, 441]. Interface [974, 2221, 2592, 462, 3843, 935, 3630, 2873]. Interfaces [2512, 956, 2594, 2889, 253]. Interfacing [1389]. Interference [115, 2424, 3511, 279, 2398, 119]. Interference-to-Signal [2424]. Interferometers [3191]. Interior [2960, 1311, 2358]. Interior-Point [1311]. Interlocked [131]. Intermodal [3571]. Internal [708, 498, 2707, 2479]. International [2318, 3580]. Internet [1908, 3286, 170, 2461, 3709, 1531, 3802]. Interpolating [102, 104]. Interpolation [484, 438, 3587, 994, 644, 2842, 2714, 3079]. Interpolations [3171]. Interpretability [3705, 3722, 1791, 3695, 960]. Interpretable [1317, 3342, 2533, 2014, 1324, 3232, 1028, 1747, 3919, 3679]. Interpretation [1529, 3877, 2807, 2204]. Interrupted [3851]. Interruptible [2956]. Intersection [3434, 702, 1925, 2595, 3690, 2492, 1336, 163, 2327, 3399]. Interval [3259, 1232, 2356, 484, 1105, 1825, 1957, 3098, 2800, 1390, 3891, 728, 3633, 1009, 475, 1848, 985, 597, 651, 525, 703]. Interval-Valued [3633, 1848, 651, 703]. Intervals [2116, 3711]. Intervention [3851]. Interventions [1939, 1854]. Intervertebral [2791]. Intra [3501, 2335, 319]. Intra- [3501]. Intra-Cluster [319]. Intra-Echelon [2335]. Intractabilities [163]. Intraday [3509]. Intranet [1738]. Intrinsic [960, 180]. Intrinsically [790, 1355, 1690, 1350]. Introduced [511]. Introduction [95, 850, 583, 367, 1422]. Introductory [3065]. Intrusion [3195, 3435, 2158, 1846, 3321, 458, 2371, 1857, 2721, 1708, 1472, 660, 3353, 2079]. Intuitionistic [499, 1096, 1848, 703]. Invariant [667, 1694, 2955]. Invariants [610, 2934]. Invasive [3347, 1686]. Inventory [2434, 2928].

Inverse [2116, 2811, 471, 3275, 2042, 2544, 3421, 209, 3094, 563, 1190, 2151, 1980, 2728, 3715, 3791, 239]. Inverses [1452, 310]. Inversion [1086, 2384, 1672, 3001]. Inversion-Based [3001]. Inverted [436, 3120]. Inverter [3248, 2320]. Inverter-Based [3248]. Inverters [2974, 623]. Investigate [50, 3380]. Investigating [2889, 1318, 2136, 2464, 3181, 2038]. Investigation [1908, 3847, 2044, 1779, 3502, 49, 3352, 1050, 1394, 3743, 1310]. Investigations [68, 3422]. Investment [3205, 779, 1241, 3361]. Investments [1815]. Investor [2077]. Involved [1555]. Involvement [1465]. Involves [2950]. IOb [1466]. IOb-Cache [1466]. IOCA2021 [2036]. IoMT [3033]. Ion [2156, 2021, 3555, 2072, 1005]. Ionized [532]. IOT [1094, 1991, 3435, 3013, 2242, 2524, 1182, 2224, 3369, 1560, 1632, 3216, 1838, 3893, 2198, 2561, 3815, 3577, 2027, 3329]. IoT-Based [2198]. IP [170]. IP-Based [170]. IPF [3594]. IPF-RADAR [3594]. IPMC [1785]. IPMC-Based [1785]. IPTV [2165]. Ireland [1869]. Iris [1509, 1722, 2749, 2170, 938]. IRONEDGE [2206]. Irradiance [1425, 2807]. Irregular [886, 3170, 1930]. Irregularly [397, 462, 3158]. Irregularly-Shaped [397, 462]. Irrigated [3202]. Irrigation [3815]. ISBN [74]. Ischemic [2828]. Isentropic [3612]. Ising [1430, 1076, 2297]. Ising-Based [2297]. Islands [3438]. Isocontouring [2740]. Isolated [702]. Isolation [1798, 3040, 2102, 467]. Isolator [939]. Isometric [1159]. Isothermal [861]. Isotope [28]. ISS [1985]. ISS-USC [1985]. Issue [698, 1922, 2189, 2228, 2619, 734, 2118, 2410, 971, 215, 2104, 556, 1481, 813, 1398, 1386, 3267, 387, 2002, 3006, 1273, 183, 880, 2271, 3450, 2504, 1210, 2451, 3268, 2055, 1213, 2205, 2596, 212, 1170, 2600, 2912, 2274, 1082, 2132, 1716, 2065, 2028, 1701, 2068, 1257, 2662, 618, 952, 1601, 2036, 2351, 2722]. Issues [2395, 3643, 1902, 1236, 3207, 3381, 634]. IT2FRAM [1105]. Italian [1808, 2177]. Italy [3438]. Item [3614, 1684, 2498, 2810]. Item-Choice [2498]. Items [117]. Itemsets [740]. Iterated [1990, 1936, 3087, 1393]. Iteration [1020, 2457, 1328, 532, 416, 488]. Iterations [310, 2147]. Iterative [1126, 353, 328, 455, 423, 1937, 324, 3820, 1349, 1947, 2246, 372, 861, 2365, 1779, 318, 343, 2143, 3154, 360, 3225, 1212, 285, 2920, 290, 413, 509, 675, 3692, 929, 686, 2556, 3163, 3691, 398, 341, 1453, 3641, 968, 2110, 2905, 299, 337, 2573, 705, 895, 294]. Itineraries [3362]. iTransformer [3548]. IWO [2772].

J [1810]. Jacket [3541]. Jacobi [1163]. Jacobian [3094, 341]. Jamming [3864]. Japanese [1885]. Jarratt [274, 1212]. Jarratt-Like [1212]. Jarratt-Type [274]. Java [1721]. JAYA [3131, 3036]. Jellyfish [3106]. JEM [874]. Jet [3683, 2208]. Jets [1072]. Job [1945, 1949, 2799, 3389, 3592, 2290, 2257, 1250, 987, 1568, 248, 2098, 3167, 2024, 1845, 3527, 947, 949, 1196, 1393, 724, 3252]. Job-Shop [1949, 2799, 2098, 3167, 947]. Jobs [574, 1664, 1561, 2480, 513, 1988]. Join [1577, 1585]. Joining [1917]. Joins [1787]. Joint [1086, 970, 2000, 3203, 1328, 3027, 696, 1007, 352, 1035, 1852, 852, 3112, 2665, 365, 3776, 1053]. JointContrast [2273]. Joints [1961, 736, 2576]. Journey [945]. Journeys [2018]. JPS [1802]. JSC [1552]. JSSP [1084]. Judgments [3759]. Jump [1612]. Just [574]. Just-In-Time [574]. JWL [3612].

K20 [1452]. Kadane [2591]. Kalman [3839, 3639, 431, 723, 2900, 983, 830, 1390, 52, 1001]. Kalman-Filter-Based [830]. KAN [3342]. Kanban [3592]. Kantorovich [482]. Kao [74]. Karush [2868]. Kazakhstan [3757]. Kd [1234]. Kd-Trees [1234]. KDAS [1385]. KDAS-ReID [1385]. Keeping [375]. Kernel [420, 3866, 2297, 3821, 3330, 3323, 1255, 70, 593, 249, 3487]. Kernels [2309, 2677, 2697]. Key [51, 2456, 1179, 1563, 1016, 3403, 1704, 270, 3669, 3046, 3570]. Key-Value [1179]. Keynesian [2363]. Keypoint [3785, 3631]. Keywords [2279, 1971]. KGs [2101]. Kibria [3421]. Kibria-Lukman [3421]. Kidney [3082, 794, 1215, 1907]. Kind [2053]. Kinematic [2785]. Kinematics [3275, 2042, 3094, 1190, 3715]. Kinematics-Augmented [3715]. Kinetic [1179, 568, 21, 3536]. Kinetics [2090, 1896]. Kiva [2446]. Kleene [1694]. KMC3 [1711]. Knapsack [1944, 2876, 3183, 3549]. Knee [3595, 2665]. KNN [3445]. KNN-Based [3445]. Knopp [1391]. Knowledge [668, 3650, 1654, 1670, 2219, 3703, 2330, 3503, 2590, 3876, 1566, 1385, 1629, 1536, 3478, 2786, 3368, 1443, 1764, 3685, 3249, 3258, 3737, 1600]. Knowledge-Based [3650, 2330]. Knowledge-Driven [1443]. Kohonen [587]. Kolmogorov [3806, 3345]. Koopman [3025]. Korovkin [3480]. Korovkin-Type [3480]. Kriging [1607, 2286, 1705, 2714, 3584, 786]. Kriging-Based [3584]. Krill [1582, 308, 456]. Krylov [2757, 3095]. Kubernetes [1833, 3576]. Kuhn [2868]. Kullback [2022, 2681]. Kung [353, 324]. Kuramoto [1019]. Kurdish [2949]. Kutta [2780, 3736, 754, 2634, 3221, 3896].

L1 [3813]. Label [2620, 426, 3793, 529, 345, 1485, 1912, 2982, 1789]. Label-Free [1912]. Label-Setting [2982]. Labeled [2099, 808]. Labeling [3203, 3014]. Labelled [320]. Labelling [2564]. Labels [2322, 1412]. Labor [1462, 3806, 1502, 3770]. Laboratories [2539]. Laboratory [2800, 3830]. LADTree [3063]. Lag [1037]. Lagged [481]. Lagrange [2103]. Lagrangian [519]. Lags [386]. Laguna [2315]. Laminates [47]. Land [5, 1130]. Landau [3253, 3728]. lands [13]. Landscape [3751, 2587, 2898, 1326, 1291, 3523, 3772, 5]. Landscapes [3566]. Landslide [3494]. Lane [2247]. Langton [1242, 106]. Language [3741, 3759, 2899, 3751, 2304, 2096, 3209, 3554, 2486, 570, 3569, 720, 1324, 1871, 3597, 2807, 770, 3396, 3110, 1808, 469, 3376, 2944, 3911, 674, 2789, 3715, 3543, 3648, 3679, 3628, 3717, 1200]. Language-Based [3679]. Language-Identification [570]. Language-Independent [770]. Language-Transformer-Enhanced [2486]. Languages [440, 2925, 355, 3697]. Laplace [3492, 155]. Laplacian [2757, 3200, 395, 2635, 3522]. Large [3741, 734, 783, 3751, 1145, 330, 126, 3209, 941, 3554, 2943, 1661, 1560, 1664, 3569, 957, 439, 2527, 1187, 1249, 1640, 3008, 2015, 3513, 2369, 3416, 979, 3021, 2804, 3396, 1396, 1475, 2507, 1914, 165, 2944, 3911, 1163, 3326, 10, 1394, 2075, 3913, 474, 1546, 1750, 3541, 3543, 3628]. Large- [1249]. Large-Scale [330, 1560, 2527, 1249, 2015, 3416, 2804, 1396, 2507, 1914, 3326, 1394, 2075, 474, 1546, 1750]. Largely [2421]. Laser [2420, 1703, 1035, 2343, 1749, 1985]. Lasso [1519, 3439, 3726]. Last [1263, 2661, 1292]. Last-Mile [1292]. Late [1080, 1786, 1497]. Latencies [3050]. Latency [975, 3709, 2286, 1634, 3050]. Latency-Bounded [975]. Latent [1290, 3360, 469, 1332, 1691, 2606, 3603, 3015]. Lattice [207, 259, 97, 2349]. Lattices [1592]. Law [3211, 3343, 2282, 2139]. Laws [447, 3644]. Layer [3847, 812, 3304, 2003, 2706, 3465, 365]. Layered [624, 3660, 52]. Layers [282, 1703]. Layout [3499, 2219, 2997, 1572, 1217, 1689, 1311, 1059, 1970, 1451, 2358, 724]. Layouts [3099, 1224]. LBMs [3207]. LC [2256]. LC-CRF [2256]. LCFRS [381]. LCMV [2468]. LCS [186]. LCZ [3899]. LCZ42 [3899]. LDPC [914]. Lead [3021, 1037, 2081]. Lead-Lag [1037]. Leader [3169]. Leaderless [2121]. Leading [3401]. Leak [3400]. Leakage [1642, 3004]. Leaking [2425]. Lean [2755]. Leapfrog [1994]. Leapfrog-Hopscotch [1994]. Leaping [248, 346]. Learn [2000, 2046, 2936]. Learnable [3415]. Learned [2342, 395, 3032]. Learner [1978]. Learnheuristic [2630, 2857]. Learnheuristics [726]. Learning [1406, 1675, 2614, 2973, 1677, 3629, 1498, 3193, 668, 2663, 3741, 3734, 2817, 3846, 2812, 3839, 2580, 2144, 1493, 2670, 2864, 2947, 3129, 2406, 3397, 1431, 2617, 2040, 1404, 1609, 1175, 3082, 2391, 3346, 2990, 2926, 3083, 3066, 2899, 1770, 3435, 1435, 915, 1395, 3259, 2621, 2301, 3132, 2622, 3852, 2171, 2292, 1678, 1429, 3133, 3013, 3409, 2304, 3333, 3768, 3061, 3085, 1061, 1581, 2434, 2888, 2624, 596, 1233, 2586, 3730, 2858, 3408, 2495, 2467, 2229, 2676, 3773, 3496, 1074, 862, 669, 2756, 2363, 1004, 1928, 2388, 2878, 1811, 922, 1127, 1628, 3413, 3709, 3140, 1526]. Learning [3205, 779, 3757, 1846, 2977, 3283, 2904, 1835, 2629, 3440, 1271, 2671, 3795, 3142, 1239, 2924, 1240, 2546, 926, 881, 1697, 1051, 3143, 1813, 2159, 2005, 2460, 711, 760, 751, 1954, 2745, 2090, 3446, 3750, 1025, 3799, 2700, 3653, 2315, 30, 2741, 1751, 3503, 3398, 1586, 1720, 3504, 3146, 1208, 1559, 3147, 1886, 2194, 3019, 1653, 2941, 619, 2969, 2449, 2709, 3557, 2289, 3321, 1144, 1925, 2094, 3508, 3214, 1959, 1488, 1075, 1872, 3307, 3744, 3217, 3453, 3218, 1754, 690, 839, 3880, 2441, 2285, 3057, 3881, 2331, 3454, 2533, 2992, 2097, 616, 3365, 2013, 3646, 2909, 2667, 2491, 715]. Learning [1173, 1278, 1905, 2371, 2505, 2694, 2776, 3611, 3565, 3683, 3886, 2369, 1060, 2856, 1063, 3746, 1044, 929, 3891, 548, 1838, 686, 2437, 3161, 680, 1323, 3102, 2259, 3068, 2639, 2426, 3062, 2350, 2555, 3269, 2742, 3824, 3892, 2329, 2557, 3011, 470, 1515, 3519, 3164, 1857, 3272, 3000, 3384, 3029, 1907, 3339, 2893, 1965, 2017, 2465, 2559, 3676, 770, 1681, 1464, 3607, 3462, 1814, 1277, 2021, 3357, 2445, 2135, 3583, 3595, 2735, 3900, 2490, 2310, 2136, 3017, 1708, 2528, 2060, 3110, 2643, 3555, 2332, 3168, 2319, 3467, 609, 2644, 3901, 1899, 2755, 2025, 1685, 1218, 1534, 1219, 2679, 2160]. Learning [982, 3358, 3662, 2210, 3112, 1082, 2132, 843, 745, 2115, 2645, 1695, 2066, 1023, 2665, 3589, 233, 3069, 865, 3577, 3237, 3687, 2602, 3684, 2939, 469, 1153, 1334, 2494, 3659, 2603, 2604, 1799, 3474, 1006, 93, 1256, 2151, 665, 2753, 2178, 1972, 2567, 3476, 1016, 492, 2727, 2680, 94, 114, 421, 1701, 2502, 2266, 2413, 973, 776, 1719, 587, 2763, 3810, 3242, 2691, 2859, 943, 1734, 1914, 1977, 3909, 1460, 3825, 966, 2033, 2798, 2744, 2323, 875, 3273, 3631, 2513, 2034, 3531, 2478, 950, 2354, 3910, 2570, 3532, 3911, 3533, 2408, 465, 2835, 2233, 1918]. Learning [2480, 2106, 1854, 2649, 3070, 3422, 2747, 2182, 2073, 2760, 2452, 2609, 3334, 3912, 3181, 1259, 1067, 3682, 1676, 1449, 674, 2152, 1679, 1688, 1826, 1738, 2241, 2717, 2748, 3400, 3353, 3811, 3788, 3485, 2836, 1137, 2871, 3914, 838, 1135, 1053, 1832, 3916, 1484, 1286, 3568, 638, 2079, 1809, 573, 806, 3919, 2487, 2971, 1198, 2180, 2361, 3187, 3648, 3188, 3771, 1638, 1690, 1712, 697, 1350, 1789, 1916, 1793, 1713, 2387, 2273, 2702, 2872, 3632, 3674, 3395, 8, 1882, 3679, 1693, 82, 1461, 2891, 1753, 3291, 1013, 1200, 3863, 2708, 3202, 3865]. Learning- [3583]. Learning-Architecture [2878]. Learning-Based [2812, 2947, 2990, 3435, 1678, 3773, 3709, 1526, 3446, 619, 1959, 2639, 3676, 3583, 2755, 2210, 2115, 2665, 3577, 2602, 3531, 1826, 838, 2971, 2180, 2702]. Learning-Driven [3188]. Learning-Enabled [1137]. Learning-Guided [1127]. Learning-Style [3187]. Learning-to-Rank [3214, 3568]. Leases [1002]. Least [113, 108, 3612, 372, 1077, 2851, 418, 1794, 2920, 2868, 3731, 541, 3676, 1715, 2232, 2220, 3486]. Least-Cost [3731]. Least-Mean-Square [1715]. Least-Squares [1077, 2851, 1794, 2920, 2868]. Leaves [2673]. Leaving [1365]. Leffler [3492]. Left [1655, 3598]. Left-Censoring [3598]. Legal [428, 738]. Legendre [3718]. Leibler [2022, 2681]. Leja [994]. Lemma [1583, 2227]. Lemma-Variations [2227]. Lemon [2196]. Lempel [116, 1409, 1587]. Lemurs [2817]. Lending [2446]. Length [3829, 868, 1585, 422, 640, 735, 3257, 1371, 2275, 118]. Length-Bounded [422]. Length-Dependent [118]. LEO [2197]. Lesion [2821, 2230, 2547, 3307]. Lesions [2060, 85, 2033]. Less [3664, 672, 3179, 733]. Lessons [1792, 395]. Lester [2987]. Lettuce [43]. Leukemia [2617]. Level [2973, 1276, 654, 152, 2508, 3860, 1169, 1880, 2462, 2219, 3678, 887, 1754, 3159, 728, 2913, 1464, 3298, 3229, 2598, 1724, 876, 1864, 665, 2964, 1844, 1765, 674, 1053, 623, 2079, 782, 724]. Leveling [949]. Levelized [3351]. Levels [2754, 10, 932]. Levenberg [406]. Leverage [2009]. Leveraging [2966, 3136, 1633, 2676, 3444, 1590, 2599, 3317]. Levitation [1143]. Lexico [2281]. Lexico-Semantic [2281]. Lexicographic [1345, 1640]. Lexicon [3296]. Lexicon-Based [3296]. LGNMNet [3178]. LGNMNet-RF [3178]. Li [2156]. Li-Ion [2156]. LIA [673]. libkww [155]. Library [3141, 1927]. License [2919]. Licensed [1328]. Lidar [1737, 3776]. Lie [2727, 1889]. Lienard [2380]. Life [783, 1702, 2795, 3514, 43, 3555, 1771, 517]. Life-Cycle [1771]. Lifecycle [3148]. Lifelogs [2113]. Lifetime [1418, 471, 1094]. Lifshitz [3253]. Lifting [1240]. Light [3714, 2673]. LightGBM [3419]. Lighting [2458]. Lightning [878]. Lights [1307]. Lightweight [3780, 3680, 2628, 2815, 3876, 3706, 3747, 134, 2310, 3802, 938, 2654, 3353, 3685, 3725, 1602, 3385]. Like [464, 305, 182, 1151, 1212, 2811, 3401]. Likelihood [2446, 167, 1396, 2509]. Likelihoods [2870]. Limb [2692, 3040, 2532]. Limit [3387, 2053, 716]. Limitations [2740, 2587, 1249, 3754]. Limited [2302, 2539, 3087, 3867, 2924, 1561, 130, 594, 3562]. Limits [2252]. Linac [2624, 2593]. Line [1798, 169, 3829, 431, 3567, 2219, 1687, 3303, 1645, 2284, 2597, 1749, 2183, 1706]. Line-Following [3303]. Line-Level [2219]. Lineages [2843, 1479]. Linear [1040, 906, 2896, 1042, 2418, 1159, 3612, 2439, 311, 1842, 147, 3136, 2588, 667, 2851, 1768, 1108, 2745, 2960, 528, 1612, 3215, 101, 1665, 185, 1663, 3058, 259, 616, 2998, 406, 1165, 2923, 3692, 2325, 929, 3624, 3519, 3310, 2804, 1852, 1255, 80, 2190, 1637, 1403, 405, 1594, 1537, 1453, 1333, 2605, 3392, 3480, 1226, 2842, 3814, 925, 1707, 270, 2530, 838, 359, 992, 2114, 3800, 3710, 524, 1372, 3489, 1547, 3015, 354, 3126, 3743, 1988]. Linear-Quadratic [992]. Linear-Time [1159, 80]. Linearity [102]. Linearization [1033]. Linearized [1141]. Linearly [1555]. Lines [2272, 1303]. Linesearch [3207]. Linguistic [1383, 536, 666]. Link [3504, 756]. Linked [517]. Linking [1767, 3473, 584]. Links [226]. Lion [1623]. Lions [63]. Liouville [1126, 3647]. Lip [2352]. Lip-Reading [2352]. Lipschitz [2563]. Liquid [2513]. Liquids [29]. List [3183]. List-Based [3183]. Listening [1951]. Listing [218]. Lists [202, 177]. Lite [3155]. Literary [2225]. Literature [3697, 2096, 2880, 3105, 2733, 3695, 2250, 3544, 3863]. Lithium [2021, 3555, 2072, 1005]. Lithium-Ion [2021, 3555, 2072, 1005]. Lithuanian [3162]. Liu [1475]. Lived [2466]. Liver [1867, 1392, 3446, 3501]. LLL [97]. LLM [3454]. LLM-Guided [3454]. LLMs [3298, 3911]. LMI [911]. LMS [723, 891]. Load [2618, 3599, 2970, 3443, 781, 138, 403, 1411, 3891, 2198, 3384, 1816, 3284, 1700, 2890, 3537, 3918, 2743, 3188, 786, 1638, 1987, 3548, 3060]. Load-Balancer [1700]. Load-Balancing [3443]. Load-Shifting [2970]. Loading [811]. Loadings [52]. Loads [2084, 1538, 1974]. Lobe [242, 1180, 2913]. Local [1990, 562, 1335, 3650, 1517, 315, 289, 388, 2085, 3372, 171, 3041, 3075, 3260, 1583, 192, 711, 483, 318, 1272, 2635, 1134, 185, 2054, 3390, 507, 1485, 640, 2556, 3163, 416, 1055, 1613, 495, 2105, 2186, 73, 2196, 1089, 3764, 968, 3370, 657, 1844, 2374, 3122, 3123, 2076, 2317, 3539, 1012, 2038, 197, 3137]. Local-Global [3041]. Local-Topology-Based [1134]. Localisation [2040, 3774]. Localised [1522]. Localises [2518]. Locality [688, 2]. Localization [522, 1560, 2094, 3216, 347, 1390, 920, 3462, 1277, 3475, 2072, 2609, 3319, 348, 231]. Localize [2639]. Locally [3356, 178, 1591, 359]. Locate [3481]. Locating [1415, 923, 2154]. Locating-Chromatic [1415]. Location [2423, 2792, 1321, 2272, 743, 3302, 1354, 1439, 1292, 986, 3118, 1296, 2, 2874, 873]. Location-Aware [2423]. Locations [2570]. Locator [3607]. Loci [1384, 190]. Lock [1262]. Locker [2661, 1152]. Log [2870]. Log-Marginal [2870]. Logarithm [2517]. Logic [2321, 1934, 2189, 424, 478, 1730, 3449, 3843, 1883, 392, 1284, 492, 254, 1296, 332]. Logic-Based [2189]. Logical [3727, 1483, 2951]. Logics [1536]. Logistic [3130, 141, 561, 2570, 2182, 1079]. Logistics [2485, 2633, 699, 2369, 2335, 1292, 1507, 1973, 3825, 3071, 1643, 3863, 3290]. Logo [2827]. Logs [1784, 3335]. Long [3436, 1519, 3857, 3749, 713, 1829, 1428, 2667, 620, 917, 3455, 640, 3161, 2813, 614, 1685, 2612, 2656]. Long-Axis [2667]. Long-Endurance [3455]. Long-Range [620]. Long-Span [1428]. Long-Term [1519, 614, 1685, 2612]. Longest [1665, 80, 186]. Longest-First [80]. LongestMatch [1316]. Longitudinal [3476, 3661, 3598]. Look [700, 3176]. Loop [981, 1524, 1086, 37, 2435, 563, 637, 1761, 3811]. Looper [2759]. Loops [2016]. LoRa [1560]. Lorenz [2296]. Loss [840, 3356, 1392, 3503, 1722, 3819, 1479, 3771, 2872]. Loss-Based [1392]. Losses [769]. Lossless [1041, 339, 1101]. Lossy [226, 451]. Lot [1080, 2737, 1845]. Lots [3427]. Lottery [1667]. Loudness [2885]. Louvain [3284]. Lovász [1583]. Low [1989, 3429, 3741, 1042, 2172, 1147, 3709, 630, 3593, 1877, 204, 371, 2286, 3373, 1564, 1024, 3344, 3386, 2556, 23, 2889, 2345, 3823, 2156, 1816, 2064, 351, 2191, 2385, 474, 2574, 2114, 1849, 1690, 873, 2083]. Low-Carbon [3344, 873]. Low-Code-No-Code [2191]. Low-Cost [1564]. Low-Dimensional [3593]. Low-Dose [3386]. Low-Energy [2556, 3823]. Low-GWP [2064]. Low-Income [2889]. Low-Latency [3709]. Low-Load [1816]. Low-Observable [3373]. Low-Orbit [2574]. Low-Order [2156]. Low-Power [1989, 1147]. Low-Rank [1042, 204, 351, 1849, 2083]. Low-Resource [3741, 474]. Low-Speed [1877, 2345]. Low-Stakes [2172]. Low-Variance [2385]. Lower [3040, 1127, 2913, 2335, 1412, 2532, 985, 2315]. Lozi [2553]. Lp [1926]. Lp-Pseudo-Norm [1926]. LPRM [2919]. LQR [539]. LR [381]. LS [224]. LS-SVM [224]. LSGO [1050]. LSM [1717]. LSM-Tree [1717]. LSTGINet [3390]. LSTM [3329, 3054, 3705, 2280, 2907, 1491, 2832, 2095, 3748, 3897, 2135, 2344, 681, 1331]. LSTM-Based [3748]. LTU [1853]. Lukman [3421]. Lumáwig [1209]. Luminance [3915]. Lumpability [2816]. Lumped [589]. Lunar [1436]. Lung [1180, 3086, 3446, 3501, 808, 188, 316]. Lungau [2502]. LungVision [2937]. Lyapunov [528, 2472]. Lyme [1259]. Lymph [1975]. Lymphatic [2251]. Lyndon [1206, 868]. LZ [427]. LZ-Compressed [427]. LZ77 [1295]. LZ78 [1295].

M [1210, 1334, 869]. M1 [1388]. M2Align [3859]. M2M [3174]. MAC [1739]. Macaque [2892]. Mach [3191]. Machine [1498, 2663, 3741, 3734, 2144, 2670, 1732, 2864, 3129, 3397, 1175, 3082, 3346, 3083, 3066, 2899, 1770, 3435, 1435, 1395, 2621, 3132, 3852, 2171, 1678, 3013, 2304, 3333, 3085, 1061, 2888, 3408, 2676, 3773, 956, 976, 1074, 2756, 457, 1127, 2724, 2394, 1119, 1526, 3757, 1835, 3440, 3795, 3142, 1240, 2546, 881, 1697, 2159, 2005, 711, 3093, 2700, 2315, 2290, 1208, 3096, 3019, 1653, 2449, 1144, 3508, 1959, 3217, 3218, 1754, 1029, 2331, 715, 2371, 2505, 3683, 3556, 3821, 2369, 3746, 1044, 548, 1838, 3161, 680, 3068, 2742, 3824, 1215, 2329, 3011, 1515, 3272, 3384, 3029, 3339, 2893, 1681]. Machine [1814, 3357, 3583, 3595, 2735, 3900, 2490, 2136, 3017, 2024, 3555, 2332, 3168, 2319, 3234, 2644, 1899, 1534, 3358, 3662, 843, 745, 2645, 1023, 3589, 3069, 3577, 3237, 3687, 2602, 3684, 2939, 2603, 2604, 2952, 1799, 2753, 1972, 2727, 2680, 1701, 2266, 2413, 776, 3810, 943, 1734, 1914, 597, 947, 2513, 2354, 3911, 3005, 3533, 2835, 2233, 2649, 2609, 3334, 3912, 611, 2076, 1259, 3682, 1676, 1449, 2152, 1679, 1688, 3400, 3485, 2871, 3914, 1544, 3916, 3540, 756, 1747, 638, 1809, 573, 3487, 2361, 3187, 3648, 3188, 1690, 549, 1350, 2387, 2176, 8, 411, 82, 2891]. Machine [1753, 3863, 2708, 3202, 3865]. Machine-Based [3540]. Machine-Learning [1208, 2449, 2490, 1023, 2891, 3202]. Machine-Learning-Based [2315, 2371, 2505, 2319]. Machine-Processable [1029]. Machinery [1692, 3096, 620, 2514, 2387]. Machines [1928, 3619, 1278, 375, 2198, 56, 3677, 587, 599]. Machining [2150]. Macrocephalus [70]. Macromolecular [36]. Macromolecules [412]. Macular [795]. Made [672, 1862]. Madre [2315]. MADRL [2693]. Magnetic [1143, 57, 1062, 2313, 3004]. Magnetization [1076]. Magneto [2389]. Magnetohydrodynamics [3044]. Magnetorheological [939]. Magnetotelluric [731]. Magnitude [3657, 1137]. Magnus [98]. Mahjong [2318]. Maintenance [2614, 654, 1127, 2290, 3663, 3068, 1153, 1334, 2480, 2428, 2137, 2874, 2417]. Major [229, 160, 3065]. Majority [2060, 3119]. Majority-Minority [3119]. Make [1058]. Maker [2578]. Makespan [844, 1054, 947]. MAKHA [269]. Making [590, 3729, 2975, 1481, 787, 3408, 3209, 3264, 1096, 1528, 1502, 2863, 3512, 2471, 1359, 684, 1283, 560, 943, 536, 2570, 533, 1738, 407, 651, 666, 2499, 703, 2245]. Making-Pareto [2863]. Malaria [1277]. Malaysia [3188]. Malfatti [3031]. Malicious [3607]. Malignancies [3086]. Malmquist [3770]. Malware [655, 1931, 1534]. MAML [3912]. Mammogram [54]. Mammograms [86]. Mammographic [140]. Mammography [2567, 53]. Man [177]. Man-Optimal [177]. Manage [2311]. Management [1549, 3851, 734, 3681, 2962, 3366, 702, 3037, 2585, 2215, 2676, 1998, 3425, 1182, 3442, 3144, 1046, 2548, 978, 3043, 3148, 3213, 3671, 1925, 1029, 2784, 2015, 3282, 1044, 1804, 3606, 3466, 1746, 1059, 564, 2222, 3634, 1454, 2535, 2030, 2928, 3403, 2440, 3591, 3361, 3574]. Managing [336, 321, 738, 187]. Mandarin [1764]. Mandatory [3362]. Maneuvering [3056]. Maneuvers [3185]. Manifestations [2828]. Manifold [930, 915, 1233, 2564]. Manifolds [2200, 1696, 1114, 973]. Manipulation [200]. Manipulator [2476, 404]. Manipulators [3275, 3094, 810]. Manoeuvres [2467]. Manual [3501]. Manufactured [1862]. Manufacturer [1979]. Manufacturing [3359, 3267, 3499, 1611, 3093, 1703, 1250, 3883, 621, 1967, 3274, 2066, 3666, 588, 599, 1600]. Many [203, 63, 406, 2928, 2944, 2037, 479]. Many-Core [203]. Many-Objective [2928, 2944, 2037, 479]. Map [303, 2774, 1183, 157, 2047, 561, 2553, 2521, 2874, 733, 1372, 2256, 918, 740]. Map-Reduce [740]. Mapping [3855, 3075, 503, 695, 893, 190, 1913, 1046, 2741, 45, 1215, 3899, 2772, 3630, 2287, 937, 642, 2905, 3383, 3913, 3538]. Mappings [360, 398, 1745, 1471]. MapReduce [905, 330, 743, 749, 910]. Maps [3494, 636, 15, 1956, 1502, 2686, 1671]. MAPSkew [749]. Margin [1722, 3771]. Margin-Aware [3771]. Margin-Based [1722]. Marginal [1910, 2870]. Marine [1087, 3783, 3068, 1766, 594, 1608, 3667, 613]. Mario [2964]. Maritime [771, 1482, 2854, 496]. Markers [1645]. Market [2621, 669, 3205, 3806, 704, 585, 2599, 3117, 3723, 453, 768]. Marketplace [1750]. Markets [1992, 1517, 2456, 732, 669, 207, 1863, 745, 3770]. Marking [3631]. Markov [490, 373, 1339, 1709, 3135, 2165, 26, 1928, 3673, 1528, 2658, 1612, 775, 3229, 2768, 79, 3659, 3764, 877, 589, 3537, 2428, 2137, 964]. Markov-Modulated [3673]. Markovian [3702, 3560, 877]. Marquardt [406]. Marriage [185]. Mars [3528]. Martensite [3368]. Mask [2375, 1819, 2078]. Mask-Based [2078]. Mask-Wearing [2375]. Masked [3578, 2514]. Masking [1413, 3398]. Masks [2717]. Mass [2109, 2208, 3901, 44]. Massive [1014, 1167, 1133, 945, 1813, 2344, 1733, 365, 2577]. Massively [3266, 1980]. Master [2469]. Match [1358]. Matching [215, 3793, 976, 192, 202, 1381, 1328, 817, 3212, 36, 1022, 3507, 110, 2429, 2992, 2009, 458, 422, 121, 909, 263, 1440, 200, 3402, 65, 1930, 1597, 712, 2822, 2860, 3126]. Matching-Iteration-Based [1328]. Matching-Widths [712]. Matchings [178, 205, 177, 214]. Material [3612, 2330, 1703, 3343, 49, 3114]. Materials [772, 863, 3211, 2465, 1283]. Mathematical [2501, 424, 1553, 3858, 2791, 2545, 1825, 3830, 7, 2335, 2166, 40, 2414, 2848]. Mathematics [3441]. Mathematics-Driven [3441]. Matheuristic [970, 1080]. Matheuristics [1556, 1196]. MATLAB [2221, 3500, 2492]. Matrices [2411, 1921, 1470]. Matrix [3254, 1788, 490, 203, 1205, 3500, 2688, 1652, 2517, 3215, 2447, 2052, 2097, 800, 1337, 1740, 2270, 1285, 3691, 2344, 1453, 310, 2569, 591, 1542, 1379, 391, 555, 2157]. Matter [3837, 2077]. Maturity [1057]. Max [1751]. Maximal [426, 386, 1206, 1045, 2923, 767, 2648, 2397]. Maximal-Sum [2397]. Maximise [2931]. Maximising [1195]. Maximization [574, 1125, 2029, 3184]. Maximizing [1418, 1011, 2940, 2837]. Maximum [166, 303, 1942, 159, 340, 178, 3867, 1813, 3872, 2591, 185, 167, 541, 193, 54, 2459, 1396, 767, 153, 1001, 504, 1193, 3723, 712]. MaxSAT [1363]. Maxwell [1938]. May [3703]. Mayfly [2320]. Mayfly-Based [2320]. Maze [3061, 2720]. MBCC [3841]. MBCC-HITS [3841]. MCBILSTM [3885]. MCDA [229, 1651, 160]. MCDA-Based [1651]. MCFN [3776]. MCMC [3878]. MCNN [2081]. MCNN-BLSTM [2081]. MCSA [1816]. MCSA-Based [1816]. MDAN [1003]. MDAN-UNet [1003]. Mead [3128, 1470]. Meals [3903]. Mean [1677, 305, 1946, 2430, 1696, 3172, 1715, 2220, 536, 533, 1500, 249, 2180]. Mean-Field [1677]. Mean-Variance [1500]. Means [1382, 3192, 1034, 1744, 3438, 1266, 687, 1304, 1100, 1506, 3381, 507, 314, 2022, 1721, 2641, 3784, 3114, 3173, 2032, 2232, 2714, 1630, 890, 2038].

Measure [1049, 476, 443, 3281, 3166, 2366]. Measured [2791, 3779]. Measurement [3629, 38, 204, 2430, 3844, 2344, 1432, 3391, 2072, 703]. Measurement-Type [3629]. Measurements [1633, 521]. Measures [1110, 3729, 3259, 2585, 2021, 1795, 2088]. Measuring [704, 3102, 2640, 1756, 2475, 3258]. MEC [1186]. Mechanical [511, 2454, 3234, 2139, 2389, 1698]. Mechanism [1551, 2419, 1440, 609, 1851, 1595, 1927, 1487, 516, 2654, 3917, 3725, 549, 1456, 497, 1604]. Mechanisms [170, 2117, 2981, 553]. Media [1744, 2621, 2107, 861, 2482, 704, 1890, 1008, 950, 2354]. Medial [3798]. Median [1357, 1378, 3257]. Medical [1662, 3686, 2581, 1790, 2765, 2825, 3408, 3564, 3703, 3444, 2635, 3452, 1748, 1044, 1804, 3407, 843, 1710, 3376, 3329, 1405, 904, 3562, 2655]. MedicalSeg [1804]. Medicine [3797, 2800, 3070]. Medium [1641, 3263, 3831, 1426, 3368]. Medium-Carbon [3368]. Medium-Scale [1426]. Medium-Voltage [1641]. Medoid [2432]. Medoid-Based [2432]. Medoids [921]. Meet [1579]. Meeting [1583]. Meets [2245]. Mel [1672]. Melanoma [3803]. Mellin [1946]. Melting [1703]. Members [1601, 2722]. Membership [355, 1853]. Membrane [2528]. Membranes [2010]. Memetic [1073, 2489, 172, 1287]. Memory [3650, 3436, 2720, 139, 3749, 503, 713, 1829, 1428, 967, 917, 2813, 1419, 3602, 1599, 2275, 2656, 2860, 317]. Memory-Based [3650]. Memory-Efficient [3602]. Memory-Enhanced [2720]. Memristive [3875]. MemristiveAdamW [3875]. Mendell [1525]. Menopause [2670]. Mental [3671, 889, 1155]. mer [1380]. Mergers [1396]. Merging [3099, 1674, 1589, 1196]. Mesh [636, 409, 1606, 2254, 2875, 923, 3823, 2905, 1347]. Meshed [2221, 1964]. Meshes [2007, 2214, 2826, 1574, 3114, 731]. Meshfree [3072]. Meshing [76]. Message [2704, 2592, 402]. Messages [1540]. Messy [2967]. Meta [478, 25, 236, 1004, 2666, 2378, 2289, 1434, 3886, 1684, 2459, 2027, 3403, 1630, 3482, 206]. Meta-Analyses [1630]. Meta-Analysis [3886, 3403]. Meta-Ensemble [1004]. Meta-Heuristic [236, 2666, 2459, 2027, 3482]. Meta-Heuristics [478, 25, 206]. Meta-Learning [2289]. Meta-Path [1684]. meta.shrinkage [1630]. Metabolic [2993]. Metabolite [1062]. Metabolomics [44]. Metaclassifier [3049, 2771]. Metadata [3742]. Metaheuristic [3347, 803, 1517, 3287, 2416, 225, 3089, 3074, 1237, 2799, 2496, 3097, 3153, 1187, 834, 2866, 959, 1900, 1761, 749, 1059, 1749, 2691, 2070, 1478]. Metaheuristics [1010, 2993, 329, 2476, 1561, 1562, 1596, 3764, 605, 3531, 3568]. Metal [3883, 1505]. Metamodeling [2698]. Metamorphic [2185]. Metatheorems [367]. Metaverse [2882]. METE [3316]. Meteor [2531]. Meteorological [2505, 3748]. Meter [1564]. Method [1289, 455, 423, 3193, 3614, 1935, 1493, 291, 2518, 315, 289, 454, 1938, 283, 3729, 2821, 298, 1519, 50, 2840, 2280, 2950, 431, 322, 3749, 1670, 1998, 243, 1004, 519, 3138, 1784, 280, 3304, 3680, 3378, 713, 1183, 3866, 2904, 1305, 2664, 1834, 1238, 673, 1450, 2671, 3720, 2477, 1607, 3735, 2359, 754, 274, 1801, 2547, 3712, 2322, 2827, 3263, 3443, 323, 472, 2454, 2960, 2519, 572, 1207, 3145, 3678, 2549, 1661, 648, 2175, 2590, 3095, 3504, 2517, 1610, 1794, 3097, 333, 2286, 3874, 1047, 3149, 3321, 3213, 267, 3702, 1134, 944, 3253, 839, 2764, 3879, 3022, 3882, 2783, 3766]. Method [3224, 438, 2729, 39, 923, 927, 1477, 1411, 1905, 1775, 2505, 2776, 3455, 3513, 3556, 3821, 290, 2181, 3420, 406, 413, 675, 736, 621, 828, 3625, 2277, 1070, 3458, 11, 481, 2198, 3310, 1966, 2019, 3897, 626, 744, 730, 3654, 1113, 1311, 2332, 358, 652, 3466, 3902, 1866, 539, 900, 1120, 2565, 3904, 2119, 341, 2494, 2122, 3475, 1475, 2349, 936, 594, 968, 351, 2779, 2110, 1225, 1820, 644, 487, 1303, 3552, 877, 1033, 1163, 3813, 2890, 2154, 2073, 3775, 2905, 2472, 445, 2392, 3319, 3584, 3661, 3791, 904, 407, 2874, 488, 593, 3540, 3186, 3724, 332, 2574, 3685, 3710, 573, 666, 3486, 2854, 2971, 402]. Method [2379, 1331, 3618, 1712, 899, 1789, 2387, 2613, 3395, 1421, 664, 841, 3189, 3127, 3251, 1066, 1693, 2725, 3300, 1230]. Method-Signal [3475]. Methodological [2084, 59]. Methodologies [3286]. Methodology [3432, 3500, 919, 3209, 2049, 78, 1144, 1102, 1651, 3468, 2389, 1781, 2648, 3537]. Methods [3316, 2321, 2184, 353, 328, 2580, 3129, 1759, 464, 302, 2085, 2899, 305, 324, 1052, 3198, 2740, 3051, 264, 2229, 1947, 1074, 1027, 511, 3647, 2370, 3207, 3736, 1686, 409, 2546, 1281, 2865, 3301, 2634, 3878, 1919, 832, 3744, 3216, 3218, 3831, 2331, 3221, 1434, 1212, 1812, 3700, 2218, 1923, 3103, 2170, 2846, 3337, 3896, 3837, 2248, 2730, 3461, 3462, 2516, 2208, 1618, 3721, 2177, 2186, 2274, 2063, 2665, 1537, 341, 3239, 3115, 2566, 1416, 2151, 2178, 3641, 1668, 3177, 2142, 2484, 3535, 299, 337, 1427, 2575, 872, 3489, 1690, 1882, 3250, 2708]. Methods-Based [1618]. Methylation [1167]. Methylcytosine [1543]. Metric [303, 3409, 2279, 410, 460, 1889, 101, 1075, 1348, 1332, 3402, 292, 1541]. Metric-Driven [3409]. Metrical [1337]. Metrics [1142, 2281, 1321, 2791, 2378, 144, 1540]. Metro [580]. Metropolis [1591, 3728]. Mexico [3440, 855, 2979]. MF [3072]. MF-VPINN [3072]. MGKGR [3249]. MHD [3771]. MHD-Protonet [3771]. MIA-LSTM [2135]. Mice [2909]. Micro [1798, 3366, 42, 1714, 38, 1749, 3178, 2931]. Micro-Bubbles [1714]. Micro-cracked [42]. Micro-Expression [3178]. Micro-Scale [1749, 38]. Microalgae [993]. Microarray [867, 125]. Microblog [364]. Microcalcification [879]. Microgrid [3681, 3363, 3288, 3265, 3905, 2797]. Microgrids [3394, 3783, 2908, 1987]. Microphotograph [1867]. Microrelief [1297]. MicroRNAs [1380]. Microscope [2581]. Microscopic [1104, 3755]. Microscopy [2695, 1912]. Microservices [2524]. Microwave [71, 2407, 1405]. Midpoint [3690, 3001]. Might [2077]. Migrating [987]. Migration [3027, 3373, 380, 759, 869]. Migratory [3299]. Mild [2460, 3632]. Mildly [355]. Mile [2661, 1292]. Military [1998]. Milling [2861]. MILP [949]. MIMO [3688, 1905, 2344, 890, 384]. Min [3179]. MINC [1624]. MINC-NRL [1624]. Mind [2415, 2151]. Mineralization [3320]. Ming [74]. Ming-Yang [74].

Mini [38, 2531]. Mini- [38]. Mini-EUSO [2531]. Miniature [1868]. Minimal [95, 120, 386, 667, 3575, 3364, 3521]. Minimalist [3868]. Minimally [1269]. Minimax [3378]. Minimising [1195]. Minimization [3612, 853, 1923, 1470]. Minimize [1644, 947, 513]. Minimizing [1430, 1945, 225, 2424, 2903, 3619, 3889, 1634, 3050, 2797]. Minimum [1327, 311, 2669, 473, 978, 3872, 177, 2915, 210, 1562, 2729, 3365, 338, 2998, 541, 276, 3107, 3778, 111, 3167, 3723, 3539]. Minimum-Cost [2669]. Minimum-Energy [3167]. Mining [1757, 89, 2765, 1581, 2165, 1133, 1784, 1103, 2685, 3795, 434, 1273, 3296, 2486, 356, 704, 738, 3881, 1490, 3220, 1800, 3901, 1848, 2816, 1777, 1065, 198, 628, 2608, 1427, 2833, 3635, 3920, 3771]. Minipig [2199]. Minkowski [1305]. Minority [3049, 2771, 3119]. MinSum [3871]. miRNA [527]. MIRROR [2084]. Misalignment [467, 999]. Mish [3350]. Mislabeled [3569]. Mislabeling [1713]. Mismatches [263]. MISO [746, 296]. Missile [892]. Missiles [3530]. Missing [3371, 1794, 957, 2505, 2095, 2560, 1851, 3603]. Missing-Data [3371]. Mission [2834, 1097, 2574, 1436]. Mitigate [2537]. Mitigating [3085, 3307, 3511, 2887]. Mitigation [71, 115, 3864, 119]. MITRE [1913]. Mittag [3492]. Mittag-Leffler [3492]. Mixed [42, 921, 3051, 1169, 667, 329, 2503, 1477, 2956, 3798, 1432, 1494, 1707, 1620, 3833, 467, 3487, 1725, 2083]. Mixed-Domain [467]. Mixed-Integer [1169, 2503, 1725]. Mixers [1806]. MixFormer [2254]. Mixing [1620, 719]. Mixture [312, 2328, 2601, 1222, 2278, 2647, 3827, 1986, 300]. Mixtures [1558, 2838]. Mixup [1764]. Mixup-Based [1764]. MKD [1764]. mKdV [1372]. ML [1549, 3577, 3911, 3620]. ML-Powered [1549]. ML-RASPF [3577]. MLlib [3326]. MLMVN [3018]. MLMVN-Based [3018]. MLP [3047]. MM* [2057]. MMD [2844]. MMD-MSD [2844]. MMN [454]. MOABC [959]. Mobile [2966, 1442, 3061, 478, 3861, 3027, 1635, 182, 1444, 2595, 3303, 1802, 3517, 1277, 3524, 647, 3280, 1676, 656, 1451, 575, 1602, 2082, 497, 792]. Mobile-Aware [1277]. MobileNet [2753, 2525]. MobileNetV2 [3847]. Mobility [1480, 3213, 1221, 2823, 3907, 3293, 1513]. Modal [3678, 347, 3885, 462, 2018, 3424, 461, 3776]. Modalflow [1221]. Mode [3196, 632, 3199, 386, 1262, 3277, 717, 1246, 1047, 2272, 462, 3891, 2312, 180, 993, 3120, 3786, 2976, 719, 2770, 729, 565]. Model [3491, 2579, 1991, 224, 2375, 1575, 981, 2670, 1143, 803, 2188, 3432, 2352, 2108, 3082, 2470, 2446, 2620, 3066, 3434, 3351, 3436, 2260, 2501, 1757, 2481, 1553, 3063, 544, 2910, 1867, 3199, 1842, 2692, 791, 2907, 3780, 3419, 3639, 3599, 3439, 3610, 3861, 191, 1670, 540, 2792, 2363, 1236, 1004, 2281, 2000, 3279, 3090, 2711, 1462, 2140, 2685, 1105, 1496, 1018, 2791, 2420, 3795, 2004, 926, 2503, 1697, 1076, 1741, 1825, 2167, 1877, 2632, 2315, 1244, 2175, 1869, 2500, 2407, 1783, 1267, 1632, 2636, 1780, 2285, 1251, 2053, 834, 1566, 2055, 2013, 2796, 3755, 737, 917, 1000, 1590]. Model [2371, 3332, 3747, 3556, 3821, 2369, 3420, 296, 809, 1786, 2867, 3327, 2057, 2749, 1188, 3731, 3701, 1323, 2687, 3104, 21, 3331, 3011, 884, 265, 1573, 3021, 2804, 1966, 1622, 2135, 3583, 637, 1270, 3396, 2335, 1772, 2463, 1359, 998, 1617, 2932, 1095, 2421, 1724, 1593, 960, 1120, 2565, 1755, 3526, 3069, 2985, 865, 3350, 3684, 1190, 3600, 1927, 1222, 1332, 2293, 3473, 1057, 2930, 1691, 986, 3352, 223, 2232, 3560, 2606, 1280, 953, 3908, 2238, 1659, 1225, 2414, 2223, 1981, 996, 3733, 774, 3667, 3786, 877, 628, 2570, 3532, 368, 3341, 3329, 3380, 653, 3723, 641, 2789]. Model [232, 284, 741, 1301, 1819, 2193, 2933, 2955, 3400, 3655, 3827, 2428, 2137, 2275, 1260, 819, 999, 477, 2782, 766, 3603, 782, 883, 705, 231, 2514, 786, 552, 2346, 2656, 3546, 3621, 1347, 1548, 3126, 3548, 3717, 768, 2725, 1342, 1862, 3572]. Model-Agnostic [1004, 1573]. Model-Based [2632, 2175, 2407, 3396, 2985, 1190, 2606, 953, 653]. Model-Driven [1575]. Model-Free [3199, 3546]. Model-NSGA-II [786]. Model-Robust [2293]. Modelica [3574]. Modelica-Based [3574]. Modeling [2184, 414, 1232, 661, 3796, 2155, 2086, 885, 862, 1858, 3204, 3413, 2922, 678, 2526, 2718, 1607, 209, 1611, 2251, 2322, 2283, 2700, 2678, 138, 1649, 967, 1891, 3218, 704, 543, 2778, 1115, 1736, 3888, 43, 3758, 288, 1298, 1254, 2733, 530, 546, 1808, 2177, 2166, 1401, 2064, 2566, 1749, 2122, 3764, 3906, 2278, 5, 1975, 1981, 3910, 731, 1785, 3536, 2130, 284, 3827, 3675, 1546, 3125, 2861, 1643, 453, 939]. Modelling [3397, 3817, 3858, 2249, 2234, 1792, 2378, 3830, 738, 1324, 60, 3048, 1864, 2677, 2232, 2339, 1098]. Models [2973, 2341, 2485, 3741, 3686, 3651, 2188, 3081, 490, 3759, 3130, 3681, 531, 1040, 1770, 2410, 471, 3372, 2993, 3409, 3751, 3333, 1481, 2474, 2888, 147, 3136, 2895, 2870, 26, 1929, 156, 617, 2669, 3088, 3338, 2001, 2683, 1880, 2365, 2295, 2629, 3078, 2851, 1271, 1263, 3292, 1917, 2924, 3209, 3264, 1801, 3554, 3870, 3067, 2090, 49, 3447, 3289, 1886, 2247, 3569, 3321, 2328, 3452, 2550, 3217, 3830, 1754, 1903, 2761, 1483, 509, 3158, 1063, 3622, 3102, 1800, 1963, 1907, 3339, 2248, 3229, 3110, 3555, 2601, 1520, 1166, 1036, 73, 2959, 2026, 3237, 3687, 399, 79, 1121, 2529]. Models [2293, 2810, 1705, 2753, 634, 3033, 2606, 2647, 1710, 3810, 2979, 2185, 2034, 91, 2263, 3376, 2944, 3911, 2848, 3326, 2824, 2649, 1620, 2385, 3620, 3682, 2777, 501, 710, 533, 2521, 3914, 3919, 3486, 2936, 3543, 3648, 2346, 3015, 3598, 3126, 3628, 3202]. Models-Based [3321]. Moderate [2938]. Modern [1759, 440, 2411, 295, 1881, 2668, 2979]. Modernising [2336]. Modes [3414]. Modification [3551, 933, 2484]. Modifications [3871, 1978]. Modified [3080, 1933, 3579, 2975, 242, 1090, 3421, 1779, 1619, 2751, 1267, 2094, 2635, 222, 401, 987, 736, 561, 297, 3458, 2807, 398, 609, 3784, 2902, 1475, 869, 759, 605, 3661, 1505, 559, 518, 664]. Modifying [345]. Modular [2039, 3852, 97, 2887]. Modularity [486, 396]. Modulated [3673, 608]. Modulation [2049, 3349, 485]. Module [3386, 527, 2657]. Modules [3507, 3518, 2954]. Modulo [2299]. MOEA [1565, 479]. MOEA/D [1565, 479]. Mold [2542, 1916]. Molding [2601]. Molecular [2089, 931, 1314, 3799, 532]. Molecules [3605, 2556]. MoM [2283]. MoM-Based [2283]. Moment [610]. Moments [2007, 2153, 2790, 1820]. Momentum [2783, 2054, 715]. Momentum-Based [2054]. Monarch [500, 602]. Monetary [881, 2479]. Money [881]. MongoDB [2161]. Monitor [3629, 2873, 3065]. Monitoring [107, 1798, 3493, 1609, 3346, 1678, 3336, 3279, 1803, 1730, 2224, 1087, 2438, 3318, 3451, 2051, 3883, 1791, 3843, 143, 20, 3897, 2023, 3466, 1023, 2092, 2266, 2383, 2034, 2073, 643, 2237, 1012, 3185, 3603, 3250, 3570]. Monkey [2892]. Mono [2954]. Monocular [3842]. Monopolar [1876]. Monotone [1138, 1614, 131]. Monotonicity [1600]. Montage [1773]. Montages [3854]. Monte [373, 3878, 964, 22, 1052, 2086, 3498, 2420, 2881, 568, 3150, 3659, 573, 2991]. Moore [1980]. MOPSO [3261, 3530]. MOr [1998]. More-for-Less [3664]. Moreau [1555]. Morphological [2841]. Morphology [3651, 1580, 2412, 3465]. Morse [1589]. Most [1822]. Moth [2767, 3551, 3106]. Moth-Flame [2767, 3551]. Motif [201, 194]. Motility [3477]. Motion [3271, 1818, 683, 976, 48, 38, 2313, 1889, 2637, 737, 2332, 3169, 1190, 3476, 3178, 2946, 1341, 899, 2872]. Motion-Based [2313]. Motor [2671, 690, 686, 2866, 520, 421, 1441, 3070]. Motorcycle [1824, 1870]. Motorcyclist [2859]. Motorway [1977]. Mould [1961]. Mount [1504, 1870]. Mountain [2566]. Mouse [2625]. Move [1211]. Movement [661, 2165, 3059, 669, 3451, 314, 187]. Movements [1260]. Moves [2616, 277, 349]. Moving [3200, 2903, 409, 2920, 2552, 2714, 1067, 750, 294]. MOWFA [1864]. MOX [1955]. MPC [2326, 1085, 1285, 3331, 2908, 3800]. MPC-Based [2908]. MPM [2256]. Mpox [2501]. MPPT [1284, 254, 3745]. MPVF [3424]. MR [2624, 3446, 3501, 2593, 1847, 2259]. MR-Linac [2624, 2593]. MRI [3415, 2199, 2787, 157, 2460, 77, 2667, 2559, 3909, 1460, 2872, 1497]. MRI-Based [3415]. mRNA [3319, 527]. MRP [1644]. MRSO [3080]. MS [1439]. MSAAD [3829]. MSD [2844]. MSEANet [3294]. MSFA [3143]. MSPSO [959]. MSV [3636]. MU [890]. MU-MIMO [890]. Much [3754]. Multi [3191, 1382, 1289, 2578, 3566, 3714, 3713, 654, 3012, 627, 1325, 981, 2988, 2947, 803, 974, 3195, 1388, 3797, 1551, 2620, 1910, 3261, 1940, 3262, 2622, 203, 2821, 544, 3804, 1384, 1943, 1417, 220, 2041, 2508, 600, 3781, 386, 791, 1514, 2586, 3041, 374, 506, 3359, 235, 372, 3024, 2803, 1509, 3203, 3693, 1811, 1161, 2457, 2424, 2145, 2433, 2462, 1182, 529, 2904, 3499, 3208, 3735, 1239, 190, 3264, 1801, 3144, 2827, 751, 3590, 2519, 3210, 3750, 3678, 817, 2330, 2698, 1457, 78, 172, 3753, 1022, 2257, 2709, 2390, 887, 3321, 2008, 1502, 1674, 1075, 2863, 306, 1855, 1780]. Multi [3335, 2052, 360, 3794, 834, 3008, 2013, 458, 1884, 39, 279, 800, 1172, 1516, 1775, 3660, 3885, 3325, 3256, 3886, 1723, 3785, 3157, 694, 1003, 1786, 2978, 3294, 3276, 3327, 3282, 3416, 3844, 2277, 3581, 1634, 2398, 1838, 1657, 3624, 3597, 977, 174, 193, 2889, 1189, 2956, 3265, 2885, 2017, 2018, 1967, 3583, 1354, 1439, 1897, 275, 2598, 3465, 404, 998, 3169, 3522, 2746, 1932, 1724, 2064, 843, 3904, 1465, 3113, 1292, 3350, 2389, 256, 603, 757, 3114, 1845, 1705, 665, 986, 1454, 2316, 2928, 869, 759, 2366, 3119, 1192, 3317, 3242, 3424, 943, 938, 198, 1508, 1758, 461, 3121, 605, 875, 3733, 2071, 3530, 3609]. Multi [1194, 1765, 2448, 1735, 2848, 1195, 2890, 653, 2747, 2982, 2571, 1471, 1315, 643, 805, 1660, 1826, 2748, 3308, 3319, 3293, 3788, 2275, 3776, 1650, 407, 651, 1053, 477, 2573, 3724, 249, 1056, 1356, 3918, 806, 1849, 1878, 3834, 2931, 724, 549, 814, 1131, 1789, 2702, 3046, 3550, 1230, 2245, 3285]. Multi-Access [3750]. Multi-Actor [1465]. Multi-Agent [803, 3144, 1457, 3886, 3522, 3242, 3121, 1826, 814]. Multi-Algorithm [938]. Multi-Area [256, 2890]. Multi-Arm [3191]. Multi-Armed [544, 2316, 3308]. Multi-Attack [1161]. Multi-Attribute [1801, 651]. Multi-Augmentation-Based [2748]. Multi-Band [39]. Multi-Bit [1932]. Multi-Branch [1239, 2709, 694, 998]. Multi-Capacity [1634]. Multi-Cell [2424]. Multi-Channel [2848]. Multi-Chaotic [3046]. Multi-Choice [887]. Multi-Class [2622, 1514, 751, 2257, 3321]. Multi-Classifier [1723]. Multi-Color [1780]. Multi-Component [3285]. Multi-Constraints [603]. Multi-Controller [2947]. Multi-Core [203, 458, 198, 1660]. Multi-Criteria [1388, 3208, 3264, 172, 1502, 2863, 2746, 3904, 1465, 1315, 407, 2931]. Multi-Cut [1417]. Multi-Dataset [2747]. Multi-Depot [605]. Multi-Destination [2982]. Multi-Dimensional [220, 3024, 3203, 2052, 1884, 3609, 1230]. Multi-Directional [1075]. Multi-Document [627]. Multi-DOF [404]. Multi-Facility [1354, 1439]. Multi-Factor [986]. Multi-Fault [3325]. Multi-Feature [3113, 3550]. Multi-Feedback [279]. Multi-Fidelity [1705, 1192, 1735, 1849]. Multi-Genome [2448]. Multi-Granularity [3581]. Multi-Graph [3041]. Multi-Guide [2145]. Multi-Head [3885, 2978]. Multi-Hop [306]. Multi-Improved [1056]. Multi-Innovation [477]. Multi-Input [372, 2457]. Multi-Instance [1765]. Multi-Interface [974]. Multi-Label [2620, 529, 1789]. Multi-Layer [3465]. Multi-Layered [3660]. Multi-Level [2508, 3678, 2598, 1724, 665, 1053]. Multi-Loop [981]. Multi-Marginal [1910]. Multi-Material [2330, 3114]. Multi-Metaheuristic [834]. Multi-Modal [2018, 3424, 461]. Multi-Mode [386]. Multi-Model [2013, 3583]. Multi-Object [1022, 2064]. Multi-Objective [1289, 2578, 654, 3012, 1325, 1551, 3261, 2041, 2508, 1514, 2586, 374, 2803, 2145, 2433, 2462, 190, 3590, 2519, 2698, 2390, 887, 1674, 2863, 3794, 3008, 800, 1172, 1786, 3416, 2277, 3624, 3265, 1897, 275, 843, 1845, 2366, 3119, 1508, 2571, 805, 3293, 1650, 3724, 249, 1356, 1878, 724, 1131, 2245, 1195]. Multi-Objective-Based [1855]. Multi-Occurrence [3335]. Multi-Omics [3797]. Multi-One-Class [1382]. Multi-Output [2448]. Multi-Parametric [875]. Multi-Period [977, 1292]. Multi-Person [3785]. Multi-Perspective [3317]. Multi-Phase [3530]. Multi-Physics [3262, 1657]. Multi-Physics-Based [2389]. Multi-Physiological [3885]. Multi-Population [1940, 869, 759]. Multi-Precision [198]. Multi-Process [2889]. Multi-Processor [2956]. Multi-Product [3499, 1786]. Multi-Project [3566]. Multi-Qubit [3753]. Multi-Resolution [3157]. Multi-RIS-Assisted [2702]. Multi-Roots [3834]. Multi-Row [724]. Multi-Scale [3714, 2821, 3781, 1003, 3294, 3350, 3733, 643, 3776]. Multi-Scenario [3293, 3918]. Multi-Sensor [235, 1182, 78]. Multi-Sequence [3597]. Multi-Server [3276]. Multi-Sided [174].

Multi-Site [2398, 3319]. Multi-SNP [2573]. Multi-Source [3041, 1509, 3282, 757, 1454]. Multi-Spectral [2071]. Multi-Stacking [3195]. Multi-Stage [600, 3693, 2827, 1194, 549]. Multi-Step [3735, 360]. Multi-Storage [2928]. Multi-Strategy [3359, 3210]. Multi-Surrogate [3256]. Multi-Swarm [1758]. Multi-Task [3804, 1516, 3844, 2017]. Multi-Tasking [3256]. Multi-Template [653]. Multi-Threading [506, 193]. Multi-Time [1967]. Multi-Track [817, 2885]. Multi-Trip [2008, 3788]. Multi-Type [1943, 2275]. Multi-UAV [791]. Multi-User [2702]. Multi-Valued [3713, 360, 1471]. Multi-Vector [1838]. Multi-Vehicle [2988, 3169]. Multi-View [1811, 2904, 1775, 806]. Multi-Wavelength [3327]. Multi-Way [1384]. Multi-Winner [1189, 943]. Multiagent [929, 2908, 1346]. Multiband [339]. Multicast [75, 319, 762]. Multicasting [365]. Multichannel [746, 2995]. Multicommodity [168]. Multicore [201, 196, 2951]. Multicriteria [1998, 1458]. Multidimensional [3191, 1176, 2211, 3428, 1607, 1319, 2739, 3625, 2642, 1120, 997]. Multiextremal [1978]. Multifield [42]. Multifractal [1858]. Multigrid [2133]. MultiKOC [1382]. Multilabel [1025]. Multilayer [3713, 503, 1437, 2407, 3047, 1971, 3071]. Multilevel [3658, 1052, 942]. Multilinear [2526, 2187]. Multilingual [1629]. Multimaterial [3500]. Multimedia [812, 3407, 1602]. Multimodal [3858, 3040, 2433, 3593, 800, 3763, 1723, 2856, 3344, 3824, 2844, 2269, 3247, 1750, 3765, 3249, 3250, 3717]. Multimodality [3110]. Multinomial [1375]. Multiobjective [3859, 2879, 446, 1187, 257, 3884, 3633, 3180, 1294]. Multiomics [2670]. Multiorgan [2624, 2593]. Multiparameter [2073]. Multiparametric [1847, 3585]. Multipartite [571]. Multipath [2802]. Multiple [2896, 3351, 1384, 1944, 1481, 1581, 3859, 3820, 1670, 695, 194, 1096, 1703, 379, 333, 172, 2239, 1100, 1506, 2441, 2773, 2425, 290, 3349, 2951, 1318, 2255, 2560, 1359, 1095, 1403, 2120, 2645, 2862, 2293, 3392, 560, 536, 3620, 533, 1543, 819, 1451, 1742, 666, 3543, 3546, 245]. Multiple-Attribute [536]. Multiple-Group [2293]. Multiple-Input [3349]. Multiple-Instance [1581]. Multiple-Output [3349]. Multiple-Vehicle [1095]. Multiplexing [3349]. Multiplication [2688, 69, 1446]. Multiplicative [684]. Multiplicity [328, 423, 333, 290, 455]. Multiplier [519]. Multipliers [2103, 351]. Multiply [1405]. Multipoint [968, 317]. Multiprocessor [1017, 3028, 2542, 1494]. MultiRes [1817]. MultiRes-RCF [1817]. Multiscale [3829, 928, 535, 1812, 3457, 3234, 1203]. Multisensor [3899]. Multisensorial [64]. Multisensorial/Multitemporal [64]. Multisensory [2844]. Multiset [2253]. Multiset-Trie [2253]. Multisets [1383]. Multisource [1733]. Multispectral [537, 2562, 2680]. Multistep [2063]. Multitask [1720]. Multitemporal [64]. Multithreaded [2382]. Multithreading [3038]. Multithreading-Based [3038]. Multivariable [1270, 705]. Multivariate [1525, 449, 1830, 418, 344, 1254, 3015]. Multiyear [3363]. muMAB [544]. Municipal [1315]. Murine [2251]. Muscle [241]. Muscles [3684]. Musculoskeletal [2844]. Music [1885, 1750]. Mutation [1276, 3551, 885, 1378, 448, 1376, 2829]. Mutations [1518]. Mutual [50, 2693, 3233, 3464, 847, 384]. MVACLNet [2856]. myDATA [3591]. Myocardial [1497]. Myocardium [2559].

N [3211]. N4 [1543]. N4-Methylcytosine [1543]. Nadaraya [2309]. NaI [28]. Nakagami [2360]. Nakagami- [2360]. Named [3763]. Names [2969]. Nanning [3419, 3752]. Nanobeam [3316]. Nanobeams [1755]. Nanoparticles [31]. Nanosensors [2164]. Nanotubes [17]. Naples [3438]. Narrative [3566]. Narrow [2609]. Narrow-Down [2609]. NARX [1143]. Nash [1576, 1275]. Nashville [2577]. Natalie [313]. Natural [2304, 1834, 2453, 2052, 2311, 1871, 2807, 1808, 1465, 674, 2789]. Natural-Cultural [1465]. Naturalize [2615]. Nature [1231, 556, 511, 2496, 1008, 1647, 2838]. Nature-Based [511]. Nature-Inspired [1231, 556, 2496, 1008, 1647, 2838]. Navier [1877, 1141]. Navigating [3588, 2884, 2686]. Navigation [2720, 127, 2495, 3495, 123, 9, 2796, 937, 1235]. Navy [1388]. Nayar [606]. NB15 [2721]. NBME [3848, 3116]. NBME-Standard [3848, 3116]. NBTI [487, 332]. NBTI-Aware [332]. NDARTS [2631]. Near [574, 314, 1844]. Near-Infrared [314]. Near-Optimal [574]. Nearest [2226, 924, 2111, 258, 550, 1597]. Nearly [608]. Need [2785, 1029, 2210, 2502, 1460]. Needed [63]. Needs [3293]. Negamax [151]. Negations [1096, 833]. Negative [1740, 1857, 1222, 1361]. Negatively [167]. Negatives [2552]. Negotiation [3037]. Negotiation-Based [3037]. NEH [1054]. Neighbor [2111, 550, 3183]. Neighborhood [1936, 783, 1187, 331, 1796, 569, 173, 3290]. Neighborhoods [164]. Neighbors [258]. Neighbour [2125]. Neighbour-Based [2125]. Neighbours [2226, 1597]. Nelder [3128, 1470]. Nemesis [2180]. Nephrolithotomy [3810]. Nero [2121]. Nerve [3854]. Nervousness [2469]. Nested [924, 940, 1003]. Nesterov [1610]. Net [3780, 1880, 3152, 2636, 1003, 2837, 3317, 814, 1987, 2482, 3009, 3840, 3323, 3317, 1407, 2914]. Net-Zero [3152]. Nets [695, 3739, 1229]. Network [1406, 66, 3025, 2973, 1498, 3713, 3849, 3194, 3839, 2302, 2581, 3665, 871, 1797, 3427, 373, 2149, 2470, 1010, 905, 840, 2427, 2466, 1860, 1268, 544, 13, 449, 3035, 866, 1069, 1147, 1068, 2825, 2919, 3041, 3749, 2879, 540, 862, 1097, 1837, 503, 2444, 313, 473, 1558, 2462, 2628, 1952, 1428, 3369, 2340, 1239, 1873, 2251, 2547, 2207, 760, 1241, 2815, 3263, 2519, 1621, 234, 1491, 184, 707, 1687, 1244, 2254, 2590, 2809, 3146, 2194, 2493, 1267, 2504, 3876, 2257, 1444, 1615, 508, 1186, 2143, 2285, 1302, 2761, 458, 3694, 3390, 1007, 706, 917, 1115, 2371, 2209, 3706, 3512, 3885, 3821, 43, 650, 535]. Network [2997, 694, 1312, 2978, 2856, 2867, 3294, 3327, 3836, 3844, 1570, 3355, 2057, 2277, 2749, 2398, 1531, 1657, 1365, 2687, 3819, 3322, 2813, 1216, 1857, 585, 2886, 3323, 2986, 2314, 2706, 2597, 2598, 2335, 12, 2463, 2643, 1440, 2563, 2062, 2679, 3468, 1401, 233, 1927, 1821, 1971, 1416, 1256, 292, 3240, 2772, 2258, 2030, 2464, 3175, 1733, 1472, 587, 3404, 3789, 2859, 3733, 2914, 3529, 1094, 2686, 3341, 3608, 580, 2452, 2074, 2338, 1107, 2651, 2317, 2654, 2981, 3353, 3788, 3571, 3182, 3776, 1111, 1443, 3071, 3841, 3248, 408, 3540, 357, 807, 1355, 1545, 3377, 3917, 527, 559, 2079, 2114, 3541, 3725, 806, 3542, 859]. Network [872, 1712, 697, 552, 826, 963, 2081, 1763, 2576, 2577, 2702, 2847, 2860, 939, 1603, 2770, 3490, 841, 1347, 1987, 3300, 1342, 2531]. Network-Based [2986, 2463, 3341]. Network-Enhanced [2576]. Network-Generated [2859]. Network-VGG16 [1355]. Network-Wide [1267]. Networked [247]. Networking [3213, 3534]. Networks [1154, 2579, 1641, 3018, 1680, 3054, 2947, 1850, 974, 3261, 1418, 2360, 3262, 2221, 414, 3072, 2292, 1275, 1578, 2506, 1327, 2511, 2993, 2623, 3275, 2540, 3356, 3085, 25, 241, 236, 266, 633, 2820, 2858, 3861, 1386, 1948, 970, 945, 3864, 1499, 975, 2787, 663, 503, 75, 1840, 3413, 2898, 2424, 2544, 812, 1895, 2883, 301, 1399, 3561, 2664, 2003, 2705, 2819, 2900, 919, 2049, 2162, 1087, 3636, 1392, 1328, 3055, 978, 2419, 59, 238, 379, 1410, 1635, 3447, 1437, 1610, 1560, 2407, 795, 3874, 3875, 3373, 19, 2449, 2709, 210, 2527, 2094, 3806, 3215, 306, 3782]. Networks [437, 1748, 3219, 1767, 2593, 3100, 1861, 1302, 2637, 2011, 1290, 2729, 3303, 1213, 1330, 1411, 1847, 3884, 3514, 3226, 793, 2596, 829, 3889, 1140, 62, 226, 3573, 2259, 1174, 979, 3823, 1876, 1924, 2733, 1814, 1015, 18, 161, 3520, 2208, 2105, 1218, 2344, 2754, 1125, 463, 647, 3526, 1283, 3472, 56, 1927, 1971, 3634, 2030, 2308, 2316, 869, 759, 2405, 2507, 1827, 2681, 1974, 3004, 1223, 3716, 3807, 2569, 40, 2288, 1338, 2071, 2354, 2607, 1906, 2243, 2154, 3122, 3481, 3422, 3586, 2182, 2074, 2402, 70, 898, 3345, 3050, 674, 3775, 286, 847, 1522, 1511, 1819]. Networks [2392, 3319, 3368, 907, 545, 1197, 2572, 3724, 656, 1447, 396, 348, 2575, 364, 1638, 1131, 1026, 1602, 2877, 497, 679, 792, 2381, 228, 1066, 1235, 1497]. Neumann [2631, 2194, 3171]. Neural [1406, 3025, 2973, 2579, 3018, 3713, 3849, 3839, 2302, 2581, 3665, 2470, 1010, 2360, 3262, 2427, 2466, 3072, 1578, 13, 2506, 449, 2511, 3275, 3035, 2540, 3356, 241, 236, 266, 633, 1147, 2820, 2107, 1948, 540, 1702, 1499, 2787, 2444, 1840, 3413, 2898, 2628, 2046, 1952, 3561, 1428, 2003, 2705, 2819, 2340, 2900, 1873, 2049, 983, 2547, 2207, 2476, 2162, 1087, 760, 1241, 3263, 3636, 1392, 1621, 3055, 1687, 59, 379, 1610, 1560, 2194, 2493, 2407, 795, 2504, 3874, 3875, 2257, 1444, 2527, 2697, 1754, 1748, 2143, 2285, 2593, 3100, 1302, 2761, 2011, 1290, 3303, 3694, 917, 1115]. Neural [1411, 1847, 2371, 2209, 43, 535, 2596, 2867, 829, 1570, 3355, 716, 1140, 2277, 2398, 1657, 2687, 3819, 2813, 2558, 2886, 2733, 2986, 2314, 2706, 3520, 2208, 12, 2463, 2643, 1440, 1218, 2062, 2679, 2754, 3468, 463, 3526, 3472, 56, 1927, 3175, 1733, 2405, 3004, 3716, 3807, 2569, 2071, 2354, 2607, 580, 1542, 2243, 3122, 3422, 3586, 2182, 2074, 2402, 2338, 1107, 898, 3775, 847, 1511, 3308, 3319, 3368, 3913, 3571, 907, 2572, 357, 756, 807, 1355, 3917, 559, 2575, 2180, 364, 859, 872, 1638, 1712, 552, 826, 2576, 2877, 679, 939, 1603, 2381, 2770, 841, 1987, 1066, 1342, 2531, 1235, 1497]. Neural-Network-Assisted [2493]. Neural-Network-Based [2470, 2427, 2466, 2194]. Neuro [2234, 564]. Neuro-Dynamic [564]. Neurodegenerative [3033]. Neurodynamic [882].

Neuroevolution [1376, 1726]. Neuroevolutionary [1862]. Neurohybrid [2909]. Neuroimaging [2199]. Neurological [2341]. Neuronal [24, 3100]. Neurons [3713, 3582, 2016, 1772, 1520, 1107]. Neurophysiology [3777]. Neutral [3026, 1974]. Neutrality [881]. Neutrosophic [590, 563, 560, 533, 524]. News [2580, 142, 3088, 3678, 704, 3672, 3298, 2059, 2727, 2789, 2891]. Newton [1020, 291, 302, 2811, 305, 283, 280, 323, 2454, 1610, 3831, 1132, 299, 445, 2575]. Newton-Like [305, 2811]. Newton-Type [291, 302, 445]. Next [1093, 3730, 1625]. Next-Generation [1093, 1625]. NFT [3148]. NFT-Based [3148]. NICA [2395]. Niching [3105]. NIDS [2812]. Nine [1504]. Ninth [1089]. NIPALS [2560]. NIRS [604]. NK [2587]. NLP [3129, 440, 2941, 3660]. NMF [1896]. NN [727, 1124, 238]. NNetEn [2338]. No [512, 410, 460, 1655, 2191, 1393]. No-Wait [512, 1393]. Node [1954, 1560, 1632, 1330, 2209, 143, 1975, 2317, 1131, 1602]. Nodes [1145, 3496, 473, 2094, 3156, 1330, 650, 3889, 1971, 1223, 3608, 3841, 756, 691]. Nodule [2818, 260]. Nodules [91]. Noise [3191, 1662, 1937, 1892, 1366, 2457, 2851, 3052, 751, 379, 3505, 385, 3349, 265, 1030, 6, 180, 1453, 3244, 2914, 10, 2300, 710, 3308, 2379, 300, 377]. Noise-Adding [1892]. Noises [2311, 294]. Noising [705]. Noisy [1306, 3829, 2322, 154]. Nomex [2861]. Nomological [3558]. Non [3192, 3347, 1040, 3582, 1275, 3857, 3612, 3439, 3497, 3777, 861, 1686, 2546, 1953, 2632, 528, 951, 3702, 3707, 1295, 3831, 822, 913, 507, 965, 1740, 1905, 1568, 582, 167, 2016, 3519, 1255, 3233, 1403, 1932, 2768, 3007, 1332, 2605, 2677, 1756, 1361, 3370, 857, 1099, 877, 3814, 731, 1620, 2905, 2653, 838, 477, 1451, 513, 377, 3570, 3202, 1474]. Non-Atomic [3233]. Non-Attendance [1332]. Non-Coding [3857]. Non-Conjugate [1620]. Non-Convex [3439, 2653]. Non-Cooperative [1275]. Non-Destructive [3497]. Non-Deterministic [1474]. Non-Dominated [1568]. Non-English [2546]. Non-Food [3777]. Non-Gaussian [377]. Non-Gradient [2905]. Non-Ideal [582]. Non-Identical [3582]. Non-IID [3707, 3370]. Non-Interactive [1932]. Non-Invasive [3347, 1686]. Non-Irrigated [3202]. Non-Isothermal [861]. Non-Iterative [2905]. Non-Linear [1040, 3612, 528, 3519, 1255, 1403, 2605, 3814, 838]. Non-Local [507]. Non-Markovian [3702, 877]. Non-Negative [1740, 1361]. Non-Negatively [167]. Non-Normal [1099]. Non-Orthogonal [1905]. Non-Overlapping [1295]. Non-Parametric [2632, 2768, 3570]. Non-Penetrating [3831]. Non-Periodic [2016]. Non-Preemptible [513]. Non-Singular [2677]. Non-Smooth [951, 3007, 857]. Non-Standard [913]. Non-Stationary [1953]. Non-Traditional [1451]. Non-Transitivity [1756]. Non-Uniform [3192, 731]. Non-Uniformly [477]. Non-Uniquely [822]. Noncircular [384]. Nonconvex [882, 2764]. Nondestructive [2824]. Nondifferentiable [3721]. Nonexpansive [360, 398, 1471]. Noninvasive [3327]. Nonlinear [1020, 2039, 328, 455, 423, 454, 2811, 305, 1429, 2506, 3820, 322, 1349, 2246, 1204, 3382, 482, 3207, 2526, 1496, 409, 274, 2355, 717, 3712, 1046, 543, 1212, 3694, 2961, 290, 1188, 481, 1367, 2131, 3837, 1009, 2963, 3170, 900, 180, 341, 865, 993, 1616, 325, 3241, 3641, 1132, 3243, 3244, 2147, 2523, 299, 337, 445, 2822, 2743, 3834, 317, 294]. Nonlinearity [1302, 1859, 270]. Nonlocal [3617, 1594, 2119, 249]. Nonnegative [2270]. Nonparametric [3340, 3829, 1271, 2393, 1106, 465, 1203]. Nonsmooth [322, 3378, 2477, 3006, 1210, 2527, 406, 1132, 2110, 3180]. Nonstationary [3193, 180]. Nonuniform [355]. Norm [340, 1555, 3500, 751, 1592, 625, 1926, 1918]. Norm-Based [3500]. Normal [1525, 732, 2112, 2460, 3740, 1099]. Normalization [2938, 2969, 1099]. Normalized [1715, 2220]. Normothermic [1215]. North [2979]. North-Eastern [2979]. Norwegian [1792]. NoSQL [2161, 850]. Note [126, 2936]. Notes [1258]. Novel [2614, 2949, 2617, 440, 1940, 14, 1062, 3596, 1179, 2211, 2298, 230, 3599, 2970, 537, 3382, 874, 2093, 3709, 926, 2496, 3263, 1621, 3503, 36, 1047, 1506, 2942, 3766, 3224, 557, 1847, 2181, 621, 547, 3327, 327, 1464, 626, 1967, 2135, 2490, 3166, 730, 1651, 1095, 1059, 3117, 1300, 2287, 3404, 378, 918, 2914, 2075, 1446, 3345, 643, 805, 2933, 904, 2275, 578, 244, 608, 3725, 1198, 2743, 762, 1793, 2656, 2872, 3737, 1066, 1230]. Novelty [2144, 1943, 3165, 2070]. Novice [2136]. NP [512, 2633]. NP-Hard [512, 2633]. NRL [1624]. NSBR [2914]. NSBR-Net [2914]. NSCT [507]. NSGA [3261, 521, 2277, 959, 1650, 786]. NSGA-II [3261, 959, 1650]. NSGA-PINN [2277]. Nuclear [1580, 1127, 625]. Null [1661]. Number [2757, 1549, 3411, 3139, 473, 3867, 2814, 2496, 3067, 3374, 2632, 1415, 2390, 2008, 2153, 259, 3889, 2668, 1218, 1315, 741, 1141, 1131]. Numberlink [133]. Numbers [2306, 1946, 2852, 2727]. Numerals [903]. Numeration [2712]. Numerical [1040, 3579, 1553, 661, 261, 1072, 2984, 980, 1714, 861, 954, 1557, 2420, 1779, 323, 3873, 2194, 2493, 3150, 3044, 3830, 3223, 1736, 2133, 2109, 3103, 2020, 2139, 2063, 3007, 900, 1537, 1432, 3600, 842, 3475, 2267, 1361, 1445, 3405, 3814, 2701, 2707, 1495, 2743, 3125]. Numerical-Semantic [3873]. Numerically [1085, 1285]. Numerous [3654]. NURBS [1024]. Nurse [173]. NuSTAR [3437]. Nutrition [2113].

OADMs [1399]. Obfuscation [3306]. Object [2540, 3610, 1670, 540, 38, 707, 2254, 1022, 3877, 3747, 2552, 1109, 1253, 2846, 3330, 2310, 2064, 3170, 681, 3424, 3789, 3273, 91, 2987, 653, 2760, 2933, 1443, 2766, 733, 2613, 2657, 958, 3190]. Object-Oriented [1670, 540]. Objective [1289, 1017, 3028, 2578, 654, 3012, 1325, 1551, 3261, 3579, 2041, 2508, 1514, 2586, 374, 2792, 2803, 2145, 3074, 2433, 2462, 190, 3590, 2519, 2698, 2390, 887, 1674, 2863, 2800, 1855, 3794, 3008, 800, 1172, 1786, 3416, 2277, 3624, 3265, 1897, 275, 843, 1845, 2928, 2366, 3392, 3119, 1508, 597, 2944, 2571, 805, 3293, 1650, 3724, 249, 1356, 599, 2037, 1878, 724, 1131, 479, 2245, 1195]. Objectives [611]. Objects [3279, 1635, 1120, 2376, 1146, 2905, 3385]. Oblique [2365]. Observability [2439, 1009, 3522, 710]. Observable [3373, 3112, 2428, 2137]. Observation [1730, 3278, 2505, 3426, 2197, 3670]. Observations [2896, 1889, 2807]. Observe [3191, 1284]. Observed [1138].

Observer [2927, 431, 3501, 1367, 1453, 1336, 719]. Observer-Based [719]. Observer/Kalman [431]. Observers [667]. Obstacle [3552, 2303, 3652]. Obstacles [3275, 3561, 3077]. Obstetric [3844]. Obtained [2973, 323]. Obtaining [3654]. OC [2384]. OC-SMART [2384]. Occlusion [2258, 836]. Occultation [13]. Occupancy [2223]. Occupants [964]. Occupational [3506]. Occurrence [3335, 158]. Occurrences [2837]. Ocean [2384, 3236, 3683]. Oceanographic [64]. OCSI [1320]. Octagonal [23]. Octapeptide [22]. Octonions [1554]. Ocular [3062]. ODE [1874, 1909]. ODEs [341]. OFDM [279, 965]. Off [1946, 1825, 1767, 1923, 2325, 744, 3813, 553]. Off-Road [1767, 553]. Off-Target [3813]. Offensive [3648]. Office [1354, 964]. Official [2827, 2318, 3920]. Offline [976]. Offloading [792]. Offset [325]. Offset-Assisted [325]. Offshore [3081, 3441]. OGGWO [3265]. OIKAN [3919]. Oil [2437, 3011, 1822, 1375, 3801, 3609, 2428, 2137, 1012]. Oil-Impregnated [1375]. Oils [89]. OLAP [1513]. Old [3128, 2020, 3648]. Older [3040, 356, 1982]. Olive [2426]. Omega [1258]. Omics [3797]. Omnichannel [1171]. On-Board [2197]. On-Device [3013]. On-Edge [2937]. On-Line [431]. On-Tiny-Device [2679]. Once [2925, 3444, 3176]. Oncological [1891]. One [1382, 287, 3143, 1566, 2706, 2186, 3257, 2031, 2333, 3586, 2076, 501, 870, 3791]. One-Bit [287]. One-Day-Ahead [501]. One-Dimensional [2031, 3586, 870, 3791]. One-Machine [2076]. One-Shot [3143]. One-Sided [3257]. Ones [1558]. Online [1498, 2894, 1321, 1184, 781, 2148, 1022, 690, 920, 134, 58, 2498, 2902, 1125, 1593, 2645, 564, 1001, 1821, 1691, 3240, 1162, 3176, 2169, 1826, 2036, 545, 2307, 718, 682, 3252]. Only [315, 388, 3176]. onto [695, 1921, 2772]. Ontologies [1915]. Ontology [1002, 926, 1352]. Oort [2709]. Open [1017, 3028, 169, 3084, 1943, 1833, 1846, 3128, 55, 2778, 3337, 1761, 725, 1466, 2151, 1038, 2264, 155]. Open-Loop [1761]. Open-Shop [3028]. Open-Source [169, 1833, 1466, 155]. OpenACC [721]. OpenCL [893]. OpenMp [2157]. Operating [2336]. Operation [2467, 2693, 3468, 2028, 2339, 1376, 3005, 2874, 1512, 3835, 3305]. Operational [2865, 1293, 693]. Operations [1388, 2990, 3037, 2585, 3279, 1834, 3638, 2491, 1423, 1777, 1821, 1026]. Operator [3025, 2107, 1628, 2207, 778, 747, 1960, 1736, 1924, 3123, 3183]. Operators [291, 1096, 3619, 3145, 2841, 1336, 560, 3480, 536, 533, 651]. OPGs [3002]. Opinion [1649, 3560]. Opinions [3734, 82]. Opportunistic [650, 656]. Opportunities [2587, 2512, 3808, 3604, 2440]. Opportunity [2486, 2569]. Opposite [3157]. Opposition [366, 3811, 578, 244, 1916, 415]. Opposition-Based [366, 3811, 578, 244, 1916, 415]. OPT [1211]. OPTCON [1429]. Optic [3022]. Optical [1275, 2086, 1399, 648, 619, 795, 1003, 2687, 1532, 2731, 2528, 2166, 2343, 660, 3244, 836, 3353]. Optics [3804]. Optimal [66, 1641, 3847, 3427, 289, 1910, 2896, 3434, 3262, 955, 2221, 1306, 1429, 2172, 3363, 1947, 970, 617, 3423, 1123, 1297, 473, 2627, 2510, 2850, 574, 2664, 2112, 2503, 3871, 892, 1437, 177, 333, 3318, 2431, 2257, 267, 1615, 3023, 3153, 2214, 1909, 1775, 2961, 2868, 3890, 2164, 3518, 3104, 3106, 728, 1964, 1318, 1769, 1897, 2598, 1900, 1761, 3358, 612, 3257, 3414, 3600, 1538, 3402, 1412, 1448, 986, 3476, 2308, 3392, 1445, 1976, 3626, 1441, 2945, 966, 588, 1391, 2954, 3791, 545, 3918, 3835, 1870, 2417, 3488, 1725, 2702, 3270, 317, 479, 729, 3060]. Optimality [200, 1416, 2029, 634, 3180]. Optimality-Based [2029]. Optimally [909, 1541]. Optimisation [3768, 3693, 1840, 1291, 3105, 1467, 2402, 3916, 396]. Optimised [1426]. Optimistic [2357]. Optimization [1289, 3080, 1231, 2578, 1990, 562, 1523, 3412, 1325, 3650, 2616, 1164, 803, 3809, 3081, 3850, 1517, 3130, 1797, 3605, 3681, 726, 2455, 2391, 1442, 3718, 3083, 1818, 3261, 363, 1995, 169, 3131, 1940, 3287, 3084, 424, 1860, 512, 2993, 2129, 510, 799, 2041, 2508, 478, 1888, 3363, 753, 3051, 2767, 171, 242, 500, 1514, 2965, 3551, 3495, 1702, 1081, 554, 1582, 1027, 2970, 1389, 2803, 457, 519, 3204, 31, 1462, 2394, 521, 3371, 2683, 511, 2140, 2145, 3205, 3206, 3074, 812, 2433, 2200, 3378, 1806, 3369, 2999, 3500, 1238, 2396, 3795, 3627, 335, 2482, 3209, 926]. Optimization [1801, 2048, 1953, 3006, 1108, 3593, 2496, 711, 472, 3590, 778, 2454, 3636, 2751, 1868, 3410, 3446, 3027, 1527, 1703, 1692, 602, 747, 2175, 2128, 1437, 3873, 3147, 2500, 213, 211, 1480, 2832, 882, 446, 3875, 1623, 951, 2390, 2852, 1210, 2527, 1293, 1157, 2094, 3805, 1674, 2863, 269, 2690, 3707, 3453, 437, 2764, 3594, 3219, 543, 2783, 3221, 3794, 834, 1434, 581, 257, 285, 401, 800, 987, 1172, 1411, 1774, 2015, 1961, 2920, 3660, 3455, 3884, 3760, 3256, 3887, 3514, 2997, 397, 675, 547, 828, 2596, 2867, 3344, 3282, 3625, 3416, 3890, 2277, 3456, 327, 1802, 3891, 548, 3517, 2534, 2556, 3104, 2640, 3106, 1216]. Optimization [884, 2291, 2198, 2913, 1298, 3288, 3265, 3384, 2345, 3823, 1760, 3339, 1613, 3461, 1689, 1279, 2459, 946, 2737, 1354, 1439, 1533, 275, 2103, 3553, 2463, 382, 3167, 494, 1618, 1095, 2682, 3633, 3721, 1900, 1761, 1008, 569, 2755, 1219, 2344, 1535, 2274, 475, 3007, 2064, 1593, 3170, 3469, 1059, 851, 3769, 237, 1292, 2939, 3073, 603, 3114, 1222, 3659, 2952, 1616, 2029, 1482, 3475, 2772, 1475, 3477, 2928, 869, 759, 2366, 3119, 1192, 3120, 480, 2763, 1441, 538, 3789, 3176, 3036, 3825, 827, 1487, 2100, 1978, 1038, 3745, 2744, 1758, 605, 2710, 346, 1303, 2842, 3530]. Optimization [1478, 3245, 3667, 516, 1033, 2374, 1735, 1501, 2835, 1918, 2154, 3122, 2747, 2571, 3536, 3124, 1394, 2075, 1370, 2838, 821, 262, 326, 425, 878, 824, 805, 1826, 2193, 3584, 3811, 3484, 3788, 2137, 486, 1650, 838, 1544, 1197, 3248, 578, 3540, 3724, 992, 502, 1356, 1809, 2114, 782, 599, 493, 2037, 2612, 691, 1198, 1878, 1870, 2379, 3545, 3188, 3771, 658, 786, 697, 762, 1131, 1916, 1793, 1643, 2991, 415, 792, 2829, 3921, 3922, 228, 479, 664, 1548, 3548, 553, 2797, 3399]. Optimization-Application [2048]. Optimization-Based [2527, 1157, 437, 1172, 2867, 548, 2682, 2374]. Optimization-Jaya [3036]. Optimizations [3579]. Optimized [2375, 1550, 2618, 2620, 777, 3063, 3781, 3136, 3279, 723, 521, 472, 3097, 3362, 795, 2094, 3223, 271, 2386, 3828, 375, 3891, 2019, 3460, 3526, 891, 1972, 1597, 1765, 2484, 425, 1762]. Optimizer [2817, 3003, 3084, 2112, 1623, 1434, 2348, 468, 2571, 559]. Optimizing [1010, 1463, 3854, 2938, 2921, 3698, 2326, 3712, 1527, 2739, 3154, 1736, 2694, 1815, 722, 2643, 3905, 3659, 3326, 2106, 2747, 3361]. OPTIMUS [885]. Option [2711, 3413, 3067, 3057]. Options [1263]. Optoelectronic [766]. Oracle [3039, 2831]. Oracles [1033]. Orbit [2574, 1436]. Orbital [3896]. Orchard [3815]. Orchestration [1186]. Order [3254, 2780, 1143, 291, 464, 2618, 315, 289, 388, 1455, 2085, 1442, 283, 2501, 240, 2418, 1232, 632, 1201, 2586, 1112, 1947, 1999, 2711, 667, 503, 754, 717, 2751, 318, 1457, 333, 3876, 2634, 267, 2239, 3221, 1640, 616, 1212, 290, 621, 1165, 2715, 3163, 1058, 2156, 2019, 2730, 626, 1921, 2314, 637, 1120, 2026, 865, 1089, 603, 634, 3641, 1099, 1033, 445, 606, 1408, 3616, 493, 1926, 2743, 228]. Order-Based [2314]. Ordered [2767, 3711, 2956]. Ordering [2685, 128]. Orderings [198]. Orders [3038, 940]. Ordinal [1388, 1306, 961]. Ordinary [754, 1919]. Ore [3005]. Oren [606]. Organic [568]. Organising [1322]. Organism [3012, 1448]. Organization [30, 24]. Organizations [3425, 1057]. Organize [3703]. Organizing [1183]. Orientation [733]. Oriented [531, 2896, 1670, 540, 831, 2330, 1631, 1055, 2022, 2561, 2492, 3174, 1224, 1856, 3916]. Orienteering [627, 3561, 329, 2857]. Origami [1415]. Origin [2668, 1221, 231]. Origin-Based [231]. OrthoDETR [2655]. Orthoframes [1140]. Orthogonal [848, 1090, 2403, 3149, 345, 2380, 338, 1905, 397, 3349, 3330, 3486]. Orthogonalized [2923]. Orthogonalized/Affine [2923]. Orthopedic [2655]. Orthorectify [390]. Oscillating [1504]. Oscillation [102]. Oscillations [3045, 3165]. Oscillators [2778]. Oscillatory [2820, 3896]. Osteoarthritis [3595, 2665]. Other [2723]. Out-of-Domain [2512]. Outbreak [1175]. Outcome [2159, 1791]. Outcome-Based [1791]. Outcomes [1395, 3879, 1016]. Outdoor [1560, 3462, 937, 3635]. Outer [3411]. Outer-Independent [3411]. Outerplanar [166]. Outlier [2916, 1314, 2014, 3406, 480, 2979]. Outliers [1656, 2649]. Outlook [2388, 2378]. Outpatient [1997, 1689]. Outperforming [2484]. Output [1236, 372, 418, 839, 3349, 539, 865, 2232, 3807, 2448, 2652, 1512, 1742]. Output-Error [418]. Outputs [3558]. Outranking [3259]. Outsmarting [1746]. Ovarian [2413]. Over-Constrained [275]. Over-Determined [2868]. Over-Relaxation [3097]. Overall [2661, 1644]. Overallocation [2887]. Overcome [904]. Overcoming [2258]. Overcurrent [3724]. Overdetermined [2325]. Overexpression [2806]. Overhead [2272, 1195]. Overlap [303, 331, 1320]. Overlap-Based [331]. Overlapping [1295, 1485, 2019, 1926]. Overlaps [1489]. Overlay [706]. Overlays [202]. Overrelaxed [1391]. Oversampling [3049, 2771]. Overtaking [1970]. Overview [1021, 1580, 170, 2879, 2461, 2830, 2596, 2284, 1814, 2274, 3056, 3814, 286, 391]. OVSF [58]. Owl [2618]. Ownership [2826]. Oxide [1018, 1001]. Oxy [601]. Oxy- [601]. Oxygen [1550].

P [1702]. P6 [2887]. Package [1895, 1630, 2338]. Packages [3524].

Packaging [3622]. Packet [1836, 1884, 422, 2079]. Packet-Level [2079]. Packets [650, 419]. Packing [2855, 3367, 3619, 3754, 1593, 3170, 2212]. Pad [2921]. Padé [3340]. Page [2342, 380]. pages [74]. Pain [2883]. Painting [535]. Pair [1265, 376, 2337, 2876, 2317]. Paired [3067]. Pairs [3812, 3778]. Pairwise [775, 1470]. Palm [230, 1822]. Palpation [3798]. PAMA [3507]. Pamplona [1280]. Pancreatic [2892]. Pandemic [2700, 1492]. Panel [3890, 82]. Panels [3745]. Panoptic [3323]. Pansharpening [760]. Pantograph [3251]. Paper [1716]. Papers [2002, 3580, 1601, 2722]. Parabola [2764]. Parabolic [1191, 3791, 231]. Paraclinical [2828]. Paraclique [1314]. Paradigm [709]. Paradigms [1881, 1016]. Paradox [3703, 3664]. Parallel [3271, 1517, 930, 1552, 298, 3035, 1554, 3051, 1639, 848, 2774, 636, 116, 427, 3393, 2367, 3266, 3736, 2048, 3443, 3575, 1185, 196, 721, 450, 1836, 2325, 3103, 3011, 1787, 2123, 770, 3463, 1626, 1721, 2024, 1617, 3048, 1539, 3241, 538, 281, 938, 198, 1980, 2070, 3005, 1894, 1405, 273, 643, 2163, 488, 1546, 1712, 1510, 3399]. Parallel-in-Time [3736]. Parallelism [1139, 934, 184, 2193]. Parallelizable [1768]. Parallelization [3853, 1589, 3160]. Parallelized [1084, 3784]. Parallelizing [3281, 237]. Parameter [1139, 1261, 424, 3612, 761, 2347, 372, 1123, 631, 2683, 287, 1619, 2745, 3505, 2128, 1783, 1623, 696, 2690, 2143, 616, 285, 401, 1173, 1411, 3617, 509, 3890, 716, 994, 801, 1772, 2362, 2186, 392, 1453, 256, 845, 1914, 2383, 1226, 1726, 644, 798, 1706, 326, 710, 3791, 216, 1870, 2379, 815, 354, 718, 1421, 3548]. Parameterisation [2324]. Parameterised [933]. Parameterization [252, 3227]. Parameterized [1088, 251, 2052, 794, 1170, 946, 932]. Parameters [3491, 22, 478, 1297, 2510, 1611, 2454, 67, 1661, 1246, 887, 3154, 2783, 2272, 561, 2866, 1897, 2643, 2562, 1535, 2768, 1333, 1608, 2154, 3481, 3536, 3775, 579]. Parametric [2246, 2632, 338, 737, 2563, 2768, 1844, 875, 3570]. Parametrized [1999, 1106]. ParaNet [2518]. Parasites [1277]. Parathyroid [2040, 2518]. Parcel [2661]. Parents [3235]. Pareto [2375, 335, 211, 2863, 3624, 3104, 2029, 213]. Pareto-Optimal [3104]. Pareto-Optimized [2375]. Parity [2985]. Parking [3427, 2965, 3551, 2737, 2115]. Parkinson [2582, 1609, 2043, 3451, 3779]. Parkinsonisms [1300]. Parks [2425, 1000]. Parliament [1666]. Parsing [2925, 120, 381, 1587, 1199]. Part [3093, 3450, 2649]. Partial [3254, 955, 1942, 3428, 3138, 409, 941, 2493, 3149, 2429, 1812, 3837, 3676, 3654, 2232, 1162, 3243]. Partially [1377, 412, 2956, 2459, 3112, 2428, 2137]. Participant [1340]. Participants [1754]. Participation [1651]. Particle [562, 3130, 2455, 2466, 512, 1580, 171, 554, 457, 2145, 3795, 3593, 711, 2454, 446, 2852, 2875, 257, 268, 620, 923, 775, 2867, 3890, 3456, 2291, 494, 2682, 1535, 237, 3120, 3036, 2710, 2374, 2154, 3536, 1599, 3484, 3540, 1198, 309, 228, 1510, 2797]. Particle-Locating [923]. Particle-Mesh [2875]. Particle-Swarm-Optimization-Based [2797]. Particleboard [719]. Particleboards [750]. Particles [3453, 401]. Partite [767]. Partition [1261, 3853, 1843, 1234]. Partitioned [1788, 1708, 1414]. Partitioning [1680, 906, 921, 3393, 931, 2883, 1245, 2393, 749, 1668]. Partitions [908, 942, 486]. Parts [3796, 3883, 2026, 1862]. Party [1666]. Passage [1263]. Passengers [2830, 2445]. Passing [2592, 1977, 402]. Passivated [2954]. Passive [894, 3352]. Past [356]. Patch [1611, 2567, 657, 3079]. Patent [149]. Patents [1686]. Path [995, 3328, 1936, 2879, 1133, 207, 2805, 3638, 747, 3398, 369, 1562, 2390, 1444, 2592, 3077, 2595, 3156, 1035, 1775, 1684, 2056, 1802, 3517, 2713, 884, 979, 3358, 2841, 3640, 2130, 3293, 3655, 2303, 2854, 3542, 658, 575, 2082, 3652, 2245, 1294]. Path-Connected [3328]. Path-Planning [3293]. Path-Width [3638]. Pathfinding [1585]. Pathogen [27]. Pathological [3348]. Pathology [151, 619, 2559]. Pathology-Free [151]. Pathology-Specific [2559]. Paths [1136, 159, 301, 1281, 1957, 3098, 1665, 259, 3812, 3344, 1683, 2349, 2982]. Pathway [3797, 44]. Pathways [2732]. Patient [3003, 2732, 3727, 3446, 735, 1689, 3229, 2835, 1259]. Patient-Level [3229]. Patient-Specific [3446]. Patients [1289, 2341, 2663, 2518, 3346, 2938, 1236, 2159, 2460, 3501, 3563]. Pattern [195, 278, 2497, 556, 3138, 2543, 38, 817, 3212, 1615, 508, 1151, 1483, 458, 3888, 422, 121, 1752, 2880, 2713, 3837, 3460, 3904, 198, 44, 1718]. Pattern-Guided [195]. Pattern-Matching [458]. Patterns [3582, 2301, 3063, 176, 434, 2435, 2265, 1065, 3187]. Pavement [1140, 1023, 2535, 3621]. Paxos [672]. Payload [2356]. Payment [3240]. Payments [2837]. PBL [3892]. PC [1988]. PCA [2211, 2907, 1278, 3285]. PCA-LSTM [2907]. PD [991, 983, 2945]. PDASTSGAT [2802]. PDE [1903]. PDE-Based [1903]. PDEC [2786]. PDEs [2411, 2194, 341]. PDI [966]. PE [1931]. Peak [3795, 614, 154, 3788, 3675, 513, 3739]. Peaking [3483]. Peaks [1680]. Pedestrian [184, 2637, 3113, 742, 2696]. Pedestrian/Bike [184]. Pedestrians [1239]. Pediatric [3350, 3070, 2828]. Peelle [113, 108]. Peer [2446, 688, 1909]. Peer-to-Peer [2446, 688]. Pelton [3287]. Pelvic [2593]. PEM [1972]. PEMFC [3574]. PEN [1691]. Penalized [167]. Penalties [1637]. Penalty [2347, 701, 256]. Pendulum [1818, 3120, 3547]. Penetrating [1834, 3831]. Penetration [453]. Peninsular [3188]. Penrose [1980]. Pepper [751]. Peptide [2904]. Peptides [3066]. Perception [2821, 3283, 410, 2995, 460]. Perceptron [2407, 3047]. Perceptual [537]. Perch [1761]. Percutaneous [3810]. Perfect [797, 3227, 2372, 1626, 608]. Performance [1901, 3038, 981, 3493, 531, 888, 3259, 1463, 2161, 236, 3136, 3439, 1262, 2895, 746, 1072, 617, 1833, 1039, 2520, 1240, 2814, 1955, 2167, 2666, 3289, 1559, 138, 1144, 3058, 1302, 3794, 3883, 1252, 621, 1931, 3824, 3103, 1800, 2098, 2640, 174, 1907, 2721, 1348, 3298, 3463, 801, 2113, 1721, 2562, 1899, 2963, 725, 3905, 1466, 3352, 2031, 2366, 1441, 378, 3370, 2853, 3910, 3326, 2786, 1707, 2300, 1050, 2610, 326, 501, 1474]. Performance-Constraint [2853]. Perfusion [1215]. Perimeter [3108]. Period [3605, 977, 1292, 949, 2612]. Period-Aggregated [949]. Periodic [343, 2940, 2016, 3171, 154, 1976].

Periodically [477]. Periodicity [2202, 2758]. Peripheral [3854]. Perishable [699]. Permit [3179]. Permutation [1360, 928, 1458, 844]. Permutations [1423]. Permuted [817]. Perpetual [1263]. Persistence [2538, 1156]. Persistent [962, 1228, 1643, 2244]. Person [2406, 3009, 1385, 3785, 1253, 1724, 1135, 1053, 3645, 1604]. Personal [3324, 1713]. Personalised [3858]. Personality [1691, 3527]. Personalized [3503, 3362, 2885, 2017, 495, 3235, 2684, 421, 1676]. Perspective [1902, 3014, 3000, 1681, 3317, 286]. Perspectives [2536, 1021, 2740, 3588, 3808, 1248, 2183, 53]. PERT [2784]. PERT/CPM [2784]. Pertinent [113, 108]. Perturb [3191, 1284]. Perturb-and-Observe [3191]. Perturbation [285, 3902, 2939]. Perturbations [3392]. Perturbed [717]. Pessimistic [607]. Pest [1046]. Pesticide [3203]. Pests [1531]. PET [3086, 3338, 1071, 3175]. PET-CT-Derived [3338]. PET/CT [3086]. Petri [695]. PFCP [3701]. PFMEA [3594]. PFSegIris [1509]. Phantom [3175, 3125]. Pharmacology [3848, 3116]. Phase [3191, 2659, 510, 1366, 521, 861, 2004, 648, 1622, 3230, 2974, 1755, 1412, 811, 1149, 3244, 173, 996, 1930, 3530, 3827, 2871, 1986, 3547, 1395]. Phase-Sensitive [648]. Phase-Type [2004, 3827, 1986]. Phases [3493]. PHEFT [607]. Phenology [5]. Phenomena [1018]. Phenomenon [510]. Pheromone [2334]. Phishing [2449, 3911]. Photoacoustic [2207]. Photobioreactor [993]. Photographic [2613]. Photomatrix [984]. Photometric [2458]. Photoplethysmography [3347]. Photovoltaic [3366, 1942, 1678, 3090, 2683, 3076, 1876, 2459, 1284, 3745, 2954, 2931]. Phylogeny [2416, 797]. Phymastichus [3889]. Phymastichus-Hypothenemus [3889]. Physeter [70]. Physical [2973, 29, 1939, 3090, 1730, 2051, 3836, 2168, 365, 2936]. Physicist [2429]. Physics [3262, 3072, 2511, 2388, 2869, 2277, 1657, 3897, 2643, 1655, 2389, 2333, 3586, 2074, 3913, 2936]. Physics-Based [1655]. Physics-Enhanced [3897]. Physics-Guided [2936]. Physics-Informed [2511, 2388, 2277, 1657, 2643, 3586, 2074]. Physiological [3822, 3885, 3250]. Pi [3020, 3643, 617, 1457, 1367, 626, 2945, 798, 3762]. PI-D [3643, 2945]. PI3K [3797]. PI3K/AKT [3797]. Piano [1750]. Picard [416]. Pick [1171, 1593]. Pick-up [1171]. Picker [2026]. Picker-to-Parts [2026]. Picking [1442, 1593, 2026, 1718]. Pickup [2988, 902]. Picture [2646]. PID [1143, 692, 2618, 2118, 3643, 3315, 696, 690, 2715, 1426, 293, 1219, 612, 1441, 425, 579, 3484, 1421]. PID-Based [3315, 1426]. PIDD [2520]. Piecewise [3136, 76, 1108, 101, 1919, 3587, 2804, 2316, 270]. Piecewise-Linear [270]. Piecewise-Stationary [2316]. Piezoelectric [1365]. Pigeon [747]. Pigeon-Inspired [747]. Pigeonhole [2069]. PigSNIPE [2199]. Pile [1495]. Pile-Group [1495]. Pilot [3040, 3446, 3892]. Pinhole [3004]. PINN [2277, 1657]. Pinned [3889]. Pinning [3889]. PINNs [2074]. Pipe [996]. Pipeline [3852, 1047, 2997, 1140, 773, 3175]. Pipelines [3004, 2428, 2137]. Pipes [3115]. PIPET [3175]. PISI [3836]. Piston [2203]. Pitfalls [2993, 598]. Pivot [675]. Pixel [1880, 2125]. Pixel-Level [1880]. Pixel-Wise [2125]. Placement [1732, 2524, 1237, 345, 3805, 2952, 911, 691, 1131]. Placements [1467]. Plagiarism [469]. Planar [2042, 162, 3118, 2981, 3539]. Planarity [1136]. Plane [1999, 1768, 3627, 2492, 3547]. Planned [3662, 1644]. Planner [3461, 2445]. Planning [788, 1264, 3275, 2127, 2767, 2834, 683, 945, 1097, 1127, 747, 3398, 887, 1468, 2595, 1035, 1961, 1775, 2961, 1786, 3517, 959, 2841, 1059, 2028, 2469, 2197, 2570, 3640, 3293, 3655, 2574, 2303, 2854, 658, 3305, 575, 864, 2082, 3632, 3547, 3490]. Plans [2012, 577, 3490]. Plant [393, 3336, 3076, 3514, 1219, 2525, 3685]. Plant-Integrated [3514]. Plantar [3142]. Plantower [2475]. Plants [25, 3639, 3497, 1127, 2520]. Plasma [2427]. Plastic [3773, 2601]. Plate [1549, 2919, 1450, 3712, 3263, 736, 641]. Plates [3211, 3343]. Plates/Cylindrical [3211, 3343]. Platform [3013, 1670, 1044, 2788, 1369, 273, 766, 1713]. Platforms [1833, 1249, 3240, 2354]. Plausibility [2292]. Play [2653]. Player [1268]. Players [3313]. Playground [2318]. Plenoptic [1463]. Plentiful [440]. Plethysmography [2918]. Pliant [1258]. PLM [2801]. PLM-ICD [2801]. PLS [2232]. PLS-ECOC [2232]. Plug [2653]. Plug-and-Play [2653]. Plugging [3628]. Plum [2217]. Plus [617]. PM [3419, 1103, 2135, 2475]. PM2.5 [1456]. PMOD [3309]. PMS6MC [201]. PMSM [798, 3811, 2487]. PMU [2848]. Pneumonia [3350, 3334]. PnP [653]. POD [3618]. POI [3628]. Poincaré [3172]. Point [3429, 1020, 1942, 1180, 3730, 3774, 1583, 1450, 635, 2815, 3869, 2960, 102, 944, 338, 1024, 2554, 1070, 2109, 52, 2459, 744, 1311, 2364, 3007, 1001, 2138, 1745, 3118, 2399, 3404, 996, 1471, 3913, 1985, 1728, 3598, 1693, 2815]. Point-of-Interest [3730]. Point-Sim [2815]. Point-Wise [3913]. Pointing [1633]. Points [1234, 2685, 1480, 1767, 360, 3624, 2492, 2874, 1545, 682]. Pointwise [3424]. Poisoning [2812, 2623]. Poisson [858, 3421, 1620]. Poisson-Inverse-Gaussian [3421]. Polak [664]. Polar [936, 72]. Polarization [3026, 3560]. Polarized [3026]. Polarizing [33]. POLCA [3592]. POLCA-Constrained [3592]. Pole [911, 634]. Policies [3147, 1465]. Policy [2975, 3332, 1174, 2408, 1346]. Policy-Based [1174]. Polishing [3305]. Political [2260, 3560]. Pollination [2751, 1978, 1478]. Pollution [1741, 3565, 3895, 2383]. Poly [1919]. Poly-Sinc [1919]. Polyacrylamide [2806]. Polyak [664]. Polychrony [2276]. Polygon [948, 2364]. Polygonal [66, 2682]. Polygons [131, 2794]. Polyharmonic [3171]. Polyharmonic-Neumann [3171]. Polyhedral [1108, 923]. Polymatrix [1576]. Polymer [1929, 3883, 1616]. Polynomial [166, 1201, 1205, 853, 136, 2227, 3098, 1032, 2053, 145, 3458, 1705, 1098, 932]. Polynomial-Correction [2053]. Polynomial-Time [166, 136, 2227]. Polynomials [3038, 1788, 3718, 2373, 2380, 3224, 3330, 3046]. Polyomino [1669]. Polyominoes [1768]. Polyp [3294, 3542]. Polyps [84]. Polytopic [764, 1453]. Pooling [2003, 2739, 1026]. Poor [1574]. Population [803, 1940, 2767, 2906, 602, 1044, 869, 759, 1771, 1982, 1393, 1627]. Population-Based [1393, 1627]. Populations [181, 3403]. Porcupine [3891, 2366]. Pore [1619, 3320]. Pore-Scale [1619]. Pores [2125]. Poroelastic [3831]. Porous [2107, 861, 1008]. Portable [1546]. Portfolio [169, 521, 2145, 3205, 443, 3636, 2015, 603, 1500, 493]. Portfolios [3361]. Portrait [3395]. Ports [3145]. PoS [2548]. Pose [2043, 1239, 3457, 2034, 653, 3645]. Pose-Based [2043]. Posed [754]. Posedness [1112]. Position [1031, 991, 799, 3567, 75, 2791, 19, 293, 961, 1512, 1742]. Position-Based [19]. Positional [2668, 436]. Positioning [3429, 1551, 2010, 2205, 3690, 3269, 1311, 2983, 3922, 3251]. Positions [50, 3077]. Positive [1646, 1327, 908, 3295, 3480]. Positive-Instance [1646]. Possibilities [531, 3448]. Possibility [3689]. Possible [2053]. Post [3086, 2798, 2610, 486]. Post-Processing [2798, 486]. Post-Quantum [2610]. Post-Radiation [3086]. Posterior [2768]. Postoperative [3810]. Postprandial [3900]. Posture [3822, 2023]. Potential [3063, 2089, 2903, 1183, 2249, 3444, 3217, 20, 1799, 3116]. Potra [267]. PoW [2548]. Power [1989, 2536, 3054, 2618, 1937, 1642, 3262, 1860, 1463, 1942, 1147, 3639, 970, 3394, 3090, 1127, 678, 1018, 2112, 3076, 1687, 3448, 3301, 3783, 2050, 1855, 3154, 3030, 2272, 3511, 1411, 3747, 3514, 929, 3891, 2058, 3573, 3748, 3106, 20, 1964, 3897, 2459, 801, 2282, 1375, 1618, 652, 1403, 233, 3414, 1001, 1538, 2028, 1539, 1454, 1974, 254, 3121, 1099, 837, 487, 2890, 1501, 3179, 2571, 2651, 3483, 3248, 623, 3835, 2417, 3574, 453, 769, 3548, 768, 2891, 3060]. Power-Aware [678]. Power-Law [2282]. Power-Spectrum [1099]. Powered [1549, 3455, 3907]. Powerful [2618]. PPG [2213, 3327]. PPP [3642]. Practical [175, 2252, 916, 386, 1045, 2335, 3754, 1920, 2191]. Practicality [1431]. Practice [1757, 2970, 3014, 1468]. Pre [2973, 1580, 1068, 2412, 2941, 1854, 3182, 2346, 3717]. Pre-Detonation [1580]. Pre-Evaluation [3182]. Pre-Implementation [1854]. Pre-Specified [2412]. Pre-Trained [2973, 1068, 2941, 2346, 3717]. Preamble [3349]. Preamble-Based [3349]. Precedence [2669, 3002]. Precedence-Constrained [2669]. preCICE [1414]. Precipitable [3429]. Precipitation [457, 663]. Precise [3429, 1509, 2500]. Precision [3797, 2001, 2916, 3690, 809, 3700, 198, 1659, 2655]. Precision-Based [1659]. Precision-Driven [2916]. Precoding [890]. Preconditioned [2757, 455, 423]. Preconditioners [1042, 1141]. Preconditioning [2411, 1606, 1452, 3559]. Precursor [2966, 1380]. Predator [803, 2690]. Predator-Prey [2690]. Predators [3783]. Predatory [1647]. Predefined [2927, 1373]. Predefined-Time [2927]. Predicate [2786]. Predicates [76]. Predict [3066, 3333, 2474, 791, 1697, 1559, 138, 3557, 2942, 43, 2490, 3235, 144, 1691, 3117, 3065, 1342]. Predicted [991]. Predicting [2663, 2446, 2623, 139, 236, 2820, 1400, 3413, 2791, 2340, 1239, 2090, 2700, 77, 2449, 3657, 2637, 1684, 548, 2465, 2498, 2599, 2645, 2602, 3563, 3801, 2823, 3810, 3368, 82]. Prediction [3347, 3397, 1175, 3082, 2391, 1435, 89, 3852, 2171, 1699, 241, 2888, 1202, 3041, 3419, 3599, 3439, 243, 1702, 669, 400, 1499, 3338, 3693, 1811, 457, 1462, 2683, 1526, 3757, 529, 1951, 126, 2159, 3799, 790, 3055, 781, 1244, 1208, 1128, 2500, 2832, 2673, 3744, 2550, 2909, 3390, 620, 917, 1723, 2369, 2978, 3624, 2638, 1085, 1907, 2917, 1681, 1822, 2735, 3900, 3284, 3901, 982, 3470, 3237, 757, 1540, 2405, 2339, 2223, 2070, 3667, 676, 1982, 3609, 3533, 247, 580, 2835, 3723, 3536, 1449, 227, 1427, 2955, 3319, 3571, 3585, 118, 999, 756, 964, 1436, 1690, 963, 1456, 1350, 2381, 3679, 1548, 3717]. Prediction [2891]. Prediction-Based [3624]. Predictions [3090, 1116, 3813]. Predictive [995, 2470, 2990, 3063, 1092, 3639, 3342, 1039, 2045, 1803, 1496, 3722, 2055, 1791, 3332, 3310, 3897, 1270, 1122, 1153, 1334, 3033, 3906, 3814, 3326, 623, 3710, 2656, 3546]. Predicts [1395]. Preemptible [513]. Preemptive [130, 985]. Preference [1154, 2578, 202, 177, 2902, 1332, 703, 3507]. Preference-Driven [2578]. Preferences [215, 202, 896, 200, 685, 565]. Prefix [955]. Pregnancy [2095]. Preliminary [3440, 49, 1657, 2122]. Premium [1058]. Preparation [2789]. Preprocessing [2375, 2302, 1014, 350, 1784, 2547, 2227, 2776, 2135, 2989]. Preptimize [2989]. Prescribed [2240]. Presence [1662, 3689, 3142, 2853]. Present [2837, 2884]. Presentation [1318]. Presentations [3777]. Presented [1383]. Preservation [2882, 408]. Preserved [2567]. Preserving [1842, 727, 1124, 1806, 3732, 1134, 822, 2533, 1828, 961, 2327]. Press [807]. Pressure [3629, 224, 3059, 1858, 2370, 2785, 3142, 472, 3472, 3801, 425]. Pressure-Correction [2370]. Pretrained [2091]. Pretraining [3832]. Prevention [2555, 1472]. Preventive [2290]. Preview [764]. Prey [803, 2690]. Price [2621, 139, 1499, 2050, 2008, 1859, 74, 985, 3609, 501, 2307, 2891]. Prices [2260, 3333, 1061, 3413, 1829, 2599, 3723, 2077]. Pricing [1268, 699, 1776, 2115, 1541, 768]. PRIM [3232]. Primal [1311]. Primal-Dual [1311]. Primary [2518, 3198, 611, 348]. Primavera [2887]. Prime [2814, 2948]. Primitives [1546]. Principal [1948, 3866, 3381, 347, 1571, 2484, 435, 2397]. Principle [1058, 256, 2069]. Principles [24]. Printable [3756]. Printed [2067, 2243]. Printing [1862]. Prior [3809, 988]. Priori [2126, 3053]. Prioritization [3196, 709]. Prioritizing [1502]. Priority [818, 3673, 3043, 1250, 1658, 3512, 1125, 703]. Privacy [2087, 1842, 2950, 2461, 727, 1124, 3732, 447, 822, 2533, 2882, 1828, 1831, 2327, 408, 3545]. Privacy-Preserving [1842, 3732, 822, 2533, 1828, 2327]. Private [3284, 2327]. Proactive [3514, 3033]. Probabilistic [1086, 1425, 2870, 662, 2435, 49, 437, 1571, 18, 1536, 651, 357, 3490]. Probabilities [2473, 534, 1810, 1986]. Probability [3340, 2953, 1583, 775, 1531, 2393, 3467, 1106, 1258, 332]. Probability-Based [3467]. Probing [2588]. Problem [1289, 2404, 562, 2485, 1575, 2116, 627, 1732, 2988, 3026, 1936, 1360, 2173, 1430, 2455, 1442, 1818, 2894, 783, 3084, 2584, 1327, 1944, 1945, 2127, 120, 600, 311, 2720, 2434, 386, 500, 3031, 2541, 1169, 225, 3087, 818, 2792, 2669, 631, 1127, 1556, 695, 2544, 1299, 3207, 443, 567, 1951, 1399, 3561, 2799, 329, 190, 688, 754, 1240, 586, 1080, 2355, 2630, 1458, 1184, 3255, 896, 1825, 3410, 978, 1585, 2290, 3872, 251, 3417, 164, 136, 1561, 1562, 2591, 887, 295, 1157, 1287, 1378, 380, 1091, 944, 844, 221, 1187, 1251, 1658, 581, 2729, 3755]. Problem [4, 987, 794, 248, 1657, 297, 977, 1171, 162, 2098, 2640, 1572, 3228, 1515, 3302, 2131, 58, 2123, 3778, 2660, 1592, 2348, 1354, 1439, 902, 2598, 382, 1311, 2723, 948, 361, 1095, 1900, 2265, 684, 855, 2857, 3007, 2876, 2299, 1152, 1507, 256, 1845, 1371, 1541, 68, 2031, 3118, 1054, 985, 1973, 173, 3479, 1038, 597, 947, 1494, 857, 605, 2570, 949, 3005, 394, 2374, 925, 2848, 1195, 1296, 3053, 2982, 1676, 2130, 870, 3539, 904, 3183, 217, 493, 197, 873, 762, 789, 3549, 186, 3252, 3137]. Problem-Driven [2541]. Problems [2780, 698, 1990, 1523, 3412, 3650, 726, 3131, 3, 152, 512, 1429, 510, 2439, 3200, 171, 2541, 322, 261, 2879, 933, 1117, 1949, 615, 3206, 1557, 2489, 1343, 430, 3078, 1669, 2477, 1607, 192, 1917, 209, 3210, 1527, 2549, 55, 1272, 343, 2334, 2194, 1084, 951, 969, 576, 3153, 2869, 1663, 3278, 2778, 1483, 1909, 3256, 406, 2868, 3416, 1634, 1188, 541, 3103, 2329, 3896, 2019, 1613, 2560, 3754, 398, 1796, 3633, 2025, 2186, 2837, 200, 3170, 1594, 1292, 2389, 2151, 1705, 1475, 2032, 1973, 3241, 3641, 1132, 1978, 2744, 346, 3245, 2728, 1501, 3180, 1050, 1394]. Problems [2075, 47, 2707, 821, 805, 3584, 3791, 618, 244, 1809, 524, 3618, 206, 549, 1916, 479, 664, 932, 1294]. Procedure [2341, 1936, 252, 386, 853, 3141, 3445, 3223, 1318, 1594, 399, 993, 1099, 2072, 3481, 878]. Process [3262, 1757, 3401, 661, 449, 1425, 514, 1202, 2165, 1234, 1803, 1103, 2140, 2520, 1611, 1273, 2845, 2330, 1631, 2809, 2658, 3301, 2378, 2759, 2447, 3220, 3335, 1791, 3694, 397, 1800, 3331, 2889, 935, 3339, 3320, 3109, 2601, 2061, 2902, 1883, 1777, 2122, 3634, 2266, 973, 594, 1731, 1977, 2073, 1706, 2873, 3827, 2428, 2137, 3917, 3399, 1229]. Process-Based [1202]. Process-Oriented [1631]. Processable [1029]. Processes [1339, 1092, 1902, 2246, 3382, 3204, 3139, 1784, 2140, 2791, 318, 3448, 1958, 3882, 1070, 419, 2282, 3464, 2139, 3764, 1194, 1856, 1778]. Processing [3271, 3849, 1744, 3697, 2304, 2199, 336, 636, 3774, 1124, 607, 1205, 2046, 2827, 3732, 598, 648, 238, 2175, 321, 3022, 1250, 621, 1871, 2807, 3894, 2788, 54, 3274, 2269, 2343, 2603, 2067, 2798, 597, 1494, 1257, 2206, 1402, 286, 391, 3247, 486, 1546, 3570, 2619]. Processing-Based [3022]. Processor [2956]. Processors [458]. Producing [740, 3151, 763, 833]. Product [139, 147, 2898, 3499, 1917, 1696, 2052, 1786, 977, 2469, 2716, 474, 2157]. Product-Driven [2469]. Production [2295, 3619, 887, 1287, 1250, 1786, 3104, 3000, 2755, 2939, 2469, 1644, 2183, 1979, 3916, 864]. Productivity [1462, 1502, 2703, 1777, 3770]. Products [2093, 2178]. Professional [2659, 3659]. Professionals [3065]. Professions [3324]. Profile [1450, 3537]. Profiles [1873, 2484]. Profiling [13]. Profit [271, 1508, 1301]. Profit-Based [1508]. Profitable [2988]. Prognosis [1698, 3487]. Prognostics [1007, 2021]. Program [1928, 2719, 1249, 3392, 2944]. Programing [1188]. Programmable [3013, 2774]. Programming [1154, 820, 916, 526, 3061, 311, 629, 2541, 2724, 3207, 1329, 1768, 335, 2477, 209, 2745, 3410, 2262, 2368, 1438, 887, 132, 1663, 7, 2961, 1365, 2123, 2190, 2837, 564, 237, 2299, 93, 2222, 2605, 1482, 94, 40, 2944, 1467, 1707, 3180, 1155, 724, 1066]. Programs [1805, 2960, 2592, 88]. Progress [3204, 1005, 599, 3190]. Progressive [2510, 2773]. Project [3566, 386, 1105, 1825, 2832, 1250, 2784, 2837, 757, 2478]. Projected [884]. Projection [2996, 3718, 746, 2370, 338, 1024, 2923, 1921, 1749, 112]. Projections [1075, 2083]. Projects [2431, 3008, 2386, 1254, 2887, 368, 2134]. PROMETHEE [1388, 3862]. PROMETHEE-SAPEVO-M1 [1388]. Promote [2446, 94]. Prompt [3869, 3444, 2935]. Prompt-Based [3869]. Prompted [2149]. Prone [945]. Proof [1654, 1153, 1334, 937]. Proof-of-Concept [937]. Propagating [23]. Propagation [3714, 1241, 369, 3874, 1485, 1802, 11, 1300, 996, 2877]. Proper [1576, 1090, 1271]. Properties [1867, 1730, 323, 1703, 563, 2231, 2386, 2240, 253, 2020, 3464, 223, 1893, 2649, 1066]. Property [207, 2006, 3799, 1489, 69]. Property-Based [1489]. Prophet [3419]. Proportional [981, 3197, 2355, 3694, 2963, 2816]. Proportional-Integral-Derivative [3197, 3694]. Proportionality [3736]. Proposal [3681, 3141, 3815, 3468, 1970, 2067, 2392]. Proposed [2962, 758]. Propulsion [378]. Prospect [787]. Prospective [2911, 1359, 676]. Prospects [3726, 2245]. Prostate [1395, 3446, 3501, 1653, 2893]. Prosthesis [2692]. Prosthetic [3894, 3904]. Protecting [3430]. Protection [2006, 447, 827, 3248, 3724, 3545]. Protein [303, 3066, 176, 15, 32, 67, 2806, 1288, 253, 645, 1416, 1355, 2346, 3717, 1310]. Protein-Binding [3066]. Protein-Protein [32, 253]. Proteins [3859, 790, 2010, 1690, 1350]. Protocol [51, 1654, 3304, 2865, 1457, 306, 814]. Protocols [182, 2192]. Protograph [914]. Protonet [3771]. Prototype [103, 498, 2700, 608]. Prototype-Based [498]. Prototyping [2203]. Provable [1592]. Provably [1412]. Provide [1791, 2976]. Providing [3013, 3711]. Proving [3464, 1745]. Provisioning [3155, 3454, 744]. Proximity [3437, 646]. Pruning [151, 2481, 2705, 851, 1827, 2736, 1667]. Pseudo [1218, 1260, 1926, 3270]. Pseudo-Images [1260]. Pseudo-Spectral [3270]. Pseudoinverse [3094]. Pseudorandom [2496]. PSO [2852, 3427, 610, 3035, 266, 457, 1462, 2326, 2390, 1565, 293, 47, 1762]. PSO-Based [3427]. PSO-CS [266]. PSO-MPC [2326]. PSO-SVM [610, 457]. PSPACE [106]. PSPACE-Hard [106]. Psychiatric [2038]. Psychological [1939]. Psychosomatic [2986]. Pták [267]. PTAS [4, 162]. PUB [1068]. PUB-SalNet [1068]. Public [2781, 2234, 2816, 1927, 2237, 1356, 3046]. Publication [1763]. Publications [1049]. Publishing [2950]. Puffin [3206]. Pulmonary [2818, 91, 819, 260]. Pulse [1147, 1230]. Pulsed [1820]. Pump [1030]. Pumping [859]. Pumping-Unit [859]. Pumps [2203]. Pure [3386]. Purification [2915]. Purpose [2870, 2128, 1253]. Purposes [976, 3774, 3468, 2405]. Pursuit [2457, 112, 435, 705, 780]. Pushing [2252]. Put [3413]. Puzzle [124, 113, 108, 136, 277]. Puzzles [137]. PV [3890, 3748]. Pyraformer [3382]. Pyraformer-Deep [3382]. Pyramid [3529, 3542]. Pyroelectric [245]. PySpark [3878]. Pythagorean [536]. Python [2994, 3564, 3141, 2338]. Python-Driven [3564].

Q [3454, 3512, 3458, 3402, 2847]. Q-Convergence [3458]. Q-Gram [3402]. Q-Learning [3454]. Q-Network [3512, 2847]. Q8S [3576]. QAM [3244]. QAOA [1527]. QB4MobOLAP [1513]. QEMU [3576]. QFC [1894]. QoE [1826]. QoS [3359, 812]. QoS-Aware [3359]. QPSO [2983]. QR [3787, 2243]. qRobot [1442]. QSAR [1368]. Quad [2921, 1606]. Quad-Tree [2921]. Quadratic [353, 2616, 324, 3768, 313, 3207, 2231, 3228, 2187, 992, 3015]. Quadrature [3316, 152, 1976, 3626]. Quadrilateral [2373]. Quadrilaterals [72]. Quadrotor [717, 2205, 2312]. Quadrotors [718]. Qualify [1036]. Qualitative [3459, 2421, 2939, 2521]. Quality [1049, 2781, 3862, 1181, 874, 410, 460, 2795, 3264, 1825, 2378, 3569, 3511, 2013, 3843, 3746, 1324, 3845, 144, 227, 2179, 3490]. Quantification [107, 1138, 1072, 1032, 994, 3684, 2566, 1849, 2971, 1098, 1497]. Quantify [1062, 1518]. Quantifying [3501, 3886, 2384]. Quantiles [2510, 2881]. Quantitative [779, 190, 3502, 1278, 2521, 3309]. Quantization [2045, 287, 1852, 451, 469, 641, 2114]. Quantized [2049, 936]. Quantum [1666, 1442, 1435, 51, 3428, 2820, 3408, 2984, 1204, 931, 1806, 2396, 778, 3732, 747, 3449, 3753, 3826, 3688, 2739, 882, 2415, 2690, 1960, 2687, 2324, 3039, 2737, 3406, 3525, 1758, 2233, 2610, 821, 2871, 1916, 3743]. Quantum-Behaved [747, 882]. Quantum-Classical [3525]. Quantum-Inspired [2690, 2687, 2737]. Quantum-Resistant [3732]. Quark [2427]. Quarter [3352]. Quartically [274]. Quasi [1052, 1353, 1076, 2454, 1610, 2852, 3587, 2022, 398, 3721, 3171, 154, 2575]. Quasi-Arithmetic [2022]. Quasi-Convex [3721]. Quasi-Hole [1353]. Quasi-Interpolation [3587]. Quasi-Monte [1052]. Quasi-Newton [2454, 1610, 2575]. Quasi-Nonexpansive [398]. Quasi-Periodic [3171, 154]. Quasi-Random [2852]. Quasi-Static [1076]. QUATRE [1421]. Quay [3145]. Qubit [3753, 3826, 2871]. QUBO [1637, 1796]. Qudit [2233]. Queries [1158, 1306, 238, 1151, 1295, 452, 1274, 910, 2686]. Query [3136, 1124, 2903, 572, 2011, 429, 743, 725, 851, 1412, 3376, 1397]. Query-Efficient [2011]. Querying [336, 321]. Quest [2911]. Question [2935, 3396]. Questionnaires [2795]. Questions [3848, 3128, 3116]. Queue [3750, 3276]. Queueing [3139, 877, 3482]. Queues [3673]. Queuing [1864]. Quick [1309, 219]. Quintic [438]. Quotient [3146]. Quotients [3146].

R [614, 742, 3273, 1630, 3739]. R-CNN [3273]. R-CNN-Based [742]. R-Peak [3739]. R2D2 [1151]. Rabbit [3359]. Radar [1834, 2453, 1905, 2126, 1012, 384, 3594]. Rademacher [1065]. Radial [2221, 1261, 2544, 3694, 62, 1964, 3047, 3481, 3791]. Radiation [29, 1994, 2624, 3086, 683, 2287, 857]. Radio [13, 28, 115, 3861, 119, 348, 496, 228]. Radio-Frequency [115]. Radio-Isotope [28]. Radiocommunication [2738]. Radioelectric [405].

Radiograph [2665]. Radiological [82]. Radiometers [115]. Radiometry [71]. Radiomic [3338]. Radiopharmaceuticals [3309]. Radiotherapy [3501]. Radish [2673]. Radius [416, 2080]. Radon [3001]. RAG [3797]. RAG-Driven [3797]. Rail [1569, 3571, 1512, 1742]. Railroad [3024]. Railway [2636, 3029]. Railways [1552]. Rainbow [251]. RainForest [3354]. Rallies [3560]. Random [1823, 3718, 1578, 2172, 278, 139, 928, 914, 156, 554, 320, 3088, 1558, 3369, 1076, 49, 2658, 1559, 1560, 2852, 1518, 3232, 3320, 2641, 1218, 1292, 3906, 3392, 3370, 2376, 1303, 505, 2035, 2905, 3603, 1986, 3598, 3921, 1718]. Random-Field [1076]. Randomized [3, 1888, 1794, 951, 1658, 794, 977, 902, 2703, 612]. Randomly [2679]. Randomness [786]. Range [620, 3518, 2964]. Rank [3614, 1042, 240, 204, 371, 1610, 3214, 285, 1921, 351, 1099, 3568, 712, 1849, 2083]. Rank-1 [1610]. Rank-Constrained [1921]. Rank-Order-Normalization [1099]. Rank/Select [240]. Ranking [2301, 3862, 2234, 3096, 130, 1330, 3660, 2917, 3186, 1763]. Rankings [763]. RANSAC [216]. Rape [3680]. Rapid [2172, 2203]. Rapidly [3921]. Rapidly-Exploring [3921]. Rasch [3473]. Raspberry [3020, 3762]. RASPF [3577]. Raster [2385]. Rat [3080]. Rate [3347, 3338, 2666, 272, 1857, 728, 2516, 12, 614, 1851, 3577, 2216, 270, 332, 3752]. Rate-Adaptive [3577]. Rateless [895]. Rates [1633, 1244, 3702, 3426]. Rating [883]. Ratings [1116, 3235, 3865]. Ratio [3453, 3349, 58]. Rational [1661, 2677]. Rationality [585]. Ratios [2424, 1316, 2934]. Rats [1867]. Ray [3437, 1178, 3386, 539, 2937, 1431, 2322, 1570, 2557, 54, 260, 99, 100]. Ray-Shooting [539]. Rays [808, 3334, 2629]. Razgon [1334]. RBF [1444, 3047, 2569, 2154, 1789]. RBFNN [1512]. RCF [1817]. RDF [330, 2101]. RDF/S [2101]. Re [3009, 1389, 3091, 586, 1265, 1385, 2923, 1253, 1724, 3113, 1135, 1053, 3621, 1604]. Re-Embedding [3621]. Re-Entry [1389]. Re-Identification [3009, 3091, 1385, 1253, 1724, 3113, 1135, 1053, 1604]. Re-Orthogonalized [2923]. Re-Orthogonalized/Affine [2923]. Re-Pair [1265]. Re-Scheduling [586]. Reachability [1011]. Reaction [1040, 1619, 2198]. Reaction-Diffusion [1040]. Reaction-Dominated [1619]. Reactive [630, 2112, 1619, 2713, 1964]. Reactor [3694, 1085]. Readable [3558, 2486]. Readers [2589]. Readiness [3892, 2568]. Reading [2352, 427, 1965, 2275]. Real [3642, 3429, 1549, 1498, 2144, 2966, 1732, 3493, 1744, 783, 556, 2356, 791, 3610, 683, 3495, 2629, 2224, 1351, 586, 2438, 2704, 3783, 1022, 2690, 1754, 1563, 401, 3282, 3843, 2742, 1800, 3893, 3894, 2203, 3460, 2599, 2483, 3274, 614, 2177, 852, 2186, 1190, 1333, 3402, 1972, 2772, 2535, 3176, 2072, 628, 3345, 782, 518, 789]. Real-Estate [3345]. Real-Life [783]. Real-Parameter [2690, 401]. Real-Time [3642, 3429, 1549, 1498, 2144, 2966, 3493, 1744, 2356, 3610, 683, 3495, 2224, 1351, 2438, 3783, 1022, 3843, 1800, 3893, 3894, 2203, 2599, 2483, 3274, 614, 852, 1190, 3402, 1972, 2772, 2535, 3176, 628, 518, 791, 782].

Real-World [1732, 556, 2704, 2742, 2186, 789]. Realistic [320, 3727, 3109, 3537]. Reality [2692, 3145, 884, 1341]. Realization [2098]. Realizations [667]. Realize [841]. Realized [131]. Reallocation [3466]. Rear [2954]. Rearrangements [1417]. Reasoning [2188, 1941, 2242, 1670, 1805, 3558, 1631, 1438, 2786]. Reasons [2292]. Receive [3688]. Receiver [127, 2336, 3244]. Receivers [2659]. Recency [2479]. Recently [511]. Recently-Introduced [511]. Recipes [1318]. Reciprocal [706]. Reciprocal-Selection-Based [706]. Reclaiming [3838]. Recognition [3629, 1549, 2375, 3846, 2144, 1493, 2173, 2899, 2481, 278, 2497, 556, 3781, 2215, 2957, 2919, 2895, 2367, 3260, 2878, 3138, 2444, 1183, 2628, 2295, 2219, 2049, 2865, 2254, 1615, 508, 1754, 2761, 1722, 1516, 1775, 3763, 3885, 2091, 3226, 1812, 1214, 2201, 755, 3159, 1531, 1109, 1752, 250, 2846, 3330, 3460, 1464, 3462, 3110, 3042, 852, 2160, 3524, 3904, 463, 233, 3350, 3239, 383, 1733, 3907, 1893, 2513, 91, 1765, 1984, 1904, 3246, 2747, 3247, 2981, 3715, 2275, 3182, 1111, 1764, 3184, 485, 357, 3725, 1713, 2273, 2758, 3189, 3250, 2531, 44, 1013, 1200]. Recognize [3157]. Recognizing [1353, 26, 106, 3645]. Recombinant [1550]. Recombination [1384]. Recommendation [668, 1790, 2585, 370, 3730, 2543, 3208, 743, 2912, 2901, 3280, 2568, 1542, 1511, 3628]. Recommendations [2257, 3527, 997, 2607]. Recommender [1118, 2298, 2141, 2515, 2311, 1871, 2265, 3235, 114, 943]. Reconciliation [1159, 2843, 419, 1479]. Reconfigurable [893, 3525, 3476, 898, 1424]. Reconfiguration [698, 551, 583, 671]. Reconstruction [2395, 1138, 220, 3045, 15, 1794, 2143, 3154, 1563, 1035, 2920, 3683, 3836, 167, 1008, 2353, 3175, 2287, 918, 1906, 606, 870, 3915, 608, 1408]. Record [3402]. Record-Matching [3402]. Recording [3187]. Recordings [3021, 2885, 2081]. Records [2801]. Recovering [2333]. Recovery [204, 3504, 3887, 296, 413, 3011, 351, 870, 1522, 1408, 780]. Recreational [2608]. RecSys [2568]. Rectangular [2468, 129, 1820]. Recurrences [3401]. Recurrent [3356, 3086, 2820, 862, 1499, 3055, 3100, 2761, 1411, 2867, 829, 2314, 1218, 1927]. Recursive [372, 1077, 3872, 1684, 634, 88, 3661]. Recursively [831, 3339]. Recycling [3037, 2198]. Red [3052]. Reddit [2507]. Redesign [431]. Redesigning [2174]. Redistribution [886]. Redox [3649]. Reduce [2084, 1674, 740]. Reduced [3196, 1232, 2730, 1921, 3111, 223]. Reduced-Diameter [3111]. Reduced-Order [1232, 1921]. Reducing [2326, 693, 2693, 1857, 2469, 2798, 679]. Reduct [1309]. Reduction [3195, 2953, 1090, 1625, 3840, 3445, 2313, 136, 1134, 285, 385, 97, 2324, 2886, 1965, 1166, 3406, 1120, 180, 504, 1224, 1226, 2070, 487, 2300, 2179, 2611, 3618, 899]. Redundancy [3043]. Reeb [1185]. Reed [1130]. Refactoring [1540]. Reference [1908, 981, 1143, 2965, 1161, 410, 460, 2800, 637, 865]. Refined [1052]. Refinement [1557, 73, 1844, 2914, 872]. Refining [1635]. Reflected [2282]. Reflecting [1586]. Reflectometry [648, 2343]. Refraction [1793]. Refrigerants [2064]. Refrigeration [2032]. Refueling [1127, 1821].

Regardless [3002]. Regeneration [3214]. Regenerative [543, 2345]. Regime [1528, 1863, 1731]. Regime-Changing [1528]. Regime-Switching [1863]. Regimes [732]. Region [66, 1180, 2367, 3394, 3207, 2167, 1687, 3508, 2593, 2502, 1402, 2392, 260, 2575]. Region-Based [1402]. Region-Convolutional [1687]. Regional [3075, 2141, 628, 3775, 1356, 733]. Regionalized [1181]. Regions [848, 2510, 34, 1768, 1799, 911, 1355]. Registers [3233]. Registration [1662, 3203, 64, 843, 1985]. Regression [1935, 2732, 3262, 1519, 1425, 1234, 3421, 2881, 781, 3289, 2247, 141, 2527, 2309, 3794, 2992, 2998, 1324, 3624, 3893, 2804, 3676, 2498, 2601, 2768, 3004, 1620, 2182, 2530, 3710, 1079, 1203]. Regret [3538]. Regret-Enhanced [3538]. Regrinding [1798]. Regular [1297, 1381, 720]. Regularization [1160, 3439, 944, 1663, 701, 413, 3813, 2653, 1926, 1079]. Regularizations [751]. Regularized [152, 3707, 2529, 2810, 1391, 1918, 870, 2514]. Regulating [3605]. Regulation [803, 3514, 798, 527]. Regulatory [414]. Rehabilitation [3003, 3070, 2608]. ReID [1385]. Reinforced [967, 1616, 1987]. Reinforcement [2614, 1677, 2406, 2926, 3133, 596, 2586, 2495, 2467, 2363, 1628, 3205, 2977, 2671, 3750, 3398, 3504, 3147, 2334, 3150, 1925, 1488, 3880, 3057, 2491, 1173, 2694, 3886, 1060, 2437, 3000, 2021, 2445, 2310, 3110, 2755, 1685, 1218, 3112, 2115, 1695, 2151, 665, 3242, 2744, 2478, 2408, 2106, 1854, 3422, 1738, 2717, 3788, 1286, 2079, 2487, 2971, 1638, 1712, 3679]. Reinforcing [2480]. Reinforcing-Learning-Driven [2480]. Rejection [1143, 888, 1623, 1762, 3818, 3060]. Rejuvenation [678]. Relate [1520]. Related [2306, 3613, 2968, 2439, 1583, 1183, 2795, 829, 2131, 2728, 2152, 658, 3717]. Relation [442, 3203, 773, 1973, 2185, 961]. Relational [3288, 2611]. Relations [2396, 158, 3258, 703]. Relationship [1110, 2712]. Relationships [820, 656]. Relative [1311]. Relative-Positioning [1311]. Relatively [360]. Relaxation [3097]. Relaxed [1153, 1334, 592]. Relay [1403, 319]. Release [2290, 3005]. Relevance [3166, 3487]. Relevant [35, 3833, 3917]. Reliability [2583, 3729, 471, 3613, 3095, 1267, 1632, 3664, 1477, 1569, 2103, 1195, 3483, 2971, 3487]. Reliability-Based [1569]. Reliable [3605, 3689, 443, 3414, 1875, 1129]. Relief [516]. Relocation [3189]. Relying [1670]. Remaining [1702, 3555]. Remanufacturing [3217]. Remeshing [3114]. Remodel [2084]. Remote [3491, 252, 3346, 540, 2839, 507, 2808, 1532, 2562, 2566, 3789, 1130]. Remotely [20]. Removal [3614, 3720, 751, 314]. Removals [1110]. Remove [506]. Rendezvous [1645, 1436]. Renewable [693, 2103, 2024, 1538, 3905, 3538, 453, 1987]. Renovation [3047]. Rental [1692, 2737]. RePair [1316]. Repeat [2047]. Repeatable [106]. Repeated [3777]. Repeats [1045]. Repetitive [764, 3008]. Replacement [3711]. Replacing [1136]. Replay [3512]. Replenishment [3479]. Reporting [1653, 272]. Reports [142]. Representation [1609, 2042, 442, 3496, 1373, 1239, 1954, 181, 1584, 2430, 371, 822, 835, 1476, 1030, 1464, 1220, 1536, 281, 657, 875, 918, 1598, 2452, 756, 806, 864, 1026, 2273, 3249, 384, 1693]. Representation-Focused [1239]. Representations [2584, 2672, 2342, 1848, 674, 2241, 1199]. Representative [1317, 3126]. Represented [1928]. Representing [1290, 2577]. Reproducibility [3886]. Reproducible [1873]. Reproduction [2632]. Reputation [1632]. Reputation-Driven [1632]. Requirements [2486, 1530, 709, 114]. Resampling [3289, 2286, 2779, 1599]. Rescheduling [2799, 172, 1247]. Rescue [2341, 3087, 1569, 594, 516, 2854]. Research [2213, 3435, 3751, 3855, 2587, 2907, 3564, 1027, 1837, 779, 1868, 2412, 1692, 648, 2590, 1516, 1775, 2808, 3556, 3821, 3887, 2318, 1879, 2844, 1532, 2727, 2191, 1727, 3786, 676, 3070, 3742, 1738, 2736, 3400, 3655, 3811, 3484, 613, 467, 1356, 806, 2822, 3543, 3670, 552, 2082, 3251]. Researchers [2698]. Resemblance [3259]. Reservation [1184]. Reserve [1105]. Reservoir [3081, 567, 721]. Reservoir-Generation [567]. Residential [3749, 1172, 3426, 3752]. Residual [3086, 2590, 3821, 694, 3021, 2220, 2452, 3353, 1111]. Residual/Inception [3021]. Residual/Inception-Based [3021]. Residual/Recurrent [3086]. Resilience [3887, 2248]. Resilient [3864, 3043, 1966, 2248]. Resistance [3693, 3281, 2117]. Resistant [3732, 863]. Resistivity [1086]. ResNet [3304, 3146]. ResNet18 [2981]. Resolution [791, 1234, 199, 3706, 3785, 3157, 694, 2749, 1657, 1071, 1532, 3114, 2567, 3789, 2069, 1484, 1839, 3001]. Resolution-Preserved [2567]. Resolve [3437]. Resolved [21, 38]. Resonance [57, 1062, 2313, 33]. Resonances [1834, 2453]. Resource [3566, 1991, 3741, 734, 3132, 386, 787, 1328, 3027, 781, 1586, 2809, 3880, 3155, 3454, 3008, 130, 352, 1569, 2058, 1298, 3607, 744, 2887, 2024, 998, 2837, 3577, 949, 1196, 2610, 232, 474, 545, 2702]. Resource-Bounded [1991]. Resource-Constrained [3132, 386, 3008, 2610]. Resources [440, 1998, 3264, 172, 130, 2952, 1987]. Respect [1261, 1399, 2661]. Respiratory [2918]. Responding [1778]. Response [3412, 3699, 3822, 2172, 2086, 3698, 3143, 77, 2810, 925, 752, 3126]. Response-Aided [3126]. Responses [3822, 731, 3181]. Responsive [1172, 2490, 2597]. Rest [1362]. Resting [1609, 1387]. Resting-State [1609, 1387]. Restoration [944, 1780, 918]. Restraint [2555]. Restraint-Free [2555]. Restricted [3139, 615, 2056, 1189, 2498]. Resulting [1703, 1651]. Results [2984, 1583, 2420, 3128, 102, 3381, 3278, 2807, 946, 2563, 2532, 2267, 1471]. Retail [3176, 2479]. Retailing [1593]. Rethinking [3360]. Retinal [619, 2972]. Retinopathy [2821, 2230, 758]. Retraining [2446, 2209]. Retrial [3276]. Retrials [3139]. Retrieval [3429, 1177, 1233, 1185, 428, 35, 1629, 3581, 3230, 2703, 2562, 3407, 1608, 646, 961, 2327, 3919, 53, 2708]. Retrievals [2384]. Retrieved [2549]. Returns [1859]. Reuse [1715]. Reveals [2202]. Revenue [1746]. Reversal [715]. Reverse [3551, 75, 1507, 3917, 3863]. Reversed [1409]. Reversible [106]. Review [1675, 3566, 3686, 2406, 1759, 3681, 3132, 1942, 3198, 3286, 3697, 3275, 3199, 2439, 3408, 3336, 3497, 3201, 1400, 37, 1103, 1846, 1686, 2814, 342, 3442, 2005, 27, 3055, 2378, 1128, 3019, 1438, 3808, 182, 3506, 3671, 832, 2051, 74, 3282, 2218, 3311, 2880, 3105, 3000, 6, 3604, 2733, 1822, 3039, 2735, 2516, 2294, 3110, 2319, 3606, 3695, 2250, 2117, 2066, 2092, 3687, 2918, 1539, 2178, 68, 2464, 3909, 3403, 2191, 362, 2728, 1707, 1005, 2760, 2946, 3124, 1679, 3065, 474, 3185, 1809, 3544, 1728, 2176, 1882, 3570, 3863]. Reviewers [972, 208, 246, 334, 417, 542, 765, 1648, 2146]. Reviews [3129, 1144, 2716]. Revised [444]. Revisit [1931]. Revisited [2588, 1345]. Revisiting [1787, 645]. Revolution [2615]. Reward [521, 1685, 1627]. Rewards [2316, 3788]. Reweighted [300]. Rewrite [209]. Rewriting [1397]. Reynolds [1141]. RF [1462, 2672, 1837, 3178]. RFI [71]. RFSVM [3088]. RGB [3646, 3645]. RGB-D [3646]. RGloVe [442]. RHP [846]. Rhythm [3605, 862]. Ribiére [664]. Ribo [1873]. Ribo-Seq [1873]. Rice [1123]. Rich [1080]. Ride [783, 1634]. Ride-Sharing [1634]. Ridesharing [2666]. Riemann [1126, 3647, 1682]. Riesz [1673]. Right [398]. Rights [2006]. Rigid [412, 899]. Ring [2305, 1083, 3414]. Ringing [2179]. Ripple [686, 2797]. RIS [2702]. Risk [2663, 3196, 2585, 521, 1105, 2249, 1825, 3145, 1410, 3722, 3161, 3573, 2471, 3459, 1254, 730, 3470, 1057, 1915, 591, 3533, 2408, 2835, 3775, 685, 3361]. Risk-Sensitive [2408]. Risks [3133]. Risky [1016]. RL [2365]. RL1 [300]. RMFRASL [2097]. RMSprop [3603]. RMSprop-Incorporated [3603]. RNA [3857, 1625, 21, 118]. RNAs [3857]. RNN [1699, 1684]. RNN-Based [1699]. RNNs [3055]. RO [1219]. Road [3379, 639, 2340, 1267, 1767, 520, 1023, 1821, 2535, 1338, 3631, 3552, 2686, 1026, 553]. Road-Marking [3631]. Roadmap [1536]. Roads [991]. Robot [1442, 478, 2495, 1868, 30, 696, 2390, 2595, 3303, 1035, 1961, 3517, 626, 538, 1785, 1521, 810, 3305, 575, 1547, 2082, 2576, 2770]. Robotic [3132, 3275, 2476, 3094, 2255, 404, 3112, 1451]. Robotics [2660]. Robots [1313, 2042, 1444, 3747, 1802, 1307, 1643, 3270]. Robust [1662, 3846, 888, 2410, 632, 1944, 386, 1233, 1514, 727, 1124, 1950, 2520, 1730, 2819, 3787, 3097, 3783, 967, 2875, 2097, 764, 285, 557, 1571, 1116, 1216, 801, 2492, 2312, 3042, 647, 2603, 2293, 993, 3677, 911, 2532, 3476, 1361, 1192, 480, 3317, 3370, 85, 657, 857, 1598, 1904, 3180, 359, 3917, 1512, 1870, 3361, 377, 752, 3767, 3739, 3252]. Robustness [2812, 2689, 2629, 2819, 3884, 1534, 2405, 2142, 2233, 402]. ROC [3196, 2336]. Rock [2125, 3320, 3109]. Role [2938, 867, 863, 1518, 2735, 3376]. Roll [830]. Roll-to-Roll [830]. Rolled [3609]. Rolling [928, 2759, 701, 3310, 1969, 782, 3487, 1331]. Rolls [3677]. Roman [3227, 3648]. Romanian [1968, 2059]. Romberg [2517]. RONO [1099]. Rooms [3747]. Root [1020, 3133, 3263, 323, 2759, 2239]. Root-Finding [1020, 323]. Root-Findings [2239]. Rooted [179]. RooTri [2492]. Roots [3820, 333, 290, 3834]. Rose [1783]. Rostering [173]. RoSummary [2059]. Rotary [1297]. Rotated [3273]. Rotating [3096, 620, 3234, 2387]. Rotation [412, 1916]. Rotatory [3860]. Rotorcraft [1426]. Rotors [342]. Rotoscope [2987]. Rough [590, 1021, 1006, 1545]. Round [1923, 2325, 98]. Round-Off [1923, 2325]. Roundabout [2247]. Roundup [2003]. Route [3495, 2664, 1468, 2345, 3907, 1727, 1906]. Routes [2456]. Routing [562, 726, 3084, 600, 500, 3087, 1556, 75, 2093, 1299, 896, 2419, 3417, 2368, 19, 182, 2008, 1157, 1287, 1187, 1767, 650, 2056, 3159, 1171, 1015, 902, 18, 161, 855, 1152, 1507, 1821, 2464, 2316, 3479, 1038, 605, 1338, 3528, 1094, 1296, 3053, 2130, 3788, 1197, 1447, 873, 762, 789, 3137]. Row [724]. RRR [538]. rs [1799]. rs-fMRI [1799]. RSA [3471]. RSB [1950]. RSSI [3922]. RSU [491]. RT3102 [288]. RTN [864]. RTT [2044]. RU [3648]. RU-OLD [3648]. Rubric [3848]. RUemo [2152]. RUL [2978]. Rule [283, 1578, 484, 3838, 472, 2435, 199, 3557, 1791, 1568, 1153, 1334, 1227, 1461]. Rule-Based [3838, 199, 3557, 1791, 1153, 1334, 1461]. Rules [3401, 1317, 3051, 2724, 2368, 356, 1250]. Rumor [2856]. Run [868, 3257, 577, 2169]. Run-Length [868]. Runge [2780, 3736, 754, 2634, 3221, 3896]. Runners [2608]. Running [3229]. Runtime [1991, 2930]. Rural [2834]. Russia [2152]. Russian [1552, 2096]. Russian-Language [2096]. RVFL [580].

S [1661, 235, 3593, 2240, 2101]. S-Boxes [2240]. S-EPSO [3593]. S-Parameters [1661]. S4 [2548]. SA [3317]. SA-Net [3317]. SaDE [613]. SaDE-ELM [613]. Safe [1790, 3689, 1097, 1299, 771, 575]. Safeguarding [3895]. Safety [3759, 3279, 1730, 3224, 2483, 2859, 2080, 1436]. Safety-Critical [3279]. SAFFRON [1993]. Saharan [3403]. SAIN [3858]. Sales [3495, 3682]. Salesman [2485, 2127, 695, 1240, 164, 2640, 1515, 2123, 948, 1095, 1371, 217]. Salesmen [2120, 2174]. Salient [1068, 2766]. Saline [3081]. Salinity [3491]. SalNet [1068]. Salp [1219]. Salt [751]. SAM [3444]. Same [1455]. Sample [2286, 3373, 3625, 2382, 904]. Sampled [270, 477]. Sampled-Data [477]. Sampler [3266]. Samples [1142, 3816, 2125, 2573]. Sampling [2840, 3711, 1889, 412, 1279, 1769, 2768, 577, 628, 3728]. SAND [1819]. SAND-Mask [1819]. SAPBERT [2091]. SAPEVO [1388]. SAR [249, 3771]. Sarcophagus [2287]. SARIMA [3419, 530]. SARIMA-Based [530]. SARS [3665, 2632, 1598]. SARS-CoV-2 [3665, 2632, 1598]. SAT [1901, 1823, 1251, 3802, 2069]. SAT-Based [3802]. Satellite [390, 127, 64, 199, 3784, 2062, 2564, 2197, 937, 1137, 895]. Satellites [3670]. Satisfaction [1941, 1917]. Satisfiability [1272]. Satisfy [1236]. Sato [869]. Saturation [3054, 2247]. Savings [3668]. Say [3324]. Say-Computing [3324]. SBM [3770]. SBWO [3818]. SCA [2210]. Scaffolding [2431]. Scalability [3103, 3532, 3545]. Scalable [2199, 491, 931, 1266, 2592, 3454, 3274, 990, 1414, 3715, 1141]. Scalar [1126, 2492]. Scalarized [3873]. Scale [3714, 1823, 734, 2481, 2821, 3781, 330, 287, 941, 1619, 1661, 1560, 439, 2527, 1249, 2015, 3513, 1003, 3294, 3416, 2804, 1426, 358, 1396, 3350, 2299, 1749, 1475, 2507, 504, 1914, 3733, 1163, 3326, 1394, 2075, 643, 2736, 474, 3776, 1546, 1750, 3685, 1986, 239, 38]. Scale-Free [1823]. Scaled [1661, 1591, 3314]. Scaling [127, 2000, 1607, 781, 1134, 2642, 1487]. Scaling-Topic [2000]. Scan [1706]. Scanned [2827]. Scanner [1985]. Scans [3665, 2518, 2818, 91]. Scarce [2765]. Scatter [3028, 2629, 978]. Scattered [3079]. Scatterers [3423]. Scattering [29]. Scenario [2341, 2260, 2541, 3625, 3293, 3918]. Scenario-Based [2260]. Scenarios [3139, 2916, 2633, 3460, 1359, 114, 1280, 1711, 2275]. Scene [2846, 2513]. Scenes [1531, 2790, 2696, 2933]. Schedule [2940, 2314, 2478, 2417]. Scheduling [1017, 3028, 3566, 1325, 1360, 3434, 512, 1945, 3637, 386, 491, 2586, 1997, 225, 970, 2542, 3673, 3267, 607, 3698, 631, 1949, 2489, 567, 574, 2799, 3078, 1351, 586, 2048, 1080, 3389, 1458, 3592, 3410, 2290, 1561, 210, 844, 581, 3008, 2784, 886, 987, 3332, 547, 1569, 3625, 248, 2437, 2098, 2956, 3288, 3265, 3000, 2123, 2348, 2255, 2887, 3167, 2024, 3708, 2025, 1920, 855, 2837, 2027, 1845, 2124, 1448, 1459, 1054, 81, 588, 597, 947, 1494, 2183, 1979, 3005, 394, 1196, 2480, 611, 2076, 1660, 618, 952, 2351, 1393, 2802, 1356, 3918, 513, 2612, 2854, 3670, 724, 549, 864, 2307]. Scheduling [1510, 1987, 3189, 3252]. Schedulings [818]. Schema [1412, 1310]. Schemas [1489]. Scheme [2780, 1143, 388, 14, 127, 2921, 861, 812, 420, 969, 532, 706, 2015, 547, 1041, 3163, 3320, 1883, 3257, 3418, 1361, 774, 606, 2078, 3616, 231, 2743, 1424]. Schemes [3668, 2411, 2688, 3692, 2020, 2063, 2327, 479]. Schizophrenia [1387]. School [295, 3299, 2598, 1761, 2374]. Schools [1027]. Science [3425, 676, 2969]. Sciences [517]. Scientific [1049, 3523, 1402, 1546, 3743]. SCIM [1816]. Scintigraphic [2518]. SCMs [2959]. Scooter [2737]. Scoping [3506, 3065]. Scopus [3523]. Score [2279, 3166, 1765, 3126]. Score-Matching [3126]. Scoring [3470, 1508, 1753]. Scraper [828]. Screen [1181]. Screening [2784, 3584]. Screw [3835]. scRNA [3485]. scRNA-Seq [3485]. SDN [2280, 1186, 2316]. SDN-Based [2316]. SDPhound [50]. Se [1236]. Sea [2462, 1216, 3571, 3299]. Sea-Rail [3571]. Search [3028, 1990, 562, 3012, 3650, 1936, 2618, 1517, 1818, 363, 90, 426, 2172, 753, 761, 171, 2155, 2587, 1997, 374, 2970, 2683, 2433, 901, 420, 567, 630, 2664, 3795, 192, 3593, 2631, 978, 2175, 3505, 3417, 1272, 1187, 2272, 1385, 448, 219, 561, 558, 3106, 1613, 569, 2679, 444, 392, 475, 2117, 3526, 3764, 1448, 281, 3704, 1193, 1487, 2070, 1478, 349, 516, 2233, 2076, 1521, 326, 824, 3482, 216, 1473, 197, 1492, 1627, 2991, 3549, 3137, 3290, 3858]. Search-And-INfer [3858]. Search-Based [2433]. Search-Exponential [2683]. Searchable [452]. Searching [1646, 1614, 1204, 2769]. Seasonal [2331, 3683]. Seat [1332]. Secant [3820]. Secant-Type [3820]. Second [3254, 1201, 717, 2053, 603, 493]. Second-Order [3254, 1201, 603]. Secondary [3857, 611, 118]. Secrecy [793]. Secret [3787]. Section [819]. Sections [3151, 833]. Sector [2234, 2294]. Secure [3849, 3658, 3653, 3766, 523, 452, 3666]. Security [2947, 3430, 2195, 51, 3085, 170, 447, 3877, 3311, 2826, 1369, 1738, 1832, 3918, 3545]. Seed [1691]. Seed-Guided [1691]. Seeding [534, 1810, 2641]. Seeker [686]. Seeking [2775]. Segment [3444, 1245, 304, 97]. Segment-Based [1245]. Segmentation [3491, 2864, 2617, 1160, 140, 1911, 2821, 2342, 3134, 311, 2624, 1180, 1068, 2825, 2625, 2787, 1509, 1880, 2785, 2924, 3761, 2162, 3869, 3444, 1392, 2593, 2667, 2808, 1003, 3294, 1140, 1912, 3062, 1804, 2846, 2559, 3323, 1769, 2528, 3803, 3407, 3238, 3684, 3477, 3909, 1460, 1407, 2914, 3529, 188, 2987, 262, 2479, 249, 3635, 3542, 2860]. Segmentation-Based [3062]. Segmented [1935]. SEIR [2501, 1342]. Seismic [1526, 1736, 1535, 465, 3285]. Seizure [3744, 920, 3329]. Seizures [3150]. Select [240, 2569]. Selectable [802]. Selected [2002, 2496, 3580, 1601, 2722]. Selecting [2931]. Selection [3003, 3012, 1073, 2302, 3347, 3261, 544, 1384, 514, 2111, 26, 2347, 975, 3567, 2724, 1628, 2394, 443, 2411, 2999, 1625, 3445, 3067, 1956, 790, 2148, 969, 2451, 3268, 1563, 2594, 2097, 706, 1368, 755, 1324, 3456, 3824, 3459, 3748, 2935, 1760, 1348, 770, 1279, 1359, 1828, 2775, 3902, 2250, 3904, 3474, 1149, 594, 1508, 3370, 2513, 1478, 1597, 3532, 1598, 505, 1259, 365, 2530, 1500, 1544, 3920, 3562, 319, 300, 3126]. Selection-Based [2513]. Selective [1703, 1186, 279, 3821, 3294, 3622, 748, 792]. Selectivity [3854]. Self [1493, 1110, 924, 24, 1068, 885, 631, 1183, 2628, 2924, 30, 2666, 602, 2254, 2590, 1444, 1378, 3782, 122, 2978, 3828, 670, 808, 3457, 1288, 2282, 2190, 2908, 609, 1219, 2343, 1851, 256, 1322, 3404, 966, 1487, 3631, 589, 2075, 425, 317, 3842]. Self-Acceleration [317]. Self-Adaptive [885, 631, 2666, 602, 2590, 1444, 256, 2075]. Self-Admitted [1110]. Self-Affine [122]. Self-Attention [3828, 3457, 1851]. Self-Attention-Based [2978]. Self-Attentive [2254]. Self-Autonomous [3782]. Self-Aware [1068]. Self-Awareness [966]. Self-Configuring [1378]. Self-Distillation [2628]. Self-Labeled [808]. Self-Learning [1493, 609, 1219]. Self-Nested [924]. Self-Organising [1322]. Self-Organization [30, 24]. Self-Organizing [1183]. Self-Scaling [1487]. Self-Similarity [2282]. Self-Stabilizing [589]. Self-Supervised [3404, 3631, 3842]. Self-Supervision [1288]. Self-Sustainability [2190]. Self-Sweeping [2343]. Self-Training [670]. Self-Triggered [2908]. Self-Tuning [425]. Selfish [75, 2816]. Semantic [3759, 2281, 3869, 1489, 3873, 2808, 1102, 1769, 469, 2930, 3174, 114, 1407, 1137, 1513, 3635, 3765, 3249]. Semantics [2268, 3799, 1530, 441]. Semantics-Based [1530]. Semi [2085, 930, 3804, 1811, 1874, 312, 2924, 3209, 371, 1612, 3557, 1506, 35, 3881, 543, 3831, 130, 1740, 1847, 2209, 2270, 670, 3352, 1149, 2748]. Semi- [3209]. Semi-Active [543, 3352]. Semi-Automatic [1847]. Semi-empirical [35]. Semi-Implicit [1874]. Semi-Local [2085]. Semi-Markov [1612]. Semi-Preemptive [130]. Semi-Self-Supervised [2924]. Semi-Smooth [3831]. Semi-Supervised [930, 3804, 1811, 312, 371, 1506, 3881, 1740, 2209, 2270, 670, 1149, 2748]. Semi-Supervised-Learning-Aided [3557]. Semiconductor [621, 2701]. Seminal [227]. SENSE [1530]. Sensed [252, 1071]. Sensing [3491, 1164, 3861, 540, 489, 2839, 306, 3883, 507, 2808, 413, 582, 1532, 6, 2562, 2566, 3789, 389, 870, 3353, 1130, 348]. Sensitive [355, 1817, 648, 1743, 889, 2603, 2408, 3640, 961]. Sensitivities [1032]. Sensitivity [490, 400, 3498, 713, 1104, 2562, 2648, 3669]. Sensor [66, 1040, 1418, 2927, 3286, 25, 3864, 235, 75, 473, 812, 1182, 1955, 234, 30, 238, 1635, 78, 19, 210, 2094, 3215, 306, 226, 1365, 3823, 1015, 18, 161, 3274, 647, 3634, 2266, 378, 40, 2300, 286, 1197, 1447, 1131, 8]. Sensor-Based [3286, 30, 378]. Sensorized [2692]. Sensors [2583, 26, 3201, 3040, 27, 3451, 3022, 2880, 2788, 2882, 3239, 2523, 2475]. Sensors-Based [2880]. Sensory [2285, 3779]. Sentence [674]. SentenceLDA [1763]. SentenceLDA- [1763]. Sentences [3296, 2225]. Sentential [2699]. Sentiment [2401, 3129, 440, 3729, 439, 3298, 1968, 2716, 474, 364, 2176, 1461]. Sentiments [3168]. Separable [755, 2572]. Separate [3114]. Separated [376, 2337]. Separation [1709, 1413, 270, 255]. Separator [95]. SepFree [1896]. Septic [2159]. Seq [1873, 3485]. Sequence [3859, 176, 225, 862, 126, 2805, 3597, 212, 2862, 1543, 2275, 1310]. Sequence-Based [2275]. Sequences [2966, 1042, 1558, 194, 1239, 209, 3711, 1587, 1847, 645, 451, 2862, 2758]. Sequencing [1093, 1625, 3093, 3879, 1711]. Sequential [2042, 2543, 2961, 3269, 846, 1065, 1896, 2068, 2300, 1511, 3186, 1603, 3679]. Series [2780, 2321, 654, 1404, 3851, 2990, 1261, 414, 2506, 3829, 3749, 1948, 243, 3092, 3866, 1829, 2718, 3868, 3832, 1087, 3575, 1955, 2631, 1207, 1584, 2289, 704, 2331, 429, 1000, 1962, 1063, 3748, 314, 2733, 1440, 2601, 2932, 180, 1851, 3769, 3526, 399, 3049, 2771, 3602, 2202, 1037, 3910, 368, 505, 2989, 2338, 3682, 3314, 307, 1230]. Series-Parallel [3575]. Server [3, 3276, 3482]. Serverless [1833, 2977]. Service [2537, 3351, 3013, 2756, 3359, 3139, 3027, 1480, 130, 2471, 1864, 1856, 3591, 1638, 1643]. Service-Oriented [1856]. Services [1991, 2423, 896, 1186, 699, 2737, 1729, 1352, 407, 319]. Serving [1048]. Servo [596, 2671, 686]. SEsM [3654]. Set [2404, 1455, 152, 1327, 975, 2145, 1805, 2262, 978, 2915, 1438, 3155, 1640, 1252, 541, 3459, 162, 1420, 2956, 2018, 1921, 1848, 1401, 2299, 2138, 1006, 2327, 3539, 712, 197]. Set-Based [2145]. Set-Point [2138]. Set-Theoretic [1401]. Sets [2302, 499, 1418, 3035, 484, 659, 908, 101, 1134, 3364, 835, 2551, 1694, 429, 1282, 3021, 2432, 2399, 560, 1844, 2931]. Sets-On [2302]. Setting [2982, 579]. Settings [956, 3785, 2350]. Settlements [3668, 374]. Setup [1430, 2604, 599]. Sevcik [489]. Seventh [3224]. Seventh-Degree [3224]. Several [3104, 3654]. Severe [2938, 1222]. Severity [3847, 2941, 548, 1799]. Sex [3508]. Sexually [3403]. SFCs [1174]. SFR [352]. SGP [3652]. Shack [2420]. Shaded [2459]. Shading [1942, 606]. Shadow [3720]. Shadowed [761, 2362]. Shadows [3656]. Shadows-A [3656]. Shallow [1526, 2550, 3049, 2771, 3725]. Shannon [92]. SHAP [3705, 3121, 3813]. SHAP-Based [3705]. SHAP-Guided [3813]. Shape [1339, 915, 1261, 2107, 976, 967, 2429, 1903, 2150, 1572, 2498, 3798, 606, 3791, 1650, 3377, 260]. Shape-Aware [3377]. Shape-from-Shading [606]. Shape-Restricted [2498]. Shaped [1665, 397, 462]. Shapes [3077, 2682, 3521, 2244]. Shapley [2809, 2807, 3716]. SHAPs [2035]. Sharding [1632]. Shared [2111, 2495, 258, 2737, 2038]. Sharing [1002, 3650, 2434, 3041, 1351, 3787, 3783, 706, 135, 1634, 1298]. Shark [2965]. Shear [1737]. Shearlet [755]. Shelf [43]. Shells [1373, 3211, 3343]. Shelved [2549]. Shelved-Retrieved [2549]. Shielding [2377]. Shift [2659, 3022, 2124, 3914]. Shifting [2970]. Shifts [1731]. Ship [661, 3673, 1606, 375, 2062, 3530, 2654, 2303, 2854]. Shipborne [1012]. Shipbuilding [1450]. Shipping [2462]. Shocks [1661]. Shooters [2406]. Shooting [539]. Shop [1017, 3028, 1360, 512, 1945, 225, 1949, 2489, 574, 2799, 3389, 1458, 1561, 844, 1658, 987, 248, 2098, 2348, 3167, 1845, 947, 1393, 599, 724, 3252]. Shops [3592]. Short [1002, 3699, 3436, 2260, 2466, 3749, 1074, 2462, 713, 1951, 1829, 1428, 141, 1362, 403, 917, 3891, 2813, 1216, 2597, 1866, 577, 1541, 588, 3121, 3535, 2275, 2936, 3188, 2081, 2656, 1603, 3548]. Short-Lived [2466]. Short-Memory [2275]. Short-Run [577]. Short-Sea [1216]. Short-Term [1002, 3699, 3436, 2260, 3749, 1074, 713, 1951, 1829, 1428, 141, 403, 917, 3891, 2813, 588, 3121, 3188, 2656, 3548]. Shortages [1600]. Shortest [1158, 1168, 1133, 1281, 3098, 3077, 3812, 979, 1592, 1683, 2982, 3652, 2245, 1294]. Shortest-Path [1133]. Shortness [2841]. Shot [1004, 3143, 1751, 3581, 3819, 2935, 3912, 3771, 3674]. Shoulder [3684]. Show [3142]. Shrinkage [2590, 3821]. Shuffled [248, 346]. Shufflenet [1214]. Shunt [3649, 1855]. Shuttle [1303]. Si [3211]. Siamese [3819, 2288]. Side [2579, 2913, 2908]. Sided [207, 174, 3257]. siEDM [349]. Sieve [2948]. Sieving [2814, 1592]. SIFT [2966]. Sigma [52]. Sigma-Point [52]. Sign [2220, 3715, 377, 518]. Sign-Error [2220]. Signal [1062, 127, 393, 3342, 2424, 1205, 630, 1730, 3052, 204, 3097, 3874, 1047, 3453, 3511, 509, 621, 385, 3349, 2788, 2343, 2959, 463, 3475, 3177, 465, 2338, 286, 391, 870, 782, 780, 3570]. Signal-to-Noise [3349]. Signal/Background [3453]. Signals [2235, 3347, 2957, 1858, 3205, 3277, 2453, 1413, 648, 2892, 2909, 1567, 3885, 3327, 929, 1365, 1071, 2813, 2886, 154, 1162, 3004, 660, 644, 1984, 3833, 1698, 3353, 1230]. Signature [2680, 2610]. Signed [975, 2712, 2794]. Significance [3763]. Significance-Aware [3763]. Silica [3794]. Silicate [2090]. Silico [3854, 3104, 181]. Sim [2857, 2815]. Sim-Learnheuristic [2857]. Simhash [655]. Simhash-Based [655]. Simheuristic [1458, 1171]. Simheuristics [2633, 1283, 1644]. Similar [3146, 1473]. Similarity [393, 2281, 572, 2430, 1489, 369, 650, 773, 2282, 1432, 3563, 3402, 1224, 2801, 83, 1230]. Similarity-Driven [1224]. Simple [1818, 3401, 921, 2895, 1748, 1043, 535, 3654, 2492, 948, 3176, 2300, 2982, 186]. Simpler [1558, 214]. Simplex [2953, 3128]. Simplicial [3798]. Simplification [531, 1450, 3162]. Simplified [490, 2419, 3325, 3375, 2061, 515, 3549]. Simplify [1235]. Simply [736]. Simpson [283]. Simulated [1164, 2107, 2629, 3228, 1695, 1152, 1507, 827, 3481]. Simulating [2370, 3509, 3882]. Simulation [2695, 1040, 888, 22, 103, 3612, 2086, 1072, 243, 1714, 519, 2922, 1504, 2295, 1238, 1104, 3144, 2906, 1877, 2698, 2704, 3649, 721, 1293, 1891, 2491, 268, 771, 1216, 2733, 3229, 730, 3109, 2643, 358, 2274, 1292, 3472, 1777, 842, 2532, 1361, 1280, 3825, 3528, 1771, 3530, 1478, 516, 368, 1296, 3715, 3305]. Simulation-Based [1238, 1293, 2274, 3715]. Simulation-Optimization [1292]. Simulations [1553, 2292, 2207, 3150, 2053, 2778, 2644, 2139, 1414, 1402]. Simultaneous [2711, 3139, 2881, 3458, 902, 3902, 2186, 1608, 1544, 1988]. Simultaneously [1630]. Sin [3701]. Sin-Cos-bIAVOA [3701]. Sinc [1919]. Sine [1232]. Sinelshchikov [1019]. Single [1929, 3203, 2916, 3074, 2629, 3444, 2290, 2247, 2286, 1962, 3785, 694, 1041, 2019, 2598, 2974, 811, 597, 1981, 2848, 611, 533, 1688, 549, 2081]. Single- [2629]. Single-Block [1041]. Single-Cell [1962]. Single-Channel [2848]. Single-Cluster [2916]. Single-Framework [2598]. Single-Lane [2247]. Single-Lead [2081]. Single-Machine [2290]. Single-Objective [3074]. Single-Phase [2974]. Single-Sample [2286]. Single-Stage [3203]. Single-Step [2019]. Single-Valued [533]. Singular [2689, 152, 2436, 3052, 1612, 546, 2122, 2677, 368]. Sink [3872]. Sinkhorn [1910, 1391]. SIRC [1225]. Site [3261, 34, 2398, 1059, 3319]. Sites [41, 1543]. Siting [3731, 2308]. Situationally [3640]. Six [2634]. Sixth [2085, 1212, 445]. Sixth-Order [1212, 445]. Sixty [3128]. Size [922, 1874, 3518, 891, 1715, 3531, 1296, 904, 104]. Sizing [3363, 1080, 2308]. Skateboard [3251]. Skeletal [1239, 1260]. Skeleton [852, 3246, 2273]. Skeleton-Based [3246, 2273]. Skeptical [1713]. Sketch [535]. Skew [749]. Skid [1762]. Skill [595]. Skills [3806, 3898, 421]. Skin [3010, 2005, 2547, 3307, 3231, 2060]. Skip [3578]. Skyband [910]. Skyline [743, 3281, 910]. Skylines [3393]. Slab [736, 3677]. SLAM [3646, 2157]. Sleep [3832, 2886, 3172, 2038]. Sleepers [2636]. Slicing [1186]. Slide [2911, 2754, 2033]. Sliding [632, 649, 3199, 1262, 717, 1246, 1562, 3100, 3518, 2312, 1656, 993, 3120, 117, 2976, 719, 2770, 729]. Sliding-Mode [2976, 719]. Slime [2542, 1961, 1916]. Slip [3414]. Slippage [1444]. Slippery [639, 991]. Slitherlink [133]. SLM [1703]. SLMS [300]. SLMS-RL1 [300]. SLoPCloud [688]. Sloppy [802]. Sludge [3223]. SMAC [3886]. Small [3605, 1086, 2950, 3610, 2107, 3279, 868, 483, 1664, 1265, 1640, 1252, 3625, 541, 1965, 1820, 904, 712, 3628]. Small-Bin [2950]. Small-Loop [1086]. Small-Sample [3625]. Small-Shape [2107]. Smallest [120]. Smart [1908, 2536, 3681, 1182, 234, 3653, 1490, 1564, 2371, 3282, 2164, 2198, 3384, 2730, 3165, 2450, 2561, 3815, 3418, 3577, 3474, 869, 759, 3908, 2568, 232, 3383, 1056, 3835, 1643, 3628, 2384, 1308, 1769]. SmartBuild [2568]. Smartphone [1904]. Smartphones [2659]. Smartwatches [3671]. SMC [3266]. SME [1979]. Smeaheia [3081]. Smears [1277]. Smell [2965, 2710]. SMILES [3799]. Smooth [76, 951, 3831, 3587, 3007, 2810, 857, 3827]. Smooth-Delayed [3827]. Smoothed [268]. Smoothing [322]. SMOTE [3289]. SMOTEENN [3289]. SMT [673]. SNE [2181]. Snow [2562]. SNP [2573]. SNR [965]. So2Sat [3899]. Sobolev [3626]. SOC [2195, 2224]. SoC-FPGA [2224]. SOC4AS [2195]. Social [2446, 820, 1744, 905, 2621, 1327, 3085, 1902, 1081, 3040, 1182, 919, 2482, 704, 1861, 1189, 1890, 979, 1125, 2124, 3477, 950, 2354, 2607, 408, 656, 415]. Social-Statistical [1902]. Societal [2152]. Socio [3593, 2602]. Socio-Demographic [2602]. Socio-Emotional [3593]. Socioeconomic [3722]. Soft [1493, 969, 2441, 3837, 1866, 1256, 244, 895]. Soft-Voting [1493, 969]. Softmax [2693]. SoftNet [1895]. Software [1575, 2947, 1386, 678, 3213, 735, 2098, 709, 3894, 1354, 2139, 1707, 1894, 3050]. Software-Defined [2947, 3213, 3050]. Soil [3491, 3855, 1115, 2386]. Soils [1526, 2550]. Solar [1404, 1942, 1633, 2683, 3152, 3890, 2807, 3265, 1876, 2459, 254, 2571, 2954, 2931]. Solar-Assisted [2571]. SOLiD [174, 1018, 2175, 1001]. Solid-Tank [2175]. Solids [11]. Solomon [1130]. Solomonoff [99, 100]. Solon [738]. Solow [3817]. Solution [3492, 2988, 113, 108, 261, 3367, 1299, 673, 688, 3872, 343, 2194, 2493, 2800, 2992, 3302, 481, 2442, 2255, 3815, 900, 1807, 1755, 1453, 325, 3118, 96, 1974, 3641, 1820, 2848, 1196]. Solutions [3642, 1201, 2042, 1019, 482, 3627, 3870, 2704, 3845, 1626, 2642, 3654, 2335, 843, 2115, 1078, 634, 3392, 133]. Solvable [1113]. Solve [3492, 124, 1936, 2455, 3643, 2669, 3561, 978, 2290, 581, 3837, 2560, 2031, 346, 1296, 47, 2707, 873]. Solved [3652]. Solver [1442, 395, 3836, 3243, 3743]. Solver-Interactive [3836]. Solvers [2757, 1901, 1874, 941, 196, 2069]. Solving [3254, 2780, 1126, 353, 328, 455, 423, 1575, 2116, 627, 3412, 3650, 1732, 1360, 454, 2811, 305, 324, 2252, 203, 512, 2993, 1944, 2127, 3428, 171, 322, 261, 1349, 3087, 2970, 3206, 2489, 409, 754, 586, 2355, 1458, 2960, 1272, 1084, 951, 2633, 532, 1157, 132, 844, 221, 1212, 290, 406, 1165, 2923, 2868, 3692, 3416, 2325, 1188, 1171, 3106, 2131, 3896, 2123, 3778, 2019, 1796, 1095, 277, 3170, 1594, 1537, 2119, 341, 2299, 1152, 1507, 1541, 1132, 2110, 173, 1978, 2147, 774, 2374, 1501, 3586, 1050, 2075, 1676, 299, 337, 445, 805, 488, 217, 3559]. Solving [317, 3549, 186]. Somatization [548]. Some [464, 29, 1994, 305, 283, 147, 3078, 2778, 1282, 1252, 2733, 946, 3471, 2349, 2110, 310, 1258]. SOMs [1051]. Sonar [777]. Sorghum [2680]. Sorting [92, 955, 1417, 3037, 2892, 1423, 2029, 2216, 2758]. Sound [3856, 268]. Sounds [1524]. Source [169, 2511, 3041, 1833, 1509, 2046, 1413, 648, 3044, 839, 2425, 3282, 2558, 2442, 757, 1466, 1540, 1454, 857, 2243, 155, 3918]. Source-Load [3918]. Sources [3437, 2589, 738, 2323, 2168, 384]. Sources-Based [384]. South [3236]. Soybeans [3202]. Space [2188, 2085, 1168, 2042, 629, 1999, 1234, 1343, 2851, 673, 1661, 2390, 380, 1265, 1290, 3617, 3622, 3349, 3103, 796, 1220, 3360, 399, 757, 292, 2267, 2606, 1976, 3626, 968, 1338, 367, 710, 646, 593, 249]. Space-Dependent [3617]. Space-Efficient [796, 1338]. Space-Time [3349]. Spacecraft [1633]. Spaced [3158]. Spaces [3809, 302, 2781, 2467, 1670, 229, 1889, 2997, 1683, 398, 1471, 2237]. SPAD [2673]. SPADE [3803]. Spain [1280, 3662]. Spam [2176, 3767]. Span [1428].

Spanning [3107, 3778, 584]. Spark [1014, 744, 3326, 273, 3035, 725, 1405, 1510]. Sparks [476, 2515]. Sparrow [2683, 3795, 3505]. Sparse [1910, 1034, 220, 2434, 1555, 3439, 3423, 313, 1183, 2718, 3292, 3870, 2688, 204, 196, 3215, 1780, 1452, 1116, 1030, 1685, 804, 1078, 1705, 2897, 281, 657, 875, 918, 1918, 870, 2653, 608, 1079, 377, 780, 384]. Sparse-Reward [1685]. Sparsity [2955, 1926]. Sparsity-Invariant [2955]. Spatial [3018, 1325, 3854, 1717, 379, 3505, 3222, 2637, 338, 2808, 2196, 3317, 3280, 2714, 2686, 2237, 3377, 573, 752]. Spatial-Aware [3317]. Spatial-Domain [3018]. Spatial-Temporal [752]. Spatially [1662, 1160, 2138]. Spatially-Varying [1662]. Spatio [1951, 3390, 3182]. Spatio-Temporal [1951, 3390, 3182]. Spatiotemporal [3136, 727, 2943, 982, 2566, 3635, 963]. SPD [2395]. Speaker [1892, 2236, 2091, 2180]. Speaker-Aware [2091]. Speaker-Independent [2236]. Speaking [3726]. Special [698, 1922, 2189, 2228, 2619, 734, 2118, 2410, 971, 215, 2104, 556, 1481, 813, 1398, 1386, 3267, 313, 387, 2002, 754, 3006, 1273, 183, 880, 2271, 3450, 2504, 1210, 2451, 3268, 2055, 1213, 2596, 212, 1170, 1967, 2600, 2912, 2274, 1082, 2132, 1716, 2065, 2028, 1970, 1701, 1795, 2068, 1257, 2662, 618, 952, 1601, 2036, 2351, 2722]. Specialisation [2324]. Specialist [3703]. Specialized [3877]. Species [3357, 56, 1361, 1984, 1543, 2836, 1380]. Species-Concentration [1361]. Specific [3697, 2000, 434, 3446, 125, 1272, 1047, 459, 2935, 2559, 2717]. Specification [2685]. Specificity [50]. Specificity-Determining [50]. Specifics [1781]. Specified [2412]. Spectator [3698]. Spectra [28, 67]. Spectral [1798, 2236, 1148, 3497, 3143, 258, 39, 1278, 2731, 3837, 3676, 2196, 3173, 3630, 3244, 2376, 2071, 1698, 3270]. Spectral-Spatial [2196]. Spectrogram [119]. Spectrograms [3360, 3738]. Spectrometry [44]. Spectroscopic [220, 1896]. Spectroscopy [1062, 1178, 314, 3380]. Spectrum [2184, 1937, 1770, 3286, 3861, 489, 1205, 2436, 3052, 1813, 1328, 234, 304, 706, 547, 558, 582, 2995, 546, 405, 2122, 1099, 368, 496, 228]. Spectrum-Adapted [1205]. Spectrum-Sharing [706]. Speech [2949, 3430, 2173, 2582, 2481, 2236, 1413, 823, 39, 3515, 1890, 3583, 3042, 2747, 1764, 3544, 3765, 1672, 1013]. Speed [3193, 788, 596, 1074, 1877, 690, 917, 3692, 686, 2345, 2674, 2210, 1078, 798, 3484, 1545, 3818, 3546]. Speeding [3812]. Speeds [1562, 1645, 3414, 3005]. Sperm [70]. Speyer [387]. SPH [2875]. Sphere [1119, 682]. SphereHarmonic [1885]. Spheres [1885, 1320]. Spherical [687, 1749]. Spider [415]. Spike [2892, 2813]. Spike-Weighted [2813]. Spiking [1702, 3875, 3100, 2813, 3231, 3526, 1107]. Spill [1012]. Spillover [1863]. Spindle [3556, 3821]. Spine [2791, 3467]. Spinous [2791]. Spiral [3058]. SPL [3680]. SPL-YOLOv8 [3680]. Spline [438, 3587, 1165, 104]. Splines [102, 1673, 104]. Split [3417, 250, 398, 3786, 3053, 2766]. Splitting [3692, 2763, 3053]. Spoke [2462]. Spoofing [2400]. Sports [3698]. Spot [1129, 3185, 453]. Spotted [2348]. Spread [3194, 2108, 2474, 1946, 1869, 304, 2070]. Spreading [1861]. Springer [74]. Springer-Verlag [74]. Sputum [188]. SPWM [2974]. SQL [3376]. Square [3612, 1374, 1715, 2232]. Square-Free [1374]. Squared [4]. Squares [113, 108, 372, 1077, 2851, 418, 1794, 2920, 2868, 3676, 2220, 3486]. Squeezing [1751]. Squirrel [824]. SR [1484]. SR-Inpaint [1484]. SS [2873]. SSA [1964]. SSVEP [2594]. SSVEP-Based [2594]. Stability [3492, 1641, 2039, 3197, 1938, 1232, 1349, 1947, 2246, 3736, 207, 3750, 3783, 967, 3221, 271, 288, 652, 2166, 200, 634, 525]. Stability-Constrained [3750]. Stabilization [3191, 632, 3870]. Stabilized [1877]. Stabilizing [589]. Stable [252, 1306, 178, 189, 205, 192, 2007, 177, 185, 168, 1909, 2442, 214, 2353, 1445, 3641, 435]. Stacked [1790, 1636, 1476, 652, 1200]. Stacked-GRU [652]. Stackelberg [1268, 768]. Stacking [3195, 3723]. Staff [1920, 2124]. Stage [1770, 600, 3203, 3693, 3255, 2827, 3619, 3321, 3830, 3455, 3416, 1845, 2798, 1194, 1707, 488, 402, 549]. Stages [2886]. Staging [3832]. Staining [3062]. Stakeholder [1465]. Stakeholders [3588]. Stakes [2172]. Standard [3848, 3197, 1263, 125, 913, 3116]. Standardization [2992, 129]. Standardized [3595]. Star [2774, 3034, 2997, 3721, 16]. Stars [3328]. Start [2265, 1676, 3368]. STATCOMs [2221, 2112, 1964]. State [1609, 2213, 2439, 3639, 1387, 746, 3773, 1582, 2388, 2851, 2396, 3398, 3373, 3278, 616, 2150, 3622, 3331, 193, 20, 1367, 2156, 399, 1336, 1766, 3242, 938, 996, 2513, 2484, 2074, 3124, 710, 2717, 2428, 357]. State-Masking [3398]. State-of-the-Art [193, 2484]. State-Space [710]. State-Specific [2717]. States [2140, 3099, 532, 1155, 3250, 1122]. Static [2444, 1076, 969, 539, 1988]. Static-Output-Feedback [539]. Station [3731, 1970, 986, 3541]. Stationarity [3092]. Stationary [3193, 3135, 1953, 2442, 2316, 1361, 1162]. Stations [3441, 580]. Statistical [3397, 449, 1902, 1074, 2281, 3727, 1248, 2869, 145, 3886, 3746, 253, 3272, 614, 2790, 1731, 3480, 2070, 286]. Statistics [3699, 1058]. Status [2670, 3346, 53]. Stay [735]. Steady [746, 2150, 3331, 996]. Steady-State [746, 2150, 3331, 996]. Steam [2181]. Steel [1450, 3712, 2590, 385, 2703, 358, 3529]. Steels [3368]. Steering [639, 991, 799, 801, 3655, 553]. Stefan [68]. Steffensen [264, 3834]. Steffensen-Genetic [3834]. Steffensen-Type [264]. Steganography [3658, 2985]. Stem [1036, 3470]. Stenosis [2776]. Stent [2359]. Step [3191, 1014, 1874, 2373, 1917, 3735, 2960, 2783, 3510, 360, 2019, 891, 1715, 3531, 3616]. Step-by-Step [1917]. Step-Size [891]. Stepping [3253, 3103]. Stepsize [22]. Stepwise [2996, 3141, 3323]. Stereo [2458, 2657]. Stiefel [3633]. Stiefel-Type [3633]. Stiffness [3029]. Still [2415]. Stimulation [3854, 2934, 2594, 3181, 1773]. Stimuli [3777]. Stirling [2306]. Stirred [3694]. Stitching [2393]. Stochastic [2039, 1922, 1360, 3582, 3401, 1429, 3135, 3051, 431, 2541, 1169, 1997, 3498, 1953, 1458, 3067, 3095, 984, 1250, 406, 1171, 2016, 1591, 1924, 3900, 2105, 3902, 1036, 1655, 1453, 603, 1864, 96, 173, 628, 1707, 118, 216, 477, 992, 559, 493, 2080, 2575, 1422]. Stochastic-Machine [3900]. Stochastically [689]. Stochastically-Evolving [689]. Stock [1430, 3333, 1499, 3255, 704, 2471, 3117, 2031, 2928, 3723, 2077, 1548, 2891]. Stock-Price [2891]. Stocks [3723]. Stoichiometric [2993].

Stoichiometric-Based [2993]. Stokes [1877, 1141]. Stomach [2929]. Stopped [1952]. Storage [3081, 3366, 1179, 2048, 2438, 2688, 3502, 3755, 135, 212, 3265, 2703, 452, 3905, 2928, 1987]. Store [1043, 2369, 1593]. Storey [1475]. Storing [1420]. Storm [2433, 2315, 2291, 869, 759, 1487, 2829]. Story [3420]. Storytelling [2218]. Stranded [2285]. Strategic [3910, 2835]. Strategies [1517, 3853, 3135, 885, 3393, 934, 779, 409, 2334, 1502, 3222, 699, 3553, 2561, 3606, 1593, 1821, 2469]. Strategy [2896, 714, 3072, 3656, 596, 3359, 3203, 595, 3210, 1243, 2290, 308, 338, 800, 1776, 226, 98, 1113, 2908, 1920, 1192, 3242, 281, 3370, 346, 2842, 3833, 3482, 2358, 2847, 3737, 2829, 3679, 768]. Strategy-Solvable [1113]. Straw [3536]. Stream [2342, 1843, 1469, 781, 1787, 3244, 117, 2206, 1111, 1101, 638]. Stream-Based [1101]. Streaming [990, 1845, 665, 1162, 1826, 2361]. Streaming-Based [990]. Streamlined [3469, 2655]. Streams [3429, 2704, 143]. Street [3752]. Streetlights [1056]. Strength [471, 1330, 3281, 1862]. Strengthening [3019]. Stress [1939, 471, 3671, 20, 2490]. Stress-Responsive [2490]. Stress-Strength [471]. Stretched [155]. String [46, 1206, 136, 263, 1841, 990, 65, 349]. Strings [868, 817, 1316]. Strip [1526, 2590, 3529]. Stroke [3003, 2004, 2500, 3218, 2828]. Strong [3792, 2117, 2646]. Strongly [3787, 2783, 1251, 398, 2389]. Structural [3433, 3601, 46, 1181, 3074, 34, 1606, 67, 3095, 3318, 1959, 3883, 913, 347, 927, 2529, 2293, 657, 2101, 643, 1650, 3603, 3251, 3570]. Structure [2253, 2537, 303, 373, 3857, 3730, 862, 519, 2685, 1051, 1243, 1410, 1959, 2097, 835, 145, 1007, 3836, 1288, 2730, 1255, 1796, 1401, 3802, 3909, 1844, 2317, 118, 2572, 396, 2861, 786, 697, 3559, 768]. Structured [278, 1148, 3209, 2688, 3215, 3104, 1921, 1033, 3910, 1146, 3620, 3489]. Structures [2184, 2624, 15, 3627, 1703, 1245, 880, 4, 3165, 2463, 1535, 2065, 1704, 2361, 3618, 72]. STSGAT [2802]. STSGAT-Based [2802]. Stubborn [1589]. Student [3852, 2539, 3768, 236, 1559, 1566, 3102, 2645, 3237]. Students [2957, 1467]. Studies [3075, 2313, 3099, 3299, 3537, 1130]. Study [1289, 3316, 3741, 654, 3847, 3699, 871, 1430, 440, 2899, 2501, 2456, 3438, 3061, 753, 3040, 1027, 663, 2394, 3371, 511, 3140, 3205, 1557, 797, 2046, 2420, 3440, 3142, 3209, 434, 1076, 2476, 3590, 1392, 3446, 3289, 2892, 2678, 2449, 295, 3218, 3220, 1565, 3883, 3646, 1252, 1172, 3299, 1044, 1452, 1657, 2133, 3068, 3163, 1752, 162, 2951, 3461, 3595, 3396, 959, 494, 2723, 293, 2562, 3784, 1808, 2177, 3236, 3111, 3358, 855, 2645, 153, 742, 889, 3801, 3239, 1507, 3115, 1971, 1719, 254, 1914, 3243, 2979, 2223, 2233, 937, 1315, 1341, 3485, 3188, 3752]. Study [2577, 1737]. Studying [2507]. Style [3420, 3733, 3187]. Style-Transfer [3733]. Sub [3207, 2839, 737, 846, 634, 3403]. Sub-Band [2839]. Sub-Fractional [737]. Sub-Optimality [634]. Sub-Problem [3207]. Sub-Saharan [3403]. Sub-Systems [846]. Subarray [2591]. Subcellular [3319]. Subclasses [2754]. Subcubic [1211]. Subdivision [2411]. Subgradient [3378, 3721]. Subgraph [166, 3793, 2111, 901, 251, 541, 218]. Subgroup [2357, 2329, 3232]. Subhedge [2996]. Subject [1817, 3276, 52, 1535, 3598]. Subject-Sensitive [1817]. Subjective [3734, 105]. Subjectivity [3729]. Sublinear [194]. Submatrix [2397]. Subnetworks [344]. Suboptimal [845]. Subpath [1274]. Subregional [3761]. Subsampling [3158]. Subsea [786]. Subsequence [186]. Subset [519, 3567]. Subspace [1579, 204, 3095, 3661, 3915]. Substation [1642, 2860]. Substitutability [1541]. Substitution [80]. Substitutions [1379]. Substring [1158, 1168, 3623, 1295, 186]. Substrings [1206]. Subtle [3017]. Subtraction [1319, 39, 2790, 859, 872]. Subtractor [1446]. Subtypes [2806]. Subtyping [1314]. Subway [1154, 742]. Success [1031, 2666]. Successful [3466]. Successive [1950, 129, 2100]. Succinct [2672]. Sudan [3672]. Sudoku [1113]. Sufficient [664]. Suffix [175, 92, 649, 1295, 1409]. Suffix-Sorting [92]. Suicide [950]. Suitability [2453, 1791]. Suitable [2865, 2902, 3125]. Suite [1326]. Sum [213, 211, 2591, 2131, 1405, 2397]. Summability [3480]. Summarisation [1504]. Summarization [627, 3824, 2059, 2101]. Summarized [2589]. Summary [144]. Summation [2905]. SumMER [2101]. Sums [3030, 1986]. Sun [1633]. Super [3647, 3706, 694, 2749, 3770, 1545, 2964]. Super-Exponential [3647]. Super-Resolution [3706, 694, 2749]. Super-SBM [3770]. Superbubble [825]. Superconductor [2465]. Supercritical [3639]. Supergraph [1473]. Supergrid [1665, 1996]. Superimposed [3017]. Superiorities [3419]. Supermarket [2032]. Superpixel [3803]. Supersonic [1072, 3343, 1849]. Supervised [2228, 930, 3804, 2756, 1811, 312, 2924, 371, 3557, 1506, 3881, 1302, 1740, 2209, 2270, 670, 2932, 3113, 2494, 2566, 1149, 2266, 2647, 3404, 3631, 1449, 2748, 3842]. Supervision [1741, 1288]. Supplier [594]. Supply [2623, 1488, 566, 2437, 1596, 1864, 1459, 1482, 2928, 487, 232]. Supply-Chain [566]. Support [1388, 84, 2129, 1670, 2589, 457, 1046, 781, 3096, 2431, 1249, 735, 1044, 375, 2844, 2600, 3234, 1808, 3358, 1283, 56, 1057, 421, 1734, 1979, 3591, 1544, 3540, 2931, 3137, 3572]. Supported [301, 736]. Supporting [2585, 956, 93]. Supports [1970]. Suppressed [909]. Suppression [29, 696, 686, 2730, 2220, 2914]. Supremum [1126]. Surface [3781, 1180, 1714, 3795, 3740, 2283, 1703, 2007, 2590, 2214, 1035, 3109, 842, 1749, 2983, 5, 3529, 187, 719, 750, 1505, 3127]. Surfaces [2089, 1297, 1880, 1586, 1024]. Surge [2315]. Surpasses [3768]. Surprisal [2404]. Surprisal-Based [2404]. Surrogate [3081, 2407, 3256, 2463, 1705, 857, 1650, 1878, 786]. Surrogate-Based [857, 1650]. Surveillance [3371, 2555, 1322, 859]. Survey [1989, 1158, 1139, 1992, 1797, 1042, 355, 3756, 2740, 1614, 3637, 2858, 2541, 2388, 1039, 2543, 1583, 1088, 2006, 356, 19, 3217, 3646, 2553, 1291, 2660, 1015, 946, 2902, 1969, 1274, 2258, 2069, 2183, 16, 1259, 898, 2833, 3290]. Surveys [2271, 3450, 763]. SURVISHNO [976]. Survival [2279, 3087, 2004, 1800]. SUS304 [3211]. Susceptibility [3494]. Suspension [543, 3352, 953, 729]. Sustainability [1267, 2190]. Sustainable [2817, 3849, 1551, 2585, 374, 3267, 3726, 2755, 2928, 1915, 3916, 3290]. Sustaining [184]. SVD [1014, 14, 1794, 433]. SVM [224, 610, 3088, 457, 2023, 3400, 1012]. SVMD [2387]. SVR [2506, 593, 1548]. SVR-Empowered [2506]. SVSL [1493]. Swap [2641]. Swarm [3080, 562, 2988, 3130, 2455, 1010, 512, 171, 2155, 2587, 554, 457, 2145, 2476, 3593, 711, 2454, 3147, 1084, 446, 2852, 1157, 269, 308, 2867, 3890, 3456, 2291, 728, 2660, 382, 494, 2682, 1219, 1535, 237, 3073, 3120, 3036, 1758, 2710, 2374, 2154, 3536, 47, 805, 3484, 216, 3728, 3540, 691, 1198, 1643, 228, 1510, 2797]. Swarm-Based [3147]. Sway [3547]. Sweeping [2343]. Swimming [1785]. Swin [3621, 2749]. Swing [3472]. Switch [3050]. Switch-Controller [3050]. Switching [1863, 2974, 2890, 3489, 2320]. SXP [3437]. Sybil [3085, 2117]. SybilSocNet [3085]. Sylvester [488]. Symbiotic [3012, 1448]. Symbol [3877]. Symbolic [2045, 1207, 1324, 2119, 3919]. Symmetric [1042, 1610, 69]. Symmetrical [2320]. Symmetry [3874, 601, 2954]. Symmetry-Based [3874]. Symmetry-Inspired [2954]. Symptom [3451]. Symptoms [2327]. Synaptic [3125]. Synchronization [2296, 3582, 1888, 1246, 1488, 3278, 304, 2016, 2263, 286, 807]. Synchronized [2715]. Synchronous [3276, 3807, 1660]. Syndrome [1697, 3533]. Synergic [2066]. Synergistic [1789]. Synesth [2843]. Synoptic [1653]. Syntactical [614]. Syntenic [2843]. Syntenies [1400]. Synthesis [2949, 3262, 3024, 3449, 462, 2913, 2730, 3039, 3407, 2944, 3765]. Synthesisable [3597]. Synthesizing [3513, 3620]. Synthetic [115, 3727, 3832, 2906, 3032, 3881, 3225, 3049, 2771, 3115, 3177, 1831, 3537, 2650, 2204, 3562, 1474]. SynthSecureNet [3291]. System [3492, 1549, 1143, 692, 803, 1021, 2296, 2896, 1860, 1790, 1942, 1179, 2585, 2298, 127, 370, 2624, 3796, 3059, 431, 282, 772, 2589, 3089, 235, 2787, 3394, 3698, 511, 1182, 2224, 2141, 274, 2234, 2251, 1046, 2438, 3014, 2712, 3696, 2262, 3871, 1586, 830, 2431, 1438, 3702, 532, 3806, 2942, 3782, 1754, 2380, 130, 3325, 3887, 849, 3690, 2164, 1871, 1109, 3062, 3349, 2555, 2715, 2170, 20, 2889, 288, 1318, 3583, 2442, 801, 2882, 2643, 2972, 3042, 1618, 652, 2344, 2674, 2269, 990, 469, 3801, 1284, 2675, 1864, 2930, 1322, 1454, 3174, 383, 2032, 845, 1644]. System [2339, 3120, 3907, 953, 3280, 2568, 85, 1979, 461, 2070, 784, 2072, 316, 877, 676, 247, 2890, 3591, 2608, 2571, 1782, 719, 1341, 2327, 3353, 3483, 613, 1832, 1887, 1762, 1451, 883, 1436, 2743, 2931, 496, 814, 3395, 3574, 769, 729, 1988, 1461, 3060]. System-Optimal [3871]. Systematic [1908, 2406, 3681, 3198, 3286, 3697, 1998, 3497, 1103, 3205, 2814, 2005, 2378, 3019, 3671, 2761, 2218, 2880, 3105, 3604, 2319, 3695, 2959, 2918, 2469, 3602, 3403, 2174, 1904, 2472, 3544, 1728, 1882, 3863]. Systemic [1410]. Systems [3492, 328, 455, 423, 1798, 3849, 1850, 3195, 3699, 2618, 1797, 454, 57, 3366, 305, 1118, 3132, 1042, 2418, 1232, 24, 2268, 3045, 2524, 761, 2439, 1317, 370, 596, 1892, 2820, 1262, 1670, 2246, 3495, 1702, 3838, 116, 372, 2457, 667, 3139, 2158, 2543, 678, 2526, 1846, 1730, 1686, 3208, 2112, 1351, 688, 209, 2503, 2048, 717, 2661, 3619, 2454, 2751, 3750, 3093, 3374, 528, 2666, 418, 1457, 3448, 78, 343, 199, 1612, 1246, 3043, 3321, 1032, 984, 1958, 532, 616, 764, 1694, 458, 1212, 1007, 279, 965, 296, 2923, 2868, 3692, 716, 994, 135, 1881, 1776, 2325]. Systems [929, 2311, 2995, 1109, 2668, 3311, 1800, 846, 265, 481, 1857, 728, 3894, 1298, 2721, 2248, 2951, 1255, 2459, 2703, 2255, 2103, 3553, 2600, 2912, 2023, 3708, 1403, 3234, 1761, 2265, 3235, 1009, 3169, 3522, 444, 900, 2343, 2565, 1419, 2261, 1865, 2026, 341, 865, 3472, 1453, 1333, 1538, 3048, 2028, 1539, 325, 665, 492, 223, 114, 1361, 2333, 943, 1734, 3825, 588, 2147, 1980, 1028, 2440, 2106, 3070, 2168, 1228, 2073, 1676, 2472, 1379, 270, 337, 445, 1660, 3571, 2738, 477, 623, 2612, 3835, 705, 3489, 525, 895, 1987, 3548, 294, 3137, 1235]. Systems-Based [3132].

T [2624, 2593, 1450, 2181, 623]. T-Profile [1450]. T-SNE [2181]. T-Type [623]. T2 [970]. Table [46, 3002, 1518, 2935, 2067, 1704]. tables [74]. Tablet [2785]. Tablet-Acquired [2785]. Tablets [3127]. Tabu [3028, 630, 1193]. Tabular [3371, 149, 3881, 3225, 909, 2691, 2264, 2204]. Tabular-Based [149]. Tackle [2257]. Tackling [3185]. Tactics [2948]. Tagged [2699]. Tail [2372, 1986]. Tailhook [3159]. Tailored [1960, 1776]. Taken [1867]. Taking [2325]. Tale [3816]. Tall [1959]. Tampering [2896]. Tandem [2047]. Tangent [239]. Tank [2175, 3694, 842]. Tanks [2963]. Tantrix [132]. Tapered [701]. Tapers [1036]. Tardiness [1945]. Target [282, 975, 2903, 9, 1532, 3298, 2684, 281, 3813, 3377, 228]. Target-Level [3298]. Targets [750]. Tarifit [3741]. Task [1325, 3804, 595, 892, 3782, 1516, 2205, 2201, 3844, 2017, 2027, 1448, 3174, 2648, 3917, 814, 3490]. Task-Driven [2201]. Task-Oriented [3174]. Task-Relevant [3917]. Tasking [3256]. Tasks [2940, 2956, 1270, 1494, 2142]. Taxonomy [2248, 3357, 2025]. Taylor [2780, 1261, 1652, 3876]. TBRNet [1111]. TC [3279]. Tchebichef [3038, 610]. TD3 [2487]. Teacher [3768, 2957, 1566, 3659, 1978, 2180]. Teacher-Learner [1978]. Teaching [2094, 838]. Teaching-Learning-Based [2094]. Team [3561, 329, 3594, 2857, 3036]. Technical [1110, 9, 3877, 3573, 419, 2565]. Technician [1556]. Technique [1578, 3438, 230, 2895, 482, 2784, 2641, 3902, 2974, 1048, 2523, 1315, 270, 907]. Techniques [2375, 2864, 3397, 3195, 1797, 1021, 2391, 89, 2171, 1942, 2765, 2910, 3637, 867, 3861, 26, 3201, 1835, 2420, 3440, 3142, 1913, 125, 1631, 3753, 2449, 1144, 3508, 1674, 704, 428, 3511, 3883, 886, 3692, 1881, 2557, 2284, 2886, 3607, 1814, 3691, 2136, 2450, 3110, 3168, 3662, 3172, 1766, 2178, 2727, 2031, 504, 3241, 40, 3122, 3912, 88, 2521, 3485, 2192, 359, 3570, 2891]. Technologies [3286, 2977, 3442, 2865, 3426, 1879, 94]. Technology [692, 3822, 858, 235, 2051, 2909, 1871, 2844, 1122, 3659, 2238, 676, 3065, 841]. Telecommunication [3261, 2888, 2879, 2424, 2664]. Telegraph [3254]. Teleoperators [3197]. Telescope [2531]. Temari [1885]. Temperature [2470, 3142, 1385, 1115, 2465, 3368, 3710]. Temperatures [2139]. Tempered [1682, 3559]. Template [976, 2992, 1930, 653]. Template-Based [2992]. Template-Matching [1930]. Templates [3631]. Temporal [1670, 1145, 1730, 1951, 2805, 2430, 1957, 3098, 2289, 3365, 3390, 3888, 1962, 1324, 2109, 2269, 3240, 2648, 3246, 3353, 3182, 3914, 752, 3285]. Temporal-Correlation [3285]. Temporary [3099]. Tendency [2088].

Tennessee [2577]. Tennis [1048]. Tensile [1862]. Tension [2260, 830]. Tensor [2093, 2526, 3870, 3322, 351, 2484, 625, 3603]. Tensor-Based [2526]. Tensorial [2187]. Terfenol [3211, 3343]. Terfenol-D [3211, 3343]. Term [1002, 3699, 3436, 3643, 2260, 1519, 2511, 3749, 1074, 713, 1951, 1829, 1428, 141, 403, 917, 3891, 2813, 614, 1685, 588, 3121, 2612, 3188, 2656, 3548]. Terminal [3530]. Terminals [1293]. Termination [3051]. Terminus [2566]. Termite [3514]. Terms [2851]. Terrains [575]. Territorial [1647]. Terrorism [854]. Tertiary [3220]. Tesla [1452]. Tessellation [2983]. Test [3850, 2792, 1234, 397, 2329, 2980, 3179, 1650, 402]. Tested [22]. Testing [1993, 2108, 1994, 3601, 156, 739, 1869, 3019, 1859, 2796, 288, 1433, 577, 2185, 2475]. Tests [3527, 112]. Tetrahedral [3114]. Text [1231, 2949, 1073, 2149, 2899, 191, 2546, 926, 1381, 570, 3212, 2009, 1043, 148, 1102, 1464, 80, 2319, 1866, 3407, 144, 2801, 2142, 3065, 2833, 3920, 1603]. Text-Based [3920]. Text-Guided [3407]. Text-Image [2009]. Text-to-Speech [2949]. Textile [967]. Textile-Reinforced [967]. Texts [2899, 1928, 2281, 3162, 3535, 674, 2241]. Textual [364]. Texture [1867, 262, 3915, 733]. Texture-Adaptive [3915]. Texture-Less [733]. TFHE [1932]. Their [1675, 1922, 820, 971, 2042, 850, 2281, 2003, 67, 3448, 563, 1032, 1958, 1212, 3606, 684, 3589, 3626, 1731, 588, 441, 1771, 2848, 137, 651]. Theorems [1745, 3480]. Theoretic [922, 2533, 1401]. Theoretical [2584, 2712, 1248, 2794, 846, 2335, 3111, 3600]. Theories [1805]. Theorietage [387]. Theory [3656, 1201, 787, 235, 3267, 1741, 3870, 823, 105, 2271, 3450, 3878, 1468, 1774, 2329, 3288, 3662, 2151, 1094, 3537, 100, 391, 646, 502]. Theory-Based [1094]. Theory-Inspired [502]. Therapeutic [984, 3758]. Therapy [2624, 683, 3506, 3598]. Thermal [252, 1657, 2064, 580]. Thermo [2139]. Thermo-Mechanical [2139]. Thermodynamic [2012, 2649]. Thermograms [3089, 1835]. Thermography [3115, 1706]. Thermomechanical [2644]. Thermoplastic [3913]. Theses [3742]. Theta [2648]. Theta* [1802, 2082]. Theta-Gamma [2648]. Thick [1277, 1336]. Thickness [1373]. Things [2461, 1908, 3286, 170, 1531, 3802]. Thinning [2502]. Third [2780, 291, 283, 318, 2239]. Third-Order [2780, 2239]. Thoracic [2624, 91]. Threading [506, 193]. Threat [2410]. Threats [3409]. Three [2855, 2416, 229, 15, 2128, 3826, 2875, 2488, 1377, 1140, 3330, 3803, 3007, 2862, 1864, 3404, 2202, 3644, 3530, 3186, 623, 3309, 2861]. Three-Class [3803]. Three-Dimensional [2488, 1377, 3330, 3404, 3530, 2861, 15]. Three-Level [1864, 623]. Three-Point [3007]. Three-Qubit [3826]. Three-Way [2862, 3186]. Threshold [2973, 2172, 489, 3369, 1774, 143, 851, 3183, 1012]. Threshold-Based [851]. Thresholding [1161, 989, 3477, 780]. Thresholds [3023]. Thrift [81]. Throughput [2784, 1711]. Ticket [1667]. Tickets [2129]. Tidal [809]. Tiling [1669, 1768]. Time [3642, 3429, 2321, 2341, 1549, 1498, 166, 654, 2116, 3194, 2144, 2966, 2188, 3493, 2659, 1404, 1993, 3851, 3197, 1938, 2990, 1168, 3351, 1744, 2927, 414, 1699, 2506, 1159, 2172, 3829, 2539, 2356, 3387, 600, 1888, 386, 2165, 3610, 683, 3749, 225, 1948, 189, 3495, 243, 3087, 1133, 617, 2000, 2457, 2711, 667, 2903, 3092, 3736, 3866, 430, 1829, 2718, 2224, 194, 574, 3868, 3292, 329, 1351, 1240, 2948, 586, 3832, 38, 1087, 2438, 1825, 1955, 1281, 1207, 1584, 648, 3783, 343, 136, 2227, 3098, 1889, 1022, 1612, 1246, 2709, 210, 1562, 2289, 1032, 2008, 1665, 3744, 3253]. Time [185, 1754, 704, 1187, 1468, 1563, 2331, 3100, 259, 689, 764, 1211, 1658, 145, 581, 429, 1645, 268, 1000, 1477, 1962, 3690, 296, 1569, 1063, 3843, 1634, 2639, 2109, 3349, 3103, 1800, 21, 3893, 3748, 1815, 1055, 314, 3894, 2203, 2951, 2733, 1967, 80, 2599, 2483, 2561, 2312, 3708, 1440, 2601, 3274, 2932, 614, 1900, 852, 1220, 855, 180, 1851, 3769, 3526, 399, 1190, 1333, 1336, 3049, 2771, 911, 3402, 3634, 860, 1972, 2772, 2267, 1541, 2535, 1307, 1668, 3176, 3602, 2202, 597, 1338, 1037, 628, 3910, 368, 247, 106, 505, 2989, 2076, 2338, 3682, 2651, 1782]. Time [685, 3314, 3788, 1500, 599, 705, 525, 518, 3632, 3570, 3001, 1230, 791, 782]. Time-Constrained [1825, 2561]. Time-Delay [247, 525]. Time-Delays [2457, 705]. Time-Dependent [1240, 1281, 1468, 1338, 2076, 2651]. Time-Differenced [2659]. Time-Division [2951]. Time-Domain [648]. Time-Efficient [2341]. Time-Evolving [1133]. Time-Frame [386]. Time-Frequency [3744, 2639, 3570]. Time-Invariant [667]. Time-Limited [2539]. Time-of-Arrival [3690]. Time-Resolved [21, 38]. Time-Reversible [106]. Time-Series [2718, 2331, 1063, 3748, 2601, 399, 3910]. Time-Space [1168]. Time-Specific [2000]. Time-Stepping [3253]. Time-Triggered [3708]. Time-Universal [860]. Time-Variant [1477]. Time-Varying [3194, 3197, 2109, 1333, 3788]. Timeline [572, 3365]. Timely [2700]. Times [2260, 1945, 2669, 2373, 1263, 3389, 2290, 1561, 1250, 2369, 1965, 597, 1494, 3005, 589]. Timetabling [171, 295, 221, 2374]. Timing [896, 3598]. Tiny [3132, 2679, 2613]. Tire [3760, 3661]. Tissue [1215, 3909]. TLD [958]. TMP [3859]. TMP-M2Align [3859]. TOCA [3369]. TOCA-IoT [3369]. Tokens [2059]. Tolerance [3736, 1866, 988]. Tolerance-Based [1866]. Tolerant [3278, 2056, 2232, 2853]. Tomographic [167]. Tomography [1937, 2544, 2396, 1392, 619, 795, 832, 2143, 1003, 3386, 2528, 3684, 2604, 2126, 2728]. Tomtit [1900]. Tone [642]. Too [3754]. Tool [888, 823, 2412, 78, 2052, 3556, 3821, 2986, 2564, 1894, 1130, 8]. Toolbox [1618, 1896, 604]. Toolkit [2169]. Tools [2401, 1386, 3496, 123, 2911, 447, 1208, 2218, 3898, 93, 94, 3489]. Top [1145, 1420, 2917, 1129, 961, 1497]. Top- [1145]. Top-Hat [1497]. Top-Ranking [2917]. Topic [2422, 2000, 2932, 1808, 2177, 628]. Topic-Based [2422]. Topics [2000]. Topological [1209, 3150, 1962]. Topologies [2093, 21, 1376]. Topology [3859, 1840, 3500, 1703, 1134, 1035, 1312, 2314, 3114, 1162]. Topology-Aware [3859]. TOPSIS [3726, 1315]. TOREADA [3150]. TORNADO [2522]. Torque [2734]. Torrefaction [3536]. Tortilla [599]. Torus [124, 1242]. Total [3411, 225, 759, 597, 394, 677, 3050, 869]. Tour [111, 1371]. Tourism [2298, 2141, 1491, 1465, 3628]. Tours [2445]. Tower [3279]. Towing [1549]. Townhouses [3384]. Toys [3758]. Trace [912, 223]. Tracing [2785, 2759]. Track [991, 2395, 817, 3029, 2885, 1970, 309]. Track-Before-Detect [309]. Tracker [431]. Tracking [2966, 888, 2540, 242, 1069, 683, 976, 1119, 717, 48, 38, 2438, 707, 648, 1022, 9, 3148, 1444, 2488, 2425, 929, 1912, 1340, 2555, 2459, 2310, 18, 1655, 865, 3238, 1190, 1001, 281, 3176, 657, 2987, 653, 750, 3818, 752, 2860, 3270, 2770, 958]. Tracks [2531]. Tractabilities [163]. Tractability [159]. Tractor [2394]. Trade [1946, 1825, 744]. Trade-Off [1946, 1825, 744]. Tradeoff [276]. Tradeoffs [1168, 629, 3229]. Trading [714, 715, 1283, 2106]. Traditional [1144, 3461, 1284, 2760, 1451]. Traffic [1498, 2485, 103, 491, 2388, 1837, 3709, 1399, 2664, 2340, 586, 3055, 3871, 3447, 1437, 2504, 3099, 3515, 1838, 3731, 530, 982, 1121, 2859, 591, 1906, 2074, 1156, 3185, 1839, 524, 518, 963, 1026, 2577, 2381]. Traffic-Aware [2485]. Trail [1595]. Train [2692, 586, 1247, 917, 1512]. Trainable [2441]. Trained [2973, 1068, 983, 2941, 2309, 2346, 3717]. Training [2949, 1335, 2302, 3493, 1757, 3072, 3035, 3356, 236, 266, 2900, 2476, 1610, 3874, 2709, 3663, 969, 670, 2277, 3322, 1966, 1681, 1924, 2113, 1851, 3175, 2763, 3602, 2079, 2575, 1712, 1882]. Trains [3818, 1742, 2847, 3546]. Trait [190]. Traits [3497, 45, 2038]. Trajectories [2497, 3448, 2488, 2637, 773]. Trajectory [478, 2127, 1202, 2155, 2965, 3551, 727, 1124, 717, 1868, 2290, 1961, 2961, 538, 3530, 1906, 2981, 1839, 3305, 3270, 3547, 2770]. Trajectory-Based [2290]. Transactional [1419]. Transactions [3240]. Transcription [34, 41]. Transcutaneous [3181]. Transducers [720, 2734]. Transfer [1609, 1860, 2229, 934, 1628, 3500, 2460, 1886, 839, 1566, 2208, 2160, 982, 2744, 3733, 2513, 1479, 332, 1604]. Transferable [2533]. Transform [2360, 14, 2689, 2211, 1946, 2787, 3709, 3140, 342, 535, 755, 3062, 2160, 3468, 155, 555, 1310, 2995]. Transformation [3013, 1944, 484, 3442, 1886, 3023, 2231, 1590, 2294, 1333, 3674, 354]. Transformation-Based [484, 3674]. Transformations [3638]. Transformed [249]. Transformer [2949, 2235, 3054, 2581, 3665, 2481, 3868, 2943, 2486, 3799, 3313, 2009, 2694, 3888, 3420, 2749, 3386, 3602, 3341, 1511, 3648, 2655, 3621, 3385, 3190]. Transformer-Based [2235, 3888, 2655, 3190]. Transformers [3054, 2046, 2941, 2554, 1375, 3247, 3542]. Transforming [3666, 3376]. Transforms [3492, 69, 2605, 1497, 1910]. Transient [2418, 652, 2267, 332]. Transients [2976]. Transit [945, 1157, 1402]. Transition [510, 2015, 1651, 3242]. Transitions [3865]. Transitive [3812]. Transitivity [1756]. Translating [1229]. Translation [3741, 2001, 412]. Translation-Based [2001]. Translation/Rotation [412]. Translator [2694]. Transmembrane [3859]. Transmission [3658, 2503, 3750, 1687, 3447, 3753, 3826, 3688, 523, 1820, 656, 3670, 769]. Transmit [3688]. Transmitted [3403]. Transparent [3237]. TransPCGC [2554]. Transport [1910, 1619, 1437, 2257, 3782, 2214, 1569, 3344, 11, 2674, 1927, 1391]. Transportation [3642, 1680, 1850, 2962, 2455, 2462, 1267, 2633, 1465, 3786, 3608, 1356, 206]. Transporters [1664]. Transversal [1282, 3016]. Traub [353, 324]. Traumatic [2938]. Travel [3362, 2369, 1569, 1634, 3182]. Traveling [2485, 2127, 695, 164, 2640, 1515, 2123, 948, 1095, 2120, 1371, 217]. Travelling [1240, 2174]. Traversals [330]. Treatment [1939, 2975, 980, 37, 2249, 77, 2697, 2309, 3223, 2012, 1568, 2109, 1898, 2828]. Treatments [3563, 3600]. TRed [3917]. TRed-GNN [3917]. Tree [2537, 916, 2925, 426, 2672, 649, 2921, 797, 1329, 3638, 942, 1717, 2741, 1245, 2441, 2997, 1931, 3893, 3302, 1217, 2217, 3778, 1396, 3522, 1844, 1782, 2392, 2836, 3669, 2991, 1424, 1243]. Tree-Based [1245, 1931, 1217, 1424]. Tree-Crown [2392]. Tree-Width [3638]. Trees [175, 924, 2306, 526, 2975, 1699, 1578, 1159, 1234, 3266, 2365, 897, 2004, 209, 1559, 1248, 1295, 1409, 3513, 2426, 276, 1622, 998, 1932, 584, 2349, 3064, 3921, 1229]. Trend [3142, 2436, 620, 546]. Trends [1002, 3435, 2879, 3201, 1567, 1879, 2733, 2136, 2912, 3825, 3682]. Tri [2462, 2223, 2263]. Tri-Level [2462]. Tri-Model [2223]. Tri-Trophic [2263]. Trial [3543]. Trials [1395]. Triangle [522, 635, 3539]. Triangle-Free [3539]. Triangular [1574, 3079]. Triangulation [2283]. Triangulations [1114]. Tricalcium [2090]. Trie [2253]. Tries [1626]. Tries-Based [1626]. Trigger [2427]. Triggered [2285, 3708, 2908]. Trigonometric [3312, 69, 3030]. Trilinear [1077]. Trimming [818, 3517]. Trinity [1712]. Trio [3879]. Trip [2008, 3788]. Triple [1151, 3120, 625]. Triple-Pattern [1151]. Triplet [840, 563, 3819, 179]. Triplet-Based [3819]. Triplets [1909]. TRMs [3207]. Trophic [2263]. tropical [13]. Tropospheric [13]. Truckload [2443]. True [3652]. Truly [669, 1730]. Truncated [1531]. Truncations [1422]. Trunk [2597]. Truss [519, 2463]. Trust [3207, 2167, 447, 3893, 1223, 2575]. Trust-Region [3207, 2167, 2575]. Trust-Value [1223]. Trusted [1787]. Trustworthy [2584, 3558, 3660]. Tryptophan [67, 3104, 2939]. Tsetlin [1747, 2176]. TSK [492]. TSMPC [3760]. TSP [1281, 1084, 1668, 244]. Tsunami [1588]. Tubal [625]. Tube [2674]. Tuberculosis [3600]. Tubular [1683]. Tucker [2868]. Tukey [2777]. Tumor [1867, 683, 2416, 2878, 1392, 3446, 3501, 2259, 2914, 3341, 3912]. Tumors [77]. Tuned [2963]. TUNI [3405, 3814]. Tuning [3650, 692, 888, 617, 690, 3828, 1426, 626, 1270, 293, 468, 1914, 798, 425, 1421]. Tunnel [654, 1839]. Tunnels [1977]. Tuple [536]. Tuples [3281]. Turbine [3830, 3484, 2417, 2656]. Turbines [2618, 3287, 2181, 2734, 3414, 467, 999]. Turbofan [2978]. Turbulence [261, 3830]. Turnouts [1970]. Turns [259]. Tutors [3507]. TVF [2379]. TVF-EMD [2379]. TVL2 [944]. Tweaks [2948]. Tweets [2891]. Twenty [1425, 2319]. Twenty-Four-Hour [1425]. Twin [2692, 3148, 3236, 3916]. Twinning [2584]. Twins [1404, 2470, 3653, 3663]. TWISS [3904]. Twisted [1242]. Twister [858]. Twitter [2622, 2474, 439, 1340, 1968, 3168, 2152]. Two [654, 2116, 1014, 1666, 1430, 2085, 3, 236, 282, 2588, 980, 861, 207, 3255, 3619, 2960, 1825, 3871, 3447, 1652, 2591, 2634, 3321, 3830, 2214, 2783, 2551, 1590, 3455, 3276, 3692, 3416, 3896, 2432, 2723, 361, 684, 1755, 2026, 1152, 1089, 1507, 1121, 1845, 811, 1054, 1149, 173, 947, 996, 160, 1707, 337, 579, 870, 1111, 488, 3616, 782, 2037, 402, 1372, 1066, 3137]. Two-Archive [2037]. Two-Blocks [2026]. Two-Body [2723]. Two-Component [1121]. Two-Degree-of-Freedom [579]. Two-Derivative [2634, 3896]. Two-Dimensional [1430, 980, 1590, 870, 1372]. Two-Echelon [1152, 3137]. Two-Level [654, 782]. Two-Machine [947]. Two-Party [1666]. Two-Phase [861, 1755, 811, 1149, 173, 996]. Two-Sided [207]. Two-Stage [3255, 3619, 3321, 3830, 3455, 3416, 1845, 1707, 488, 402]. Two-Step [1014, 2960, 2783, 3616]. Two-Stream [1111]. Two-Way [2588, 3447, 3276]. Tympanic [2528]. Type [3629, 291, 302, 1943, 2356, 478, 761, 484, 264, 3820, 1019, 1112, 37, 1105, 2004, 754, 274, 2053, 737, 406, 728, 3900, 2362, 494, 2332, 3633, 1883, 475, 1594, 1807, 2119, 1865, 1745, 2897, 492, 3480, 362, 3405, 3814, 1107, 299, 445, 3827, 2275, 623, 1986, 664, 499]. Type-1 [475, 499]. Type-2 [2356, 478, 761, 484, 1105, 728, 2362, 494, 1883, 475, 492, 362]. Types [2615, 2042, 3091, 3335, 1776, 1920, 592, 3309]. Typhoon [1331]. Typical [581].

U [1407, 1880, 2482, 3840, 1003, 3739]. U-NET [2482, 3840, 1880, 1003]. U-Nets [3739]. U.S. [2456, 1410]. U2Net [2766, 3127]. U2Net-Based [3127]. UAS [2548]. UAS-S4 [2548]. UAV [1551, 1339, 791, 747, 3318, 965, 2961, 3332, 2426, 1367, 2310, 2312, 3469, 3238, 2680, 3534]. UAV-Based [3318]. UAVs [3193, 2834, 1237, 3455]. Ubiquitous [184]. UCINet [3844]. UFaceNet [1516]. UGV [2834]. UGVs [3315]. UK [3682]. UKF [2476]. UKF-Based [2476]. Ukraine [2152]. Ulcer [3062]. Ultimate [2550]. Ultra [3429, 3513, 3541]. Ultra-Large [3541]. Ultra-Large-Scale [3513]. Ultra-Low-Cost [3429]. Ultrasonic [3223, 2824, 2861]. Ultrasound [3844, 2914]. UMAP [1837, 3225]. UMAP-Based [3225]. UMAP-RF [1837]. Umbilical [3844]. Unannounced [1939]. Unary [1316]. Unavailability [2651]. Unbalanced [1974]. Unbiased [1944, 1433]. Unbounded [2757]. Uncertain [1730, 321, 764, 3344, 946, 1009, 1220, 1536, 1333, 947, 810]. Uncertainties [3054, 1840, 1032, 1477, 2384, 864]. Uncertainty [3566, 1138, 3363, 2793, 1072, 3394, 3091, 1611, 2633, 967, 994, 2471, 745, 1453, 996, 2070, 591, 2478, 1296, 3180, 3918, 1849, 1461, 1098]. Unconstrained [2616, 3768, 2785, 2783, 2682, 1475, 664]. Uncontrolled [3829]. Underdetermined [870]. Underdeveloped [3752]. Underground [3897, 3901, 1547]. Underlying [441]. Understandable [468]. Understanding [3822, 1343, 708, 1570, 2846, 1194, 3765]. Underwater [3714, 927, 1390, 1341, 1197]. Undirected [796]. Unequal [1572, 1298]. Unequal-Area [1572]. UNet [1003]. Unfolded [3230]. Unfolding [2353]. Unified [3785, 1537, 2073, 3645]. Uniform [3192, 2306, 1942, 355, 984, 380, 689, 3607, 731, 1988]. Uniformity [2709]. Uniformly [477, 682]. Unifying [2955]. Union [2188, 1577, 1336, 3222]. Unipolar [250, 2974]. Unique [1158, 1168, 2911]. Uniquely [822]. Unit [3025, 2663, 147, 1169, 2898, 2004, 3563, 1541, 3483, 2, 3918, 859]. Unit-Commitment [1169]. Unit-Demand [1541]. United [1122]. Units [3366, 2047, 3675]. Unity [1261]. Univariate [2506, 2916, 3868, 1830, 102, 145, 104]. Universal [1553, 55, 1748, 860, 3733, 3405, 3814]. University [171, 3410, 1559]. Unknown [328, 455, 423, 2458, 2457, 287, 1585, 333, 1246, 865]. Unlabeled [2782]. Unleashing [2891]. Unlicensed [1328]. Unlocking [3217]. Unmanned [2356, 2127, 1389, 983, 2741, 2857, 3238, 1821, 3476, 3534, 2933, 2613]. Unpredictable [3185]. Unranked [2925]. Unranking [1345, 1640]. Unraveling [3419]. Unrelated [2724, 2024, 599]. Unrooted [1159]. Unsampled [2843, 1479]. Unseen [2819]. Unsignalized [3434]. Unsteady [3343, 938]. Unstructured [967, 680]. Unsupervised [3614, 3012, 1073, 2235, 2228, 840, 3134, 866, 1068, 2858, 2111, 2756, 1954, 3445, 2943, 2892, 2421, 2288, 3246, 1135]. Unsuspected [2603]. UNSW [2721]. UNSW-NB15 [2721]. Until [3508]. Unveiling [2615, 3656, 2884, 3187]. Up-Downwind [774]. Update [3323]. Updated [3339]. Updates [1042]. UPGMA [2080]. Uplift [2249]. upon [1069]. Upper [2692, 985]. Upper-Limb [2692]. Upsampling [1728]. Upscaling [49]. UR10 [626]. Urban [2485, 2849, 3438, 3419, 3599, 1560, 1157, 3214, 2425, 3514, 3426, 3282, 3463, 982, 1221, 1280, 3124, 2651, 3293, 2237, 1512, 1742, 1026]. Urbanism [184]. Urdu [3648]. Urgent [1561]. URLs [3607]. Usage [1790, 546]. USC [1985]. Use [2173, 726, 1049, 2692, 2394, 2678, 3324, 951, 141, 1891, 3223, 385, 1518, 1708, 2746, 3662, 1332, 2202, 1193, 1130, 964]. Used [2642, 2405, 2204]. Useful [1702, 3555]. User [3431, 2084, 1328, 1025, 1144, 403, 1684, 2889, 2882, 2826, 3527, 890, 1676, 348, 2702, 1988]. User-Defined [2826]. User-Generated [1025]. User-Item [1684]. User-PC [1988]. Users [1776, 1691]. Using [3429, 2973, 3340, 3379, 2341, 3629, 1549, 1498, 2663, 1935, 2817, 3003, 654, 1073, 3686, 1493, 3347, 3054, 692, 1850, 3848, 803, 3397, 1788, 3430, 3658, 3081, 3665, 303, 2659, 3130, 3822, 2582, 1609, 315, 388, 1524, 3718, 3083, 930, 3261, 1995, 2360, 1709, 3262, 1435, 3436, 140, 2221, 14, 46, 2689, 905, 2622, 424, 2171, 1306, 3063, 1232, 13, 449, 2304, 1580, 2342, 3438, 3045, 1945, 71, 3035, 478, 2474, 753, 2434, 1317, 1425, 2888, 3428, 514, 171, 484, 928, 1180, 1946, 2155, 2165, 3599, 431, 683, 2858, 2111, 26, 1387, 3564, 746, 3342, 1074, 669]. Using [3040, 3576, 1383, 3089, 931, 922, 2444, 3371, 1880, 893, 1182, 3757, 1835, 1829, 1428, 2791, 2999, 2420, 2249, 1271, 2396, 854, 2482, 2546, 2503, 157, 1801, 3840, 3761, 3052, 2326, 1611, 3143, 2547, 1619, 2460, 3712, 3554, 2943, 2881, 1830, 2454, 1314, 1955, 2751, 1621, 2090, 3410, 2283, 3145, 2412, 1631, 2741, 49, 781, 1243, 2809, 3095, 2658, 1208, 1559, 3096, 1560, 1794, 1272, 1886, 3449, 1245, 199, 172, 2407, 1653, 3569, 1783, 356, 369, 1623, 2759, 1438, 2633, 2289, 3671, 1144, 969, 1925, 1362, 1157, 2094, 3806, 3508, 1502, 1488, 1674, 508, 2636, 3451, 1506]. Using [3744, 306, 944, 1754, 2214, 3154, 704, 3594, 1151, 3880, 3657, 3881, 3831, 2331, 3454, 3100, 3794, 2992, 3224, 1434, 2012, 222, 3646, 3303, 2909, 458, 3694, 3156, 3664, 268, 456, 620, 743, 1115, 1278, 1847, 2694, 3885, 2369, 3226, 3690, 1812, 3700, 3828, 994, 1923, 375, 582, 1657, 2133, 1871, 2995, 680, 1365, 3102, 3573, 2259, 3597, 2638, 2426, 3062, 2555, 2742, 3819, 3824, 1116, 3104, 1215, 977, 1800, 1476, 3331, 3011, 2098, 3348, 3106, 97, 1890, 1055, 1288, 412, 2198, 728, 2913, 2886, 3330, 1964, 3837, 1965, 2465, 1689, 3607, 3388, 495, 1354, 1439, 1897, 2136]. Using [2483, 2463, 2528, 3762, 2060, 3555, 2332, 3168, 1095, 3467, 1009, 852, 2061, 2062, 2564, 1534, 1883, 392, 475, 3468, 3112, 2837, 1605, 2196, 1023, 463, 341, 3350, 3237, 3172, 1190, 469, 1927, 1507, 79, 3115, 2494, 1970, 2566, 2790, 1616, 3764, 2677, 2178, 1972, 2930, 1322, 1691, 3779, 1482, 2567, 2684, 3117, 3352, 2308, 2680, 2502, 2266, 2413, 845, 776, 1710, 2349, 936, 953, 2333, 1915, 3004, 254, 2691, 2859, 2779, 3716, 2801, 3807, 3370, 3403, 2798, 1844, 2100, 2569, 3745, 605, 1930, 487, 2513, 3177, 950, 91, 2607, 877, 628, 1982, 2570, 2944]. Using [1856, 3533, 798, 3178, 3608, 2448, 1904, 1467, 465, 3380, 2154, 3422, 3723, 3586, 589, 3535, 1228, 505, 2608, 2182, 2571, 1405, 2452, 3334, 2102, 2716, 3912, 2838, 1107, 1259, 3345, 1449, 1521, 2873, 3775, 2777, 641, 1402, 3065, 227, 270, 389, 646, 606, 3585, 2428, 2836, 1156, 3738, 2192, 2163, 907, 2871, 1544, 819, 485, 2509, 1012, 1505, 1393, 3568, 1747, 249, 1750, 666, 2487, 3309, 1198, 1870, 231, 364, 3487, 1331, 1155, 2514, 3648, 2038, 1690, 433, 1131, 1350, 2081, 2872, 3598, 856, 2381, 2657, 53, 228, 1988, 3001, 82, 2891, 1718, 1200, 3572, 3202, 2244, 1497, 180]. Utility [595, 1058]. Utilization [3019, 145, 1122]. Utilized [2734]. Utilizing [991, 373, 3643, 3091, 3447, 1754, 829, 3272, 2601, 1535, 3481, 2613]. UWB [3690, 784]. UWB-Based [3690].

V [2195, 2701]. V&V [531]. V-SOC4AS [2195]. V100 [1452]. V2 [2525]. v8 [3251]. Vacancy [3426, 3752]. Vacation [3276]. Vaccination [2501]. Vaccine [1818]. Vacuum [2674]. Valence [1200]. Valet [2307]. Validating [1869]. Validation [3197, 3338, 1496, 708, 2359, 498, 1955, 3150, 626, 3595, 2882, 3684, 399, 1766, 3239, 3173, 1016, 2100, 1037, 1495, 1650, 3675, 1408]. Valuation [3057]. Value [2780, 2116, 2689, 1179, 2355, 2741, 2673, 271, 1774, 2693, 2095, 2019, 1685, 2837, 1594, 2684, 1223, 2478, 2479]. Valued [2757, 3713, 360, 327, 3633, 1848, 1471, 533, 651, 703]. Values [554, 2809, 3324, 2505, 3716, 1521, 2905, 2509, 3258]. Valve [224, 3400]. Vanadium [3649]. VANET [2464]. Vanishing [713, 3180]. Vapor [3429]. Variabilities [3501]. Variability [107, 3347, 3683, 2516, 614, 2873]. Variability-Based [3347]. Variable [2757, 3718, 2481, 921, 2390, 2008, 844, 3520, 569, 3414, 1715, 173, 774, 949, 394, 505, 270, 3584, 2275, 3015, 3290]. Variable-Length [2275]. Variables [510, 1317, 1661, 379, 2231, 3748, 2603, 254, 1986, 666]. Variance [1963, 2385, 1500]. Variance-Component [1963]. Variant [283, 1621, 1477, 2868]. Variants [298, 2003, 1491, 3016, 3687, 3471, 910, 1981, 1598]. Variation [3100, 1567, 2339]. Variational [2226, 3072, 42, 3277, 3208, 3263, 892, 2739, 1047, 3881, 561, 1743, 2103, 2563, 628, 2204, 1887, 1203]. Variational-Physics-Informed [3072]. Variations [2455, 2227, 1282, 2566, 2088]. Variety [3142]. Various [2511, 35, 3030, 1920]. VARMAX [2447]. Varying [1662, 3194, 3197, 2109, 1333, 1037, 1620, 3788]. Vase [3064]. Vasicek [2711, 737]. Vasicek-Type [737]. Vault [452]. VC [1065]. VC-Dimension [1065]. Vector [2342, 1670, 457, 3074, 1834, 157, 781, 3096, 1652, 1589, 285, 375, 1838, 1592, 3234, 451, 469, 56, 3404, 487, 1544, 3540, 2487, 3487]. Vectorization [3168]. Vectorized [2707]. Vectors [240, 442]. Vehicle [562, 3193, 2485, 2988, 3427, 726, 2195, 3084, 2356, 2127, 600, 2767, 491, 500, 1389, 2158, 3078, 983, 3144, 896, 3502, 2741, 3417, 2368, 2008, 1187, 3760, 1171, 288, 2345, 902, 1095, 3169, 2674, 3358, 3238, 1152, 1507, 1821, 2823, 953, 3479, 1038, 605, 3053, 3682, 2130, 2933, 3788, 2509, 1839, 2613, 3574, 789, 3189, 3137, 1235]. Vehicle-Routing [3084]. Vehicle-SOC [2195]. Vehicles [995, 3271, 991, 3434, 788, 3610, 1097, 3398, 9, 1362, 2488, 3460, 3606, 2857, 520, 564, 3476, 2072, 3534, 2760, 1341, 553]. Vehicular [3283]. Vein [1765]. Velarde [1019]. Velarde-Type [1019]. Velocity [2659, 3795, 3211, 1521, 2409]. Venous [1757]. Venue [3698]. Venues [2422]. Verbal [709]. Verification [2149, 1892, 2930]. Verifying [3233]. Verlag [74]. Version [3784, 3176, 2110]. Versioning [1419]. Versus [966, 3911, 2283, 1877, 1812, 2562, 2499]. Vertebral [2791, 3467]. Vertex [506, 3297, 3364, 2729, 551, 197, 3652]. Vertex- [506]. Vertex-Weighted [3297]. Vertical [2965, 1872, 3801]. Vertically [1708]. Verticillium [2426]. Very [199, 1249, 3833]. Vessel [2497, 591, 231, 2409]. Vessels [730]. VGG [2895]. VGG16 [1355]. via [92, 3347, 1993, 2470, 906, 1339, 1910, 3133, 3275, 2586, 3551, 3496, 1204, 2367, 3024, 3647, 1952, 2112, 181, 204, 3313, 184, 3096, 3032, 3876, 3151, 720, 2488, 1385, 1774, 2693, 3888, 1723, 2868, 3891, 1070, 1531, 1752, 2016, 1189, 2935, 2123, 1876, 3320, 1536, 539, 1749, 657, 1487, 833, 3719, 465, 1315, 1706, 847, 2955, 3913, 3915, 3184, 3771, 1602, 2797]. Vibration [3379, 3211, 3343, 648, 696, 1883, 1698, 2861]. Vibration-Assisted [2861]. Vibrations [3316]. Vibratory [3277]. Vibrotactile [421]. Video [2973, 1933, 2966, 3658, 3009, 1843, 3260, 874, 3371, 1308, 854, 2943, 3688, 1754, 3578, 3355, 3824, 1965, 982, 1322, 628, 2987, 1826, 2237, 1111, 976]. Video-Based [3009]. Videos [3645]. View [1811, 2904, 1083, 2429, 715, 1775, 2353, 806, 3488, 3752]. VIKOR [3186]. VIKOR-Based [3186]. Vineyard [2676, 2071]. Violations [2859, 3185]. Violence [2973, 3291]. Viral [3554]. Virtual [1732, 84, 2781, 702, 3013, 2692, 491, 3145, 1243, 1186, 3514, 2856, 1174, 3519, 6, 2952, 1341, 904]. Virus [3696]. Viruses [3796]. Viscoplastic [2133]. Viscosity [268]. Viscous [842]. Visibility [2279, 3327, 1307]. Vision [2581, 3665, 3759, 3132, 3024, 3020, 3313, 2636, 2051, 2796, 3386, 3348, 2310, 2208, 2494, 3542, 2409]. Vision-Based [3024, 2636, 2796]. Vision-Language [3759]. Visitation [1685]. Visiting [228]. Visits [3362]. Visual [3856, 1581, 2793, 282, 3777, 2444, 410, 460, 598, 9, 3646, 3706, 2639, 1109, 3762, 2088, 657, 3070, 364, 2157, 752, 958, 1693, 1461]. Visual-Motor [3070]. Visualization [867, 1051, 771, 1221, 128, 2073]. Visualizations [2901]. Visually [2934]. ViT [2581]. VITS [3765]. VITS-Based [3765]. vivo [220]. VLSD [2357]. VMP [2952]. VMP-ER [2952]. VNS [1796]. Vocabulary [1513]. Vocal [2582]. Vocalization [79]. VOE [3646]. VOF [3882]. Voice [3590]. Voiceprint [3226]. Void [1526]. Vol [2600]. Volatility [2711, 1863, 700, 1591, 774]. Volcanic [141]. Voltage [1641, 3783, 3514, 3807, 487, 332, 2797]. Volterra [2039, 1807, 2119]. Volume [673, 466, 1568, 2203, 3655, 3571]. Volumetric [1373, 3798]. Voting [1493, 969, 808, 2060, 943, 1028]. Voxels [3601]. Voxelwise [3424]. Voyage [741]. VPINN [3072]. vs [355, 427, 3289, 725, 2686]. Vulnerability [3071]. Vultures [2999]. VX [786].

WACNN [3571]. Wagons [2008, 3111]. Wait [512, 1393]. Waiting [2669, 1561]. Walk [1204, 3320, 2233]. Walk-Based [3320]. Walker [1867]. Walks [1957, 1718]. Walsh [2605]. War [2152]. War-Related [2152]. Warehouse [1238, 2491, 2471, 2026, 2299, 1301]. Warm [1533]. Warm-Up [1533]. Warning [927]. Warranty [3442]. Warshall [3812]. Warship [3887]. Wasserstein [1579, 2214, 1146]. Waste [3037, 3564]. Watching [20]. Water [3429, 3264, 35, 3794, 3843, 3746, 1215, 3463, 3553, 876, 5]. Waterborne [1697]. Watermarking [14, 2689, 433]. Watson [2309]. Wave [2757, 661, 2207, 1903, 1588, 1736, 11, 2343, 1230]. WaveCluster [140]. Waveform [393, 1147]. Waveforms [1905]. Waveguide [2109]. Wavelength [3327]. Wavelet [3054, 2360, 2689, 3709, 342, 3744, 385, 2160, 3468, 3475, 389, 555]. Wavelet-Based [3475]. Wavelets [71, 419, 72]. Waves [1714, 268, 23]. Way [1384, 2588, 3447, 1043, 3276, 1120, 2862, 3186]. Weak [289, 701, 3581, 2646, 3833]. Weakly [528, 2549, 3113, 811]. Weakness [1704]. Wear [1798, 3111]. Wearable [3822, 2224, 3451, 3239, 3065]. Wearables [3829, 429]. Wearing [2375]. Weather [3076, 2599, 2760, 3185]. Weave [508]. Web [191, 1177, 2969, 1871, 3555, 3591, 1513]. Web-Based [3555]. Weed [2676, 2001]. Weibull [471, 1943, 2510]. Weibull-Open-World [1943]. Weierstrass [3458]. Weight [1995, 1272, 1363, 829, 1465, 2405, 961, 3571, 733, 703]. Weight-Constrained [829]. Weighted [3614, 1418, 1945, 1555, 1011, 506, 320, 2970, 2851, 3297, 3503, 77, 213, 211, 808, 994, 2813, 2217, 1439, 2060, 200, 647, 3626, 1659, 1597, 642, 838, 593, 3685, 691, 1198]. Weighting [675, 1432, 255]. Weights [361, 907]. WEKA [1449]. Welfare [2124]. Well [1112, 3671, 376, 2337]. Well-Being [3671]. Well-Posedness [1112]. Well-Separated [376, 2337]. Wellbeing [2224, 1899]. Wellbeing-Monitoring [2224]. Wells [3801]. WENO [3616]. WES [174]. WGAN [1851]. Whale [926, 1961, 3769, 3659, 2772, 70]. Whale-Driven [3769]. Wheat [2924, 3536]. Wheel [3111, 2174]. Wheelchair [378]. Wheeled [2488]. Where [3417, 440]. Which [411]. Whispered [1672]. White [2581, 1366, 3052, 1251, 1277, 960]. White-Box [960]. Whole [2911, 3879, 935, 2754, 2033]. Whole-Genome [3879]. Wi [2044]. Wi-Fi [2044]. Wide [1267, 1186, 665]. Wide-Area [1186, 665]. Widely [1354]. Widest [2752]. Widow [2308]. Width [974, 629, 3638, 3575, 2324, 2114]. Widths [712]. Wiener [294]. Wikipedia [1049]. Wildfire [2070]. Will [105]. Win [706]. Wind [2618, 3436, 1074, 1855, 3265, 1254, 2734, 3414, 3484, 467, 999, 623, 2417, 2656, 1737]. Wind-Solar [3265]. Wind/Battery [1855]. Wind/Battery-Connected [1855]. Window [3087, 3100, 2714, 117, 3914, 104]. Windowed [644]. Windows [1945, 329, 102, 2008, 1187, 1656, 855]. Wing [2467, 1389, 3469]. Winglet [2522]. Wings [3469, 2522]. Winner [1189, 943]. Winter [1362]. Wire [2283, 849, 1974, 3614]. Wire-Grid [2283]. Wireless [2536, 1551, 1418, 1860, 544, 3864, 75, 473, 812, 234, 978, 1635, 19, 210, 2094, 306, 793, 226, 3823, 2017, 1015, 18, 161, 2882, 3634, 286, 391, 1197, 1447, 1131, 769]. Wise [2125, 3913]. Within [3578, 2425, 2010, 2755, 2823, 2212, 2287, 1193]. Without [3275, 3678, 76, 2936]. WLAN [2983]. Wolf [3794, 327, 468, 603, 2747, 578]. Wolf-Based [3794]. Word [2352, 459]. Word2Vec [3799]. Words [3324, 1374]. Work [2840, 3569, 3099, 3179, 599]. Work-Sampling [2840]. Worker [1238]. Workers [2124, 1988]. Workflow [3351, 607, 3782, 2386, 2573, 1229]. Workflows [172]. Working [3276]. Workloads [889]. Workpiece [1930]. Workshop [3580, 1510]. Workstation [3499]. Workstations [1451]. World [1732, 1943, 556, 2704, 2742, 2186, 789]. WorldView [390]. WorldView-2 [390]. WOW [1943]. Wrapper [3904]. Wrapper-Based [3904]. Writer [2713]. WSN [2983, 1094, 691]. WSN-Assisted [1094]. Wyrtki [3683].

X [2937, 1431, 3437, 2629, 2322, 808, 1570, 3386, 2557, 54, 3334, 260]. X-Ray [3437, 3386, 2937, 1431, 2322, 1570, 2557, 54, 260]. X-Rays [808, 3334, 2629]. XAI [1982, 3838, 1803, 1890, 3329]. XFEM [2373]. XGB4mcPred [1543]. XGBoost [2906, 2181, 2725].

Yang [74]. Yarn [508]. Year [1559]. Years [3128, 2319]. Yellow [3299]. Yield [2326]. Yoga [2034]. YOLO [3279, 3747, 3700, 3159, 2426, 1505, 3251]. YOLO-Based [3700]. YOLO-TC [3279]. YOLOv2 [518]. YOLOv3 [1056]. YOLOv5 [1362, 2535, 3552, 2654]. YOLOv5-Based [2535]. YOLOv7 [2483, 2062, 2613]. YOLOv7-Tiny [2613]. YOLOv8 [3680, 3164, 3762, 2859, 2933, 2831]. YOLOv8-Based [3762]. YOLOv8n [3680]. YOLOv8s [3610].

Z [3758]. ZDDs [1420, 158, 133]. Zehnder [3191]. ZenoPS [1832]. Zernike [2007]. Zero [1654, 1004, 1209, 3152, 846, 2131, 634, 3598]. Zero-Inflated [3598]. Zero-Knowledge [1654]. Zero-Shot [1004]. Zero-Sum [2131]. Zeros [1788, 3458]. Zeta [2976]. Zigzag [1228]. Ziv [116, 1409, 1587]. Zone [3569, 2639]. Zones [1551, 3698, 3099]. Zoning [2433].

References

Iwama:2008:EF

[1] Kazuo Iwama. Editor’s foreword. Algorithms (Basel), 1(1):1, September 2008. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/1/1/1.

Wiese:2008:ILL

[2] Andreas Wiese and Evangelos Kranakis. Impact of locality on location aware unit disk graphs. Algorithms (Basel), 1(1):2–29, September 2008. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/1/1/2.

Bein:2008:RCA

[3] Wolfgang Bein, Kazuo Iwama, and Jun Kawahara. Randomized competitive analysis for two server problems. Algorithms (Basel), 1(1):30–42, September 2008. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/1/1/30.

Li:2008:PCS

[4] Shuai Cheng Li, Yen Kaow Ng, and Louxin Zhang. A PTAS for the k-consensus structures problem under squared Euclidean distance. Algorithms (Basel), 1(2):43–51, December 2008. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/1/2/
43.

Senay:2008:MLE

[5] Gabriel B. Senay. Modeling landscape evapotranspiration by integrating land surface phenology and a water balance algorithm. Algorithms (Basel), 1(2):52–68, December 2008. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/1/2/52.

Moreau:2008:RVS

[6] Danielle Moreau, Ben Cazzolato, Anthony Zander, and Cornelis Petersen. A review of virtual sensing algorithms for active noise control. Algorithms (Basel), 1(2):69–99, December 2008. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
1/2/69.

Lancia:2008:MPC

[7] Giuseppe Lancia. Mathematical programming in computational biology: an annotated bibliography. Algorithms (Basel), 1(2):100–129, December 2008. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/1/2/100.

Zhao:2008:MLC

[8] Weixiang Zhao, Abhinav Bhushan, Anthony D. Santamaria, Melinda G. Simon, and Cristina E. Davis. Machine learning: a crucial tool for sensor design. Algorithms (Basel), 1(2):130–152, December 2008. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/1/2/130.

Jia:2008:AVN

[9] Zhen Jia, Arjuna Balasuriya, and Subhash Challa. Autonomous vehicles navigation with visual target tracking: Technical approaches. Algorithms (Basel), 1(2):153–182, December 2008. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/1/2/153.

Trepalin:2008:HCL

[10] Sergei V. Trepalin and Alexander V. Yarkov. Hierarchical clustering of large databases and classification of antibiotics at high noise levels. Algorithms (Basel), 1(2):183–200, December 2008. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
1/2/183.

Mariani:2009:FEF

[11] Stefano Mariani, Roberto Martini, and Aldo Ghisi. A finite element flux-corrected transport method for wave propagation in heterogeneous solids. Algorithms (Basel), 2(1):1–18, March 2009. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
2/1/1.

Noguchi:2009:NNA

[12] Yasuaki Noguchi, Fujihiko Matsumoto, Kazuo Maeda, and Takashi Nagasawa. Neural network analysis and evaluation of the fetal heart rate. Algorithms (Basel), 2(1):19–30, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic).

Bonafoni:2009:CDN

[13] Stefania Bonafoni, Fabrizio Pelliccia, and Roberta Anniballe. Comparison of different neural network approaches for the tropospheric profiling over the inter-tropical lands using GPS radio occultation data. Algorithms (Basel), 2(1):31–45, March 2009. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/2/1/
31.

Basso:2009:NBB

[14] Alessandro Basso, Francesco Bergadano, Davide Cavagnino, Victor Pomponiu, and Annamaria Vernone. A novel block-based watermarking scheme using the SVD transform. Algorithms (Basel), 2(1):46–75, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/1/46.

DiLena:2009:RTD

[15] Pietro Di Lena, Marco Vassura, Luciano Margara, Piero Fariselli, and Rita Casadio. On the reconstruction of three-dimensional protein structures from contact maps. Algorithms (Basel), 2(1):76–92, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/1/76.

Spratling:2009:SSI

[16] Benjamin B. Spratling IV and Daniele Mortari. A survey on star identification algorithms. Algorithms (Basel), 2(1):93–107, March 2009. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/2/1/93.

Contreras:2009:ANC

[17] M. Leonor Contreras, Eliseo Benítez, José Alvarez, and Roberto Rozas. Algorithm for nanotubes computer generation with different configurations. Algorithms (Basel), 2(1):108–120, March 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/108.

Nikoletseas:2009:PDA

[18] Sotiris Nikoletseas and Paul G. Spirakis. Probabilistic distributed algorithms for energy efficient routing and tracking in wireless sensor networks. Algorithms (Basel), 2(1):121–157, March 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/121.

Jin:2009:SPB

[19] Zhang Jin, Yu Jian-Ping, Zhou Si-Wang, Lin Ya-Ping, and Li Guang. A survey on position-based routing algorithms in wireless sensor networks. Algorithms (Basel), 2(1):158–182, March 2009. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/2/1/
158.

McGillis:2009:ASI

[20] Donald McGillis, Reginald Brearley, Khalil El Arroudi, and Geza Joos. The autonomous stress indicator for remotely monitoring power system state and watching for potential instability. Algorithms (Basel), 2(1): 183–199, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/2/1/183.

Martin:2009:EEK

[21] Joshua S. Martin, Katrina Simmons, and Alain Laederach. Exhaustive enumeration of kinetic model topologies for the analysis of time-resolved RNA folding. Algorithms (Basel), 2(1):200–214, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/200.

Banfelder:2009:ADS

[22] Jason R. Banfelder, Joshua A. Speidel, and Mihaly Mezei. Automatic determination of stepsize parameters in Monte Carlo simulation tested on a bromodomain-binding octapeptide. Algorithms (Basel), 2(1):215– 226, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/2/1/215.

Marzani:2009:HFW

[23] Alessandro Marzani and Ivan Bartoli. High frequency waves propagating in octagonal bars: a low cost computation algorithm. Algorithms (Basel), 2(1):227–246, March 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/1/227.

Burmakin:2009:SOD

[24] Eugene Burmakin, Alexander A. Fingelkurts, and Andrew A. Fingelkurts. Self-organization of dynamic distributed computational systems applying principles of integrative activity of brain neuronal assemblies. Algorithms (Basel), 2(1):247–258, March 2009. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/2/1/

Carnero:2009:DSN

[25] Mercedes Carnero, José L. Hernández, and Mabel C. Sánchez. Design of sensor networks for chemical plants based on meta-heuristics. Algorithms (Basel), 2(1):259–281, March 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/1/259.

Cilla:2009:RHA

[26] Rodrigo Cilla, Miguel A. Patricio, Jesús García, Antonio Berlanga, and Jose M. Molina. Recognizing human activities from sensors using hidden Markov models constructed by feature selection techniques. Algorithms (Basel), 2(1):282–300, March 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/1/282.

Hahn:2009:APD

[27] Federico Hahn. Actual pathogen detection: Sensors and algorithms — a review. Algorithms (Basel), 2(1):301–338, March 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/301.

Burr:2009:RII

[28] Tom Burr and Michael Hamada. Radio-isotope identification algorithms for NaI γ spectra. Algorithms (Basel), 2(1):339–360, March 2009. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/2/1/339.

Arabadzhi:2009:AAS

[29] Vladimir V. Arabadzhi. Algorithm for active suppression of radiation and acoustical scattering fields by some physical bodies in liquids. Algorithms (Basel), 2(1):361–397, March 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/1/361.

Hesse:2009:SBL

[30] Frank Hesse, Georg Martius, Ralf Der, and J. Michael Herrmann. A sensor-based learning algorithm for the self-organization of robot behavior. Algorithms (Basel), 2(1):398–409, March 2009. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
2/1/398.

Dugan:2009:GAA

[31] Nazım Dugan and Şakir Erkoç. Genetic algorithms in application to the geometry optimization of nanoparticles. Algorithms (Basel), 2(1):410– 428, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/2/1/410.

Fink:2009:PPI

[32] Florian Fink, Stephan Ederer, and Wolfram Gronwald. Protein-protein interaction analysis by docking. Algorithms (Basel), 2(1):429–436, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/1/429.

Torrens:2009:RII

[33] Francisco Torrens and Gloria Castellano. Resonance in interacting induced-dipole polarizing force fields: Application to force-field derivatives. Algorithms (Basel), 2(1):437–447, March 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/437.

Gardiner:2009:SFT

[34] Eleanor J. J. Gardiner, Christopher A. Hunter, and Peter Willett. Structural fingerprints of transcription factor binding site regions. Algorithms (Basel), 2(1):448–469, March 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/1/448.

Korosov:2009:SEA

[35] Anton A. Korosov, Dmitry V. Pozdnyakov, Are Folkestad, Lasse H. Pettersson, Kai Sørensen, and Robert Shuchman. Semi-empirical algorithm for the retrieval of ecology-relevant water constituents in various aquatic environments. Algorithms (Basel), 2(1):470–497, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/470.

Jakuschev:2009:NAM

[36] Stanislav Jakuschev and Daniel Hoffmann. A novel algorithm for macromolecular epitope matching. Algorithms (Basel), 2(1):498–517, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/1/498.

ElYoussef:2009:RCL

[37] Joseph El Youssef, Jessica Castle, and W. Kenneth Ward. A review of closed-loop algorithms for glycemic control in the treatment of type 1 diabetes. Algorithms (Basel), 2(1):518–532, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/518.

Gui:2009:IPT

[38] Lichuan Gui and John M. Seiner. An image pattern tracking algorithm for time-resolved measurement of mini- and micro-scale motion of complex object. Algorithms (Basel), 2(1):533–549, March 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/533.

Li:2009:MBS

[39] Sheng Li, MingXi Wan, and SuPin Wang. Multi-band spectral subtraction method for electrolarynx speech enhancement. Algorithms (Basel), 2(1):550–564, March 2009. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/2/1/550.

Sorokin:2009:MPT

[40] Alexey Sorokin, Nikita Boyko, Vladimir Boginski, Stan Uryasev, and Panos M. Pardalos. Mathematical programming techniques for sensor networks. Algorithms (Basel), 2(1):565–581, March 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/1/565.

Nguyen:2009:RAC

[41] Tung T. Nguyen and Ioannis P. Androulakis. Recent advances in the computational discovery of transcription factor binding sites. Algorithms (Basel), 2(1):582–605, March 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/1/582.

Bruggi:2009:MVF

[42] Matteo Bruggi and Paolo Venini. Mixed variational formulations for micro-cracked continua in the multifield framework. Algorithms (Basel), 2(1):606–622, March 2009. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/2/1/606.

Lin:2009:NNM

[43] Wei-Chin Lin and Glen S. Block. Neural network modeling to predict shelf life of greenhouse lettuce. Algorithms (Basel), 2(2):623–637, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/2/623.

Zou:2009:PRP

[44] Wei Zou and Vladimir V. Tolstikov. Pattern recognition and pathway analysis with genetic algorithms in mass spectrometry based metabolomics. Algorithms (Basel), 2(2):638–666, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/2/638.

Liu:2009:BAF

[45] Tian Liu and Rongling Wu. A Bayesian algorithm for functional mapping of dynamic complex traits. Algorithms (Basel), 2(2):667–691, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/2/667.

Bauer:2009:FSA

[46] Raphael André Bauer, Kristian Rother, Peter Moor, Knut Reinert, Thomas Steinke, Janusz M. Bujnicki, and Robert Preissner. Fast structural alignment of biomolecules using a hash table, N-grams and string descriptors. Algorithms (Basel), 2(2):692–709, June 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/2/692.

Vannucci:2009:APA

[47] Paolo Vannucci. ALE-PSO: an adaptive swarm algorithm to solve design problems of laminates. Algorithms (Basel), 2(2):710–734, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/2/710.

Gui:2009:AIT

[48] Lichuan Gui and John M. Seiner. Application of an image tracking algorithm in fire ant motion experiment. Algorithms (Basel), 2(2):735– 749, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/2/2/735.

Hu:2009:PUM

[49] Keqiang Hu and X. Frank Xu. Probabilistic upscaling of material failure using random field models — a preliminary investigation. Algorithms (Basel), 2(2):750–763, June 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/2/750.

Bonella:2009:SMI

[50] Sara Bonella, Walter Rocchia, Pietro Amat, Riccardo Nifosí, and Valentina Tozzini. SDPhound, a mutual information-based method to investigate specificity-determining positions. Algorithms (Basel), 2(2): 764–789, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/2/2/764.

Boyer:2009:SBB

[51] Michel Boyer, Ran Gelles, and Tal Mor. Security of the Bennett–Brassard quantum key distribution protocol against collective attacks. Algorithms (Basel), 2(2):790–807, June 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/2/790.

Mariani:2009:FAL

[52] Stefano Mariani. Failure assessment of layered composites subject to impact loadings: a finite element, sigma-point Kalman filter approach. Algorithms (Basel), 2(2):808–827, June 2009. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/2/2/
808.

Zheng:2009:CAD

[53] Bin Zheng. Computer-aided diagnosis in mammography using contentbased image retrieval approaches: Current status and future perspectives. Algorithms (Basel), 2(2):828–849, June 2009. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
2/2/828.

Mutihac:2009:BME

[54] Radu Mutihac. Bayesian maximum entropy based algorithm for digital X-ray mammogram processing. Algorithms (Basel), 2(2):850–878, June 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/2/850.

Hutter:2009:OPU

[55] Marcus Hutter. Open problems in universal induction & intelligence. Algorithms (Basel), 2(3):879–906, September 2009. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
2/3/879.

Redgwell:2009:CEC

[56] Robert D. Redgwell, Joseph M. Szewczak, Gareth Jones, and Stuart Parsons. Classification of echolocation calls from 14 species of bat by support vector machines and ensembles of neural networks. Algorithms (Basel), 2(3):907–924, September 2009. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/2/3/907.

Arimura:2009:CAD

[57] Hidetaka Arimura, Taiki Magome, Yasuo Yamashita, and Daisuke Yamamoto. Computer-aided diagnosis systems for brain diseases in magnetic resonance images. Algorithms (Basel), 2(3):925–952, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/3/925.

Miyazaki:2009:ICR

[58] Shuichi Miyazaki and Kazuya Okamoto. Improving the competitive ratio of the online OVSF code assignment problem. Algorithms (Basel), 2(3): 953–972, September 2009. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/2/3/953.

Huang:2009:AAN

[59] Yanbo Huang. Advances in artificial neural networks — methodological development and application. Algorithms (Basel), 2(3):973–1007, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/2/3/973.

Nikbakht:2009:AME

[60] Mehdi Nikbakht and Garth N. Wells. Automated modelling of evolving discontinuities. Algorithms (Basel), 2(3):1008–1030, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/3/1008.

Apostolico:2009:GCB

[61] Alberto Apostolico and Guido Drovandi. Graph compression by BFS. Algorithms (Basel), 2(3):1031–1044, September 2009. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
2/3/1031.

Lu:2009:RBF

[62] Weiying Lu and Peter de B. Harrington. Radial basis function cascade correlation networks. Algorithms (Basel), 2(3):1045–1068, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/2/3/1045.

Berger:2009:HML

[63] Florian Berger, Alexander Gilbers, Ansgar Grüne, and Rolf Klein. How many lions are needed to clear a grid? Algorithms (Basel), 2(3):1069– 1086, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/2/3/1069.

Eugenio:2009:FBA

[64] Francisco Eugenio and Javier Marcello. Featured-based algorithm for the automated registration of multisensorial/multitemporal oceanographic satellite imagery. Algorithms (Basel), 2(3):1087–1104, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/3/1087.

Russo:2009:ASM

[65] Luís M. S. Russo, Gonzalo Navarro, Arlindo L. Oliveira, and Pedro Morales. Approximate string matching with compressed indexes. Algorithms (Basel), 2(3):1105–1136, September 2009. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
2/3/1105.

Abellanas:2009:OCP

[66] Manuel Abellanas, Antonio L. Bajuelos, and Inês Matos. Optimal 2coverage of a polygonal region in a sensor network. Algorithms (Basel), 2(3):1137–1154, September 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/3/1137.

Hixon:2009:AAT

[67] John Hixon and Yana K. Reshetnyak. Algorithm for the analysis of tryptophan fluorescence spectra and their correlation with protein structural parameters. Algorithms (Basel), 2(3):1155–1176, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/3/1155.

Salvatori:2009:SPT

[68] Luca Salvatori and Niccolò Tosi. Stefan problem through extended finite elements: Review and further investigations. Algorithms (Basel), 2(3):1177–1220, September 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/3/1177.

Kim:2009:MSC

[69] Do Nyeon Kim and K. R. Rao. Multiplication symmetric convolution property for discrete trigonometric transforms. Algorithms (Basel), 2 (3):1221–1231, September 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/3/1221.

VanderSchaar:2009:CSW

[70] Mike Van der Schaar, Eric Delory, and Michel André. Classification of sperm whale clicks (Physeter Macrocephalus) with Gaussian-kernel-based networks. Algorithms (Basel), 2(3):1232–1247, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/2/3/1232.

Camps:2009:RMM

[71] Adriano Camps and José Miguel Tarongí. RFI mitigation in microwave radiometry using wavelets. Algorithms (Basel), 2(3):1248–1262, September 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/2/3/1248.

Zhao:2009:CBW

[72] Chong Zhao, Hanqiu Sun, Huawei Wang, and Kaihuai Qin. Compound biorthogonal wavelets on quadrilaterals and polar structures. Algorithms (Basel), 2(4):1263–1280, December 2009. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/2/4/
1263.

Pitt:2009:ARA

[73] Jonathan S. Pitt and Francesco Costanzo. An adaptive h-refinement algorithm for local damage models. Algorithms (Basel), 2(4):1281–1300, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/2/4/1281.

Lin:2009:BRE

[74] Shu-Kun Lin. Book review: Encyclopedia of Algorithms. Edited by Kao, Ming-Yang, Springer-Verlag GmbH, 2008; 1220 pages, 183 figures, 38 tables; Hard Cover. Price: EUR 309.- / CHF 479.50.- ISBN 978-0-387-30770-1. Algorithms (Basel), 2(4):1301–1302, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1301.

Eidenbenz:2009:ICG

[75] Stephan Eidenbenz, Gunes Ercal-Ozkaya, Adam Meyerson, Allon Percus, and Sarvesh Varatharajan. Incentive compatible and globally efficient position based routing for selfish reverse multicast in wireless sensor networks. Algorithms (Basel), 2(4):1303–1326, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1303.

Dey:2009:DMP

[76] Tamal K. Dey and Joshua A. Levine. Delaunay meshing of piecewise smooth complexes without expensive predicates. Algorithms (Basel), 2 (4):1327–1349, December 2009. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/2/4/1327.

Huo:2009:CGB

[77] Jing Huo, Kazunori Okada, Hyun J. Kim, Whitney B. Pope, Jonathan G. Goldin, Jeffrey R. Alger, and Matthew S. Brown. CADrx for GBM brain tumors: Predicting treatment response from changes in diffusionweighted MRI. Algorithms (Basel), 2(4):1350–1367, December 2009. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/2/4/1350.

Iswandy:2009:MAE

[78] Kuncup Iswandy and Andreas König. Methodology, algorithms, and emerging tool for automated design of intelligent integrated multi-sensor systems. Algorithms (Basel), 2(4):1368–1409, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1368.

Ren:2009:FBV

[79] Yao Ren, Michael T. Johnson, Patrick J. Clemins, Michael Darre, Sharon Stuart Glaeser, Tomasz S. Osiejuk, and Ebenezer Out-Nyarko. A framework for bioacoustic vocalization analysis using hidden Markov models. Algorithms (Basel), 2(4):1410–1428, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1410.

Nakamura:2009:LTT

[80] Ryosuke Nakamura, Shunsuke Inenaga, Hideo Bannai, Takashi Funamoto, Masayuki Takeda, and Ayumi Shinohara. Linear-time text compression by longest-first substitution. Algorithms (Basel), 2(4):1429– 1448, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/2/4/1429.

Short:2009:EHA

[81] Michael J. Short. Exact and heuristic algorithms for thrift cyclic scheduling. Algorithms (Basel), 2(4):1449–1472, December 2009. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1449.

Zinovev:2009:PRP

[82] Dmitriy Zinovev, Daniela Raicu, Jacob Furst, and Samuel G. Armato III. Predicting radiological panel opinions using a panel of machine learning classifiers. Algorithms (Basel), 2(4):1473–1502, December 2009. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/2/4/1473.

Tahmoush:2009:ISI

[83] Dave Tahmoush. Image similarity to improve the classification of breast cancer images. Algorithms (Basel), 2(4):1503–1525, December 2009. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/2/4/1503.

Awad:2010:CDS

[84] Mariette Awad, Yuichi Motai, Janne Näppi, and Hiroyuki Yoshida. A clinical decision support framework for incremental polyps classification in virtual colonoscopy. Algorithms (Basel), 3(1):1–20, March 2010. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/3/1/1.

Slabaugh:2010:RFS

[85] Greg Slabaugh, Xiaoyun Yang, Xujiong Ye, Richard Boyes, and Gareth Beddoe. A robust and fast system for CTC computer-aided detection of colorectal lesions. Algorithms (Basel), 3(1):21–43, March 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/3/1/21.

Zheng:2010:BCD

[86] Yufeng Zheng. Breast cancer detection with Gabor features from digital mammograms. Algorithms (Basel), 3(1):44–62, March 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/3/1/44.

Carpentieri:2010:ICD

[87] Bruno Carpentieri. Interactive compression of digital data. Algorithms (Basel), 3(1):63–75, March 2010. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/3/1/63.

Velazquez-Iturbide:2010:IIT

[88] J. Ángel Velázquez-Iturbide and Antonio Pérez-Carrasco. InfoVis interaction techniques in animation of recursive programs. Algorithms (Basel), 3(1):76–91, March 2010. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/3/1/76.

BenAbdelmelek:2010:BOB

[89] Sihem Ben Abdelmelek, Saloua Saidane, and Malika Trabelsi. Base oils biodegradability prediction with data mining techniques. Algorithms (Basel), 3(1):92–99, March 2010. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/3/1/92.

Berry:2010:GED

[90] Anne Berry, Jean R. S. Blair, Jean-Paul Bordat, and Geneviève Simonet. Graph extremities defined by search algorithms. Algorithms (Basel), 3(2):100–124, June 2010. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/3/2/100.

Takizawa:2010:RPN

[91] Hotaka Takizawa, Shinji Yamamoto, and Tsuyoshi Shiina. Recognition of pulmonary nodules in thoracic CT scans using 3-D deformable object models of different classes. Algorithms (Basel), 3(2):125–144, June 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/3/2/125.

Adjeroh:2010:SSS

[92] Donald Adjeroh and Fei Nan. Suffix-sorting via Shannon–Fano–Elias codes. Algorithms (Basel), 3(2):145–167, June 2010. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
3/2/145.

Rossling:2010:FTS

[93] Guido Rößling. A family of tools for supporting the learning of programming. Algorithms (Basel), 3(2):168–182, June 2010. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
3/2/168.

Santos:2010:INT

[94] Álvaro Santos, Anabela Gomes, and António José Mendes. Integrating new technologies and existing tools to promote programming learning. Algorithms (Basel), 3(2):183–196, June 2010. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/3/2/
183.

Berry:2010:ICM

[95] Anne Berry, Romain Pogorelcnik, and Geneviève Simonet. An introduction to clique minimal separator decomposition. Algorithms (Basel), 3(2): 197–215, June 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/3/2/197.

Schurz:2010:ASS

[96] Henri Schurz. Algorithmic solution of stochastic differential equations. Algorithms (Basel), 3(3):216–223, September 2010. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
3/3/216.

Mehrotra:2010:SLR

[97] Sanjay Mehrotra and Zhifeng Li. Segment LLL reduction of lattice bases using modular arithmetic. Algorithms (Basel), 3(3):224–243, September 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/3/3/224.

Nedev:2010:RSM

[98] Zhivko Nedev. An O(n)-round strategy for the Magnus–Derek game. Algorithms (Basel), 3(3):244–254, September 2010. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
3/3/244.

Solomonoff:2010:RS

[99] Grace Solomonoff. Ray Solomonoff (1926–2009). Algorithms (Basel), 3(3):255–259, September 2010. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/3/3/255.

Vitanyi:2010:RSF

[100] Paul M. B. Vitanyi. Ray Solomonoff, founding father of algorithmic information theory. Algorithms (Basel), 3(3):260–264, September 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/3/3/260.

Kels:2010:CMA

[101] Shay Kels, Nira Dyn, and Evgeny Lipovetsky. Computation of the metric average of 2D sets with piecewise linear boundaries. Algorithms (Basel), 3(3):265–275, September 2010. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/3/3/265.

Jin:2010:UCI

[102] Qingwei Jin, John E. Lavery, and Shu-Cherng Fang. Univariate cubic L_{1} interpolating splines: Analytical results for linearity, convexity and oscillation on 5-point windows. Algorithms (Basel), 3(3):276–293, September 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/3/3/276.

Caramia:2010:FCT

[103] Massimiliano Caramia, Ciro D’Apice, Benedetto Piccoli, and Antonino Sgalambro. Fluidsim: a car traffic simulation prototype based on Fluid-Dynamic. Algorithms (Basel), 3(3):294–310, September 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/3/3/294.

Yu:2010:UCI

[104] Lu Yu, Qingwei Jin, John E. Lavery, and Shu-Cherng Fang. Univariate cubic L_{1} interpolating splines: Spline functional, window size and analysis-based algorithm. Algorithms (Basel), 3(3):311–328, September 2010. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/3/3/311.

Hutter:2010:CTE

[105] Marcus Hutter. A complete theory of everything (will be subjective). Algorithms (Basel), 3(4):329–350, December 2010. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
3/4/329.

Tsukiji:2011:RRC

[106] Tatsuie Tsukiji and Takeo Hagiwara. Recognizing the repeatable configurations of time-reversible generalized Langton’s ant is PSPACE-hard. Algorithms (Basel), 4(1):1–15, March 2011. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/4/1/
1.

Aboufadel:2011:QVC

[107] Edward Aboufadel, Robert Castellano, and Derek Olson. Quantification of the variability of continuous glucose monitoring data. Algorithms (Basel), 4(1):16–27, March 2011. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/4/1/16.

Burr:2011:DLS

[108] Tom Burr, Toshihiko Kawano, Patrick Talou, Feng Pan, and Nicolas Hengartner. Defense of the least squares solution to Peelle’s pertinent puzzle. Algorithms (Basel), 4(1):28–39, March 2011. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
4/1/28.

Shapira:2011:EDB

[109] Dana Shapira and James A. Storer. Edit distance with block deletions. Algorithms (Basel), 4(1):40–60, March 2011. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/4/1/
40.

Klein:2011:CMD

[110] Shmuel T. Klein and Dana Shapira. Compressed matching in dictionaries. Algorithms (Basel), 4(1):61–74, March 2011. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
4/1/61.

Nguyen:2011:AMT

[111] Viet Hung Nguyen. Approximating the minimum tour cover of a digraph. Algorithms (Basel), 4(2):75–86, June 2011. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/4/2/

Touboul:2011:GFT

[112] Jacques Touboul. Goodness-of-fit tests for elliptical and independent copulas through projection pursuit. Algorithms (Basel), 4(2):87–114, June 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/4/2/87.

Burr:2011:ALS

[113] Tom Burr, Todd Graves, Nicolas Hengartner, Toshihiko Kawano, Feng Pan, and Patrick Talou. Alternatives to the least squares solution to Peelle’s pertinent puzzle. Algorithms (Basel), 4(2):115–130, June 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/4/2/115.

Santos:2011:RSE

[114] Olga C. Santos and Jesus G. Boticario. Requirements for semantic educational recommender systems in formal e-learning scenarios. Algorithms (Basel), 4(2):131–154, June 2011. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/4/2/131.

Camps:2011:RFI

[115] Adriano Camps, Jerome Gourrion, Jose Miguel Tarongi, Mercedes Vall Llossera, Antonio Gutierrez, Jose Barbosa, and Rita Castro. Radiofrequency interference detection and mitigation algorithms for synthetic aperture radiometers. Algorithms (Basel), 4(3):155–182, September 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/4/3/155.

DeAgostino:2011:LZD

[116] Sergio De Agostino. Lempel–Ziv data compression on parallel and distributed systems. Algorithms (Basel), 4(3):183–199, September 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/4/3/183.

Ting:2011:AFI

[117] Hing-Fung Ting, Lap-Kei Lee, Ho-Leung Chan, and Tak-Wah Lam. Approximating frequent items in asynchronous data stream over a sliding window. Algorithms (Basel), 4(3):200–222, September 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/4/3/200.

Weinberg:2011:ALD

[118] Frank Weinberg and Markus E. Nebel. Applying length-dependent stochastic context-free grammars to RNA secondary structure prediction.

Algorithms (Basel), 4(4):223–238, December 2011. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
4/4/223.

Tarongi:2011:RFI

[119] Jose Miguel Tarongi and Adriano Camps. Radio frequency interference detection and mitigation algorithms based on spectrogram analysis. Algorithms (Basel), 4(4):239–261, December 2011. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
4/4/239.

Carrascosa:2011:SGP

[120] Rafael Carrascosa, François Coste, Matthias Gallé, and Gabriel Infante-Lopez. The smallest grammar problem as constituents choice and minimal grammar parsing. Algorithms (Basel), 4(4):262–284, December 2011. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/4/4/262.

Marschall:2011:ACC

[121] Tobias Marschall and Sven Rahmann. An algorithm to compute the character access count distribution for pattern matching algorithms. Algorithms (Basel), 4(4):285–306, December 2011. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
4/4/285.

Kieffer:2011:CSA

[122] John Kieffer. A catalog of self-affine hierarchical entropy functions. Algorithms (Basel), 4(4):307–333, December 2011. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/4/4/
307.

DiLillo:2012:CBT

[123] Antonella Di Lillo, Ajay Daptardar, Kevin Thomas, James A. Storer, and Giovanni Motta. Compression-based tools for navigation with an image database. Algorithms (Basel), 5(1):1–17, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/5/1/1.

Amano:2012:HST

[124] Kazuyuki Amano, Yuta Kojima, Toshiya Kurabayashi, Keita Kurihara, Masahiro Nakamura, Ayaka Omi, Toshiyuki Tanaka, and Koichi Yamazaki. How to solve the torus puzzle. Algorithms (Basel), 5(1):18–29, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/5/1/18.

Hernandez-Cabronero:2012:SSC

[125] Miguel Hernández-Cabronero, Ian Blanes, Michael W. Marcellin, and Joan Serra-Sagristà. Standard and specific compression techniques for DNA microarray images. Algorithms (Basel), 5(1):30–49, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/5/1/30.

Gagie:2012:NSP

[126] Travis Gagie. A note on sequence prediction over large alphabets. Algorithms (Basel), 5(1):50–55, March 2012. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/5/1/
50.

Chang:2012:AGC

[127] Chung-Liang Chang and Ho-Nien Shou. Application of genetic control with adaptive scaling scheme to signal acquisition in global navigation satellite system receiver. Algorithms (Basel), 5(1):56–75, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/5/1/56.

Pizzolante:2012:VBO

[128] Raffaele Pizzolante and Bruno Carpentieri. Visualization, band ordering and compression of hyperspectral images. Algorithms (Basel), 5(1):76– 97, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/5/1/76.

Olshen:2012:SSR

[129] Richard A. Olshen and Bala Rajaratnam. Successive standardization of rectangular arrays. Algorithms (Basel), 5(1):98–112, March 2012. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/5/1/98.

Leu:2012:SPC

[130] Fang-Yie Leu, Keng-Yen Chao, Ming-Chang Lee, and Jia-Chun Lin. A semi-preemptive computational service system with limited resources and dynamic resource ranking. Algorithms (Basel), 5(1):113–147, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/5/1/113.

Demaine:2012:MFR

[131] Erik D. Demaine, Martin L. Demaine, and Ryuhei Uehara. Any monotone function is realized by interlocked polygons. Algorithms (Basel), 5 (1):148–157, March 2012. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/5/1/148.

Kino:2012:IPA

[132] Fumika Kino and Yushi Uno. An integer programming approach to solving Tantrix on fixed boards. Algorithms (Basel), 5(1):158–175, March 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/5/1/158.

Yoshinaka:2012:FAS

[133] Ryo Yoshinaka, Toshiki Saitoh, Jun Kawahara, Koji Tsuruma, Hiroaki Iwashita, and Shin ichi Minato. Finding all solutions and instances of Numberlink and Slitherlink by ZDDs. Algorithms (Basel), 5(2):176–213, June 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/5/2/176.

Maruyama:2012:OAL

[134] Shirou Maruyama, Hiroshi Sakamoto, and Masayuki Takeda. An online algorithm for lightweight grammar-based compression. Algorithms (Basel), 5(2):214–235, June 2012. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/5/2/214.

Lu:2012:CSG

[135] Maohua Lu, Cornel Constantinescu, and Prasenjit Sarkar. Content sharing graphs for deduplication-enabled storage systems. Algorithms (Basel), 5(2):236–260, June 2012. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/5/2/236.

Iwamoto:2012:PTR

[136] Chuzo Iwamoto, Kento Sasaki, and Kenichi Morita. A polynomial-time reduction from the 3SAT problem to the generalized string puzzle problem. Algorithms (Basel), 5(2):261–272, June 2012. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
5/2/261.

Tsuiki:2012:ICT

[137] Hideki Tsuiki. Imaginary cubes and their puzzles. Algorithms (Basel), 5(2):273–288, June 2012. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/5/2/273.

Ismail:2012:MPA

[138] Leila Ismail and Liren Zhang. Modeling and performance analysis to predict the behavior of a divisible load application in a cloud computing environment. Algorithms (Basel), 5(2):289–303, June 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/5/2/289.

Chen:2012:ABF

[139] Toly Chen. An agent-based fuzzy collaborative intelligence approach for predicting the price of a dynamic random access memory (DRAM) product. Algorithms (Basel), 5(2):304–317, June 2012. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
5/2/304.

Barnathan:2012:MSU

[140] Michael Barnathan. Mammographic segmentation using WaveCluster. Algorithms (Basel), 5(3):318–329, September 2012. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
5/3/318.

Junek:2012:ULR

[141] William N. Junek, Linwood W. Jones, and Mark T. Woods. Use of logistic regression for forecasting short-term volcanic activity. Algorithms (Basel), 5(3):330–363, September 2012. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/5/3/330.

Azzopardi:2012:ICN

[142] Joel Azzopardi and Christopher Staff. Incremental clustering of news reports. Algorithms (Basel), 5(3):364–378, September 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/5/3/364.

Malinovsky:2012:MTF

[143] Yaakov Malinovsky and Jacob Kogan. Monitoring threshold functions over distributed data streams with node dependent constraints. Algorithms (Basel), 5(3):379–397, September 2012. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/5/3/
379.

Rankel:2012:BMA

[144] Peter A. Rankel, John M. Conroy, and Judith D. Schlesinger. Better metrics to automatically predict the quality of a text summary. Algorithms (Basel), 5(4):398–420, December 2012. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/5/4/

Lavery:2012:UAP

[145] John E. Lavery. Univariate L^{p} and ℓ^{p} averaging, 0 < p < 1, in polynomial time by utilization of statistical structure. Algorithms (Basel), 5(4):421– 432, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/5/4/421.

Lin:2012:IEI

[146] Jun-Lin Lin, Yu-Hsiang Tsai, Chun-Ying Yu, and Meng-Shiou Li. Interaction enhanced imperialist competitive algorithms. Algorithms (Basel), 5(4):433–448, December 2012. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/5/4/433.

Chen:2012:FUC

[147] Toly Chen. Forecasting the unit cost of a product with some linear fuzzy collaborative forecasting models. Algorithms (Basel), 5(4):449– 468, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/5/4/449.

Mahapatra:2012:CAD

[148] Amogh Mahapatra, Nisheeth Srivastava, and Jaideep Srivastava. Contextual anomaly detection in text data. Algorithms (Basel), 5(4):469–489, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/5/4/469.

Koessler:2012:ETB

[149] Denise R. Koessler, Benjamin W. Martin, Bruce E. Kiefer, and Michael W. Berry. The effects of tabular-based content extraction on patent document clustering. Algorithms (Basel), 5(4):490–505, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/5/4/490.

Sapozhnikov:2012:EHD

[150] German Sapozhnikov and Alexander Ulanov. Extracting hierarchies from data clusters for better classification. Algorithms (Basel), 5(4):506–520, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/5/4/506.

Abdelbar:2012:ABP

[151] Ashraf M. Abdelbar. Alpha–beta pruning and Althöfer’s pathology-free negamax algorithm. Algorithms (Basel), 5(4):521–528, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/5/4/521.

Benvenuti:2012:FEQ

[152] Elena Benvenuti, Giulio Ventura, and Nicola Ponara. Finite element quadrature of regularized discontinuous and singular level set functions in 3D problems. Algorithms (Basel), 5(4):529–544, December 2012. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/5/4/529.

Prosser:2012:EAM

[153] Patrick Prosser. Exact algorithms for maximum clique: a computational study. Algorithms (Basel), 5(4):545–587, December 2012. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/5/4/545.

Scholkmann:2012:EAA

[154] Felix Scholkmann, Jens Boss, and Martin Wolf. An efficient algorithm for automatic peak detection in noisy periodic and quasi-periodic signals. Algorithms (Basel), 5(4):588–603, December 2012. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
5/4/588.

Wuttke:2012:LFT

[155] Joachim Wuttke. Laplace–Fourier transform of the stretched exponential function: Analytic error bounds, double exponential transform, and open-source implementation “libkww”. Algorithms (Basel), 5(4):604– 628, December 2012. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/5/4/604.

Csiszar:2012:TGF

[156] Villõ Csiszár, Péter Hussami, János Komlós, Tamás F. Móri, Lídia Rejtõ, and Gábor Tusnády. Testing goodness of fit of random graph models. Algorithms (Basel), 5(4):629–635, December 2012. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
5/4/629.

Giuliani:2012:EDM

[157] Donatella Giuliani. Edge detection from MRI and DTI images with an anisotropic vector field flow using a divergence map. Algorithms (Basel), 5(4):636–653, December 2012. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/5/4/636.

Toda:2012:ECO

[158] Takahisa Toda. Extracting co-occurrence relations from ZDDs. Algorithms (Basel), 5(4):654–667, December 2012. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/5/4/
654.

Bonizzoni:2013:MDP

[159] Paola Bonizzoni, Riccardo Dondi, and Yuri Pirola. Maximum disjoint paths on edge-colored graphs: Approximability and tractability. Algorithms (Basel), 6(1):1–11, March 2013. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/6/1/1.

Tian:2013:MCD

[160] Ye Tian, Qingwei Jin, John E. Lavery, and Shu-Cherng Fang. ℓ^{1} major component detection and analysis (ℓ^{1} MCDA): Foundations in two dimensions. Algorithms (Basel), 6(1):12–28, March 2013. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/1/12.

Nikolidakis:2013:EER

[161] Stefanos A. Nikolidakis, Dionisis Kandris, Dimitrios D. Vergados, and Christos Douligeris. Energy efficient routing in wireless sensor networks through balanced clustering. Algorithms (Basel), 6(1):29–42, March 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/6/1/29.

Marzban:2013:CSP

[162] Marjan Marzban and Qian-Ping Gu. Computational study on a PTAS for planar dominating set problem. Algorithms (Basel), 6(1):43–59, March 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/6/1/43.

Uehara:2013:TIG

[163] Ryuhei Uehara. Tractabilities and intractabilities on geometric intersection graphs. Algorithms (Basel), 6(1):60–83, March 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/1/60.

Isaacs:2013:DTS

[164] Jason T. Isaacs and João P. Hespanha. Dubins traveling salesman problem with neighborhoods: a graph-based approach. Algorithms (Basel), 6 (1):84–99, March 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/6/1/84.

Takes:2013:CED

[165] Frank W. Takes and Walter A. Kosters. Computing the eccentricity distribution of large graphs. Algorithms (Basel), 6(1):100–118, March 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/6/1/100.

Akutsu:2013:PTA

[166] Tatsuya Akutsu and Takeyuki Tamura. A polynomial-time algorithm for computing the maximum common connected edge subgraph of outerplanar graphs of bounded degree. Algorithms (Basel), 6(1):119–135, March 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/6/1/119.

Ma:2013:ANN

[167] Jun Ma. Algorithms for non-negatively constrained maximum penalized likelihood reconstruction in tomographic imaging. Algorithms (Basel), 6(1):136–160, March 2013. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/6/1/136.

Kiraly:2013:SMF

[168] Tamás Király and Júlia Pap. Stable multicommodity flows. Algorithms (Basel), 6(1):161–168, March 2013. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/6/1/161.

Bailey:2013:OSI

[169] David H. Bailey and Marcos López de Prado. An open-source implementation of the critical-line algorithm for portfolio optimization. Algorithms (Basel), 6(1):169–196, March 2013. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/6/1/169.

Cirani:2013:ESM

[170] Simone Cirani, Gianluigi Ferrari, and Luca Veltri. Enforcing security mechanisms in the IP-based Internet of Things: an algorithmic overview. Algorithms (Basel), 6(2):197–226, June 2013. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/6/2/
197.

Chen:2013:SUC

[171] Ruey-Maw Chen and Hsiao-Fang Shih. Solving university course timetabling problems using constriction particle swarm optimization with local search. Algorithms (Basel), 6(2):227–244, June 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/2/227.

Jakob:2013:FRM

[172] Wilfried Jakob, Sylvia Strack, Alexander Quinte, Günther Bengel, Karl-Uwe Stucky, and Wolfgang Süß. Fast rescheduling of multiple workflows to constrained heterogeneous resources using multi-criteria memetic computing. Algorithms (Basel), 6(2):245–277, June 2013. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/2/245.

Solos:2013:GTP

[173] Ioannis P. Solos, Ioannis X. Tassopoulos, and Grigorios N. Beligiannis. A generic two-phase stochastic variable neighborhood approach for effectively solving the nurse rostering problem. Algorithms (Basel), 6(2): 278–308, June 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/6/2/278.

Mazza:2013:MSC

[174] Tommaso Mazza and Stefano Castellana. Multi-sided compression performance assessment of ABI SOLiD WES data. Algorithms (Basel), 6(2): 309–318, June 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/6/2/309.

Abeliuk:2013:PCS

[175] Andrés Abeliuk, Rodrigo Cánovas, and Gonzalo Navarro. Practical compressed suffix trees. Algorithms (Basel), 6(2):319–351, June 2013. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/6/2/319.

Comin:2013:FDP

[176] Matteo Comin and Davide Verzotto. Filtering degenerate patterns with application to protein sequence analysis. Algorithms (Basel), 6(2):352– 370, June 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/6/2/352.

Inoshita:2013:IMO

[177] Takao Inoshita, Robert W. Irving, Kazuo Iwama, Shuichi Miyazaki, and Takashi Nagase. Improving man-optimal stable matchings by minimum change of preference lists. Algorithms (Basel), 6(2):371–382, June 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/2/371.

Cheng:2013:MLS

[178] Christine T. Cheng and Eric McDermid. Maximum locally stable matchings. Algorithms (Basel), 6(3):383–395, September 2013. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/3/383.

Tazehkand:2013:NHR

[179] Soheil Jahangiri Tazehkand, Seyed Naser Hashemi, and Hadi Poormohammadi. New heuristics for rooted triplet consistency. Algorithms (Basel), 6(3):396–406, September 2013. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/6/3/396.

Premanode:2013:NRN

[180] Bhusana Premanode, Jumlong Vongprasert, and Christofer Toumazou. Noise reduction for nonlinear nonstationary time series data using averaging intrinsic mode function. Algorithms (Basel), 6(3):407–429, September 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/6/3/407.

Haiminen:2013:ESC

[181] Niina Haiminen, Filippo Utro, Claude Lebreton, Pascal Flament, Zivan Karaman, and Laxmi Parida. Efficient in silico chromosomal representation of populations via indexing ancestral genomes. Algorithms (Basel), 6(3):430–441, September 2013. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/6/3/430.

Kambayashi:2013:RRP

[182] Yasushi Kambayashi. A review of routing protocols based on ant-like mobile agents. Algorithms (Basel), 6(3):442–456, September 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/3/442.

Jansson:2013:SIG

[183] Jesper Jansson. Special issue on graph algorithms. Algorithms (Basel), 6(3):457–458, September 2013. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/6/3/457.

Hsu:2013:UIN

[184] Li-Yen Hsu. Ubiquitous integrity via network integration and parallelism — sustaining pedestrian/bike urbanism. Algorithms (Basel), 6(3):459– 470, September 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/6/3/459.

Kiraly:2013:LTL

[185] Zoltán Király. Linear time local approximation algorithm for maximum stable marriage. Algorithms (Basel), 6(3):471–484, September 2013. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/6/3/471.

Zhu:2013:SAS

[186] Daxin Zhu and Xiaodong Wang. A simple algorithm for solving for the generalized longest common subsequence (LCS) problem with a substring exclusion constraint. Algorithms (Basel), 6(3):485–493, September 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/3/485.

Tancredi:2013:AMA

[187] Urbano Tancredi, Domenico Accardo, Giancarmine Fasano, Alfredo Renga, Giancarlo Rufino, and Giuseppe Maresca. An algorithm for managing aircraft movement on an airport surface. Algorithms (Basel), 6(3): 494–511, September 2013. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/6/3/494.

Taher:2013:ESS

[188] Fatma Taher, Naoufel Werghi, Hussain Al-Ahmad, and Christian Donner. Extraction and segmentation of sputum cells for lung cancer early diagnosis. Algorithms (Basel), 6(3):512–531, September 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/3/512.

Cseh:2013:SFT

[189] Ágnes Cseh, Jannik Matuschke, and Martin Skutella. Stable flows over time. Algorithms (Basel), 6(3):532–545, September 2013. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/6/3/532.

Ghaffarizadeh:2013:QTL

[190] Ahmadreza Ghaffarizadeh, Mehdi Eftekhari, Ali K. Esmailizadeh, and Nicholas S. Flann. Quantitative trait loci mapping problem: An extinction-based multi-objective evolutionary algorithm approach. Algorithms (Basel), 6(3):546–564, September 2013. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
6/3/546.

Cimino:2013:EAT

[191] Mario G. C. A. Cimino and Gigliola Vaglini. An emergent approach to text analysis based on a connectionist model and the web. Algorithms (Basel), 6(3):565–590, September 2013. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/6/3/565.

Gelain:2013:LSA

[192] Mirco Gelain, Maria Silvia Pini, Francesca Rossi, K. Brent Venable, and Toby Walsh. Local search approaches in stable matching problems. Algorithms (Basel), 6(4):591–617, December 2013. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/6/4/
591.

McCreesh:2013:MTS

[193] Ciaran McCreesh and Patrick Prosser. Multi-threading a state-of-the-art maximum clique algorithm. Algorithms (Basel), 6(4):618–635, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/6/4/618.

Fu:2013:STM

[194] Bin Fu, Yunhui Fu, and Yuan Xue. Sublinear time motif discovery from multiple sequences. Algorithms (Basel), 6(4):636–677, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/4/636.

Bredereck:2013:PGA

[195] Robert Bredereck, André Nichterlein, and Rolf Niedermeier. Patternguided k-anonymity. Algorithms (Basel), 6(4):678–701, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/4/678.

Hogg:2013:NPS

[196] Jonathan Hogg and Jennifer Scott. New parallel sparse direct solvers for multicore architectures. Algorithms (Basel), 6(4):702–725, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/6/4/702.

Zhang:2013:ELS

[197] Zhiqiang Zhang, Ansheng Ye, Xiaoqing Zhou, and Zehui Shao. An efficient local search for the feedback vertex set problem. Algorithms (Basel), 6(4):726–746, December 2013. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/6/4/726.

Sicard:2013:MCP

[198] Nicolas Sicard, Yogi Satrya Aryadinata, Federico Del Razo Lopez, Anne Laurent, and Perfecto Malaquias Quintero Flores. Multi-core parallel gradual pattern mining based on multi-precision fuzzy orderings. Algorithms (Basel), 6(4):747–761, December 2013. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/6/4/
747.

Jabari:2013:VHR

[199] Shabnam Jabari and Yun Zhang. Very high resolution satellite image classification using fuzzy rule-based systems. Algorithms (Basel), 6(4): 762–781, December 2013. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/6/4/762.

Pini:2013:SOM

[200] Maria Silvia Pini, Francesca Rossi, K. Brent Venable, and Toby Walsh. Stability, optimality and manipulation in matching problems with weighted preferences. Algorithms (Basel), 6(4):782–804, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/4/782.

Bandyopadhyay:2013:PMA

[201] Shibdas Bandyopadhyay, Sartaj Sahni, and Sanguthevar Rajasekaran. PMS6MC: a multicore algorithm for motif discovery. Algorithms (Basel), 6(4):805–823, December 2013. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/6/4/805.

Georgiadis:2013:OPD

[202] Giorgos Georgiadis and Marina Papatriantafilou. Overlays with preferences: Distributed, adaptive approximation algorithms for matching with preference lists. Algorithms (Basel), 6(4):824–856, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/4/824.

Benner:2013:SME

[203] Peter Benner, Pablo Ezzatti, Hermann Mena, Enrique S. Quintana-Ortí, and Alfredo Remón. Solving matrix equations on multi-core and manycore architectures. Algorithms (Basel), 6(4):857–870, December 2013. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/6/4/857.

He:2013:SSR

[204] Jun He, Ming-Wei Gao, Lei Zhang, and Hao Wu. Sparse signal recovery from fixed low-rank subspace via compressive measurement. Algorithms (Basel), 6(4):871–882, December 2013. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/6/4/871.

Fleiner:2014:SMF

[205] Tamás Fleiner. On stable matchings and flows. Algorithms (Basel), 7 (1):1–14, March 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/7/1/1.

Zhang:2014:BIM

[206] Min-Xia Zhang, Bei Zhang, and Yu-Jun Zheng. Bio-inspired metaheuristics for emergency transportation problems. Algorithms (Basel), 7 (1):15–31, March 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/7/1/15.

Fleiner:2014:CFB

[207] Tamàs Fleiner and Zsuzsanna Jankó. Choice function-based two-sided markets: Stability, lattice property, path independence and algorithms. Algorithms (Basel), 7(1):32–59, March 2014. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/7/1/
32.

Office:2014:ARA

[208] Algorithms Editorial Office. Acknowledgement to reviewers of algorithms in 2013. Algorithms (Basel), 7(1):60–61, March 2014. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/1/60.

Giegerich:2014:MDP

[209] Robert Giegerich and Héléne Touzet. Modeling dynamic programming problems over sequences and trees with inverse coupled rewrite systems. Algorithms (Basel), 7(1):62–144, March 2014. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/7/1/
62.

Jung:2014:MST

[210] Changyong Jung, Suk Jin Lee, and Vijay Bhuse. The minimum scheduling time for convergecast in wireless sensor networks. Algorithms (Basel), 7(1):145–165, March 2014. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/7/1/145.

Jakob:2014:POC

[211] Wilfried Jakob and Christian Blume. Pareto optimization or cascaded weighted sum: A comparison of concepts. Algorithms (Basel), 7(1):166– 185, March 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/7/1/166. See correction [213].

Makinen:2014:ESI

[212] Veli Mäkinen. Editorial: Special issue on algorithms for sequence analysis and storage. Algorithms (Basel), 7(1):186–187, March 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/1/186.

Iwama:2014:CPO

[213] Kazuo Iwama. Correction: Pareto Optimization or Cascaded Weighted Sum: a Comparison of Concepts. Algorithms 2014, 7, 166–185. Algorithms (Basel), 7(2):188, June 2014. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/7/2/188.

See [211].

Paluch:2014:FSA

[214] Katarzyna Paluch. Faster and simpler approximation of stable matchings. Algorithms (Basel), 7(2):189–202, June 2014. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
7/2/189.

Biro:2014:ESI

[215] Péter Biró and David F. Manlove. Editorial: Special issue on matching under preferences. Algorithms (Basel), 7(2):203–205, June 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/2/203.

Williams:2014:SDS

[216] Howard Williams and Mark Bishop. Stochastic diffusion search: a comparison of swarm intelligence parameter estimation algorithms with RANSAC. Algorithms (Basel), 7(2):206–228, June 2014. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/2/206.

Xu:2014:AIC

[217] Shuhui Xu, Yong Wang, and Aiqin Huang. Application of imperialist competitive algorithm on solving the traveling salesman problem. Algorithms (Basel), 7(2):229–242, June 2014. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/7/2/
229.

Zechner:2014:EAS

[218] Niklas Zechner and Andrzej Lingas. Efficient algorithms for subgraph listing. Algorithms (Basel), 7(2):243–252, June 2014. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
7/2/243.

Lin:2014:FQS

[219] Jie Lin, Donald Adjeroh, and Yue Jiang. A faster quick search algorithm. Algorithms (Basel), 7(2):253–275, June 2014. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/7/2/
253.

Burns:2014:GSR

[220] Brian L. Burns, Neil E. Wilson, and M. Albert Thomas. Group sparse reconstruction of multi-dimensional spectroscopic imaging in human brain in vivo. Algorithms (Basel), 7(3):276–294, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/3/276.

Kolonias:2014:SET

[221] Vasileios Kolonias, George Goulas, Christos Gogos, Panayiotis Alefragis, and Efthymios Housos. Solving the examination timetabling problem in GPUs. Algorithms (Basel), 7(3):295–327, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/3/295.

Latif:2014:EDU

[222] Aadil Latif and Peter Palensky. Economic dispatch using modified bat algorithm. Algorithms (Basel), 7(3):328–338, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/3/328.

Santone:2014:MCP

[223] Antonella Santone and Gigliola Vaglini. Model checking properties on reduced trace systems. Algorithms (Basel), 7(3):339–362, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/7/3/339.

Aiqin:2014:PMC

[224] Huang Aiqin and Wang Yong. Pressure model of control valve based on LS-SVM with the fruit fly algorithm. Algorithms (Basel), 7(3):363–375, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/7/3/363.

Costa:2014:HMA

[225] Antonio Costa, Fulvio Antonio Cappadonna, and Sergio Fichera. A hybrid metaheuristic approach for minimizing the total flow time in a flow shop sequence dependent group scheduling problem. Algorithms (Basel), 7(3):376–396, September 2014. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/7/3/376.

Ma:2014:ABH

[226] Wen-Qing Ma and Jing Zhang. Algorithm based on heuristic strategy to infer lossy links in wireless sensor networks. Algorithms (Basel), 7(3):397– 404, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/7/3/397.

Wang:2014:SQP

[227] Hong Wang, Qingsong Xu, and Lifeng Zhou. Seminal quality prediction using clustering-based decision forests. Algorithms (Basel), 7(3):405–417, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/7/3/405.

Zheng:2014:TCV

[228] Shilian Zheng, Zhijin Zhao, Changlin Luo, and Xiaoniu Yang. Target channel visiting order design using particle swarm optimization for spectrum handoff in cognitive radio networks. Algorithms (Basel), 7(3):418– 428, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/7/3/418.

Deng:2014:MCD

[229] Zhibin Deng, John E. Lavery, Shu-Cherng Fang, and Jian Luo. ℓ_{1} major component detection and analysis (ℓ_{1} MCDA) in three and higher dimensional spaces. Algorithms (Basel), 7(3):429–443, September 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/7/3/429.

Chang:2014:NCE

[230] Yung-Tsang Chang, Jen-Tse Wang, and Wang-Hsai Yang. A novel contrast enhancement technique on palm bone images. Algorithms (Basel), 7(3):444–455, September 2014. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/7/3/444.

Yu:2014:FLS

[231] Chun-Yuan Yu, Chen-Chung Liu, and Shyr-Shen Yu. A fovea localization scheme using vessel origin-based parabolic model. Algorithms (Basel), 7(3):456–470, September 2014. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/7/3/456.

Wan:2014:ADR

[232] Kaiyu Wan, Yuji Dong, Qian Chang, and Tengfei Qian. Applying a dynamic resource supply model in a smart grid. Algorithms (Basel), 7(3):471–491, September 2014. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/7/3/471.

Qin:2014:PCF

[233] Xuebin Qin, Mei Wang, Jzau-Sheng Lin, and Xiaowei Li. Power cable fault recognition based on an annealed chaotic competitive learning network. Algorithms (Basel), 7(4):492–509, December 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/7/4/492.

He:2014:CSA

[234] Wei He, Ke Li, Qiang Zhou, and Songnong Li. A CR spectrum allocation algorithm in smart grid wireless sensor network. Algorithms (Basel), 7 (4):510–522, December 2014. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/7/4/510.

Ding:2014:MSB

[235] Qian Ding, Zhenghong Peng, Tianzhen Liu, and Qiaohui Tong. Multisensor building fire alarm system with information fusion technology based on D-S evidence theory. Algorithms (Basel), 7(4):523–537, December 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/7/4/523.

Chen:2014:PSA

[236] Jeng-Fung Chen, Ho-Nien Hsieh, and Quang Hung Do. Predicting student academic performance: a comparison of two meta-heuristic algorithms inspired by cuckoo birds for training neural networks. Algorithms (Basel), 7(4):538–553, December 2014. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/7/4/538.

Rabanal:2014:PPS

[237] Pablo Rabanal, Ismael Rodríguez, and Fernando Rubio. Parallelizing particle swarm optimization in a functional programming environment. Algorithms (Basel), 7(4):554–581, December 2014. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
7/4/554.

Huang:2014:PNQ

[238] Yuan-Ko Huang. Processing K NN queries in grid-based sensor networks. Algorithms (Basel), 7(4):582–596, December 2014. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
7/4/582.

Yu:2014:ESI

[239] Cheng-Yi Yu, Chi-Yuan Lin, Sheng-Chih Yang, and Hsueh-Yi Lin. Eightscale image contrast enhancement based on adaptive inverse hyperbolic tangent algorithm. Algorithms (Basel), 7(4):597–607, December 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/7/4/597.

Beskers:2014:HOE

[240] Kai Beskers and Johannes Fischer. High-order entropy compressed bit vectors with rank/select. Algorithms (Basel), 7(4):608–620, December 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/7/4/608.

Cecchini:2014:NNM

[241] Giulio Cecchini, Gabriele Maria Lozito, Maurizio Schmid, Silvia Conforto, Francesco Riganti Fulginei, and Daniele Bibbo. Neural networks for muscle forces prediction in cycling. Algorithms (Basel), 7(4):621–634, December 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/7/4/621.

Chen:2014:LFT

[242] Chii-Jen Chen, You-Wei Wang, Wei-Chih Shen, Chih-Yi Chen, and Wen-Pinn Fang. The lobe fissure tracking by the modified ant colony optimization framework in CT images. Algorithms (Basel), 7(4):635–649, December 2014. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/7/4/635.

Cui:2014:TSP

[243] Wen-Hua Cui, Jie-Sheng Wang, and Chen-Xu Ning. Time series prediction method of bank cash flow and simulation comparison. Algorithms (Basel), 7(4):650–662, December 2014. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/7/4/650.

Xu:2014:CNO

[244] Qingzheng Xu, Lemeng Guo, Na Wang, and Yongjian He. COOBBO: a novel opposition-based soft computing algorithm for TSP problems. Algorithms (Basel), 7(4):663–684, December 2014. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
7/4/663.

Zhou:2014:FMP

[245] Wanchun Zhou, Ji Xiong, Fangmin Li, Na Jiang, and Ning Zhao. Fusion of multiple pyroelectric characteristics for human body identification. Algorithms (Basel), 7(4):685–702, December 2014. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
7/4/685.

Office:2015:ARA

[246] Algorithms Editorial Office. Acknowledgement to reviewers of algorithms in 2014. Algorithms (Basel), 8(1):1–2, March 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/1/1.

Tian:2015:NCS

[247] Zhong-Da Tian, Shu-Jiang Li, Yan-Hong Wang, and Hong-Xia Yu. Networked control system time-delay compensation based on time-delay prediction and improved implicit GPC. Algorithms (Basel), 8(1):3–18, March 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/1/3.

Lu:2015:ISF

[248] Kong Lu, Li Ting, Wang Keming, Zhu Hanbing, Takano Makoto, and Yu Bin. An improved shuffled frog-leaping algorithm for flexible job shop scheduling problem. Algorithms (Basel), 8(1):19–31, March 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/8/1/19.

Yang:2015:ESI

[249] Dongdong Yang, Hui Yang, and Rong Fei. An efficient SAR image segmentation framework using transformed nonlocal mean and multiobjective clustering in kernel space. Algorithms (Basel), 8(1):32–45, March 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/1/32.

McDiarmid:2015:RUG

[250] Colin McDiarmid and Nikola Yolov. Recognition of unipolar and generalised split graphs. Algorithms (Basel), 8(1):46–59, March 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/1/46.

Huffner:2015:PCR

[251] Falk Hüffner, Christian Komusiewicz, Rolf Niedermeier, and Martin Rötzschke. The parameterized complexity of the rainbow subgraph problem. Algorithms (Basel), 8(1):60–81, March 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/1/60.

Anniballe:2015:SGF

[252] Roberta Anniballe and Stefania Bonafoni. A stable Gaussian fitting procedure for the parameterization of remote sensed thermal images. Algorithms (Basel), 8(2):82–91, June 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/2/
82.

Mezei:2015:SPP

[253] Mihaly Mezei. Statistical properties of protein-protein interfaces. Algorithms (Basel), 8(2):92–99, June 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/2/92.

Shiau:2015:SFL

[254] Jaw-Kuen Shiau, Yu-Chen Wei, and Bo-Chih Chen. A study on the fuzzy-logic-based solar power MPPT algorithms using different fuzzy input variables. Algorithms (Basel), 8(2):100–127, June 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/2/100.

Zhou:2015:CAB

[255] Yuan Zhou, Hong fu Zuo, and Jiao Feng. A clustering algorithm based on feature weighting fuzzy compactness and separation. Algorithms (Basel), 8(2):128–143, June 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/2/128.

Ren:2015:APP

[256] Yaming Ren and Shumin Fei. The auxiliary problem principle with selfadaptive penalty parameter for multi-area economic dispatch problem. Algorithms (Basel), 8(2):144–156, June 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/2/
144.

Li:2015:MCP

[257] Wei Li, Lei Wang, Qiaoyong Jiang, Xinhong Hei, and Bin Wang. Multiobjective cloud particle optimization algorithm based on decomposition. Algorithms (Basel), 8(2):157–176, June 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/2/

He:2015:ASC

[258] Xiaoqi He, Sheng Zhang, and Yangguang Liu. An adaptive spectral clustering algorithm based on the importance of shared nearest neighbors. Algorithms (Basel), 8(2):177–189, June 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/2/
177.

Kuo:2015:EGL

[259] Ting Kuo. From enumerating to generating: a linear time algorithm for generating 2D lattice paths with a given number of turns. Algorithms (Basel), 8(2):190–208, June 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/2/190.

Yanagihara:2015:PND

[260] Takanobu Yanagihara and Hotaka Takizawa. Pulmonary nodule detection from X-ray CT images based on region shape analysis and appearance-based clustering. Algorithms (Basel), 8(2):209–223, June 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/8/2/209.

Cordero:2015:NST

[261] Alicia Cordero, Antonio Franques, and Juan R. Torregrosa. Numerical solution of turbulence problems by solving Burgers’ equation. Algorithms (Basel), 8(2):224–233, June 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/2/224.

Wang:2015:OCA

[262] Gaihua Wang, Yang Liu, and Caiquan Xiong. An optimization clustering algorithm based on texture feature fusion for color image segmentation. Algorithms (Basel), 8(2):234–247, June 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/2/
234.

Nicolae:2015:SMM

[263] Marius Nicolae and Sanguthevar Rajasekaran. On string matching with mismatches. Algorithms (Basel), 8(2):248–270, June 2015. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/2/248.

Chicharro:2015:DFD

[264] Francisco I. Chicharro, Alicia Cordero, and Juan R. Torregrosa. Dynamics and fractal dimension of Steffensen-type methods. Algorithms (Basel), 8(2):271–279, June 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/2/271.

Meng:2015:MEB

[265] Dandan Meng and Feng Ding. Model equivalence-based identification algorithm for equation-error systems with colored noise. Algorithms (Basel), 8(2):280–291, June 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/2/280.

Chen:2015:TAN

[266] Jeng-Fung Chen, Quang Hung Do, and Ho-Nien Hsieh. Training artificial neural networks by a hybrid PSO-CS algorithm. Algorithms (Basel), 8(2): 292–308, June 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/8/2/292.

Kansal:2015:OEO

[267] Munish Kansal, Vinay Kanwar, and Saurabh Bhatia. An optimal eighthorder derivative-free family of Potra–Pták’s method. Algorithms (Basel), 8(2):309–320, June 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/2/309.

Li:2015:TDS

[268] Xu Li, Tao Zhang, and Yong Ou Zhang. Time domain simulation of sound waves using smoothed particle hydrodynamics algorithm with artificial viscosity. Algorithms (Basel), 8(2):321–335, June 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/2/321.

Khalil:2015:MNH

[269] Ahmed M. E. Khalil, Seif-Eddeen K. Fateen, and Adrián Bonilla-Petriciolet. MAKHA — a new hybrid swarm intelligence global optimization algorithm. Algorithms (Basel), 8(2):336–365, June 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/2/336.

Wang:2015:IDR

[270] Ying-Ying Wang, Xiang-Dong Wang, and Dong-Qing Wang. Identification of dual-rate sampled Hammerstein systems with a piecewise-linear nonlinearity using the key variable separation technique. Algorithms (Basel), 8(3):366–379, September 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/3/366.

Li:2015:ICP

[271] Yefeng Li, Jiajin Le, and Mei Wang. Improving CLOPE’s profit value and stability with an optimized agglomerative approach. Algorithms (Basel), 8(3):380–394, September 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/3/380.

Maeda:2015:ACF

[272] Kazuo Maeda, Yasuaki Noguchi, Masaji Utsu, and Takashi Nagassawa. Algorithms for computerized fetal heart rate diagnosis with direct reporting. Algorithms (Basel), 8(3):395–406, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/395.

Wang:2015:IPA

[273] Longhui Wang, Yong Wang, and Yudong Xie. Implementation of a parallel algorithm based on a spark cloud computing platform. Algorithms (Basel), 8(3):407–414, September 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/3/407.

Ghorbanzadeh:2015:QCJ

[274] Mohammad Ghorbanzadeh and Fazlollah Soleymani. A quartically convergent Jarratt-type method for nonlinear system of equations. Algorithms (Basel), 8(3):415–423, September 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/3/
415.

Ni:2015:MOD

[275] Yuan Ni, Zongquan Deng, Junbao Li, Xiang Wu, and Long Li. Multiobjective design optimization of an over-constrained flexure-based amplifier. Algorithms (Basel), 8(3):424–434, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/424.

Meghanathan:2015:BAD

[276] Natarajan Meghanathan. A benchmarking algorithm to determine minimum aggregation delay for data gathering trees and an analysis of the diameter-aggregation delay tradeoff. Algorithms (Basel), 8(3):435–458, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/3/435.

Parberry:2015:SPE

[277] Ian Parberry. Solving the (n^{2} − 1)-puzzle with (8/3)pn^{3} expected moves. Algorithms (Basel), 8(3):459–465, September 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/3/459.

Burr:2015:CRF

[278] Tom Burr and Alexei Skurikhin. Conditional random fields for pattern recognition applied to structured data. Algorithms (Basel), 8(3):466–483, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/3/466.

Li:2015:MFI

[279] Peng Li, Min Chen, Li (Alex) Li, and Jiao Feng. Multi-feedback interference cancellation algorithms for OFDM systems over doubly-selective channels. Algorithms (Basel), 8(3):484–513, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/484.

Ezquerro:2015:ANM

[280] José Antonio Ezquerro and Miguel Ángel Hernández-Verón. On the accessibility of Newton’s method under a Hölder condition on the first derivative. Algorithms (Basel), 8(3):514–528, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/514.

Shi:2015:PSS

[281] Zhen Shi, Chang’an Wei, Ping Fu, and Shouda Jiang. A parallel search strategy based on sparse representation for infrared target tracking. Algorithms (Basel), 8(3):529–540, September 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/3/529.

Cui:2015:TDA

[282] Zheng Cui, Jingli Yang, Shouda Jiang, and Changan Wei. Target detection algorithm based on two layers human visual system. Algorithms (Basel), 8(3):541–551, September 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/3/541.

Babajee:2015:SIT

[283] Diyashvir Kreetee Rajiv Babajee. Some improvements to a third order variant of Newton’s method from Simpson’s rule. Algorithms (Basel), 8(3):552–561, September 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/3/552.

Wang:2015:MDE

[284] Longhui Wang, Guoguang Zhao, and Donghong Sun. Modeling documents with event model. Algorithms (Basel), 8(3):562–572, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/8/3/562.

Li:2015:RRR

[285] Peng Li, Jiao Feng, and Rodrigo C. De Lamare. Robust rank reduction algorithm with iterative parameter optimization and vector perturbation. Algorithms (Basel), 8(3):573–589, September 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/3/573.

Wang:2015:OCC

[286] Xu Wang, Daniel Jeske, and Erchin Serpedin. An overview of a class of clock synchronization algorithms for wireless sensor networks: a statistical signal processing perspective. Algorithms (Basel), 8(3):590–620, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/3/590.

Gao:2015:OBQ

[287] Fei Gao, Lili Guo, Hongbin Li, and Jun Fang. One-bit quantization and distributed detection with an unknown scale parameter. Algorithms (Basel), 8(3):621–631, September 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/3/621.

Miao:2015:DFM

[288] Zhibin Miao and Hongtian Zhang. Data fusion modeling for an RT3102 and Dewetron system application in hybrid vehicle stability testing. Algorithms (Basel), 8(3):632–644, September 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/3/632.

Argyros:2015:LCO

[289] Ioannis K. Argyros, Ramandeep Behl, and S. S. Motsa. Local convergence of an optimal eighth order method under weak conditions. Algorithms (Basel), 8(3):645–655, September 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/3/
645.

Liang:2015:FOI

[290] Juan Liang, Xiaowu Li, Zhinan Wu, Mingsheng Zhang, Lin Wang, and Feng Pan. Fifth-order iterative method for solving multiple roots of the highest multiplicity of nonlinear equation. Algorithms (Basel), 8(3):656– 668, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/8/3/656.

Amat:2015:EAT

[291] Sergio Amat, Sonia Busquier, Concepción Bermúdez, and Ángel Alberto Magreñán. Expanding the applicability of a third order Newton-type method free of bilinear operators. Algorithms (Basel), 8(3):669–679, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/3/669.

Saha:2015:NCD

[292] Suman Saha and Satya P. Ghrera. Network community detection on metric space. Algorithms (Basel), 8(3):680–696, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/680.

Ouyang:2015:CSP

[293] Puren Ouyang and Vangjel Pano. Comparative study of DE, PSO and GA for position domain PID controller tuning. Algorithms (Basel), 8(3): 697–711, September 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/3/697.

Zhou:2015:GBI

[294] Lincheng Zhou, Xiangli Li, Huigang Xu, and Peiyi Zhu. Gradient-based iterative identification for Wiener nonlinear dynamic systems with moving average noises. Algorithms (Basel), 8(3):712–722, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/8/3/712.

Katsaragakis:2015:CSM

[295] Iosif V. Katsaragakis, Ioannis X. Tassopoulos, and Grigorios N. Beligiannis. A comparative study of modern heuristics on the school timetabling problem. Algorithms (Basel), 8(3):723–742, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/723.

Liu:2015:CRA

[296] Yanjun Liu and Taiyang Tao. A CS recovery algorithm for model and time delay identification of MISO–FIR systems. Algorithms (Basel), 8 (3):743–753, September 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/3/743.

Mallen-Fullerton:2015:MCG

[297] Guillermo M. Mallén-Fullerton, J. Emilio Quiroz-Ibarra, Antonio Miranda, and Guillermo Fernández-Anaya. Modified classical graph algorithms for the DNA fragment assembly problem. Algorithms (Basel), 8 (3):754–773, September 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/3/754.

Bistran:2015:PVB

[298] Ioan Bistran, Stefan Maruster, and Liviu Octavian Mafteiu-Scai. Parallel variants of Broyden’s method. Algorithms (Basel), 8(3):774–785, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/3/774.

Wang:2015:FNT

[299] Xiaofeng Wang, Yuping Qin, Weiyi Qian, Sheng Zhang, and Xiaodong Fan. A family of Newton type iterative methods for solving nonlinear equations. Algorithms (Basel), 8(3):786–798, September 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/3/786.

Zhang:2015:RFS

[300] Tingping Zhang and Guan Gui. Reweighted factor selection for SLMS-RL1 algorithm under Gaussian mixture noise environments. Algorithms (Basel), 8(4):799–809, December 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/4/799.

Fertin:2015:FSP

[301] Guillaume Fertin, Christian Komusiewicz, Hafedh Mohamed-Babou, and Irena Rusu. Finding supported paths in heterogeneous networks. Algorithms (Basel), 8(4):810–831, December 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/4/
810.

Anastassiou:2015:NTM

[302] George A. Anastassiou and Ioannis K. Argyros. Newton-type methods on generalized Banach spaces and applications in fractional calculus. Algorithms (Basel), 8(4):832–849, December 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/4/
832.

Andonov:2015:ACP

[303] Rumen Andonov, Hristo Djidjev, Gunnar W. Klau, Mathilde Le Boudic-Jamin, and Inken Wohlers. Automatic classification of protein structure using the maximum contact map overlap metric. Algorithms (Basel), 8 (4):850–869, December 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/4/850.

Li:2015:CSA

[304] Aohan Li, Ziheng Yang, Renji Qi, Feng Zhou, and Guangjie Han. Code synchronization algorithm based on segment correlation in spread spectrum communication. Algorithms (Basel), 8(4):870–894, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/8/4/870.

Babajee:2015:SIH

[305] Diyashvir Kreetee Rajiv Babajee, Kalyanasundaram Madhu, and Jayakumar Jayaraman. On some improved harmonic mean Newton-like methods for solving systems of nonlinear equations. Algorithms (Basel), 8(4):895–909, December 2015. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/8/4/895.

Khedr:2015:EDA

[306] Ahmed M. Khedr. Effective data acquisition protocol for multi-hop heterogeneous wireless sensor networks using compressive sensing. Algorithms (Basel), 8(4):910–928, December 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/4/
910.

Yang:2015:SAF

[307] Kai Yang, Rencheng Zhang, Shouhong Chen, Fujiang Zhang, Jianhong Yang, and Xingbin Zhang. Series arc fault detection algorithm based on autoregressive bispectrum analysis. Algorithms (Basel), 8(4):929–950, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/4/929.

Li:2015:NSI

[308] Zhi-Yong Li, Jiao-Hong Yi, and Gai-Ge Wang. A new swarm intelligence approach for clustering based on krill herd with elitism strategy. Algorithms (Basel), 8(4):951–964, December 2015. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/8/4/
951.

Zhang:2015:PFT

[309] Chaozhu Zhang, Lin Li, and Yu Wang. A particle filter track-before-detect algorithm based on hybrid differential evolution. Algorithms (Basel), 8(4):965–981, December 2015. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/8/4/965.

Soleimani:2015:SMI

[310] Farahnaz Soleimani, Predrag S. Stanimirovíč, and Fazlollah Soleymani. Some matrix iterations for computing generalized inverses and balancing chemical equations. Algorithms (Basel), 8(4):982–998, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/8/4/982.

Catanzaro:2015:ILP

[311] Daniele Catanzaro and Céline Engelbeen. An integer linear programming formulation for the minimum cardinality segmentation problem. Algorithms (Basel), 8(4):999–1020, December 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/4/999.

Feng:2015:SSC

[312] Lei Feng and Guoxian Yu. Semi-supervised classification based on mixture graph. Algorithms (Basel), 8(4):1021–1034, December 2015. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/8/4/1021.

El-Kebir:2015:NSG

[313] Mohammed El-Kebir, Jaap Heringa, and Gunnar W. Klau. Natalie 2.0: Sparse global network alignment as a special case of quadratic assignment. Algorithms (Basel), 8(4):1035–1051, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/4/1035.

Metz:2015:NAA

[314] Andreas Jaakko Metz, Martin Wolf, Peter Achermann, and Felix Scholkmann. A new approach for automatic removal of movement artifacts in near-infrared spectroscopy time series by means of acceleration data. Algorithms (Basel), 8(4):1052–1075, December 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/4/1052.

Argyros:2015:LCE

[315] Ioannis K. Argyros, Ramandeep Behl, and S. S. Motsa. Local convergence of an efficient high convergence order method using hypothesis only on the first derivative. Algorithms (Basel), 8(4):1076–1087, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/8/4/1076.

Taher:2015:CAD

[316] Fatma Taher, Naoufel Werghi, and Hussain Al-Ahmad. Computer aided diagnosis system for early lung cancer detection. Algorithms (Basel), 8(4):1088–1110, December 2015. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/8/4/1088.

Zheng:2015:OBM

[317] Quan Zheng, Xin Zhao, and Yufeng Liu. An optimal biparametric multipoint family and its self-acceleration with memory for solving nonlinear equations. Algorithms (Basel), 8(4):1111–1120, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/4/1111.

Hernandez-Veron:2015:LCT

[318] M. A. Hernández-Verón and N. Romero. On the local convergence of a third order family of iterative processes. Algorithms (Basel), 8(4):1121– 1128, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/8/4/1121.

Zhang:2015:EIC

[319] Yong Zhang and Fangmin Li. Efficiency intra-cluster device-to-device relay selection for multicast services based on combinatorial auction. Algorithms (Basel), 8(4):1129–1142, December 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/4/1129.

Davis:2015:GRL

[320] Michael Charles Davis, Zhanyu Ma, Weiru Liu, Paul Miller, Ruth Hunter, and Frank Kee. Generating realistic labelled, weighted random graphs. Algorithms (Basel), 8(4):1143–1174, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/4/1143.

Jiang:2015:DAA

[321] Fan Jiang and Carson K. Leung. A data analytic algorithm for managing, querying, and processing uncertain big data in cloud environments. Algorithms (Basel), 8(4):1175–1194, December 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/4/1175.

Chu:2015:NSC

[322] Ajie Chu, Shouqiang Du, and Yixiao Su. A new smoothing conjugate gradient method for solving nonlinear nonsmooth complementarity problems. Algorithms (Basel), 8(4):1195–1209, December 2015. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/8/4/1195.

Gutierrez:2015:NPD

[323] José M. Gutiérrez. Numerical properties of different root-finding algorithms obtained for approximating continuous Newton’s method. Algorithms (Basel), 8(4):1210–1218, December 2015. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
8/4/1210.

Babajee:2016:KTC

[324] Diyashvir Kreetee Rajiv Babajee. On the Kung–Traub conjecture for iterative methods for solving quadratic equations. Algorithms (Basel), 9 (1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/1/1. See comment [353].

Ruiz-Oltra:2016:OAF

[325] José M. Ruiz-Oltra, Catalina Gómez-Quiles, and Antonio Gómez-Expósito. Offset-assisted factored solution of nonlinear systems. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/1/2.

Wang:2016:FOP

[326] Jie-Sheng Wang and Jiang-Di Song. Function optimization and parameter performance analysis based on gravitation search algorithm. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/1/3.

Luo:2016:NCV

[327] Qifang Luo, Sen Zhang, Zhiming Li, and Yongquan Zhou. A novel complex-valued encoding grey wolf optimization algorithm. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/9/1/4.

Ahmad:2016:FIM

[328] Fayyaz Ahmad, S. Serra-Capizzano, Malik Zaka Ullah, and A. S. Al-Fhaid. A family of iterative methods for solving systems of nonlinear equations having unknown multiplicity. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/1/5.

Gavalas:2016:EMM

[329] Damianos Gavalas, Charalampos Konstantopoulos, Konstantinos Mastakas, Grammati Pantziou, and Nikolaos Vathis. Efficient metaheuristics for the mixed team orienteering problem with time windows. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/9/1/6.

Cuzzocrea:2016:EEM

[330] Alfredo Cuzzocrea, Mirel Cosulschi, and Roberto De Virgilio. An effective and efficient MapReduce algorithm for computing BFS-based traversals of large-scale RDF graphs. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/1/7.

Meghanathan:2016:GAN

[331] Natarajan Meghanathan. A greedy algorithm for neighborhood overlapbased community detection. Algorithms (Basel), 9(1), March 2016. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/9/1/8.

Yang:2016:NAT

[332] Zhiming Yang, Junbao Li, Yang Yu, and Xiyuan Peng. NBTI-aware transient fault rate analysis method for logic circuit based on probability voltage transfer characteristics. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/1/9.

Jaiswal:2016:OOM

[333] Jai Prakash Jaiswal. An optimal order method for multiple roots in case of unknown multiplicity. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/1/10.

Office:2016:ARA

[334] Algorithms Editorial Office. Acknowledgement to reviewers of algorithms in 2015. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/9/1/
11.

Gatter:2016:IPO

[335] Thomas Gatter, Robert Giegerich, and Cédric Saule. Integrating Pareto optimization into dynamic programming. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/1/12.

Cuzzocrea:2016:AMQ

[336] Alfredo Cuzzocrea. Algorithms for managing, querying and processing big data in cloud environments. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/1/13.

Wang:2016:TED

[337] Xiaofeng Wang and Xiaodong Fan. Two efficient derivative-free iterative methods for solving nonlinear systems. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/1/14.

Li:2016:GOP

[338] Xiaowu Li, Zhinan Wu, Linke Hou, Lin Wang, Chunguang Yue, and Qiao Xin. A geometric orthogonal projection strategy for computing the minimum distance between a point and a spatial parametric curve. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/1/15.

Pizzolante:2016:MLC

[339] Raffaele Pizzolante and Bruno Carpentieri. Multiband and lossless compression of hyperspectral images. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/1/16.

Bulteau:2016:CCU

[340] Laurent Bulteau, Vincent Froese, Sepp Hartung, and Rolf Niedermeier. Co-clustering under the maximum norm. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/1/17.

Qasim:2016:CFJ

[341] Uswah Qasim, Zulifqar Ali, Fayyaz Ahmad, Stefano Serra-Capizzano, Malik Zaka Ullah, and Mir Asma. Constructing frozen Jacobian iterative methods for solving systems of nonlinear equations, associated with ODEs and PDEs using the homotopy method. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/1/18.

Gomez:2016:RRA

[342] María J. Gómez, Cristina Castejón, and Juan C. García-Prada. Review of recent advances in the application of the wavelet transform to diagnose cracked rotors. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/1/19.

Ivanov:2016:ISD

[343] Ivan G. Ivanov and Boryana C. Bogdanova. The iterative solution to discrete-time H_{∞} control problems for periodic systems. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/9/1/20.

Komusiewicz:2016:MAF

[344] Christian Komusiewicz. Multivariate algorithmics for finding cohesive subnetworks. Algorithms (Basel), 9(1), March 2016. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
9/1/21.

Kakoulis:2016:MOD

[345] Konstantinos G. Kakoulis and Ioannis G. Tollis. Modifying orthogonal drawings for label placement. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/2/22.

Sun:2016:IFA

[346] Yu-Feng Sun, Jie-Sheng Wang, and Jiang-Di Song. An improved fireworks algorithm based on grouping strategy of the shuffled frog leaping algorithm to solve function optimization problems. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/2/23.

Li:2016:SDL

[347] Cui-Hong Li, Qiu-Wei Yang, and Bing-Xiang Sun. Structural damage localization by the principal eigenvector of modal flexibility change. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/2/24.

Ye:2016:PUL

[348] Fang Ye, Xun Zhang, Yibing Li, and Hui Huang. Primary user localization algorithm based on compressive sensing in cognitive radio networks. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/2/25.

Takabatake:2016:SES

[349] Yoshimasa Takabatake, Kenta Nakashima, Tetsuji Kuboyama, Yasuo Tabei, and Hiroshi Sakamoto. siEDM: an efficient string index and search algorithm for edit distance with moves. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/2/26.

Ayedh:2016:EPA

[350] Abdullah Ayedh, Guanzheng Tan, Khaled Alwesabi, and Hamdi Rajeh. The effect of preprocessing on Arabic document categorization. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/2/27.

Shi:2016:ADM

[351] Jiarong Shi, Qingyan Yin, Xiuyun Zheng, and Wei Yang. Alternating direction method of multipliers for generalized low-rank tensor recovery. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/2/28.

Li:2016:IDJ

[352] Yibing Li, Xueying Diao, Ge Dong, and Fang Ye. An improved dynamic joint resource allocation algorithm based on SFR. Algorithms (Basel), 9 (2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/2/29.

Ahmad:2016:CKT

[353] Fayyaz Ahmad. Comment on: On the Kung–Traub Conjecture for Iterative Methods for Solving Quadratic Equations. Algorithms 2016, 9,

1. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/9/2/

30. See [324].

Zhao:2016:IDL

[354] Zhenqing Zhao, Dong Ye, Xin Zhang, Gang Chen, and Bin Zhang. Improved direct linear transformation for parameter decoupling in camera calibration. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
9/2/31.

Bjorklund:2016:UVN

[355] Henrik Björklund, Martin Berglund, and Petter Ericson. Uniform vs. nonuniform membership for mildly context-sensitive languages: a brief survey. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/9/2/
32.

Jeske:2016:MBR

[356] Daniel R. Jeske, Jeffrey Longmate, Vani Katheria, and Arti Hurria. Mining branching rules from past survey data with an illustration using a geriatric assessment survey for older adults with cancer. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/9/2/33.

Xu:2016:SRA

[357] Jing Xu, Zhongbin Wang, Chao Tan, and Xinhua Liu. A state recognition approach for complex equipment based on a fuzzy probabilistic neural network. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/9/2/
34.

Pan:2016:AEC

[358] Tianlin Pan, Bin Wu, Yongsheng Chen, and Guoshan Xu. Application of the energy-conserving integration method to hybrid simulation of a full-scale steel frame. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/2/35.

Xing:2016:RHL

[359] Xianglei Xing, Sidan Du, and Kejun Wang. Robust Hessian locally linear embedding techniques for high-dimensional data. Algorithms (Basel), 9 (2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/2/36.

Kumam:2016:NMS

[360] Wiyada Kumam, Pongsakorn Sunthrayuth, Phond Phunchongharn, Khajonpong Akkarajitsakul, Parinya Sa Ngiamsunthorn, and Poom Kumam. A new multi-step iterative algorithm for approximating common fixed points of a finite family of multi-valued Bregman relatively nonexpansive mappings. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/2/37.

Page:2016:AAG

[361] Daniel R. Page and Roberto Solis-Oba. A 3/2-approximation algorithm for the graph balancing problem with two weights. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/2/38.

Tai:2016:RRT

[362] Kevin Tai, Abdul-Rahman El-Sayed, Mohammad Biglarbegian, Claudia I. Gonzalez, Oscar Castillo, and Shohel Mahmud. Review of recent type-2 fuzzy controller applications. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/2/39.

Baeyens:2016:DSA

[363] Enrique Baeyens, Alberto Herreros, and José R. Perán. A direct search algorithm for global optimization. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/2/40.

Yu:2016:VTS

[364] Yuhai Yu, Hongfei Lin, Jiana Meng, and Zhehuan Zhao. Visual and textual sentiment analysis of a microblog using deep convolutional neural networks. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/9/2/
41.

Wang:2016:JAS

[365] Xinhua Wang and Jinlu Sheng. Joint antenna selection and beamforming algorithms for physical layer multicasting with massive antennas. Algorithms (Basel), 9(2), June 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/2/42.

Gong:2016:OBA

[366] Chibing Gong. Opposition-based adaptive fireworks algorithm. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/3/43.

Tantau:2016:GIA

[367] Till Tantau. A gentle introduction to applications of algorithmic metatheorems for space and circuit classes. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/3/44.

Tavassoli:2016:DFI

[368] Zahra Hojjati Tavassoli, Seyed Hossein Iranmanesh, and Ahmad Tavassoli Hojjati. Designing a framework to improve time series data of construction projects: Application of a simulation model and singular spectrum analysis. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/3/45.

Jiang:2016:APC

[369] Yuan Jiang, Yuliang Liao, and Guoxian Yu. Affinity propagation clustering using path based similarity. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/3/46.

Chang:2016:HCR

[370] Pei-Chann Chang, Cheng-Hui Lin, and Meng-Hui Chen. A hybrid course recommendation system by integrating collaborative filtering and artificial immune systems. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/3/47.

Hou:2016:SSC

[371] Xuan Hou, Guangjun Yao, and Jun Wang. Semi-supervised classification based on low rank representation. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/3/48.

Ding:2016:DFB

[372] Jiling Ding. Data filtering based recursive and iterative least squares algorithms for parameter estimation of multi-input output systems. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/3/49.

Askarian:2016:UNS

[373] Ahmad Askarian, Rupei Xu, and András Faragó. Utilizing network structure to accelerate Markov Chain Monte Carlo algorithms. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/3/50.

Cubukcuoglu:2016:MOH

[374] Cemre Cubukcuoglu, Ioannis Chatzikonstantinou, Mehmet Fatih Tasgetiren, I. Sevil Sariyildiz, and Quan-Ke Pan. A multi-objective harmony search algorithm for sustainable design of floating settlements. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/3/51.

Luo:2016:CSC

[375] Weilin Luo and Hongchao Cong. Control for ship course-keeping using optimized support vector machines. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/3/52.

Lipp:2016:FFD

[376] Fabian Lipp, Alexander Wolff, and Johannes Zink. Faster force-directed graph drawing with the well-separated pair decomposition. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/3/53.

Zhang:2016:SFB

[377] Tingping Zhang and Guan Gui. Sign function based sparse adaptive filtering algorithms for robust channel estimation under non-Gaussian noise environments. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/3/54.

Shepherd:2016:NAI

[378] Jonathan Bruce Shepherd, Tomohito Wada, David Rowlands, and Daniel Arthur James. A novel AHRS inertial sensor-based algorithm for wheelchair propulsion performance analysis. Algorithms (Basel), 9 (3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/9/3/55.

Huang:2016:MAN

[379] Xiang Huang and Zhizhong Wang. Multiple artificial neural networks with interaction noise for estimation of spatial categorical variables. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/3/56.

Khorramian:2016:UPM

[380] Amanj Khorramian and Akira Matsubayashi. Uniform page migration problem in Euclidean space. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/3/57.

Kallmeyer:2016:LPL

[381] Laura Kallmeyer and Wolfgang Maier. LR parsing for LCFRS. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/3/58.

Obagbuwa:2016:BCS

[382] Ibidun Christiana Obagbuwa and Ademola Philips Abidoye. Binary cockroach swarm optimization for combinatorial optimization problem. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/3/59.

San-Segundo:2016:HAI

[383] Rubén San-Segundo, Juan M. Montero, José Moreno-Pimentel, and José M. Pardo. HMM adaptation for improving a human activity recognition system. Algorithms (Basel), 9(3), September 2016. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/3/60.

Zhou:2016:NSB

[384] Weidong Zhou, Jing Liu, Pengxiang Zhu, Wenhe Gong, and Jiaxin Hou. Noncircular sources-based sparse representation algorithm for direction of arrival estimation in MIMO radar with mutual coupling. Algorithms (Basel), 9(3), September 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/3/61.

Ma:2016:NRS

[385] Hong-Wei Ma, Hong-Wei Fan, Qing-Hua Mao, Xu-Hui Zhang, and Wang Xing. Noise reduction of steel cord conveyor belt defect electromagnetic signal by combined use of improved wavelet and EMD. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/9/4/62.

Chen:2016:PRE

[386] Angela Hsiang-Ling Chen, Yun-Chia Liang, and Jose David Padilla. A practical and robust execution time-frame procedure for the multi-mode resource-constrained project scheduling problem with minimal and maximal time lags. Algorithms (Basel), 9(4), December 2016. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/4/63.

Fernau:2016:TGI

[387] Henning Fernau. Theorietage der Gesellschaft für Informatik in Speyer 2015 — special issue. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/4/64.

Argyros:2016:LCA

[388] Ioannis K. Argyros, Ramandeep Behl, and Sandile S. Motsa. Local convergence analysis of an eighth order scheme using hypothesis only on the first derivative. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/4/65.

Wang:2016:FSU

[389] Mei Wang, Liang Zhu, and Yanan Guo. Fault sensing using fractal dimension and wavelet. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/4/66.

Belfiore:2016:CDA

[390] Oscar Rosario Belfiore and Claudio Parente. Comparison of different algorithms to orthorectify WorldView-2 satellite imagery. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/4/67.

Wang:2016:OAM

[391] Xu Wang and Erchin Serpedin. An overview on the applications of matrix theory in wireless communications and signal processing. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/4/68.

Peraza:2016:NFH

[392] Cinthia Peraza, Fevrier Valdez, Mario Garcia, Patricia Melin, and Oscar Castillo. A new fuzzy harmony search algorithm using fuzzy logic for dynamic parameter adaptation. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/4/69.

Chen:2016:PES

[393] Yang Chen, Dong-Jie Zhao, Zi-Yang Wang, Zhong-Yi Wang, Guiliang Tang, and Lan Huang. Plant electrical signal classification based on waveform similarity. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/4/70.

Tasgetiren:2016:VBI

[394] Mehmet Fatih Tasgetiren, Quan-Ke Pan, Damla Kizilay, and Kaizhou Gao. A variable block insertion heuristic for the blocking flowshop scheduling problem with total flowtime criterion. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/9/4/71.

Hoske:2016:ECL

[395] Daniel Hoske, Dimitar Lukarski, Henning Meyerhenke, and Michael Wegner. Engineering a combinatorial Laplacian solver: Lessons learned. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/9/4/72.

Yang:2016:CSD

[396] Lingjian Yang, Jonathan C. Silva, Lazaros G. Papageorgiou, and Sophia Tsoka. Community structure detection for directed networks through modularity optimisation. Algorithms (Basel), 9(4), December 2016. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/9/4/73.

Liu:2016:OAD

[397] Hanbing Liu, Xin He, Xianqiang Wang, Yubo Jiao, and Gang Song. An optimization algorithm for the design of an irregularly-shaped bridge based on the orthogonal test and analytic hierarchy process. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/4/74.

Padcharoen:2016:MIA

[398] Anantachai Padcharoen, Poom Kumam, Yeol Je Cho, and Phatiphat Thounthong. A modified iterative algorithm for split feasibility problems of right Bregman strongly quasi-nonexpansive mappings in Banach spaces with applications. Algorithms (Basel), 9(4), December 2016. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/9/4/75.

Ramos:2016:PIA

[399] Patrícia Ramos and José Manuel Oliveira. A procedure for identification of appropriate state space and ARIMA models based on time-series cross-validation. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/9/4/76.

DeNiz:2016:ADS

[400] Carlos De Niz, Raziur Rahman, Xiangyuan Zhao, and Ranadip Pal. Algorithms for drug sensitivity prediction. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/4/77.

Li:2016:MCP

[401] Wei Li. A modified cloud particles differential evolution algorithm for real-parameter optimization. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/9/4/78.

You:2016:TSM

[402] Qingshan You, Yongjie Luo, and Qun Wan. A two-stage method to test the robustness of the generalized approximate message passing algorithm. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/9/4/
79.

Li:2016:STL

[403] Yuancheng Li, Panpan Guo, and Xiang Li. Short-term load forecasting based on the analysis of user electricity behavior. Algorithms (Basel), 9 (4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/9/4/80.

Ouyang:2016:CCC

[404] Puren Ouyang, Yuqi Hu, Wenhui Yue, and Deshun Liu. Cross-coupled contouring control of multi-DOF robotic manipulator. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/4/81.

Pedraza:2016:LAR

[405] Luis F. Pedraza, Cesar A. Hernandez, Ingrid P. Paez, Jorge E. Ortiz, and E. Rodriguez-Colina. Linear algorithms for radioelectric spectrum forecast. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
9/4/82.

Liu:2016:NLM

[406] Zhimin Liu, Shouqiang Du, and Ruiying Wang. Nonsmooth Levenberg– Marquardt type method for solving a class of stochastic linear complementarity problems with finitely many elements. Algorithms (Basel), 9 (4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/9/4/83.

Wibowo:2016:ECS

[407] Santoso Wibowo, Hepu Deng, and Wei Xu. Evaluation of cloud services: a fuzzy multi-criteria group decision making method. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/9/4/84.

Xie:2016:DAA

[408] Yuqin Xie and Mingchun Zheng. A differentiated anonymity algorithm for social network privacy preservation. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/4/85.

Gao:2016:MMS

[409] Qinjiao Gao and Shenggang Zhang. Moving mesh strategies of adaptive methods for solving nonlinear partial differential equations. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/9/4/86.

Fu:2016:NRI

[410] Yan Fu and Shengchun Wang. A no reference image quality assessment metric based on visual perception. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/9/4/87. See correction [460].

Zheng:2016:WWH

[411] Huanyang Zheng and Jie Wu. Which, when, and how: Hierarchical clustering with human–machine cooperation. Algorithms (Basel), 9(4), December 2016. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/9/4/88.

Mezei:2017:UFF

[412] Mihaly Mezei. Using force-field grids for sampling translation/rotation of partially rigid macromolecules. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/1/6.

Liu:2017:BBI

[413] Lingjun Liu, Zhonghua Xie, and Jiuchao Feng. Backtracking-based iterative regularization method for image compressive sensing recovery. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/1/7.

BenAbdallah:2017:MDD

[414] Emna Ben Abdallah, Tony Ribeiro, Morgan Magnin, Olivier Roux, and Katsumi Inoue. Modeling delayed dynamics in biological regulatory networks from time series data. Algorithms (Basel), 10(1), March 2017.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/1/8.

Zhao:2017:EOB

[415] Ruxin Zhao, Qifang Luo, and Yongquan Zhou. Elite opposition-based social spider optimization algorithm for global function optimization. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/1/9.

Maruster:2017:ELR

[416] Ştefan Măruşter. Estimating the local radius of convergence for Picard iteration. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/1/10.

Office:2017:ARA

[417] Algorithms Editorial Office. Acknowledgement to reviewers of algorithms in 2016. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/1/11.

Huang:2017:CLS

[418] Wu Huang and Feng Ding. Coupled least squares identification algorithms for multivariate output-error systems. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/1/12.

Mercorelli:2017:FDD

[419] Paolo Mercorelli. A fault detection and data reconciliation algorithm in technical processes with the help of Haar wavelets packets. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/1/13.

Feng:2017:KCD

[420] Yu Feng, Jianzhong Zhou, and Muhammad Tayyab. Kernel clustering with a differential harmony search algorithm for scheme classification. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/1/14.

Santos:2017:TPV

[421] Olga C. Santos. Toward personalized vibrotactile support when learning motor skills. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/1/15.

Lin:2017:LBH

[422] Yi-Shan Lin, Chun-Liang Lee, and Yaw-Chung Chen. Length-bounded hybrid CPU/GPU pattern matching algorithm for deep packet inspection. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/1/
16.

Ahmad:2017:PIM

[423] Fayyaz Ahmad, Toseef Akhter Bhutta, Umar Shoaib, Malik Zaka Ullah, Ali Saleh Alshomrani, Shamshad Ahmad, and Shahid Ahmad. A preconditioned iterative method for solving systems of nonlinear equations having unknown multiplicity. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/1/17. See erratum [455].

Bernal:2017:ICA

[424] Emer Bernal, Oscar Castillo, José Soria, and Fevrier Valdez. Imperialist competitive algorithm with dynamic parameter adaptation using fuzzy logic applied to the optimization of mathematical functions. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/1/18.

Wang:2017:PCH

[425] Ru Wang, Chao Tan, Jing Xu, Zhongbin Wang, Jingfei Jin, and Yiqiao Man. Pressure control for a hydraulic cylinder based on a self-tuning PID controller optimized by a hybrid optimization algorithm. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/1/19.

Berry:2017:CCT

[426] Anne Berry and Geneviève Simonet. Computing a clique tree with the algorithm maximal label search. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/1/20.

DeAgostino:2017:CVE

[427] Sergio De Agostino, Bruno Carpentieri, and Raffaele Pizzolante. Concurrent vs. exclusive reading in parallel decoding of LZ-compressed files. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/1/21.

Koniaris:2017:EDT

[428] Marios Koniaris, Ioannis Anagnostopoulos, and Yannis Vassiliou. Evaluation of diversification techniques for legal information retrieval. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/1/22.

Lee:2017:ABF

[429] James Lee, David Rowlands, Nicholas Jackson, Raymond Leadbetter, Tomohito Wada, and Daniel A. James. An architectural based framework for the distributed collection, analysis and query from inhomogeneous time series data sets and wearables for biofeedback applications. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/1/23.

Fernau:2017:PFA

[430] Henning Fernau and Andreas Krebs. Problems on finite automata and the exponential time hypothesis. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/1/24.

Chien:2017:LTS

[431] Tseng-Hsu Chien and Yeong-Chin Chen. An on-line tracker for a stochastic chaotic system using observer/Kalman filter identification combined with digital redesign method. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/1/25.

Li:2017:AIF

[432] Xi-Guang Li, Shou-Fei Han, and Chang-Qing Gong. Analysis and improvement of fireworks algorithm. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/1/26.

Zhang:2017:FWI

[433] Heng Zhang, Chengyou Wang, and Xiao Zhou. Fragile watermarking for image authentication using the characteristic of SVD. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/1/27.

Gkantouna:2017:MDS

[434] Vassiliki Gkantouna and Giannis Tzimas. Mining domain-specific design patterns: a case study. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/1/28.

You:2017:SAC

[435] Qingshan You and Qun Wan. Stable analysis of compressive principal component pursuit. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/1/29.

Prochazka:2017:TEP

[436] Petr Procházka and Jan Holub. Towards efficient positional inverted index. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/1/
30.

Kobayashi:2017:OBA

[437] Koichi Kobayashi and Kunihiko Hiraishi. Optimization-based approaches to control of probabilistic Boolean networks. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/1/31.

Lang:2017:NQS

[438] Feng-Gong Lang. A new quintic spline method for integro interpolation and its error analysis. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/1/32.

Kanavos:2017:LSI

[439] Andreas Kanavos, Nikolaos Nodarakis, Spyros Sioutas, Athanasios Tsakalidis, Dimitrios Tsolis, and Giannis Tzimas. Large scale implementations for Twitter sentiment classification. Algorithms (Basel), 10 (1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/1/33.

Athanasiou:2017:NGB

[440] Vasileios Athanasiou and Manolis Maragoudakis. A novel, gradient boosting framework for sentiment analysis in languages where NLP resources are not plentiful: a case study for modern Greek. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/1/34.

Spyrou:2017:GCA

[441] Evaggelos Spyrou, Michalis Korakakis, Vasileios Charalampidis, Apostolos Psallas, and Phivos Mylonas. A geo-clustering approach for the detection of areas-of-interest and their underlying semantics. Algorithms (Basel), 10(1), March 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/1/35.

Chen:2017:RIA

[442] Ziyan Chen, Yu Huang, Yuexian Liang, Yang Wang, Xingyu Fu, and Kun Fu. RGloVe: an improved approach of global vectors for distributional entity relation representation. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/2/42.

Feng:2017:RPS

[443] Yuan Feng, Li Wang, and Xinhong Liu. Reliable portfolio selection problem in fuzzy environment: an m_{λ} measure based approach. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/2/43.

Pelusi:2017:RGS

[444] Danilo Pelusi, Raffaele Mascella, and Luca Tallini. Revised gravitational search algorithms based on evolutionary-fuzzy systems. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/2/44.

Wang:2017:ESO

[445] Xiaofeng Wang and Yang Li. An efficient sixth-order Newton-type method for solving nonlinear systems. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/2/45.

Jia:2017:IMP

[446] Chunhua Jia and Hong Zhu. An improved multiobjective particle swarm optimization based on culture algorithms. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/2/46.

Hofman:2017:TBD

[447] Darra Hofman, Luciana Duranti, and Elissa How. Trust in the balance: Data protection laws as tools for privacy and security in the cloud. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/2/47.

Li:2017:AMD

[448] Xi-Guang Li, Shou-Fei Han, Liang Zhao, Chang-Qing Gong, and Xiao-Jing Liu. Adaptive mutation dynamic search fireworks algorithm. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/2/48.

Bouzenad:2017:MSP

[449] Khaled Bouzenad and Messaoud Ramdani. Multivariate statistical process control using enhanced bottleneck neural network. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/2/49.

Klein:2017:HPE

[450] Shmuel T. Klein and Dana Shapira. Hierarchical parallel evaluation of a Hamming code. Algorithms (Basel), 10(2), June 2017. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/2/50.

Pizzolante:2017:AVQ

[451] Raffaele Pizzolante, Bruno Carpentieri, and Sergio De Agostino. Adaptive vector quantization for lossy compression of image sequences. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/2/51.

Poh:2017:SDV

[452] Geong Sen Poh, Vishnu Monn Baskaran, Ji-Jian Chin, Moesfa Soeheila Mohamad, Kay Win Lee, Dharmadharshni Maniam, and Muhammad Reza Z’aba. Searchable data vault: Encrypted queries in secure distributed cloud storage. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/2/52.

Zhao:2017:AGD

[453] Huiru Zhao, Yuwei Wang, Mingrui Zhao, Chuyu Sun, and Qingkun Tan. Application of gradient descent continuous actor-critic algorithm for bilateral spot electricity market modeling considering renewable power penetration. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/2/53.

Argyros:2017:EAM

[454] Ioannis K. Argyros, Janak Raj Sharma, and Deepak Kumar. Extending the applicability of the MMN–HSS method for solving systems of nonlinear equations under generalized conditions. Algorithms (Basel), 10 (2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/2/54.

Ahmad:2017:EAF

[455] Fayyaz Ahmad, Toseef Akhter Bhutta, Umar Shoaib, Malik Zaka Ullah, Ali Saleh Alshomrani, Shamshad Ahmad, and Shahid Ahmad. Erratum: Ahmad, F., et al. A Preconditioned Iterative Method for Solving Systems of Nonlinear Equations Having Unknown Multiplicity. Algorithms 2017, 10, 17. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/2/

55. See [423].

Li:2017:CUI

[456] Qin Li and Bo Liu. Clustering using an improved krill herd algorithm. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/2/56.

Du:2017:PPD

[457] Jinglin Du, Yayun Liu, Yanan Yu, and Weilan Yan. A prediction of precipitation data based on support vector machine and particle swarm optimization (PSO-SVM) algorithms. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/2/57.

Lee:2017:FPM

[458] Chun-Liang Lee and Tzu-Hao Yang. A flexible pattern-matching algorithm for network intrusion detection systems using multi-core processors. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/2/
58.

Li:2017:CDC

[459] Luyang Li, Bing Qin, and Ting Liu. Contradiction detection with contradiction-specific word embedding. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/2/59.

Fu:2017:CNR

[460] Yan Fu and Shengchun Wang. Correction: A No Reference Image Quality Assessment Metric Based on Visual Perception. Algorithms 2016, 9, 87.

Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/2/60.

See [410].

Stefani:2017:DIM

[461] Elisabetta Stefani and Carlo Ferrari. Design and implementation of a multi-modal biometric system for company access control. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/2/61.

Liu:2017:IFA

[462] Hanbing Liu, Mengsu Zhang, Xianqiang Wang, Shuai Tian, and Yubo Jiao. Influence factors analysis on the modal characteristics of irregularlyshaped bridges based on a free-interface mode synthesis algorithm. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/2/62.

Purnamasari:2017:DFB

[463] Prima Dewi Purnamasari, Anak Agung Putri Ratna, and Benyamin Kusumoputro. Development of filtered bispectrum for EEG signal feature extraction in automatic emotion recognition using artificial neural networks. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/2/63.

Amat:2017:EAS

[464] Sergio Amat, Ioannis K. Argyros, Miguel A. Hernández-Verón, and Natalia Romero. Expanding the applicability of some high order Householder-like methods. Algorithms (Basel), 10(2), June 2017. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/10/2/64.

Tian:2017:SSC

[465] Xin Tian and Song Li. Seismic signal compression using nonparametric Bayesian dictionary learning via clustering. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/2/65.

Hassannejad:2017:NAI

[466] Hamid Hassannejad, Guido Matrella, Paolo Ciampolini, Ilaria De Munari, Monica Mordonini, and Stefano Cagnoni. A new approach to imagebased estimation of food volume. Algorithms (Basel), 10(2), June 2017.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/2/66.

Xiao:2017:RMF

[467] Yancai Xiao, Yujia Wang, Huan Mu, and Na Kang. Research on misalignment fault isolation of wind turbines based on the mixed-domain features. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/2/67.

Precup:2017:EUG

[468] Radu-Emil Precup, Radu-Codrut David, Alexandra-Iulia Szedlak-Stinean, Emil M. Petriu, and Florin Dragan. An easily understandable grey wolf optimizer and its application to fuzzy controller tuning. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/2/68.

Ratna:2017:LPD

[469] Anak Agung Putri Ratna, Prima Dewi Purnamasari, Boma Anantasatya Adhi, F. Astha Ekadiyanto, Muhammad Salman, Mardiyah Mardiyah, and Darien Jonathan Winata. ^{Cross−} language plagiarism detection system using latent semantic analysis and learning vector quantization. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/2/69.

Mei:2017:IBI

[470] Ying Mei, Guanzheng Tan, and Zhentao Liu. An improved brain-inspired emotional learning algorithm for fast classification. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/2/70.

Bi:2017:BCE

[471] Qixuan Bi and Wenhao Gui. Bayesian and classical estimation of stress-strength reliability for inverse Weibull lifetime models. Algorithms (Basel), 10(2), June 2017. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/10/2/71.

Guzman:2017:DOF

[472] Juan Carlos Guzman, Patricia Melin, and German Prado-Arechiga. Design of an optimized fuzzy classifier for the diagnosis of blood pressure with a new computational method for expert rule optimization. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/3/79.

ElKhamlichi:2017:HAO

[473] Yasser El Khamlichi, Abderrahim Tahiri, Anouar Abtoy, Inmaculada Medina-Bulo, and Francisco Palomo-Lozano. A hybrid algorithm for optimal wireless sensor network deployment with the minimum number of sensor nodes. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/80.

Wei:2017:LRC

[474] Xiaocong Wei, Hongfei Lin, Yuhai Yu, and Liang Yang. Low-resource cross-domain product review sentiment classification based on a CNN with an auxiliary large-scale corpus. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/10/3/81.

Peraza:2017:OIC

[475] Cinthia Peraza, Fevrier Valdez, and Patricia Melin. Optimization of intelligent controllers using a type-1 and interval type-2 fuzzy harmony search algorithm. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/82.

Barraza:2017:FFA

[476] Juan Barraza, Patricia Melin, Fevrier Valdez, and Claudia I. Gonzalez. Fuzzy fireworks algorithm based on a sparks dispersion measure. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/3/
83.

Xie:2017:AMB

[477] Li Xie and Huizhong Yang. Auxiliary model based multi-innovation stochastic gradient identification algorithm for periodically nonuniformly sampled-data Hammerstein systems. Algorithms (Basel), 10 (3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/10/3/84.

Caraveo:2017:NMH

[478] Camilo Caraveo, Fevrier Valdez, and Oscar Castillo. A new metaheuristics of optimization with dynamic adaptation of parameters using type-2 fuzzy logic for trajectory control of a mobile robot. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/3/85.

Zheng:2017:IMD

[479] Wei Zheng, Yanyan Tan, Xiaonan Fang, and Shengtao Li. An improved MOEA/D with optimal DE schemes for many-objective optimization problems. Algorithms (Basel), 10(3), September 2017. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/86.

Shahbazi:2017:EOR

[480] Mozhdeh Shahbazi, Gunho Sohn, and Jérôme Théau. Evolutionary optimization for robust epipolar-geometry estimation and outlier detection. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/3/
87.

Mezzadri:2017:LDM

[481] Francesco Mezzadri and Emanuele Galligani. On the lagged diffusivity method for the solution of nonlinear finite difference systems. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/3/88.

Ezquerro:2017:ESN

[482] José Antonio Ezquerro and Miguel Ángel Hernández-Verón. On the existence of solutions of nonlinear Fredholm integral equations from Kantorovich’s technique. Algorithms (Basel), 10(3), September 2017. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/10/3/89.

Hamann:2017:LCD

[483] Michael Hamann, Eike Röhrs, and Dorothea Wagner. Local community detection based on small cliques. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/3/90.

Chen:2017:TBF

[484] Chengyuan Chen and Qiang Shen. Transformation-based fuzzy rule interpolation using interval type-2 fuzzy sets. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/3/91.

Xie:2017:AMR

[485] Lijin Xie and Qun Wan. Automatic modulation recognition using compressive cyclic features. Algorithms (Basel), 10(3), September 2017. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/10/3/92.

Weir:2017:PPP

[486] William H. Weir, Scott Emmons, Ryan Gibson, Dane Taylor, and Peter J. Mucha. Post-processing partitions to identify domains of modularity optimization. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/93.

Sun:2017:NPR

[487] Peng Sun, Zhiming Yang, Yang Yu, Junbao Li, and Xiyuan Peng. NBTI and power reduction using an input vector control and supply voltage assignment method. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/94.

Xiao:2017:PTS

[488] Manyu Xiao, Quanyi Lv, Zhuo Xing, and Yingchun Zhang. A parallel two-stage iteration method for solving continuous Sylvester equations. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/3/
95.

Diao:2017:DTC

[489] Xueying Diao, Qianhui Dong, Zijian Yang, and Yibing Li. Doublethreshold cooperative spectrum sensing algorithm based on Sevcik fractal dimension. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/96.

Amsalu:2017:SMF

[490] Seifemichael B. Amsalu, Abdollah Homaifar, and Albert C. Esterline. A simplified matrix formulation for sensitivity analysis of hidden Markov models. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/3/97.

Chen:2017:AVR

[491] Fei Chen, Xiaohong Bi, Ruimin Lyu, Zhongwei Hua, Yuan Liu, and Xiaoting Zhang. Adaptive virtual RSU scheduling for scalable coverage under bidirectional vehicle traffic flow. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/3/98.

Sanchez:2017:HLG

[492] Mauricio A. Sanchez, Juan R. Castro, Violeta Ocegueda-Miramontes, and Leticia Cervantes. Hybrid learning for general type-2 TSK fuzzy logic systems. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/99.

Ye:2017:BBO

[493] Tao Ye, Ziqiang Yang, and Siling Feng. Biogeography-based optimization of the portfolio optimization problem with second order stochastic dominance constraints. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/100.

Olivas:2017:CST

[494] Frumen Olivas, Leticia Amador-Angulo, Jonathan Perez, Camilo Caraveo, Fevrier Valdez, and Oscar Castillo. Comparative study of type-2 fuzzy particle swarm, bee colony and bat algorithms in optimization of fuzzy controllers. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/3/101.

Nathan:2017:LCD

[495] Eisha Nathan, Anita Zakrzewska, Jason Riedy, and David A. Bader. Local community detection in dynamic graphs using personalized centrality. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/3/
102.

Zhang:2017:EDS

[496] Jingbo Zhang, Henan Yu, and Shufang Zhang. An enhanced dynamic spectrum allocation algorithm based on Cournot game in maritime cognitive radio communication system. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/3/103.

Zhao:2017:CBI

[497] Nan Zhao, Menglin Fan, Chao Tian, and Pengfei Fan. Contract-based incentive mechanism for mobile crowdsourcing networks. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/3/104.

Hamalainen:2017:CIC

[498] Joonas Hämäläinen, Susanne Jauhiainen, and Tommi Kärkkäinen. Comparison of internal clustering validation indices for prototype-based clustering. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/3/105.

Atanassov:2017:TFS

[499] Krassimir T. Atanassov. Type-1 fuzzy sets and intuitionistic fuzzy sets. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/3/
106.

Chen:2017:MBO

[500] Shifeng Chen, Rong Chen, and Jian Gao. A monarch butterfly optimization for the dynamic vehicle routing problem. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/3/107.

Wang:2017:PAF

[501] Deyun Wang, Chenqiang Yue, Shuai Wei, and Jun Lv. Performance analysis of four decomposition-ensemble models for one-day-ahead agricultural commodity futures price forecasting. Algorithms (Basel), 10(3), September 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/3/108.

Yang:2017:GTI

[502] Guanci Yang. Game theory-inspired evolutionary algorithm for global optimization. Algorithms (Basel), 10(4), December 2017. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/4/111.

Edler:2017:MHO

[503] Daniel Edler, Ludvig Bohlin, and Martin Rosvall. Mapping higher-order network flows in memory and multilayer networks with infomap. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/10/4/112.

Shahinpour:2017:SRT

[504] Shahram Shahinpour, Shirin Shirvani, Zeynep Ertem, and Sergiy Butenko. Scale reduction techniques for computing maximum induced bicliques. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/4/113.

Tyralis:2017:VST

[505] Hristos Tyralis and Georgia Papacharalampous. Variable selection in time series forecasting using random forests. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/4/114.

Cui:2017:MTA

[506] Huanqing Cui, Jian Niu, Chuanai Zhou, and Minglei Shu. A multithreading algorithm to detect and remove cycles in vertex- and arcweighted digraph. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/4/115.

Li:2017:RSI

[507] Liangliang Li, Yujuan Si, and Zhenhong Jia. Remote sensing image enhancement based on non-local means filter in NSCT domain. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/10/4/116.

Khan:2017:FWP

[508] Babar Khan, Zhijie Wang, Fang Han, Ather Iqbal, and Rana Javed Masood. Fabric weave pattern and yarn color recognition and classification using a deep ELM network. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/10/4/117.

Liu:2017:IPE

[509] Siyu Liu, Ling Xu, and Feng Ding. Iterative parameter estimation algorithms for dual-frequency signal models. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/4/118.

Cao:2017:OAI

[510] Zijian Cao and Lei Wang. An optimization algorithm inspired by the phase transition phenomenon for global optimization problems with continuous variables. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/4/119.

ElHadiSaad:2017:CSR

[511] Abdulbaset El Hadi Saad, Zuomin Dong, and Meysam Karimi. A comparative study on recently-introduced nature-based global optimization methods in complex mechanical system design. Algorithms (Basel), 10 (4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/10/4/120.

Bewoor:2017:EHP

[512] Laxmi A. Bewoor, V. Chandra Prakash, and Sagar U. Sapkal. Evolutionary hybrid particle swarm optimization algorithm for solving NP-hard no-wait flow shop scheduling problems. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/4/121.

Yaw:2017:SNP

[513] Sean Yaw and Brendan Mumey. Scheduling non-preemptible jobs to minimize peak demand. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/4/122.

Chehouri:2017:SPG

[514] Adam Chehouri, Rafic Younes, Jihan Khoder, Jean Perron, and Adrian Ilinca. A selection process for genetic algorithm using clustering analysis. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/4/
123.

Wang:2017:IIA

[515] Yingying Wang, Yibin Li, Yong Song, Xuewen Rong, and Shuaishuai Zhang. Improvement of ID3 algorithm based on simplified information entropy and coordination degree. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/4/124.

Tang:2017:SOS

[516] Jian Tang, Kejun Zhu, Haixiang Guo, Can Liao, and Shuwen Zhang. Simulation optimization of search and rescue in disaster relief based on distributed auction mechanism. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/4/125.

Zaveri:2017:LDL

[517] Amrapali Zaveri and Gökhan Ertaylan. Linked data for life sciences. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/4/
126.

Zhang:2017:RTC

[518] Jianming Zhang, Manting Huang, Xiaokang Jin, and Xudong Li. A realtime Chinese traffic sign detection algorithm based on modified YOLOv2. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/10/4/
127.

Du:2017:TSO

[519] Feng Du, Qiao-Yue Dong, and Hong-Shuang Li. Truss structure optimization with subset simulation and augmented Lagrangian multiplier method. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/4/128.

Podusenko:2017:CAC

[520] Albert Podusenko, Vsevolod Nikulin, Ivan Tanev, and Katsunori Shimohara. Comparative analysis of classifiers for classification of emergency braking of road motor vehicles. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/4/129.

Eftekharian:2017:PNI

[521] Seyedeh Elham Eftekharian, Mohammad Shojafar, and Shahaboddin Shamshirband. 2-phase NSGA II: an optimized reward and risk measurements algorithm in portfolio optimization. Algorithms (Basel), 10 (4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/10/4/130.

Chen:2017:ICC

[522] Su-Ting Chen, Chuang Zhang, Peng Li, Yan-Yan Zhang, and Liang-Bao Jiao. An indoor collaborative coefficient-triangle APIT localization algorithm. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
10/4/131.

Pizzolante:2017:HDE

[523] Raffaele Pizzolante and Bruno Carpentieri. Hyperspectral data: Efficient and secure transmission. Algorithms (Basel), 10(4), December 2017. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/10/4/132.

Ye:2017:NLE

[524] Jun Ye. Neutrosophic linear equations and application in traffic flow problems. Algorithms (Basel), 10(4), December 2017. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/4/133.

Zhang:2017:III

[525] Shuai Zhang and Xiao Qi. Improved integral inequalities for stability analysis of interval time-delay systems. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/4/134.

Berkemer:2017:ADP

[526] Sarah J. Berkemer, Christian Höner zu Siederdissen, and Peter F. Stadler. Algebraic dynamic programming on trees. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/10/4/135.

Yang:2017:DCF

[527] Yi Yang and Chu Pan. Detecting composite functional module in miRNA regulation and mRNA interaction network. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/10/4/136.

Hernandez-Gomez:2017:WCD

[528] Jorge J. Hernández-Gómez, Carlos Couder-Castañeda, Israel E. Herrera-Díaz, Norberto Flores-Guzmán, and Enrique Gómez-Cruz. Weakly coupled distributed calculation of Lyapunov exponents for non-linear dynamical systems. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/10/4/137.

Feng:2017:HML

[529] Shou Feng, Ping Fu, and Wenbin Zheng. A hierarchical multi-label classification algorithm for gene function prediction. Algorithms (Basel), 10 (4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/10/4/138.

Nai:2017:ESB

[530] Wei Nai, Lu Liu, Shaoyin Wang, and Decun Dong. An EMD–SARIMAbased modeling approach for air traffic forecasting. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/10/4/139.

Auer:2017:COM

[531] Ekaterina Auer, Luise Senkel, Stefan Kiel, and Andreas Rauh. Controloriented models for SO fuel cells from the angle of V&V: Analysis, simplification possibilities, performance. Algorithms (Basel), 10(4), December 2017. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/10/4/140.

Kasprzhitskii:2018:ISS

[532] Anton Kasprzhitskii, Georgy Lazorenko, and Victor Yavna. Iteration scheme for solving the system of coupled integro-differential equations for excited and ionized states of molecular systems. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/1/1.

Wang:2018:MMA

[533] Jie Wang, Xiyue Tang, and Guiwu Wei. Models for multiple attribute decision-making with dual generalized single-valued neutrosophic Bonferroni mean operators. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/1/2.

Filion:2018:ACC

[534] Guillaume J. Filion. Analytic combinatorics for computing seeding probabilities. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/

11/1/3. See correction [1810].

Lin:2018:TSS

[535] Daoyu Lin, Yang Wang, Guangluan Xu, Jun Li, and Kun Fu. Transform a simple sketch to a Chinese painting by a multiscale deep neural network. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/1/4.

Tang:2018:AMA

[536] Xiyue Tang, Yuhan Huang, and Guiwu Wei. Approaches to multipleattribute decision-making based on Pythagorean 2-tuple linguistic Bonferroni mean operators. Algorithms (Basel), 11(1), January 2018. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/11/1/5.

Ding:2018:NPH

[537] Kaimeng Ding, Shiping Chen, and Fan Meng. A novel perceptual hash algorithm for multispectral image authentication. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/1/6.

Sheng:2018:ODG

[538] Lianchao Sheng and Wei Li. Optimization design by genetic algorithm controller for trajectory control of a 3-RRR parallel robot. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/1/7.

Peretz:2018:ARS

[539] Yossi Peretz. On application of the ray-shooting method for LQR via static-output-feedback. Algorithms (Basel), 11(1), January 2018. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/11/1/8.

Cui:2018:AHM

[540] Wei Cui, Qi Zhou, and Zhendong Zheng. Application of a hybrid model based on a convolutional auto-encoder and convolutional neural network in object-oriented remote sensing classification. Algorithms (Basel), 11 (1), January 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/1/9.

Manurangsi:2018:IMB

[541] Pasin Manurangsi. Inapproximability of maximum biclique problems, minimum k-cut and densest at-least-k-subgraph from the small set expansion hypothesis. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/1/10.

Office:2018:ARA

[542] Algorithms Editorial Office. Acknowledgement to reviewers of algorithms in 2017. Algorithms (Basel), 11(1), January 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/1/11.

Kou:2018:NMC

[543] Farong Kou, Jiafeng Du, Zhe Wang, Dong Li, and Jianan Xu. Nonlinear modeling and coordinate optimization of a semi-active energy regenerative suspension with an electro-hydraulic actuator. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/2/12.

Boldrini:2018:MMA

[544] Stefano Boldrini, Luca De Nardis, Giuseppe Caso, Mai T. P. Le, Jocelyn Fiorina, and Maria-Gabriella Di Benedetto. muMAB: a multi-armed bandit model for wireless network selection. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/2/13.

Wu:2018:OOR

[545] Guangyuan Wu, Zhigang Chen, Lin Guo, and Jia Wu. An optimal online resource allocation algorithm for energy harvesting body area networks. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
2/14.

Nai:2018:MTC

[546] Wei Nai, Lu Liu, Shaoyin Wang, and Decun Dong. Modeling the trend of credit card usage behavior for different age groups based on singular spectrum analysis. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/2/15.

Liu:2018:NSS

[547] Liping Liu, Ning Wang, Zhigang Chen, and Lin Guo. A novel spectrum scheduling scheme with ant colony optimization algorithm. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/2/16.

Lv:2018:IBF

[548] Xinen Lv, Huiling Chen, Qian Zhang, Xujie Li, Hui Huang, and Gang Wang. An improved bacterial-foraging optimization-based machine learning framework for predicting the severity of somatization disorder. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
2/17.

Zhang:2018:NGI

[549] Hongliang Zhang, Youcai Fang, Ruilin Pan, and Chuanming Ge. A new greedy insertion heuristic algorithm with a multi-stage filtering mechanism for energy-efficient single machine scheduling problems. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/2/18.

Lemke:2018:CNN

[550] Oliver Lemke and Bettina G. Keller. Common nearest neighbor clustering — a benchmark. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/2/19.

Mouawad:2018:VCR

[551] Amer E. Mouawad, Naomi Nishimura, Venkatesh Raman, and Sebastian Siebertz. Vertex cover reconfiguration and beyond. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/2/20.

Zhang:2018:RDM

[552] Lijun Zhang and Junyu Tao. Research on degeneration model of neural network for deep groove ball bearing based on feature fusion. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/2/21.

Zhou:2018:DOS

[553] Chen Zhou, Xinhui Liu, and Feixiang Xu. Design optimization of steering mechanisms for articulated off-road vehicles based on genetic algorithms. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
2/22.

Dai:2018:ERV

[554] Hou-Ping Dai, Dong-Dong Chen, and Zhou-Shun Zheng. Effects of random values for particle swarm optimization algorithm. Algorithms (Basel), 11(2), February 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/2/23.

Yi:2018:CAC

[555] Hua Yi, Shi-You Xin, and Jun-Feng Yin. A class of algorithms for continuous wavelet transform based on the circulant matrix. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/3/24.

Cagnoni:2018:SIC

[556] Stefano Cagnoni and Mauro Castelli. Special issue on computational intelligence and nature-inspired algorithms for real-world data analytics and pattern recognition. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/3/25.

Li:2018:NEA

[557] Shaobo Li, Wang Zou, and Jianjun Hu. A novel evolutionary algorithm for designing robust analog filters. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/3/26.

Liu:2018:SAB

[558] Liping Liu, Ning Wang, Zhigang Chen, and Lin Guo. Spectrum allocation based on an improved gravitational search algorithm. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/3/27.

Yang:2018:MCN

[559] Jing Yang and Guanci Yang. Modified convolutional neural network based on dropout and the stochastic gradient descent optimizer. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/3/28.

Shi:2018:DAO

[560] Lilian Shi and Jun Ye. Dombi aggregation operators of neutrosophic cubic sets for multiple attribute decision-making. Algorithms (Basel), 11 (3), March 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/3/29.

Liu:2018:MCS

[561] Liping Liu, Xiaobo Liu, Ning Wang, and Peijun Zou. Modified cuckoo search algorithm with variational parameters and logistic map. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/3/30.

Ahmed:2018:BLS

[562] A. K. M. Foysal Ahmed and Ji Ung Sun. Bilayer local search enhanced particle swarm optimization for the capacitated vehicle routing problem. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/3/31.

Jaiyeola:2018:IPN

[563] Temitope Gbolahan Jaiyeola and Florentin Smarandache. Inverse properties in neutrosophic triplet loop and their application to cryptography. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/3/32.

Qin:2018:OEM

[564] Feiyan Qin, Weimin Li, Yue Hu, and Guoqing Xu. An online energy management control for hybrid electric vehicles based on neuro-dynamic programming. Algorithms (Basel), 11(3), March 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/3/33.

Zhu:2018:FME

[565] Jianghong Zhu, Rui Wang, and Yanlai Li. Failure mode and effects analysis considering consensus and preferences interdependence. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/4/34.

Kriheli:2018:EBA

[566] Boris Kriheli and Eugene Levner. Entropy-based algorithm for supplychain complexity assessment. Algorithms (Basel), 11(4), April 2018. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/11/4/35.

Feng:2018:GBC

[567] Yu Feng, Jianzhong Zhou, Li Mo, Chao Wang, Zhe Yuan, and Jiang Wu. A gradient-based cuckoo search algorithm for a reservoir-generation scheduling problem. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/4/36.

Kaiser:2018:GKM

[568] Waldemar Kaiser, Johannes Popp, Michael Rinderle, Tim Albes, and Alessio Gagliardi. Generalized kinetic Monte Carlo framework for organic electronics. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/4/37.

Papalitsas:2018:CGG

[569] Christos Papalitsas, Panayiotis Karakostas, Theodore Andronikos, Spyros Sioutas, and Konstantinos Giannakis. Combinatorial GVNS (general variable neighborhood search) optimization for dynamic garbage collection. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/4/
38.

Hasimu:2018:HTL

[570] Maimaitiyiming Hasimu and Wushour Silamu. On hierarchical text language-identification algorithms. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/4/39.

Haas:2018:CHC

[571] Ruth Haas and Gary MacGillivray. Connectivity and Hamiltonicity of canonical colouring graphs of bipartite and complete multipartite graphs. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/4/40.

He:2018:DIM

[572] Zhenwen He and Xiaogang Ma. A distributed indexing method for timeline similarity query. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/4/41.

Yasuda:2018:LAB

[573] Muneki Yasuda. Learning algorithm of Boltzmann machine based on spatial Monte Carlo integration method. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/4/42.

Fuchigami:2018:NOH

[574] Helio Yochihiro Fuchigami, Ruhul Sarker, and Socorro Rangel. Nearoptimal heuristics for just-in-time jobs maximization in flow shop scheduling. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/4/43.

Zhang:2018:SPP

[575] Hong-Mei Zhang, Ming-Long Li, and Le Yang. Safe path planning of mobile robot based on improved A∗ algorithm in complex terrains. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/4/44.

Klein:2018:AAG

[576] Rolf Klein, Christos Levcopoulos, and Andrzej Lingas. Approximation algorithms for the geometric firefighter and budget fence problems. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/4/45.

Rodriguez:2018:SRC

[577] Mirella Rodriguez and Daniel R. Jeske. Short-run contexts and imperfect testing for continuous sampling plans. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/4/46.

Xing:2018:NDG

[578] Yanzhen Xing, Donghui Wang, and Leiou Wang. A novel dynamic generalized opposition-based grey wolf optimization algorithm. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/4/47.

Wang:2018:ASP

[579] Xinxin Wang, Xiaoqiang Yan, Donghai Li, and Li Sun. An approach for setting parameters for two-degree-of-freedom PID controllers. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/4/48.

Tian:2018:TEP

[580] Qing Tian, Weihang Zhao, Yun Wei, and Liping Pang. Thermal environment prediction for metro stations based on an RVFL neural network. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/4/49.

Lazarev:2018:ETA

[581] Alexander A. Lazarev, Ivan Nekrasov, and Nikolay Pravdivets. Evaluating typical algorithms of combinatorial optimization to solve continuoustime based scheduling problem. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/4/50.

Lv:2018:CCS

[582] Xinxin Lv and Qi Zhu. A crowd cooperative spectrum sensing algorithm using a non-ideal channel. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/4/51.

Nishimura:2018:IR

[583] Naomi Nishimura. Introduction to reconfiguration. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/4/52.

Russo:2018:LCS

[584] Luís M. S. Russo, Andreia Sofia Teixeira, and Alexandre P. Francisco. Linking and cutting spanning trees. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/4/53.

Mogale:2018:DMF

[585] D. G. Mogale, Geet Lahoti, Shashi Bhushan Jha, Manish Shukla, Narasimha Kamath, and Manoj Kumar Tiwari. Dual market facility network design under bounded rationality. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/4/54.

Gholami:2018:HAS

[586] Omid Gholami and Johanna Törnquist Krasemann. A heuristic approach to solving the train traffic re-scheduling problem in real time. Algorithms (Basel), 11(4), April 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/4/55.

Senthilnath:2018:BBE

[587] J. Senthilnath, Sumanth Simha C, Nagaraj G, Meenakumari Thapa, and Indiramma M. BELMKN: Bayesian extreme learning machines Kohonen network. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/5/
56.

Sokolov:2018:OCA

[588] Boris Sokolov, Alexandre Dolgui, and Dmitry Ivanov. Optimal control algorithms and their analysis for short-term scheduling in manufacturing systems. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/5/
57.

Turau:2018:CFC

[589] Volker Turau. Computing fault-containment times of self-stabilizing algorithms using lumped Markov chains. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/58.

Akram:2018:DMA

[590] Muhammad Akram, Nabeela Ishfaq, Sidra Sayed, and Florentin Smarandache. Decision-making approach based on neutrosophic rough information. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/5/
59.

Sun:2018:VTR

[591] Molin Sun and Zhongyi Zheng. Vessel traffic risk assessment based on uncertainty analysis in the risk matrix. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/60.

Talmage:2018:RDT

[592] Edward Talmage and Jennifer L. Welch. Relaxed data types as consistency conditions. Algorithms (Basel), 11(5), May 2018. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/5/61.

Xie:2018:FWS

[593] Minghua Xie, Decheng Wang, and Lili Xie. A feature-weighted SVR method based on kernel space feature. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/62.

Shao:2018:SSM

[594] Minghui Shao, Yan Song, Biao Wu, and Yanjie Chang. The supplier selection of the marine rescue equipment based on the analytic hierarchy process (AHP)-limited diversity factors method. Algorithms (Basel), 11 (5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/5/63.

Fu:2018:UDS

[595] Ming Lan Fu, Hao Wang, and Bao Fu Fang. Utility distribution strategy of the task agents in coalition skill games. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/64.

Chen:2018:CSS

[596] Pengzhan Chen, Zhiqiang He, Chuanxi Chen, and Jiahong Xu. Control strategy of speed servo systems based on deep reinforcement learning. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/5/65.

Sotskov:2018:SMS

[597] Yuri N. Sotskov and Natalja G. Egorova. Single machine scheduling problem with interval processing times and total completion time objective. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/5/66.

Hocke:2018:APF

[598] Lia M. Hocke, Ibukunoluwa K. Oni, Chris C. Duszynski, Alex V. Corrigan, Blaise DeB. Frederick, and Jeff F. Dunn. Automated processing of fNIRS data — a visual guide to the pitfalls and consequences. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/5/67.

Yaurima-Basaldua:2018:HFS

[599] Victor Hugo Yaurima-Basaldua, Andrei Tchernykh, Francisco Villalobos-Rodríguez, and Ricardo Salomon-Torres. Hybrid flow shop with unrelated machines, setup time, and work in progress buffers for bi-objective optimization of tortilla manufacturing. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/68.

Cassettari:2018:MSA

[600] Lucia Cassettari, Melissa Demartini, Roberto Mosca, Roberto Revetria, and Flavio Tonelli. A multi-stage algorithm for a capacitated vehicle routing problem with time constraints. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/69.

Montero-Hernandez:2018:EFC

[601] Samuel Montero-Hernandez, Felipe Orihuela-Espina, Luis Enrique Sucar, Paola Pinti, Antonia Hamilton, Paul Burgess, and Ilias Tachtsidis. Estimating functional connectivity symmetry between oxy- and deoxyhaemoglobin: Implications for fNIRS connectivity analysis. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/5/70.

Hu:2018:IMB

[602] Hui Hu, Zhaoquan Cai, Song Hu, Yingxue Cai, Jia Chen, and Sibo Huang. Improving monarch butterfly optimization algorithm with selfadaptive population. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/5/71.

Ren:2018:GWO

[603] Yixuan Ren, Tao Ye, Mengxing Huang, and Siling Feng. Gray wolf optimization algorithm for multi-constraints second-order stochastic dominance portfolio optimization. Algorithms (Basel), 11(5), May 2018. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/11/5/72.

Santosa:2018:NBA

[604] Hendrik Santosa, Xuetong Zhai, Frank Fishburn, and Theodore Huppert. The NIRS Brain AnalyzIR Toolbox. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/5/73.

Stodola:2018:UMM

[605] Petr Stodola. Using metaheuristics on the multi-depot vehicle routing problem with modified optimization criterion. Algorithms (Basel), 11 (5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/5/74.

Wang:2018:NON

[606] Guohui Wang and Yuanbo Chu. A new Oren–Nayar shape-from-shading approach for 3D reconstruction using high-order Godunov-based scheme. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/5/75.

Drozdov:2018:PPI

[607] Alexander Yu. Drozdov, Andrei Tchernykh, Sergey V. Novikov, Victor E. Vladislavlev, and Raul Rivera-Rodriguez. PHEFT: Pessimistic image processing workflow scheduling for DSP clusters. Algorithms (Basel), 11 (5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/5/76.

Xu:2018:NDS

[608] Wei Xu, Yi Li, Jinghong Miao, Jiaxiang Zhao, and Xin Gao. A novel design of sparse prototype filter for nearly perfect reconstruction cosinemodulated filter banks. Algorithms (Basel), 11(5), May 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/5/77.

Pang:2018:MAB

[609] Bao Pang, Yong Song, Chengjin Zhang, Hongling Wang, and Runtao Yang. A modified artificial bee colony algorithm based on the selflearning mechanism. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/6/78.

Bian:2018:FDA

[610] Yongming Bian, Meng Yang, Xuying Fan, and Yuchao Liu. A fire detection algorithm based on Tchebichef moment invariants and PSO-SVM. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/6/79.

Vakhania:2018:SSM

[611] Nodari Vakhania. Scheduling a single machine with primary and secondary objectives. Algorithms (Basel), 11(6), June 2018. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/6/80.

Peretz:2018:RAO

[612] Yossi Peretz. A randomized algorithm for optimal PID controllers. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/6/81.

Wei:2018:RFD

[613] Yi Wei and Yaokun Yue. Research on fault diagnosis of a marine fuel system based on the SaDE-ELM algorithm. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/6/82.

Pang:2018:CSS

[614] David Pang and Tomohiko Igasaki. A combined syntactical and statistical approach for R peak detection in real-time long-term heart rate variability analysis. Algorithms (Basel), 11(6), June 2018. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/6/83.

Epasto:2018:EAR

[615] Alessandro Epasto and Eli Upfal. Efficient approximation for restricted biclique cover problems. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/6/84.

Lan:2018:IIS

[616] Yong-Hong Lan and Zhe-Min Cui. ILC with initial state learning for fractional order linear distributed parameter systems. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/6/85.

Dalen:2018:POP

[617] Christer Dalen and David Di Ruscio. Performance optimal PI controller tuning based on integrating plus time delay models. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/6/86.

Werner:2018:SIA

[618] Frank Werner, Larysa Burtseva, and Yuri N. Sotskov. Special issue on algorithms for scheduling problems. Algorithms (Basel), 11(6), June 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/6/87.

Ji:2018:EDL

[619] Qingge Ji, Wenjie He, Jie Huang, and Yankui Sun. Efficient deep learning-based automated pathology identification in retinal optical coherence tomography images. Algorithms (Basel), 11(6), June 2018. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/11/6/88.

Li:2018:DTP

[620] Qing Li and Steven Y. Liang. Degradation trend prediction for rotating machinery using long-range dependence and particle filter approach. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/7/89.

Liu:2018:NMC

[621] Kai Liu, YangQuan Chen, Paweł D. Domański, and Xi Zhang. A novel method for control performance assessment with fractional order signal processing and its application to semiconductor manufacturing. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/7/90.

Himpe:2018:EEG

[622] Christian Himpe. emgr — the empirical Gramian framework. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/7/91.

Yang:2018:PCC

[623] Guoliang Yang, Haitao Yi, Chunhua Chai, Bingxu Huang, Yuna Zhang, and Zhe Chen. Predictive current control of boost three-level and T-type inverters cascaded in wind power generation systems. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/7/92.

Chitturi:2018:LGA

[624] Bhadrachalam Chitturi, Srijith Balachander, Sandeep Satheesh, and Krithic Puthiyoppil. Layered graphs: Applications and algorithms. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/7/93.

Wei:2018:TCB

[625] Dongxu Wei, Andong Wang, Xiaoqin Feng, Boyu Wang, and Bo Wang. Tensor completion based on triple tubal nuclear norm. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/7/94.

Muresan:2018:EVN

[626] Cristina I. Muresan, Cosmin Copot, Isabela Birs, Robin De Keyser, Steve Vanlanduit, and Clara M. Ionescu. Experimental validation of a novel auto-tuning method for a fractional order PI controller on an UR10 robot. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/7/95.

Al-Saleh:2018:SMD

[627] Asma Al-Saleh and Mohamed El Bachir Menai. Solving multi-document summarization as an orienteering problem. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/7/96.

Tang:2018:RTM

[628] Lin Tang, Lin Liu, and Jianhou Gan. A regional topic model using hybrid stochastic variational Gibbs sampling for real-time video mining. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/7/97.

Chen:2018:WDS

[629] Li-Hsuan Chen, Felix Reidl, Peter Rossmanith, and Fernando Sánchez Villaamil. Width, depth, and space: Tradeoffs between branching and dynamic programming. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/7/98.

Feng:2018:LCR

[630] Jiao Feng, Xiaofei Zhang, Peng Li, and Dongshun Hu. A low complexity reactive tabu search based constellation constraints in signal detection. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/7/99.

Dulebenets:2018:SAE

[631] Maxim A. Dulebenets, Masoud Kavoosi, Olumide Abioye, and Junayed Pasha. A self-adaptive evolutionary algorithm for the berth scheduling problem: Towards efficient parameter control. Algorithms (Basel), 11 (7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/7/100.

Bourouba:2018:RFA

[632] Bachir Bourouba and Samir Ladaci. Robust fuzzy adaptive sliding mode stabilization for fractional-order chaos. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/7/101.

Chen:2018:AAN

[633] Tin-Chih Toly Chen, Cheng-Li Liu, and Hong-Dar Lin. Advanced artificial neural networks. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/7/102.

Sabatier:2018:SSO

[634] Jocelyn Sabatier. Solutions to the sub-optimality and stability issues of recursive pole and zero distribution algorithms for the approximation of fractional order models. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/7/103.

Gribanov:2018:GHD

[635] Igor Gribanov, Rocky Taylor, and Robert Sarracino. The gradient and the Hessian of the distance between point and triangle in 3D. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/7/104.

Damiand:2018:DCM

[636] Guillaume Damiand, Aldo Gonzalez-Lorenzo, Florence Zara, and Florent Dupont. Distributed combinatorial maps for parallel mesh processing. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/7/105.

Navarro-Guerrero:2018:FOC

[637] Gerardo Navarro-Guerrero and Yu Tang. Fractional-order closed-loop model reference adaptive control for anesthesia. Algorithms (Basel), 11 (7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/7/106.

Yang:2018:EEL

[638] Rui Yang, Shuliang Xu, and Lin Feng. An ensemble extreme learning machine for data stream classification. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/7/107.

Alekseeva:2018:ECA

[639] Natalia Alekseeva, Ivan Tanev, and Katsunori Shimohara. Evolving the controller of automated steering of a car in slippery road conditions. Algorithms (Basel), 11(7), July 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/7/108.

Liu:2018:LLD

[640] Liu Liu, Kaile Liu, Zhenghai Cong, Jiali Zhao, Yefei Ji, and Jun He. Long length document classification by local convolutional feature aggregation. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/8/109.

Volgyes:2018:IQU

[641] David Völgyes, Anne Catrine Trægde Martinsen, Arne Stray-Pedersen, Dag Waaler, and Marius Pedersen. Image de-quantization using plate bending model. Algorithms (Basel), 11(8), August 2018. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/8/110.

Volgyes:2018:WHB

[642] David Völgyes, Anne Catrine Trægde Martinsen, Arne Stray-Pedersen, Dag Waaler, and Marius Pedersen. A weighted histogram-based tone mapping algorithm for CT images. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/8/111.

Wang:2018:NPA

[643] Ruhua Wang, Ling Li, and Jun Li. A novel parallel auto-encoder framework for multi-scale data in civil structural health monitoring. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/8/112.

Su:2018:IPI

[644] Xiangfeng Su, Huaiqing Zhang, Lin Chen, Ling Qin, and Lili Yu. Improved parameter identification method for envelope current signals based on windowed interpolation FFT and DE algorithm. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/8/113.

Mezei:2018:RCS

[645] Mihaly Mezei. Revisiting chameleon sequences in the protein data bank. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/8/114.

Wang:2018:CBI

[646] Jing Wang, Lidong Wang, Xiaodong Liu, Yan Ren, and Ye Yuan. Colorbased image retrieval using proximity space theory. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/8/115.

Qi:2018:REE

[647] Huamei Qi, Fengqi Liu, Tailong Xiao, and Jiang Su. A robust and energyefficient weighted clustering algorithm on mobile ad hoc sensor networks. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/8/116.

Hu:2018:RVS

[648] Yanzhu Hu, Song Wang, and Xinbo Ai. Research of the vibration source tracking in phase-sensitive optical time-domain reflectometry signals based by image processing method. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/11/8/117.

Brodnik:2018:SST

[649] Andrej Brodnik and Matevž Jekovec. Sliding suffix tree. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/11/8/118.

Lin:2018:ONR

[650] Yucheng Lin, Zhigang Chen, Jia Wu, and Leilei Wang. An opportunistic network routing algorithm based on cosine similarity of data packets between nodes. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/8/119.

Wu:2018:PIV

[651] Wenying Wu, Ying Li, Zhiwei Ni, Feifei Jin, and Xuhui Zhu. Probabilistic interval-valued hesitant fuzzy information aggregation operators and their application to multi-attribute decision making. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/8/120.

Pan:2018:SGB

[652] Feilai Pan, Jun Li, Bendong Tan, Ciling Zeng, Xinfan Jiang, Li Liu, and Jun Yang. Stacked-GRU based power system transient stability assessment method. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/8/121.

Tsai:2018:EMB

[653] Chi-Yi Tsai, Kuang-Jui Hsu, and Humaira Nisar. Efficient model-based object pose estimation based on multi-template tracking and PnP algorithms. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/8/122.

Al-Douri:2018:TSF

[654] Yamur K. Al-Douri, Hussan Hamodi, and Jan Lundberg. Time series forecasting using a two-level multi-objective genetic algorithm: a case study of maintenance cost data for tunnel fans. Algorithms (Basel), 11 (8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/8/123.

Li:2018:SBI

[655] Yihong Li, Fangzheng Liu, Zhenyu Du, and Dubing Zhang. A Simhashbased integrative features extraction algorithm for malware detection. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/8/124.

Yan:2018:EDT

[656] Yeqing Yan, Zhigang Chen, Jia Wu, and Leilei Wang. An effective data transmission algorithm based on social relationships in opportunistic mobile social networks. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/8/125.

Song:2018:RVT

[657] Zhiguo Song, Jifeng Sun, Jialin Yu, and Shengqing Liu. Robust visual tracking via patch descriptor and structural local sparse representation. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/8/126.

Zeng:2018:AAR

[658] Mingbin Zeng, Xu Yang, Mengxing Wang, and Bangjiang Xu. Application of angle related cost function optimization for dynamic path planning algorithm. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/8/127.

Denzumi:2018:DCF

[659] Shuhei Denzumi, Jun Kawahara, Koji Tsuda, Hiroki Arimura, Shin ichi Minato, and Kunihiko Sadakane. DenseZDD: a compact and fast index for families of sets. Algorithms (Basel), 11(8), August 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/8/128.

Sheng:2018:FDI

[660] Zhiyong Sheng, Dandan Qu, Yuan Zhang, and Dan Yang. The fast detection and identification algorithm of optical fiber intrusion signals. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/9/
129.

Borkowski:2018:NMW

[661] Piotr Borkowski. Numerical modeling of wave disturbances in the process of ship movement control. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/9/130.

Groote:2018:EAD

[662] Jan Friso Groote, Jao Rivera Verduzco, and Erik P. De Vink. An efficient algorithm to determine probabilistic bisimulation. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/9/131.

Du:2018:SPF

[663] Jinglin Du, Yayun Liu, and Zhijun Liu. Study of precipitation forecast based on deep belief networks. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/9/132.

Zheng:2018:MSD

[664] Xiuyun Zheng and Jiarong Shi. A modified sufficient descent5 Polak– Ribiére–Polyak type conjugate gradient method for unconstrained optimization problems. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/9/133.

Russo:2018:MLE

[665] Gabriele Russo Russo, Matteo Nardelli, Valeria Cardellini, and Francesco Lo Presti. Multi-level elasticity for wide-area data streaming systems: a reinforcement learning approach. Algorithms (Basel), 11 (9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/9/134.

Ye:2018:MAD

[666] Jun Ye and Wenhua Cui. Multiple attribute decision-making method using linguistic cubic hesitant variables. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/11/9/135.

Duarte-Mermoud:2018:MOF

[667] Manuel A. Duarte-Mermoud, Javier A. Gallegos, Norelys Aguila-Camacho, and Rafael Castro-Linares. Mixed order fractional observers for minimal realizations of linear time-invariant systems. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/9/136.

Ai:2018:LHK

[668] Qingyao Ai, Vahid Azizi, Xu Chen, and Yongfeng Zhang. Learning heterogeneous knowledge base embeddings for explainable recommendation. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/9/
137.

Das:2018:MTE

[669] Sanjiv R. Das, Karthik Mokashi, and Robbie Culkin. Are markets truly efficient? Experiments using deep learning algorithms for market movement prediction. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/9/138.

Livieris:2018:AAS

[670] Ioannis E. Livieris, Andreas Kanavos, Vassilis Tampakas, and Panagiotis Pintelas. An auto-adjustable semi-supervised self-training algorithm. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/9/
139.

Takaoka:2018:CHC

[671] Asahi Takaoka. Complexity of Hamiltonian cycle reconfiguration. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/9/140.

Pires:2018:GPM

[672] Miguel Pires, Srivatsan Ravi, and Rodrigo Rodrigues. Generalized Paxos made Byzantine (and less complex). Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/11/9/141.

Gao:2018:EVS

[673] Wei Gao, Hengyi Lv, Qiang Zhang, and Dunbo Cai. Estimating the volume of the solution space of SMT(LIA) constraints by a flat histogram method. Algorithms (Basel), 11(9), September 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/9/142.

Vlachostergiou:2018:LRN

[674] Aggeliki Vlachostergiou, George Caridakis, Phivos Mylonas, and Andreas Stafylopatis. Learning representations of natural language texts with generative adversarial networks at document, sentence, and aspect level. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/10/164.

Liu:2018:AOD

[675] Xinqiang Liu and Weiliang He. Airfoil optimization design based on the pivot element weighting iterative method. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/10/163.

Tang:2018:ADS

[676] Haijing Tang, Guo Chen, Yu Kang, and Xu Yang. Application of data science technology on research of circulatory system disease prediction based on a prospective cohort. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/10/162.

Vignesh:2018:TCC

[677] R. Vignesh, J. Geetha, and K. Somasundaram. Total coloring conjecture for certain classes of graphs. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/10/161.

Fakhrolmobasheri:2018:MEP

[678] Sharifeh Fakhrolmobasheri, Ehsan Ataie, and Ali Movaghar. Modeling and evaluation of power-aware software rejuvenation in cloud systems. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
10/160.

Zhao:2018:FAR

[679] Yulin Zhao, Donghui Wang, Leiou Wang, and Peng Liu. A faster algorithm for reducing the computational complexity of convolutional neural networks. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/10/159.

Madhusudhanan:2018:ILC

[680] Sathya Madhusudhanan, Suresh Jaganathan, and Jayashree L. S. Incremental learning for classification of unstructured data using extreme learning machine. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/10/158.

Savvopoulos:2018:LAC

[681] Alkiviadis Savvopoulos, Andreas Kanavos, Phivos Mylonas, and Spyros Sioutas. LSTM accelerator for convolutional object identification. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/11/10/157.

Zhou:2018:OUI

[682] Rong Zhou, Chun Chen, Liqun Sun, Francis C. M. Lau, Sheung-Hung Poon, and Yong Zhang. Online uniformly inserting points on the sphere. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
10/156.

Choi:2018:RTT

[683] Jang-Hwan Choi and Sooyeul Lee. Real-time tumor motion tracking in 3D using planning 4D CT images during image-guided radiation therapy. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
10/155.

Pei:2018:THM

[684] Lidan Pei and Feifei Jin. Two hesitant multiplicative decision-making algorithms and their application to fog-haze factor assessment problem. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
10/154.

Wang:2018:CRT

[685] Di Wang, Frank McGroarty, and Eng-Tuck Cheah. Chronotype, risk and time preferences, and financial behaviour. Algorithms (Basel), 11 (10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/10/153.

Ma:2018:AIL

[686] Dongqi Ma and Hui Lin. Accelerated iterative learning control of speed ripple suppression for a seeker servo motor. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/10/152.

Hedar:2018:MCA

[687] Abdel-Rahman Hedar, Abdel-Monem M. Ibrahim, Alaa E. Abdel-Hakim, and Adel A. Sewisy. K-means cloning: Adaptive spherical K-means clustering. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/10/151.

Gharib:2018:SES

[688] Mohammed Gharib, Marzieh Malekimajd, and Ali Movaghar. SLoP-Cloud: an efficient solution for locality problem in peer-to-peer cloud systems. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/10/150.

Lamprou:2018:CTE

[689] Ioannis Lamprou, Russell Martin, and Paul Spirakis. Cover time in edge-uniform stochastically-evolving graphs. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/10/149.

Kofinas:2018:FLA

[690] Panagiotis Kofinas and Anastasios I. Dounis. Fuzzy Q-learning agent for online tuning of PID controller for DC motor speed control. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/10/148.

Yin:2018:WNP

[691] Hong Yin, Ying Zhang, and Xu He. WSN nodes placement optimization based on a weighted centroid artificial fish swarm algorithm. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/10/147.

Almabrok:2018:FTP

[692] Abdoalnasir Almabrok, Mihalis Psarakis, and Anastasios Dounis. Fast tuning of the PID controller in an HVAC system using the Big Bang–Big Crunch algorithm and FPGA technology. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/10/146.

Lagana:2018:ROC

[693] Demetrio Laganà, Carlo Mastroianni, Michela Meo, and Daniela Renga. Reducing the operational cost of cloud data centers through renewable energy. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/10/145.

Liu:2018:MBD

[694] Peng Liu, Ying Hong, and Yan Liu. Multi-branch deep residual network for single image super-resolution. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/10/144.

Essani:2018:AMA

[695] Furqan Hussain Essani and Sajjad Haider. An algorithm for mapping the asymmetric multiple traveling salesman problem onto colored petri nets. Algorithms (Basel), 11(10), October 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/10/143.

Ju:2018:VSF

[696] Jinyong Ju, Yongrui Zhao, Chunrui Zhang, and Yufei Liu. Vibration suppression of a flexible-joint robot based on parameter identification and fuzzy PID control. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/11/189.

Zhang:2018:DEB

[697] Xiangyin Zhang, Yuying Xue, Xingyang Lu, and Songmin Jia. Differential-evolution-based coevolution ant colony optimization algorithm for Bayesian network structure learning. Algorithms (Basel), 11 (11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/11/188.

Abu-Khzam:2018:SIR

[698] Faisal Abu-Khzam, Henning Fernau, and Ryuhei Uehara. Special issue on reconfiguration problems. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/11/187.

Li:2018:PSL

[699] Tao Li, Yan Chen, and Taoying Li. Pricing strategies of logistics distribution services for perishable commodities. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/11/186.

Ma:2018:ALF

[700] Lin Ma and Jean-Paul Delahaye. An algorithmic look at financial volatility. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/11/185.

Li:2018:WFD

[701] Qing Li and Steven Y. Liang. Weak fault detection of tapered rolling bearing based on penalty regularization approach. Algorithms (Basel), 11 (11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/11/184.

Bian:2018:VBA

[702] Chentong Bian, Guodong Yin, Liwei Xu, and Ning Zhang. Virtual belt algorithm for the management of isolated autonomous intersection. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
11/183.

Zhuang:2018:AIV

[703] Hua Zhuang, Yanzhao Tang, and Meijuan Li. An algorithm for intervalvalued intuitionistic fuzzy preference relations in group decision making based on acceptability measurement and priority weight determination. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
11/182.

Kollintza-Kyriakoulia:2018:MIF

[704] Foteini Kollintza-Kyriakoulia, Manolis Maragoudakis, and Anastasia Krithara. Measuring the impact of financial news and social media on stock market modeling using time series mining techniques. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/11/181.

You:2018:IIM

[705] Junyao You and Yanjun Liu. Iterative identification for multivariable systems with time-delays based on basis pursuit de-noising and auxiliary model. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/11/180.

Li:2018:RSB

[706] Peng Li, Chenchen Shu, and Jiao Feng. A reciprocal-selection-based ‘win–win’ overlay spectrum-sharing scheme for device-to-device-enabled cellular network. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/11/179.

Hu:2018:DDN

[707] Zhaohua Hu and Xiaoyi Shi. Deep directional network for object tracking. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
11/178.

Gao:2018:UEI

[708] Xuedong Gao and Minghan Yang. Understanding and enhancement of internal clustering validation indexes for categorical data. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/11/177.

MeloBarbosa:2018:TVD

[709] Paulo Alberto Melo Barbosa, Plácido Rogério Pinheiro, and Francisca Raquel De Vasconcelos Silveira. Towards the verbal decision analysis paradigm for implementable prioritization of software requirements. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
11/176.

Wang:2018:BCB

[710] Xuehai Wang, Feng Ding, Qingsheng Liu, and Chuntao Jiang. The bias compensation based parameter and state estimation for observability canonical state-space models with colored noise. Algorithms (Basel), 11 (11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/11/175.

Guo:2018:LCE

[711] Hongli Guo, Bin Li, Wei Li, Fengjuan Qiao, Xuewen Rong, and Yibin Li. Local coupled extreme learning machine based on particle swarm optimization. Algorithms (Basel), 11(11), November 2018. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/11/174.

Yamazaki:2018:IRC

[712] Koichi Yamazaki. Inapproximability of rank, clique, Boolean, and maximum induced matching-widths under small set expansion hypothesis. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
11/173.

Fei:2018:BGL

[713] Hongxiao Fei and Fengyun Tan. Bidirectional grid long short-term memory (BiGridLSTM): a method to address context-sensitivity and vanishing gradient. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/11/172.

Bakhach:2018:IDB

[714] Amer Bakhach, Venkata L. Raju Chinthalapati, Edward P. K. Tsang, and Abdul Rahman El Sayed. Intelligent dynamic backlash agent: a trading strategy based on the directional change framework. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/11/171.

Li:2018:MLV

[715] Zhixi Li and Vincent Tam. A machine learning view on momentum and reversal trading. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/11/170.

Lou:2018:PEC

[716] Xuyang Lou, Xu Cai, and Baotong Cui. Parameter estimation of a class of neural systems with limit cycles. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/11/11/169.

Govea-Vargas:2018:FOS

[717] Arturo Govea-Vargas, Rafael Castro-Linares, Manuel A. Duarte-Mermoud, Norelys Aguila-Camacho, and Gustavo E. Ceballos-Benavides. Fractional order sliding mode control of a class of second order perturbed nonlinear systems: Application to the trajectory tracking of a quadrotor. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/11/168.

Zhao:2018:OAP

[718] Jun Zhao, Xian Wang, Guanbin Gao, Jing Na, Hongping Liu, and Fujin Luan. Online adaptive parameter estimation for quadrotors. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/11/167.

Wang:2018:HGO

[719] Peiyu Wang, Chunrui Zhang, Liangkuan Zhu, and Chengcheng Wang. High-gain observer-based sliding-mode dynamic surface control for particleboard glue mixing and dosing system. Algorithms (Basel), 11(11), November 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/11/166.

Kari:2018:EAC

[720] Lila Kari, Stavros Konstantinidis, Steffen Kopecki, and Meng Yang. Efficient algorithms for computing the inner edit distance of a regular language via transducers. Algorithms (Basel), 11(11), November 2018. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/11/11/165.

Kang:2018:PRS

[721] Zhijiang Kang, Ze Deng, Wei Han, and Dongmei Zhang. Parallel reservoir simulation with OpenACC and domain decomposition. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/12/213.

Mramba:2018:EAE

[722] Lazarus K. Mramba and Salvador A. Gezan. Evaluating algorithm efficiency for optimizing experimental designs with correlated data. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
12/212.

Dogariu:2018:CBK

[723] Laura-Maria Dogariu, Silviu Ciochină, Constantin Paleologu, and Jacob Benesty. A connection between the Kalman filter and an optimized LMS algorithm for bilinear forms. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/11/12/211.

Zhang:2018:MOB

[724] Hongliang Zhang, Haijiang Ge, Ruilin Pan, and Yujuan Wu. Multiobjective bi-level programming for the energy-aware integration of flexible job shop scheduling and multi-row layout. Algorithms (Basel), 11 (12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/12/210.

Pelucchi:2018:HVS

[725] Mauro Pelucchi, Giuseppe Psaila, and Maurizio Toccu. Hadoop vs. Spark: Impact on performance of the Hammer query engine for open data corpora. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/12/209.

Arnau:2018:ULV

[726] Quim Arnau, Angel A. Juan, and Isabel Serra. On the use of learnheuristics in vehicle routing optimization problems with dynamic inputs. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
12/208.

Dritsas:2018:TCN

[727] Elias Dritsas, Maria Trigka, Panagiotis Gerolymatos, and Spyros Sioutas. Trajectory clustering and k-NN for robust privacy preserving spatiotemporal databases. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/12/207.

Miramontes:2018:ODI

[728] Ivette Miramontes, Juan Carlos Guzman, Patricia Melin, and German Prado-Arechiga. Optimal design of interval type-2 fuzzy heart rate level classification systems using the bird swarm algorithm. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/11/12/206.

Zhou:2018:OSM

[729] Chen Zhou, Xinhui Liu, Wei Chen, Feixiang Xu, and Bingwei Cao. Optimal sliding mode control for an active suspension system based on a genetic algorithm. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/12/205.

Nguyen:2018:NMR

[730] ManhCuong Nguyen, Shufang Zhang, and Xiaoye Wang. A novel method for risk assessment and simulation of collision avoidance for vessels based on AIS. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/12/204.

Tong:2018:FDA

[731] Xiaozhong Tong, Yujun Guo, and Wei Xie. Finite difference algorithm on non-uniform meshes for modeling 2D magnetotelluric responses. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
12/203.

Chen:2018:CNA

[732] Jun Chen and Edward P. K. Tsang. Classification of normal and abnormal regimes in financial markets. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/12/202.

Yu:2018:FAT

[733] Hancheng Yu, Haibao Qin, and Maoting Peng. A fast approach to texture-less object detection based on orientation compressing map and discriminative regional weight. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/12/201.

Ardagna:2018:SIA

[734] Danilo Ardagna, Claudia Canali, and Riccardo Lancellotti. Special issue on algorithms for the resource management of large scale infrastructures. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/11/
12/200.

Livieris:2018:DSS

[735] Ioannis E. Livieris, Theodore Kotsilieris, Ioannis Dimopoulos, and Panagiotis Pintelas. Decision support software for forecasting patient’s length of stay. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/12/199.

Liu:2018:DIA

[736] Hanbing Liu, Xin He, and Yubo Jiao. Damage identification algorithm of hinged joints for simply supported slab bridges based on modified hinge plate method and artificial bee colony algorithms. Algorithms (Basel), 11 (12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/12/198.

Li:2018:PEV

[737] Shengfeng Li and Yi Dong. Parametric estimation in the Vasicek-type model driven by sub-fractional Brownian motion. Algorithms (Basel), 11 (12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/12/197.

Koniaris:2018:SHA

[738] Marios Koniaris, George Papastefanatos, and Ioannis Anagnostopoulos. Solon: a holistic approach for modelling, managing and mining legal sources. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/12/196.

Gamby:2018:CHA

[739] Ask Neve Gamby and Jyrki Katajainen. Convex-hull algorithms: Implementation, testing, and experimentation. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/11/12/195.

Gonen:2018:NEA

[740] Yaron Gonen, Ehud Gudes, and Kirill Kandalov. New and efficient algorithms for producing frequent itemsets with the Map-Reduce framework. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/12/194.

Wang:2018:FMN

[741] Yuchuang Wang, Guoyou Shi, and Xiaotong Sun. A forecast model of the number of containers for containership voyage. Algorithms (Basel), 11 (12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/11/12/193.

Qu:2018:SFR

[742] Hongquan Qu, Meihan Wang, Changnian Zhang, and Yun Wei. A study on faster R-CNN-based subway pedestrian detection with ACE enhancement. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
11/12/192.

Li:2018:MAL

[743] Chen Li, Annisa Annisa, Asif Zaman, Mahboob Qaosar, Saleh Ahmed, and Yasuhiko Morimoto. MapReduce algorithm for location recommendation by using area skyline query. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/11/12/191.

Nghiem:2018:BTP

[744] Peter P. Nghiem. Best trade-off point method for efficient resource provisioning in spark. Algorithms (Basel), 11(12), December 2018. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/11/12/190.

Plakandaras:2019:EGU

[745] Vasilios Plakandaras, Periklis Gogas, and Theophilos Papadimitriou. The effects of geopolitical uncertainty in forecasting financial markets: a machine learning approach. Algorithms (Basel), 12(1), January 2019. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/12/1/1.

Comminiello:2019:SSP

[746] Danilo Comminiello, Michele Scarpiniti, Luis A. Azpicueta-Ruiz, and Aurelio Uncini. Steady-state performance of an adaptive combined MISO filter using the multichannel affine projection algorithm. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/1/2.

Hu:2019:AOQ

[747] Chunhe Hu, Yu Xia, and Junguo Zhang. Adaptive operator quantumbehaved pigeon-inspired optimization algorithm with application to UAV path planning. Algorithms (Basel), 12(1), January 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/1/3.

Mendonca:2019:FCD

[748] Marcele O. K. Mendonça, Jonathas O. Ferreira, Christos G. Tsinos, Paulo S. R. Diniz, and Tadeu N. Ferreira. On fast converging dataselective adaptive filtering. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/1/4.

Pericini:2019:MMA

[749] Matheus H. M. Pericini, Lucas G. M. Leite, Francisco H. De Carvalho-Junior, Javam C. Machado, and Cenez A. Rezende. MAPSkew: Metaheuristic approaches for partitioning skew in MapReduce. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/1/5.

Wang:2019:EDS

[750] Chengcheng Wang, Yaqiu Liu, and Peiyu Wang. Extraction and detection of surface defects in particleboards by tracking moving targets. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/1/6.

Guo:2019:SPN

[751] Di Guo, Zhangren Tu, Jiechao Wang, Min Xiao, Xiaofeng Du, and Xiaobo Qu. Salt and pepper noise removal with multi-class dictionary learning and L_{0} norm regularizations. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/1/7.

Zhang:2019:RVT

[752] Wancheng Zhang, Yanmin Luo, Zhi Chen, Yongzhao Du, Daxin Zhu, and Peizhong Liu. A robust visual tracking algorithm based on spatialtemporal context hierarchical response fusion. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/1/8.

Castillo:2019:CSF

[753] Oscar Castillo, Fevrier Valdez, José Soria, Leticia Amador-Angulo, Patricia Ochoa, and Cinthia Peraza. Comparative study in fuzzy controller optimization using bee colony, differential evolution, and harmony search algorithms. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/1/9.

Ghawadri:2019:DIR

[754] Nizam Ghawadri, Norazak Senu, Firas Adel Fawzi, Fudziah Ismail, and Zarina Bibi Ibrahim. Diagonally implicit Runge–Kutta type method for directly solving special fourth-order ordinary differential equations with ill-posed problem of a beam on elastic foundation. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/1/10.

Lu:2019:FER

[755] Yang Lu, Shigang Wang, and Wenting Zhao. Facial expression recognition based on discrete separable shearlet transform and feature selection. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/1/11.

Xu:2019:ENR

[756] Guangluan Xu, Xiaoke Wang, Yang Wang, Daoyu Lin, Xian Sun, and Kun Fu. Edge-nodes representation neural machine for link prediction. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/1/12.

Ren:2019:DSB

[757] Shengbing Ren, Wanying Zhang, Hafiz Shahbaz Munir, and Lei Xia. Dissimilarity space based multi-source cross-project defect prediction. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/
1/13.

Qureshi:2019:HPF

[758] Imran Qureshi, Jun Ma, and Kashif Shaheed. A hybrid proposed fundus image enhancement framework for diabetic retinopathy. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/1/14.

Sato:2019:TOE

[759] Mayuko Sato, Yoshikazu Fukuyama, Tatsuya Iizaka, and Tetsuro Matsui. Total optimization of energy networks in a smart city by multipopulation global-best modified brain storm optimization with migration. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/1/

15. See correction [869].

Guo:2019:LEC

[760] Yecai Guo, Fei Ye, and Hao Gong. Learning an efficient convolution neural network for pansharpening. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/1/16.

Castillo:2019:STF

[761] Oscar Castillo, Patricia Melin, Fevrier Valdez, Jose Soria, Emanuel Ontiveros-Robles, Cinthia Peraza, and Patricia Ochoa. Shadowed type-2 fuzzy systems for dynamic parameter adaptation in harmony search and differential evolution algorithms. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/1/17.

Zhang:2019:NHA

[762] Xiaoxia Zhang, Xin Shen, and Ziqiao Yu. A novel hybrid ant colony optimization for a multicast routing problem. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/1/18.

Overland:2019:APR

[763] Indra Overland and Javlon Juraev. Algorithm for producing rankings based on expert surveys. Algorithms (Basel), 12(1), January 2019. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/12/1/19.

Lan:2019:RGC

[764] Yong-Hong Lan, Jun-Jun Xia, and Yue-Xiang Shi. Robust guaranteedcost preview repetitive control for polytopic uncertain discrete-time systems. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/1/20.

Office:2019:ARA

[765] Algorithms Editorial Office. Acknowledgement to reviewers of algorithms in 2018. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/1/21.

Xu:2019:GEC

[766] Xingkui Xu, Chunfeng Wu, Qingyu Hou, and Zhigang Fan. Gyro error compensation in optoelectronic platform based on a hybrid ARIMA– Elman model. Algorithms (Basel), 12(1), January 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/1/22.

Phillips:2019:FEM

[767] Charles A. Phillips, Kai Wang, Erich J. Baker, Jason A. Bubier, Elissa J. Chesler, and Michael A. Langston. On finding and enumerating maximal and maximum k-partite cliques in k-partite graphs. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/1/23.

Zhu:2019:PSC

[768] Ling Zhu and Jie Lin. A pricing strategy of e-commerce advertising cooperation in the Stackelberg game model with different market power structure. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/1/24.

Zhao:2019:PAA

[769] Su Zhao, Gang Huang, and Qi Zhu. Power allocation algorithm for an energy-harvesting wireless transmission system considering energy losses. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/1/25.

Mouratidis:2019:EDL

[770] Despoina Mouratidis and Katia Lida Kermanidis. Ensemble and deep learning for language-independent automatic selection of parallel data.

Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/1/26.

Maragoudakis:2019:DAS

[771] Manolis Maragoudakis. Data analysis, simulation and visualization for environmentally safe maritime data. Algorithms (Basel), 12(1), January 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/1/27.

Deabes:2019:FIE

[772] Wael Deabes. FPGA implementation of ECT digital system for imaging conductive materials. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/2/28.

Maguerra:2019:DEP

[773] Soufiane Maguerra, Azedine Boulmakoul, Lamia Karim, and Hassan Badir. A distributed execution pipeline for clustering trajectories based on a fuzzy similarity relation. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/2/29.

Sun:2019:EBD

[774] Chong Sun and Qin Sheng. An exploration of a balanced up-downwind scheme for solving Heston volatility model equations on variable grids. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/
2/30.

Liu:2019:PPH

[775] Jiangyi Liu, Chunping Wang, Wei Wang, and Zheng Li. Particle probability hypothesis density filter based on pairwise Markov chains. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/2/31.

Scire:2019:FCB

[776] Alessandro Scirè, Fabrizio Tropeano, Aris Anagnostopoulos, and Ioannis Chatzigiannakis. Fog-computing-based heartbeat detection and arrhythmia classification using machine learning. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/2/32.

Bi:2019:OSB

[777] Yang Bi, Xi’an Feng, and Yangmei Zhang. Optimized sonar broadband focused beamforming algorithm. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/2/33.

Hadfield:2019:QAO

[778] Stuart Hadfield, Zhihui Wang, Bryan O’Gorman, Eleanor G. Rieffel, Davide Venturelli, and Rupak Biswas. From the quantum approximate optimization algorithm to a quantum alternating operator ansatz. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/2/34.

Fang:2019:RQI

[779] Yujie Fang, Juan Chen, and Zhengxuan Xue. Research on quantitative investment strategies based on deep learning. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/2/35.

Zhang:2019:CGH

[780] Yanfeng Zhang, Yunbao Huang, Haiyan Li, Pu Li, and Xi’an Fan. Conjugate gradient hard thresholding pursuit algorithm for sparse signal recovery. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/2/36.

Hu:2019:SDL

[781] Zhigang Hu, Hui Kang, and Meiguang Zheng. Stream data load prediction for resource scaling using online support vector regression. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/2/37.

Yao:2019:TLR

[782] Zhihong Yao, Yibing Wang, Wei Xiao, Bin Zhao, and Bo Peng. A twolevel rolling optimization model for real-time adaptive signal control. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/
2/38.

Belhaiza:2019:HAL

[783] Slim Belhaiza. A hybrid adaptive large neighborhood heuristic for a real-life dial-a-ride problem. Algorithms (Basel), 12(2), February 2019.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/2/39.

Sun:2019:IUB

[784] Shunkai Sun, Jianping Hu, Jie Li, Ruidong Liu, Meng Shu, and Yang Yang. An INS-UWB based collision avoidance system for AGV. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/2/40.

Chen:2019:CCD

[785] Jie Chen, Gang Yang, and Meng Yang. Computation of compact distributions of discrete elements. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/2/41.

Zeng:2019:DOV

[786] Wei Zeng, Tao Ren, Lijun Yu, and Jingjing Huang. Design optimization of a VX gasket structure for a subsea connector based on the kriging surrogate model-NSGA-II algorithm considering the load randomness. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/
2/42.

Chen:2019:IGA

[787] Leiwen Chen, Yingming Wang, and Geng Guo. An improved genetic algorithm for emergency decision making under resource constraints based on prospect theory. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/2/43.

Bian:2019:ISP

[788] Chentong Bian, Tong Zhu, Guodong Yin, and Liwei Xu. Integrated speed planning and friction coefficient estimation algorithm for intelligent electric vehicles. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/2/44.

Zhao:2019:HAR

[789] Mengting Zhao and Yuwei Lu. A heuristic approach for a real-world electric vehicle routing problem. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/2/45.

He:2019:PID

[790] Hao He, Jiaxiang Zhao, and Guiling Sun. The prediction of intrinsically disordered proteins based on feature selection. Algorithms (Basel), 12 (2), February 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/2/46.

Chen:2019:RTC

[791] Hao-Xiang Chen, Ying Nan, and Yi Yang. Real-time conflict resolution algorithm for multi-UAV based on model predict control. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/2/47.

Zhao:2019:SOE

[792] Ming Zhao and Ke Zhou. Selective offloading by exploiting ARIMA–BP for energy optimization in mobile edge computing networks. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/2/48.

Liu:2019:SCW

[793] Qiuming Liu, Shumin Liu, Chunshui Zeng, Xiaohong Qiu, and He Xiao. Secrecy control of wireless networks with finite encoding blocklength. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/
2/49.

Lin:2019:RPA

[794] Mugang Lin, Jianxin Wang, Qilong Feng, and Bin Fu. Randomized parameterized algorithms for the kidney exchange problem. Algorithms (Basel), 12(2), February 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/2/50.

Ji:2019:ODC

[795] Qingge Ji, Jie Huang, Wenjie He, and Yankui Sun. Optimized deep convolutional neural networks for identification of macular diseases from optical coherence tomography images. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/3/51.

Nakamura:2019:SEF

[796] Kengo Nakamura and Kunihiko Sadakane. Space-efficient fully dynamic DFS in undirected graphs. Algorithms (Basel), 12(3), March 2019. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/12/3/52.

Fernandez-Baca:2019:TCI

[797] David Fernández-Baca and Lei Liu. Tree compatibility, incomplete directed perfect phylogeny, and dynamic graph connectivity: An experimental study. Algorithms (Basel), 12(3), March 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/3/53.

Templos-Santos:2019:PTP

[798] Juan Luis Templos-Santos, Omar Aguilar-Mejia, Edgar Peralta-Sanchez, and Raul Sosa-Cortez. Parameter tuning of PI control for speed regulation of a PMSM using bio-inspired algorithms. Algorithms (Basel), 12 (3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/3/54.

Cao:2019:DOA

[799] Bing wei Cao, Xin hui Liu, Wei Chen, Yong Zhang, and Ai min Li. Depth optimization analysis of articulated steering hinge position based on genetic algorithm. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/3/55.

Li:2019:MAE

[800] Wei Li. Matrix adaptation evolution strategy with multi-objective optimization for multimodal optimization. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/3/56.

Nguyen:2019:PER

[801] Van Giao Nguyen, Xuexun Guo, Chengcai Zhang, and Xuan Khoa Tran. Parameter estimation, robust controller design and performance analysis for an electric power steering system. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/3/57.

Dumitrescu:2019:SSH

[802] Adrian Dumitrescu. A selectable sloppy heap. Algorithms (Basel), 12 (3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/3/58.

Almonacid:2019:APR

[803] Boris Almonacid, Fabián Aspée, and Francisco Yimes. Autonomous population regulation using a multi-agent system in a prey–predator model that integrates cellular automata and the African buffalo optimization metaheuristic. Algorithms (Basel), 12(3), March 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/3/59.

Pfander:2019:HDB

[804] David Pfander, Gregor Daiß, and Dirk Pflüger. Heterogeneous distributed big data clustering on sparse grids. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/3/60.

Wang:2019:NCA

[805] Yechuang Wang, Zhihua Cui, and Wuchao Li. A novel coupling algorithm based on glowworm swarm optimization and bacterial foraging algorithm for solving multi-objective optimization problems. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/3/61.

Ye:2019:MVN

[806] Zhonglin Ye, Haixing Zhao, Ke Zhang, and Yu Zhu. Multi-view network representation learning algorithm research. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/3/62.

Xu:2019:SCA

[807] Xiaodan Xu, Zhifeng Bai, and Yuanyuan Shao. Synchronization control algorithm of double-cylinder forging hydraulic press based on fuzzy neural network. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/3/63.

Livieris:2019:WVE

[808] Ioannis E. Livieris, Andreas Kanavos, Vassilis Tampakas, and Panagiotis Pintelas. A weighted voting ensemble self-labeled algorithm for the detection of lung abnormalities from X-rays. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/3/64.

Liu:2019:HPC

[809] Jiao Liu, Guoyou Shi, and Kaige Zhu. High-precision combined tidal forecasting model. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/3/65.

Yang:2019:ADC

[810] Xiaohui Yang, Xiaolong Zhang, Shaoping Xu, Yihui Ding, Kun Zhu, and Peter Xiaoping Liu. An approach to the dynamics and control of uncertain robot manipulators. Algorithms (Basel), 12(3), March 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/3/66.

Saraiva:2019:TPA

[811] Rommel Dias Saraiva, Napoleão Nepomuceno, and Plácido Rogério Pinheiro. A two-phase approach for single container loading with weakly heterogeneous boxes. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/4/67.

Fan:2019:CLO

[812] Shu Fan. A cross-layer optimization QoS scheme in wireless multimedia sensor networks. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/4/68.

Chinthalapati:2019:SIA

[813] V. L. Raju Chinthalapati and Edward Tsang. Special issue on algorithms in computational finance. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/4/69.

Zhang:2019:TAI

[814] Jiarui Zhang, Gang Wang, and Yafei Song. Task assignment of the improved contract net protocol under a multi-agent system. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/4/70.

Zhang:2019:PCF

[815] Jinghua Zhang and Ze Dong. Parameter combination framework for the differential evolution algorithm. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/4/71.

Chen:2019:IAA

[816] Weijia Chen and Yancai Xiao. An improved ABC algorithm and its application in bearing fault diagnosis with EEMD. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/4/72.

Hendrian:2019:PPM

[817] Diptarama Hendrian, Yohei Ueki, Kazuyuki Narisawa, Ryo Yoshinaka, and Ayumi Shinohara. Permuted pattern matching algorithms on multitrack strings. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/4/73.

DEmidio:2019:BGT

[818] Mattia D’Emidio, Gabriele Di Stefano, and Alfredo Navarra. Bamboo garden trimming problem: Priority schedulings. Algorithms (Basel), 12 (4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/4/74.

Xiao:2019:PFD

[819] Runing Xiao and Jinzhi Zhou. Pulmonary fissure detection in 3D CT images using a multiple section model. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/4/75.

Baldoni:2019:PAT

[820] Matteo Baldoni, Cristina Baroglio, Roberto Micalizio, and Stefano Tedeschi. Programming agents by their social relationships: A commitment-based approach. Algorithms (Basel), 12(4), April 2019. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/12/4/76.

Vyskocil:2019:EEC

[821] Tomas Vyskocil and Hristo Djidjev. Embedding equality constraints of optimization problems into a quantum annealer. Algorithms (Basel), 12 (4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/4/77.

Kulekci:2019:ANU

[822] Muhammed Oğuzhan Külekci and Yasin Öztürk. Applications of nonuniquely decodable codes to privacy-preserving high-entropy data representation. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/4/78.

Hanna:2019:SAT

[823] Nader Hanna and Deborah Richards. Speech act theory as an evaluation tool for human–agent communication. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/4/79.

Wang:2019:ISS

[824] Yanjiao Wang and Tianlin Du. An improved squirrel search algorithm for global function optimization. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/4/80.

Gartner:2019:DSD

[825] Fabian Gärtner and Peter F. Stadler. Direct superbubble detection. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/4/81.

Zhang:2019:IEC

[826] Zhiqiang Zhang, Rong Huang, Fang Han, and Zhijie Wang. Image error concealment based on deep neural network. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/4/82.

Song:2019:OAF

[827] Chunfeng Song, Mei Wang, Xuebin Qin, Pai Wang, and Bao Liu. The optimization algorithm of the forced current cathodic protection base on simulated annealing. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/4/83.

Liu:2019:HBA

[828] Ting Liu, Chao Tan, Zhongbin Wang, Jing Xu, Yiqiao Man, and Tuo Wang. Horizontal bending angle optimization method for scraper conveyor based on improved bat algorithm. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/4/84.

Livieris:2019:FER

[829] Ioannis E. Livieris. Forecasting economy-related data utilizing weightconstrained recurrent neural networks. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/4/85.

Hwang:2019:KFB

[830] Hyeongjin Hwang, Jehwon Lee, Sangjune Eum, and Kanghyun Nam. Kalman-filter-based tension control design for industrial roll-to-roll system. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/4/
86.

Gurski:2019:OCR

[831] Frank Gurski, Dominique Komander, and Carolin Rehs. Oriented coloring on recursively defined digraphs. Algorithms (Basel), 12(4), April 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/4/87.

Khan:2019:REI

[832] Talha Ali Khan and Sai Ho Ling. Review on electrical impedance tomography: Artificial intelligence methods and its applications. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/5/88.

Souliotis:2019:APF

[833] Georgios Souliotis and Basil Papadopoulos. An algorithm for producing fuzzy negations via conical sections. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/5/89.

Lagunes:2019:MMC

[834] Marylu L. Lagunes, Oscar Castillo, Fevrier Valdez, and Jose Soria. Multimetaheuristic competitive model for optimization of fuzzy controllers. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/5/90.

Lancia:2019:FFD

[835] Giuseppe Lancia and Marcello Dalpasso. FASTSET: a fast data structure for the representation of sets of integers. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/5/91.

Wang:2019:OFE

[836] Song Wang and Zengfu Wang. Optical flow estimation with occlusion detection. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/5/92.

Su:2019:PCC

[837] Na Su and Qi Zhu. Power control and channel allocation algorithm for energy harvesting D2D communications. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/5/93.

Wu:2019:CNL

[838] Zongsheng Wu and Ru Xue. A cyclical non-linear inertia-weighted teaching–learning-based optimization algorithm. Algorithms (Basel), 12 (5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/5/94.

Koishi:2019:SDE

[839] Yasutake Koishi, Shuichi Ishida, Tatsuo Tabaru, and Hiroyuki Miyamoto. A source domain extension method for inductive transfer learning based on flipping output. Algorithms (Basel), 12(5), May 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/5/95.

Bekkouch:2019:TLN

[840] Imad Eddine Ibrahim Bekkouch, Youssef Youssry, Rustam Gafarov, Adil Khan, and Asad Masood Khattak. Triplet loss network for unsupervised domain adaptation. Algorithms (Basel), 12(5), May 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/5/96.

Zheng:2019:NMA

[841] Song Zheng, Chao Bi, and Yilin Song. A new method of applying data engine technology to realize neural network control. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/5/97.

Ren:2019:FSF

[842] Xingyue Ren, Fangjie Xiong, Ke Qu, and Norimi Mizutani. Free surface flow simulation by a viscous numerical cylindrical tank. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/5/98.

Pirpinia:2019:EML

[843] Kleopatra Pirpinia, Peter A. N. Bosman, Jan-Jakob Sonke, Marcel van Herk, and Tanja Alderliesten. Evolutionary machine learning for multiobjective class solutions in medical deformable image registration. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/5/99.

Kizilay:2019:VBI

[844] Damla Kizilay, Mehmet Fatih Tasgetiren, Quan-Ke Pan, and Liang Gao. A variable block insertion heuristic for solving permutation flow shop scheduling problem with makespan criterion. Algorithms (Basel), 12 (5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/5/100.

Schimmack:2019:ADE

[845] Manuel Schimmack and Paolo Mercorelli. An adaptive derivative estimator for fault-detection using a dynamic system with a suboptimal parameter. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/5/101.

Maxim:2019:TFD

[846] Anca Maxim, Riccardo Ferracuti, and Clara M. Ionescu. A theoretical framework to determine RHP zero dynamics in sequential interacting sub-systems. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/5/102.

Wang:2019:INN

[847] Meng Wang, Chuang-Bai Xiao, Zhen-Hu Ning, Jing Yu, Ya-Hao Zhang, and Jin Pang. Improved neural networks based on mutual information via information geometry. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/5/103.

Clements:2019:BPE

[848] Wyatt Clements, Costas Busch, Limeng Pu, Daniel Smith, and Hsiao-Chun Wu. Balanced parallel exploration of orthogonal regions. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/5/104.

Lin:2019:AAC

[849] Hsiung-Cheng Lin and Chung-Hao Cheng. Achievement of automatic copper wire elongation system. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/5/105.

Chen:2019:IND

[850] Jeang-Kuo Chen and Wei-Zhe Lee. An introduction of NoSQL databases based on their categories and application industries. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/5/106.

Qin:2019:POT

[851] Jiwei Qin, Liangli Ma, and Qing Liu. Pruning optimization over threshold-based historical continuous query. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/5/107.

Paraskevopoulos:2019:RTA

[852] Georgios Paraskevopoulos, Evaggelos Spyrou, Dimitrios Sgouropoulos, Theodoros Giannakopoulos, and Phivos Mylonas. Real-time arm gesture recognition using 3D skeleton joint data. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/5/108.

Favati:2019:APG

[853] Paola Favati, Grazia Lotti, Ornella Menchi, and Francesco Romani. An adaptive procedure for the global minimization of a class of polynomial functions. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/5/109.

Garcia-Retuerta:2019:CTV

[854] David García-Retuerta, Álvaro Bartolomé, Pablo Chamoso, and Juan Manuel Corchado. Counter-terrorism video analysis using hashbased algorithms. Algorithms (Basel), 12(5), May 2019. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/5/110.

Perez:2019:HAR

[855] Marco Antonio Juárez Pérez, Rodolfo Eleazar Pérez Loaiza, Perfecto Malaquias Quintero Flores, Oscar Atriano Ponce, and Carolina Flores Peralta. A heuristic algorithm for the routing and scheduling problem with time windows: a case study of the automotive industry in Mexico. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/5/111.

Zhao:2019:CAD

[856] Yulin Zhao, Donghui Wang, and Leiou Wang. Convolution accelerator designs using fast algorithms. Algorithms (Basel), 12(5), May 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/5/112.

Stefanescu:2019:SBR

[857] Razvan Stefanescu, Jason Hite, Jared Cook, Ralph C. Smith, and John Mattingly. Surrogate-based robust design for a non-smooth radiation source detection problem. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/6/113.

Deon:2019:PTG

[858] Aleksei F. Deon and Yulian A. Menyaev. Poisson twister generator by cumulative frequency technology. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/6/114.

Yu:2019:CBS

[859] Tianming Yu, Jianhua Yang, and Wei Lu. Combining background subtraction and convolutional neural network for anomaly detection in pumping-unit surveillance. Algorithms (Basel), 12(6), June 2019. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/12/6/115.

Ryabko:2019:TUD

[860] Boris Ryabko. Time-universal data compression. Algorithms (Basel), 12 (6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/6/116.

El-Amin:2019:INS

[861] Mohamed F. El-Amin. Iterative numerical scheme for non-isothermal two-phase flow in heterogeneous porous media. Algorithms (Basel), 12 (6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/6/117.

Darmawahyuni:2019:DLR

[862] Annisa Darmawahyuni, Siti Nurmaini, Sukemi, Wahyu Caesarendra, Vicko Bhayyu, M. Naufal Rachmatullah, and Firdaus. Deep learning with a recurrent network structure in the sequence modeling of imbalanced data for ECG-rhythm classifier. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/6/118.

Hasheminasab:2019:RFM

[863] Hamidreza Hasheminasab, Sarfaraz Hashemkhani Zolfani, Mahdi Bitarafan, Prasenjit Chatterjee, and Alireza Abhaji Ezabadi. The role of façade materials in blast-resistant buildings: an evaluation based on fuzzy Delphi and fuzzy EDAS. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/6/119.

Zhang:2019:IPP

[864] Tao Zhang, Yue Wang, Xin Jin, and Shan Lu. Integration of production planning and scheduling based on RTN representation under uncertainties. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/6/
120.

Radac:2019:LOR

[865] Mircea-Bogdan Radac and Timotei Lala. Learning output reference model tracking for higher-order nonlinear systems with unknown dynamics. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/6/
121.

Chen:2019:HAN

[866] Pei-Yin Chen and Jih-Jeng Huang. A hybrid autoencoder network for unsupervised image clustering. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/6/122.

Ciaramella:2019:RCV

[867] Angelo Ciaramella and Antonino Staiano. On the role of clustering and visualization techniques in gene microarray data. Algorithms (Basel), 12 (6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/6/123.

Ghuman:2019:LFA

[868] Sukhpal Singh Ghuman, Emanuele Giaquinta, and Jorma Tarhio. Lyndon factorization algorithms for small alphabets and run-length encoded strings. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/6/
124.

Sato:2019:CSM

[869] Mayuko Sato, Yoshikazu Fukuyama, Tatsuya Iizaka, and Tetsuro Matsui. Correction: Sato, M., et al. Total Optimization of Energy Networks in a Smart City by Multi-Population Global-Best Modified Brain Storm Optimization with Migration, Algorithms 2019, 12, 15. Algorithms (Basel), 12(6), June 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/6/125. See [759].

Wang:2019:NRR

[870] Bin Wang, Li Wang, Hao Yu, and Fengming Xin. A new regularized reconstruction algorithm based on compressed sensing for the sparse underdetermined problem and applications of one-dimensional and twodimensional signal recovery. Algorithms (Basel), 12(7), July 2019. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/12/7/126.

Angriman:2019:GEA

[871] Eugenio Angriman, Alexander van der Grinten, Moritz von Looz, Henning Meyerhenke, Martin Nöllenburg, Maria Predari, and Charilaos Tzovas. Guidelines for experimental algorithmics: a case study in network analysis. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/7/
127.

Yu:2019:RBS

[872] Tianming Yu, Jianhua Yang, and Wei Lu. Refinement of backgroundsubtraction methods based on convolutional neural network features for dynamic background. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/7/128.

Zhang:2019:HHA

[873] Chunmiao Zhang, Yanwei Zhao, and Longlong Leng. A hyper heuristic algorithm to solve the low-carbon location routing problem. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/7/129.

Duong:2019:NCQ

[874] Dinh Trieu Duong, Huy Phi Cong, and Xiem Hoang Van. A novel consistent quality driven for JEM based distributed video coding. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/7/130.

Stoican:2019:ADL

[875] Florin Stoican and Paul Irofti. Aiding dictionary learning through multiparametric sparse representation. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/7/131.

Pu:2019:DWL

[876] Cuiping Pu, Yicheng Zhu, and Jianbo Su. Drum water level control based on improved ADRC. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/7/132.

Tanackov:2019:NMM

[877] Ilija Tanackov, Olegas Prentkovskis, Zarko Jevtić, Gordan Stojić, and Pamela Ercegovac. A new method for Markovian adaptation of the non-Markovian queueing system using the hidden Markov model. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/7/133.

Wang:2019:ELA

[878] Yanjiao Wang and Xintian Jiang. An enhanced lightning attachment procedure optimization algorithm. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/7/134.

Cai:2019:BMD

[879] Sheng Cai, Pei-Zhong Liu, Yan-Min Luo, Yong-Zhao Du, and Jia-Neng Tang. Breast microcalcification detection algorithm based on contourlet and ASVM. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/7/135.

Jansson:2019:ESI

[880] Jesper Jansson. Editorial: Special issue on efficient data structures. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/7/136.

Gogas:2019:MNM

[881] Periklis Gogas, Theophilos Papadimitriou, and Emmanouil Sofianos. Money neutrality, monetary aggregates and machine learning. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/7/137.

Ji:2019:QBN

[882] Zheng Ji, Xu Cai, and Xuyang Lou. A quantum-behaved neurodynamic approach for nonconvex optimization with constraints. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/7/138.

Yazdi:2019:CRM

[883] Amir Karbassi Yazdi, Thomas Hanne, Yong J. Wang, and Hui-Ming Wee. A credit rating model in a fuzzy inference system environment. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/7/139.

Mendes:2019:PAR

[884] Mateus Mendes, Jorge Almeida, Hajji Mohamed, and Rudi Giot. Projected augmented reality intelligent model of a city area with path optimization. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/7/140.

Cubukcuoglu:2019:OSA

[885] Cemre Cubukcuoglu, Berk Ekici, Mehmet Fatih Tasgetiren, and Sevil Sariyildiz. OPTIMUS: Self-adaptive differential evolution with ensemble of mutation strategies for grasshopper algorithmic modeling. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/12/7/141.

Li:2019:NBG

[886] Qinghai Li and Chang Wu Yu. New bipartite graph techniques for irregular data redistribution scheduling. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/7/142.

Kamal:2019:BLM

[887] Murshid Kamal, Srikant Gupta, Prasenjit Chatterjee, Dragan Pamucar, and Zeljko Stevic. Bi-level multi-objective production planning problem with multi-choice parameters: a fuzzy goal programming algorithm. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/7/143.

Bagyaveereswaran:2019:STT

[888] Veeramani Bagyaveereswaran, Subramaniam Umashankar, and Pachiyappan Arulmozhivarman. Simulation tool for tuning and performance analysis of robust, tracking, disturbance rejection and aggressiveness controller. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/7/144.

Qu:2019:SSB

[889] Hongquan Qu, Zhanli Fan, Shuqin Cao, Liping Pang, Hao Wang, and Jie Zhang. A study on sensitive bands of EEG data under different mental workloads. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/7/145.

Trifan:2019:HMM

[890] Razvan-Florentin Trifan, Andrei-Alexandru Enescu, and Constantin Paleologu. Hybrid MU-MIMO precoding based on K-means user clustering. Algorithms (Basel), 12(7), July 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/7/146.

Rusu:2019:ODS

[891] Alexandru-George Rusu, Silviu Ciochina, Constantin Paleologu, and Jacob Benesty. An optimized differential step-size LMS algorithm. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/8/147.

Horla:2019:VCA

[892] Dariusz Horla. Variational calculus approach to optimal interception task of a ballistic missile in 1D and 2D cases. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/8/148.

Faict:2019:MGI

[893] Thomas Faict, Erik H. D’Hollander, and Bart Goossens. Mapping a guided image filter on the HARP reconfigurable architecture using OpenCL. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/8/149.

Bergadano:2019:DDP

[894] Francesco Bergadano, Fabio Carretto, Fabio Cogno, and Dario Ragno. Defacement detection with passive adversaries. Algorithms (Basel), 12 (8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/8/150.

Zhang:2019:SID

[895] Meixiang Zhang, Satya Chan, and Sooyoung Kim. Soft iterative decoding algorithms for rateless codes in satellite systems. Algorithms (Basel), 12 (8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/8/151.

Haddadene:2019:BVR

[896] Syrine Roufaida Ait Haddadene, Nacima Labadie, and Caroline Prodhon. Bicriteria vehicle routing problem with preferences and timing constraints in home health care services. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/8/152.

Finbow:2019:GT

[897] Stephen Finbow and Christopher M. van Bommel. γ-graphs of trees. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/8/153.

Vestias:2019:SCN

[898] Mário P. Véstias. A survey of convolutional neural networks on edge with reconfigurable computing. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/8/154.

Zhang:2019:RMA

[899] Yuan Zhang and Liyi Zhang. A rigid motion artifact reduction method for CT based on blind deconvolution. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/8/155.

Pitolli:2019:CMN

[900] Francesca Pitolli. A collocation method for the numerical solution of nonlinear fractional dynamical systems. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/8/156.

Farago:2019:SDS

[901] András Faragó and Zohre R. Mojaveri. In search of the densest subgraph. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/8/157.

Nepomuceno:2019:FRA

[902] Napoleão Nepomuceno, Ricardo Barboza Saboia, and Plácido Rogério Pinheiro. A fast randomized algorithm for the heterogeneous vehicle routing problem with simultaneous pickup and delivery. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/8/158.

Furuya:2019:CCN

[903] Isamu Furuya and Takuya Kida. Compaction of Church numerals. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/8/159.

Wedyan:2019:NVS

[904] Mohammad Wedyan, Alessandro Crippa, and Adel Al-Jumaily. A novel virtual sample generation method to overcome the small sample size problem in computer aided medical diagnosing. Algorithms (Basel), 12 (8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/8/160.

Behera:2019:DCA

[905] Ranjan Kumar Behera, Santanu Kumar Rath, Sanjay Misra, Robertas Damasevicius, and Rytis Maskeliunas. Distributed centrality analysis of social network data using MapReduce. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/8/161.

Aydin:2019:DBP

[906] Kevin Aydin, MohammadHossein Bateni, and Vahab Mirrokni. Distributed balanced partitioning via linear embedding. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/8/162.

Winnicka:2019:IUB

[907] Alicja Winnicka and Karolina Kesik. Idea of using blockchain technique for choosing the best configuration of weights in neural networks. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/8/163.

Eggleton:2019:EPS

[908] Roger B. Eggleton. Equisum partitions of sets of positive integers. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/8/164.

Minami:2019:AMA

[909] Kazuhiro Minami and Yutaka Abe. Algorithmic matching attacks on optimally suppressed tabular data. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/8/165.

Siddique:2019:MAV

[910] Md. Anisuzzaman Siddique, Hao Tian, Mahboob Qaosar, and Yasuhiko Morimoto. MapReduce algorithm for variants of skyline queries: Skyband and dominating queries. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/8/166.

Rosinova:2019:LPR

[911] Danica Rosinová and Mária Hypiusová. LMI pole regions for a robust discrete-time pole placement controller design. Algorithms (Basel), 12 (8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/8/167.

Crnkovic:2019:CTC

[912] Dean Crnković, Andrea Svob, and Vladimir D. Tonchev. Cyclotomic trace codes. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/8/168.

Lei:2019:SAA

[913] Zhao Lei, Hu Lai, Zhang Hua, and Chen Hua. Structural analysis and application of non-standard components based on genetic algorithm. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/8/169.

Clazzer:2019:PLC

[914] Federico Clazzer, Balázs Matuz, Sachini Jayasooriya, Mahyar Shirvanimoghaddam, and Sarah J. Johnson. Protograph LDPC code design for asynchronous random access. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/8/170.

Bashiri:2019:AML

[915] Fereshteh S. Bashiri, Reihaneh Rostami, Peggy Peissig, Roshan M. D’Souza, and Zeyun Yu. An application of manifold learning in global shape descriptors. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/8/171.

Bannach:2019:PAD

[916] Max Bannach and Sebastian Berndt. Practical access to dynamic programming on tree decompositions. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/8/172.

Li:2019:LST

[917] Zhen Li, Tao Tang, and Chunhai Gao. Long short-term memory neural network applied to train dynamic model and speed prediction. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/8/173.

Sun:2019:NBR

[918] Yunshan Sun, Liyi Zhang, Yanqin Li, and Juan Meng. A novel blind restoration and reconstruction approach for CT images based on sparse representation and hierarchical Bayesian-MAP. Algorithms (Basel), 12 (8), August 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/8/174.

Georgiou:2019:DHC

[919] Konstantinos Georgiou, Christos Makris, and Georgios Pispirigos. A distributed hybrid community detection methodology for social networks. Algorithms (Basel), 12(8), August 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/8/175.

Mansouri:2019:OES

[920] Amirsalar Mansouri, Sanjay P. Singh, and Khalid Sayood. Online EEG seizure detection and localization. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/9/176.

Budiaji:2019:SMP

[921] Weksi Budiaji and Friedrich Leisch. Simple K-medoids partitioning algorithm for mixed variable data. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/9/177.

Dumitrescu:2019:ASD

[922] Bogdan Dumitrescu and Ciprian Doru Giurcaneanu. Adaptive-size dictionary learning using information theoretic criteria. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/9/178.

Li:2019:FPL

[923] Zongyang Li, Yefei Wang, and Le Wang. A fast particle-locating method for the arbitrary polyhedral mesh. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/12/9/179.

Azais:2019:NEE

[924] Romain Azaïs. Nearest embedded and embedding self-nested trees. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/12/9/180.

Teng:2019:FAL

[925] Zhongming Teng and Linzhang Lu. A FEAST algorithm for the linear response eigenvalue problem. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/12/9/181.

Ghoniem:2019:NHG

[926] Rania M. Ghoniem, Nawal Alhelwa, and Khaled Shaalan. A novel hybrid genetic-whale optimization model for ontology learning from Arabic text. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/12/9/
182.

Li:2019:IWM

[927] Kexin Li, Jun Wang, and Dawei Qi. An intelligent warning method for diagnosing underwater structural damage. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/9/183.

Chen:2019:FDR

[928] Yinsheng Chen, Tinghao Zhang, Wenjie Zhao, Zhongming Luo, and Kun Sun. Fault diagnosis of rolling bearing using multiscale amplitudeaware permutation entropy and random forest. Algorithms (Basel), 12 (9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/9/184.

Luo:2019:CTI

[929] Yu-Juan Luo, Cheng-Lin Liu, and Guang-Ye Liu. Consensus tracking by iterative learning control for linear heterogeneous multiagent systems based on fractional-power error signals. Algorithms (Basel), 12 (9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/9/185.

Aziz:2019:SSM

[930] Fayeem Aziz, Aaron S. W. Wong, and Stephan Chalup. Semi-supervised manifold alignment using parallel deep autoencoders. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/9/186.

Djidjev:2019:UGP

[931] Hristo N. Djidjev, Georg Hahn, Susan M. Mniszewski, Christian F. A. Negre, and Anders M. N. Niklasson. Using graph partitioning for scalable distributed quantum molecular dynamics. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/9/187.

deHaan:2019:CPP

[932] Ronald de Haan and Stefan Szeider. A compendium of parameterized problems at higher levels of the polynomial hierarchy. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/9/188.

Creignou:2019:PEM

[933] Nadia Creignou, Raïda Ktari, Arne Meier, Julian-Steffen Müller, Frédéric Olive, and Heribert Vollmer. Parameterised enumeration for modification problems. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/9/189.

Dourado:2019:PSB

[934] Jonas R. Dourado, Jordão Natal de Oliveira Júnior, and Carlos D. Maciel. Parallelism strategies for big data delayed transfer entropy evaluation. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
12/9/190.

Meindl:2019:FBI

[935] Bernhard Meindl and Matthias Templ. Feedback-based integration of the whole process of data anonymization in a graphical interface. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/9/191.

Shah:2019:CQD

[936] Syed Aizaz Ali Shah, Maximilian Stark, and Gerhard Bauch. Coarsely quantized decoding and construction of polar codes using the information bottleneck method. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/9/192.

Torres-Sospedra:2019:CSI

[937] Joaquín Torres-Sospedra and Patricio Nebot. Combining satellite images and cadastral information for outdoor autonomous mapping and navigation: a proof-of-concept study in citric groves. Algorithms (Basel), 12 (9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/12/9/193.

Shuai:2019:USL

[938] Liu Shuai, Liu Yuanning, Zhu Xiaodong, Zhang Kuo, Ding Tong, Li Xinlong, and Wang Chaoqun. Unsteady state lightweight Iris certification based on multi-algorithm parallel integration. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/9/194.

Zhao:2019:IAN

[939] Shiping Zhao, Yong Ma, and Dingxin Leng. An intelligent artificial neural network modeling of a magnetorheological elastomer isolator. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/12/9/195.

Gottesburen:2019:FBN

[940] Lars Gottesbüren, Michael Hamann, Tim Niklas Uhl, and Dorothea Wagner. Faster and better nested dissection orders for customizable contraction hierarchies. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/9/196.

Gotschel:2019:CCL

[941] Sebastian Götschel and Martin Weiser. Compression challenges in large scale partial differential equation solvers. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/9/197.

Hamann:2019:CBM

[942] Michael Hamann and Ben Strasser. Correspondence between multilevel graph partitions and tree decompositions. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/12/9/198.

Shrestha:2019:FAD

[943] Yash Raj Shrestha and Yongjie Yang. Fairness in algorithmic decisionmaking: Applications in multi-winner voting, machine learning, and recommender systems. Algorithms (Basel), 12(9), September 2019. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/12/9/199.

Kim:2020:IRU

[944] Kyoum Sun Kim and Jae Heon Yun. Image restoration using a fixedpoint method for a TVL2 regularization problem. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/1/1.

DEmidio:2020:JPA

[945] Mattia D’Emidio, Imran Khan, and Daniele Frigioni. Journey planning algorithms for massive delay-prone transit networks. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/1/2.

Narayanaswamy:2020:POU

[946] N. S. Narayanaswamy and R. Vijayaragunathan. Parameterized optimization in uncertain graphs — a survey and some results. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/1/3.

Sotskov:2020:TMJ

[947] Yuri N. Sotskov, Natalja M. Matsveichuk, and Vadzim D. Hatsura. Twomachine job-shop scheduling problem to minimize the makespan with uncertain job durations. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/4.

Pacheco-Valencia:2020:SCI

[948] Víctor Pacheco-Valencia, José Alberto Hernández, José María Sigarreta, and Nodari Vakhania. Simple constructive, insertion, and improvement heuristics based on the girding polygon for the Euclidean traveling salesman problem. Algorithms (Basel), 13(1), January 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/5.

Tarasov:2020:GMF

[949] Ilia Tarasov, Alain Haït, and Olga Battaïa. A generalized MILP formulation for the period-aggregated resource leveling problem with variable job duration. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/1/6.

Tadesse:2020:DSI

[950] Michael Mesfin Tadesse, Hongfei Lin, Bo Xu, and Liang Yang. Detection of suicide ideation in social media forums using deep learning. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/1/7.

Juan:2020:UBR

[951] Angel Alejandro Juan, Canan Gunes Corlu, Rafael David Tordecilla, Rocio de la Torre, and Albert Ferrer. On the use of biased-randomized algorithms for solving non-smooth optimization problems. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/1/8.

Werner:2020:SIE

[952] Frank Werner, Larysa Burtseva, and Yuri N. Sotskov. Special issue on exact and heuristic scheduling algorithms. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/1/9.

Shahid:2020:CAD

[953] Yumna Shahid and Minxiang Wei. Comparative analysis of different model-based controllers using active vehicle suspension system. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/1/10.

Espinoza:2020:NAF

[954] Jesús F. Espinoza, Rosalía Hernández-Amador, Héctor A. Hernández-Hernández, and Beatriz Ramonetti-Valencia. A numerical approach for the filtered generalized Cech complex. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/1/11.

Barbay:2020:OPF

[955] Jérémy Barbay. Optimal prefix free codes with partial sorting. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/1/12.

Cuzzocrea:2020:EEG

[956] Alfredo Cuzzocrea, Enzo Mumolo, and Giorgio Mario Grasso. An effective and efficient genetic-fuzzy algorithm for supporting advanced human-machine interfaces in big data settings. Algorithms (Basel), 13 (1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/1/13.

Ji:2020:ICL

[957] Jianjian Ji and Gang Yang. Image completion with large or edge-missing areas. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/1/14.

Zhen:2020:VOT

[958] Xinxin Zhen, Shumin Fei, Yinmin Wang, and Wei Du. A visual object tracking algorithm based on improved TLD. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/1/15.

Niyomubyeyi:2020:CSF

[959] Olive Niyomubyeyi, Tome Eduardo Sicuaio, José Ignacio Díaz González, Petter Pilesjö, and Ali Mansourian. A comparative study of four metaheuristic algorithms, AMOSA, MOABC, MSPSO, and NSGA-II for evacuation planning. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/16.

Pintelas:2020:GBE

[960] Emmanuel Pintelas, Ioannis E. Livieris, and Panagiotis Pintelas. A greybox ensemble model exploiting black-box accuracy and white-box intrinsic interpretability. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/17.

Wang:2020:TPS

[961] Zhen Wang, Fuzhen Sun, Longbo Zhang, Lei Wang, and Pingping Liu. Top position sensitive ordinal relation preserving bitwise weight for image retrieval. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/1/18.

Lutgehetmann:2020:CPH

[962] Daniel Lütgehetmann, Dejan Govc, Jason P. Smith, and Ran Levi. Computing persistent homology of directed flag complexes. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/1/19.

Zhang:2020:CCT

[963] Dehai Zhang, Linan Liu, Cheng Xie, Bing Yang, and Qing Liu. Citywide cellular traffic prediction based on a hybrid spatiotemporal network. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/1/20.

Yan:2020:MCM

[964] Qiao Yan, Xiaoqian Liu, Xiaoping Deng, Wei Peng, and Guiqing Zhang. Markov Chain Monte Carlo based energy use behaviors prediction of office occupants. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/21.

Li:2020:NDA

[965] Junfang Li, Mingqian Liu, Ningjie Tang, and Bodong Shang. Non dataaided SNR estimation for UAV OFDM systems. Algorithms (Basel), 13 (1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/1/22.

Smeresky:2020:OLS

[966] Brendon Smeresky, Alex Rizzo, and Timothy Sands. Optimal learning and self-awareness versus PDI. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/1/23.

Keshtkar:2020:UUB

[967] Najmeh Keshtkar and Klaus Röbenack. Unstructured uncertainty based modeling and robust stability analysis of textile-reinforced composites with embedded shape memory alloys. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/1/24.

Sharma:2020:LCE

[968] Janak Raj Sharma, Sunil Kumar, and Ioannis K. Argyros. Local convergence of an efficient multipoint iterative method in Banach space. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/1/25.

Karlos:2020:SVE

[969] Stamatis Karlos, Georgios Kostopoulos, and Sotiris Kotsiantis. A softvoting ensemble based co-training scheme using static selection for binary classification problems. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/1/26.

DAndreagiovanni:2020:MJO

[970] Fabio D’Andreagiovanni, Hicham Lakhlef, and Antonella Nardin. A matheuristic for joint optimal power and scheduling assignment in DVB-T2 networks. Algorithms (Basel), 13(1), January 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/27.

Bille:2020:ESI

[971] Philip Bille. Editorial: Special issue on data compression algorithms and their applications. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/1/28.

AEO:2020:ARA

[972] Algorithms Editorial Office. Acknowledgement to reviewers of Algorithms in 2019. Algorithms (Basel), 13(1), January 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/1/29.

Schoeneman:2020:LMD

[973] Frank Schoeneman, Varun Chandola, Nils Napp, Olga Wodo, and Jaroslaw Zola. Learning manifolds from dynamic process data. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/2/30.

Aloisio:2020:CCB

[974] Alessandro Aloisio and Alfredo Navarra. Constrained connectivity in bounded X-width multi-interface networks. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/2/31.

DiIanni:2020:LBT

[975] Miriam Di Ianni and Giovanna Varricchio. Latency-bounded target set selection in signed networks. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/2/32.

Daga:2020:GAT

[976] Alessandro Paolo Daga and Luigi Garibaldi. GA-adaptive template matching for offline shape motion tracking based on edge detection: IAS estimation from the SURVISHNO 2019 Challenge Video for machine diagnostics purposes. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/2/33.

Marmol:2020:UBR

[977] Mage Marmol, Leandro do C. Martins, Sara Hatami, Angel A. Juan, and Vicenc Fernandez. Using biased-randomized algorithms for the multiperiod product display problem with dynamic attractiveness. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/2/34.

Hedar:2020:ASS

[978] Abdel-Rahman Hedar, Shada N. Abdulaziz, Adel A. Sewisy, and Gamal A. El-Sayed. Adaptive scatter search to solve the minimum connected dominating set problem for efficient management of wireless networks. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/2/35.

Mensah:2020:AAS

[979] Dennis Nii Ayeh Mensah, Hui Gao, and Liang Wei Yang. Approximation algorithm for shortest path in large social networks. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/2/36.

Darweesh:2020:NNT

[980] Amer Darweesh, Marwan Alquran, and Khawla Aghzawi. New numerical treatment for a family of two-dimensional fractional Fredholm integrodifferential equations. Algorithms (Basel), 13(2), February 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/13/2/37.

Alagoz:2020:MLM

[981] Baris Baykant Alagoz, Aleksei Tepljakov, Eduard Petlenkov, and Celaleddin Yeroglu. Multi-loop model reference proportional integral derivative controls: Design and performance evaluations. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/2/38.

Pavlyuk:2020:TLV

[982] Dmitry Pavlyuk. Transfer learning: Video prediction and spatiotemporal urban traffic forecasting. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/2/39.

Gomez-Avila:2020:NPC

[983] Javier Gomez-Avila, Carlos Villaseñor, Jesus Hernandez-Barragan, Nancy Arana-Daniel, Alma Y. Alanis, and Carlos Lopez-Franco. Neural PD controller for an unmanned aerial vehicle trained with extended Kalman filter. Algorithms (Basel), 13(2), February 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/2/40.

Karaduta:2020:DUS

[984] Oleg K. Karaduta, Aleksei F. Deon, and Yulian A. Menyaev. Designing the uniform stochastic photomatrix therapeutic systems. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/2/41.

Schwarz:2020:LUB

[985] Michael Schwarz. Lower and upper bounds for the discrete bi-directional preemptive conversion problem with a constant price interval. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/2/42.

Sai:2020:OMC

[986] Qiuyue Sai, Jun Bi, and Jinxian Chai. Optimal model for carsharing station location based on multi-factor constraints. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/2/43.

Li:2020:MMB

[987] Hongchan Li, Haodong Zhu, and Tianhua Jiang. Modified migrating birds optimization for energy-aware flexible job shop scheduling problem. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
2/44.

Yang:2020:ATD

[988] Fan Yang and ShouLian Tang. Adaptive tolerance dehazing algorithm based on dark channel prior. Algorithms (Basel), 13(2), February 2020.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/2/45.

Min:2020:FFD

[989] Yufang Min and Yaonan Zhang. FADIT: Fast document image thresholding. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/2/46.

Pilz:2020:ABS

[990] Sarah Pilz, Florian Porrmann, Martin Kaiser, Jens Hagemeyer, James M. Hogan, and Ulrich Rückert. Accelerating binary string comparisons with a scalable, streaming-based system architecture based on FPGAs. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/2/47.

Alekseeva:2020:PSC

[991] Natalia Alekseeva, Ivan Tanev, and Katsunori Shimohara. PD steering controller utilizing the predicted position on track for autonomous vehicles driven on slippery roads. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/2/48.

Xue:2020:OCS

[992] Ruobing Xue, Xiangshen Ye, and Weiping Wu. Optimization of constrained stochastic linear-quadratic control on an infinite horizon: a direct-comparison based approach. Algorithms (Basel), 13(2), February 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/2/49.

Rodriguez-Mata:2020:FSM

[993] Abraham Efraim Rodríguez-Mata, Ricardo Luna, Jose Ricardo Pérez-Correa, Alejandro Gonzalez-Huitrón, Rafael Castro-Linares, and Manuel A. Duarte-Mermoud. Fractional sliding mode nonlinear procedure for robust control of an eutrophying microalgae photobioreactor. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/3/50.

Loukrezis:2020:AUQ

[994] Dimitrios Loukrezis and Herbert De Gersem. Approximation and uncertainty quantification of systems with arbitrary parameter distributions using weighted Leja interpolation. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/3/51.

Abdelaal:2020:PPF

[995] Mohamed Abdelaal and Steffen Schön. Predictive path following and collision avoidance of autonomous connected vehicles. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/3/52.

Strand:2020:UPT

[996] Andreas Strand, Ivar Eskerud Smith, Tor Erling Unander, Ingelin Steinsland, and Leif Rune Hellevik. Uncertainty propagation through a point model for steady-state two-phase pipe flow. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/53.

Stratigi:2020:MGR

[997] Maria Stratigi, Haridimos Kondylakis, and Kostas Stefanidis. Multidimensional group recommendations in the health domain. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/3/54.

Oyamaguchi:2020:MMB

[998] Natsumi Oyamaguchi, Hiroyuki Tajima, and Isamu Okada. Model of multi-branch trees for efficient resource allocation. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/3/55.

Xiao:2020:MFP

[999] Yancai Xiao and Zhe Hua. Misalignment fault prediction of wind turbines based on combined forecasting model. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/3/56.

Li:2020:TSC

[1000] Taoying Li, Xu Wu, and Junhe Zhang. Time series clustering model based on DTW for classifying car parks. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/57.

Rauh:2020:KFB

[1001] Andreas Rauh, Wiebke Frenkel, and Julia Kersten. Kalman filter-based online identification of the electric power characteristic of solid oxide fuel cells aiming at maximum power point tracking. Algorithms (Basel), 13 (3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/58.

Alexandridis:2020:GAD

[1002] Georgios Alexandridis, Yorghos Voutos, Phivos Mylonas, and George Caridakis. A geolocation analytics-driven ontology for short-term leases: Inferring current sharing economy trends. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/59.

Liu:2020:MUM

[1003] Wen Liu, Yankui Sun, and Qingge Ji. MDAN-UNet: Multi-scale and dual attention enhanced nested U-net architecture for segmentation of optical coherence tomography images. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/3/60.

Demertzis:2020:GMA

[1004] Konstantinos Demertzis and Lazaros Iliadis. GeoAI: a model-agnostic meta-ensemble zero-shot learning method for hyperspectral image analysis and classification. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/3/61.

Tran:2020:RLI

[1005] Manh-Kien Tran and Michael Fowler. A review of lithium-ion battery fault diagnostic algorithms: Current progress and future challenges. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/3/62.

Ropiak:2020:HDL

[1006] Krzysztof Ropiak and Piotr Artiemjew. On a hybridization of deep learning and rough set based granular computing. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/63.

Lewis:2020:DBN

[1007] Austin D. Lewis and Katrina M. Groth. A dynamic Bayesian network structure for joint diagnostics and prognostics of complex engineering systems. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/3/64.

Papakostas:2020:NIO

[1008] George A. Papakostas, John W. Nolan, and Athanasios C. Mitropoulos. Nature-inspired optimization algorithms for the 3D reconstruction of porous media. Algorithms (Basel), 13(3), March 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/3/65.

Paradowski:2020:OUN

[1009] Thomas Paradowski, Sabine Lerch, Michelle Damaszek, Robert Dehnert, and Bernd Tibken. Observability of uncertain nonlinear systems using interval analysis. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/3/66.

Bacanin:2020:OCN

[1010] Nebojsa Bacanin, Timea Bezdan, Eva Tuba, Ivana Strumberger, and Milan Tuba. Optimizing convolutional neural network hyperparameters by enhanced swarm intelligence metaheuristics. Algorithms (Basel), 13 (3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/67.

Coro:2020:AEM

[1011] Federico Corò, Gianlorenzo D’Angelo, and Cristina M. Pinotti. Adding edges for maximizing weighted reachability. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/3/68.

Xu:2020:OSM

[1012] Jin Xu, Haixia Wang, Can Cui, Baigang Zhao, and Bo Li. Oil spill monitoring of shipborne radar image features using SVM and local adaptive threshold. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/3/69.

Zvarevashe:2020:ELH

[1013] Kudakwashe Zvarevashe and Oludayo Olugbara. Ensemble learning of hybrid acoustic features for speech emotion recognition. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/3/70.

Alexopoulos:2020:TSC

[1014] Athanasios Alexopoulos, Georgios Drakopoulos, Andreas Kanavos, Phivos Mylonas, and Gerasimos Vonitsanos. Two-step classification with SVD preprocessing of distributed massive datasets in Apache spark. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/3/71.

Nakas:2020:EER

[1015] Christos Nakas, Dionisis Kandris, and Georgios Visvardis. Energy efficient routing in wireless sensor networks: a comprehensive survey. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/3/72.

Samee:2020:CDL

[1016] Nagwan M. Abdel Samee. Classical and deep learning paradigms for detection and validation of key genes of risky outcomes of HCV. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/3/73.

Abdelmaguid:2020:BOD

[1017] Tamer F. Abdelmaguid. Bi-objective dynamic multiprocessor open shop scheduling: an exact algorithm. Algorithms (Basel), 13(3), March 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/3/74.

Frenkel:2020:EBC

[1018] Wiebke Frenkel, Andreas Rauh, Julia Kersten, and Harald Aschemann. Experiments-based comparison of different power controllers for a solid oxide fuel cell against model imperfections and delay phenomena. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/4/76.

Coclite:2020:CSK

[1019] Giuseppe Maria Coclite and Lorenzo di Ruvo. On classical solutions for a Kuramoto–Sinelshchikov–Velarde-type equation. Algorithms (Basel), 13 (4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/4/77.

Aggarwal:2020:BNN

[1020] Ankush Aggarwal and Sanjay Pant. Beyond Newton: a new root-finding fixed-point iteration for nonlinear equations. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/4/78.

Artiemjew:2020:AGR

[1021] Piotr Artiemjew. About granular rough computing-overview of decision system approximation techniques and future perspectives. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/4/79.

Ji:2020:HMB

[1022] Qingge Ji, Haoqiang Yu, and Xiao Wu. Hierarchical-matching-based online and real-time multi-object tracking with deep appearance features. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/4/80.

Pratico:2020:DMB

[1023] Filippo Giammaria Praticò, Rosario Fedele, Vitalii Naumov, and Tomas Sauer. Detection and monitoring of bottom-up cracks in road pavement using a machine-learning approach. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/4/81.

Liao:2020:APP

[1024] Huanyu Liao, Pavan Kumar Vaitheeswaran, Tao Song, and Ganesh Subbarayan. Algebraic point projection for immersed boundary analysis on low degree NURBS curves and surfaces. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/4/82.

Haralabopoulos:2020:EDL

[1025] Giannis Haralabopoulos, Ioannis Anagnostopoulos, and Derek McAuley. Ensemble deep learning for multilabel binary classification of usergenerated content. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/83.

Zhang:2020:RTC

[1026] Sen Zhang, Shaobo Li, Xiang Li, and Yong Yao. Representation of traffic congestion data for urban road traffic networks based on pooling operations. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/4/
84.

Demidova:2020:RSH

[1027] Liliya A. Demidova and Artyom V. Gorchakov. Research and study of the hybrid algorithms based on the collective behavior of fish schools and classical optimization methods. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/4/85.

Stanovov:2020:CBV

[1028] Vladimir Stanovov, Shakhnaz Akhmedova, and Yukihiro Kamiya. Confidence-based voting for the design of interpretable ensembles with fuzzy systems. Algorithms (Basel), 13(4), April 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/86.

Konstantinidis:2020:NMP

[1029] George Konstantinidis, Adriane Chapman, Mark J. Weal, Ahmed Alzubaidi, Lisa M. Ballard, and Anneke M. Lucassen. The need for machine-processable agreements in health data management. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/4/87.

Miertoiu:2020:FPA

[1030] Florin Ilarion Miertoiu and Bogdan Dumitrescu. Feasibility pump algorithm for sparse representation under Gaussian noise. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/4/88.

Akhmedova:2020:SHB

[1031] Shakhnaz Akhmedova, Vladimir Stanovov, Danil Erokhin, and Olga Semenkina. Success history-based position adaptation in fuzzy-controlled ensemble of biology-inspired algorithms. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/4/89.

Kantarakias:2020:AGP

[1032] Kyriakos Dimitrios Kantarakias and George Papadakis. Application of generalized polynomial chaos for quantification of uncertainties of time averages and their sensitivities in chaotic systems. Algorithms (Basel), 13 (4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/4/90.

Tang:2020:GAL

[1033] Chunming Tang, Yanni Li, Xiaoxia Dong, and Bo He. A generalized alternating linearization bundle method for structured convex optimization with inexact first-order oracles. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/4/91.

Barger:2020:DCM

[1034] Artem Barger and Dan Feldman. Deterministic coresets for k-means of big sparse data. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/92.

Li:2020:PPL

[1035] Yuanjin Li, Tao Chen, and Defu Liu. Path planning for laser cladding robot on artificial joint surface based on topology reconstruction. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/4/93.

Petrauskas:2020:SMQ

[1036] Edmundas Petrauskas, Petras Rupsys, Martynas Narmontas, Marius Aleinikovas, Lina Beniusiene, and Benas Silinskas. Stochastic models to qualify stem tapers. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/94.

Stubinger:2020:HIV

[1037] Johannes Stübinger and Katharina Adler. How to identify varying leadlag effects in time series data: Implementation, validation, and application of the generalized causality algorithm. Algorithms (Basel), 13 (4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/4/95.

Soto-Mendoza:2020:HGO

[1038] Valeria Soto-Mendoza, Irma García-Calvillo, Efraín Ruiz y Ruiz, and Jaime Pérez-Terrazas. A hybrid grasshopper optimization algorithm applied to the open vehicle routing problem. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/4/96.

Domanski:2020:PAP

[1039] Paweł D. Domański. Performance assessment of predictive control — a survey. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/4/

Averos:2020:NSN

[1040] Joan Cecilia Averós, Jaume Puy Llorens, and Ramiro Uribe-Kaffure. Numerical simulation of non-linear models of reaction-diffusion for a DGT sensor. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/4/
98.

Mansouri:2020:NLD

[1041] Deloula Mansouri, Xiaohui Yuan, and Abdeldjalil Saidani. A new lossless DNA compression algorithm based on a single-block encoding scheme. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/4/99.

Bergamaschi:2020:SLR

[1042] Luca Bergamaschi. A survey of low-rank updates of preconditioners for sequences of symmetric linear systems. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/4/100.

Lawnik:2020:NWS

[1043] Marcin Lawnik, Artur Pełka, and Adrian Kapczyński. A new way to store simple text files. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/101.

Lopez-Martinez:2020:CSB

[1044] Fernando López-Martínez, Edward Rolando Núñez-Valdez, Vicente García-Díaz, and Zoran Bursac. A case study for a big data and machine learning platform to improve medical decision support in population health management. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/102.

Furuya:2020:PGC

[1045] Isamu Furuya, Takuya Takagi, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Takuya Kida. Practical grammar compression based on maximal repeats. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/103.

Guimapi:2020:DSS

[1046] Ritter A. Guimapi, Samira A. Mohamed, Lisa Biber-Freudenberger, Waweru Mwangi, Sunday Ekesi, Christian Borgemeister, and Henri E. Z.

Tonnang. Decision support system for fitting and mapping nonlinear functions with application to insect pest management in the biological control context. Algorithms (Basel), 13(4), April 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/4/104.

Ju:2020:VSM

[1047] Haiyang Ju, Xinhua Wang, and Yizhen Zhao. Variational specific mode extraction: a novel method for defect signal detection of ferromagnetic pipeline. Algorithms (Basel), 13(4), April 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/4/105.

Roanes-Lozano:2020:DTS

[1048] Eugenio Roanes-Lozano, Eduardo A. Casella, Fernando Sánchez, and Antonio Hernando. Diagnosis in tennis serving technique. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/5/106.

Azeroual:2020:HIM

[1049] Otmane Azeroual and Włodzimierz Lewoniewski. How to inspect and measure data quality about scientific publications: Use case of Wikipedia and CRIS databases. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/5/107.

Vakhnin:2020:IIF

[1050] Alexey Vakhnin and Evgenii Sopov. Investigation of the iCC framework performance for solving constrained LSGO problems. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/5/108.

Gorzalczany:2020:ESS

[1051] Marian B. Gorzałczany and Filip Rudziński. Evolution of SOMs’ structure and learning algorithm: From visualization of high-dimensional data to clustering of complex data. Algorithms (Basel), 13(5), May 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/13/5/109.

Blondeel:2020:RMQ

[1052] Philippe Blondeel, Pieterjan Robbe, Cédric Van hoorickx, Stijn François, Geert Lombaert, and Stefan Vandewalle. p-refined multilevel quasi-Monte Carlo for Galerkin finite element methods with applications in civil engineering. Algorithms (Basel), 13(5), May 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/5/110.

Wu:2020:MLJ

[1053] Shaojun Wu and Ling Gao. Multi-level joint feature learning for person re-identification. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/5/111.

Sauvey:2020:TNH

[1054] Christophe Sauvey and Nathalie Sauer. Two NEH heuristic improvements for flowshop scheduling problem with makespan criterion. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/5/112.

Meisrimel:2020:GOT

[1055] Peter Meisrimel and Philipp Birken. Goal oriented time adaptivity using local error estimates. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/5/113.

Yang:2020:AFG

[1056] Fan Yang, Deming Yang, Zhiming He, Yuanhua Fu, and Kui Jiang. Automobile fine-grained detection algorithm based on multi-improved YOLOv3 in smart streetlights. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/5/114.

Roghabadi:2020:FBD

[1057] Mohammadjavad Arabpour Roghabadi and Osama Moselhi. A fuzzybased decision support model for risk maturity evaluation of construction organizations. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/5/115.

Mazzoccoli:2020:EUI

[1058] Alessandro Mazzoccoli and Maurizio Naldi. The expected utility insurance premium principle with fourth-order statistics: Does it make a difference? Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/5/116.

Prayogo:2020:NHM

[1059] Doddy Prayogo, Min-Yuan Cheng, Yu-Wei Wu, A. A. N. Perwira Redi, Vincent F. Yu, Satria Fadil Persada, and Reny Nadlifatin. A novel hybrid metaheuristic algorithm for optimization of construction management site layout planning. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/5/117.

Lindenberg:2020:DRL

[1060] Björn Lindenberg, Jonas Nordqvist, and Karl-Olof Lindahl. Distributional reinforcement learning with ensembles. Algorithms (Basel), 13 (5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/5/118.

Castelli:2020:FEP

[1061] Mauro Castelli, Ales Groznik, and Ales Popovic. Forecasting electricity prices: a machine learning approach. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/5/119.

Bazgir:2020:NDD

[1062] Omid Bazgir, Eric Walden, Brian Nutter, and Sunanda Mitra. A novel data-driven magnetic resonance spectroscopy signal analysis framework to quantify metabolite concentration. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/5/120.

Livieris:2020:EDL

[1063] Ioannis E. Livieris, Emmanuel Pintelas, Stavros Stavroyiannis, and Panagiotis Pintelas. Ensemble deep learning models for forecasting cryptocurrency time-series. Algorithms (Basel), 13(5), May 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/5/121.

Meier:2020:IFD

[1064] Arne Meier. Incremental FPT delay. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/5/122.

Santoro:2020:MSP

[1065] Diego Santoro, Andrea Tonon, and Fabio Vandin. Mining sequential patterns with VC-dimension and Rademacher complexity. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/5/123.

Zhu:2020:NMI

[1066] Jianshen Zhu, Chenxi Wang, Aleksandar Shurbevski, Hiroshi Nagamochi, and Tatsuya Akutsu. A novel method for inference of chemical compounds of cycle index two with desired properties based on artificial neural networks and integer programming. Algorithms (Basel), 13 (5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/5/124.

Vestias:2020:MDL

[1067] Mário P. Véstias, Rui Policarpo Duarte, José T. de Sousa, and Horácio C. Neto. Moving deep learning to the edge. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/5/125.

Chen:2020:PSP

[1068] Feiyang Chen, Ying Jiang, Xiangrui Zeng, Jing Zhang, Xin Gao, and Min Xu. PUB-SalNet: a pre-trained unsupervised self-aware backpropagation network for biomedical salient segmentation. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/5/126.

Chen:2020:EDD

[1069] Hui-Hui Chen, Bor-Jiunn Hwang, Jung-Shyr Wu, and Po-Ting Liu. The effect of different deep network architectures upon CNN-based gaze tracking. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
5/127.

Ma:2020:CPD

[1070] Lijing Ma and Georgy Sofronov. Change-point detection in autoregressive processes via the cross-entropy method. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/5/128.

Malczewski:2020:IRE

[1071] Krzysztof Malczewski. Image resolution enhancement of highly compressively sensed CT/PET signals. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/5/129.

Cravero:2020:UQA

[1072] Carlo Cravero, Davide De Domenico, and Andrea Ottonello. Uncertainty quantification approach on numerical simulation for supersonic jets performance. Algorithms (Basel), 13(5), May 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/5/130.

Al-Jadir:2020:UTF

[1073] Ibraheem Al-Jadir, Kok Wai Wong, Chun Che Fung, and Hong Xie. Unsupervised text feature selection using memetic dichotomous differential evolution. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/6/131.

Daniel:2020:STW

[1074] Lucky O. Daniel, Caston Sigauke, Colin Chibaya, and Rendani Mbuvha. Short-term wind speed forecasting using statistical and machine learning methods. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/6/132.

Kertesz:2020:MEL

[1075] Gábor Kertész. Metric embedding learning on multi-directional projections. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/6/
133.

Gozdur:2020:SQS

[1076] Roman Gozdur. Study of quasi-static magnetization with the randomfield Ising model. Algorithms (Basel), 13(6), June 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/6/134.

Elisei-Iliescu:2020:RLS

[1077] Camelia Elisei-Iliescu, Laura-Maria Dogariu, Constantin Paleologu, Jacob Benesty, Andrei-Alexandru Enescu, and Silviu Ciochina. A recursive least-squares algorithm for the identification of trilinear forms. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/6/135.

Quan:2020:ICS

[1078] Zhi Quan and Shuhua Lv. Improved convergence speed of a DCD-based algorithm for sparse solutions. Algorithms (Basel), 13(6), June 2020.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/6/136.

Zanon:2020:SLR

[1079] Mattia Zanon, Giuliano Zambonin, Gian Antonio Susto, and Seán McLoone. Sparse logistic regression: Comparison of regularization and Bayesian implementations. Algorithms (Basel), 13(6), June 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-48931/3/6/137.

Goerler:2020:LAH

[1080] Andreas Goerler, Eduardo Lalla-Ruiz, and Stefan Voß. Late acceptance hill-climbing matheuristic for the general lot sizing and scheduling problem with rich constraints. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/6/138.

Cutello:2020:OAD

[1081] Vincenzo Cutello, Georgia Fargetta, Mario Pavone, and Rocco A. Scollo. Optimization algorithms for detection of social interactions. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/6/139.

Pintelas:2020:SIE

[1082] Panagiotis Pintelas and Ioannis E. Livieris. Special issue on ensemble learning and applications. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/6/140.

Gotoh:2020:DRE

[1083] Tsuyoshi Gotoh, Yuichi Sudo, Fukuhito Ooshita, and Toshimitsu Masuzawa. Dynamic ring exploration with (H, S) view. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/6/141.

Jedrzejowicz:2020:PSI

[1084] Piotr Jedrzejowicz and Izabela Wierzbowska. Parallelized swarm intelligence approach for solving TSP and JSSP problems. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/6/142.

Marusak:2020:NEF

[1085] Piotr M. Marusak. Numerically efficient fuzzy MPC algorithm with advanced generation of prediction — application to a chemical reactor. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/6/143.

Bobe:2020:EPJ

[1086] Christin Bobe, Daan Hanssens, Thomas Hermans, and Ellen Van De Vijver. Efficient probabilistic joint inversion of direct current resistivity and small-loop electromagnetic data. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/6/144.

Gundersen:2020:BTS

[1087] Kristian Gundersen, Guttorm Alendal, Anna Oleynik, and Nello Blaser. Binary time series classification with Bayesian convolutional neural networks when monitoring for marine gas discharges. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/6/145.

Feldmann:2020:SAP

[1088] Andreas Emil Feldmann, Karthik C. S., Euiwoong Lee, and Pasin Manurangsi. A survey on approximation in parameterized complexity: Hardness and algorithms. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/6/146.

Regmi:2020:LCB

[1089] Samundra Regmi, Ioannis K. Argyros, and Santhosh George. Local comparison between two ninth convergence order algorithms for equations. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/6/147.

Dechanubeksa:2020:AMG

[1090] Chutipong Dechanubeksa and Saifon Chaturantabut. An application of a modified gappy proper orthogonal decomposition on complexity reduction of Allen–Cahn equation. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/6/148.

Khryashchev:2020:DAE

[1091] Denis Khryashchev, Jie Chu, Mikael Vejdemo-Johansson, and Ping Ji. A distributed approach to the evasion problem. Algorithms (Basel), 13 (6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/6/149.

Carlsson:2020:FFC

[1092] Leo S. Carlsson, Mikael Vejdemo-Johansson, Gunnar Carlsson, and Pär G. Jönsson. Fibers of failure: Classifying errors in predictive processes. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/6/
150.

Carpentieri:2020:CNG

[1093] Bruno Carpentieri. Compression of next-generation sequencing data and of DNA digital files. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/6/151.

Tang:2020:DET

[1094] Liangrui Tang and Zhilin Lu. DS evidence theory-based energy balanced routing algorithm for network lifetime enhancement in WSN-assisted IOT. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/6/
152.

Pamosoaji:2020:NGM

[1095] Anugrah K. Pamosoaji and Djoko Budiyanto Setyohadi. Novel graph model for solving collision-free multiple-vehicle traveling salesman problem using ant colony optimization. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/6/153.

Giakoumakis:2020:AFN

[1096] Stylianos Giakoumakis and Basil Papadopoulos. An algorithm for fuzzy negations based-intuitionistic fuzzy copula aggregation operators in multiple attribute decision making. Algorithms (Basel), 13(6), June 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/6/154.

Dezan:2020:EBN

[1097] Catherine Dezan, Sara Zermani, and Chabha Hireche. Embedded Bayesian network contribution for a safe mission planning of autonomous vehicles. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/7/

Zoccarato:2020:GPC

[1098] Claudia Zoccarato, Laura Gazzola, Massimiliano Ferronato, and Pietro Teatini. Generalized polynomial chaos expansion for fast and accurate uncertainty quantification in geomechanical modelling. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/7/156.

Sturrock:2020:RRO

[1099] Peter Sturrock and Felix Scholkmann. The RONO (rank-order-normalization) procedure for power-spectrum analysis of datasets with non-normal distributions. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/7/157.

Khang:2020:FMC

[1100] Tran Dinh Khang, Nguyen Duc Vuong, Manh-Kien Tran, and Michael Fowler. Fuzzy C-means clustering algorithm with multiple fuzzification coefficients. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/7/158.

Yamagiwa:2020:SBL

[1101] Shinichi Yamagiwa, Eisaku Hayakawa, and Koichi Marumo. Streambased lossless data compression applying adaptive entropy coding for hardware-based implementation. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/7/159.

Makris:2020:TSA

[1102] Christos Makris, Georgios Pispirigos, and Michael Angelos Simos. Text semantic annotation: a distributed methodology based on community coherence. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/7/160.

Elkhawaga:2020:CPS

[1103] Ghada Elkhawaga, Mervat Abuelkheir, Sherif I. Barakat, Alaa M. Riad, and Manfred Reichert. CONDA-PM — a systematic review and framework for concept drift analysis in process mining. Algorithms (Basel), 13 (7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/7/161.

Godel:2020:SAM

[1104] Marion Gödel, Rainer Fischer, and Gerta Köster. Sensitivity analysis for microscopic crowd simulation. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/7/162.

Fateminia:2020:ITF

[1105] Seyed Hamed Fateminia, Vuppuluri Sumati, and Aminah Robinson Fayek. An interval type-2 fuzzy risk analysis model (IT2FRAM) for determining construction project contingency reserve. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/7/163.

Rafajlowicz:2020:NEC

[1106] Wojciech Rafajłowicz. Nonparametric estimation of continuously parametrized families of probability density functions — computational aspects. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/7/
164.

Vemuru:2020:IED

[1107] Krishnamurthy V. Vemuru. Image edge detector with Gabor type filters using a spiking neural network of biologically inspired neurons. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/7/165.

Griewank:2020:PDD

[1108] Andreas Griewank and Andrea Walther. Polyhedral DC decomposition and DCA optimization of piecewise linear functions. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/7/166.

Malowany:2020:BIV

[1109] Dan Malowany and Hugo Guterman. Biologically inspired visual system architecture for object recognition in autonomous systems. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/7/167.

Aversano:2020:RBS

[1110] Lerina Aversano, Martina Iammarino, Mimmo Carapella, Andrea Del Vecchio, and Laura Nardi. On the relationship between self-admitted technical debt removals and technical debt measures. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/7/168.

Wu:2020:TTS

[1111] Xiao Wu and Qingge Ji. TBRNet: Two-stream BiLSTM residual network for video action recognition. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/7/169.

Coclite:2020:WPH

[1112] Giuseppe Maria Coclite and Lorenzo di Ruvo. On the well-posedness of a high order convective Cahn–Hilliard type equations. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/7/170.

Nishikawa:2020:EMG

[1113] Kohei Nishikawa and Takahisa Toda. Exact method for generating strategy-solvable Sudoku clues. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/7/171.

Knudson:2020:ATG

[1114] Kevin P. Knudson. Approximate triangulations of Grassmann manifolds. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/7/172.

Li:2020:MHS

[1115] Cong Li, Yaonan Zhang, and Xupeng Ren. Modeling hourly soil temperature using deep BiLSTM neural network. Algorithms (Basel), 13 (7), July 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/7/173.

Margaris:2020:ADE

[1116] Dionisis Margaris, Dimitris Spiliotopoulos, Gregory Karagiorgos, and Costas Vassilakis. An algorithm for density enrichment of sparse collaborative filtering datasets using robust predictions as derived ratings. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/7/174.

Dietrich:2020:EFH

[1117] Eric Dietrich and Chris Fields. Equivalence of the frame and halting problems. Algorithms (Basel), 13(7), July 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/7/175.

Beheshti:2020:TCR

[1118] Amin Beheshti, Shahpar Yakhchi, Salman Mousaeirad, Seyed Mohssen Ghafari, Srinivasa Reddy Goluguri, and Mohammad Amin Edrisi. Towards cognitive recommender systems. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/13/8/176.

Epstein:2020:SFA

[1119] Dror Epstein and Dan Feldman. Sphere fitting with applications to machine tracking. Algorithms (Basel), 13(8), August 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/8/177.

Plamowski:2020:MOR

[1120] Sebastian Plamowski and Richard W. Kephart. The model order reduction method as an effective way to implement GPC controller for multidimensional objects. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/8/178.

Rezapour:2020:TCB

[1121] Mahdi Rezapour and Khaled Ksaibati. Two-component Bayesian hierarchical models for cost-benefit analysis of traffic barrier crash count. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/8/179.

Qi:2020:PAE

[1122] Bing Qi, Shuyu Qian, and Aaron Costin. A predictive analysis on emerging technology utilization in industrialized construction in the United States and China. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/8/180.

Donado:2020:OCR

[1123] Fernando Solano Donado. On the optimal calculation of the Rice coding parameter. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/8/181.

Dritsas:2020:TCN

[1124] Elias Dritsas, Andreas Kanavos, Maria Trigka, Gerasimos Vonitsanos, Spyros Sioutas, and Athanasios Tsakalidis. Trajectory clustering and k-NN for robust privacy preserving k-NN query processing in GeoSpark. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/8/182.

Pham:2020:IMP

[1125] Canh V. Pham, Dung K. T. Ha, Quang C. Vu, Anh N. Su, and Huan X. Hoang. Influence maximization with priority in online social networks. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/8/183.

Agarwal:2020:IAS

[1126] Ravi Agarwal, Snezhana Hristova, Donal O’Regan, and Kremena Stefanova. Iterative algorithm for solving scalar fractional differential equations with Riemann–Liouville derivative and supremum. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/8/184.

Dupin:2020:MLG

[1127] Nicolas Dupin and El-Ghazali Talbi. Machine learning-guided dual heuristics and new lower bounds for the refueling and maintenance planning problem of nuclear power plants. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/8/185.

Islam:2020:RRA

[1128] Md. Saiful Islam, Emam Hossain, Abdur Rahman, Mohammad Shahadat Hossain, and Karl Andersson. A review on recent advancements in FOREX currency prediction. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/8/186.

Sampaio:2020:CRC

[1129] Altino M. Sampaio and Jorge G. Barbosa. Constructing reliable computing environments on top of Amazon EC2 spot instances. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/8/187.

Werner:2020:ARS

[1130] Piotr A. Werner. Application of the Reed–Solomon algorithm as a remote sensing data fusion tool for land use studies. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/8/188.

Zhang:2020:NPO

[1131] Yijie Zhang and Mandan Liu. Node placement optimization of wireless sensor networks using multi-objective adaptive degressive ary number encoded genetic algorithm. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/8/189.

Smietanski:2020:NGN

[1132] Marek J. Śmietański. On a nonsmooth Gauss–Newton algorithms for solving nonlinear complementarity problems. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/8/190.

DEmidio:2020:FAM

[1133] Mattia D’Emidio. Faster algorithms for mining shortest-path distances from massive time-evolving graphs. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/8/191.

Khosla:2020:LTB

[1134] Karaj Khosla, Indra Prakash Jha, Ajit Kumar, and Vibhor Kumar. Local-topology-based scaling for distance preserving dimension reduction method to improve classification of biomedical data-sets. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/8/192.

Wu:2020:CCE

[1135] Shaojun Wu and Ling Gao. Cross-camera erased feature learning for unsupervised person re-identification. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/8/193.

Angelini:2020:GPR

[1136] Patrizio Angelini, Peter Eades, Seok-Hee Hong, Karsten Klein, Stephen Kobourov, Giuseppe Liotta, Alfredo Navarra, and Alessandra Tappini. Graph planarity by replacing cliques with paths. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/8/194.

Wheeler:2020:DLE

[1137] Bradley J. Wheeler and Hassan A. Karimi. Deep learning-enabled semantic inference of individual building damage magnitude from satellite images. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/8/195.

Bonnet:2020:ARI

[1138] Luc Bonnet, Jean-Luc Akian, Éric Savin, and T. J. Sullivan. Adaptive reconstruction of imperfectly observed monotone functions, with applications to uncertainty quantification. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/8/196.

Abu-Khzam:2020:BSF

[1139] Faisal N. Abu-Khzam and Karam Al Kontar. A brief survey of fixedparameter parallelism. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/8/197.

Louk:2020:PDS

[1140] Roland Lõuk, Andri Riid, René Pihlak, and Aleksei Tepljakov. Pavement defect segmentation in orthoframes with a pipeline of three convolutional neural networks. Algorithms (Basel), 13(8), August 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/8/198.

Zanetti:2020:SBP

[1141] Filippo Zanetti and Luca Bergamaschi. Scalable block preconditioners for linearized Navier–Stokes equations at high Reynolds number. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/8/199.

Cavanna:2020:AMA

[1142] Nicholas J. Cavanna and Donald R. Sheehy. Adaptive metrics for adaptive samples. Algorithms (Basel), 13(8), August 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/8/200.

Alimohammadi:2020:NMR

[1143] Hossein Alimohammadi, Baris Baykant Alagoz, Aleksei Tepljakov, Kristina Vassiljeva, and Eduard Petlenkov. A NARX model reference adaptive control scheme: Improved disturbance rejection fractional-order PID control of an experimental magnetic levitation system. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/8/201.

Karim:2020:MAT

[1144] Abdul Karim, Azhari Azhari, Samir Brahim Belhaouri, Ali Adil Qureshi, and Maqsood Ahmad. Methodology for analyzing the traditional algorithms performance of user reviews using machine learning techniques. Algorithms (Basel), 13(8), August 2020. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/13/8/202.

Crescenzi:2020:FTN

[1145] Pierluigi Crescenzi, Clémence Magnien, and Andrea Marino. Finding top-k nodes for temporal closeness in large temporal graphs. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/9/211.

Vayer:2020:FGW

[1146] Titouan Vayer, Laetitia Chapel, Remi Flamary, Romain Tavenard, and Nicolas Courty. Fused Gromov–Wasserstein distance for structured objects. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/9/212.

Chen:2020:LPF

[1147] Chuanglu Chen, Zhiqiang Li, Yitao Zhang, Shaolong Zhang, Jiena Hou, and Haiying Zhang. Low-power FPGA implementation of convolution neural network accelerator for pulse waveform classification. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/9/213.

Coutino:2020:FSA

[1148] Mario Coutino, Sundeep Prabhakar Chepuri, Takanori Maehara, and Geert Leus. Fast spectral approximation of structured graphs with applications to graph filtering. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/9/214.

Saxena:2020:TPA

[1149] Amit Saxena, Shreya Pare, Mahendra Singh Meena, Deepak Gupta, Akshansh Gupta, Imran Razzak, Chin-Teng Lin, and Mukesh Prasad. A two-phase approach for semi-supervised feature selection. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/9/215.

Ceccarello:2020:DGD

[1150] Matteo Ceccarello, Andrea Pietracaprina, Geppino Pucci, and Eli Upfal. Distributed graph diameter approximation. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/9/216.

Kondylakis:2020:RDC

[1151] Haridimos Kondylakis, Dimitrios Tsirigotakis, Giorgos Fragkiadakis, Emmanouela Panteri, Alexandros Papadakis, Alexandros Fragkakis, Eleytherios Tzagkarakis, Ioannis Rallis, Zacharias Saridakis, Apostolos Trampas, Giorgos Pirounakis, and Nikolaos Papadakis. R2D2: a Dbpedia chatbot using triple-pattern like queries. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/9/217.

Redi:2020:SAA

[1152] A. A. N. Perwira Redi, Parida Jewpanya, Adji Candra Kurniawan, Satria Fadil Persada, Reny Nadlifatin, and Oki Anita Candra Dewi. A simulated annealing algorithm for solving two-echelon vehicle routing problem with locker facilities. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/9/218.

Razgon:2020:RRB

[1153] Margarita Razgon and Alireza Mousavi. Relaxed rule-based learning for automated predictive maintenance: Proof of concept. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/9/219. See

correction [1334].

Abouhamad:2020:FPP

[1154] Mona Abouhamad and Tarek Zayed. Fuzzy preference programming framework for functional assessment of subway networks. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/9/220.

Z-Flores:2020:EFE

[1155] Emigdio Z-Flores, Leonardo Trujillo, Pierrick Legrand, and Frédérique Faïta-Aïnseba. EEG feature extraction using genetic programming for the classification of mental states. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/9/221.

Weber:2020:DTI

[1156] Eric S. Weber, Steven N. Harding, and Lee Przybylski. Detecting traffic incidents using persistence diagrams. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/13/9/222.

Katsaragakis:2020:SUT

[1157] Iosif V. Katsaragakis, Ioannis X. Tassopoulos, and Grigorios N. Beligiannis. Solving the urban transit routing problem using a cat swarm optimization-based algorithm. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/9/223.

Abedin:2020:SSU

[1158] Paniz Abedin, M. Oguzhan Külekci, and Shama V. Thankachan. A survey on shortest unique substring queries. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/9/224.

Brejova:2020:LTA

[1159] Brona Brejová and Rastislav Královic. A linear-time algorithm for the isometric reconciliation of unrooted trees. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/9/225.

Antonelli:2020:SAR

[1160] Laura Antonelli, Valentina De Simone, and Daniela di Serafino. Spatially adaptive regularization in image segmentation. Algorithms (Basel), 13 (9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/9/226.

Du:2020:IHA

[1161] Ling Du, Zehong He, Yijing Wang, Xiaochao Wang, and Anthony T. S. Ho. An image hashing algorithm for authentication with multi-attack reference generation and adaptive thresholding. Algorithms (Basel), 13 (9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/9/227.

Shafipour:2020:OTI

[1162] Rasoul Shafipour and Gonzalo Mateos. Online topology inference from streaming stationary graph signals with partial connectivity information. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/9/
228.

Teng:2020:JDM

[1163] Zhongming Teng and Xiaowei Zhang. A Jacobi–Davidson method for large scale canonical correlation analysis. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/9/229.

Almarashi:2020:SAE

[1164] Majid Almarashi, Wael Deabes, Hesham H. Amin, and Abdel-Rahman Hedar. Simulated annealing with exploratory sensing for global optimization. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/9/230.

Liu:2020:CSF

[1165] Chengzhi Liu, Xuli Han, and Juncheng Li. A class of spline functions for solving 2-order linear differential equations with boundary conditions. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/9/
231.

Pendharkar:2020:CED

[1166] Parag C. Pendharkar. A comparison of ensemble and dimensionality reduction DEA models based on entropy criterion. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/9/232.

Cumbo:2020:BIH

[1167] Fabio Cumbo, Eleonora Cappelli, and Emanuel Weitschek. A braininspired hyperdimensional computing approach for classifying massive DNA methylation data of cancer. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/9/233.

Bannai:2020:MTS

[1168] Hideo Bannai, Travis Gagie, Gary Hoppenworth, Simon J. Puglisi, and Luís M. S. Russo. More time-space tradeoffs for finding a shortest unique substring. Algorithms (Basel), 13(9), September 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/9/234.

Colonetti:2020:MIA

[1169] Bruno Colonetti, Erlon Cristian Finardi, and Welington de Oliveira. A mixed-integer and asynchronous level decomposition with application to the stochastic hydrothermal unit-commitment problem. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/9/235.

Misra:2020:SIN

[1170] Neeldhara Misra, Frances Rosamond, and Meirav Zehavi. Special issue “New Frontiers in Parameterized Complexity and Algorithms”: Foreward by the Guest Editors. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/9/236.

Martins:2020:SAS

[1171] Leandro do C. Martins, Christopher Bayliss, Pedro J. Copado-Méndez, Javier Panadero, and Angel A. Juan. A simheuristic algorithm for solving the stochastic omnichannel vehicle routing problem with pick-up and delivery. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/9/237.

Li:2020:SMO

[1172] Zhixing Li, Paolo Vincenzo Genovese, and Yafei Zhao. Study on multiobjective optimization-based climate responsive design of residential building. Algorithms (Basel), 13(9), September 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/9/238.

Li:2020:FAA

[1173] Menglin Li, Xueqiang Gu, Chengyi Zeng, and Yuan Feng. Feasibility analysis and application of reinforcement learning algorithm based on dynamic parameter adjustment. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/9/239.

Mandarawi:2020:PBC

[1174] Waseem Mandarawi, Jürgen Rottmeier, Milad Rezaeighale, and Hermann de Meer. Policy-based composition and embedding of extended virtual networks and SFCs for IIoT. Algorithms (Basel), 13(9), September 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/13/9/240.

Ardabili:2020:COP

[1175] Sina F. Ardabili, Amir Mosavi, Pedram Ghamisi, Filip Ferdinand, Annamaria R. Varkonyi-Koczy, Uwe Reuter, Timon Rabczuk, and Peter M. Atkinson. COVID-19 outbreak prediction with machine learning. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/249.

Bilo:2020:MCG

[1176] Vittorio Bilò, Michele Flammini, Vasco Gallotti, and Cosimo Vinci. On multidimensional congestion games. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/10/261.

Bryniarska:2020:ADG

[1177] Anna Bryniarska. The auto-diagnosis of granulation of information retrieval on the Web. Algorithms (Basel), 13(10):??, October 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/13/10/264.

Burr:2020:AAB

[1178] Tom Burr, Andrea Favalli, Marcie Lombardi, and Jacob Stinnett. Application of the approximate Bayesian computation algorithm to gammaray spectroscopy. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/10/265.

Cao:2020:NGK

[1179] Xiang Cao and Cheng Li. A novel global key-value storage system based on kinetic drives. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/10/247.

Chen:2020:LLS

[1180] Xin Chen, Hong Zhao, and Ping Zhou. Lung lobe segmentation based on lung fissure surface classification using a point cloud region growing approach. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/10/263.

Dong:2020:BQE

[1181] Wu Dong, Hongxia Bie, Likun Lu, and Yeli Li. Blind quality evaluation for screen content images based on regionalized structural features. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/257.

Fedele:2020:ISS

[1182] Rosario Fedele and Massimo Merenda. An IoT system for social distancing and emergency management in smart cities using multi-sensor data. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/10/254.

Feng:2020:EFE

[1183] Shang Feng, Haifeng Li, Lin Ma, and Zhongliang Xu. An EEG feature extraction method based on sparse dictionary self-organizing map for event-related potential recognition. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/10/259.

Goyal:2020:ORP

[1184] Shashank Goyal and Diwakar Gupta. The online reservation problem. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/241.

Hajij:2020:EDR

[1185] Mustafa Hajij and Paul Rosen. An efficient data retrieval parallel Reeb graph algorithm. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/10/258.

Kim:2020:DVN

[1186] Dongkyun Kim and Yong-Hwan Kim. Dynamic virtual network slicing and orchestration for selective MEC services over wide-area SDN. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/245.

Konstantakopoulos:2020:MLN

[1187] Grigorios D. Konstantakopoulos, Sotiris P. Gayialis, Evripidis P. Kechagias, Georgios A. Papadopoulos, and Ilias P. Tatsiopoulos. A multiobjective large neighborhood search metaheuristic for the vehicle routing problem with time windows. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/10/243.

Ma:2020:SAN

[1188] Linmao Ma and Guangmin Wang. A solving algorithm for nonlinear bilevel programing problems based on human evolutionary model. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/260.

Mehrizi:2020:MWE

[1189] Mohammad Abouei Mehrizi and Gianlorenzo D’Angelo. Multi-winner election control via social influence: Hardness and algorithms for restricted cases. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/10/251.

Rapetti:2020:MBR

[1190] Lorenzo Rapetti, Yeshasvi Tirupachuri, Kourosh Darvish, Stefano Dafarra, Gabriele Nava, Claudia Latella, and Daniele Pucci. Model-based real-time motion tracking using dynamical inverse kinematics. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/266.

Reddy:2020:CFI

[1191] Gujji Murali Mohan Reddy, Alan B. Seitenfuss, Débora de Oliveira Medeiros, Luca Meacci, Milton Assunção, and Michael Vynnycky. A compact FEM implementation for parabolic integro-differential equations in 2D. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/242.

Serafino:2020:MFG

[1192] Aldo Serafino, Benoit Obert, and Paola Cinnella. Multi-fidelity gradientbased strategy for robust optimization in computational fluid dynamics. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/248.

Smith:2020:UEA

[1193] Derek H. Smith, Roberto Montemanni, and Stephanie Perkins. The use of an exact algorithm within a tabu search maximum clique algorithm. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/253.

Tariq:2020:UCB

[1194] Zeeshan Tariq, Naveed Khan, Darryl Charles, Sally McClean, Ian Mc-Chesney, and Paul Taylor. Understanding contrail business processes through hierarchical clustering: a multi-stage framework. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/10/244.

Thiruvady:2020:MOB

[1195] Dhananjay Thiruvady, Asef Nazari, and Aldeida Aleti. Multi-objective beam-ACO for maximising reliability and minimising communication overhead in the component deployment problem. Algorithms (Basel), 13 (10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/10/252.

Thiruvady:2020:SMM

[1196] Dhananjay Thiruvady, Christian Blum, and Andreas T. Ernst. Solution merging in matheuristics for resource constrained job scheduling. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/256.

Xiao:2020:CRA

[1197] Xingxing Xiao and Haining Huang. A clustering routing algorithm based on improved ant colony optimization algorithms for underwater wireless sensor networks. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/10/250.

You:2020:WEL

[1198] Gui-Rong You, Yeou-Ren Shiue, Wei-Chang Yeh, Xi-Li Chen, and Chih-Ming Chen. A weighted ensemble learning algorithm based on diversity using a novel particle swarm optimization approach. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/10/255.

Zanzotto:2020:CPD

[1199] Fabio Massimo Zanzotto, Giorgio Satta, and Giordano Cristini. CYK parsing over distributed representations. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/10/262.

Zvarevashe:2020:RCL

[1200] Kudakwashe Zvarevashe and Oludayo O. Olugbara. Recognition of crosslanguage acoustic emotional valence using stacked ensemble learning. Algorithms (Basel), 13(10):??, October 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
10/246.

Bryenton:2020:SSO

[1201] Kyle R. Bryenton, Andrew R. Cameron, Keegan L. A. Kirk, Nasser Saad, Patrick Strongman, and Nikita Volodin. On the solutions of second-order differential equations with polynomial coefficients: Theory, algorithm, application. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/286.

Chen:2020:DGP

[1202] Zhengmao Chen, Dongyue Guo, and Yi Lin. A deep Gaussian processbased flight trajectory prediction approach and its application on conflict detection. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/293.

delAlamo:2020:VMN

[1203] Miguel del Alamo, Housen Li, Axel Munk, and Frank Werner. Variational multiscale nonparametric regression: Algorithms and implementation. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
11/296.

DiMolfetta:2020:SNQ

[1204] Giuseppe Di Molfetta. Searching via nonlinear quantum walk on the 2Dgrid. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/11/305.

Fan:2020:SAP

[1205] Tiffany Fan, David I. Shuman, Shashanka Ubaru, and Yousef Saad. Spectrum-adapted polynomial approximation for matrix functions with applications in graph signal processing. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/11/295.

Franek:2020:CML

[1206] Frantisek Franek and Michael Liut. Computing maximal Lyndon substrings of a string. Algorithms (Basel), 13(11):??, November 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/13/11/294.

He:2020:BDB

[1207] Zhenwen He, Shirong Long, Xiaogang Ma, and Hong Zhao. A boundary distance-based symbolic aggregate approximation method for time series data. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/11/284.

Hussein:2020:GPU

[1208] Eslam A. Hussein, Christopher Thron, Mehrdad Ghaziasgar, Antoine Bagula, and Mattia Vaccari. Groundwater prediction using machinelearning tools. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/300.

Ignacio:2020:LEA

[1209] Paul Samuel Ignacio, Jay-Anne Bulauan, and David Uminsky. Lumáwig: an efficient algorithm for dimension zero bottleneck distance computation in topological data analysis. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/11/291.

Karmitsa:2020:SIN

[1210] Napsu Karmitsa and Sona Taheri. Special issue “Nonsmooth Optimization in Honor of the 60th Birthday of Adil M. Bagirov”: Foreword by Guest Editors. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/282.

Lancia:2020:FBO

[1211] Giuseppe Lancia and Marcello Dalpasso. Finding the best 3-OPT move in subcubic time. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/306.

Lee:2020:DFJ

[1212] Min-Young Lee and Young Ik Kim. Development of a family of Jarrattlike sixth-order iterative methods for solving nonlinear systems with their basins of attraction. Algorithms (Basel), 13(11):??, November 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/13/11/303.

Lewis:2020:ESI

[1213] Rhyd Lewis. Editorial for the special issue on “Algorithms for Graphs and Networks”. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/292.

Liu:2020:ISB

[1214] Guocheng Liu, Caixia Zhang, Qingyang Xu, Ruoshi Cheng, Yong Song, Xianfeng Yuan, and Jie Sun. I3D-shufflenet based human action recognition. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/11/301.

Markgraf:2020:AMK

[1215] Wenke Markgraf, Jannis Lilienthal, Philipp Feistel, Christine Thiele, and Hagen Malberg. Algorithm for mapping kidney tissue water content during normothermic machine perfusion using hyperspectral imaging. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
11/289.

Medboen:2020:COS

[1216] Carl Axel Benjamin Medbøen, Magnus Bolstad Holm, Mohamed Kais Msakni, Kjetil Fagerholt, and Peter Schütz. Combining optimization and simulation for designing a robust short-sea feeder network. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/11/304.

Minghim:2020:GFT

[1217] Rosane Minghim, Liz Huancapaza, Erasmo Artur, Guilherme P. Telles, and Ivar V. Belizario. Graphs from features: Tree-based graph layout for feature analysis. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/302.

Pasqualini:2020:PRN

[1218] Luca Pasqualini. Pseudo random number generation through reinforcement learning and recurrent neural networks. Algorithms (Basel), 13(11): ??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/13/11/307.

Patnana:2020:SLS

[1219] Naresh Patnana. Self-learning salp swarm optimization based PID design of Doha RO plant. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/11/287.

Pegoraro:2020:ETS

[1220] Marco Pegoraro. Efficient time and space representation of uncertain event data. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/285.

Perez-Messina:2020:MCO

[1221] Ignacio Pérez-Messina. Modalflow: Cross-origin flow data visualization for urban mobility. Algorithms (Basel), 13(11):??, November 2020. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/13/11/298.

Rezapour:2020:ABH

[1222] Mahdi Rezapour. Application of Bayesian hierarchical negative binomial finite mixture model for cost–benefit analysis of barriers optimization, accounting for severe heterogeneity. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/11/288.

Sheng:2020:IIN

[1223] Jinfang Sheng. Identifying influential nodes of complex networks based on trust-value. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/280.

Sikansi:2020:SDE

[1224] Fabio Sikansi. Similarity-driven edge bundling: Data-oriented clutter reduction in graphs layouts. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/11/290.

Stachowiak:2020:CEM

[1225] Maria Katarzyna Stachowiak. Cross-entropy method in application to the SIRC model. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/281.

Stanovov:2020:DEL

[1226] Vladimir Stanovov. Differential evolution with linear bias reduction in parameter adaptation. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/11/283.

Thanajiranthorn:2020:ERG

[1227] Chartwut Thanajiranthorn. Efficient rule generation for associative classification. Algorithms (Basel), 13(11):??, November 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/11/299.

Tymochko:2020:UZP

[1228] Sarah Tymochko. Using zigzag persistent homology to detect Hopf bifurcations in dynamical systems. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/11/278.

vanZelst:2020:TWN

[1229] Sebastiaan J. van Zelst. Translating workflow nets to process trees: an algorithmic approach. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/11/279.

Zou:2020:NMD

[1230] Hongjie Zou. A novel multi-dimensional composition method based on time series similarity for array pulse wave signals detecting. Algorithms (Basel), 13(11):??, November 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/11/297.

Abualigah:2020:NIO

[1231] Laith Abualigah, Amir H. Gandomi, Mohamed Abd Elaziz, Abdelazim G. Hussien, Ahmad M. Khasawneh, Mohammad Alshinwan, and Essam H. Houssein. Nature-inspired optimization algorithms for text document clustering — a comprehensive analysis. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/12/345.

Bokam:2020:SCA

[1232] Jagadish Kumar Bokam, Naresh Patnana, Tarun Varshney, and Vinay Pratap Singh. Sine cosine algorithm assisted FOPID controller design for interval systems using reduced-order modeling ensuring stability. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
13/12/317.

Chen:2020:DFL

[1233] Xin Chen and Ying Li. Deep feature learning with manifold embedding for robust image retrieval. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/12/318.

DeBoi:2020:FKT

[1234] Ivan De Boi, Bart Ribbens, Pieter Jorissen, and Rudi Penne. Feasibility of Kd-trees in Gaussian process regression to partition test points in high resolution input space. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/12/327.

deCelis:2020:ANN

[1235] Raúl de Celis, Pablo Solano, and Luis Cadarso. Applying neural networks in aerial vehicle guidance to simplify navigation systems. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/12/333.

DelaSen:2020:CSI

[1236] Manuel De la Sen and Asier Ibeas. On a controlled Se(Is)(Ih)(Iicu)AR epidemic model with output controllability issues to satisfy hospital constraints on hospitalized patients. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/13/12/322.

Fernandez:2020:HMA

[1237] Stephanie Alvarez Fernandez, Marcelo M. Carvalho, and Daniel G. Silva. A hybrid metaheuristic algorithm for the efficient placement of UAVs. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/323.

Ganbold:2020:SBO

[1238] Odkhishig Ganbold, Kaustav Kundu, Haobin Li, and Wei Zhang. A simulation-based optimization method for warehouse worker assignment. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/326.

Gesnouin:2020:PIP

[1239] Joseph Gesnouin, Steve Pechberti, Guillaume Bresson, Bogdan Stanciulescu, and Fabien Moutarde. Predicting intentions of pedestrians from 2D skeletal pose sequences with a representation-focused multi-branch deep learning network. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/12/331.

Ghiani:2020:LPH

[1240] Gianpaolo Ghiani, Tommaso Adamo, Pierpaolo Greco, and Emanuela Guerriero. Lifting the performance of a heuristic for the time-dependent travelling salesman problem through machine learning. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/12/340.

Guo:2020:EAI

[1241] Ru Guo, Xiaodong Qiu, and Yiyi He. Evaluation of agricultural investment climate in CEE countries: The application of back propagation neural network. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/12/336.

Hagiwara:2020:HAL

[1242] Takeo Hagiwara and Tatsuie Tsukiji. Hardness of approximation for Langton’s ant on a twisted torus. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/13/12/344.

Huang:2020:HTE

[1243] Ting Huang, Zhengping Weng, Gang Liu, and Zhenwen He. HD-Tree: an efficient high-dimensional virtual index structure using a half decomposition strategy. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/12/338.

Huang:2020:NCT

[1244] Guojing Huang, Qingliang Chen, and Congjian Deng. A new clickthrough rates prediction model based on Deep&Cross network. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/342.

Ismail:2020:SBC

[1245] Mohamed Ismail and Milica Orlandić. Segment-based clustering of hyperspectral images using tree-based data partitioning structures. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/330.

Jia:2020:ISM

[1246] Hao Jia, Chen Guo, Lina Zhao, and Zhao Xu. Improved sliding mode finite-time synchronization of chaotic systems with unknown parameters. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/346.

Josyula:2020:EFA

[1247] Sai Prashanth Josyula, Johanna Törnquist Krasemann, and Lars Lundberg. An evaluation framework and algorithms for train rescheduling. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/332.

Kanari:2020:TBB

[1248] Lida Kanari, Adélie Garin, and Kathryn Hess. From trees to barcodes and back again: Theoretical and statistical perspectives. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/12/335.

Koukoutsis:2020:DLE

[1249] Elias Koukoutsis, Constantin Papaodysseus, George Tsavdaridis, Nikolaos V. Karadimas, Athanasios Ballis, Eirini Mamatsi, and Athanasios Rafail Mamatsis. Design limitations, errors and hazards in creating decision support platforms with large- and very large-scale data and program cores. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/12/341.

Kuhn:2020:AEG

[1250] Mathias Kühn, Michael Völker, and Thorsten Schmidt. An algorithm for efficient generation of customized priority rules for production control in project manufacturing with stochastic job processing times. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/12/337.

Kusper:2020:CSC

[1251] Gábor Kusper and Csaba Biró. Convert a strongly connected directed graph to a black-and-white 3-SAT problem by the Balatonboglár model. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/321.

Li:2020:PSS

[1252] Jonathan Li, Rohan Potru, and Farhad Shahrokhi. A performance study of some approximation algorithms for computing a small dominating set in a graph. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/12/339.

Mihaescu:2020:PRI

[1253] Roxana-Elena Mihaescu, Mihai Chindea, Constantin Paleologu, Serban Carata, and Marian Ghenescu. Person re-identification across data distributions based on general purpose DNN object detector. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/13/12/343.

Mohamed:2020:FBM

[1254] Emad Mohamed, Parinaz Jafari, and Simaan AbouRizk. Fuzzy-based multivariate analysis for input modeling of risk assessment in wind farm projects. Algorithms (Basel), 13(12):??, December 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/12/325.

Mzyk:2020:KIN

[1255] Grzegorz Mzyk, Zygmunt Hasiewicz, and Paweł Mielcarek. Kernel identification of non-linear systems with general structure. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/13/12/328.

Ruggieri:2020:HSE

[1256] Andrea Ruggieri, Francesco Stranieri, Fabio Stella, and Marco Scutari. Hard and soft EM in Bayesian network learning from incomplete data. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/13/
12/329.

Szenasi:2020:SIB

[1257] Sándor Szénási and Gábor Kertész. Special issue on bio-inspired algorithms for image processing. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/12/320.

Vasileva:2020:SNO

[1258] Maria T. Vasileva. Some notes on the Omega distribution and the pliant probability distribution family. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/13/12/324.

Vendrow:2020:FSL

[1259] Joshua Vendrow, Jamie Haddock, Deanna Needell, and Lorraine Johnson. Feature selection from Lyme disease patient survey using machine learning. Algorithms (Basel), 13(12):??, December 2020. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/13/12/334.

Xi:2020:GMS

[1260] Wang Xi, Guillaume Devineau, Fabien Moutarde, and Jie Yang. Generative model for skeletal human movements based on conditional DC-GAN applied to pseudo-images. Algorithms (Basel), 13(12):??, December 2020. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/13/12/319.

Bawazeer:2021:NAR

[1261] Saleh A. Bawazeer, Saleh S. Baakeem, and Abdulmajeed A. Mohamad. New approach for radial basis function based on partition of unity of Taylor series expansion with respect to shape parameter. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/1.

Chereji:2021:SMC

[1262] Emanuel Chereji, Mircea-Bogdan Radac, and Alexandra-Iulia Szedlak-Stinean. Sliding mode control algorithms for anti-lock braking systems with performance comparisons. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/1/2.

Gapeev:2021:PAC

[1263] Pavel V. Gapeev, Libo Li, and Zhuoshu Wu. Perpetual American cancellable standard options in models with last passage times. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/1/3.

Borndorfer:2021:DCA

[1264] Ralf Borndörfer, Fabian Danecker, and Martin Weiser. A discrete– continuous algorithm for free flight planning. Algorithms (Basel), 14 (1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/4.

Koppl:2021:RPS

[1265] Dominik Köppl, Tomohiro I, Isamu Furuya, Yoshimasa Takabatake, Kensuke Sakai, and Keisuke Goto. Re-pair in small space. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/5.

Hamalainen:2021:ISK

[1266] Joonas Hämäläinen, Tommi Kärkkäinen, and Tuomo Rossi. Improving scalable K-means++. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/1/6.

Ji:2021:MNW

[1267] Kui Ji and Jianxiao Ma. A modified network-wide road capacity reliability analysis model for improving transportation sustainability. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/1/7.

Bilo:2021:HAP

[1268] Davide Bilò, Luciano Gualà, and Guido Proietti. Hardness of an asymmetric 2-player Stackelberg network pricing game. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/8.

Costalonga:2021:CMC

[1269] João Paulo Costalonga, Robert J. Kingan, and Sandra R. Kingan. Constructing minimally 3-connected graphs. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/1/9.

Nebeluk:2021:TMM

[1270] Robert Nebeluk and Maciej Lawryńczuk. Tuning of multivariable model predictive control for industrial tasks. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/1/10.

Galvani:2021:BNL

[1271] Marta Galvani, Chiara Bardelli, Silvia Figini, and Pietro Muliere. A Bayesian nonparametric learning approach to ensemble models using the proper Bayesian bootstrap. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/1/11.

Ishtaiwi:2021:DCS

[1272] Abdelraouf Ishtaiwi, Feda Alshahwan, Naser Jamal, Wael Hadi, and Muhammad AbuArqoub. A dynamic clause specific initial weight assignment for solving satisfiability problems using local search. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/1/12.

Guzzo:2021:SIP

[1273] Antonella Guzzo. Special issue on process mining and emerging applications. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/1/13.

Prezza:2021:SQC

[1274] Nicola Prezza. Subpath queries on compressed graphs: a survey. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/1/14.

Bilo:2021:NEN

[1275] Vittorio Bilò, Michele Flammini, and Luca Moscardelli. On Nash equilibria in non-cooperative all-optical networks. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/1/15.

Al-Afandi:2021:AGL

[1276] Jalal Al-Afandi and András Horváth. Adaptive gene level mutation. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
1/16.

Nakasi:2021:MAD

[1277] Rose Nakasi, Ernest Mwebaze, and Aminah Zawedde. Mobile-aware deep learning algorithms for malaria parasites and white blood cells localization in thick blood smears. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/1/17.

Li:2021:QSD

[1278] Michael Li, Santoso Wibowo, Wei Li, and Lily D. Li. Quantitative spectral data analysis using extreme learning machines algorithm incorporated with PCA. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/1/18.

Munoz:2021:SEA

[1279] Mario Andrés Muñoz and Michael Kirley. Sampling effects on algorithm selection for continuous black-box optimization. Algorithms (Basel), 14 (1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/19.

Serrano-Hernandez:2021:UGD

[1280] Adrian Serrano-Hernandez, Rocio de la Torre, Luis Cadarso, and Javier Faulin. Urban e-grocery distribution design in Pamplona (Spain) applying an agent-based simulation model with horizontal cooperation scenarios. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/1/
20.

Hansknecht:2021:DSP

[1281] Christoph Hansknecht, Imke Joormann, and Sebastian Stiller. Dynamic shortest paths methods for the time-dependent TSP. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/21.

Lee:2021:AAS

[1282] Chuan-Min Lee. Algorithmic aspects of some variations of clique transversal and clique independent sets on graphs. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/1/22.

Rabe:2021:SAE

[1283] Markus Rabe, Majsa Ammouriova, Dominik Schmitt, and Felix Dross. Simheuristics approaches for efficient decision-making support in materials trading networks. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/1/23.

Remoaldo:2021:ATF

[1284] Diogo Remoaldo and Isabel Jesus. Analysis of a traditional and a fuzzy logic enhanced perturb and observe algorithm for the MPPT of a photovoltaic system. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/1/24.

Marusak:2021:ACD

[1285] Piotr M. Marusak. Advanced construction of the dynamic matrix in numerically efficient fuzzy MPC algorithms. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/1/25.

Xue:2021:CEG

[1286] Yiran Xue, Rui Wu, Jiafeng Liu, and Xianglong Tang. Crowd evacuation guidance based on combined action reinforcement learning. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/1/26.

Kaye:2021:MAE

[1287] Bi Kouaï Bertin Kayé, Moustapha Diaby, Moussa Koivogui, and Souleymane Oumtanaga. A memetic algorithm for an external depot production routing problem. Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/1/27.

Melidis:2021:CPD

[1288] Damianos P. Melidis and Wolfgang Nejdl. Capturing protein domain structure and function using self-supervision on domain architectures.

Algorithms (Basel), 14(1), January 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
1/28.

AbdelAziz:2021:MOO

[1289] Amr Mohamed AbdelAziz, Louai Alarabi, Saleh Basalamah, and Abdeltawab Hendawi. A multi-objective optimization method for hospital admission problem —- a case study on Covid-19 patients. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/2/38.

Lassance:2021:RDN

[1290] Carlos Lassance, Vincent Gripon, and Antonio Ortega. Representing deep neural networks latent space geometries with graphs. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/2/39.

Malan:2021:SAL

[1291] Katherine Mary Malan. A survey of advances in landscape analysis for optimisation. Algorithms (Basel), 14(2), February 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/40.

Rabe:2021:SOA

[1292] Markus Rabe, Jesus Gonzalez-Feliu, Jorge Chicaiza-Vaca, and Rafael D. Tordecilla. Simulation-optimization approach for multi-period facility location problems with forecasted and random demands in a last-mile logistics application. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/41.

Kastner:2021:ISB

[1293] Marvin Kastner, Nicole Nellen, Anne Schwientek, and Carlos Jahn. Integrated simulation-based optimization of operational decisions at container terminals. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/42.

delasCasas:2021:FDM

[1294] Pedro Maristany de las Casas, Ralf Borndörfer, Luitgard Kraus, and Antonio Sedeño-Noda. An FPTAS for dynamic multiobjective shortest path problems. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/2/43.

Koppl:2021:NOL

[1295] Dominik Köppl. Non-overlapping LZ77 factorization and LZ78 substring compression queries with suffix trees. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/14/2/44.

Tordecilla:2021:CHS

[1296] Rafael D. Tordecilla, Pedro J. Copado-Méndez, Javier Panadero, Carlos L. Quintero-Araujo, Jairo R. Montoya-Torres, and Angel A. Juan. Combining heuristics with simulation and fuzzy logic to solve a flexiblesize location routing problem under uncertainty. Algorithms (Basel), 14 (2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/2/45.

Dzyura:2021:DOP

[1297] Volodymyr Dzyura, Pavlo Maruschak, and Olegas Prentkovskis. Determining optimal parameters of regular microrelief formed on the end surfaces of rotary bodies. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/2/46.

Mo:2021:MOR

[1298] Xiaoting Mo, Xinglu Liu, and Wai Kin (Victor) Chan. Modeling and optimization in resource sharing systems: Application to bike-sharing with unequal demands. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/47.

Farago:2021:SAE

[1299] András Faragó and Zohre R. Mojaveri. Safe approximation — an efficient solution for a hard routing problem. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/2/48.

Sarica:2021:NAC

[1300] Alessia Sarica, Maria Grazia Vaccaro, Andrea Quattrone, and Aldo Quattrone. A novel approach for cognitive clustering of Parkinsonisms through affinity propagation. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/2/49.

Wang:2021:CWM

[1301] Li Wang, Xi Wang, and Xilong Cai. The common warehouse model and profit distribution of the express industry. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/2/50.

Kulathunga:2021:ENN

[1302] Nalinda Kulathunga, Nishath Rajiv Ranasinghe, Daniel Vrinceanu, Zackary Kinsman, Lei Huang, and Yunjiao Wang. Effects of nonlinearity and network architecture on the performance of supervised neural networks. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/2/51.

Sun:2021:OMC

[1303] Zhichao Sun, Kang Zhou, Xinzheng Yang, Xiao Peng, and Rui Song. Optimization method of customized shuttle bus lines under random condition. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/2/52.

Jin:2021:KMC

[1304] Qibing Jin, Nan Lin, and Yuming Zhang. K-means clustering algorithm based on chaotic adaptive artificial bee colony. Algorithms (Basel), 14 (2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/2/53.

Fu:2021:GCI

[1305] Chen Fu and Jianhua Yang. Granular classification for imbalanced datasets: a Minkowski distance-based method. Algorithms (Basel), 14 (2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/2/54.

Bianchi:2021:OCS

[1306] Enrico Bianchi and Paolo Penna. Optimal clustering in stable instances using combinations of exact and noisy ordinal queries. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/2/55.

Sharma:2021:CTC

[1307] Gokarna Sharma, Ramachandran Vaidyanathan, and Jerry L. Trahan. Constant-time complete visibility for robots with lights: The asynchronous case. Algorithms (Basel), 14(2), February 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/56.

Feng:2021:SBB

[1308] Ryan Feng, Yu Yao, and Ella Atkins. Smart Black Box 2.0: Efficient highbandwidth driving data collection based on video anomalies. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/2/57.

Ferone:2021:AQR

[1309] Alessio Ferone and Antonio Maratea. Adaptive quick reduct for feature drift detection. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/58.

Zoun:2021:IAT

[1310] Roman Zoun, Kay Schallert, David Broneske, Ivayla Trifonova, Xiao Chen, Robert Heyer, Dirk Benndorf, and Gunter Saake. An investigation of alternatives to transform protein sequence databases to a columnar index schema. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/59.

Ohmori:2021:PDI

[1311] Shunichi Ohmori and Kazuho Yoshimoto. A primal-dual interior-point method for facility layout problem with relative-positioning constraints. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
2/60.

Liu:2021:NIG

[1312] Kuan Liu, Haiyuan Liu, Dongyan Sun, and Lei Zhang. Network inference from gene expression data with distance correlation and network topology centrality. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/2/61.

Bhagat:2021:CFE

[1313] Subhash Bhagat, Bibhuti Das, Abhinav Chakraborty, and Krishnendu Mukhopadhyaya. k-circle formation and k-epf by asynchronous robots. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
2/62.

Hagan:2021:MSO

[1314] Ronald D. Hagan and Michael A. Langston. Molecular subtyping and outlier detection in human disease using the paraclique algorithm. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/2/63.

Vavrek:2021:NFI

[1315] Roman Vavrek, Jirí Becica, Viera Papcunová, Petra Gundová, and Jana Mitríková. Number of financial indicators as a factor of multi-criteria analysis via the TOPSIS technique: a municipal case study. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/2/64.

Hucke:2021:ARR

[1316] Danny Hucke and Carl Philipp Reh. Approximation ratios of RePair, LongestMatch and Greedy on unary strings. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/2/65.

Champion:2021:DRV

[1317] Camille Champion, Anne-Claire Brunet, Rémy Burcelin, Jean-Michel Loubes, and Laurent Risser. Detection of representative variables in complex systems with interpretable rules using core-clusters. Algorithms (Basel), 14(2), February 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/2/66.

Nakabe:2021:OCP

[1318] Jin Nakabe, Teruhiro Mizumoto, Hirohiko Suwa, and Keiichi Yasumoto. Optimal cooking procedure presentation system for multiple recipes and investigating its effect. Algorithms (Basel), 14(2), February 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/2/67.

Gurvich:2021:CHM

[1319] Vladimir Gurvich and Mikhail Vyalyi. On computational hardness of multidimensional subtraction games. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/3/71.

Tonti:2021:FOD

[1320] Luca Tonti and Alessandro Patti. Fast overlap detection between hardcore colloidal cuboids and spheres. The OCSI algorithm. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/3/72.

Fotakis:2021:OFL

[1321] Dimitris Fotakis, Loukas Kavouras, and Lydia Zakynthinou. Online facility location in evolving metrics. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/3/73.

Saffre:2021:ABS

[1322] Fabrice Saffre and Hanno Hildmann. Adaptive behaviour for a selforganising video surveillance system using a genetic algorithm. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/3/74.

Mahmood:2021:DLM

[1323] Usman Mahmood, Zening Fu, Vince D. Calhoun, and Sergey Plis. A deep learning model for data-driven discovery of functional connectivity. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/3/75.

Lucena-Sanchez:2021:FLS

[1324] Estrella Lucena-Sánchez, Guido Sciavicco, and Ionel Eduard Stan. Feature and language selection in temporal symbolic regression for interpretable air quality modelling. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/3/76.

Alabbadi:2021:MOT

[1325] Afra A. Alabbadi and Maysoon F. Abulkhair. Multi-objective task scheduling optimization in spatial crowdsourcing. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/3/77.

Lang:2021:ELA

[1326] Ryan Dieter Lang and Andries Petrus Engelbrecht. An exploratory landscape analysis-based benchmark suite. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/3/78.

Bouamama:2021:IGH

[1327] Salim Bouamama and Christian Blum. An improved greedy heuristic for the minimum positive influence dominating set problem in social networks. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/3/
79.

Han:2021:DAC

[1328] Qiuqi Han, Guangyuan Zheng, and Chen Xu. D2D assisted cellular networks in licensed and unlicensed spectrum: Matching-iteration-based joint user access and resource allocation. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/3/80.

Fichte:2021:DDP

[1329] Johannes K. Fichte, Markus Hecher, Michael Morak, and Stefan Woltran. DynASP2.5: Dynamic programming on tree decompositions in action. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/3/81.

Li:2021:IRI

[1330] Xu Li and Qiming Sun. Identifying and ranking influential nodes in complex networks based on dynamic node strength. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/3/82.

Yuan:2021:TIF

[1331] Shijin Yuan, Cheng Wang, Bin Mu, Feifan Zhou, and Wansuo Duan. Typhoon intensity forecasting based on LSTM using the rolling forecast method. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/3/83.

Rezapour:2021:AAN

[1332] Mahdi Rezapour and Khaled Ksaibati. Accounting for attribute nonattendance and common-metric aggregation in the choice of seat belt use, a latent class model with preference heterogeneity. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/3/84.

Rauh:2021:TUL

[1333] Andreas Rauh and Julia Kersten. Transformation of uncertain linear systems with real eigenvalues into cooperative form: The case of constant and time-varying bounded parameters. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/3/85.

Razgon:2021:CRM

[1334] Margarita Razgon and Alireza Mousavi. Correction: Razgon, M., et al. Relaxed Rule-Based Learning for Automated Predictive Maintenance: Proof of Concept. Algorithms 2020, 13, 219. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/3/86. See [1153].

Aivodji:2021:LDD

[1335] Ulrich Aïvodji, François Bidet, Sébastien Gambs, Rosin Claude Ngueveu, and Alain Tapp. Local data debiasing for fairness based on generative adversarial training. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/3/87.

Rauh:2021:UIO

[1336] Andreas Rauh, Auguste Bourgois, and Luc Jaulin. Union and intersection operators for thick ellipsoid state enclosures: Application to bounded-error discrete-time state observer design. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/3/88.

Li:2021:FMC

[1337] Xinsheng Li and Xuedong Yuan. Fundamental matrix computing based on 3D metrical distance. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/3/89.

Strasser:2021:SEF

[1338] Ben Strasser, Dorothea Wagner, and Tim Zeitz. Space-efficient, fast and exact routing in time-dependent road networks. Algorithms (Basel), 14 (3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/3/90.

Azam:2021:UFS

[1339] Md Ali Azam, Hans D. Mittelmann, and Shankarachary Ragi. UAV formation shape control via decentralized Markov decision processes. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/3/91.

Mamo:2021:APD

[1340] Nicholas Mamo, Joel Azzopardi, and Colin Layfield. An automatic participant detection framework for event tracking on Twitter. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/3/92.

Wang:2021:SMC

[1341] Minghui Wang, Bi Zeng, and Qiujie Wang. Study of motion control and a virtual reality system for autonomous underwater vehicles. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/3/93.

Zisad:2021:INN

[1342] Sharif Noor Zisad, Mohammad Shahadat Hossain, Mohammed Sazzad Hossain, and Karl Andersson. An integrated neural network and SEIR model to predict COVID-19. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/3/94.

Fernandes:2021:TUC

[1343] Luiz Henrique dos Santos Fernandes, Ana Carolina Lorena, and Kate Smith-Miles. Towards understanding clustering problems and algorithms: an instance space analysis. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/3/95.

Bannach:2021:DCC

[1344] Max Bannach and Till Tantau. On the descriptive complexity of color coding. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/3/96.

Genitrini:2021:LUC

[1345] Antoine Genitrini and Martin Pépin. Lexicographic unranking of combinations revisited. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/3/97.

Wang:2021:MHC

[1346] Huimu Wang, Zhen Liu, Jianqiang Yi, and Zhiqiang Pu. Multiagent hierarchical cognition difference policy for multiagent cooperation. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/3/98.

Zheng:2021:MMC

[1347] Yang Zheng, Jieyu Zhao, Yu Chen, Chen Tang, and Shushi Yu. 3D mesh model classification with a capsule network. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/3/99.

Mostert:2021:FSA

[1348] Werner Mostert, Katherine M. Malan, and Andries P. Engelbrecht. A feature selection algorithm performance metric for comparative analysis. Algorithms (Basel), 14(3), March 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/3/100.

Cordero:2021:CSN

[1349] Alicia Cordero, Marlon Moscoso-Martínez, and Juan R. Torregrosa. Chaos and stability in a new iterative family for solving nonlinear equations. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/4/
101.

Zhang:2021:PID

[1350] Lin Zhang, Haiyuan Liu, and Hao He. Prediction of intrinsically disordered proteins using machine learning algorithms based on fuzzy entropy feature. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/4/
102.

Georgoulaki:2021:EIC

[1351] Eirini Georgoulaki, Kostas Kollias, and Tami Tamir. Equilibrium inefficiency and computation in cost-sharing games in real-time scheduling systems. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/4/103.

Tzagkarakis:2021:OBG

[1352] Eleftherios Tzagkarakis, Haridimos Kondylakis, George Vardakis, and Nikolaos Papadakis. Ontology based governance for employee services. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/4/104.

Cicerone:2021:QHD

[1353] Serafino Cicerone. A quasi-hole detection algorithm for recognizing kdistance-hereditary graphs, with k < 2. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/4/105.

Nemec:2021:OMF

[1354] Petr Nemec and Petr Stodola. Optimization of the multi-facility location problem using widely available office software. Algorithms (Basel), 14 (4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/4/106.

Xu:2021:IID

[1355] Pengchang Xu, Jiaxiang Zhao, and Jie Zhang. Identification of intrinsically disordered protein regions based on deep neural network-VGG16. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/4/107.

Yang:2021:ROM

[1356] Xinfeng Yang and Yicheng Qi. Research on optimization of multiobjective regional public transportation scheduling. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/4/108.

Dey:2021:MFA

[1357] Subhrajit Dey, Rajdeep Bhattacharya, Friedhelm Schwenker, and Ram Sarkar. Median filter aided CNN based image denoising: An ensemble approach. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/4/109.

Schaller:2021:ACC

[1358] David Schaller, Manuela Geiß, Marc Hellmuth, and Peter F. Stadler. Arc-completion of 2-colored best match graphs to binary-explainable best match graphs. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/4/110.

Oliveira:2021:MCD

[1359] Altina S. Oliveira, Carlos F. S. Gomes, Camilla T. Clarkson, Adriana M. Sanseverino, Mara R. S. Barcelos, Igor P. A. Costa, and Marcos Santos. Multiple criteria decision making and prospective scenarios model for selection of companies to be incubated. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/4/111.

Amirghasemi:2021:EDB

[1360] Mehrdad Amirghasemi. An effective decomposition-based stochastic algorithm for solving the permutation flow-shop scheduling problem. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/4/112.

Schwoebel:2021:REN

[1361] Stephan Daniel Schwoebel, Thomas Mehner, and Thomas Lampke. On a robust and efficient numerical scheme for the simulation of stationary 3-component systems with non-negative species-concentration with an application to the Cu deposition from a Cu-(β-alanine)-electrolyte. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/4/113.

Kasper-Eulaers:2021:SCD

[1362] Margrit Kasper-Eulaers, Nico Hahn, Stian Berger, Tom Sebulonsen, Øystein Myrland, and Per Egil Kummervold. Short communication: Detecting heavy goods vehicles in rest areas in winter conditions using YOLOv5. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/4/114.

Ishtaiwi:2021:DIW

[1363] Abdelraouf Ishtaiwi and Qasem Abu Al-Haija. Dynamic initial weight assignment for MaxSAT. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/4/115.

Mitsuya:2021:CCC

[1364] Shiori Mitsuya, Yuto Nakashima, Shunsuke Inenaga, Hideo Bannai, and Masayuki Takeda. Compressed communication complexity of Hamming distance. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/4/116.

Madokoro:2021:ACP

[1365] Hirokazu Madokoro, Stephanie Nix, and Kazuhito Sato. Automatic calibration of piezoelectric bed-leaving sensor signals using genetic network programming algorithms. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/4/117.

Deon:2021:PCW

[1366] Aleksei F. Deon, Oleg K. Karaduta, and Yulian A. Menyaev. Phase congruential white noise generator. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/4/118.

Miao:2021:FDA

[1367] Qing Miao, Juhui Wei, Jiongqi Wang, and Yuyun Chen. Fault diagnosis algorithm based on adjustable nonlinear PI state observer and its application in UAV fault diagnosis. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/4/119.

Lin:2021:IAB

[1368] Yanhong Lin, Jing Wang, Xiaolin Li, Yuanzi Zhang, and Shiguo Huang. An improved artificial bee colony for feature selection in QSAR. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/4/120.

Stodt:2021:SAB

[1369] Jan Stodt, Daniel Schönle, Christoph Reich, Fatemeh Ghovanlooy Ghajar, Dominik Welte, and Axel Sikora. Security audit of a blockchain-based industrial application platform. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/4/121.

Valdez:2021:BIA

[1370] Fevrier Valdez, Oscar Castillo, and Patricia Melin. Bio-inspired algorithms and its applications for optimization in fuzzy clustering. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/4/122.

Roy:2021:ETL

[1371] Debdatta Sinha Roy, Bruce Golden, Xingyin Wang, and Edward Wasil. Estimating the tour length for the close enough traveling salesman problem. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/4/
123.

Zakaria:2021:TDM

[1372] La Zakaria, Endah Yuliani, and Asmiati Asmiati. A two-dimensional mKdV linear map and its application in digital image cryptography. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/4/124.

Geier:2021:CIR

[1373] Martin Geier and Hussein Alihussein. Computation of implicit representation of volumetric shells with predefined thickness. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/4/125.

Petrova:2021:BDC

[1374] Elena A. Petrova and Arseny M. Shur. Branching densities of cube-free and square-free words. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/4/126.

Odongo:2021:MDC

[1375] George Odongo, Richard Musabe, and Damien Hanyurwimfura. A multinomial DGA classifier for incipient fault detection in oil-impregnated power transformers. Algorithms (Basel), 14(4), April 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/4/128.

Stanovov:2021:DBM

[1376] Vladimir Stanovov, Shakhnaz Akhmedova, and Eugene Semenkin. Difference-based mutation operation for neuroevolution of augmented topologies. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/5/127.

Li:2021:TDE

[1377] Yuan Li, Ni Zhang, Yuejiao Gong, Wentao Mao, and Shiguang Zhang. Three-dimensional elastodynamic analysis employing partially discontinuous boundary elements. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/5/129.

Kazakovtsev:2021:SCE

[1378] Lev Kazakovtsev, Ivan Rozhnov, and Guzel Shkaberina. Self-configuring (1+1)-evolutionary algorithm for the continuous p-median problem with agglomerative mutation. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/5/130.

Vu:2021:AMC

[1379] Martin Vu and Henning Fernau. Adding matrix control: Insertion– deletion systems with substitutions III. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/5/131.

Yousef:2021:CPM

[1380] Malik Yousef and Jens Allmer. Classification of precursor MicroRNAs from different species based on K-mer distance features. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/5/132.

Gibney:2021:TIR

[1381] Daniel Gibney and Sharma V. Thankachan. Text indexing for regular expression matching. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/5/133.

Abdallah:2021:MMO

[1382] Loai Abdallah, Murad Badarna, Waleed Khalifa, and Malik Yousef. MultiKOC: Multi-one-class classifier based K-means clustering. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/5/134.

Demidova:2021:ADP

[1383] Liliya A. Demidova and Julia S. Sokolova. Analysis of data presented by multisets using a linguistic approach. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/5/135.

Bose:2021:MLS

[1384] Aritra Bose, Filippo Utro, Daniel E. Platt, and Laxmi Parida. Multiple loci selection with multi-way epistasis in coalescence with recombination. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/5/136.

Lei:2021:KRA

[1385] Zhou Lei, Kangkang Yang, Kai Jiang, and Shengbo Chen. KDAS-ReID: Architecture search for person re-identification via distilled knowledge with dynamic temperature. Algorithms (Basel), 14(5), May 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/5/137.

Codetta-Raiteri:2021:ESI

[1386] Daniele Codetta-Raiteri. Editorial for the special issue on “Bayesian Networks: Inference Algorithms, Applications, and Software Tools”. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/5/138.

Ciprian:2021:DSU

[1387] Claudio Ciprian, Kirill Masychev, Maryam Ravan, Akshaya Manimaran, and AnkitaAmol Deshmukh. Diagnosing schizophrenia using effective connectivity of resting-state EEG data. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/5/139.

AngeloLellisMoreira:2021:PSM

[1388] Miguel Ângelo Lellis Moreira, Igor Pinheiro de Araújo Costa, Maria Teresa Pereira, Marcos dos Santos, Carlos Francisco Simões Gomes, and Fernando Martins Muradas. PROMETHEE-SAPEVO-M1: a hybrid approach based on ordinal and cardinal inputs: Multi-criteria evaluation of helicopters to support Brazilian Navy operations. Algorithms (Basel), 14 (5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/5/140.

DiCaprio:2021:DOI

[1389] Francesco Di Caprio, Roberto Scigliano, Roberto Fauci, and Domenico Tescione. Design optimization of interfacing attachments for the deployable wing of an unmanned re-entry vehicle. Algorithms (Basel), 14 (5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/5/141.

Louedec:2021:IEK

[1390] Morgan Louédec and Luc Jaulin. Interval extended Kalman filter — application to underwater localization and control. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/5/142.

Thibault:2021:OSK

[1391] Alexis Thibault, Lénaïc Chizat, Charles Dossal, and Nicolas Papadakis. Overrelaxed Sinkhorn–Knopp algorithm for regularized optimal transport. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/5/

Han:2021:BLB

[1392] Yuexing Han, Xiaolong Li, Bing Wang, and Lu Wang. Boundary lossbased 2.5D fully convolutional neural networks approach for segmentation: a case study of the liver and tumor on computed tomography. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/5/144.

Xu:2021:NWJ

[1393] Mingming Xu, Shuning Zhang, and Guanlong Deng. No-wait job shop scheduling using a population-based iterated greedy algorithm. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/5/145.

Vakhnin:2021:IIC

[1394] Aleksei Vakhnin and Evgenii Sopov. Investigation of improved cooperative coevolution for large-scale global optimization problems. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/5/146.

Beacher:2021:MLP

[1395] Felix D. Beacher, Lilianne R. Mujica-Parodi, Shreyash Gupta, and Leonardo A. Ancora. Machine learning predicts outcomes of Phase III clinical trials for prostate cancer. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/5/147.

Park:2021:DTM

[1396] Minhyuk Park, Paul Zaharias, and Tandy Warnow. Disjoint tree mergers for large-scale maximum likelihood tree estimation. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/5/148.

Zervoudakis:2021:QRI

[1397] Petros Zervoudakis, Haridimos Kondylakis, Nicolas Spyratos, and Dimitris Plexousakis. Query rewriting for incremental continuous query evaluation in HIFUN. Algorithms (Basel), 14(5), May 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/5/149.

Cicerone:2021:SIG

[1398] Serafino Cicerone and Gabriele Di Stefano. Special issue on “Graph Algorithms and Applications”. Algorithms (Basel), 14(5), May 2021.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/5/150.

Flammini:2021:TGP

[1399] Michele Flammini, Gianpiero Monaco, Luca Moscardelli, Mordechai Shalom, and Shmuel Zaks. The traffic grooming problem in optical networks with respect to ADMs and OADMs: Complexity and approximation. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/5/
151.

El-Mabrouk:2021:PES

[1400] Nadia El-Mabrouk. Predicting the evolution of syntenies — an algorithmic review. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/5/152.

Pfaltz:2021:STA

[1401] John L. Pfaltz. A set-theoretic approach to modeling network structure. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/5/153.

Wallden:2021:ATC

[1402] Marcus Walldén, Masao Okita, Fumihiko Ino, Dimitris Drikakis, and Ioannis Kokkinakis. Accelerating in-transit co-processing for scientific simulations using region-based data-driven analysis. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/5/154.

Pan:2021:EEP

[1403] Huifang Pan and Qi Zhu. Energy-efficient power allocation in non-linear energy harvesting multiple relay systems. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/5/155.

Arafet:2021:DTS

[1404] Kamel Arafet and Rafael Berlanga. Digital twins in solar farms: an approach through time series and deep learning. Algorithms (Basel), 14 (5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/5/156.

Ullah:2021:PDM

[1405] Rahmat Ullah and Tughrul Arslan. Parallel delay multiply and sum algorithm for microwave medical imaging using Spark Big Data framework.

Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/5/157.

AbdEl-MoamenMohamed:2021:NCC

[1406] Soha Abd El-Moamen Mohamed, Marghany Hassan Mohamed, and Mohammed F. Farghally. A new cascade-correlation growing deep learning neural network algorithm. Algorithms (Basel), 14(5), May 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/5/158.

Sun:2021:CUN

[1407] Feng Sun, Ajith Kumar V, Guanci Yang, Ansi Zhang, and Yiyun Zhang. Circle-U-Net: an efficient architecture for semantic segmentation. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/6/159.

Xu:2021:VAC

[1408] Qiaoji Xu, Lingling Jin, James H. Leebens-Mack, and David Sankoff. Validation of automated chromosome recovery in the reconstruction of ancestral gene order. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/6/160.

Koppl:2021:RLZ

[1409] Dominik Köppl. Reversed Lempel–Ziv factorization with suffix trees. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/6/161.

Huang:2021:CSS

[1410] Yajing Huang and Feng Chen. Community structure and systemic risk of bank correlation networks based on the U.S. financial crisis in 2008. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/6/162.

Li:2021:NHP

[1411] Yaru Li, Yulai Zhang, and Yongping Cai. A new hyper-parameter optimization method for power load forecast based on recurrent neural networks. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/6/
163.

Rupp:2021:LBQ

[1412] Tobias Rupp and Stefan Funke. A lower bound for the query phase of contraction hierarchies and hub labels and a provably optimal instancebased schema. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/6/164.

Guo:2021:ECB

[1413] Shiyu Guo, Mengna Shi, Yanqi Zhou, Jiayin Yu, and Erfu Wang. An efficient convolutional blind source separation algorithm for speech signals under chaotic masking. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/6/165.

Totounferoush:2021:ESI

[1414] Amin Totounferoush, Frédéric Simonis, Benjamin Uekermann, and Miriam Schulte. Efficient and scalable initialization of partitioned coupled simulations with preCICE. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/6/166.

Irawan:2021:LCN

[1415] Agus Irawan, Asmiati Asmiati, La Zakaria, and Kurnia Muludi. The locating-chromatic number of origami graphs. Algorithms (Basel), 14 (6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/6/167.

Ruffini:2021:GDO

[1416] Manon Ruffini, Jelena Vucinic, Simon de Givry, George Katsirelos, Sophie Barbe, and Thomas Schiex. Guaranteed diversity and optimality in cost function network based computational protein design methods. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/6/168.

Bulteau:2021:SMC

[1417] Laurent Bulteau, Guillaume Fertin, Géraldine Jean, and Christian Komusiewicz. Sorting by multi-cut rearrangements. Algorithms (Basel), 14 (6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/6/169.

Balbal:2021:GHM

[1418] Samir Balbal, Salim Bouamama, and Christian Blum. A greedy heuristic for maximizing the lifetime of wireless sensor networks based on disjoint weighted dominating sets. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/6/170.

Poudel:2021:AVT

[1419] Pavan Poudel and Gokarna Sharma. Adaptive versioning in transactional memory systems. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/6/171.

Matsuda:2021:SSF

[1420] Kotaro Matsuda, Shuhei Denzumi, and Kunihiko Sadakane. Storing set families more compactly with top ZDDs. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/6/172.

Zhao:2021:PPT

[1421] Zhuo-Qiang Zhao, Shi-Jian Liu, and Jeng-Shyang Pan. A PID parameter tuning method based on the improved QUATRE algorithm. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/6/173.

Zhao:2021:IDC

[1422] Wenxiao Zhao. An introduction to development of centralized and distributed stochastic approximation algorithm with expanding truncations. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/6/174.

Miranda:2021:AAS

[1423] Guilherme Henrique Santos Miranda, Alexsandro Oliveira Alexandrino, Carla Negri Lintzmayer, and Zanoni Dias. Approximation algorithms for sorting λ-permutations by λ-operations. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/6/175.

Zhu:2021:DTB

[1424] Wei Zhu and Xiaoyang Zeng. Decision tree-based adaptive reconfigurable cache scheme. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/6/176.

Chandiwana:2021:TFH

[1425] Edina Chandiwana, Caston Sigauke, and Alphonce Bere. Twenty-four-hour ahead probabilistic global horizontal irradiance forecasting using Gaussian process regression. Algorithms (Basel), 14(6), June 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/6/177.

Mpanza:2021:OTP

[1426] Lindokuhle J. Mpanza and Jimoh Olarewaju Pedro. Optimised tuning of a PID-based flight controller for a medium-scale rotorcraft. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/6/178.

Wang:2021:APC

[1427] Chunxia Wang, Jun Bi, Qiuyue Sai, and Zun Yuan. Analysis and prediction of carsharing demand based on data mining methods. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/6/179.

Fu:2021:DIL

[1428] Lei Fu, Qizhi Tang, Peng Gao, Jingzhou Xin, and Jianting Zhou. Damage identification of long-span bridges using the hybrid of convolutional neural network and long short-term memory network. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/6/180.

Blueschke:2021:AOC

[1429] Dmitri Blueschke, Viktoria Blueschke-Nikolaeva, and Reinhard Neck. Approximately optimal control of nonlinear dynamic stochastic problems with learning: The OPTCON algorithm. Algorithms (Basel), 14 (6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/6/181.

Arai:2021:SIF

[1430] Hiroshi Arai and Harumi Haraguchi. A study of Ising formulations for minimizing setup cost in the two-dimensional cutting stock problem. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/6/182.

Alorf:2021:PDL

[1431] Abdulaziz Alorf. The practicality of deep learning algorithms in COVID-19 detection: Application to chest X-ray images. Algorithms (Basel), 14 (6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/6/183.

Que:2021:SME

[1432] Xia Que, Siyuan Jiang, Jiaoyun Yang, and Ning An. A similarity measurement with entropy-based weighting for clustering mixed numerical and categorical datasets. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/6/184.

Mylonas:2021:UFE

[1433] Nikos Mylonas and Basil Papadopoulos. Unbiased fuzzy estimators in fuzzy hypothesis testing. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/6/185.

Landro:2021:COM

[1434] Nicola Landro, Ignazio Gallo, and Riccardo La Grassa. Combining optimization methods using an adaptive meta optimizer. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/6/186.

Barbosa:2021:UML

[1435] Aaron Barbosa, Elijah Pelofske, Georg Hahn, and Hristo N. Djidjev. Using machine learning for quantum annealing accuracy prediction. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/6/187.

Yu:2021:SPS

[1436] Dan Yu, Peng Liu, Dezhi Qiao, and Xianglong Tang. A safety prediction system for lunar orbit rendezvous and docking mission. Algorithms (Basel), 14(6), June 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/6/188.

Ibrahim:2021:OTM

[1437] Abdullahi Adinoyi Ibrahim, Alessandro Lonardi, and Caterina De Bacco. Optimal transport in multilayer networks for traffic flow optimization. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/189.

Jnoub:2021:FCR

[1438] Nour Jnoub, Admir Brankovic, and Wolfgang Klas. Fact-checking reasoning system for fake review detection using answer set programming. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/190.

Nemec:2021:OWM

[1439] Petr Nemec, Petr Stodola, Miroslav Pecina, Jirí Neubauer, and Martin Blaha. Optimization of the weighted multi-facility location problem using MS Excel. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/7/191.

Ouyang:2021:CNN

[1440] Kewei Ouyang, Yi Hou, Shilin Zhou, and Ye Zhang. Convolutional neural network with an elastic matching mechanism for time series classification. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/192.

Shamseldin:2021:OCO

[1441] Mohamed A. Shamseldin. Optimal coronavirus optimization algorithm based PID controller for high performance brushless DC motor. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/7/193.

Atchade-Adelomou:2021:QQC

[1442] Parfait Atchade-Adelomou, Guillermo Alonso-Linaje, Jordi Albo-Canals, and Daniel Casado-Fauli. qRobot: a quantum computing approach in mobile robot order picking and batching problem solver optimization. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/194.

Wu:2021:KDN

[1443] Yundong Wu, Jiajia Liao, Yujun Liu, Kaiming Ding, Shimin Li, Zhilin Zhang, Guorong Cai, and Jinhe Su. Knowledge-driven network for object detection. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/7/195.

Kang:2021:SAP

[1444] Yiting Kang, Biao Xue, and Riya Zeng. Self-adaptive path tracking control for mobile robots under slippage conditions based on an RBF neural network. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/
196.

Seman:2021:OSA

[1445] Ali Seman and Azizian Mohd Sapawi. An optimal and stable algorithm for clustering numerical data. Algorithms (Basel), 14(7), July 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/7/197.

Vestias:2021:DMF

[1446] Mário P. Véstias and Horácio C. Neto. Decimal multiplication in FPGA with a novel decimal adder/subtractor. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/7/198.

Yan:2021:CCA

[1447] Jiangyu Yan and Bing Qi. CARA: a congestion-aware routing algorithm for wireless sensor networks. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/7/199.

Saad:2021:EDS

[1448] Suleiman Sa’ad, Abdullah Muhammed, Mohammed Abdullahi, Azizol Abdullah, and Fahrul Hakim Ayob. An enhanced discrete symbiotic organism search algorithm for optimal task scheduling in the cloud. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/7/200.

Villavicencio:2021:CPA

[1449] Charlyn Nayve Villavicencio, Julio Jerison Escudero Macrohon, Xavier Alphonse Inbaraj, Jyh-Horng Jeng, and Jer-Guang Hsieh. COVID-19 prediction applying supervised machine learning algorithms with comparative analysis using WEKA. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/7/201.

Gao:2021:SMP

[1450] Yanfeng Gao, Cicao Ping, Ling Wang, and Binrui Wang. A simplification method for point cloud of T-profile steel plate for shipbuilding. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/202.

Yang:2021:NTL

[1451] Xiuqing Yang, Xinglu Liu, Lijuan Feng, Jianquan Zhang, and Mingyao Qi. Non-traditional layout design for robotic mobile fulfillment system with multiple workstations. Algorithms (Basel), 14(7), July 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/7/203.

Ma:2021:CSB

[1452] Wenpeng Ma, Wu Yuan, and Xiazhen Liu. A comparative study of block incomplete sparse approximate inverses preconditioning on Tesla K20 and V100 GPUs. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/7/204.

Rauh:2021:ISL

[1453] Andreas Rauh, Robert Dehnert, Swantje Romig, Sabine Lerch, and Bernd Tibken. Iterative solution of linear matrix inequalities for the combined control and observer design of systems with polytopic parameter uncertainty and stochastic noise. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/7/205.

Salah:2021:EMM

[1454] Omar Salah, Abdulrahim Shamayleh, and Shayok Mukhopadhyay. Energy management of a multi-source power system. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/7/206.

Argyros:2021:EHO

[1455] Ioannis K. Argyros, Debasis Sharma, Christopher I. Argyros, Sanjaya Kumar Parhi, Shanta Kumari Sunanda, and Michael I. Argyros. Extended high order algorithms for equations under the same set of conditions. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/
207.

Zhang:2021:PCP

[1456] Jinsong Zhang, Yongtao Peng, Bo Ren, and Taoying Li. PM2.5 concentration prediction based on CNN-BiLSTM and attention mechanism. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/208.

Huang:2021:CCF

[1457] Mingyang Huang, Chenglin Liu, and Liang Shan. Containment control of first-order multi-agent systems under PI coordination protocol. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/7/209.

Gonzalez-Neira:2021:MSA

[1458] Eliana Maria Gonzalez-Neira, Jairo R. Montoya-Torres, and Jose-Fernando Jimenez. A multicriteria simheuristic approach for solving a stochastic permutation flow shop scheduling problem. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/7/210.

Sadawi:2021:EDC

[1459] Alia Al Sadawi, Abdulrahim Shamayleh, and Malick Ndiaye. Efficient dynamic cost scheduling algorithm for financial data supply chain. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/7/211.

Skandarani:2021:DLB

[1460] Youssef Skandarani, Pierre-Marc Jodoin, and Alain Lalande. Deep learning based cardiac MRI segmentation: Do we need experts? Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/7/212.

Zisad:2021:IDL

[1461] Sharif Noor Zisad, Etu Chowdhury, Mohammad Shahadat Hossain, Raihan Ul Islam, and Karl Andersson. An integrated deep learning and belief rule-based expert system for visual sentiment analysis under uncertainty. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/7/213.

Ebrahimi:2021:HAI

[1462] Sara Ebrahimi, Aminah Robinson Fayek, and Vuppuluri Sumati. Hybrid artificial intelligence HFS-RF-PSO model for construction labor productivity prediction and optimization. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/7/214.

Bhatti:2021:DFH

[1463] Faraz Bhatti and Thomas Greiner. Design of an FPGA hardware optimizing the performance and power consumption of a plenoptic camera depth estimation algorithm. Algorithms (Basel), 14(7), July 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/7/215.

Muaad:2021:AND

[1464] Abdullah Y. Muaad, Hanumanthappa Jayappa, Mugahed A. Al-antari, and Sungyoung Lee. ArCAR: a novel deep learning computer-aided recognition for character-level Arabic text representation and recognition. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/7/
216.

Putro:2021:DMA

[1465] Heru Purboyo Hidayat Putro, Pradono Pradono, and Titus Hari Setiawan. Development of multi-actor multi-criteria analysis based on the weight of stakeholder involvement in the assessment of natural-cultural tourism area transportation policies. Algorithms (Basel), 14(7), July 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/7/217.

Roque:2021:ICH

[1466] João V. Roque, João D. Lopes, Mário P. Véstias, and José T. de Sousa. IOb-cache: a high-performance configurable open-source cache. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/8/218.

Thiruvady:2021:ASI

[1467] Dhananjay Thiruvady, Kerri Morgan, Susan Bedingfield, and Asef Nazari. Allocating students to industry placements using integer programming and ant colony optimisation. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/14/8/219.

Kontogiannis:2021:TDA

[1468] Spyros Kontogiannis, Andreas Paraskevopoulos, and Christos Zaroliagis. Time-dependent alternative route planning: Theory and practice. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/8/220.

Du:2021:IGS

[1469] Zhihui Du, Oliver Alvarado Rodriguez, Joseph Patchett, and David A. Bader. Interactive graph stream analytics in arkouda. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/221.

Tekile:2021:CEM

[1470] Hailemariam Abebe Tekile, Michele Fedrizzi, and Matteo Brunelli. Constrained eigenvalue minimization of incomplete pairwise comparison matrices by Nelder–Mead algorithm. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/8/222.

Ullah:2021:FPR

[1471] Kifayat Ullah, Muhammad Safi Ullah Khan, and Manuel de la Sen. Fixed point results on multi-valued generalized (α, β)-nonexpansive mappings in Banach spaces. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/8/223.

Selva:2021:INI

[1472] Deepaa Selva, Balakrishnan Nagaraj, Danil Pelusi, Rajendran Arunkumar, and Ajay Nair. Intelligent network intrusion prevention feature collection and classification algorithms. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/8/224.

Yamada:2021:SSS

[1473] Masataka Yamada and Akihiro Inokuchi. Similar supergraph search based on graph edit distance. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/8/225.

vonPilchau:2021:SEA

[1474] Wenzel Pilar von Pilchau, Anthony Stein, and Jörg Hähner. Synthetic experiences for accelerating DQN performance in discrete non-deterministic environments. Algorithms (Basel), 14(8), August 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/8/226.

Salleh:2021:MLS

[1475] Zabidin Salleh, Ghaliah Alhamzi, Ibitsam Masmali, and Ahmad Alhawarat. A modified Liu and Storey conjugate gradient method for large scale unconstrained optimization problems. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/8/227.

Martey:2021:IRU

[1476] Ezekiel Mensah Martey, Hang Lei, Xiaoyu Li, and Obed Appiah. Image representation using stacked colour histogram. Algorithms (Basel), 14 (8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/8/228.

Li:2021:ETV

[1477] Fangyi Li, Yufei Yan, Jianhua Rong, and Houyao Zhu. An efficient timevariant reliability analysis method with mixed uncertainties. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/229.

Syafruddin:2021:BSM

[1478] Willa Ariela Syafruddin, Rio Mukhtarom Paweroi, and Mario Köppen. Behavior selection metaheuristic search algorithm for the pollination optimization: a simulation case of cocoa flowers. Algorithms (Basel), 14 (8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/8/230.

Weiner:2021:IDT

[1479] Samson Weiner and Mukul S. Bansal. Improved duplication-transfer-loss reconciliation with extinct and unsampled lineages. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/8/231.

Jatschka:2021:GCO

[1480] Thomas Jatschka, Günther R. Raidl, and Tobias Rodemann. A general cooperative optimization approach for distributing service points in mobility applications. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/8/232.

Caprio:2021:SIA

[1481] Debora Di Caprio and Francisco Javier Santos Arteaga. Special issue on algorithms and models for dynamic multiple criteria decision making. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/8/233.

Sahin:2021:MSC

[1482] Bekir Sahin, Devran Yazir, Abdelsalam Adam Hamid, and Noorul Shaiful Fitri Abdul Rahman. Maritime supply chain optimization by using fuzzy goal programming. Algorithms (Basel), 14(8), August 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/8/234.

Lancia:2021:CCI

[1483] Giuseppe Lancia and Paolo Serafini. Computational complexity and ILP models for pattern problems in the logical analysis of data. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/235.

Xu:2021:SIG

[1484] Haoran Xu, Xinya Li, Kaiyi Zhang, Yanbai He, Haoran Fan, Sijiang Liu, Chuanyan Hao, and Bo Jiang. SR-Inpaint: a general deep learning framework for high resolution image inpainting. Algorithms (Basel), 14 (8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/8/236.

Li:2021:DOC

[1485] Xu Li and Qiming Sun. Detect overlapping community based on the combination of local expansion and label propagation. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/8/237.

Ouardi:2021:ECE

[1486] Faissal Ouardi, Zineb Lotfi, and Bilal Elghadyry. Efficient construction of the equation automaton. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/8/238.

Song:2021:ASS

[1487] Zhenyu Song, Xuemei Yan, Lvxing Zhao, Luyi Fan, Cheng Tang, and Junkai Ji. Adaptive self-scaling brain-storm optimization via a chaotic search mechanism. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/8/239.

Kegenbekov:2021:ASC

[1488] Zhandos Kegenbekov and Ilya Jackson. Adaptive supply chain: Demandsupply synchronization using deep reinforcement learning. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/240.

Huang:2021:PBS

[1489] Lan Huang, Yuanwei Zhao, Bo Wang, Dongxu Zhang, Rui Zhang, Subhashis Das, Simone Bocca, and Fausto Giunchiglia. Property-based semantic similarity criteria to evaluate the overlaps of schemas. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/241.

Kousis:2021:DMA

[1490] Anestis Kousis and Christos Tjortjis. Data mining algorithms for smart cities: a bibliometric analysis. Algorithms (Basel), 14(8), August 2021.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/8/242.

Hsieh:2021:TDF

[1491] Shun-Chieh Hsieh. Tourism demand forecasting based on an LSTM network and its variants. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/8/243.

Zhang:2021:EGS

[1492] Zhanhao Zhang and Qifan Huang. An efficient geometric search algorithm of pandemic boundary detection. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/8/244.

Albeshri:2021:SHA

[1493] Aiiad Albeshri. SVSL: a human activity recognition method using softvoting and self-learning. Algorithms (Basel), 14(8), August 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/8/245.

Sotskov:2021:SMT

[1494] Yuri N. Sotskov and Evangelina I. Mihova. Scheduling multiprocessor tasks with equal processing times as a mixed graph coloring problem. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/8/246.

Wang:2021:NAD

[1495] Chenyu Wang, Hong Qiao, Yi Wang, and Xianting Du. Numerical algorithm for dynamic impedance of bridge pile-group foundation and its validation. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/8/247.

Fnadi:2021:EVG

[1496] Mohamed Fnadi and Julien Alexandre dit Sandretto. Experimental validation of a guaranteed nonlinear model predictive control. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/248.

delaRosa:2021:MIQ

[1497] Ezequiel de la Rosa, Désiré Sidibé, Thomas Decourselle, Thibault Leclercq, Alexandre Cochet, and Alain Lalande. Myocardial infarction quantification from late gadolinium enhancement MRI using tophat transforms and neural networks. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/8/249.

Ahmed:2021:RTN

[1498] Ahmed Abdelmoamen Ahmed and Gbenga Agunsoye. A real-time network traffic classifier for online applications using machine learning. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/8/250.

Dey:2021:CAR

[1499] Polash Dey, Emam Hossain, Md. Ishtiaque Hossain, Mohammed Armanuzzaman Chowdhury, Md. Shariful Alam, Mohammad Shahadat Hossain, and Karl Andersson. Comparative analysis of recurrent neural networks in stock price prediction for different frequency domains. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/8/251.

Wu:2021:CDM

[1500] Weiping Wu, Lifen Wu, Ruobing Xue, and Shan Pang. Constrained dynamic mean-variance portfolio selection in continuous-time. Algorithms (Basel), 14(8), August 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/8/252.

Toivonen:2021:PHA

[1501] Tapani Toivonen and Markku Tukiainen. The power of human-algorithm collaboration in solving combinatorial optimization problems. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/9/253.

Kazerooni:2021:PCL

[1502] Matin Kazerooni, Phuong Nguyen, and Aminah Robinson Fayek. Prioritizing construction labor productivity improvement strategies using fuzzy multi-criteria decision making and fuzzy cognitive maps. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/9/254.

Gerbet:2021:AAI

[1503] Daniel Gerbet and Klaus Röbenack. An algebraic approach to identifiability. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/9/255.

Fan:2021:SSC

[1504] Rang-Lin Fan, Pu Wang, Chen Han, Li-Jun Wei, Zi-Jian Liu, and Pei-Ju Yuan. Summarisation, simulation and comparison of nine control algorithms for an active control mount with an oscillating coil actuator. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/9/
256.

Xu:2021:MSD

[1505] Yiming Xu, Kai Zhang, and Li Wang. Metal surface defect detection using modified YOLO. Algorithms (Basel), 14(9), September 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/9/257.

Khang:2021:NSS

[1506] Tran Dinh Khang, Manh-Kien Tran, and Michael Fowler. A novel semisupervised fuzzy C-means clustering algorithm using multiple fuzzification coefficients. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/9/258.

Reinaldi:2021:STE

[1507] Marco Reinaldi, Anak Agung Ngurah Perwira Redi, Dio Fawwaz Prakoso, Arrie Wicaksono Widodo, Mochammad Rizal Wibisono, Agus Supranartha, Rahmad Inca Liperda, Reny Nadlifatin, Yogi Tri Prasetyo, and Sekar Sakti. Solving the two echelon vehicle routing problem using simulated annealing algorithm considering drop box facilities and emission cost: a case study of reverse logistics application in Indonesia. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/9/
259.

Simumba:2021:CPB

[1508] Naomi Simumba, Suguru Okami, Akira Kodaka, and Naohiko Kohtake. Comparison of profit-based multi-objective approaches for feature selection in credit scoring. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/9/260.

Dong:2021:PPF

[1509] Lin Dong, Yuanning Liu, and Xiaodong Zhu. PFSegIris: Precise and fast segmentation algorithm for multi-source heterogeneous iris. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/9/261.

Zheng:2021:PHP

[1510] Tianhua Zheng, Jiabin Wang, and Yuxiang Cai. Parallel hybrid particle swarm algorithm for workshop scheduling based on Spark. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/9/262.

Wang:2021:SRT

[1511] Shuli Wang, Xuewen Li, Xiaomeng Kou, Jin Zhang, Shaojie Zheng, Jinlong Wang, and Jibing Gong. Sequential recommendation through graph neural networks and transformer encoder with degree encoding. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/9/263.

Yang:2021:FAO

[1512] Junxia Yang, Youpeng Zhang, and Yuxiang Jin. Fully automatic operation algorithm of urban rail train based on RBFNN position output constrained robust adaptive control. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/9/264.

Wisnubhadra:2021:QVE

[1513] Irya Wisnubhadra, Safiza Kamal Baharin, Nurul A. Emran, and Djoko Budiyanto Setyohadi. QB4MobOLAP: a vocabulary extension for mobility OLAP on the semantic Web. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/9/265.

Chen:2021:MCF

[1514] Juan Chen, Qinxuan Feng, and Qi Guo. Multi-class freeway congestion and emission based on robust dynamic multi-objective optimization. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
9/266.

Mele:2021:NCH

[1515] Umberto Junior Mele, Luca Maria Gambardella, and Roberto Montemanni. A new constructive heuristic driven by machine learning for the traveling salesman problem. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/9/267.

Li:2021:URM

[1516] Huoyou Li, Jianshiun Hu, Jingwen Yu, Ning Yu, and Qingqiang Wu. UFaceNet: Research on multi-task face recognition algorithm based on CNN. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/9/268.

Angulo:2021:ABS

[1517] Andrés Angulo, Diego Rodríguez, Wilmer Garzón, Diego F. Gómez, Ameena Al Sumaiti, and Sergio Rivera. Algorithms for bidding strategies in local energy markets: Exhaustive search through parallel computing and metaheuristic optimization. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/9/269.

Mezei:2021:UCT

[1518] Mihaly Mezei. Use of the codon table to quantify the evolutionary role of random mutations. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/9/270.

Bo:2021:LTE

[1519] Hongjian Bo, Haifeng Li, Boying Wu, Hongwei Li, and Lin Ma. Longterm EEG component analysis method based on lasso regression. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/9/271.

Pavone:2021:HND

[1520] Arianna Pavone and Alessio Plebe. How neurons in deep models relate with neurons in the brain. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/9/272.

Villegas:2021:ISV

[1521] José M. Villegas, Camilo Caraveo, David A. Mejía, José L. Rodríguez, Yuridia Vega, Leticia Cervantes, and Alejandro Medina-Santiago. Intelligent search of values for a controller using the artificial bee colony algorithm to control the velocity of displacement of a robot. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/9/273.

Wang:2021:DRC

[1522] Fan Wang, Gaogao Dong, and Lixin Tian. Dynamical recovery of complex networks under a localised attack. Algorithms (Basel), 14(9), September 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/9/274.

Ahmid:2021:EHC

[1523] Ali Ahmid, Thien-My Dao, and Ngan Van Le. Enhanced hyper-cube framework ant colony optimization for combinatorial optimization problems. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/10/286.

Aroudi:2021:CLC

[1524] Ali Aroudi, Eghart Fischer, Maja Serman, Henning Puder, and Simon Doclo. Closed-loop cognitive-driven gain control of competing sounds using auditory attention decoding. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/10/287.

Blondell:2021:GME

[1525] Lucy Blondell, Mark Z. Kos, John Blangero, and Harald H. H. Göring. Genz and Mendell–Elston estimation of the high-dimensional multivariate normal distribution. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/10/296.

Es-haghi:2021:MLB

[1526] Mohammad Sadegh Es-haghi, Mohsen Abbaspour, Hamidreza Abbasianjahromi, and Stefano Mariani. Machine learning-based prediction of the seismic bearing capacity of a shallow strip footing over a void in heterogeneous soils. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/10/288.

Herrman:2021:GOQ

[1527] Rebekah Herrman, Lorna Treffert, James Ostrowski, Phillip C. Lotshaw, Travis S. Humble, and George Siopsis. Globally optimizing QAOA circuit depth for constrained optimization problems. Algorithms (Basel), 14(10): ??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/10/294.

Hinz:2021:AMR

[1528] Juri Hinz. An algorithm for making regime-changing Markov decisions. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
10/291.

Kallipolitis:2021:EEC

[1529] Athanasios Kallipolitis, Kyriakos Revelos, and Ilias Maglogiannis. Ensembling EfficientNets for the classification and interpretation of histopathology images. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/10/278.

Kravari:2021:SFS

[1530] Kalliopi Kravari, Christina Antoniou, and Nick Bassiliades. SENSE: a flow-down semantics-based requirements engineering framework. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/14/10/298.

Ma:2021:FGP

[1531] Kai Ma, Ming-Jun Nie, Sen Lin, Jianlei Kong, Cheng-Cai Yang, and Jinhao Liu. Fine-grained pests recognition based on truncated probability fusion network via Internet of Things in forestry and agricultural scenes. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/10/290.

Mei:2021:RBT

[1532] Yong Mei, Hao Chen, and Shuting Yang. Research on building target detection based on high-resolution optical remote sensing imagery. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
10/300.

Neroni:2021:ACO

[1533] Mattia Neroni. Ant colony optimization with warm-up. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/10/295.

Patil:2021:IRA

[1534] Shruti Patil, Vijayakumar Varadarajan, Devika Walimbe, Siddharth Gulechha, Sushant Shenoy, Aditya Raina, and Ketan Kotecha. Improving the robustness of AI-based malware detection using adversarial machine learning. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/10/297.

Peckens:2021:UPS

[1535] Courtney A. Peckens, Andrea Alsgaard, Camille Fogg, Mary C. Ngoma, and Clara Voskuil. Utilizing the particle swarm optimization algorithm for determining control parameters for civil structures subject to seismic excitation. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/10/292.

Penaloza:2021:BRU

[1536] Rafael Peñaloza. A brief roadmap into uncertain knowledge representation via probabilistic description logics. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/10/280.

Providas:2021:UFA

[1537] Efthimios Providas. A unified formulation of analytical and numerical methods for solving linear Fredholm integral equations. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/10/293.

Reyes:2021:AOP

[1538] Elkin D. Reyes and Sergio Rivera. Algorithms for optimal power flow extended to controllable renewable systems and loads. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/10/276.

Rodriguez:2021:RPH

[1539] Diego Rodriguez, Diego Gomez, David Alvarez, and Sergio Rivera. A review of parallel heterogeneous computing algorithms in power systems. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
10/275.

Sagar:2021:CCM

[1540] Priyadarshni Suresh Sagar, Eman Abdulah AlOmar, Mohamed Wiem Mkaouer, Ali Ouni, and Christian D. Newman. Comparing commit messages and source code metrics for the prediction refactoring activities. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
10/289.

Salvatierra:2021:SCO

[1541] Marcos M. Salvatierra, Mario Salvatierra, Jr., and Juan G. Colonna. Short communication: Optimally solving the unit-demand envy-free pricing problem with metric substitutability in cubic time. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/10/279.

Tian:2021:IFB

[1542] Zhen Tian, Lamei Pan, Pu Yin, and Rui Wang. Information fusion-based deep neural attentive matrix factorization recommendation. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/10/281.

Wang:2021:XID

[1543] Xiao Wang, Xi Lin, Rong Wang, Kai-Qi Fan, Li-Jun Han, and Zhao-Yuan Ding. XGB4mcPred: Identification of DNA N4-Methylcytosine sites in multiple species based on an eXtreme gradient boosting algorithm and DNA sequence information. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/10/283.

Wu:2021:SFS

[1544] Di Wu, Wanying Zhang, Heming Jia, and Xin Leng. Simultaneous feature selection and support vector machine optimization using an enhanced chimp optimization algorithm. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/10/282.

Xu:2021:REB

[1545] Jie Xu and Wei Ding. Rough estimator based asynchronous distributed super points detection on high speed network edge. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/10/277.

Yang:2021:EPD

[1546] Hao-Yi Yang, Zhi-Rong Lin, and Ko-Chih Wang. Efficient and portable distribution modeling for large-scale scientific data processing with dataparallel primitives. Algorithms (Basel), 14(10):??, October 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/10/285.

Zhang:2021:FBL

[1547] Chuanwei Zhang, Shirui Chen, Lu Zhao, Xianghe Li, and Xiaowen Ma. FPGA-based linear detection algorithm of an underground inspection robot. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/10/284.

Zheng:2021:SIP

[1548] Jianguo Zheng, Yilin Wang, Shihan Li, and Hancong Chen. The stock index prediction based on SVR model with bat optimization algorithm. Algorithms (Basel), 14(10):??, October 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
10/299.

Ahmed:2021:RTC

[1549] Ahmed Abdelmoamen Ahmed and Sheikh Ahmed. A real-time car towing management system using ML-powered automatic number plate recognition. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/11/317.

Akisue:2021:ODO

[1550] Rafael Akira Akisue, Matheus Lopes Harth, Antonio Carlos Luperni Horta, and Ruy de Sousa Junior. Optimized dissolved oxygen fuzzy control for recombinant Escherichia coli cultivations. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/11/326.

Atli:2021:MOU

[1551] Ibrahim Atli, Metin Ozturk, Gianluca C. Valastro, and Muhammad Zeeshan Asghar. Multi-objective UAV positioning mechanism for sustainable wireless connectivity in environments with forbidden flying zones. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
11/302.

Baranovskiy:2021:PIA

[1552] Nikolay Viktorovich Baranovskiy, Aleksey Podorovskiy, and Aleksey Malinin. Parallel implementation of the algorithm to compute forest fire impact on infrastructure facilities of JSC Russian railways. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/11/333.

Bertotti:2021:MMU

[1553] Maria Letizia Bertotti. A mathematical model of universal basic income and its numerical simulations. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/11/331.

Cariow:2021:PAD

[1554] Aleksandr Cariow and Janusz P. Paplinski. A parallel algorithm for dividing octonions. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/11/309.

Chen:2021:LIG

[1555] Yang Chen, Masao Yamagishi, and Isao Yamada. A linearly involved generalized Moreau enhancement of l_{2,1}-norm with application to weighted group sparse classification. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/11/312.

Dupin:2021:MCG

[1556] Nicolas Dupin, Rémi Parize, and El-Ghazali Talbi. Matheuristics and column generation for a basic technician routing problem. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/11/313.

Falini:2021:ARA

[1557] Antonella Falini and Maria Lucia Sampoli. Adaptive refinement in advection-diffusion problems by anomaly detection: a numerical study. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
11/328.

Farago:2021:DRS

[1558] András Faragó. Decomposition of random sequences into mixtures of simpler ones and its application in network analysis. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/11/336.

Huynh-Cam:2021:UDT

[1559] Thao-Trang Huynh-Cam, Long-Sheng Chen, and Huynh Le. Using decision trees and random forest algorithms to predict and determine factors contributing to first-year university students’ learning performance. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
11/318.

Ingabire:2021:ONL

[1560] Winfred Ingabire, Hadi Larijani, Ryan M. Gibson, and Ayyaz-Ul-Haq Qureshi. Outdoor node localization using random neural networks for large-scale urban IoT LoRa networks. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/11/307.

Jeong:2021:MFS

[1561] BongJoo Jeong, Jun-Hee Han, and Ju-Yong Lee. Metaheuristics for a flow shop scheduling problem with urgent jobs and limited waiting times. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
11/323.

Junior:2021:MMT

[1562] Bonfim Amaro Junior, Marcio Costa Santos, Guilherme Nepomuceno de Carvalho, Luiz Jonatã Pires de Araújo, and Placido Rogerio Pinheiro. Metaheuristics for the minimum time cut path problem with different cutting and sliding speeds. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/11/305.

Koschel:2021:SKF

[1563] Alan Koschel, Christoph Müller, and Alexander Reiterer. Selection of key frames for 3D reconstruction in real time. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/11/303.

Koutroumpina:2021:EFG

[1564] Christina Koutroumpina, Spyros Sioutas, Stelios Koutroubinas, and Kostas Tsichlas. Evaluation of features generated by a high-end lowcost electrical smart meter. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/11/311.

Lai:2021:ESC

[1565] Daphne Teck Ching Lai and Yuji Sato. An empirical study of clusterbased MOEA/D bare bones PSO for data clustering. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/11/338.

Landro:2021:OTM

[1566] Nicola Landro, Ignazio Gallo, and Riccardo La Grassa. Is one teacher model enough to transfer knowledge to a student model? Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/11/334.

Li:2021:VTF

[1567] Zhiwei Li, Jun Li, Yousheng Xia, Pingfa Feng, and Feng Feng. Variation trends of fractal dimension in epileptic EEG signals. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/11/316.

Liang:2021:NDG

[1568] Yan Liang and Qingdong Zhang. A non-dominated genetic algorithm based on decoding rule of heat treatment equipment volume and job delivery date. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/11/310.

Liu:2021:TTR

[1569] Lanfen Liu and Xinfeng Yang. Travel time reliability-based rescue resource scheduling for accidents concerning transport of dangerous goods by rail. Algorithms (Basel), 14(11):??, November 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/11/325.

Lo:2021:IBC

[1570] Shaw-Hwa Lo and Yiqiao Yin. An interaction-based convolutional neural network (ICNN) toward a better understanding of COVID-19 X-ray images. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
14/11/337.

Lu:2021:RBP

[1571] Yaohang Lu and Zhongming Teng. Robust bilinear probabilistic principal component analysis. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/11/322.

McKendall:2021:AUA

[1572] Alan McKendall and Artak Hakobyan. An application of an unequal-area facilities layout problem with fixed-shape facilities. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/11/306.

Michelucci:2021:MAA

[1573] Umberto Michelucci, Michela Sperti, Dario Piga, Francesca Venturini, and Marco A. Deriu. A model-agnostic algorithm for Bayes error determination in binary classification. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/11/301.

Rendon-Cardona:2021:AFF

[1574] Cristian Rendon-Cardona, Jorge Correa, Diego A. Acosta, and Oscar Ruiz-Salguero. Analytic form fitting in poor triangular meshes. Algorithms (Basel), 14(11):??, November 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
11/304.

Al-Azzoni:2021:MDA

[1575] Issam Al-Azzoni, Julian Blank, and Nenad Petrović. A model-driven approach for solving the software component allocation problem. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
12/354.

Belhaiza:2021:BBA

[1576] Slim Belhaiza. A branch-and-bound algorithm for polymatrix games ϵ-proper Nash equilibria computation. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/12/365.

Berry:2021:CAG

[1577] Anne Berry and Geneviève Simonet. Computing the atom graph of a graph and the union join graph of a hypergraph. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/12/347.

Bologna:2021:RET

[1578] Guido Bologna. A rule extraction technique applied to ensembles of neural networks, random forests, and gradient-boosted trees. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/12/339.

Bonet:2021:SDM

[1579] Clément Bonet, Titouan Vayer, Nicolas Courty, François Septier, and Lucas Drumetz. Subspace detours meet Gromov–Wasserstein. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/12/366.

Burr:2021:OAU

[1580] Tom Burr, Ian Schwerdt, Kari Sentz, Luther McDonald, and Marianne Wilkerson. Overview of algorithms for using particle morphology in predetonation nuclear forensics. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/12/340.

Castelo:2021:VMA

[1581] Sonia Castelo, Moacir Ponti, and Rosane Minghim. A visual mining approach to improved multiple-instance learning. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/12/344.

Damasevicius:2021:ASF

[1582] Robertas Damasevicius and Rytis Maskeliunas. Agent state flipping based hybridization of heuristic optimization algorithms: a case of bat algorithm and krill herd hybrid algorithm. Algorithms (Basel), 14(12): ??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/14/12/358.

Farago:2021:MPP

[1583] András Faragó. A meeting point of probability, graphs, and algorithms: The Lovász local lemma and related results — a survey. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/12/355.

He:2021:HAA

[1584] Zhenwen He, Chunfeng Zhang, Xiaogang Ma, and Gang Liu. Hexadecimal aggregate approximation representation and classification of time series data. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/12/353.

Hiraishi:2021:PPF

[1585] Kunihiko Hiraishi. A pathfinding problem for fork-join directed acyclic graphs with unknown edge length. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/14/12/367.

Huang:2021:RAI

[1586] Wei Huang, Zhiren Han, Li Zhao, Hongbo Xu, Zhongnian Li, and Ze Wang. Resource allocation for intelligent reflecting surfaces assisted federated learning system with imperfect CSI. Algorithms (Basel), 14 (12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/12/363.

Kosolobov:2021:LZP

[1587] Dmitry Kosolobov and Daniel Valenzuela. Lempel–Ziv parsing for sequences of blocks. Algorithms (Basel), 14(12):??, December 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/12/359.

Lavrentiev:2021:ADF

[1588] Mikhail Lavrentiev, Konstantin Lysakov, Andrey Marchuk, Konstantin Oblaukhov, and Mikhail Shadrin. Algorithmic design of an FPGA-based calculator for fast evaluation of tsunami wave danger. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/12/343.

Lenseth:2021:MDM

[1589] Douglas Lenseth and Boris Goldfarb. Merging discrete Morse vector fields: a case of stubborn geometric parallelization. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/12/360.

Li:2021:GDD

[1590] Xutong Li, Taoying Li, and Yan Wang. GW-DC: a deep clustering model leveraging two-dimensional image transformation and enhancement. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
12/349.

Mongwe:2021:LSS

[1591] Wilson Tsakane Mongwe, Rendani Mbuvha, and Tshilidzi Marwala. Locally scaled and stochastic volatility Metropolis–Hastings algorithms. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
12/351.

Mukhopadhyay:2021:FPS

[1592] Priyanka Mukhopadhyay. Faster provable sieving algorithms for the shortest vector problem and the closest vector problem on lattices in l_{p} norm. Algorithms (Basel), 14(12):??, December 2021. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/12/362.

Pietri:2021:BOP

[1593] Nicola Ognibene Pietri, Xiaochen Chou, Dominic Loske, Matthias Klumpp, and Roberto Montemanni. The buy-online-pick-up-in-store retailing model: Optimization strategies for in-store picking and packing. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/14/
12/350.

Providas:2021:PFS

[1594] Efthimios Providas and Ioannis Nestorios Parasidis. A procedure for factoring and solving nonlocal boundary value problems for a type of linear integro-differential equations. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/12/346.

Regueiro:2021:BBA

[1595] Cristina Regueiro, Iñaki Seco, Iván Gutiérrez-Agüero, Borja Urquizu, and Jason Mansell. A blockchain-based audit trail mechanism: Design and implementation. Algorithms (Basel), 14(12):??, December 2021. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/14/12/341.

Robles:2021:MHS

[1596] Francisca Santana Robles, Eva Selene Hernández-Gress, Neil Hernández-Gress, and Rafael Granillo Macias. Metaheuristics in the humanitarian supply chain. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/12/364.

Szeker:2021:OWN

[1597] Szabolcs Szekér and Ágnes Vathy-Fogarassy. Optimized weighted nearest neighbours matching algorithm for control group selection. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/12/356.

Tayebi:2021:RRE

[1598] Zahra Tayebi, Sarwan Ali, and Murray Patterson. Robust representation and efficient feature selection allows for effective clustering of SARS-CoV-2 variants. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/12/348.

Varsi:2021:LFB

[1599] Alessandro Varsi, Simon Maskell, and Paul G. Spirakis. An O(log_{2} N) fully-balanced resampling algorithm for particle filters on distributed memory architectures. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/14/12/342.

vonKurnatowski:2021:CDS

[1600] Martin von Kurnatowski, Jochen Schmid, Patrick Link, Rebekka Zache, Lukas Morand, Torsten Kraft, Ingo Schmidt, Jan Schwientek, and Anke Stoll. Compensating data shortages in manufacturing with monotonicity knowledge. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/12/345.

Werner:2021:SIS

[1601] Frank Werner. Special issue “2021 Selected Papers from Algorithms’ Editorial Board Members”. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/14/12/357.

Zhang:2021:ALA

[1602] Yajing Zhang, Kai Wang, and Jinghui Zhang. Adaptive and lightweight abnormal node detection via biological immune game in mobile multimedia networks. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/14/12/368.

Zhao:2021:SGN

[1603] Ke Zhao, Lan Huang, Rui Song, Qiang Shen, and Hao Xu. A sequential graph neural network for short text classification. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/14/12/352.

Zhong:2021:DAP

[1604] Chengyan Zhong, Guanqiu Qi, Neal Mazur, Sarbani Banerjee, Devanshi Malaviya, and Gang Hu. A domain adaptive person re-identification based on dual attention mechanism and camstyle transfer. Algorithms (Basel), 14(12):??, December 2021. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/14/12/361.

Pinto:2022:TBI

[1605] Carlos Pinto, Rui Pinto, and Gil Gonçalves. Towards bio-inspired anomaly detection using the cursory dendritic cell algorithm. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/1/1.

Grubisic:2022:PQD

[1606] Luka Grubisić, Domagoj Lacmanović, and Josip Tambaca. Preconditioning the quad dominant mesh generator for ship structural analysis. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/1/2.

Ge:2022:EKM

[1607] Yu Ge, Junjun Shi, Yaohui Li, and Jingfang Shen. An efficient kriging modeling method based on multidimensional scaling for high-dimensional problems. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/1/3.

Ssenyonga:2022:NAS

[1608] Taddeo Ssenyonga, Øyvind Frette, Børge Hamre, Knut Stamnes, Dennis Muyimbwa, Nicolausi Ssebiyonga, and Jakob J. Stamnes. A new algorithm for simultaneous retrieval of aerosols and marine parameters. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/1/4.

Arasteh:2022:DTL

[1609] Emad Arasteh, Ailar Mahdizadeh, Maryam S. Mirian, Soojin Lee, and Martin J. McKeown. Deep transfer learning for Parkinson’s disease monitoring by image-based representation of resting-state EEG using directional connectivity. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/1/5.

Indrapriyadarsini:2022:ASR

[1610] S. Indrapriyadarsini, Shahrzad Mahboubi, Hiroshi Ninomiya, Takeshi Kamio, and Hideki Asai. Accelerating symmetric rank-1 quasi-Newton method with Nesterov’s gradient for training neural networks. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/1/6.

Gorniak:2022:MBP

[1611] Piotr Górniak. Modeling of the 5G-band patch antennas using ANNs under the uncertainty of the geometrical design parameters associated with the manufacturing process. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/1/7.

Ji:2022:FTC

[1612] Xiaofu Ji and Xuehua Liu. Finite-time control of singular linear semi-Markov jump systems. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/1/8.

Muller:2022:HHB

[1613] Felipe Martins Müller and Iaê Santos Bonilha. Hyper-heuristic based on ACO and local search for dynamic optimization problems. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/1/9.

Cappelle:2022:SMA

[1614] Márcia R. Cappelle, Les R. Foulds, and Humberto J. Longo. Searching monotone arrays: a survey. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/1/10.

Keddous:2022:OCH

[1615] Fekhr Eddine Keddous and Amir Nakib. Optimal CNN-Hopfield network for pattern recognition based on a genetic algorithm. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/1/11.

Roudsari:2022:OFR

[1616] Sajjad Sayyar Roudsari, Liviu Marian Ungureanu, Soheil Soroushnia, Taher Abu-Lebdeh, and Florian Ion Tiberiu Petrescu. Optimization of fiber-reinforced polymer bars for reinforced concrete column using nonlinear finite element algorithms. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/1/12.

Padalko:2022:PCE

[1617] Mikhail Alexandrovich Padalko, Yuriy Andreevich Shevchenko, Vitalii Yurievich Kapitan, and Konstantin Valentinovich Nefedev. Parallel computing of Edwards–Anderson model. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/1/13.

Ozlu:2022:HMB

[1618] Ismail Alperen Özlü, Olzhas Baimakhanov, Almaz Saukhimov, and Oguzhan Ceylan. A heuristic methods-based power distribution system optimization toolbox. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/1/14.

Grigoriev:2022:PIR

[1619] Vasiliy V. Grigoriev, Oleg Iliev, and Petr N. Vabishchevich. On parameter identification for reaction-dominated pore-scale reactive transport using modified bee colony algorithm. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/1/15.

Tzougas:2022:MPR

[1620] George Tzougas, Natalia Hong, and Ryan Ho. Mixed Poisson regression models with varying dispersion arising from non-conjugate mixing distributions. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/1/16.

Han:2022:ACF

[1621] Liang Han, Feng Liu, and Kaifeng Chen. Analog circuit fault diagnosis using a novel variant of a convolutional neural network. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/1/17.

Mukhamedov:2022:EDP

[1622] Farrukh Mukhamedov. Extremality of disordered phase of λ-model on Cayley trees. Algorithms (Basel), 15(1), January 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/1/18.

Jin:2022:POA

[1623] Qibing Jin and Yuming Zhang. Parameter optimization of active disturbance rejection controller using adaptive differential ant-lion optimizer. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/1/19.

Chen:2022:MNI

[1624] Yinan Chen, Chuanpeng Wang, and Dong Li. MINC-NRL: an information-based approach for community detection. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/1/20.

Gakii:2022:GBF

[1625] Consolata Gakii, Paul O. Mireji, and Richard Rimiru. Graph based feature selection for reduction of dimensionality in next-generation RNA sequencing datasets. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/1/21.

Niculescu:2022:TBP

[1626] Virginia Niculescu and Robert Manuel Stefanica. Tries-based parallel solutions for generating perfect crosswords grids. Algorithms (Basel), 15 (1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/1/22.

Zhang:2022:RPB

[1627] Yang Zhang, Jiacheng Li, and Lei Li. A reward population-based differential genetic harmony search algorithm. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/1/23.

Durgut:2022:TLO

[1628] Rafet Durgut, Mehmet Emin Aydin, and Abdur Rakib. Transfer learning for operator selection: a reinforcement learning approach. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/1/24.

Li:2022:KDB

[1629] Wen Li, Junfei Xu, and Qi Chen. Knowledge distillation-based multilingual code retrieval. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/1/25.

Taketomi:2022:MSR

[1630] Nanami Taketomi, Hirofumi Michimae, Yuan-Tsung Chang, and Takeshi Emura. meta.shrinkage: an R package for meta-analyses for simultaneously estimating individual means. Algorithms (Basel), 15(1), January 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/1/26.

Hoffmann:2022:UGE

[1631] Maximilian Hoffmann and Ralph Bergmann. Using graph embedding techniques in process-oriented case-based reasoning. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/27.

Jiang:2022:RDD

[1632] Nianqi Jiang, Fenhua Bai, Lin Huang, Zhengyuan An, and Tao Shen. Reputation-driven dynamic node consensus and reliability sharding model in IoT blockchain. Algorithms (Basel), 15(2), February 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/2/28.

Colagrossi:2022:SAD

[1633] Andrea Colagrossi and Michèle Lavagna. A spacecraft attitude determination and control algorithm for solar arrays pointing leveraging Sun angle and angular rates measurements. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/2/29.

Luo:2022:MTT

[1634] Kelin Luo and Frits C. R. Spieksma. Minimizing travel time and latency in multi-capacity ride-sharing problems. Algorithms (Basel), 15 (2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/30.

Huang:2022:ADR

[1635] Shih-Chang Huang and Cong-Han Huang. Algorithms for detecting and refining the area of intangible continuous objects for mobile wireless sensor networks. Algorithms (Basel), 15(2), February 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/31.

Dai:2022:EAA

[1636] Jiazhu Dai and Siwei Xiong. An evasion attack against stacked capsule autoencoder. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/32.

Pakhomchik:2022:CBE

[1637] Aleksey I. Pakhomchik, Vladimir V. Voloshinov, Valerii M. Vinokur, and Gordey B. Lesovik. Converting of Boolean expression to linear equations, inequalities and QUBO penalties for cryptanalysis. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/33.

Zeng:2022:CNN

[1638] Fancheng Zeng, Guanqiu Qi, Zhiqin Zhu, Jian Sun, Gang Hu, and Matthew Haner. Convex neural networks based reinforcement learning for load frequency control under denial of service attacks. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/2/34.

Chen:2022:ABP

[1639] Yidong Chen, Chen Li, and Zhonghua Lu. An ADMM based parallel approach for fund of fund construction. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/2/35.

Kruchinin:2022:USC

[1640] Vladimir Kruchinin, Yuriy Shablya, Dmitry Kruchinin, and Victor Rulevskiy. Unranking small combinations of a large set in colexicographic order. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/36.

Aguirre-Angulo:2022:OID

[1641] Brayan Enrique Aguirre-Angulo, Lady Carolina Giraldo-Bello, Oscar Danilo Montoya, and Francisco David Moya. Optimal integration of dispersed generation in medium-voltage distribution networks for voltage stability enhancement. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/2/37.

Aranda:2022:AES

[1642] Ferley Castro Aranda, Rodolfo García Sierra, Andrés Felipe Cerón Piamba, Benjamin Mailhé, and Luis Miguel León Gil. An algorithm for estimation of SF_{6} leakage on power substation assets. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/38.

Zhang:2022:SRC

[1643] Meng Zhang and Bin Yang. Swarm robots cooperative and persistent distribution modeling and optimization based on the smart community logistics service framework. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/2/39.

Seiringer:2022:ASM

[1644] Wolfgang Seiringer, Juliana Castaneda, Klaus Altendorfer, Javier Panadero, and Angel A. Juan. Applying simheuristics to minimize overall costs of an MRP planned production system. Algorithms (Basel), 15 (2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/40.

Leone:2022:RLD

[1645] Pierre Leone and Nathan Cohen. Rendezvous on the line with different speeds and markers that can be dropped at chosen time. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/2/41.

Bannach:2022:RAP

[1646] Max Bannach and Sebastian Berndt. Recent advances in positiveinstance driven graph searching. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/2/42.

Trzcinski:2022:CNI

[1647] Maciej Trzciński, Piotr A. Kowalski, and Szymon Lukasik. Clustering with nature-inspired algorithm based on territorial behavior of predatory animals. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/2/43.

Office:2022:ARA

[1648] Algorithms Editorial Office. Acknowledgment to reviewers of Algorithms in 2021. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/2/44.

Johnson:2022:CAG

[1649] Kara Layne Johnson and Nicole Bohme Carnegie. Calibration of an adaptive genetic algorithm for modeling opinion diffusion. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/2/45.

Werner:2022:TVS

[1650] Yannis Werner, Tim van Hout, Vijey Subramani Raja Gopalan, and Thomas Vietor. Test and validation of the surrogate-based, multiobjective GOMORS algorithm against the NSGA-II algorithm in structural shape optimization. Algorithms (Basel), 15(2), February 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/2/46.

Ottenburger:2022:NMB

[1651] Sadeeb Simon Ottenburger, Stella Möhrle, Tim Oliver Müller, and Wolfgang Raskob. A novel MCDA-based methodology dealing with dynamics and ambiguities resulting from citizen participation in the context of the energy transition. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/47.

Ibanez:2022:TTA

[1652] Javier Ibáñez, José M. Alonso, Pedro Alonso-Jordá, Emilio Defez, and Jorge Sastre. Two Taylor algorithms for computing the action of the matrix exponential on a vector. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/2/48.

Janssen:2022:UEM

[1653] Femke M. Janssen, Katja K. H. Aben, Berdine L. Heesterman, Quirinus J. M. Voorham, Paul A. Seegers, and Arturo Moncada-Torres. Using explainable machine learning to explore the impact of synoptic reporting on prostate cancer. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/49.

Chistousov:2022:AAP

[1654] Nikita Konstantinovich Chistousov, Igor Anatolyevich Kalmykov, Daniil Vyacheslavovich Dukhovnyj, Maksim Igorevich Kalmykov, and Aleksandr Anatolyevich Olenev. Adaptive authentication protocol based on zero-knowledge proof. Algorithms (Basel), 15(2), February 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/2/50.

Pham:2022:NCL

[1655] Huy Pham, Emile R. Shehada, Shawna Stahlheber, Kushagra Pandey, and Wayne B. Hayes. No cell left behind: Automated, stochastic, physicsbased tracking of every cell in a dense, growing colony. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/2/51.

Pellizzoni:2022:CCO

[1656] Paolo Pellizzoni, Andrea Pietracaprina, and Geppino Pucci. k-center clustering with outliers in sliding windows. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/2/52.

Ma:2022:PSR

[1657] Yaoyao Ma, Xiaoyu Xu, Shuai Yan, and Zhuoxiang Ren. A preliminary study on the resolution of electro-thermal multi-physics coupling problem using physics-informed neural network (PINN). Algorithms (Basel), 15 (2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/53.

Laroque:2022:BRD

[1658] Christoph Laroque, Madlene Leißau, Pedro Copado, Christin Schumacher, Javier Panadero, and Angel A. Juan. A biased-randomized discrete-event algorithm for the hybrid flow shop problem with time dependencies and priority constraints. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/2/54.

Soman:2022:PBW

[1659] Gayathri Soman, M. V. Vivek, M. V. Judy, Elpiniki Papageorgiou, and Vassilis C. Gerogiannis. Precision-based weighted blending distributed ensemble model for emotion classification. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/2/55.

Wang:2022:ISA

[1660] Lei Wang, Chenguang Wang, and Huabing Wang. Improved scheduling algorithm for synchronous data flow graphs on a homogeneous multi-core systems. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/2/56.

Hoxha:2022:DFR

[1661] Indrit Hoxha and Taoufik Meklachi. Data fitting with rational functions: Scaled null space method with applications of fitting large scale shocks on economic variables and s-parameters. Algorithms (Basel), 15 (2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/57.

Abbasi-Asl:2022:RRM

[1662] Reza Abbasi-Asl, Aboozar Ghaffari, and Emad Fatemizadeh. Robust registration of medical images in the presence of spatially-varying noise. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
2/58.

Kostyukova:2022:RAL

[1663] Olga Kostyukova and Tatiana Tchemisova. Regularization algorithms for linear copositive programming problems: an approach based on the concept of immobile indices. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/2/59.

Jami:2022:AST

[1664] Neil Jami, Neele Leithäuser, and Christian Weiß. Allocating small transporters to large jobs. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/60.

Keshavarz-Kohjerdi:2022:FHL

[1665] Fatemeh Keshavarz-Kohjerdi and Ruo-Wei Hung. Finding Hamiltonian and longest (s, t)-paths of C-shaped supergrid graphs in linear time. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/2/61.

Andronikos:2022:TPQ

[1666] Theodore Andronikos and Michail Stefanidakis. A two-party quantum parliament. Algorithms (Basel), 15(2), February 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/62.

Wu:2022:PAL

[1667] Jiarun Wu and Qingliang Chen. Pruning adapters with lottery ticket. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
2/63.

Sharma:2022:AIT

[1668] Shalini Sharma and Jerry Chou. Accelerate incremental TSP algorithms on time evolving graphs with partitioning methods. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/64.

Garvie:2022:NAB

[1669] Marcus R. Garvie and John Burkardt. A new algorithm based on colouring arguments for identifying impossible polyomino tiling problems. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/2/65.

Colloc:2022:TCB

[1670] Joël Colloc, Relwendé Aristide Yameogo, Peter Summons, Lilian Loubet, Jean-Bernard Cavelier, and Paul Bridier. A temporal case-based reasoning platform relying on a fuzzy vector spaces object-oriented model and a method to design knowledge bases and decision support systems in multiple domains. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/66.

Valiente:2022:AME

[1671] Gabriel Valiente. Adjacency maps and efficient graph algorithms. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/2/67.

Zhu:2022:WSC

[1672] Qiang Zhu, Zhong Wang, Yunfeng Dou, and Jian Zhou. Whispered speech conversion based on the inversion of mel frequency cepstral coefficient features. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/2/68.

Pitolli:2022:ARC

[1673] Francesca Pitolli, Chiara Sorgentone, and Enza Pellegrino. Approximation of the Riesz–Caputo derivative by cubic splines. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/2/69.

Kengegowda:2022:CMT

[1674] Dhanalakshmi Bettahalli Kengegowda, Srikantaiah Kamidoddi Chowdaiah, Gururaj Harinahalli Lokesh, and Francesco Flammini. Classification and merging techniques to reduce brokerage using multi-objective optimization. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/2/70.

Abdel-Jaber:2022:RDL

[1675] Hussein Abdel-Jaber, Disha Devassy, Azhar Al Salam, Lamya Hidaytallah, and Malak El-Amir. A review of deep learning algorithms and their applications in healthcare. Algorithms (Basel), 15(2), February 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/2/71.

Viktoratos:2022:MLA

[1676] Iosif Viktoratos and Athanasios Tsadiras. A machine learning approach for solving the frozen user cold-start problem in personalized mobile advertising systems. Algorithms (Basel), 15(3), March 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/3/72.

Agarwal:2022:RLM

[1677] Mridul Agarwal, Vaneet Aggarwal, Arnob Ghosh, and Nilay Tiwari. Reinforcement learning for mean-field game. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/3/73.

Bindi:2022:MLB

[1678] Marco Bindi, Fabio Corti, Igor Aizenberg, Francesco Grasso, Gabriele Maria Lozito, Antonio Luchetta, Maria Cristina Piccirilli, and Alberto Reatti. Machine learning-based monitoring of DC–DC converters in photovoltaic applications. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/3/74.

Waldamichael:2022:MLC

[1679] Fraol Gelana Waldamichael, Taye Girma Debelee, Friedhelm Schwenker, Yehualashet Megersa Ayano, and Samuel Rahimeto Kebede. Machine learning in cereal crops disease detection: a review. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/3/75.

AlAlam:2022:PTN

[1680] Pamela Al Alam, Joseph Constantin, Ibtissam Constantin, and Clelia Lopez. Partitioning of transportation networks by efficient evolutionary clustering and density peaks. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/3/76.

Moustakidis:2022:PIC

[1681] Serafeim Moustakidis, Athanasios Siouras, Konstantinos Vassis, Ioannis Misiris, Elpiniki Papageorgiou, and Dimitrios Tsaopoulos. Prediction of injuries in CrossFit training: a machine learning perspective. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/3/77.

Lartigue:2022:DAE

[1682] Thomas Lartigue, Stanley Durrleman, and Stéphanie Allassonnière. Deterministic approximate EM algorithm; application to the Riemann approximation EM and the tempered EM. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/3/78.

Nguyen:2022:EAF

[1683] Dang-Viet-Anh Nguyen, Jérôme Szewczyk, and Kanty Rabenorosoa. An effective algorithm for finding shortest paths in tubular spaces. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/3/79.

Liu:2022:PDU

[1684] Yi Liu, Chengyu Yin, Jingwei Li, Fang Wang, and Senzhang Wang. Predicting dynamic user-item interaction with meta-path guided recursive RNN. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/3/
80.

Parisi:2022:LTV

[1685] Simone Parisi, Davide Tateo, Maximilian Hensel, Carlo D’Eramo, Jan Peters, and Joni Pajarinen. Long-term visitation value for deep exploration in sparse-reward reinforcement learning. Algorithms (Basel), 15 (3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/3/81.

Flores:2022:NIS

[1686] Nellyzeth Flores, Marco A. Reyna, Roberto L. Avitia, Jose Antonio Cardenas-Haro, and Conrado Garcia-Gonzalez. Non-invasive systems and methods patents review based on electrocardiogram for diagnosis of cardiovascular diseases. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/3/82.

Hu:2022:DIP

[1687] Haijian Hu, Yicen Liu, and Haina Rong. Detection of insulators on power transmission line based on an improved faster region-convolutional neural network. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/3/83.

Wang:2022:EFD

[1688] Yuqi Wang, Lijun Zhang, and Zhen Fang. Eye fatigue detection through machine learning based on single channel electrooculography. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/3/84.

Munavalli:2022:DLD

[1689] Jyoti R. Munavalli, Shyam Vasudeva Rao, Aravind Srinivasan, and Frits Van Merode. Dynamic layout design optimization to improve patient flow in outpatient clinics using genetic algorithms. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/3/85.

Zeng:2022:PID

[1690] Xingming Zeng, Haiyuan Liu, and Hao He. Prediction of intrinsically disordered proteins using machine learning based on low complexity methods. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/3/
86.

Sagadevan:2022:SGL

[1691] Saravanan Sagadevan, Nurul Hashimah Ahamed Hassain Malim, and Mohd Heikal Husin. A seed-guided latent Dirichlet allocation approach to predict the personality of online users using the PEN model. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/3/87.

Hou:2022:RAM

[1692] Jialin Hou, Jingtao Zhang, Wanying Wu, Tianguo Jin, and Kai Zhou. Research on agricultural machinery rental optimization based on the dynamic artificial bee-ant colony algorithm. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/3/88.

Zhu:2022:CLM

[1693] Feng Zhu, Jieyu Zhao, and Zhengyi Cai. A contrastive learning method for the visual representation of 3D point clouds. Algorithms (Basel), 15 (3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/3/89.

LeMezo:2022:KAC

[1694] Thomas Le Mézo, Luc Jaulin, Damien Massé, and Benoit Zerr. Kleene algebra to compute invariant sets of dynamical systems. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/3/90.

Portugal:2022:AEG

[1695] Ernesto Portugal, Francisco Cruz, Angel Ayala, and Bruno Fernandes. Analysis of explainable goal-driven reinforcement learning in a continuous simulated environment. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/3/91.

Jensen:2022:MED

[1696] Mathias Højgaard Jensen and Stefan Sommer. Mean estimation on the diagonal of product manifolds. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/3/92.

Gollapalli:2022:EML

[1697] Mohammed Gollapalli. Ensemble machine learning model to predict the waterborne syndrome. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/3/93.

Wang:2022:MFP

[1698] Kang Wang, Zhi-Jiang Xu, Yi Gong, and Ke-Lin Du. Mechanical fault prognosis through spectral analysis of vibration signals. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/3/94.

Boechel:2022:PHT

[1699] Tiago Boechel, Lucas Micol Policarpo, Gabriel de Oliveira Ramos, Rodrigo da Rosa Righi, and Dhananjay Singh. Prediction of harvest time of apple trees: an RNN-based approach. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/3/95.

Sioutas:2022:DDD

[1700] Spyros Sioutas, Efrosini Sourla, Kostas Tsichlas, Gerasimos Vonitsanos, and Christos Zaroliagis. A dynamic distributed deterministic load-balancer for decentralized hierarchical infrastructures. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/3/96.

Sarica:2022:ESI

[1701] Alessia Sarica. Editorial for the special issue on “Machine Learning in Healthcare and Biomedical Application”. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/3/97.

Custode:2022:EOS

[1702] Leonardo Lucio Custode, Hyunho Mo, Andrea Ferigo, and Giovanni Iacca. Evolutionary optimization of spiking neural P systems for remaining useful life prediction. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/3/98.

Holoch:2022:DSL

[1703] Jan Holoch, Sven Lenhardt, Sven Revfi, and Albert Albers. Design of selective laser melting (SLM) structures: Consideration of different material properties in multiple surface layers resulting from the manufacturing in a topology optimization. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/3/99.

Tapia-Fernandez:2022:KCW

[1704] Santiago Tapia-Fernández, Daniel García-García, and Pablo García-Hernandez. Key concepts, weakness and benchmark on hash table data structures. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/3/100.

Rumpfkeil:2022:MFS

[1705] Markus P. Rumpfkeil, Dean Bryson, and Phil Beran. Multi-fidelity sparse polynomial chaos and kriging surrogate models applied to analytical benchmark problems. Algorithms (Basel), 15(3), March 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/3/101.

Verspeek:2022:DLS

[1706] Simon Verspeek, Ivan De Boi, Xavier Maldague, Rudi Penne, and Gunther Steenackers. Dynamic line scan thermography parameter design via Gaussian process emulation. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/4/102.

Torres:2022:RPL

[1707] Juan J. Torres, Can Li, Robert M. Apap, and Ignacio E. Grossmann. A review on the performance of linear and mixed integer two-stage stochastic programming software. Algorithms (Basel), 15(4), April 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/4/103.

Novikova:2022:FLI

[1708] Evgenia Novikova, Elena Doynikova, and Sergey Golubev. Federated learning for intrusion detection in the critical infrastructures: Vertically partitioned data use case. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/4/104.

Bancioiu:2022:AMB

[1709] Camil Bancioiu and Remus Brad. Analyzing Markov boundary discovery algorithms in ideal conditions using the d-separation criterion. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/105.

Serrano:2022:FMC

[1710] Daniel Ricardo Sandoval Serrano, Juan Carlos Rincón, Julián Mejía-Restrepo, Edward Rolando Núñez-Valdez, and Vicente García-Díaz. Forecast of medical costs in health companies using models based on advanced analytics. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/4/106.

Tang:2022:KCB

[1711] Deyou Tang, Daqiang Tan, Weihao Xiao, Jiabin Lin, and Juan Fu. KMC3 and CHTKC: Best scenarios, deficiencies, and challenges in highthroughput sequencing data analysis. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/107.

Zeng:2022:TNN

[1712] Yan Zeng, Jiyang Wu, Jilin Zhang, Yongjian Ren, and Yunquan Zhang. Trinity: Neural network adaptive distributed parallel training method based on reinforcement learning. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/4/108.

Zhang:2022:SLA

[1713] Wanyi Zhang, Mattia Zeni, Andrea Passerini, and Fausto Giunchiglia. Skeptical learning — an algorithm and a platform for dealing with mislabeling in personal context recognition. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/109.

Druzhinin:2022:NSM

[1714] Oleg A. Druzhinin and Wu-Ting Tsai. Numerical simulation of microbubbles dispersion by surface waves. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/110.

Rusu:2022:VSS

[1715] Alexandru-George Rusu, Constantin Paleologu, Jacob Benesty, and Silviu Ciochina. A variable step size normalized least-mean-square algorithm based on data reuse. Algorithms (Basel), 15(4), April 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/4/111.

Pizzolante:2022:EPS

[1716] Raffaele Pizzolante. Editorial paper for the special issue “Algorithms in Hyperspectral Data Analysis”. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/4/112.

He:2022:LTI

[1717] Junjun He and Huahui Chen. An LSM-tree index for spatial data. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/113.

Zuckerman:2022:EPC

[1718] Inon Zuckerman, Dor Mizrahi, and Ilan Laufer. EEG pattern classification of picking and coordination using anonymous random walks. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/114.

Sela:2022:DLS

[1719] Elihu Sela, Shan Huang, and David Horn. Deep learning study of an electromagnetic calorimeter. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/4/115.

Huang:2022:MLF

[1720] Yucheng Huang, Rui Song, Fausto Giunchiglia, and Hao Xu. A multitask learning framework for abuse detection and emotion classification. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/116.

Nigro:2022:PPK

[1721] Libero Nigro. Performance of parallel K-means algorithms in Java. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/117.

Lat:2022:BIR

[1722] Reihan Alinia Lat, Sebelan Danishvar, Hamed Heravi, and Morad Danishvar. Boosting iris recognition by margin-based loss functions. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/118.

Liang:2022:FID

[1723] Yi Liang, Turdi Tohti, and Askar Hamdulla. False information detection via multimodal feature fusion and multi-classifier hybrid prediction. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/119.

Pei:2022:MLF

[1724] Shengyu Pei and Xiaoping Fan. Multi-level fusion model for person reidentification by attribute awareness. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/120.

Zeile:2022:CIA

[1725] Clemens Zeile, Tobias Weber, and Sebastian Sager. Combinatorial integral approximation decompositions for mixed-integer optimal control. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/4/121.

Stanovov:2022:NPA

[1726] Vladimir Stanovov, Shakhnaz Akhmedova, and Eugene Semenkin. Neuroevolution for parameter adaptation in differential evolution. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/4/122.

Szabo:2022:CCF

[1727] Peter Szabó, Miroslava Ferencová, and Vladimír Zelezník. Cloud computing in free route airspace research. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/123.

Zhang:2022:PCU

[1728] Yan Zhang, Wenhan Zhao, Bo Sun, Ying Zhang, and Wen Wen. Point cloud upsampling algorithm: a systematic review. Algorithms (Basel), 15 (4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/4/124.

Truden:2022:CAG

[1729] Christian Truden, Kerstin Maier, Anna Jellen, and Philipp Hungerländer. Computational approaches for grocery home delivery services. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/4/125.

Finkbeiner:2022:TRS

[1730] Bernd Finkbeiner, Martin Fränzle, Florian Kohn, and Paul Kröger. A truly robust signal temporal logic: Monitoring safety properties of interacting cyber-physical systems under uncertain observation. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/4/126.

Siddiqui:2022:SAD

[1731] Atiq W. Siddiqui and Syed Arshad Raza. A statistical approach to discovering process regime shifts and their determinants. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/4/127.

Alharbe:2022:FGG

[1732] Nawaf Alharbe, Abeer Aljohani, and Mohamed Ali Rakrouki. A fuzzy grouping genetic algorithm for solving a real-world virtual machine placement problem in a healthcare-cloud. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/128.

Saqib:2022:CNN

[1733] Nazmus Saqib, Khandaker Foysal Haque, Venkata Prasanth Yanambaka, and Ahmed Abdelgawad. Convolutional-neural-network-based handwritten character recognition: an approach with massive multisource data. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/4/129.

Silva:2022:MLA

[1734] Hugo Silva and Jorge Bernardino. Machine learning algorithms: an experimental evaluation for decision support systems. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/4/130.

Thelen:2022:MFG

[1735] Andrew S. Thelen, Dean E. Bryson, Bret K. Stanford, and Philip S. Beran. Multi-fidelity gradient-based optimization for high-dimensional aeroelastic configurations. Algorithms (Basel), 15(4), April 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/4/131.

Li:2022:OFD

[1736] Hui Li, Yuan Fang, Zhiguo Huang, Mengyao Zhang, and Qing Wei. Optimizing finite-difference operator in seismic wave numerical modeling. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/4/132.

Zhao:2022:SAW

[1737] Shijun Zhao and Yulong Shan. Study of the algorithm for wind shear detection with lidar based on shear intensity factor. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/4/133.

Wang:2022:RCR

[1738] Wenhao Wang, Dingyuanhao Sun, Feng Jiang, Xingguo Chen, and Cheng Zhu. Research and challenges of reinforcement learning in cyber defense decision-making for intranet security. Algorithms (Basel), 15(4), April 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/4/134.

Determe:2022:MAA

[1739] Jean-François Determe, Sophia Azzagnuni, François Horlin, and Philippe De Doncker. MAC address anonymization for crowd counting. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/5/135.

Li:2022:GSS

[1740] Pengyu Li, Christine Tseng, Yaxuan Zheng, Joyce A. Chew, Longxiu Huang, Benjamin Jarman, and Deanna Needell. Guided semi-supervised non-negative matrix factorization. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/5/136.

Gu:2022:BGT

[1741] Qianqian Gu, Lei Hang, and Shaorong Sun. Behavioral game theory model in pollution control with additional supervision. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/137.

Yang:2022:ACC

[1742] Junxia Yang and Youpeng Zhang. Adaptive cooperative control of multiple urban rail trains with position output constraints. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/138.

Liu:2022:CSV

[1743] Fen Liu and Quan Qian. Cost-sensitive variational autoencoding classifier for imbalanced data classification. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/5/139.

Barradas:2022:RTB

[1744] Adrian Barradas, Acela Tejeda-Gil, and Rosa-María Cantón-Croda. Real-time big data architecture for processing cryptocurrency and social media data: a clustering approach based on k-means. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/140.

Saleh:2022:PFP

[1745] Hayel N. Saleh, Mohammad Imdad, Salvatore Sessa, and Ferdinando Di Martino. Proving fixed-point theorems employing fuzzy (σ, Z)- contractive-type mappings. Algorithms (Basel), 15(5), May 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/5/141.

Pinheiro:2022:OHD

[1746] Giovanni Gatti Pinheiro, Michael Defoin-Platel, and Jean-Charles Regin. Outsmarting human design in airline revenue management. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/142.

Yadav:2022:EAE

[1747] Rohan Kumar Yadav and Dragos Constantin Nicolae. Enhancing attention’s explanation using interpretable Tsetlin machine. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/143.

Koga:2022:SBB

[1748] Kazuki Koga and Kazuhiro Takemoto. Simple black-box universal adversarial attacks on deep neural networks for medical image classification. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/5/144.

Rodriguez:2022:MSS

[1749] J. Apolinar Muñoz Rodríguez. Micro-scale spherical and cylindrical surface modeling via metaheuristic algorithms and micro laser line projection. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/5/
145.

Yang:2022:LSM

[1750] Daniel Yang, Arya Goutam, Kevin Ji, and TJ Tsai. Large-scale multimodal piano music identification using marketplace fingerprinting. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/5/146.

Hu:2022:SBF

[1751] Yuqing Hu, Stéphane Pateux, and Vincent Gripon. Squeezing backbone feature distributions to the max for efficient few-shot learning. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/5/147.

Martino:2022:IGD

[1752] Alessio Martino, Luca Baldini, and Antonello Rizzi. On information granulation via data clustering for granular computing-based pattern recognition: a graph embedding case study. Algorithms (Basel), 15 (5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/5/148.

Zou:2022:ELM

[1753] Yao Zou and Changchun Gao. Extreme learning machine enhanced gradient boosting for credit scoring. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/5/149.

Kodithuwakku:2022:EAR

[1754] Janith Kodithuwakku, Dilki Dandeniya Arachchi, and Jay Rajasekera. An emotion and attention recognition system to classify the level of engagement to a video conversation by participants in real time using machine learning models and utilizing a neural accelerator chip. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/150.

Providas:2022:CFS

[1755] Efthimios Providas. Closed-form solution of the bending two-phase integral model of Euler–Bernoulli nanobeams. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/5/151.

Sanjaya:2022:MNT

[1756] Ricky Sanjaya, Jun Wang, and Yaodong Yang. Measuring the nontransitivity in chess. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/5/152.

Battineni:2022:PMC

[1757] Gopi Battineni, Nalini Chintalapudi, and Gregory Zacharewicz. Process mining in clinical practice: Model evaluations in the central venous catheter installation training. Algorithms (Basel), 15(5), May 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/5/153.

Stanovov:2022:IQM

[1758] Vladimir Stanovov, Shakhnaz Akhmedova, Aleksei Vakhnin, Evgenii Sopov, Eugene Semenkin, and Michael Affenzeller. Improving the quantum multi-swarm optimization with adaptive differential evolution for dynamic environments. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/5/154.

Almutairi:2022:RMA

[1759] Zaynab Almutairi and Hebah Elgibreen. A review of modern audio deepfake detection methods: Challenges and future directions. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/155.

Moldovan:2022:BHO

[1760] Dorin Moldovan. Binary horse optimization algorithm for feature selection. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/5/
156.

Panteleev:2022:OOL

[1761] Andrei V. Panteleev and Anna A. Kolessa. Optimal open-loop control of discrete deterministic systems by application of the perch school metaheuristic optimization algorithm. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/5/157.

Xu:2022:POA

[1762] Fengrui Xu, Mengqiao Chen, Xuelin Liang, and Wensheng Liu. PSO optimized active disturbance rejection control for aircraft anti-skid braking system. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/5/

Zhang:2022:SCB

[1763] Jinsong Zhang, Bao Jin, Junyi Sha, Yan Chen, and Yijin Zhang. SentenceLDA- and ConNetClus-based heterogeneous academic network analysis for publication ranking. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/5/159.

Wu:2022:MMB

[1764] Xing Wu, Yifan Jin, Jianjia Wang, Quan Qian, and Yike Guo. MKD: Mixup-based knowledge distillation for Mandarin end-to-end speech recognition. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/5/160.

Teng:2022:OSL

[1765] Jackson Horlick Teng, Thian Song Ong, Tee Connie, Kalaiarasi Sonai Muthu Anbananthen, and Pa Pa Min. Optimized score level fusion for multi-instance finger vein recognition. Algorithms (Basel), 15 (5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/5/161.

Rauh:2022:EVE

[1766] Andreas Rauh, Yohann Gourret, Katell Lagattu, Bernardo Hummes, Luc Jaulin, Johannes Reuter, Stefan Wirtensohn, and Patrick Hoher. Experimental validation of ellipsoidal techniques for state estimation in marine applications. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/5/162.

Koppl:2022:LRP

[1767] Dominik Köppl. Linking off-road points to routing networks. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/163.

Garvie:2022:PIL

[1768] Marcus R. Garvie and John Burkardt. A parallelizable integer linear programming approach for tiling finite regions of the plane with polyominoes. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/5/164.

Negassi:2022:SSA

[1769] Misgana Negassi, Diane Wagner, and Alexander Reiterer. Smart(Sampling)Augment: Optimal and efficient data augmentation for semantic segmentation. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/5/165.

Bala:2022:EML

[1770] Mousumi Bala, Mohammad Hanif Ali, Md. Shahriare Satu, Khondokar Fida Hasan, and Mohammad Ali Moni. Efficient machine learning models for early stage detection of autism spectrum disorder. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/166.

Sun:2022:CLC

[1771] Peixia Sun, Shengxiong Lao, Dongyang Du, Jiqiang Peng, and Xu Yang. Construction of life-cycle simulation framework of chronic diseases and their comorbidities based on population cohort. Algorithms (Basel), 15 (5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/5/167.

Nogaret:2022:APE

[1772] Alain Nogaret. Approaches to parameter estimation from model neurons and biological neurons. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/5/168.

Wang:2022:SMA

[1773] Yushan Wang, Jonathan Brand, and Wentai Liu. Stimulation Montage achieves balanced focality and intensity. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/5/169.

Li:2022:ACT

[1774] Chunchun Li, Manuel Günther, Akshay Raj Dhamija, Steve Cruz, Mohsen Jafarzadeh, Touqeer Ahmad, and Terrance E. Boult. Agglomerative clustering with threshold optimization via extreme value theory. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/5/170.

Li:2022:ROP

[1775] Xinning Li, Qun He, Qin Yang, Neng Wang, Hu Wu, and Xianhai Yang. Research on an optimal path planning method based on A* algorithm for multi-view recognition. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/5/171.

Lu:2022:TPS

[1776] Rongqin Lu, Xiaomei Zhao, and Yingqi Wang. A tailored pricing strategy for different types of users in hybrid carsharing systems. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/5/172.

Rashid:2022:IPM

[1777] Khandakar M. Rashid and Joseph Louis. Integrating process mining with discrete-event simulation for dynamic productivity estimation in heavy civil construction operations. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/5/173.

Yang:2022:DRC

[1778] Lingkai Yang, Sally McClean, Mark Donnelly, Kevin Burke, and Kashaf Khan. Detecting and responding to concept drift in business processes. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/5/174.

Ghosh:2022:MIA

[1779] Indranil Ghosh, Muhammad Mahbubur Rashid, Shukranul Mawa, Rupal Roy, Md Manjurul Ahsan, Muhammad Ramiz Uddin, Kishor Datta Gupta, and Pallabi Ghosh. A modified iterative algorithm for numerical investigation of HIV infection dynamics. Algorithms (Basel), 15(5), May 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/5/175.

Kong:2022:MCC

[1780] Yanfen Kong, Caiyue Zhou, Chuanyong Zhang, Lin Sun, and Chongbo Zhou. Multi-color channels based group sparse model for image restoration. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/
176.

Rodionov:2022:MAD

[1781] Dmitriy Rodionov, Darya Kryzhko, Timur Tenishev, Victor Uimanov, Alsu Abdulmanova, Ani Kvikviniia, Pavel Aksenov, Mark Solovyov, Fedor Kolomenskii, and Evgenii Konnikov. Methodology for assessing the digital image of an enterprise with its industry specifics. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/6/177.

Waghen:2022:DDF

[1782] Kerelous Waghen and Mohamed-Salah Ouali. A data-driven fault tree for a time causality analysis in an aging system. Algorithms (Basel), 15 (6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/6/178.

Jauberthie:2022:BEP

[1783] Carine Jauberthie and Nathalie Verdière. Bounded-error parameter estimation using integro-differential equations for Hindmarsh–Rose model. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/179.

El-Gharib:2022:DPM

[1784] Najah Mary El-Gharib and Daniel Amyot. Data preprocessing method and API for mining processes from cloud-based application event logs. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/180.

Traver:2022:MCI

[1785] José Emilio Traver, Cristina Nuevo-Gallardo, Paloma Rodríguez, Inés Tejado, and Blas M. Vinagre. Modeling and control of IPMC-based artificial eukaryotic flagellum swimming robot: Distributed actuation. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/181.

Liu:2022:MOM

[1786] Lanfen Liu and Xinfeng Yang. A multi-objective model and algorithms of aggregate production planning of multi-product with early and late delivery. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/182.

Meftah:2022:RDP

[1787] Souhail Meftah, Shuhao Zhang, Bharadwaj Veeravalli, and Khin Mi Mi Aung. Revisiting the design of parallel stream joins on trusted execution environments. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/183.

Alomari:2022:BZP

[1788] Mohammad W. Alomari and Christophe Chesneau. Bounding the zeros of polynomials using the Frobenius companion matrix partitioned by the Cartesian decomposition. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/184.

Zhang:2022:IML

[1789] Dezheng Zhang, Peng Li, and Aziguli Wulamu. An improved multilabel learning method with ELM-RBF and a synergistic adaptive genetic algorithm. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/185.

Bhimavarapu:2022:FSU

[1790] Usharani Bhimavarapu, Nalini Chintalapudi, and Gopi Battineni. A fair and safe usage drug recommendation system in medical emergencies by a stacked ANN. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/186.

Lee:2022:ESR

[1791] Suhwan Lee, Marco Comuzzi, and Nahyun Kwon. Exploring the suitability of rule-based classification to provide interpretability in outcomebased process predictive monitoring. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/6/187.

Hoffmann:2022:LDD

[1792] Volker Hoffmann, Bendik Nybakk Torsæter, Gjert Hovland Rosenlund, and Christian Andre Andresen. Lessons for data-driven modelling from harmonics in the Norwegian grid. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/6/188.

Zhang:2022:NCO

[1793] Quan Zhang, Shiyu Du, Yiming Zhang, Hongzhuo Wu, Kai Duan, and Yanru Lin. A novel chimp optimization algorithm with refraction learning and its engineering applications. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/6/189.

Intawichai:2022:MDR

[1794] Siriwan Intawichai and Saifon Chaturantabut. A missing data reconstruction method using an accelerated least-squares approximation with randomized SVD. Algorithms (Basel), 15(6), June 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/190.

Shkaberina:2022:CAG

[1795] Guzel Shkaberina, Leonid Verenev, Elena Tovbis, Natalia Rezova, and Lev Kazakovtsev. Clustering algorithm with a greedy agglomerative heuristic and special distance measures. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/6/191.

Pambudi:2022:CNS

[1796] Dhidhi Pambudi and Masaki Kawamura. Constructing the neighborhood structure of VNS based on binomial distribution for solving QUBO problems. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/
192.

Antwi:2022:SNO

[1797] Robert Antwi, James Dzisi Gadze, Eric Tutu Tchao, Axel Sikora, Henry Nunoo-Mensah, Andrew Selasi Agbemenu, Kwame Opunie-Boachie Obour Agyekum, Justice Owusu Agyemang, Dominik Welte, and Eliel Keelson. A survey on network optimization techniques for blockchain systems. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/193.

Akpudo:2022:EES

[1798] Ugochukwu Ejike Akpudo and Jang-Wook Hur. Exploring the efficiencies of spectral isolation for intelligent wear monitoring of micro drill bit automatic regrinding in-line systems. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/6/194.

Rodrigues:2022:MLR

[1799] Igor D. Rodrigues, Emerson A. de Carvalho, Caio P. Santana, and Guilherme S. Bastos. Machine learning and rs-fMRI to identify potential brain regions associated with autism severity. Algorithms (Basel), 15 (6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/6/195.

Marshall:2022:PMP

[1800] Adele H. Marshall and Aleksandar Novakovic. Process mining the performance of a real-time healthcare 4.0 systems using conditional survival models. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/
196.

Gjoreski:2022:BDM

[1801] Martin Gjoreski, Vladimir Kuzmanovski, and Marko Bohanec. BAG-DSM: a method for generating alternatives for hierarchical multiattribute decision models using Bayesian optimization. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/6/197.

Luo:2022:IJP

[1802] Yuan Luo, Jiakai Lu, Qiong Qin, and Yanyu Liu. Improved JPS path optimization for mobile robots based on angle-propagation theta* algorithm. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/
198.

El-khawaga:2022:XCP

[1803] Ghada El-khawaga, Mervat Abu-Elkheir, and Manfred Reichert. XAI in the context of predictive process monitoring: an empirical analysis framework. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/199.

Mata:2022:MMG

[1804] Christian Mata, Josep Munuera, Alain Lalande, Gilberto Ochoa-Ruiz, and Raul Benitez. MedicalSeg: a medical GUI application for image segmentation management. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/200.

Fandinno:2022:CRC

[1805] Jorge Fandinno, David Pearce, Concepción Vidal, and Stefan Woltran. Comparing the reasoning capabilities of equilibrium theories and answer set programs. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/201.

Fuchs:2022:CPM

[1806] Franz Georg Fuchs, Kjetil Olsen Lye, Halvor Møll Nilsen, Alexander Johannes Stasik, and Giorgio Sartor. Constraint preserving mixers for the quantum approximate optimization algorithm. Algorithms (Basel), 15 (6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/6/202.

Providas:2022:ACF

[1807] Efthimios Providas. An algorithm for the closed-form solution of certain classes of Volterra–Fredholm integral equations of convolution type. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/6/203.

Papadia:2022:TMA

[1808] Gabriele Papadia, Massimo Pacella, and Vincenzo Giliberti. Topic modeling for automatic analysis of natural language: a case study in an Italian customer support center. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/6/204.

Yang:2022:RML

[1809] Xinyi Yang, Ziyi Wang, Hengxi Zhang, Nan Ma, Ning Yang, Hualin Liu, Haifeng Zhang, and Lei Yang. A review: Machine learning for combinatorial optimization problems in energy areas. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/6/205.

Filion:2022:CFG

[1810] Guillaume J. Filion. Correction: Filion, G. J. Analytic Combinatorics for Computing Seeding Probabilities. Algorithms 2018, 11, 3. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/6/206. See [534].

Dornaika:2022:MVG

[1811] Fadi Dornaika and Abdelmalik Moujahid. Multi-view graph fusion for semi-supervised learning: Application to image-based face beauty prediction. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/
207.

Lionnie:2022:EVE

[1812] Regina Lionnie, Catur Apriono, and Dadang Gunawan. Eyes versus eyebrows: a comprehensive evaluation using the multiscale analysis and curvature-based combination methods in partial face recognition. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/6/208.

Granziol:2022:MEA

[1813] Diego Granziol, Binxin Ru, Xiaowen Dong, Stefan Zohren, Michael Osborne, and Stephen Roberts. Maximum entropy approach to massive graph spectrum learning with applications. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/6/209.

Muscinelli:2022:ODM

[1814] Eugenio Muscinelli, Swapnil Sadashiv Shinde, and Daniele Tarchi. Overview of distributed machine learning techniques for 6G networks. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/6/210.

Mazzoccoli:2022:OCI

[1815] Alessandro Mazzoccoli and Maurizio Naldi. Optimizing cybersecurity investments over time. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/211.

Okwuosa:2022:CEM

[1816] Chibuzo Nwabufo Okwuosa, Ugochukwu Ejike Akpudo, and Jang-Wook Hur. A cost-efficient MCSA-based fault diagnostic framework for SCIM at low-load conditions. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/212.

Ding:2022:NSS

[1817] Kaimeng Ding, Shiping Chen, Jiming Yu, Yanan Liu, and Jie Zhu. A new subject-sensitive hashing algorithm based on MultiRes-RCF for blockchains of HRRS images. Algorithms (Basel), 15(6), June 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/6/213.

Aziz:2022:PSA

[1818] Nor Azlina Ab. Aziz and Kamarulzaman Ab. Aziz. Pendulum search algorithm: an optimization algorithm based on simple harmonic motion and its application for a vaccine distribution problem. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/6/214.

Wang:2022:DGM

[1819] Jigang Wang, Liang Chen, and Rui Wang. Domain generalization model of deep convolutional networks based on SAND-mask. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/6/215.

Stumpf:2022:PEF

[1820] Martin Stumpf. Pulsed electromagnetic field transmission through a small rectangular aperture: a solution based on the Cagniard–DeHoop method of moments. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/216.

Reyes-Rubiano:2022:OAR

[1821] Lorena Reyes-Rubiano, Jana Voegl, and Patrick Hirsch. An online algorithm for routing an unmanned aerial vehicle for road network exploration operations after disasters under different refueling strategies. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/6/217.

Nain:2022:RAI

[1822] Fatini Nadhirah Mohd Nain, Nurul Hashimah Ahamed Hassain Malim, Rosni Abdullah, Muhamad Farid Abdul Rahim, Mohd Azinuddin Ahmad Mokhtar, and Nurul Syafika Mohamad Fauzi. A review of an artificial intelligence framework for identifying the most effective palm oil prediction. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/6/218.

Ansotegui:2022:SFR

[1823] Carlos Ansótegui, Maria Luisa Bonet, and Jordi Levy. Scale-free random SAT instances. Algorithms (Basel), 15(6), June 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/6/219.

Aurenhammer:2022:ICM

[1824] Franz Aurenhammer, Christoph Ladurner, and Michael Steinkogler. Incremental construction of motorcycle graphs. Algorithms (Basel), 15 (7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/7/225.

Haghighi:2022:MMT

[1825] Mohammad Hossein Haghighi and Seyed Meysam Mousavi. A mathematical model and two fuzzy approaches based on credibility and expected interval for project cost-quality-risk trade-off problem in timeconstrained conditions. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/7/226.

Wang:2022:LBO

[1826] Yimeng Wang, Mridul Agarwal, Tian Lan, and Vaneet Aggarwal. Learning-based online QoE optimization in multi-agent video streaming. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/227.

Sawwan:2022:EPC

[1827] Abdalaziz Sawwan and Jie Wu. On edge pruning of communication networks under an age-of-information framework. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/7/228.

Ono:2022:PPF

[1828] Shinji Ono, Jun Takata, Masaharu Kataoka, Tomohiro I, Kilho Shin, and Hiroshi Sakamoto. Privacy-preserving feature selection with fully homomorphic encryption. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/7/229.

Fleischer:2022:TSA

[1829] Jacques Phillipe Fleischer, Gregor von Laszewski, Carlos Theran, and Yohn Jairo Parra Bautista. Time series analysis of cryptocurrency prices using long short-term memory. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/7/230.

Haddi:2022:AAF

[1830] Zouhair Haddi, Bouchra Ananou, Miquel Alfaras, Mustapha Ouladsine, Jean-Claude Deharo, Narcís Avellana, and Stéphane Delliaux. Automatic atrial fibrillation arrhythmia detection using univariate and multivariate data. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/7/231.

Triastcyn:2022:GHF

[1831] Aleksei Triastcyn and Boi Faltings. Generating higher-fidelity synthetic datasets with privacy guarantees. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/7/232.

Xie:2022:ZDL

[1832] Cong Xie, Oluwasanmi Koyejo, and Indranil Gupta. ZenoPS: a distributed learning system integrating communication efficiency and security. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/
233.

Decker:2022:PEO

[1833] Jonathan Decker, Piotr Kasprzak, and Julian Martin Kunkel. Performance evaluation of open-source serverless platforms for Kubernetes. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/7/234.

Gallego:2022:VFC

[1834] Andres Gallego, Francisco Roman, and Edwin Pineda. Vector fitting-Cauchy method for the extraction of complex natural resonances in ground penetrating radar operations. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/7/235.

Filipe:2022:ACF

[1835] Vítor Filipe, Pedro Teixeira, and Ana Teixeira. Automatic classification of foot thermograms using machine learning techniques. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/7/236.

Lee:2022:EPA

[1836] Chun-Liang Lee, Guan-Yu Lin, and Yaw-Chung Chen. An efficient parallel algorithm for detecting packet filter conflicts. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/7/237.

Du:2022:RNA

[1837] Xiaoyu Du, Cheng Cheng, Yujing Wang, and Zhijie Han. Research on network attack traffic detection HybridAlgorithm based on UMAP-RF. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/238.

Lysenko:2022:IMV

[1838] Sergii Lysenko, Kira Bobrovnikova, Vyacheslav Kharchenko, and Oleg Savenko. IoT multi-vector cyberattack detection based on machine learning algorithms: Traffic features analysis, experiments, and efficiency. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/7/239.

Yang:2022:TTE

[1839] Lu Yang, Chishe Wang, and Zhibin Li. Tunnel traffic evolution during capacity drop based on high-resolution vehicle trajectory data. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/7/240.

Eigel:2022:TOU

[1840] Martin Eigel, Marvin Haase, and Johannes Neumann. Topology optimisation under uncertainties with neural networks. Algorithms (Basel), 15 (7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/7/241.

Petty:2022:NSE

[1841] Taylor Petty, Jan Hannig, Tunde I. Huszar, and Hari Iyer. A new string edit distance and applications. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/7/242.

Cellamare:2022:FGL

[1842] Matteo Cellamare, Anna J. van Gestel, Hasan Alradhi, Frank Martin, and Arturo Moncada-Torres. A federated generalized linear model for privacy-preserving analysis. Algorithms (Basel), 15(7), July 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/7/243.

Cao:2022:IAA

[1843] Jin Cao, Bo Li, Mengni Fan, and Huiyu Liu. Inference acceleration with adaptive distributed DNN partition over dynamic video stream. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/244.

Song:2022:ELR

[1844] Tao Song, Huanyu Liao, and Ganesh Subbarayan. Efficient local refinement near parametric boundaries using kd-tree data structure and algebraic level sets. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/7/245.

Rooyani:2022:TSM

[1845] Danial Rooyani and Fantahun Defersha. A two-stage multi-objective genetic algorithm for a flexible job shop scheduling problem with lot streaming. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/7/246.

Fedorchenko:2022:CRI

[1846] Elena Fedorchenko, Evgenia Novikova, and Anton Shulepov. Comparative review of the intrusion detection systems based on federated learning: Advantages and open challenges. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/7/247.

Li:2022:SAM

[1847] Bochong Li, Ryo Oka, Ping Xuan, Yuichiro Yoshimura, and Toshiya Nakaguchi. Semi-automatic multiparametric MR imaging classification using novel image input sequences and 3D convolutional neural networks. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/248.

Petry:2022:IIV

[1848] Fred Petry and Ronald Yager. Intuitionistic and interval-valued fuzzy set representations for data mining. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/7/249.

Yildiz:2022:MFL

[1849] Sihmehmet Yildiz, Hayriye Pehlivan Solak, and Melike Nikbay. Multifidelity low-rank approximations for uncertainty quantification of a supersonic aircraft design. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/7/250.

Almalki:2022:AIC

[1850] Sultan Ahmed Almalki, Ahmed Abdel-Rahim, and Frederick T. Sheldon. Adaptive IDS for cooperative intelligent transportation systems using deep belief networks. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/7/251.

Qian:2022:IWH

[1851] Yufan Qian, Limei Tian, Baichen Zhai, Shufan Zhang, and Rui Wu. Informer-WGAN: High missing rate time series imputation based on adversarial training and a self-attention mechanism. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/7/252.

Muller:2022:LCC

[1852] Ralf Reiner Müller, Bernhard Martin Wilhelm Gäde, and Ali Bereyhi. Linear computation coding: a framework for joint quantization and computing. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/
253.

Pedersen:2022:LAM

[1853] Joseph Pedersen, Rafael Muñoz-Gómez, Jiangnan Huang, Haozhe Sun, Wei-Wei Tu, and Isabelle Guyon. LTU attacker for membership inference. Algorithms (Basel), 15(7), July 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/7/254.

Trella:2022:DRL

[1854] Anna L. Trella, Kelly W. Zhang, Inbal Nahum-Shani, Vivek Shetty, Finale Doshi-Velez, and Susan A. Murphy. Designing reinforcement learning algorithms for digital interventions: Pre-implementation guidelines. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/8/255.

Koganti:2022:DMO

[1855] Srilakshmi Koganti, Krishna Jyothi Koganti, and Surender Reddy Salkuti. Design of multi-objective-based artificial intelligence controller for wind/battery-connected shunt active power filter. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/8/256.

Tariq:2022:ADS

[1856] Zeeshan Tariq, Darryl Charles, Sally McClean, Ian McChesney, and Paul Taylor. Anomaly detection for service-oriented business processes using conformance analysis. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/8/257.

Mijalkovic:2022:RFN

[1857] Jovana Mijalkovic and Angelo Spognardi. Reducing the false negative rate in deep learning based network intrusion detection systems. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/8/258.

DeSantis:2022:MCM

[1858] Enrico De Santis, Parisa Naraei, Alessio Martino, Alireza Sadeghian, and Antonello Rizzi. Multifractal characterization and modeling of blood pressure signals. Algorithms (Basel), 15(8), August 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/8/259.

LeCaillec:2022:HTF

[1859] Jean-Marc Le Caillec. Hypothesis testing fusion for nonlinearity detection in hedge fund price returns. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/260.

Bertoluzzo:2022:OCN

[1860] Manuele Bertoluzzo, Paolo Di Barba, Michele Forzan, Maria Evelina Mognaschi, and Elisabetta Sieni. Optimization of compensation network for a wireless power transfer system in dynamic conditions: a circuit analysis approach. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/8/261.

Kuikka:2022:EAC

[1861] Vesa Kuikka, Henrik Aalto, Matias Ijäs, and Kimmo K. Kaski. Efficiency of algorithms for computing influence and information spreading on social networks. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/262.

daSilva:2022:NME

[1862] Matheus Alencar da Silva, Bonfim Amaro Junior, Ramon Rudá Brito Medeiros, and Plácido Rogério Pinheiro. A neuroevolutionary model to estimate the tensile strength of manufactured parts made by 3D printing. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/8/263.

Li:2022:IDV

[1863] Leon Li. Intensity and direction of volatility spillover effect in carbonenergy markets: a regime-switching approach. Algorithms (Basel), 15 (8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/8/264.

Rostami:2022:MCD

[1864] Parinaz Rostami, Soroush Avakh Darestani, and Mitra Movassaghi. Modelling cross-docking in a three-level supply chain with stochastic service and queuing system: MOWFA algorithm. Algorithms (Basel), 15 (8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/8/265.

Prykarpatsky:2022:DTD

[1865] Yarema A. Prykarpatsky, Ilona Urbaniak, Radosław A. Kycia, and Anatolij K. Prykarpatski. Dark type dynamical systems: The integrability algorithm and applications. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/266.

Patel:2022:STC

[1866] Vrushang Patel, Sheela Ramanna, Ketan Kotecha, and Rahee Walambe. Short text classification with tolerance-based soft computing method. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/8/267.

Carvalho:2022:CIW

[1867] Mateus F. T. Carvalho, Sergio A. Silva, Jr., Carla Cristina O. Bernardo, Franklin César Flores, Juliana Vanessa C. M. Perles, Jacqueline Nelisis Zanoni, and Yandre M. G. Costa. Cancer identification in Walker 256 tumor model exploring texture properties taken from microphotograph of rats liver. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/268.

Han:2022:RFA

[1868] Yan Han, Liang Shu, Ziran Wu, Xuan Chen, Gaoyan Zhang, and Zili Cai. Research of flexible assembly of miniature circuit breakers based on robot trajectory optimization. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/269.

Hunter:2022:VTA

[1869] Elizabeth Hunter and John D. Kelleher. Validating and testing an agentbased model for the spread of COVID-19 in Ireland. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/8/270.

Younis:2022:OME

[1870] Adel Younis, Fadi AlKhatib, and Zuomin Dong. Optimal motorcycle engine mount design parameter identification using robust optimization algorithms. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/271.

Macias:2022:BTR

[1871] Nathalie Campos Macias, Wilhelm Düggelin, Yesim Ruf, and Thomas Hanne. Building a technology recommender system using Web crawling and natural language processing technology. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/8/272.

Khan:2022:CEV

[1872] Afsana Khan, Marijn ten Thij, and Anna Wilbik. Communicationefficient vertical federated learning. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/8/273.

Giacomini:2022:NNA

[1873] Giorgia Giacomini, Caterina Graziani, Veronica Lachi, Pietro Bongini, Niccolò Pancino, Monica Bianchini, Davide Chiarugi, Angelo Valleriani, and Paolo Andreini. A neural network approach for the analysis of reproducible ribo-seq profiles. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/274.

Fedoseev:2022:NSS

[1874] Petr Fedoseev, Dmitriy Pesterev, Artur Karimov, and Denis Butusov. New step size control algorithm for semi-implicit composition ODE solvers. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/275.

Rauh:2022:ARE

[1875] Andreas Rauh, Luc Jaulin, and Julien Alexandre dit Sandretto. Algorithms for reliable estimation, identification and control. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/8/276.

Montoya:2022:SPI

[1876] Oscar Danilo Montoya, Walter Gil-González, and Luis Fernando Grisales-Noreña. Solar photovoltaic integration in monopolar DC networks via the GNDO algorithm. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/277.

Hauke:2022:SLS

[1877] Guillermo Hauke and Jorge Lanzarote. Simulation of low-speed buoyant flows with a stabilized compressible/incompressible formulation: The full Navier–Stokes approach versus the Boussinesq model. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/8/278.

Younis:2022:HFS

[1878] Adel Younis and Zuomin Dong. High-fidelity surrogate based multiobjective optimization algorithm. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/8/279.

Lundberg:2022:RTE

[1879] Lars Lundberg and Håkan Grahn. Research trends, enabling technologies and application areas for Big Data. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/8/280.

Elhariri:2022:APL

[1880] Esraa Elhariri, Nashwa El-Bendary, and Shereen A. Taie. Automated pixel-level deep crack segmentation on historical surfaces using U-net models. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/281.

Lu:2022:TPM

[1881] Yunlong Lu and Wenxin Li. Techniques and paradigms in modern game AI systems. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/282.

Zhao:2022:ATM

[1882] Weimin Zhao, Sanaa Alwidian, and Qusay H. Mahmoud. Adversarial training methods for deep learning: a systematic review. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/8/283.

Paul:2022:VBA

[1883] Satyam Paul, Rob Turnbull, Davood Khodadad, and Magnus Löfstrand. A vibration based automatic fault detection scheme for drilling process using type-2 fuzzy logic. Algorithms (Basel), 15(8), August 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/8/284.

Lee:2022:FCD

[1884] Chun-Liang Lee, Guan-Yu Lin, and Yaw-Chung Chen. Fast conflict detection for multi-dimensional packet filters. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/8/285.

Mannone:2022:TBS

[1885] Maria Mannone and Takashi Yoshino. Temari balls, spheres, Sphere-Harmonic: From Japanese folkcraft to music. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/8/286.

Islam:2022:CBT

[1886] Md. Monirul Islam, Md. Belal Hossain, Md. Nasim Akhtar, Mohammad Ali Moni, and Khondokar Fida Hasan. CNN based on transfer learning models using data augmentation and transformation for detection of concrete crack. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/8/287.

Xu:2022:CVS

[1887] Jianlou Xu, Yuying Guo, Yan Hao, and Leigang Huo. A coupled variational system for image decomposition along with edges detection. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/8/288.

Castaneda:2022:BRD

[1888] Juliana Castaneda, Mattia Neroni, Majsa Ammouriova, Javier Panadero, and Angel A. Juan. Biased-randomized discrete-event heuristics for dynamic optimization with time dependencies and synchronization. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/8/289.

Jensen:2022:DTO

[1889] Mathias Højgaard Jensen, Sarang Joshi, and Stefan Sommer. Discretetime observations of Brownian motion on Lie groups and homogeneous spaces: Sampling and metric estimation. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/15/8/290.

Mehta:2022:SMH

[1890] Harshkumar Mehta and Kalpdrum Passi. Social media hate speech detection using explainable artificial intelligence (XAI). Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/8/291.

Khoperskov:2022:IEO

[1891] Alexander V. Khoperskov and Maxim V. Polyakov. Improving the efficiency of oncological diagnosis of the breast based on the combined use of simulation modeling and artificial intelligence algorithms. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/8/292.

Chen:2022:DAF

[1892] Zesheng Chen, Li-Chi Chang, Chao Chen, Guoping Wang, and Zhuming Bi. Defending against FakeBob adversarial attacks in speaker verification systems with noise-adding. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/293.

Simonet:2022:PRA

[1893] Geneviève Simonet and Anne Berry. Properties and recognition of atom graphs. Algorithms (Basel), 15(8), August 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/8/294.

Tsoulos:2022:QPS

[1894] Ioannis G. Tsoulos. QFC: a parallel software tool for feature construction, based on grammatical evolution. Algorithms (Basel), 15(8), August 2022.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/8/295.

Fenu:2022:SPA

[1895] Caterina Fenu, Lothar Reichel, and Giuseppe Rodriguez. SoftNet: a package for the analysis of complex networks. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/9/296.

Sechi:2022:SNT

[1896] Renata Sechi, Konstantin Fackeldey, Surahit Chewle, and Marcus Weber. SepFree NMF: a toolbox for analyzing the kinetics of sequential spectroscopic data. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/297.

Nguyen:2022:ODP

[1897] Van-Vinh Nguyen, Nhat-Tung Nguyen, Quang-Thuan Nguyen, Van-Hai Bui, and Wencong Su. Optimal design parameters for hybrid DC circuit breakers using a multi-objective genetic algorithm. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/9/298.

Rashid:2022:AIA

[1898] Mudassir M. Rashid, Mohammad Reza Askari, Canyu Chen, Yueqing Liang, Kai Shu, and Ali Cinar. Artificial intelligence algorithms for treatment of diabetes. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/299.

Pap:2022:CAF

[1899] Jozsef Pap, Csaba Mako, Miklos Illessy, Zef Dedaj, Sina Ardabili, Bernat Torok, and Amir Mosavi. Correlation analysis of factors affecting firm performance and employees wellbeing: Application of advanced machine learning analysis. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/300.

Panteleev:2022:ATF

[1900] Andrei V. Panteleev and Anna A. Kolessa. Application of the Tomtit flock metaheuristic optimization algorithm to the optimal discrete time deterministic dynamical control problem. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/9/301.

Al-Yahya:2022:BPC

[1901] Tasniem Al-Yahya, Mohamed El Bachir Abdelkrim Menai, and Hassan Mathkour. Boosting the performance of CDCL-based SAT solvers by exploiting backbones and backdoors. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/9/302.

Castaneda:2022:DGB

[1902] Juliana Castaneda, Assumpta Jover, Laura Calvet, Sergi Yanes, Angel A. Juan, and Milagros Sainz. Dealing with gender bias issues in data-algorithmic processes: a social-statistical perspective. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/9/303.

Kohler:2022:CAP

[1903] Alexander Köhler and Michael Breuß. Computational analysis of PDEbased shape analysis models by exploring the damped wave equation. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
304.

Theodosiou:2022:SAD

[1904] Zenonas Theodosiou, Marios Thoma, Harris Partaourides, and Andreas Lanitis. A systematic approach for developing a robust artwork recognition framework using smartphone cameras. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/9/305.

Li:2022:MRI

[1905] Hongbing Li and Qunfei Zhang. MIMO radar imaging method with nonorthogonal waveforms based on deep learning. Algorithms (Basel), 15 (9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/9/306.

Tang:2022:TDE

[1906] Wenyun Tang, Jiahui Chen, Chao Sun, Hanbing Wang, and Gen Li. Traffic demand estimations considering route trajectory reconstruction in congested networks. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/307.

Mondol:2022:EPC

[1907] Chaity Mondol, F. M. Javed Mehedi Shamrat, Md. Robiul Hasan, Saidul Alam, Pronab Ghosh, Zarrin Tasnim, Kawsar Ahmed, Francis M. Bui, and Sobhy M. Ibrahim. Early prediction of chronic kidney disease: a comprehensive performance analysis of deep learning models. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/9/308.

Aguru:2022:IIR

[1908] Aswani Devi Aguru, Erukala Suresh Babu, Soumya Ranjan Nayak, Abhisek Sethy, and Amit Verma. Integrated industrial reference architecture for smart healthcare in Internet of Things: a systematic investigation. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
309.

Lang:2022:ISP

[1909] Jens Lang and Bernhard A. Schmitt. Implicit A-stable peer triplets for ODE constrained optimal control problems. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/9/310.

Ba:2022:ASA

[1910] Fatima Antarou Ba and Michael Quellmalz. Accelerating the Sinkhorn algorithm for sparse multi-marginal optimal transport via Fast Fourier Transforms. Algorithms (Basel), 15(9), September 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/311.

Belim:2022:ISB

[1911] Sergey V. Belim and Svetlana Yu. Belim. Images segmentation based on cutting the graph into communities. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/15/9/312.

Maddalena:2022:AIC

[1912] Lucia Maddalena, Laura Antonelli, Alexandra Albu, Aroj Hada, and Mario Rosario Guarracino. Artificial intelligence for cell segmentation, event detection, and tracking for label-free microscopy imaging. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/313.

Grigorescu:2022:CBB

[1913] Octavian Grigorescu, Andreea Nica, Mihai Dascalu, and Razvan Rughinis. CVE2ATT&CK: BERT-based mapping of CVEs to MITRE ATT&CK techniques. Algorithms (Basel), 15(9), September 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/9/314.

Sipper:2022:HPP

[1914] Moshe Sipper. High per parameter: a large-scale study of hyperparameter tuning for machine learning algorithms. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/9/315.

Shaked:2022:SRI

[1915] Avi Shaked and Oded Margalit. Sustainable risk identification using formal ontologies. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/316.

Zhang:2022:ISM

[1916] Yunyang Zhang, Shiyu Du, and Quan Zhang. Improved slime mold algorithm with dynamic quantum rotation gate and opposition-based learning for global optimization and engineering design problems. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/9/317.

Gembarski:2022:JCS

[1917] Paul Christoph Gembarski. Joining constraint satisfaction problems and configurable CAD product models: a step-by-step implementation guide. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
318.

Tong:2022:FON

[1918] Qianqian Tong, Guannan Liang, Jiahao Ding, Tan Zhu, Miao Pan, and Jinbo Bi. Federated optimization of l_{0}-norm regularized sparse learning. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
319.

Khalil:2022:APP

[1919] Omar Khalil, Hany El-Sharkawy, Maha Youssef, and Gerd Baumann. Adaptive piecewise poly-sinc methods for ordinary differential equations.

Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
320.

Park:2022:PSS

[1920] Chan Hee Park and Young Dae Ko. A practical staff scheduling strategy considering various types of employment in the construction industry. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
321.

Nagahara:2022:PSR

[1921] Masaaki Nagahara, Yu Iwai, and Noboru Sebe. Projection onto the set of rank-constrained structured matrices for reduced-order controller design. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
322.

Allassonniere:2022:SIS

[1922] Stéphanie Allassonnière. Special issue: Stochastic algorithms and their applications. Algorithms (Basel), 15(9), September 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/323.

Lukyanenko:2022:ARE

[1923] Dmitry Lukyanenko, Valentin Shinkarev, and Anatoly Yagola. Accounting for round-off errors when using gradient minimization methods. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/324.

Moya:2022:FDS

[1924] Christian Moya and Guang Lin. Fed-DeepONet: Stochastic gradientbased federated training of deep operator networks. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/9/325.

Karthikeyan:2022:AIM

[1925] P. Karthikeyan, Wei-Lun Chen, and Pao-Ann Hsiung. Autonomous intersection management by using reinforcement learning. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/9/326.

Ye:2022:IID

[1926] Zhen Ye, Xiaoming Ou, Juhua Huang, and Yingpin Chen. Infrared image deblurring based on lp-pseudo-norm and high-order overlapping group sparsity regularization. Algorithms (Basel), 15(9), September 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/9/327.

Reich:2022:MAP

[1927] Thilo Reich, David Hulbert, and Marcin Budka. A model architecture for public transport networks using a combination of a recurrent neural network encoder library and a attention mechanism. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/9/328.

Demidova:2022:CPT

[1928] Liliya A. Demidova and Artyom V. Gorchakov. Classification of program texts represented as Markov chains with biology-inspired algorithms — enhanced extreme learning machines. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/15/9/329.

Conte:2022:PMC

[1929] Mattia Conte, Andrea M. Chiariello, Alex Abraham, Simona Bianco, Andrea Esposito, Mario Nicodemi, Tommaso Matteuzzi, and Francesca Vercellone. Polymer models of chromatin imaging data in single cells. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/9/
330.

Su:2022:IWT

[1930] Shaohui Su, Jiadong Wang, and Dongyang Zhang. Irregular workpiece template-matching algorithm using contour phase. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/9/331.

Louk:2022:TBC

[1931] Maya Hilda Lestari Louk and Bayu Adhi Tama. Tree-based classifier ensembles for PE malware analysis: a performance revisit. Algorithms (Basel), 15(9), September 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/9/332.

Paul:2022:NID

[1932] Jestine Paul, Benjamin Hong Meng Tan, Bharadwaj Veeravalli, and Khin Mi Mi Aung. Non-interactive decision trees and applications with multibit TFHE. Algorithms (Basel), 15(9), September 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/9/333.

Adeniyi:2022:CCM

[1933] Abidemi Emmanuel Adeniyi, Sanjay Misra, Eniola Daniel, and Anthony Bokolo, Jr. Computational complexity of modified Blowfish cryptographic algorithm on video data. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/373.

Akhmedova:2022:HCA

[1934] Shakhnaz Akhmedova, Vladimir Stanovov, and Yukihiro Kamiya. A hybrid clustering approach based on fuzzy logic and evolutionary computation for anomaly detection. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/342.

Al-Azzeh:2022:MAI

[1935] Jamil Al-Azzeh, Abdelwadood Mesleh, Maksym Zaliskyi, Roman Odarchenko, and Valeriyi Kuzmin. A method of accuracy increment using segmented regression. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/378.

Ambrosino:2022:CHI

[1936] Daniela Ambrosino, Carmine Cerrone, and Anna Sciomachen. A constructive heuristics and an iterated neighborhood search procedure to solve the cost-balanced path problem. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/364.

Anam:2022:IIB

[1937] Choirul Anam, Ariij Naufal, Heri Sutanto, Kusworo Adi, and Geoff Dougherty. Impact of iterative bilateral filtering on the noise power spectrum of computed tomography images. Algorithms (Basel), 15(10): ??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/374.

Asadzadeh:2022:SCA

[1938] Mohammad Asadzadeh and Larisa Beilina. Stability and convergence analysis of a domain decomposition FE/FD method for Maxwell’s equations in the time domain. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/337.

Askari:2022:DCU

[1939] Mohammad Reza Askari, Mahmoud Abdel-Latif, Mudassir Rashid, Mert Sevil, and Ali Cinar. Detection and classification of unannounced physical activities and acute psychological stress events for interventions in diabetes treatment. Algorithms (Basel), 15(10):??, October 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/10/352.

Banerjee:2022:NAF

[1940] Amit Banerjee and Issam Abu-Mahfouz. A novel adaptive FCM with cooperative multi-population differential evolution optimization. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/15/10/380.

Baslyman:2022:RAC

[1941] Malak Baslyman, Daniel Amyot, and John Mylopoulos. Reasoning about confidence in goal satisfaction. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/343.

Bhukya:2022:EMP

[1942] Laxman Bhukya, Narender Reddy Kedika, and Surender Reddy Salkuti. Enhanced maximum power point techniques for solar photovoltaic system under uniform insolation and partial shading conditions: a review. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
10/365.

Boult:2022:WOW

[1943] Terrance E. Boult, Nicolas M. Windesheim, Steven Zhou, Christopher Pereyda, and Lawrence B. Holder. Weibull-open-world (WOW) multitype novelty detection in CartPole3D. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/381.

Buayen:2022:EMK

[1944] Patcharin Buayen and Jeeraporn Werapun. Efficient 0/1-multiple-knapsack problem solving by hybrid DP transformation and robust unbiased filtering. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/366.

Campo:2022:FJS

[1945] Emiro Antonio Campo, Jose Alejandro Cano, Rodrigo Gómez-Montoya, Elkin Rodríguez-Velásquez, and Pablo Cortés. Flexible job shop scheduling problem with fuzzy times and due-windows: Minimizing weighted tardiness and earliness using genetic algorithms. Algorithms (Basel), 15 (10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/10/334.

Chen:2022:DIN

[1946] Chin-Yi Chen and Jih-Jeng Huang. Defuzzify imprecise numbers using the Mellin transform and the trade-off between the mean and spread. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
10/355.

Cordero:2022:DSF

[1947] Alicia Cordero, Miguel A. Leonardo Sepúlveda, and Juan R. Torregrosa. Dynamics and stability on a family of optimal fourth-order iterative methods. Algorithms (Basel), 15(10):??, October 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/387.

Crepey:2022:ADF

[1948] Stéphane Crépey, Noureddine Lehdili, Nisrine Madhar, and Maud Thomas. Anomaly detection in financial time series by principal component analysis and neural networks. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/385.

Elmi:2022:BCJ

[1949] Atabak Elmi, Dhananjay R. Thiruvady, and Andreas T. Ernst. Blocking cyclic job-shop scheduling problems. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/375.

Falini:2022:RRS

[1950] Antonella Falini. RSB: Robust successive binarization for change detection in bitemporal hyperspectral images. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/340.

Fiorini:2022:LCA

[1951] Stefano Fiorini, Michele Ciavotta, and Andrea Maurino. Listening to the city, attentively: a spatio-temporal attention-boosted autoencoder for the short-term flow prediction problem. Algorithms (Basel), 15(10): ??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/376.

Franchini:2022:BIC

[1952] Giorgia Franchini, Micaela Verucchi, Ambra Catozzi, Federica Porta, and Marco Prato. Biomedical image classification via dynamically early stopped artificial neural network. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/386.

Gomez:2022:NSS

[1953] Jonatan Gomez and Andres Rivera. Non-stationary stochastic global optimization algorithms. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/362.

Guo:2022:FBU

[1954] Zhihao Guo, Shengyuan Chen, Xiao Huang, Zhiqiang Qian, Chunsing Yu, Yan Xu, and Fang Ding. Fair benchmark for unsupervised node representation learning. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/379.

Hammer:2022:AAD

[1955] Christof Hammer, Sebastian Sporrer, Johannes Warmer, Peter Kaul, Ronald Thoelen, and Norbert Jung. Algorithms for automatic data validation and performance assessment of MOX gas sensor data using time series analysis. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/360.

Haritha:2022:DFC

[1956] K. Haritha, M. V. Judy, Konstantinos Papageorgiou, Vassilis C. Georgiannis, and Elpiniki Papageorgiou. Distributed fuzzy cognitive maps for feature selection in big data classification. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/383.

Jain:2022:FWP

[1957] Anuj Jain and Sartaj Sahni. Foremost walks and paths in interval temporal graphs. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/361.

Karaseva:2022:EAI

[1958] Tatiana Karaseva and Eugene Semenkin. Evolutionary approaches to the identification of dynamic processes in the form of differential equations and their systems. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/351.

Kazemi:2022:CSB

[1959] Pooyan Kazemi, Aldo Ghisi, and Stefano Mariani. Classification of the structural behavior of tall buildings with a diagrid structure: a machine learning-based approach. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/349.

Leipold:2022:TQA

[1960] Hannes Leipold, Federico M. Spedalieri, and Eleanor Rieffel. Tailored quantum alternating operator Ansätzes for circuit fault diagnostics. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
10/356.

Li:2022:JTP

[1961] Xinning Li, Qin Yang, Hu Wu, Shuai Tan, Qun He, Neng Wang, and Xianhai Yang. Joints trajectory planning of robot based on slime mould whale optimization algorithm. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/363.

Lin:2022:TDA

[1962] Baihan Lin. Topological data analysis in time series: Temporal filtration and application to single-cell genomics. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/371.

Mathur:2022:CDV

[1963] Anant Mathur, Sarat Moka, and Zdravko Botev. Coordinate descent for variance-component models. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/354.

Mora-Burbano:2022:ASO

[1964] Javier Andrés Mora-Burbano, Cristian David Fonseca-Díaz, and Oscar Danilo Montoya. Application of the SSA for optimal reactive power compensation in radial and meshed distribution using D-STATCOMs. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
10/345.

Morera:2022:RVC

[1965] Hunter Morera, Roshan Warman, Azubuogu Anudu, Chukwudumebi Uche, Ivana Radosavljevic, Nikhil Reddy, Ahan Kayastha, Niharika Baviriseaty, Rahul Mhaskar, Andrew A. Borkowski, Patrick Brady, Satish Singh, Gerard Mullin, Jose Lezama, Lawrence O. Hall, Dmitry Goldgof, and Gitanjali Vidyarthi. Reduction of video capsule endoscopy reading times using deep learning with small data. Algorithms (Basel), 15 (10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/10/339.

Moskalenko:2022:MTM

[1966] Viacheslav Moskalenko, Vyacheslav Kharchenko, Alona Moskalenko, and Sergey Petrov. Model and training method of the resilient image classifier considering faults, concept drift, and adversarial attacks. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/10/384.

Nahar:2022:NAM

[1967] Mahbubun Nahar, A. H. M. Kamal, Md Rafiul Hassan, and Mohammad Ali Moni. Novel algorithm for multi-time data implantation in a special cyber-manufacturing architecture. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/335.

Neagu:2022:TSA

[1968] Dan Claudiu Neagu, Andrei Bogdan Rus, Mihai Grec, Mihai Augustin Boroianu, Nicolae Bogdan, and Attila Gal. Towards sentiment analysis for Romanian Twitter content. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/10/357.

Peng:2022:SFD

[1969] Bo Peng, Ying Bi, Bing Xue, Mengjie Zhang, and Shuting Wan. A survey on fault diagnosis of rolling bearings. Algorithms (Basel), 15(10): ??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/347.

Roanes-Lozano:2022:CAO

[1970] Eugenio Roanes-Lozano. A computational approach to overtaking station track layout design using graphs: an extension that supports special turnouts — an improved alternative track layout proposal. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/10/368.

Rodriguez-Sanchez:2022:CCN

[1971] Rosa María Rodriguez-Sánchez and Jorge Chamorro-Padial. Corner centrality of nodes in multilayer networks: a case study in the network analysis of keywords. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/336.

Saco:2022:ODA

[1972] Arun Saco, P. Shanmuga Sundari, Karthikeyan J, and Anand Paul. An optimized data analysis on a real-time application of PEM fuel cell design by using machine learning algorithms. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/346.

Seda:2022:APR

[1973] Milos Seda. The assignment problem and its relation to logistics problems. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
15/10/377.

Sepulveda-Garcia:2022:PFS

[1974] Simón Sepúlveda-García, Oscar Danilo Montoya, and Alejandro Garcés. Power flow solution in bipolar DC networks considering a neutral wire and unbalanced loads: a hyperbolic approximation. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/10/341.

Setukha:2022:CML

[1975] Alexey Setukha and Rufina Tretiakova. Computational modeling of lymph filtration and absorption in the lymph node by boundary integral equations. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/388.

Shadimetov:2022:OQF

[1976] Kholmat Shadimetov, Abdullo Hayotov, and Botir Abdikayimov. On an optimal quadrature formula in a Hilbert space of periodic functions. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
10/344.

Sisias:2022:DLP

[1977] George Sisias, Myrto Konstantinidou, and Sotirios Kontogiannis. Deep learning process and application for the detection of dangerous goods passing through motorway tunnels. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/370.

Sorokovikov:2022:MFP

[1978] Pavel Sorokovikov and Alexander Gornov. Modifications of flower pollination, teacher-learner and firefly algorithms for solving multiextremal optimization problems. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/359.

Spanos:2022:ADS

[1979] Athanasios C. Spanos, Sotiris P. Gayialis, Evripidis P. Kechagias, and Georgios A. Papadopoulos. An application of a decision support system enabled by a hybrid algorithmic framework for production scheduling in an SME manufacturer. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/10/372.

Stanojevic:2022:CMP

[1980] Vukasin Stanojević, Lev Kazakovtsev, Predrag S. Stanimirović, Natalya Rezova, and Guzel Shkaberina. Calculating the Moore–Penrose generalized inverse on massively parallel systems. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/348.

Stotzner:2022:MDD

[1981] Miles Stötzner, Steffen Becker, Uwe Breitenbücher, Kálmán Képes, and Frank Leymann. Modeling different deployment variants of a composite application in a single declarative deployment model. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/10/382.

Tang:2022:UEA

[1982] Yue Ting Tang and Roman Romero-Ortuno. Using Explainable AI (XAI) for the prediction of falls in the older population. Algorithms (Basel), 15 (10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/10/353.

terBurg:2022:CAE

[1983] Kaya ter Burg and Heysem Kaya. Comparing approaches for explaining DNN-based facial expression classifications. Algorithms (Basel), 15(10): ??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/367.

Tey:2022:CSR

[1984] Wan Teng Tey, Tee Connie, Kan Yeep Choo, and Michael Kah Ong Goh. Cicada species recognition based on acoustic signals. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/10/358.

Wu:2022:FPC

[1985] Aihua Wu, Yinjia Ding, Jingfeng Mao, and Xudong Zhang. A fast point clouds registration algorithm based on ISS-USC feature for the 3D laser scanner. Algorithms (Basel), 15(10):??, October 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/389.

Yao:2022:ETP

[1986] Hui Yao and Thomas Taimre. Estimating tail probabilities of random sums of phase-type scale mixture random variables. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/10/350.

Zheng:2022:GRD

[1987] Chaoran Zheng, Mohsen Eskandari, Ming Li, and Zeyue Sun. GAreinforced deep neural network for net electric load forecasting in microgrids with renewable energy resources for scheduling battery energy storage systems. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/10/338.

Zhou:2022:SAA

[1988] Xudong Zhou, Nobuo Funabiki, Hein Htet, Ariel Kamoyedji, Irin Tri Anggraini, Yuanzhi Huo, and Yan Watequlis Syaifudin. A static assignment algorithm of uniform jobs to workers in a user-PC computing system using simultaneous linear equations. Algorithms (Basel), 15(10):??, October 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/10/369.

Aaleskog:2022:RDL

[1989] Christoffer Åleskog, Håkan Grahn, and Anton Borg. Recent developments in low-power AI accelerators: a survey. Algorithms (Basel), 15 (11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/11/419.

Adubi:2022:EAB

[1990] Stephen A. Adubi, Olufunke O. Oladipupo, and Oludayo O. Olugbara. Evolutionary algorithm-based iterated local search hyper-heuristic for combinatorial optimization problems. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/11/405.

Ahmed:2022:ABF

[1991] Ahmed Abdelmoamen Ahmed. An actor-based formal model and runtime environment for resource-bounded IoT services. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/11/390.

Amirzadeh:2022:AAI

[1992] Rasoul Amirzadeh, Asef Nazari, and Dhananjay Thiruvady. Applying artificial intelligence in cryptocurrency markets: a survey. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/428.

Arasli:2022:DSD

[1993] Batuhan Arasli and Sennur Ulukus. Dynamic SAFFRON: Disease control over time via group testing. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/11/437.

Askar:2022:TSD

[1994] Ali Habeeb Askar, Issa Omle, Endre Kovács, and János Majár. Testing some different implementations of heat convection and radiation in the leapfrog-hopscotch algorithm. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/11/400.

Baharudin:2022:ICE

[1995] Farah Nur Arina Baharudin, Nor Azlina Ab. Aziz, Mohamad Razwan Abdul Malek, Anith Khairunnisa Ghazali, and Zuwairie Ibrahim. Indoor comfort and energy consumption optimization using an inertia weight artificial bee colony algorithm. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/11/395.

Chen:2022:DID

[1996] Jong-Shin Chen, Ruo-Wei Hung, Fatemeh Keshavarz-Kohjerdi, and Yung-Fa Huang. Domination and independent domination in extended Supergrid graphs. Algorithms (Basel), 15(11):??, November 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/11/402.

Corsini:2022:HHS

[1997] Roberto Rosario Corsini, Antonio Costa, Sergio Fichera, and Vincenzo Parrinello. Hybrid harmony search for stochastic scheduling of chemotherapy outpatient appointments. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/11/424.

Costa:2022:SAM

[1998] Igor Pinheiro de Araújo Costa, Adilson Vilarinho Terra, Miguel Ângelo Lellis Moreira, Maria Teresa Pereira, Luiz Paulo Lopes Fávero, Marcos dos Santos, and Carlos Francisco Simões Gomes. A systematic approach to the management of military human resources through the ELECTRE-MOr multicriteria method. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/11/422.

Crasmareanu:2022:HOC

[1999] Mircea Crasmareanu. Higher-order curvatures of plane and space parametrized curves. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/11/436.

Diaf:2022:TSJ

[2000] Sami Diaf and Ulrich Fritsche. Topic scaling: a joint document scalingtopic model approach to learn time-specific topics. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/11/430.

Divyanth:2022:IIT

[2001] L. G. Divyanth, D. S. Guru, Peeyush Soni, Rajendra Machavaram, Mohammad Nadimi, and Jitendra Paliwal. Image-to-image translationbased data augmentation for improving crop/weed classification models for precision agriculture applications. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/11/401.

Fernau:2022:SIS

[2002] Henning Fernau. Special issue “Selected Algorithmic Papers From CSR 2020”. Algorithms (Basel), 15(11):??, November 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/426.

Galanis:2022:CNN

[2003] Nikolaos-Ioannis Galanis, Panagiotis Vafiadis, Kostas-Gkouram Mirzaev, and George A. Papakostas. Convolutional neural networks: a roundup and benchmark of their pooling layer variants. Algorithms (Basel), 15 (11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/11/391.

Garg:2022:PTS

[2004] Lalit Garg, Sally McClean, Brian Meenan, Maria Barton, Ken Fullerton, Sandra C. Buttigieg, and Alexander Micallef. Phase-type survival trees to model a delayed discharge and its effect in a stroke care unit. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/414.

Grignaffini:2022:MLA

[2005] Flavia Grignaffini, Francesco Barbuto, Lorenzo Piazzo, Maurizio Troiano, Patrizio Simeoni, Fabio Mangini, Giovanni Pellacani, Carmen Cantisani, and Fabrizio Frezza. Machine learning approaches for skin cancer classification from dermoscopic images: a systematic review. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
11/438.

Hamza:2022:SIP

[2006] Rafik Hamza and Hilmil Pradana. A survey of intellectual property rights protection in big data applications. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/11/418.

Houdayer:2022:SEZ

[2007] Jérôme Houdayer and Patrice Koehl. Stable evaluation of 3D Zernike moments for surface meshes. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/11/406.

Karimi:2022:BPA

[2008] Leila Karimi and Chowdhury Nawrin Ferdous. Branch and price algorithm for multi-trip vehicle routing with a variable number of wagons and time windows. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/412.

La:2022:LBT

[2009] Tuan-Vinh La, Minh-Son Dao, Duy-Dong Le, Kim-Phung Thai, Quoc-Hung Nguyen, and Thuy-Kieu Phan-Thi. Leverage boosting and transformer on text-image matching for cheap fakes detection. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/423.

Lanrezac:2022:FIP

[2010] André Lanrezac, Benoist Laurent, Hubert Santuz, Nicolas Férey, and Marc Baaden. Fast and interactive positioning of proteins within membranes. Algorithms (Basel), 15(11):??, November 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/415.

Lapid:2022:EGF

[2011] Raz Lapid, Zvika Haramaty, and Moshe Sipper. An evolutionary, gradient-free, query-efficient, black-box algorithm for generating adversarial instances in deep convolutional neural networks. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/407.

Lari:2022:DHT

[2012] Salman Lari, Sang Wook Han, Jong Uk Kim, and Hyock Ju Kwon. Design of HIFU treatment plans using thermodynamic equilibrium algorithm. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
11/399.

Le:2022:IMM

[2013] Duy-Dong Le, Anh-Khoa Tran, Minh-Son Dao, Kieu-Chinh Nguyen-Ly, Hoang-Son Le, Xuan-Dao Nguyen-Thi, Thanh-Qui Pham, Van-Luong Nguyen, and Bach-Yen Nguyen-Thi. Insights into multi-model federated learning: an advanced approach for air quality index forecasting. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
11/434.

Li:2022:AEG

[2014] Jiamu Li, Ji Zhang, Mohamed Jaward Bah, Jian Wang, Youwen Zhu, Gaoming Yang, Lingling Li, and Kexin Zhang. An auto-encoder with genetic algorithm for high dimensional data: Towards accurate and interpretable outlier detection. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/11/429.

Li:2022:HOF

[2015] Zhenglong Li and Vincent Tam. A hybrid optimization framework with dynamic transition scheme for large-scale portfolio management. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
11/404.

Mazzetti:2022:PNP

[2016] Piero Mazzetti and Anna Carbone. Periodic and non-periodic brainwaves emerging via stochastic synchronization of closed loops of firing neurons. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
11/396.

Mortaheb:2022:PFM

[2017] Matin Mortaheb, Cemil Vahapoglu, and Sennur Ulukus. Personalized federated multi-task learning over wireless fading channels. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/421.

Mosquera:2022:AGD

[2018] Federico Mosquera, Pieter Smet, and Greet Vanden Berghe. An algorithm for generating a diverse set of multi-modal journeys. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/416.

Motsa:2022:OGB

[2019] Sandile Motsa. Overlapping grid-based optimized single-step hybrid block method for solving first-order initial value problems. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/427.

Nagy:2022:CCP

[2020] Ádám Nagy, János Majár, and Endre Kovács. Consistency and convergence properties of 20 recent and old numerical schemes for the diffusion equation. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/425.

Namdari:2022:LIB

[2021] Alireza Namdari, Maryam Asad Samani, and Tariq S. Durrani. Lithiumion battery prognostics through reinforcement learning based on entropy measures. Algorithms (Basel), 15(11):??, November 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/393.

Nielsen:2022:GAD

[2022] Frank Nielsen. Generalizing the alpha-divergences and the oriented Kullback–Leibler divergences with quasi-arithmetic means. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/435.

Ogundokun:2022:HIS

[2023] Roseline Oluwaseun Ogundokun, Rytis Maskeliunas, Sanjay Misra, and Robertas Damasevicius. Hybrid InceptionV3-SVM-based approach for human posture detection in health monitoring systems. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/410.

Olteanu:2022:UPM

[2024] Alexandru-Liviu Olteanu, Marc Sevaux, and Mohsen Ziaee. Unrelated parallel machine scheduling with job and machine acceptance and renewable resource allocation. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/11/433.

Paredes-Astudillo:2022:TSP

[2025] Yenny Alexandra Paredes-Astudillo, Jairo R. Montoya-Torres, and Valérie Botta-Genoulaz. Taxonomy of scheduling problems with learning and deterioration effects. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/11/439.

Pugliese:2022:AAO

[2026] Giulia Pugliese, Xiaochen Chou, Dominic Loske, Matthias Klumpp, and Roberto Montemanni. AMR-assisted order picking: Models for pickerto-parts systems in a two-blocks warehouse. Algorithms (Basel), 15(11): ??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/11/413.

Rahbari:2022:AMH

[2027] Dadmehr Rahbari. Analyzing meta-heuristic algorithms for task scheduling in a fog-based IoT application. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/15/11/397.

Rivera:2022:SIA

[2028] Sergio R. Rivera. Special issue on algorithms in planning and operation of power systems. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/408.

Ruppert:2022:POB

[2029] Jean Ruppert, Marharyta Aleksandrova, and Thomas Engel. k-Pareto optimality-based sorting with maximization of choice and its application to genetic optimization. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/11/420.

Sangaiah:2022:AFD

[2030] Arun Kumar Sangaiah, Samira Rezaei, Amir Javadpour, Farimasadat Miri, Weizhe Zhang, and Desheng Wang. Automatic fault detection and diagnosis in cellular networks and beyond 5G: Intelligent network management. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/432.

Santos:2022:CPE

[2031] José Victor Sá Santos and Napoleão Nepomuceno. Computational performance evaluation of column generation and generate-and-solve techniques for the one-dimensional cutting stock problem. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/394.

Sarabia:2022:IDS

[2032] Daniel Sarabia, María Cruz Ortiz, and Luis Antonio Sarabia. Integrated design of a supermarket refrigeration system by means of experimental design adapted to computational problems. Algorithms (Basel), 15(11): ??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/11/417.

Soldatov:2022:DLC

[2033] Sergey A. Soldatov, Danil M. Pashkov, Sergey A. Guda, Nikolay S. Karnaukhov, Alexander A. Guda, and Alexander V. Soldatov. Deep learning classification of colorectal lesions based on whole slide images. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/398.

Swain:2022:DLM

[2034] Debabrata Swain, Santosh Satapathy, Biswaranjan Acharya, Madhu Shukla, Vassilis C. Gerogiannis, Andreas Kanavos, and Dimitris Giakovis. Deep learning models for yoga pose monitoring. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/11/403.

Utkin:2022:ERS

[2035] Lev Utkin and Andrei Konstantinov. Ensembles of random SHAPs. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
11/431.

Werner:2022:SIS

[2036] Frank Werner. Special issue “1st Online Conference on Algorithms (IOCA2021)”. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/11/411.

Ye:2022:TAM

[2037] Na Ye, Cai Dai, and Xingsi Xue. A two-archive many-objective optimization algorithm based on d-domination and decomposition. Algorithms (Basel), 15(11):??, November 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/11/392.

Zammarchi:2022:ISG

[2038] Gianpaolo Zammarchi, Claudio Conversano, and Claudia Pisanu. Investigating shared genetic bases between psychiatric disorders, cardiometabolic and sleep traits using K-means clustering and local genetic correlation analysis. Algorithms (Basel), 15(11):??, November 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/11/409.

Ahadi:2022:MSA

[2039] Azam Ahadi, Zahra Eidinejad, Reza Saadati, and Donal O’Regan. Modular stability analysis of a nonlinear stochastic fractional Volterra IDE. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/459.

Apostolopoulos:2022:DLA

[2040] Dimitris J. Apostolopoulos, Ioannis D. Apostolopoulos, Nikolaos D. Papathanasiou, Trifon Spyridonidis, and George S. Panayiotakis. Detection and localisation of abnormal parathyroid glands: an explainable deep learning approach. Algorithms (Basel), 15(12):??, December 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/12/455.

Caramia:2022:SAM

[2041] Massimiliano Caramia. “Algorithms in multi-objective optimization”: Foreword by the Guest Editor. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/12/476.

Chavdarov:2022:ADT

[2042] Ivan Chavdarov and Bozhidar Naydenov. Algorithm for determining the types of inverse kinematics solutions for sequential planar robots and their representation in the configuration space. Algorithms (Basel), 15 (12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/12/469.

Connie:2022:PBG

[2043] Tee Connie, Timilehin B. Aderinola, Thian Song Ong, Michael Kah Ong Goh, Bayu Erfianto, and Bedy Purnama. Pose-based gait analysis for diagnosis of Parkinson’s disease. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/12/474.

Dong:2022:EIW

[2044] Yinhuan Dong, Duanxu Shi, Tughrul Arslan, and Yunjie Yang. Error investigation on Wi-Fi RTT in commercial consumer devices. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/12/464.

Dubnov:2022:PQS

[2045] Shlomo Dubnov. Predictive quantization and symbolic dynamics. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/484.

Ferretti:2022:DNT

[2046] Claudio Ferretti and Martina Saletta. Do neural transformers learn human-defined concepts? An extensive study in source code processing domain. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/449.

Gluncic:2022:GRM

[2047] Matko Gluncić, Ines Vlahović, Leo Mrsić, and Vladimir Paar. Global repeat map (GRM) application: Finding all DNA tandem repeat units. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/458.

Gobert:2022:BAP

[2048] Maxime Gobert, Jan Gmys, Jean-François Toubeau, Nouredine Melab, Daniel Tuyttens, and François Vallée. Batch acquisition for parallel Bayesian optimization-application to hydro-energy storage systems scheduling. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/446.

Goez:2022:MDQ

[2049] David Góez, Paola Soto, Steven Latré, Natalia Gaviria, and Miguel Camelo. A methodology to design quantized deep neural networks for automatic modulation recognition. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/441.

Jiang:2022:EGA

[2050] Yu Jiang, Changyu Qian, Jie Yu, Luyao Zhou, Zheng Wang, Qian Chen, Yang Wang, and Xiaole Ma. Evolutionary game analysis of power generation groups considering energy price fluctuation. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/12/456.

Khanal:2022:RCV

[2051] Salik Ram Khanal, Dennis Paulino, Jaime Sampaio, Joao Barroso, Arsénio Reis, and Vitor Filipe. A review on computer vision technology for physical exercise monitoring. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/444.

Kowaluk:2022:MDM

[2052] Mirosław Kowaluk and Andrzej Lingas. A multi-dimensional matrix product — a natural tool for parameterized graph algorithms. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/12/448.

Kyurkchiev:2022:HMS

[2053] Nikolay Kyurkchiev and Anton Iliev. On a hypothetical model with second kind Chebyshev’s polynomial-correction: Type of limit cycles, simulations, and possible applications. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/462.

Lang:2022:MBL

[2054] Dapeng Lang, Deyun Chen, Jinjie Huang, and Sizhao Li. A momentumbased local face adversarial example generation algorithm. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/12/465.

Lawrynczuk:2022:SIS

[2055] Maciej Lawryńczuk. Special issue “Model Predictive Control: Algorithms and Applications”: Foreword by the Guest Editor. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/12/452.

Lu:2022:BDP

[2056] Jialiang Lu, Xiaoyu Du, Huiping Li, and Zhijie Han. BCCC disjoint path construction algorithm and fault-tolerant routing algorithm under restricted connectivity. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/12/481.

Lu:2022:KDK

[2057] Jialiang Lu, Wei Zhao, and Jie Li. The t/k-diagnosability and a t/k diagnosis algorithm of the data center network BCCC under the MM* model. Algorithms (Basel), 15(12):??, December 2022. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/480.

Madani:2022:BPC

[2058] Kaveh Madani, Faraz Farhidi, and Sona Gholizadeh. Bargaining power in cooperative resource allocations games. Algorithms (Basel), 15(12): ??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/12/445.

Niculescu:2022:RCT

[2059] Mihai Alexandru Niculescu, Stefan Ruseti, and Mihai Dascalu. RoSummary: Control tokens for Romanian news summarization. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/12/472.

Okuboyejo:2022:CSL

[2060] Damilola A. Okuboyejo and Oludayo O. Olugbara. Classification of skin lesions using weighted majority voting ensemble deep learning. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/443.

Pascoe:2022:SAD

[2061] Sean Pascoe. A simplified algorithm for dealing with inconsistencies using the analytic hierarchy process. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/12/442.

Patel:2022:ISD

[2062] Krishna Patel, Chintan Bhatt, and Pier Luigi Mazzeo. Improved ship detection algorithm from satellite images using YOLOv7 and graph neural network. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/473.

Pesterev:2022:NIS

[2063] Dmitriy Pesterev, Olga Druzhina, Alexander Pchelintsev, Erivelton Nepomuceno, and Denis Butusov. Numerical integration schemes based on composition of adjoint multistep methods. Algorithms (Basel), 15(12): ??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/12/463.

Pierantozzi:2022:TCL

[2064] Mariano Pierantozzi, Sebastiano Tomassetti, and Giovanni Di Nicola. Thermal conductivity of low-GWP refrigerants modeling with multiobject optimization. Algorithms (Basel), 15(12):??, December 2022. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/15/12/482.

Policriti:2022:SIA

[2065] Alberto Policriti and Nicola Prezza. Special issue on algorithms and datastructures for compressed computation. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/457.

Pratap:2022:SAD

[2066] Ayush Pratap, Neha Sardana, Sapdo Utomo, John Ayeelyan, P. Karthikeyan, and Pao-Ann Hsiung. A synergic approach of deep learning towards digital additive manufacturing: a review. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/466.

Shi:2022:PPT

[2067] Chenrui Shi, Nobuo Funabiki, Yuanzhi Huo, Mustika Mentari, Kohei Suga, and Takashi Toshida. A proposal of printed table digitization algorithm with image processing. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/12/471.

Sofronov:2022:SIA

[2068] Georgy Sofronov. Special issue on algorithms for sequential analysis. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/470.

Sonobe:2022:ESE

[2069] Tomohiro Sonobe. An experimental survey of extended resolution effects for SAT solvers on the pigeonhole principle. Algorithms (Basel), 15(12): ??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/12/479.

Strappa:2022:ESS

[2070] Jan Strappa, Paola Caymes-Scutari, and Germán Bianchini. Evolutionary statistical system based on novelty search: a parallel metaheuristic for uncertainty reduction applied to wildfire spread prediction. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/478.

Sulemane:2022:VGD

[2071] Shazia Sulemane, João P. Matos-Carvalho, Dário Pedro, Filipe Moutinho, and Sérgio D. Correia. Vineyard gap detection by convolutional neural networks fed by multi-spectral images. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/12/440.

Szurke:2022:CFI

[2072] Szabolcs Kocsis Szürke, Gergő Sütheö, Antal Apagyi, István Lakatos, and Szabolcs Fischer. Cell fault identification and localization procedure for lithium-ion battery system of electric vehicles based on real measurement data. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/467.

Uglev:2022:ECM

[2073] Viktor Uglev and Oleg Sychev. Evaluation, comparison and monitoring of multiparameter systems by unified graphic visualization of activity method on the example of learning process. Algorithms (Basel), 15(12): ??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/15/12/468.

Usama:2022:PIN

[2074] Muhammad Usama, Rui Ma, Jason Hart, and Mikaela Wojcik. Physicsinformed neural networks (PINNs)-based traffic state estimation: an application to traffic network. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/15/12/447.

Vakhnin:2022:NSA

[2075] Aleksei Vakhnin and Evgenii Sopov. A novel self-adaptive cooperative coevolution algorithm for solving continuous large-scale global optimization problems. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/451.

Valouxis:2022:HEL

[2076] Christos Valouxis, Christos Gogos, Angelos Dimitsas, Petros Potikas, and Anastasios Vittas. A hybrid exact-local search approach for one-machine scheduling with time-dependent capacity. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/450.

Valsan:2022:DFS

[2077] Calin Vâlsan, Elena Druica, and Eric Eisenstat. On deep-fake stock prices and why investor behavior might not matter. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/12/475.

Xu:2022:MBA

[2078] Weizhen Xu, Chenyi Zhang, Fangzhen Zhao, and Liangda Fang. A maskbased adversarial defense scheme. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/15/12/461.

Yang:2022:PLF

[2079] Bin Yang, Muhammad Haseeb Arshad, and Qing Zhao. Packet-level and flow-level network intrusion detection based on reinforcement learning and adversarial training. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/15/12/453.

Yoshida:2022:SSR

[2080] Ruriko Yoshida, Lillian Paul, and Peter Nesbitt. Stochastic safety radius on UPGMA. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/15/12/483.

Zhang:2022:EAF

[2081] Hongpo Zhang, Hongzhuang Gu, Junli Gao, Peng Lu, Guanhe Chen, and Zongmin Wang. An effective atrial fibrillation detection from short single-lead electrocardiogram recordings using MCNN-BLSTM network. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/15/
12/454.

Zhang:2022:RPP

[2082] Yi Zhang, Yunchuan Hu, Jiakai Lu, and Zhiqiang Shi. Research on path planning of mobile robot based on improved theta* algorithm. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/15/12/477.

Zvonarev:2022:MAP

[2083] Nikita Zvonarev and Nina Golyandina. Mixed alternating projections with application to Hankel low-rank approximation. Algorithms (Basel), 15(12):??, December 2022. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/15/12/460.

Chimienti:2023:MMI

[2084] Michela Chimienti, Ivan Danzi, Donato Impedovo, Giuseppe Pirlo, Gianfranco Semeraro, and Davide Veneto. MIRROR: Methodological innovation to remodel the electric loads to reduce economic OR environmental impact of user. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/1.

Argyros:2023:SLC

[2085] Ioannis K. Argyros, Stepan Shakhno, Samundra Regmi, and Halyna Yarmola. On the semi-local convergence of two competing sixth order methods for equations in Banach space. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/1/2.

Chesalin:2023:MCS

[2086] Denis D. Chesalin, Andrei P. Razjivin, Alexey S. Dorokhov, and Roman Y. Pishchalnikov. Monte Carlo simulation affects convergence of differential evolution: a case of optical response modeling. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/1/3.

Alangot:2023:DIA

[2087] Bithin Alangot, Pawel Szalachowski, Tien Tuan Anh Dinh, Souhail Meftah, Jeff Ivanos Gana, Khin Mi Mi Aung, and Zengpeng Li. Decentralized identity authentication with auditability and privacy. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/1/4.

Shkaberina:2023:VAC

[2088] Guzel Shkaberina, Natalia Rezova, Elena Tovbis, and Lev Kazakovtsev. Visual assessment of cluster tendency with variations of distance measures. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/1/5.

Cruz:2023:ADG

[2089] Ahyssa R. Cruz and Walter C. Ermler. Algorithmic design of geometric data for molecular potential energy surfaces. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/1/6.

Han:2023:PDK

[2090] Taihao Han, Sai Akshay Ponduru, Arianit Reka, Jie Huang, Gaurav Sant, and Aditya Kumar. Predicting dissolution kinetics of tricalcium silicate using deep learning and analytical models. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/7.

Lim:2023:SSA

[2091] Seunguook Lim and Jihie Kim. SAPBERT: Speaker-aware pretrained BERT for emotion recognition in conversation. Algorithms (Basel), 16 (1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/8.

Qi:2023:RDD

[2092] Ruosen Qi, Jie Zhang, and Katy Spencer. A review on data-driven condition monitoring of industrial equipment. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/1/9.

El-Mesady:2023:BCG

[2093] Ahmed El-Mesady, Aleksandr Y. Romanov, Aleksandr A. Amerikanov, and Alexander D. Ivannikov. On bipartite circulant graph decompositions based on Cartesian and tensor products with novel topologies and deadlock-free routing. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/10.

Kaur:2023:OAL

[2094] Goldendeep Kaur, Kiran Jyoti, Nitin Mittal, Vikas Mittal, and Rohit Salgotra. Optimized approach for localization of sensor nodes in 2D wireless sensor networks using modified learning enthusiasm-based teaching-learning-based optimization algorithm. Algorithms (Basel), 16 (1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/11.

Lu:2023:IBL

[2095] Xinxi Lu, Lijuan Yuan, Ruifeng Li, Zhihuan Xing, Ning Yao, and Yichun Yu. An improved bi-LSTM-based missing value imputation approach for pregnancy examination data. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/12.

Fedotova:2023:DAA

[2096] Anastasia Fedotova, Aleksandr Romanov, Anna Kurtukova, and Alexander Shelupanov. Digital authorship attribution in Russian-language fanfiction and classical literature. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/13.

Lai:2023:RRM

[2097] Shumin Lai, Longjun Huang, Ping Li, Zhenzhen Luo, Jianzhong Wang, and Yugen Yi. RMFRASL: Robust matrix factorization with robust adaptive structure learning for feature selection. Algorithms (Basel), 16 (1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/14.

Matrenin:2023:IAC

[2098] Pavel V. Matrenin. Improvement of ant colony algorithm performance for the job-shop scheduling problem using evolutionary adaptation and software realization heuristics. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/15.

Goyal:2023:GCA

[2099] Ravi Goyal and Victor De Gruttola. A general computational approach for counting labeled graphs. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/16.

Soper:2023:HOU

[2100] Daniel S. Soper. Hyperparameter optimization using successive halving with greedy cross validation. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/17.

Trouli:2023:SSS

[2101] Georgia Eirini Trouli, Alexandros Pappas, Georgia Troullinou, Lefteris Koumakis, Nikos Papadakis, and Haridimos Kondylakis. SumMER: Structural summarization for RDF/S KGs. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/1/18.

Utkin:2023:IAD

[2102] Lev Utkin, Andrey Ageev, Andrei Konstantinov, and Vladimir Muliukha. Improved anomaly detection by using the attention-based isolation forest. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/
19.

Nikolaidis:2023:VBA

[2103] Pavlos Nikolaidis. Variational Bayes to accelerate the Lagrange multipliers towards the dual optimization of reliability and cost in renewable energy systems. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/20.

Boucher:2023:SIA

[2104] Christina Boucher. Special issue: Algorithms in bioinformatics. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/21.

Papei:2023:SLC

[2105] Hadi Papei and Yang Li. Stochastic local community detection in networks. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/1/22.

Tran:2023:OAT

[2106] Minh Tran, Duc Pham-Hi, and Marc Bui. Optimizing automated trading systems with deep reinforcement learning. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/1/23.

Choubineh:2023:FNO

[2107] Abouzar Choubineh, Jie Chen, David A. Wood, Frans Coenen, and Fei Ma. Fourier neural operator for fluid flow in small-shape 2D simulated porous media dataset. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/24.

Arasli:2023:DIS

[2108] Batuhan Arasli and Sennur Ulukus. Dynamic infection spread model based group testing. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/25.

Manolis:2023:NTT

[2109] George D. Manolis and Georgios I. Dadoulis. On the numerical treatment of the temporal discontinuity arising from a time-varying point mass attachment on a waveguide. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/26.

Smietanski:2023:SIV

[2110] Marek J. Śmietański. Some inexact version of the iterative exponential method for solving nonsmooth equations. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/1/27.

Chugh:2023:ADS

[2111] Deepesh Chugh, Himanshu Mittal, Amit Saxena, Ritu Chauhan, Eiad Yafi, and Mukesh Prasad. Augmentation of densest subgraph finding unsupervised feature selection using shared nearest neighbor clustering. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/28.

Garcia-Pineda:2023:ORP

[2112] Laura Patricia García-Pineda and Oscar Danilo Montoya. Optimal reactive power compensation via D-STATCOMs in electrical distribution systems by applying the generalized normal distribution optimizer. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/29.

Nguyen:2023:TPI

[2113] Phuc-Thinh Nguyen, Minh-Son Dao, Michael A. Riegler, Rage Uday Kiran, Thai-Thinh Dang, Duy-Dong Le, Kieu-Chinh Nguyen-Ly, Thanh-Qui Pham, and Van-Luong Nguyen. Training performance indications for amateur athletes based on nutrition and activity lifelogs. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/1/30.

Yang:2023:OLQ

[2114] Wenxin Yang, Xiaoli Zhi, and Weiqin Tong. Optimization of linear quantization for general and effective low bit-width network compression. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/31.

Poh:2023:DRL

[2115] Li Zhe Poh, Tee Connie, Thian Song Ong, and Michael Kah Ong Goh. Deep reinforcement learning-based dynamic pricing for parking solutions. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/32.

Al-Nuaimi:2023:SIB

[2116] Bashar Talib Al-Nuaimi, H. K. Al-Mahdawi, Zainalabideen Albadran, Hussein Alkattan, Mostafa Abotaleb, and El-Sayed M. El-kenawy. Solving of the inverse boundary value problem for the heat conduction equation in two intervals of time. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/33.

Platt:2023:SHC

[2117] Moritz Platt and Peter McBurney. Sybil in the haystack: a comprehensive review of blockchain consensus mechanisms in search of strong Sybil attack resistance. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/34.

Barbosa:2023:SIA

[2118] Ramiro S. Barbosa and Isabel S. Jesus. Special issue on algorithms for PID controllers 2021. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/35.

Providas:2023:SMS

[2119] Efthimios Providas and Ioannis Nestorios Parasidis. A symbolic method for solving a class of convolution-type Volterra–Fredholm–Hammerstein integro-differential equations under nonlocal boundary conditions. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/36.

Pereira:2023:ABM

[2120] Sílvia de Castro Pereira, Eduardo J. Solteiro Pires, and Paulo B. de Moura Oliveira. Ant-balanced multiple traveling salesmen: ACO-BmTSP. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/1/37.

Morais:2023:NDL

[2121] Rui Morais, Paul Crocker, and Valderi Leithardt. Nero: a deterministic leaderless consensus algorithm for DAG-based cryptocurrencies. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/38.

Rogachev:2023:FAM

[2122] Alexey F. Rogachev, Alexey B. Simonov, Natalia V. Ketko, and Natalia N. Skiter. Fuzzy algorithmic modeling of economics and innovation process dynamics based on preliminary component allocation by singular spectrum analysis method. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/39.

Montemanni:2023:SPD

[2123] Roberto Montemanni and Mauro Dell’Amico. Solving the parallel drone scheduling traveling salesman problem via constraint programming. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/40.

Ryu:2023:ESS

[2124] Hee Jun Ryu, Ye Na Jo, Won Jun Lee, Ji Won Cheong, Boo Yong Moon, and Young Dae Ko. An effective staff scheduling for shift workers in social welfare facilities for the disabled. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/1/41.

Naulia:2023:PWK

[2125] Pradeep S. Naulia, Arunava Roy, Junzo Watada, and Izzatdin B. A. Aziz. A pixel-wise k-immediate neighbour-based image analysis approach for identifying rock pores and fractures from grayscale image samples. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
1/42.

Rixen:2023:DBA

[2126] Jöran Rixen, Steffen Leonhardt, Jochen Moll, Duy Hai Nguyen, and Chuong Ngo. The D-Bar algorithm fusing electrical impedance tomography with a priori radar data: a hands-on analysis. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/43.

Cariou:2023:EAG

[2127] Christophe Cariou, Laure Moiroux-Arvis, François Pinet, and Jean-Pierre Chanet. Evolutionary algorithm with geometrical heuristics for solving the close enough traveling salesman problem: Application to the trajectory planning of an unmanned aerial vehicle. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/44.

Huo:2023:TDA

[2128] Yuanzhi Huo, Pradini Puspitaningayu, Nobuo Funabiki, Kazushi Hamazaki, Minoru Kuribayashi, Yihan Zhao, and Kazuyuki Kojima. Three diverse applications of general-purpose parameter optimization algorithm. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/1/45.

Bruni:2023:HBB

[2129] Renato Bruni, Gianpiero Bianchi, and Pasquale Papa. Hyperparameter black-box optimization to improve the automatic classification of support tickets. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/1/46.

Vitetta:2023:IMP

[2130] Antonino Vitetta. The importance of modeling path choice behavior in the vehicle routing problem. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/47.

Milic:2023:ASZ

[2131] Vladimir Milić, Josip Kasać, and Marin Lukas. An algorithm for solving zero-sum differential game related to the nonlinear H_{∞} control problem. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/48.

Pintelas:2023:SIE

[2132] Panagiotis Pintelas and Ioannis E. Livieris. Special issue on ensemble learning and/or explainability. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/49.

Maazioui:2023:NSV

[2133] Souhail Maazioui, Imad Kissami, Fayssal Benkhaldoun, and Driss Ouazar. Numerical study of viscoplastic flows using a multigrid initialization algorithm. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/50.

Yang:2023:DCF

[2134] Kai-Fu Yang, Yi-Ru Lin, and Long-Sheng Chen. Discovering critical factors in the content of crowdfunding projects. Algorithms (Basel), 16 (1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/51.

Narkhede:2023:NLD

[2135] Gaurav Narkhede, Anil Hiwale, Bharat Tidke, and Chetan Khadse. Novel MIA-LSTM deep learning hybrid model with data preprocessing for forecasting of PM_{2.5}. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/52.

Niazi:2023:IND

[2136] Tahira Niazi, Teerath Das, Ghufran Ahmed, Syed Muhammad Waqas, Sumra Khan, Suleman Khan, Ahmed Abdelaziz Abdelatif, and Shaukat Wasi. Investigating novice developers’ code commenting trends using machine learning techniques. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/1/53.

Wari:2023:DPO

[2137] Ezra Wari, Weihang Zhu, and Gino Lim. A discrete partially observable Markov decision process model for the maintenance optimization of oil and gas pipelines. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/54.

Robenack:2023:SPC

[2138] Klaus Röbenack and Stefan Palis. Set-point control of a spatially distributed Buck converter. Algorithms (Basel), 16(1), January 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/1/55.

Pantale:2023:DIA

[2139] Olivier Pantalé. Development and implementation of an ANN based flow law for numerical simulations of thermo-mechanical processes at high temperatures in FEM software. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/1/56.

Elkhovskaya:2023:EPD

[2140] Liubov O. Elkhovskaya, Alexander D. Kshenin, Marina A. Balakhontceva, Mikhail V. Ionov, and Sergey V. Kovalchuk. Extending process discovery with model complexity optimization and cyclic states identification: Application to healthcare processes. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/1/57.

Gamidullaeva:2023:DCT

[2141] Leyla Gamidullaeva, Alexey Finogeev, Mikhail Kataev, and Larisa Bulysheva. A design concept for a tourism recommender system for regional development. Algorithms (Basel), 16(1), January 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/1/58.

Tang:2023:DAM

[2142] Huidong Tang, Sayaka Kamei, and Yasuhiko Morimoto. Data augmentation methods for enhancing robustness in text classification tasks. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/1/59.

Kojima:2023:IIR

[2143] Takeshi Kojima and Tetsuya Yoshinaga. Iterative image reconstruction algorithm with parameter estimation by neural network for computed tomography. Algorithms (Basel), 16(1), January 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/1/60.

Albertin:2023:RTN

[2144] Umberto Albertin, Giuseppe Pedone, Matilde Brossa, Giovanni Squillero, and Marcello Chiaberge. A real-time novelty recognition framework based on machine learning for fault detection. Algorithms (Basel), 16 (2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/2/61.

Erwin:2023:MGS

[2145] Kyle Erwin and Andries Engelbrecht. Multi-guide set-based particle swarm optimization for multi-objective portfolio optimization. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/2/62.

Office:2023:ARA

[2146] Algorithms Editorial Office. Acknowledgment to the reviewers of Algorithms in 2022. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/63.

Stanimirovic:2023:IGD

[2147] Predrag S. Stanimirović, Bilall I. Shaini, Jamilu Sabi’u, Abdullah Shah, Milena J. Petrović, Branislav Ivanov, Xinwei Cao, Alena Stupina, and Shuai Li. Improved gradient descent iterations for solving systems of nonlinear equations. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/64.

Ioannou:2023:OBS

[2148] George Ioannou, Georgios Alexandridis, and Andreas Stafylopatis. Online batch selection for enhanced generalization in imbalanced datasets. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/65.

Atenco:2023:ABN

[2149] Juan Carlos Atenco, Juan Carlos Moreno, and Juan Manuel Ramirez. Audiovisual biometric network with deep feature fusion for identification and text prompted verification. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/2/66.

Lu:2023:CSD

[2150] Jinming Lu and Ewald A. Werner. Cathode shape design for steadystate electrochemical machining. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/2/67.

Ruiz-Serra:2023:IRL

[2151] Jaime Ruiz-Serra and Michael S. Harré. Inverse reinforcement learning as the algorithmic basis for theory of mind: Current methods and open problems. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/68.

Vyas:2023:RCF

[2152] Piyush Vyas, Gitika Vyas, and Gaurav Dhiman. RUemo — the classification framework for Russia–Ukraine war-related societal emotions on Twitter through machine learning. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/2/69.

Kim:2023:MNH

[2153] Leeseok Kim and José Luis Palacios. On the moments of the number of hires in the assistant hiring algorithm. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/2/70.

Tsoulos:2023:LPR

[2154] Ioannis G. Tsoulos and Vasileios Charilogis. Locating the parameters of RBF networks using a hybrid particle swarm optimization method. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/71.

Chen:2023:MFS

[2155] Ren-Raw Chen, Cameron D. Miller, and Puay Khoon Toh. Modeling firm search and innovation trajectory using swarm intelligence. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/72.

Moreno:2023:LOE

[2156] Higuatzi Moreno and Alexander Schaum. Low-order electrochemical state estimation for Li-ion batteries. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/2/73.

Zarringhalam:2023:COI

[2157] Amir Zarringhalam, Saeed Shiry Ghidary, Ali Mohades, and Seyed-Ali Sadegh-Zadeh. CUDA and OpenMp implementation of Boolean matrix product with applications in visual SLAM. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/2/74.

ElKashlan:2023:IDE

[2158] Mohamed ElKashlan, Heba Aslan, Mahmoud Said Elsayed, Anca D. Jurcut, and Marianne A. Azer. Intrusion detection for electric vehicle charging systems (EVCS). Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/75.

Greco:2023:MLE

[2159] Massimiliano Greco, Pier Francesco Caruso, Sofia Spano, Gianluigi Citterio, Antonio Desai, Alberto Molteni, Romina Aceto, Elena Costantini, Antonio Voza, and Maurizio Cecconi. Machine learning for early outcome prediction in septic patients in the emergency department. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/76.

Pavliuk:2023:TLA

[2160] Olena Pavliuk, Myroslav Mishchuk, and Christine Strauss. Transfer learning approach for human activity recognition based on continuous wavelet transform. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/77.

Carvalho:2023:PEN

[2161] Inês Carvalho, Filipe Sá, and Jorge Bernardino. Performance evaluation of NoSQL document databases: Couchbase, CouchDB, and MongoDB. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/78.

Guidi:2023:EHS

[2162] Tommaso Guidi, Lorenzo Python, Matteo Forasassi, Costanza Cucci, Massimiliano Franci, Fabrizio Argenti, and Andrea Barucci. Egyptian hieroglyphs segmentation with convolutional neural networks. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/79.

Windisch:2023:PAC

[2163] Dominic Windisch, Christian Kaever, Guido Juckeland, and André Bieberle. Parallel algorithm for connected-component analysis using CUDA. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/80.

Lopez:2023:DAS

[2164] Alberto Lopez and Jose Aguilar. A data analysis smart system for the optimal deployment of nanosensors in the context of an eHealth application. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/81.

Chen:2023:PMI

[2165] Zhi Chen, Shuai Zhang, Sally McClean, Fionnuala Hart, Michael Milliken, Brahim Allan, and Ian Kegel. Process mining IPTV customer eye gaze movement using discrete-time Markov chains. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/2/82.

Pervadchuk:2023:MMC

[2166] Vladimir Pervadchuk, Daria Vladimirova, and Anna Derevyankina. Mathematical modeling of capillary drawing stability for hollow optical fibers. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/83.

Hare:2023:APC

[2167] Warren Hare and Gabriel Jarry-Bolduc. About the performance of a calculus-based approach to building model functions in a derivative-free trust-region algorithm. Algorithms (Basel), 16(2), February 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/2/84.

Tushkanova:2023:DCA

[2168] Olga Tushkanova, Diana Levshun, Alexander Branitskiy, Elena Fedorchenko, Evgenia Novikova, and Igor Kotenko. Detection of cyberattacks and anomalies in cyber-physical systems: Approaches, data sources, evaluation. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/85.

Thiha:2023:EOE

[2169] Saw Thiha and Jay Rajasekera. Efficient online engagement analytics algorithm toolkit that can run on edge. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/16/2/86.

Matveev:2023:IIE

[2170] Ivan Matveev and Ilia Safonov. From iris image to embedded code: System of methods. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/87.

Bhatt:2023:EHD

[2171] Chintan M. Bhatt, Parth Patel, Tarang Ghetia, and Pier Luigi Mazzeo. Effective heart disease prediction using machine learning techniques. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/2/88.

Bulut:2023:RGL

[2172] Okan Bulut, Guher Gorgun, Tarid Wongvorachan, and Bin Tan. Rapid guessing in low-stakes assessments: Finding the optimal response time threshold with random search and genetic algorithm. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/2/89.

Andriyanov:2023:UCF

[2173] Nikita Andriyanov. The use of correlation features in the problem of speech recognition. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/90.

Strutz:2023:RWS

[2174] Tilo Strutz. Redesigning the wheel for systematic travelling salesmen. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/91.

Huang:2023:HDS

[2175] Zhongda Huang, Andy Ogilvy, Steve Collins, Warren Hare, Michelle Hilts, and Andrew Jirasek. A hybrid direct search and model-based derivative-free optimization method with dynamic decision processing and application in solid-tank design. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/16/2/92.

Zhang:2023:TMS

[2176] Xuanyu Zhang, Hao Zhou, Ke Yu, Xiaofei Wu, and Anis Yazidi. Tsetlin machine for sentiment analysis and spam review detection in Chinese. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/93.

Papadia:2023:CDT

[2177] Gabriele Papadia, Massimo Pacella, Massimiliano Perrone, and Vincenzo Giliberti. A comparison of different topic modeling methods through a real case study of Italian customer care. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/2/94.

Saberironaghi:2023:DDM

[2178] Alireza Saberironaghi, Jing Ren, and Moustafa El-Gindy. Defect detection methods for industrial products using deep learning techniques: a review. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/95.

Wang:2023:IQA

[2179] Yue Wang and John J. Healy. Image quality assessment for Gibbs ringing reduction. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/96.

Yousefi:2023:NNM

[2180] Aryan Yousefi and Kalpdrum Passi. Nemesis: Neural mean teacher learning-based emotion-centric speaker. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/2/97.

Liang:2023:NIM

[2181] Zhiguo Liang, Lijun Zhang, and Xizhe Wang. A novel intelligent method for fault diagnosis of steam turbines based on T-SNE and XGBoost. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/98.

Tzougas:2023:ELR

[2182] George Tzougas and Konstantin Kutzkov. Enhancing logistic regression using neural networks for classification in actuarial learning. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/99.

Sotskov:2023:APL

[2183] Yuri N. Sotskov. Assembly and production line designing, balancing and scheduling with inaccurate data: a survey and perspectives. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/100.

Agliullin:2023:CAM

[2184] Timur Agliullin, Vladimir Anfinogentov, Oleg Morozov, Airat Sakhabutdinov, Bulat Valeev, Ayna Niyazgulyeva, and Yagmyrguly Garovov. Comparative analysis of the methods for fiber Bragg structures spectrum modeling. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/101.

Sun:2023:MTR

[2185] Yuhe Sun, Zuohua Ding, Hongyun Huang, Senhao Zou, and Mingyue Jiang. Metamorphic testing of relation extraction models. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/102.

Pavkov:2023:LCA

[2186] Tsonyo M. Pavkov, Valentin G. Kabadzhov, Ivan K. Ivanov, and Stoil I. Ivanov. Local convergence analysis of a one parameter family of simultaneous methods with applications to real-world problems. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/103.

Minoccheri:2023:QMD

[2187] Cristian Minoccheri, Olivia Alge, Jonathan Gryak, Kayvan Najarian, and Harm Derksen. Quadratic multilinear discriminant analysis for tensorial data classification. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/104.

Alwidian:2023:UMM

[2188] Sanaa Alwidian, Daniel Amyot, and Yngve Lamo. Union models for model families: Efficient reasoning over space and time. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/2/105.

Amendola:2023:SIL

[2189] Giovanni Amendola. Special issue on logic-based artificial intelligence. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/106.

Oliveira:2023:SSA

[2190] Carlos Oliveira, José Baptista, and Adelaide Cerveira. Self-sustainability assessment for a high building based on linear programming and computational fluid dynamics. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/2/107.

Sufi:2023:ALC

[2191] Fahim Sufi. Algorithms in low-code-no-code for research applications: a practical review. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/108.

Wilbik:2023:TFA

[2192] Anna Wilbik, Irene Vanderfeesten, Dennis Bergmans, Serge Heines, Oktay Turetken, and Walther van Mook. Towards a flexible assessment of compliance with clinical protocols using fuzzy aggregation techniques. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
2/109.

Wang:2023:MPO

[2193] Jinnan Wang, Weiqin Tong, and Xiaoli Zhi. Model parallelism optimization for CNN FPGA accelerator. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/2/110.

Izsak:2023:LDN

[2194] Ferenc Izsák and Taki Eddine Djebbar. Learning data for neural-network-based numerical solution of PDEs: Application to Dirichlet-to-Neumann problems. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/111.

Barletta:2023:VSV

[2195] Vita Santa Barletta, Danilo Caivano, Mirko De Vincentiis, Azzurra Ragone, Michele Scalera, and Manuel Ángel Serrano Martín. V-SOC4AS: a vehicle-SOC for improving automotive security. Algorithms (Basel), 16 (2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/2/112.

Pourdarbani:2023:ELB

[2196] Razieh Pourdarbani, Sajad Sabzi, Mohsen Dehghankar, Mohammad H. Rohban, and Juan I. Arribas. Examination of lemon bruising using different CNN-based classifiers and local spectral-spatial hyperspectral imaging. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/2/
113.

Song:2023:BDO

[2197] Bingyu Song, Yingwu Chen, Qing Yang, Yahui Zuo, Shilong Xu, and Yuning Chen. On-board decentralized observation planning for LEO satellite constellations. Algorithms (Basel), 16(2), February 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/2/114.

Milan:2023:EAL

[2198] Sara Tabaghchi Milan, Mehdi Darbandi, Nima Jafari Navimipour, and Senay Yalcin. An energy-aware load balancing method for IoT-based smart recycling machines using an artificial chemical reaction optimization algorithm. Algorithms (Basel), 16(2), February 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/115.

Brzus:2023:PSN

[2199] Michal Brzus, Kevin Knoernschild, Jessica C. Sieren, and Hans J. Johnson. PigSNIPE: Scalable neuroimaging processing engine for minipig MRI. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/116.

Fang:2023:EBO

[2200] Yihao Fang, Mu Niu, Pokman Cheung, and Lizhen Lin. Extrinsic Bayesian optimization on manifolds. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/2/117.

Liu:2023:EDA

[2201] Qing Liu, Jianjun Hao, and Yijun Guo. EEG data augmentation for emotion recognition with a task-driven GAN. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/2/118.

Smeaton:2023:PIR

[2202] Alan F. Smeaton and Feiyan Hu. Periodicity intensity reveals insights into time series data: Three use cases. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/16/2/119.

Mitov:2023:RPH

[2203] Alexander Mitov, Tsonyo Slavov, and Jordan Kralev. Rapid prototyping of H_{∞} algorithm for real-time displacement volume control of axial piston pumps. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/2/120.

Wu:2023:IVA

[2204] Jinhong Wu, Konstantinos Plataniotis, Lucy Liu, Ehsan Amjadian, and Yuri Lawryshyn. Interpretation for variational autoencoder used to generate financial synthetic tabular data. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/16/2/121.

Li:2023:IBC

[2205] Yang-Rui Li, Chih-Chia Chen, and Chao-Chung Peng. Integral backstepping control algorithm for a quadrotor positioning flight task: a design issue discussion. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/2/122.

Vitorino:2023:ISP

[2206] João Pedro Vitorino, José Simão, Nuno Datia, and Matilde Pato. IRONEDGE: Stream processing architecture for edge applications. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/2/123.

Guan:2023:FNO

[2207] Steven Guan, Ko-Tsung Hsu, and Parag V. Chitnis. Fourier neural operator network for fast photoacoustic wave simulations. Algorithms (Basel), 16(2), February 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/2/124.

Nizovtseva:2023:AMT

[2208] Irina Nizovtseva, Vladimir Palmin, Ivan Simkin, Ilya Starodumov, Pavel Mikushin, Alexander Nozik, Timur Hamitov, Sergey Ivanov, Sergey Vikharev, Alexei Zinovev, Vladislav Svitich, Matvey Mogilev, Margarita Nikishina, Simon Kraev, Stanislav Yurchenko, Timofey Mityashin, Dmitrii Chernushkin, Anna Kalyuzhnaya, and Felix Blyakhman. Assessing the mass transfer coefficient in jet bioreactors with classical computer vision methods and neural networks algorithms. Algorithms (Basel), 16 (3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/3/125.

Li:2023:GGR

[2209] Jianhe Li and Suohai Fan. GRNN: Graph-retraining neural network for semi-supervised node classification. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/126.

Perin:2023:NSF

[2210] Guilherme Perin, Lichao Wu, and Stjepan Picek. The need for speed: a fast guessing entropy calculation for deep learning-based SCA. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/3/127.

Cardone:2023:NCA

[2211] Barbara Cardone and Ferdinando Di Martino. A novel classification algorithm based on multidimensional F^{1} fuzzy transform and PCA feature extraction. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/128.

Santini:2023:APC

[2212] Carlo Santini, Fabio Mangini, and Fabrizio Frezza. Apollonian packing of circles within ellipses. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/3/129.

Aziz:2023:AAE

[2213] Nor Azlina Ab. Aziz, Tawsif K., Sharifah Noor Masidayu Sayed Ismail, Muhammad Anas Hasnul, Kamarulzaman Ab. Aziz, Siti Zainab Ibrahim, Azlan Abd. Aziz, and J. Emerson Raja. Asian Affective and Emotional State (A2ES) dataset of ECG and PPG for affective computing research. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/130.

Koehl:2023:CGW

[2214] Patrice Koehl, Marc Delarue, and Henri Orland. Computing the Gromov–Wasserstein distance between two surface meshes using optimal transport. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/131.

Chen:2023:EGC

[2215] Jong-Shin Chen, Chun-Ming Kuo, and Ruo-Wei Hung. An efficient GNSS coordinate recognition algorithm for epidemic management. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/3/132.

Stanovov:2023:CRS

[2216] Vladimir Stanovov, Lev Kazakovtsev, and Eugene Semenkin. Crossover rate sorting in adaptive differential evolution. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/3/133.

Moldovan:2023:PTA

[2217] Dorin Moldovan. Plum tree algorithm and weighted aggregated ensembles for energy efficiency estimation. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/134.

Lotfi:2023:SID

[2218] Fariba Lotfi, Amin Beheshti, Helia Farhood, Matineh Pooshideh, Mansour Jamzad, and Hamid Beigy. Storytelling with image data: a systematic review and comparative analysis of methods and tools. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/3/135.

Fischer:2023:LLL

[2219] Norbert Fischer, Alexander Hartelt, and Frank Puppe. Line-level layout recognition of historical documents with background knowledge. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/3/136.

Shachar:2023:AEC

[2220] Eran Shachar, Israel Cohen, and Baruch Berdugo. Acoustic echo cancellation with the normalized sign-error least mean squares algorithm and deep residual echo suppression. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/3/137.

Barreto-Parra:2023:OID

[2221] German Francisco Barreto-Parra, Brandon Cortés-Caicedo, and Oscar Danilo Montoya. Optimal integration of D-STATCOMs in radial and meshed distribution networks using a MATLAB–GAMS interface. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/3/138.

Rosso:2023:DPA

[2222] Alessandra Rosso and Ezio Venturino. A dynamic programming approach to ecosystem management. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/139.

Stasinos:2023:TMP

[2223] Nikolaos Stasinos, Anestis Kousis, Vangelis Sarlis, Aristeidis Mystakidis, Dimitris Rousidis, Paraskevas Koukaras, Ioannis Kotsiopoulos, and Christos Tjortjis. A tri-model prediction approach for COVID-19 ICU bed occupancy: a case study. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/3/140.

Frutuoso:2023:EER

[2224] Maria Inês Frutuoso, Horácio C. Neto, Mário P. Véstias, and Rui Policarpo Duarte. Energy-efficient and real-time wearable for wellbeingmonitoring IoT system based on SoC-FPGA. Algorithms (Basel), 16 (3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/3/141.

Moreno-Jimenez:2023:AGL

[2225] Luis-Gil Moreno-Jiménez, Juan-Manuel Torres-Moreno, and Roseli Suzi. Wedemann. Automatic generation of literary sentences in French. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/3/142.

Arsini:2023:NNG

[2226] Lorenzo Arsini, Barbara Caccia, Andrea Ciardiello, Stefano Giagu, and Carlo Mancini Terracciano. Nearest neighbours graph variational AutoEncoder. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/143.

Jacob:2023:ELV

[2227] Ashwin Jacob, Diptapriyo Majumdar, and Venkatesh Raman. Expansion lemma-variations and applications to polynomial-time preprocessing. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/
144.

Antonelli:2023:SISa

[2228] Laura Antonelli and Mario Rosario Guarracino. Special issue on supervised and unsupervised classification algorithms — foreword from Guest Editors. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/145.

Combs:2023:TLA

[2229] Kara Combs, Hongjing Lu, and Trevor J. Bihl. Transfer learning and analogical inference: a critical comparison of algorithms, methods, and applications. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/146.

Chincholi:2023:DLD

[2230] Farheen Chincholi and Harald Koestler. Detectron2 for lesion detection in diabetic retinopathy. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/3/147.

Krutikov:2023:PQT

[2231] Vladimir Krutikov, Elena Tovbis, Anatoly Bykov, Predrag Stanimirovic, Ekaterina Chernova, and Lev Kazakovtsev. Properties of the quadratic transformation of dual variables. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/3/148.

Sarabia:2023:EGM

[2232] Pablo Sarabia, Alvaro Araujo, Luis Antonio Sarabia, and María de la Cruz Ortiz. Electromyography gesture model classifier for fault-tolerant-embedded devices by means of partial least square class modelling error correcting output codes (PLS-ECOC). Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/149.

Tonchev:2023:MLS

[2233] Hristo Tonchev and Petar Danev. A machine learning study of high robustness quantum walk search algorithm with qudit Householder coins. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/150.

Giotopoulos:2023:MAN

[2234] Konstantinos C. Giotopoulos, Dimitrios Michalopoulos, Aristeidis Karras, Christos Karras, and Spyros Sioutas. Modelling and analysis of neuro fuzzy employee ranking system in the public sector. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/3/151.

Alamr:2023:UTB

[2235] Abrar Alamr and Abdelmonim Artoli. Unsupervised transformer-based anomaly detection in ECG signals. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/152.

Cheng:2023:SIS

[2236] Liangliang Cheng, Yunfeng Dou, Jian Zhou, Huabin Wang, and Liang Tao. Speaker-independent spectral enhancement for bone-conducted speech. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/153.

Wong:2023:FCV

[2237] Vivian W. H. Wong and Kincho H. Law. Fusion of CCTV video and spatial information for automated crowd congestion monitoring in public urban spaces. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/154.

Sobhanmanesh:2023:CMT

[2238] Fariborz Sobhanmanesh, Amin Beheshti, Nicholas Nouri, Natalia Monje Chapparo, Sandya Raj, and Richard A. George. A cognitive model for technology adoption. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/3/155.

Kanwar:2023:NTO

[2239] Vinay Kanwar, Alicia Cordero, Juan R. Torregrosa, Mithil Rajput, and Ramandeep Behl. A new third-order family of multiple root-findings based on exponential fitted curve. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/156.

Marochok:2023:AGB

[2240] Stanislav Marochok and Pavol Zajac. Algorithm for generating S-boxes with prescribed differential properties. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/157.

Wang:2023:LDR

[2241] Shuang Wang, Amin Beheshti, Yufei Wang, Jianchao Lu, Quan Z. Sheng, Stephen Elbourn, and Hamid Alinejad-Rokny. Learning distributed representations and deep embedded clustering of texts. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/3/158.

Calimeri:2023:ABD

[2242] Francesco Calimeri, Nicola Leone, Giovanni Melissari, Francesco Pacenza, Simona Perri, Kristian Reale, Francesco Ricca, and Jessica Zangari. ASP-based declarative reasoning in data-intensive enterprise and IoT applications. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/3/159.

Tsai:2023:IDC

[2243] Min-Jen Tsai, Ya-Chu Lee, and Te-Ming Chen. Implementing deep convolutional neural networks for QR code-based printed source identification. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/

deRose:2023:HBS

[2244] Steve de Rose, Philippe Meyer, and Frédéric Bertrand. Human body shapes anomaly detection and classification using persistent homology. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/161.

daSilva:2023:MOD

[2245] Juarez Machado da Silva, Gabriel de Oliveira Ramos, and Jorge Luis Victória Barbosa. Multi-objective decision-making meets dynamic shortest path: Challenges and prospects. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/3/162.

Cordero:2023:CSN

[2246] Alicia Cordero, Javier G. Maimó, Antmel Rodríguez-Cabral, and Juan R. Torregrosa. Convergence and stability of a new parametric class of iterative processes for nonlinear systems. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/163.

Ivancev:2023:CSL

[2247] Ana Cudina Ivancev, Maja Ahac, Sasa Ahac, and Vesna Dragcević. Comparison of single-lane roundabout entry degree of saturation estimations from analytical and regression models. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/3/164.

Moskalenko:2023:RRS

[2248] Viacheslav Moskalenko, Vyacheslav Kharchenko, Alona Moskalenko, and Borys Kuzikov. Resilience and resilient systems of artificial intelligence: Taxonomy, models and methods. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/3/165.

Galatro:2023:FEP

[2249] Daniela Galatro, Rosario Trigo-Ferre, Allana Nakashook-Zettler, Vincenzo Costanzo-Alvarez, Melanie Jeffrey, Maria Jacome, Jason Bazylak, and Cristina H. Amon. Framework for evaluating potential causes of health risk factors using average treatment effect and uplift modelling. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/166.

Piri:2023:LRH

[2250] Jayashree Piri, Puspanjali Mohapatra, Raghunath Dey, Biswaranjan Acharya, Vassilis C. Gerogiannis, and Andreas Kanavos. Literature review on hybrid evolutionary approaches for feature selection. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/3/167.

Grebennikov:2023:NMM

[2251] Dmitry Grebennikov, Rostislav Savinkov, Ekaterina Zelenova, Gennady Lobov, and Gennady Bocharov. Network modeling of murine lymphatic system. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/3/168.

Balduccini:2023:PLC

[2252] Marcello Balduccini, Michael Barborak, and David Ferrucci. Pushing the limits of Clingo’s incremental grounding and solving capabilities in practical applications. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/3/169.

Akulich:2023:MTD

[2253] Mikita Akulich, Iztok Savnik, Matjaz Krnc, and Riste Skrekovski. Multiset-trie data structure. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/3/170.

Huang:2023:MSA

[2254] Lingfeng Huang, Jieyu Zhao, and Yu Chen. MixFormer: a self-attentive convolutional network for 3D mesh object recognition. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/3/171.

Nielsen:2023:SSR

[2255] Kasper Gaj Nielsen, Inkyung Sung, Mohamed El Yafrani, Deniz Kenan Kiliç, and Peter Nielsen. A scheduling solution for robotic arm-based batching systems with multiple conveyor belts. Algorithms (Basel), 16 (3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/3/172.

Azeraf:2023:EBL

[2256] Elie Azeraf, Emmanuel Monfrini, and Wojciech Pieczynski. Equivalence between LC-CRF and HMM, and discriminative computing of HMMbased MPM and MAP. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/3/173.

Jourdan:2023:HOT

[2257] Fanny Jourdan, Titon Tshiongo Kaninku, Nicholas Asher, Jean-Michel Loubes, and Laurent Risser. How optimal transport can tackle gender biases in multi-class neural network classifiers for job recommendations. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/3/174.

Saleh:2023:GAN

[2258] Kaziwa Saleh, Sándor Szénási, and Zoltán Vámossy. Generative adversarial network for overcoming occlusion in images: a survey. Algorithms (Basel), 16(3), March 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/3/175.

Mahmud:2023:DAB

[2259] Md Ishtyaq Mahmud, Muntasir Mamun, and Ahmed Abdelgawad. A deep analysis of brain tumor detection from MR images using deep learning networks. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/176.

Baskan:2023:SBM

[2260] Denis E. Baskan, Daniel Meyer, Sebastian Mieck, Leonhard Faubel, Benjamin Klöpper, Nika Strem, Johannes A. Wagner, and Jan J. Koltermann. A scenario-based model comparison for short-term day-ahead electricity prices in times of economic and political tension. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/4/177.

Provan:2023:TEA

[2261] Gregory Provan. Toward explainable AutoEncoder-based diagnosis of dynamical systems. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/4/178.

Haubelt:2023:ESD

[2262] Christian Haubelt, Luise Müller, Kai Neubauer, Torsten Schaub, and Philipp Wanko. Evolutionary system design with answer set programming. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/4/
179.

Takyi:2023:BDS

[2263] Eric M. Takyi, Rana D. Parshad, Ranjit Kumar Upadhyay, and Vikas Rai. Blow-up dynamics and synchronization in tri-trophic food chain models. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/4/
180.

Walters:2023:HOB

[2264] Bradley Walters, Sandra Ortega-Martorell, Ivan Olier, and Paulo J. G. Lisboa. How to open a black box classifier for tabular data. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/4/181.

Panteli:2023:ACS

[2265] Antiopi Panteli and Basilis Boutsinas. Addressing the cold-start problem in recommender systems based on frequent patterns. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/4/182.

Schiller:2023:PMH

[2266] Vivian Schiller, Sandra Klaus, Ali Bilen, and Gisela Lanza. In-process monitoring of hobbing process using an acoustic emission sensor and supervised machine learning. Algorithms (Basel), 16(4), April 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/4/183.

Saleh:2023:ANR

[2267] Mahmoud Saleh, Endre Kovács, and Imre Ferenc Barna. Analytical and numerical results for the transient diffusion equation with diffusion coefficient depending on both space and time. Algorithms (Basel), 16 (4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/4/184.

Cabalar:2023:SHA

[2268] Pedro Cabalar, Jorge Fandinno, Torsten Schaub, and Philipp Wanko. On the semantics of hybrid ASP systems based on Clingo. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/4/185.

Pester:2023:TMD

[2269] Andreas Pester, Yevgeniya Sulema, Ivan Dychka, and Olga Sulema. Temporal multimodal data-processing algorithms based on algebraic system of aggregates. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/186.

Lindstrom:2023:CSS

[2270] Michael R. Lindstrom, Xiaofu Ding, Feng Liu, Anand Somayajula, and Deanna Needell. Continuous semi-supervised nonnegative matrix factorization. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/4/
187.

Jansson:2023:ESA

[2271] Jesper Jansson. Editorial: Surveys in algorithm analysis and complexity theory (special issue). Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/4/188.

Kulikov:2023:ASA

[2272] Aleksandr Kulikov, Pavel Ilyushin, Anton Loskutov, and Sergey Filippov. Application of search algorithms in determining fault location on overhead power lines according to the emergency mode parameters. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/4/189.

Zhang:2023:JSB

[2273] Ji Zhang, Xiangze Jia, Zhen Wang, Yonglong Luo, Fulong Chen, Gaoming Yang, and Lihui Zhao. JointContrast: Skeleton-based interaction recognition with new representation and contrastive learning. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/4/190.

Pellegrini:2023:OSI

[2274] Riccardo Pellegrini and Andrea Serani. Overview on the special issue on “Simulation-Based Optimization: Methods and Applications in Engineering Design”. Algorithms (Basel), 16(4), April 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/4/191.

Wei:2023:NSM

[2275] Shenghan Wei, Xiang Li, Yong Yao, and Suixian Yang. A novel shortmemory sequence-based model for variable-length reading recognition of multi-type digital instruments in industrial scenarios. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/4/192.

Dolores-Cuenca:2023:PC

[2276] Eric Dolores-Cuenca, José Antonio Arciniega-Nevárez, Anh Nguyen, Amanda Yitong Zou, Luke Van Popering, Nathan Crock, Gordon Erlebacher, and Jose L. Mendoza-Cortes. Polychrony as Chinampas. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/4/193.

Lu:2023:NPM

[2277] Binghang Lu, Christian Moya, and Guang Lin. NSGA-PINN: a multiobjective optimization method for physics-informed neural network training. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/4/
194.

Scrucca:2023:EBA

[2278] Luca Scrucca. Entropy-based anomaly detection for Gaussian mixture modeling. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/195.

Chamorro-Padial:2023:ASS

[2279] Jorge Chamorro-Padial and Rosa Rodríguez-Sánchez. Attention-survival score: a metric to choose better keywords and improve visibility of information. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/196.

Chen:2023:ADL

[2280] Lei Chen, Zhihao Wang, Ru Huo, and Tao Huang. An adversarial DBN-LSTM method for detecting and defending against DDoS attacks in SDN environments. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/4/197.

Demidova:2023:MLS

[2281] Liliya Demidova, Dmitry Zhukov, Elena Andrianova, and Vladimir Kalinin. Model of lexico-semantic bonds between texts for creating their similarity metrics and developing statistical clustering algorithm. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/4/198.

Nigmatullin:2023:CSS

[2282] Raoul R. Nigmatullin and Jocelyn Sabatier. Can self-similarity processes be reflected by the power-law dependencies? Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/4/199.

Hasan:2023:MAU

[2283] Adnan Alhaj Hasan, Aleksey A. Kvasnikov, Dmitriy V. Klyukin, Anton A. Ivanov, Alexander V. Demakov, Dmitry M. Mochalov, and Sergei P. Kuksenko. On modeling antennas using MoM-based algorithms: Wire-grid versus surface triangulation. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/4/200.

Matthes:2023:LCO

[2284] Dimitrios Matthes and Vasileios Drakopoulos. Line clipping in 3D: Overview, techniques and algorithms. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/4/201.

Kontogiannis:2023:DLS

[2285] Sotirios Kontogiannis, Theodosios Gkamas, and Christos Pikridas. Deep learning stranded neural network model for the detection of sensory triggered events. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/202.

Jedermann:2023:FLL

[2286] Reiner Jedermann. Feasibility of low latency, single-sample delay resampling: a new kriging based method. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/4/203.

Saveliev:2023:INE

[2287] Maxim Saveliev, Maksym Pantin, Igor Skiter, Thomas B. Scott, and Peter G. Martin. Implementation of novel evolutional algorithm for 3dimensional radiation mapping and gamma-field reconstruction within the Chornobyl sarcophagus. Algorithms (Basel), 16(4), April 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/4/204.

Stallmann:2023:UCS

[2288] Dominik Stallmann and Barbara Hammer. Unsupervised cyclic Siamese networks automating cell imagery analysis. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/4/205.

Kaltsounis:2023:CTA

[2289] Anastasios Kaltsounis, Evangelos Spiliotis, and Vassilios Assimakopoulos. Conditional temporal aggregation for time series forecasting using feature-based meta-learning. Algorithms (Basel), 16(4), April 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/4/206.

Huang:2023:TBI

[2290] Shenquan Huang, Ya-Chih Tsai, and Fuh-Der Chou. A trajectory-based immigration strategy genetic algorithm to solve a single-machine scheduling problem with job release times and flexible preventive maintenance. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/4/207.

Michaloglou:2023:BSC

[2291] Alkmini Michaloglou and Nikolaos L. Tsitsas. A brain storm and chaotic accelerated particle swarm optimization hybridization. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/4/208.

Bickel:2023:DPE

[2292] Sebastian Bickel, Stefan Goetz, and Sandro Wartzack. Detection of plausibility and error reasons in finite element simulations with deep learning networks. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/209.

Robitzsch:2023:MRE

[2293] Alexander Robitzsch. Model-robust estimation of multiple-group structural equation models. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/4/210.

Nazari:2023:IDT

[2294] Zahra Nazari and Petr Musilek. Impact of digital transformation on the energy sector: a review. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/4/211.

Fischer:2023:ARS

[2295] Anne Fischer, Alexandre Beiderwellen Bedrikow, Iris D. Tommelein, Konrad Nübel, and Johannes Fottner. From activity recognition to simulation: The impact of granularity on production models in heavy civil engineering. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/4/212.

Bagnoli:2023:SCD

[2296] Franco Bagnoli and Michele Baia. Synchronization, control and data assimilation of the Lorenz system. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/4/213.

Kumagai:2023:IBK

[2297] Masahito Kumagai, Kazuhiko Komatsu, Masayuki Sato, and Hiroaki Kobayashi. Ising-based kernel clustering. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/4/214.

Chalkiadakis:2023:NHR

[2298] Georgios Chalkiadakis, Ioannis Ziogas, Michail Koutsmanis, Errikos Streviniotis, Costas Panagiotakis, and Harris Papadakis. A novel hybrid recommender system for the tourism domain. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/4/215.

Rajaratnam:2023:SIS

[2299] David Rajaratnam, Torsten Schaub, Philipp Wanko, Kai Chen, Sirui Liu, and Tran Cao Son. Solving an industrial-scale warehouse delivery problem with answer set programming modulo difference constraints. Algorithms (Basel), 16(4), April 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/4/216.

Turov:2023:EDA

[2300] Artem T. Turov, Yuri A. Konstantinov, Fedor L. Barkov, Dmitry A. Korobko, Igor O. Zolotovskii, Cesar A. Lopez-Mercado, and Andrei A.

Fotiadi. Enhancing the distributed acoustic sensors’ (DAS) performance by the simple noise reduction algorithms sequential application. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/5/217.

Belmecheri:2023:BLR

[2301] Nassim Belmecheri, Noureddine Aribi, Nadjib Lazaar, Yahia Lebbah, and Samir Loudni. Boosting the learning for ranking patterns. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/5/218.

Alekseev:2023:DPN

[2302] Aleksandr Alekseev, Leonid Kozhemyakin, Vladislav Nikitin, and Julia Bolshakova. Data preprocessing and neural network architecture selection algorithms in cases of limited training sets-on an example of diagnosing Alzheimer’s disease. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/219.

Yang:2023:IDD

[2303] Xiao Yang and Qilong Han. Improved DQN for dynamic obstacle avoidance and ship path planning. Algorithms (Basel), 16(5), May 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/5/220.

Brzic:2023:DDU

[2304] Barbara Brzic, Ivica Boticki, and Marina Bagic Babac. Detecting deception using natural language processing and machine learning in datasets on COVID-19 and climate change. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/221.

Dobrev:2023:AGD

[2305] Stefan Dobrev, Paola Flocchini, Giuseppe Prencipe, and Nicola Santoro. Asynchronous gathering in a dangerous ring. Algorithms (Basel), 16 (5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/5/222.

Barghi:2023:SNU

[2306] Amir Barghi and Daryl DeFord. Stirling numbers of uniform trees and related computational experiments. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/223.

Zhang:2023:BPA

[2307] Lei Zhang and Zhi Pei. A branch-and-price algorithm for the online scheduling of valet drivers. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/224.

Santamaria-Henao:2023:OSS

[2308] Nicolas Santamaria-Henao, Oscar Danilo Montoya, and César Leonardo Trujillo-Rodríguez. Optimal siting and sizing of FACTS in distribution networks using the black widow algorithm. Algorithms (Basel), 16 (5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/5/225.

Konstantinov:2023:HTE

[2309] Andrei Konstantinov, Stanislav Kirpichenko, and Lev Utkin. Heterogeneous treatment effect with trained kernels of the Nadaraya–Watson regression. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/5/226.

Nguyen:2023:UDO

[2310] Hy Nguyen, Srikanth Thudumu, Hung Du, Kon Mouzakis, and Rajesh Vasa. UAV dynamic object tracking with lightweight deep vision reinforcement learning. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/227.

Luo:2023:EAM

[2311] Chenhong Luo, Yong Wang, Bo Li, Hanyang Liu, Pengyu Wang, and Leo Yu Zhang. An efficient approach to manage natural noises in recommender systems. Algorithms (Basel), 16(5), May 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/228.

Olguin-Roque:2023:RFT

[2312] Jairo Olguin-Roque, Sergio Salazar, Iván González-Hernandez, and Rogelio Lozano. A robust fixed-time sliding mode control for quadrotor UAV. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/

Hocke:2023:AMB

[2313] Lia Maria Hocke, Yunjie Tong, and Blaise deBonneval Frederick. An automatic motion-based artifact reduction algorithm for fNIRS in concurrent functional magnetic resonance imaging studies (AMARA-fMRI). Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/230.

Narvaez:2023:OBS

[2314] Diego Sanchez Narvaez, Carlos Villaseñor, Carlos Lopez-Franco, and Nancy Arana-Daniel. Order-based schedule of dynamic topology for recurrent neural network. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/231.

Hernandez:2023:MLB

[2315] Cesar Davila Hernandez, Jungseok Ho, Dongchul Kim, and Abdoul Oubeidillah. Machine-learning-based model for hurricane storm surge forecasting in the Lower Laguna Madre. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/5/232.

Santana:2023:BMA

[2316] Pedro Santana and José Moura. A Bayesian multi-armed bandit algorithm for dynamic end-to-end routing in SDN-based networks with piecewise-stationary rewards. Algorithms (Basel), 16(5), May 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/5/233.

Wang:2023:DCN

[2317] Yinghui Wang, Wenjun Wang, Minglai Shao, and Yueheng Sun. Deep cross-network alignment with anchor node pair diverse local structure. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/234.

Lu:2023:OIM

[2318] Yunlong Lu, Wenxin Li, and Wenlong Li. Official international mahjong: a new playground for AI research. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/235.

Palanivinayagam:2023:TYM

[2319] Ashokkumar Palanivinayagam, Claude Ziad El-Bayeh, and Robertas Damasevicius. Twenty years of machine-learning-based text classification: a systematic review. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/236.

Zghoul:2023:MBA

[2320] Fadi Nessir Zghoul, Haneen Alteehi, and Ahmad Abuelrub. A mayflybased approach for CMOS inverter design with symmetrical switching. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/237.

Agayan:2023:TSA

[2321] Sergey M. Agayan, Dmitriy A. Kamaev, Shamil R. Bogoutdinov, Andron O. Aleksanyan, and Boris V. Dzeranov. Time series analysis by fuzzy logic methods. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/238.

Guan:2023:IHM

[2322] Qingji Guan, Qinrun Chen, and Yaping Huang. An improved heteroscedastic modeling method for chest X-ray image classification with noisy labels. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/5/239.

Srivastava:2023:CAB

[2323] Harshit Srivastava and Ravi Sankar. Cooperative attention-based learning between diverse data sources. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/240.

Moawad:2023:QCW

[2324] Youssef Moawad, Wim Vanderbauwhede, and René Steijl. Quantum circuit-width reduction through parameterisation and specialisation. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/5/241.

Lukyanenko:2023:PAS

[2325] Dmitry Lukyanenko. Parallel algorithm for solving overdetermined systems of linear equations, taking into account round-off errors. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/5/242.

Gong:2023:OCY

[2326] Liyun Gong, Miao Yu, and Stefanos Kollias. Optimizing crop yield and reducing energy consumption in greenhouse control using PSO-MPC algorithm. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/
243.

Wang:2023:PPS

[2327] Yi-Wei Wang and Ja-Ling Wu. A privacy-preserving symptoms retrieval system with the aid of homomorphic encryption and private set intersection schemes. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/5/244.

Karras:2023:CBD

[2328] Christos Karras, Aristeidis Karras, Konstantinos C. Giotopoulos, Markos Avlonitis, and Spyros Sioutas. Consensus big data clustering for Bayesian mixture models. Algorithms (Basel), 16(5), May 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/245.

Masich:2023:SDM

[2329] Igor Masich, Natalya Rezova, Guzel Shkaberina, Sergei Mironov, Mariya Bartosh, and Lev Kazakovtsev. Subgroup discovery in machine learning problems with formal concepts analysis and test theory algorithms. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/5/246.

Herrmann:2023:PCO

[2330] Kevin Herrmann, Stefan Plappert, Paul Christoph Gembarski, and Roland Lachmayer. Process chain-oriented design evaluation of multimaterial components by knowledge-based engineering. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/5/247.

Kramar:2023:TSF

[2331] Vadim Kramar and Vasiliy Alchakov. Time-series forecasting of seasonal data using machine learning methods. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/5/248.

Orlova:2023:MDD

[2332] Yulia Orlova, Alexander Gorobtsov, Oleg Sychev, Vladimir Rozaliev, Alexander Zubkov, and Anastasia Donsckaia. Method for determining the dominant type of human breathing using motion capture and machine learning. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/
249.

Shaikh:2023:RFF

[2333] Shadab Anwar Shaikh, Harish Cherukuri, and Taufiquar Khan. Recovering the forcing function in systems with one degree of freedom using ANN and physics information. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/250.

Ivkovic:2023:APR

[2334] Nikola Ivković, Robert Kudelić, and Marin Golub. Adjustable pheromone reinforcement strategies for problems with efficient heuristic information. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/5/251.

Niu:2023:EML

[2335] Zhiqiang Niu, Shengnan Wu, and Xuesong (Simon) Zhou. Efficient mathematical lower bounds for city logistics distribution network with intraechelon connection of facilities: Bridging the gap from theoretical model formulations to practical solutions. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/252.

Pendrill:2023:MRO

[2336] Leslie R. Pendrill, Jeanette Melin, Anne Stavelin, and Gunnar Nordin. Modernising receiver operating characteristic (ROC) curves. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/5/253.

Matijevic:2023:WSP

[2337] Domagoj Matijević. Well-separated pair decompositions for highdimensional datasets. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/254.

Velichko:2023:NNE

[2338] Andrei Velichko, Maksim Belyaev, Yuriy Izotov, Murugappan Murugappan, and Hanif Heidari. Neural network entropy (NNetEn): Entropybased EEG signal and chaotic time series classification, Python package for NNetEn calculation. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/255.

Shaer:2023:HMC

[2339] Ibrahim Shaer and Abdallah Shami. Hierarchical modelling for CO_{2} variation prediction for HVAC system operation. Algorithms (Basel), 16 (5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/5/256.

Gataric:2023:PRT

[2340] Dragan Gatarić, Nenad Ruskić, Branko Aleksić, Tihomir Durić, Lato Pezo, Biljana Loncar, and Milada Pezo. Predicting road traffic accidentsartificial neural network approach. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/5/257.

Ahammed:2023:TEI

[2341] Abu Shad Ahammed, Aniebiet Micheal Ezekiel, and Roman Obermaisser. Time-efficient identification procedure for neurological complications of rescue patients in an emergency scenario using hardwareaccelerated artificial intelligence models. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/5/258.

Busch:2023:UDL

[2342] Lukas Busch, Ruben van Heusden, and Maarten Marx. Using deeplearned vector representations for page stream segmentation by agglomerative clustering. Algorithms (Basel), 16(5), May 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/5/259.

Poddubrovskii:2023:SPO

[2343] Nikita R. Poddubrovskii, Ivan A. Lobach, and Sergey I. Kablukov. Signal processing in optical frequency domain reflectometry systems based on self-sweeping fiber laser with continuous-wave intensity dynamics. Algorithms (Basel), 16(5), May 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/5/260.

Pavel:2023:OCM

[2344] Saidur R. Pavel and Yimin D. Zhang. Optimization of the compressive measurement matrix in a massive MIMO system exploiting LSTM networks. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/
261.

Mohammadi:2023:AEV

[2345] Masoud Mohammadi, Poria Fajri, Reza Sabzehgar, and Farshad Harirchi. Autonomous electric vehicle route optimization considering regenerative braking dynamic low-speed boundary. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/262.

Zhang:2023:DDP

[2346] Mu Zhang, Fengqiang Wan, and Taigang Liu. DrugFinder: Druggable protein identification model based on pre-trained models and evolutionary information. Algorithms (Basel), 16(6), June 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/263.

Crisci:2023:APP

[2347] Serena Crisci, Valentina De Simone, and Marco Viola. On the adaptive penalty parameter selection in ADMM. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/6/264.

Mzili:2023:HGS

[2348] Toufik Mzili, Ilyass Mzili, Mohammed Essaid Riffi, and Gaurav Dhiman. Hybrid genetic and spotted hyena optimizer for flow shop scheduling problem. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/6/265.

Shablya:2023:CGA

[2349] Yuriy Shablya. Combinatorial generation algorithms for some lattice paths using the method based on AND/OR trees. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/266.

Maniezzo:2023:LIH

[2350] Vittorio Maniezzo and Tingting Zhou. Learning individualized hyperparameter settings. Algorithms (Basel), 16(6), June 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/267.

Werner:2023:SIS

[2351] Frank Werner. Special issue “Scheduling: Algorithms and Applications”. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/268.

Arakane:2023:EDM

[2352] Taiki Arakane and Takeshi Saitoh. Efficient DNN model for word lipreading. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/
269.

Piccolomini:2023:CAS

[2353] Elena Loli Piccolomini, Marco Prato, Margherita Scipione, and Andrea Sebastiani. CTprintNet: an accurate and stable deep unfolding approach for few-view CT reconstruction. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/270.

Taherdoost:2023:ESM

[2354] Hamed Taherdoost. Enhancing social media platforms with machine learning algorithms and neural networks. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/6/271.

Golev:2023:AAS

[2355] Angel Golev, Snezhana Hristova, and Asen Rahnev. Algorithm for approximate solving of a nonlinear boundary value problem for generalized proportional Caputo fractional differential equations. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/6/272.

Candan:2023:RTI

[2356] Fethi Candan, Omer Faruk Dik, Tufan Kumbasar, Mahdi Mahfouf, and Lyudmila Mihaylova. Real-time interval type-2 fuzzy control of an unmanned aerial vehicle with flexible cable-connected payload. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/6/273.

Lopez-Martinez-Carrasco:2023:VES

[2357] Antonio Lopez-Martinez-Carrasco, Jose M. Juarez, Manuel Campos, and Bernardo Canovas-Segura. VLSD — an efficient subgroup discovery algorithm based on equivalence classes and optimistic estimate. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/6/274.

Yu:2023:IDE

[2358] Fei Yu, Bang Liang, Bo Tang, and Hongrun Wu. An interactive differential evolution algorithm based on backtracking strategy applied in interior layout design. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/275.

Ghafari:2023:CVN

[2359] Chadi Ghafari, Khalil Houissa, Jo Dens, Claudiu Ungureanu, Peter Kayaert, Cyril Constant, and Stéphane Carlier. Clinical validation of a new enhanced stent imaging method. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/6/276.

Baldini:2023:NMF

[2360] Gianmarco Baldini and Fausto Bonavitacola. Nakagami-m fading channel identification using adaptive continuous wavelet transform and convolutional neural networks. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/277.

Yuan:2023:FFD

[2361] Zhehu Yuan, Yinqi Sun, and Dennis Shasha. Forgetful forests: Data structures for machine learning on streaming data under concept drift. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/278.

Ochoa:2023:STF

[2362] Patricia Ochoa, Cinthia Peraza, Oscar Castillo, and Zong Woo Geem. A shadowed type-2 fuzzy approach for crossover parameter adaptation in differential evolution. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/279.

Deak:2023:RLN

[2363] Szabolcs Deák, Paul Levine, Joseph Pearlman, and Bo Yang. Reinforcement learning in a new Keynesian model. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/6/280.

Phetmak:2023:FEP

[2364] Nattawut Phetmak and Jittat Fakcharoenphol. Folding every point on a polygon boundary to a point. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/281.

Engelhardt:2023:IOD

[2365] Raphael C. Engelhardt, Marc Oedingen, Moritz Lange, Laurenz Wiskott, and Wolfgang Konen. Iterative oblique decision trees deliver explainable RL models. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/6/282.

Scheepers:2023:PMC

[2366] Christiaan Scheepers and Andries Engelbrecht. The porcupine measure for comparing the performance of multi-objective optimization algorithms. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/
283.

Diaz-del-Rio:2023:FPH

[2367] Fernando Díaz del Río, Pablo Sanchez-Cuevas, María José Moron-Fernández, Daniel Cascado-Caballero, Helena Molina-Abril, and Pedro Real. Fully parallel homological region adjacency graph via frontier recognition. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/6/284.

Jakobovic:2023:EDR

[2368] Domagoj Jakobović, Marko Durasević, Karla Brkić, Juraj Fosin, Tonci Carić, , and Davor Davidović. Evolving dispatching rules for dynamic vehicle routing with genetic programming. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/285.

Lin:2023:DPM

[2369] Yang-Kuei Lin, Chien-Fu Chen, and Tien-Yin Chou. Developing prediction model of travel times of the logistics fleets of large convenience store chains using machine learning. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/286.

Fang:2023:FHP

[2370] Jiannong Fang. A fast hybrid pressure-correction algorithm for simulating incompressible flows by projection methods. Algorithms (Basel), 16 (6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/287.

Li:2023:ADL

[2371] Xue Jun Li, Maode Ma, and Yihan Sun. An adaptive deep learning neural network model to enhance machine-learning-based classifiers for intrusion detection in smart grids. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/288.

Mpanti:2023:ATC

[2372] Anna Mpanti, Stavros D. Nikolopoulos, , and Leonidas Palios. Adding a tail in classes of perfect graphs. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/289.

Fichera:2023:IPT

[2373] Sebastiano Fichera, Gregorio Mariggiò, Mauro Corrado, and Giulio Ventura. Integration of polynomials times double step function in quadrilateral domains for XFEM analysis. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/290.

Tassopoulos:2023:ELP

[2374] Ioannis X. Tassopoulos, Christina A. Iliopoulou, Iosif V. Katsaragakis, and Grigorios N. Beligiannis. An effective local particle swarm optimization-based algorithm for solving the school timetabling problem. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/

Akingbesote:2023:IAF

[2375] Damilola Akingbesote, Ying Zhan, Rytis Maskeliunas, and Robertas Damasevicius. Improving accuracy of face recognition in the era of maskwearing: an evaluation of a Pareto-optimized FaceNet model with data preprocessing techniques. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/292.

Solorio-Ramirez:2023:RFA

[2376] José-Luis Solorio-Ramírez, Raúl Jiménez-Cruz, Yenny Villuendas-Rey, , and Cornelio Yáñez-Márquez. Random forest algorithm for the classification of spectral data of astronomical objects. Algorithms (Basel), 16 (6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/293.

Ivanov:2023:GAB

[2377] Anton A. Ivanov, Aleksey A. Kvasnikov, Alexander V. Demakov, Maxim E. Komnatnov, Sergei P. Kuksenko, and Talgat R. Gazizov. Generalized algorithm based on equivalent circuits for evaluating shielding effectiveness of electronic equipment enclosures. Algorithms (Basel), 16 (6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/294.

Ireddy:2023:EOQ

[2378] Ashish T. S. Ireddy and Sergey V. Kovalchuk. An experimental outlook on quality metrics for process modelling: a systematic review and meta analysis. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/6/295.

Yu:2023:NCM

[2379] Miaomiao Yu, Hongyong Yuan, Kaiyuan Li, , and Lizheng Deng. Noise cancellation method based on TVF-EMD with Bayesian parameter optimization. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/6/296.

Kyurkchiev:2023:COP

[2380] Vesselin Kyurkchiev, Anton Iliev, Asen Rahnev, and Nikolay Kyurkchiev. On a class of orthogonal polynomials as corrections in Lienard differential system: Applications. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/6/297.

Zhao:2023:PFT

[2381] Yiming Zhao and Jing Dong-O’Brien. Prediction of freeway traffic breakdown using artificial neural networks. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/6/298.

Manis:2023:MAC

[2382] George Manis, Dimitrios Bakalis, and Roberto Sassi. A multithreaded algorithm for the computation of sample entropy. Algorithms (Basel), 16 (6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/299.

Spyrou:2023:CCE

[2383] Evangelos D. Spyrou, Chrysostomos Stylios, , and Ioannis Tsoulos. Classification of CO environmental parameter for air pollution monitoring with grammatical evolution. Algorithms (Basel), 16(6), June 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/6/300.

Pachniak:2023:QUO

[2384] Elliot Pachniak, Yongzhen Fan, Wei Li, , and Knut Stamnes. Quantifying uncertainties in OC-SMART ocean color retrievals: a Bayesian inversion algorithm. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/6/301.

Urena-Camara:2023:IHC

[2385] Manuel Antonio Ureña-Cámara and Antonio Tomás Mozas-Calvache. Identification of highlighted cells in low-variance raster data application to digital elevation models. Algorithms (Basel), 16(6), June 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/6/302.

Lindh:2023:OWF

[2386] Per Lindh and Polina Lemenkova. Optimized workflow framework in construction projects to control the environmental properties of soil. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/6/303.

Zhang:2023:FDM

[2387] Lijun Zhang, Yuejian Zhang, and Guangfeng Li. Fault-diagnosis method for rotating machinery based on SVMD entropy and machine learning.

Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/6/304.

Di:2023:PID

[2388] Xuan Di, Rongye Shi, Zhaobin Mo, and Yongjie Fu. Physics-informed deep learning for traffic state estimation: a survey and the outlook. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/6/305.

Reato:2023:AMP

[2389] Federico Maria Reato, Claudio Ricci, Jan Misfatto, Matteo Calzaferri, and Simone Cinquemani. An alternative multi-physics-based methodology for strongly coupled electro-magneto-mechanical problems. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/6/306.

Kadlec:2023:MOP

[2390] Petr Kadlec. Multi-objective PSO with variable number of dimensions for space robot path optimization. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/6/307.

Asif:2023:EHD

[2391] Daniyal Asif, Mairaj Bibi, Muhammad Shoaib Arif, and Aiman Mukheimer. Enhancing heart disease prediction through ensemble learning techniques with hyperparameter optimization. Algorithms (Basel), 16(6), June 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/6/308.

Wang:2023:DAT

[2392] Yisha Wang, Gang Yang, and Hao Lu. A domain-adaptive tree-crown detection and counting method based on cascade region proposal networks. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/
309.

Merino:2023:PDE

[2393] Zach D. Merino, Jenny Farmer, and Donald J. Jacobs. Probability density estimation through nonparametric adaptive partitioning and stitching. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
7/310.

Efremov:2023:OSU

[2394] Andrei A. Efremov, Yuri N. Sotskov, and Yulia S. Belotzkaya. Optimization of selection and use of a machine and tractor fleet in agricultural enterprises: a case study. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/7/311.

Amirkhanova:2023:AEE

[2395] Gulshat Amirkhanova, Madina Mansurova, Gennadii Ososkov, Nasurlla Burtebayev, Adai Shomanov, , and Murat Kunelbayev. Algorithm for enhancing event reconstruction efficiency by addressing false track filtering issues in the SPD NICA experiment. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/312.

Ganesan:2023:EDO

[2396] Timothy Ganesan and Irraivan Elamvazuthi. Entanglement distillation optimization using fuzzy relations for quantum state tomography. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/7/313.

Zhang:2023:EMS

[2397] Yizheng Zhang, Liuhong Luo, and Hongjun Li. The extraction of maximal-sum principal submatrix and its applications. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/314.

Lv:2023:MSA

[2398] Wentao Lv, Fan Li, Shijie Luo, and Jie Xiang. A multi-site antiinterference neural network for ASD classification. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/7/315.

Schwegler:2023:IGF

[2399] Markus Schwegler, Christoph Müller, , and Alexander Reiterer. Integrated gradients for feature assessment in point cloud-based data sets. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/316.

Zhang:2023:AAS

[2400] Jiachen Zhang, Guoqing Tu, Shubo Liu, , and Zhaohui Cai. Audio antispoofing based on audio feature fusion. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/317.

Abdelhady:2023:DAT

[2401] Naglaa Abdelhady, Ibrahim E. Elsemman, Mohammed F. Farghally, and Taysir Hassan A. Soliman. Developing analytical tools for Arabic sentiment analysis of COVID-19 data. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/7/318.

Vaidya:2023:GED

[2402] Gauri Vaidya, Meghana Kshirsagar, and Conor Ryan. Grammatical evolution-driven algorithm for efficient and automatic hyperparameter optimisation of neural networks. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/7/319.

El-Mesady:2023:ODC

[2403] Ahmed El-Mesady, Tasneem Farahat, Ramadan El-Shanawany, and Aleksandr Y. Romanov. On orthogonal double covers and decompositions of complete bipartite graphs by caterpillar graphs. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/320.

Adamo:2023:SBG

[2404] Tommaso Adamo, Gianpaolo Ghiani, Emanuela Guerriero, and Deborah Pareo. A surprisal-based greedy heuristic for the set covering problem. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/321.

Savva:2023:RAN

[2405] Andreas G. Savva, Theocharis Theocharides, , and Chrysostomos Nicopoulos. Robustness of artificial neural networks based on weight alterations used for prediction purposes. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/322.

Almeida:2023:RLA

[2406] Pedro Almeida, Vitor Carvalho, and Alberto Simões. Reinforcement learning applied to AI bots in first-person shooters: a systematic review. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/323.

Jamshidi:2023:FSM

[2407] Mohammad (Behdad) Jamshidi, Salah I. Yahya, Saeed Roshani, Muhammad Akmal Chaudhary, Yazeed Yasin Ghadi, and Sobhan Roshani. A fast surrogate model-based algorithm using multilayer perceptron neural networks for microwave circuit design. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/324.

Theate:2023:RSP

[2408] Thibaut Théate and Damien Ernst. Risk-sensitive policy with distributional reinforcement learning. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/7/325.

deAndrade:2023:VVE

[2409] João V. R. de Andrade, Bruno J. T. Fernandes, André R. L. C. Izídio, Nilson M. da Silva Filho, and Francisco Cruz. Vessel velocity estimation and docking analysis: a computer vision approach. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/7/326.

Bergadano:2023:SIS

[2410] Francesco Bergadano and Giorgio Giacinto. Special issue “AI for cybersecurity: Robust models for authentication, threat and anomaly detection”. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/
327.

Fuentes:2023:CMS

[2411] Rafael Díaz Fuentes, Stefano Serra-Capizzano, and Rosita Luisa Sormani. ω-circulant matrices: a selection of modern applications from preconditioning of approximated PDEs to subdivision schemes. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/7/328.

Hernandez:2023:TCR

[2412] Francisco-David Hernandez, Domingo Cortes, Marco Antonio Ramirez-Salinas, and Luis Alfonso Villa-Vargas. A tool for control research using evolutionary algorithm that generates controllers with a pre-specified morphology. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/7/329.

Schilling:2023:BAO

[2413] Vincent Schilling, Peter Beyerlein, and Jeremy Chien. A bioinformatics analysis of ovarian cancer data using machine learning. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/330.

Starikova:2023:MMF

[2414] Victoria A. Starikova, Yuri A. Konin, Alexandra Yu. Petukhova, Svetlana S. Aleshkina, Andrey A. Petrov, and Anatolii V. Perminov. Mathematical model of fuse effect initiation in fiber core. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/7/331.

Kerger:2023:MAB

[2415] Phillip Kerger, David E. Bernal Neira, Zoe Gonzalez Izquierdo, and Eleanor G. Rieffel. Mind the O♡: Asymptotically better, but still impractical, quantum distributed algorithms. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/332.

Ciccolella:2023:TMA

[2416] Simone Ciccolella, Gianluca Della Vedova, Vladimir Filipović, and Mauricio Soto Gomez. Three metaheuristic approaches for tumor phylogeny inference: an experimental comparison. Algorithms (Basel), 16 (7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/7/333.

Yu:2023:OMS

[2417] Quanjiang Yu, Ola Carlson, and Serik Sagitov. Optimal maintenance schedule for a wind power turbine with aging components. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/334.

Biolek:2023:BFT

[2418] Dalibor Biolek, Viera Biolková, Zdenek Kolka, and Zdenek Biolek. Basis functions for a transient analysis of linear commensurate fractional-order systems. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/
335.

Hollosi:2023:SRM

[2419] János Hollósi, Áron Ballagi, , and Claudiu Radu Pozna. Simplified routing mechanism for capsule networks. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/336.

Galaktionov:2023:HMA

[2420] Ilya Galaktionov, Julia Sheldakova, Alexander Nikitin, Vladimir Toporovsky, and Alexis Kudryashov. A hybrid model for analysis of laser beam distortions using Monte Carlo and Shack–Hartmann techniques: Numerical study and experimental results. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/7/337.

Paudel:2023:LUD

[2421] Rajiv Paudel and Arika Ligmann-Zielinska. A largely unsupervised domain-independent qualitative data extraction approach for empirical agent-based model development. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/7/338.

Bouras:2023:CCV

[2422] Vasilis Bouras, Dimitris Spiliotopoulos, Dionisis Margaris, Costas Vassilakis, Konstantinos Kotis, Angeliki Antoniou, George Lepouras, Manolis Wallace, and Vassilis Poulopoulos. Chatbots for cultural venues: a topic-based approach. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/7/339.

Almiani:2023:BBA

[2423] Khaled Almiani, Mutaz Abu Alrub, Young Choon Lee, Taha H. Rashidi, and Amirmohammad Pasdar. A blockchain-based auction framework for location-aware services. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/7/340.

Erdos:2023:MIS

[2424] Péter L. Erdős and Tamás Róbert Mezei. Minimizing interferenceto-signal ratios in multi-cell telecommunication networks. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/341.

Lang:2023:LGS

[2425] Junwei Lang, Zhenjia Zeng, Tengfei Ma, , and Sailing He. Leaking gas source tracking for multiple chemical parks within an urban city. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/7/342.

Mamalis:2023:DLD

[2426] Marios Mamalis, Evangelos Kalampokis, Ilias Kalfas, and Konstantinos Tarabanis. Deep learning for detecting verticillium fungus in olive trees: Using YOLO in UAV imagery. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/7/343.

Belousov:2023:NNBa

[2427] Artemiy Belousov, Ivan Kisel, Robin Lakos, , and Akhil Mithran. Neuralnetwork-based quark–gluon plasma trigger for the CBM experiment at FAIR. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/
344.

Wari:2023:CMM

[2428] Ezra Wari, Weihang Zhu, and Gino Lim. A corrosion maintenance model using continuous state partially observable Markov decision process for oil and gas pipelines. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/7/345.

Koehl:2023:PVP

[2429] Patrice Koehl and Henri Orland. A physicist’s view on partial 3D shape matching. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/7/346.

He:2023:SMC

[2430] Zhenwen He, Chi Zhang, and Yunhui Cheng. Similarity measurement and classification of temporal data based on double mean representation. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/347.

Jin:2023:IDS

[2431] Haifeng Jin and Paul M. Goodrum. Integrated decision support framework of optimal scaffolding system for construction projects. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/348.

Nigro:2023:TMB

[2432] Libero Nigro and Pasi Fränti. Two medoid-based algorithms for clustering sets. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/7/349.

Fan:2023:ZSB

[2433] Jiajia Fan, Wentao Huang, Qingchao Jiang, , and Qinqin Fan. A zoning search-based multimodal multi-objective brain storm optimization algorithm for multimodal multi-objective optimization. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/7/350.

Ceccarelli:2023:LID

[2434] Giovanni Ceccarelli, Guido Cantelmo, Marialisa Nigro, and Constantinos Antoniou. Learning from imbalanced datasets: The bike-sharing inventory problem using sparse information. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/351.

Hoffmann:2023:GLP

[2435] Rolf Hoffmann. Generating loop patterns with a genetic algorithm and a probabilistic cellular automata rule. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/7/352.

Golyandina:2023:IIT

[2436] Nina Golyandina, Pavel Dudnik, and Alex Shlemov. Intelligent identification of trend components in singular spectrum analysis. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/7/353.

Ma:2023:HRL

[2437] Nan Ma, Ziyi Wang, Zeyu Ba, Xinran Li, Ning Yang, Xinyi Yang, and Haifeng Zhang. Hierarchical reinforcement learning for crude oil supply chain scheduling. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/7/354.

Gursesli:2023:DCB

[2438] Mustafa Can Gursesli, Mehmet Emin Selek, Mustafa Oktay Samur, Mirko Duradoni, Kyoungju Park, Andrea Guazzini, and Antonio Lanatà.

Design of cloud-based real-time eye-tracking monitoring and storage system. Algorithms (Basel), 16(7), July 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/7/
355.

Castillo:2023:SAR

[2439] Enrique Castillo. A state of the art review of systems of linear inequalities and related observability problems. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/8/356.

Stoykova:2023:AIM

[2440] Stela Stoykova and Nikola Shakev. Artificial intelligence for management information systems: Opportunities, challenges, and future directions. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/8/357.

Konstantinov:2023:MIL

[2441] Andrei Konstantinov, Lev Utkin, and Vladimir Muliukha. Multiple instance learning with trainable soft decision tree ensembles. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/8/358.

Nefedov:2023:ACA

[2442] Nikolay Nefedov, Bogdan Tishchenko, and Natalia Levashova. An algorithm for construction of the asymptotic approximation of a stable stationary solution to a diffusion equation system with a discontinuous source function. Algorithms (Basel), 16(8), August 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/359.

Cyperski:2023:HAC

[2443] Szymon Cyperski, Paweł D. Domański, , and Michał Okulewicz. Hybrid approach to the cost estimation of external-fleet full truckload contracts. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/8/360.

Eid:2023:VSH

[2444] Ahmed Eid and Friedhelm Schwenker. Visual static hand gesture recognition using convolutional neural network. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/8/361.

Napoli:2023:DRL

[2445] Claudia Di Napoli, Giovanni Paragliola, Patrizia Ribino, and Luca Serino. Deep-reinforcement-learning-based planner for city tours for cruise passengers. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/362.

Austin:2023:MRP

[2446] Tasha Austin and Bharat S. Rawal. Model retraining: Predicting the likelihood of financial inclusion in Kiva’s peer-to-peer lending to promote social impact. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/363.

Klein:2023:AFI

[2447] André Klein and Guy Mélard. An algorithm for the Fisher information matrix of a VARMAX process. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/8/364.

Tetteh:2023:EMO

[2448] Michael Tetteh, Allan de Lima, Jack McEllin, Aidan Murphy, Douglas Mota Dias, and Conor Ryan. Evolving multi-output digital circuits using multi-genome grammatical evolution. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/8/365.

Joshi:2023:MLT

[2449] Kunj Joshi, Chintan Bhatt, Kaushal Shah, Dwireph Parmar, Juan M. Corchado, Alessandro Bruno, and Pier Luigi Mazzeo. Machine-learning techniques for predicting phishing attacks in blockchain networks: a comparative study. Algorithms (Basel), 16(8), August 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/366.

Ntawuzumunsi:2023:DDE

[2450] Elias Ntawuzumunsi, Santhi Kumaran, Louis Sibomana, and Kambombo Mtonga. Design and development of energy efficient algorithm for smart beekeeping device to device communication based on data aggregation techniques. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/367.

Khan:2023:SIS

[2451] Muhammad Adnan Khan. Special issue “Algorithms for Feature Selection”. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/8/
368.

Ullah:2023:HAR

[2452] Hayat Ullah and Arslan Munir. Human action representation learning using an attention-driven residual 3DCNN network. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/8/369.

Gallego:2023:HSC

[2453] Andres Gallego and Francisco Roman. Hardware suitability of complex natural resonances extraction algorithms in backscattered radar signals. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/8/370.

Hafez:2023:IMP

[2454] Ishaq Hafez and Rached Dhaouadi. Identification of mechanical parameters in flexible drive systems using hybrid particle swarm optimization based on the quasi-Newton method. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/8/371.

Aroniadi:2023:APS

[2455] Chrysanthi Aroniadi and Grigorios N. Beligiannis. Applying particle swarm optimization variations to solve the transportation problem effectively. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/372.

Bridgelall:2023:DDD

[2456] Raj Bridgelall. Data-driven deployment of cargo drones: a U.S. case study identifying key markets and routes. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/8/373.

Du:2023:GPH

[2457] Ruijuan Du and Taiyang Tao. A greedy pursuit hierarchical iteration algorithm for multi-input systems with colored noise and unknown timedelays. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/374.

Crabu:2023:AIP

[2458] Elisa Crabu, Federica Pes, Giuseppe Rodriguez, and Giuseppa Tanda. Ascertaining the ideality of photometric stereo datasets under unknown lighting. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/375.

Nagadurga:2023:CMH

[2459] Timmidi Nagadurga, Ramesh Devarapalli, , and Lukasz Knypiński. Comparison of meta-heuristic optimization algorithms for global maximum power point tracking of partially shaded solar photovoltaic systems. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/8/376.

Grover:2023:ETL

[2460] Pratham Grover, Kunal Chaturvedi, Xing Zi, Amit Saxena, Shiv Prakash, Tony Jan, , and Mukesh Prasad. Ensemble transfer learning for distinguishing cognitively normal and mild cognitive impairment patients using MRI. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/377.

Demertzi:2023:OPD

[2461] Vasiliki Demertzi, Stavros Demertzis, , and Konstantinos Demertzis. An overview of privacy dimensions on the industrial Internet of Things (IIoT). Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/378.

Farmakis:2023:MOT

[2462] Panagiotis Farmakis, Athanasios Chassiakos, , and Stylianos Karatzas. A multi-objective tri-level algorithm for hub-and-spoke network in short sea shipping transportation. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/8/379.

Nourian:2023:DOT

[2463] Navid Nourian, Mamdouh El-Badry, and Maziar Jamshidi. Design optimization of truss structures using a graph neural network-based surrogate model. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/380.

Sangaiah:2023:IRV

[2464] Arun Kumar Sangaiah, Amir Javadpour, Chung-Chian Hsu, Anandakumar Haldorai, and Ahmad Zeynivand. Investigating routing in the VANET network: Review and classification of approaches. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/8/381.

Moscato:2023:LEU

[2465] Pablo Moscato, Mohammad Nazmul Haque, Kevin Huang, Julia Sloan, and Jonathon Corrales de Oliveira. Learning to extrapolate using continued fractions: Predicting the critical temperature of superconductor materials. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/8/382.

Belousov:2023:NNBb

[2466] Artemiy Belousov, Ivan Kisel, and Robin Lakos. A neural-network-based competition between short-lived particle candidates in the CBM experiment at FAIR. Algorithms (Basel), 16(8), August 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/383.

Clarke:2023:RLD

[2467] Robert Clarke, Liam Fletcher, Sebastian East, , and Thomas Richardson. Reinforcement learning derived high-alpha aerobatic manoeuvres for fixed wing operation in confined spaces. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/8/384.

Curtarelli:2023:CBL

[2468] Vitor Probst Curtarelli and Israel Cohen. Constant-beamwidth LCMV beamformer with rectangular arrays. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/8/385.

Saez:2023:RNM

[2469] Patricio Sáez, Carlos Herrera, and Victor Parada. Reducing nervousness in master production planning: A systematic approach incorporating product-driven strategies. Algorithms (Basel), 16(8), August 2023.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/8/386.

Ates:2023:MPE

[2470] Cihan Ates, Dogan Bicat, Radoslav Yankov, Joel Arweiler, Rainer Koch, and Hans-Jörg Bauer. Model predictive evolutionary temperature control via neural-network-based digital twins. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/8/387.

Makarova:2023:DMU

[2471] Irina Makarova, Polina Buyvol, Larisa Gabsalikhova, Eduard Belyaev, and Eduard Mukhametdinov. Decision making under conditions of uncertainty and risk in the formation of warehouse stock of an automotive service enterprise. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/388.

Vosswinkel:2023:SAD

[2472] Rick Voßwinkel and Klaus Röbenack. Systematic analysis and design of control systems based on Lyapunov’s direct method. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/8/389.

Bota:2023:GFE

[2473] András Bóta and Lauren Gardner. A generalized framework for the estimation of edge infection probabilities. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/8/390.

Castiello:2023:UEM

[2474] Matteo Castiello, Dajana Conte, and Samira Iscaro. Using epidemiological models to predict the spread of information on Twitter. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/8/391.

Wallace:2023:TNS

[2475] Lance Wallace. Testing a new ‘decrypted’ algorithm for plantower sensors measuring PM_{2.5}: Comparison with an alternative algorithm. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/8/392.

Guerra:2023:CSS

[2476] Juan F. Guerra, Ramon Garcia-Hernandez, Miguel A. Llama, and Victor Santibañez. A comparative study of swarm intelligence metaheuristics in UKF-based neural training applied to the identification and control of robotic manipulator. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/393.

Gaudioso:2023:BEM

[2477] Manlio Gaudioso, Sona Taheri, Adil M. Bagirov, and Napsu Karmitsa. Bundle enrichment method for nonsmooth difference of convex programming problems. Algorithms (Basel), 16(8), August 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/394.

Szwarcfiter:2023:BPS

[2478] Claudio Szwarcfiter, Yale T. Herer, and Avraham Shtub. Balancing project schedule, cost, and value under uncertainty: a reinforcement learning approach. Algorithms (Basel), 16(8), August 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/8/395.

Wilbert:2023:RFM

[2479] Henrique José Wilbert, Aurélio Faustino Hoppe, Andreza Sartori, Stefano Frizzo Stefenon, and Luís Augusto Silva. Recency, frequency, monetary value, clustering, and internal and external indices for customer segmentation from retail data. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/9/396.

Touafek:2023:RLD

[2480] Nesrine Touafek, Fatima Benbouzid-Si Tayeb, and Asma Ladj. A reinforcing-learning-driven artificial bee colony algorithm for scheduling jobs and flexible maintenance under learning and deteriorating effects. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/
397.

BenLetaifa:2023:VSP

[2481] Leila Ben Letaifa and Jean-Luc Rouas. Variable scale pruning for transformer model compression in end-to-end speech recognition. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/398.

Ghannad:2023:DIF

[2482] Niousha Ghannad and Kalpdrum Passi. Detecting image forgery over social media using U-NET with grasshopper optimization. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/399.

Nkuzo:2023:CAR

[2483] Lwando Nkuzo, Malusi Sibiya, and Elisha Didam Markus. A comprehensive analysis of real-time car safety belt detection using the YOLOv7 algorithm. Algorithms (Basel), 16(9), September 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/400.

Turki:2023:OTD

[2484] Turki Turki, Sanjiban Sekhar Roy, and Y.-H. Taguchi. Optimized tensor decomposition and principal component analysis outperforming stateof-the-art methods when analyzing histone modification chromatin immunoprecipitation profiles. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/9/401.

Ahsini:2023:EVT

[2485] Yusef Ahsini, Pablo Díaz-Masa, Belén Inglés, Ana Rubio, Alba Martínez, Aina Magraner, and J. Alberto Conejero. The electric vehicle traveling salesman problem on digital elevation models for traffic-aware urban logistics. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/402.

Harth:2023:DHR

[2486] Pascal Harth, Orlando Jähde, Sophia Schneider, Nils Horn, and Rüdiger Buchkremer. From data to human-readable requirements: Advancing requirements elicitation through language-transformer-enhanced opportunity mining. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/403.

Yin:2023:VCP

[2487] Fengyuan Yin, Xiaoming Yuan, Zhiao Ma, and Xinyu Xu. Vector control of PMSM using TD3 reinforcement learning algorithm. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/404.

Krasnova:2023:GAT

[2488] Svetlana A. Krasnova, Julia G. Kokunko, Sergey A. Kochetkov, and Victor A. Utkin. Generation of achievable three-dimensional trajectories for autonomous wheeled vehicles via tracking differentiators. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/405.

Fazekas:2023:HDM

[2489] Levente Fazekas, Boldizsár Tüü-Szabó, László T. Kóczy, Olivér Hornyák, and Károly Nehéz. A hybrid discrete memetic algorithm for solving flowshop scheduling problems. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/9/406.

Nazari:2023:NML

[2490] Leyla Nazari, Vida Ghotbi, Mohammad Nadimi, and Jitendra Paliwal. A novel machine-learning approach to predict stress-responsive genes in Arabidopsis. Algorithms (Basel), 16(9), September 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/407.

Leon:2023:HSR

[2491] Jonas F. Leon, Yuda Li, Xabier A. Martin, Laura Calvet, Javier Panadero, and Angel A. Juan. A hybrid simulation and reinforcement learning algorithm for enhancing efficiency in warehouse operations. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
9/408.

Oellerich:2023:RSR

[2492] Jan Oellerich, Keno Jann Büscher, and Jan Philipp Degel. RooTri: a simple and robust function to approximate the intersection points of a 3D scalar field with an arbitrarily oriented plane in MATLAB. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/409.

Izsak:2023:NNA

[2493] Ferenc Izsák and Rudolf Izsák. Neural-network-assisted finite difference discretization for numerical solution of partial differential equations. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
9/410.

Ribeiro:2023:EEA

[2494] Inês A. Ribeiro, Tiago Ribeiro, Gil Lopes, and A. Fernando Ribeiro. End-to-end approach for autonomous driving: a supervised learning method using computer vision algorithms for dataset creation. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/411.

Cimurs:2023:YAT

[2495] Reinis Cimurs, Vilnis Turkovs, Martins Banis, and Aleksandrs Korsunovs. You are not alone: Towards cleaning robot navigation in shared environments through deep reinforcement learning. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/9/412.

Gulic:2023:EMO

[2496] Marko Gulić and Martina Zuskin. Enhancing metaheuristic optimization: a novel nature-inspired hybrid approach incorporating selected pseudorandom number generators. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/9/413.

Buscema:2023:PRA

[2497] Paolo Massimo Buscema, Giulia Massini, Giovanbattista Raimondi, Giuseppe Caporaso, Marco Breda, and Riccardo Petritoli. A pattern recognition analysis of vessel trajectories. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/9/414.

Nishimura:2023:POI

[2498] Naoki Nishimura, Noriyoshi Sukegawa, Yuichi Takano, and Jiro Iwanaga. Predicting online item-choice behavior: a shape-restricted regression approach. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/415.

Yukalov:2023:DVC

[2499] Vyacheslav I. Yukalov and Elizaveta P. Yukalova. Discrete versus continuous algorithms in dynamics of affective decision making. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/416.

Ivanov:2023:OPM

[2500] Ivan G. Ivanov, Yordan Kumchev, and Vincent James Hooper. An optimization precise model of stroke data to improve stroke prediction. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/
417.

Batiha:2023:MSF

[2501] Iqbal M. Batiha, Ahmad A. Abubaker, Iqbal H. Jebril, Suha B. Al-Shaikh, Khaled Matarneh, and Manal Almuzini. A mathematical study on a fractional-order SEIR mpox model: Analysis and vaccination influence. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/418.

Satlawa:2023:UDL

[2502] Philipp Satlawa and Robert B. Fisher. Using deep learning to detect the need for forest thinning: Application to the Lungau Region, Austria. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/
419.

Gil-Gonzalez:2023:OIF

[2503] Walter Gil-González, Oscar Danilo Montoya, and César Leonardo Trujillo-Rodríguez. Optimal integration of flexible alternating current transmission systems in electrical distribution grids using a mixed-integer convex model. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/420.

Jiang:2023:SIS

[2504] Weiwei Jiang. Special issue “Neural Network for Traffic Forecasting”. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/
421.

Li:2023:MLB

[2505] Cong Li, Xupeng Ren, and Guohui Zhao. Machine-learning-based imputation method for filling missing values in ground meteorological observation data. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/422.

Borrero:2023:EUT

[2506] Juan D. Borrero and Jesus Mariscal. Elevating univariate time series forecasting: Innovative SVR-empowered nonlinear autoregressive neural networks. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/423.

Sawicki:2023:RCS

[2507] Jan Sawicki, Maria Ganzha, Marcin Paprzycki, and Yutaka Watanobe. Reddit CrosspostNet — studying Reddit communities with large-scale crosspost graph networks. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/9/424.

Caramia:2023:SMO

[2508] Massimiliano Caramia. “Multi-Objective and Multi-Level Optimization: Algorithms and Applications”: Foreword by the Guest Editor. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/425.

Xie:2023:FDC

[2509] Qian Xie and Tae J. Kwon. A framework for determining collision likelihood using continuous friction values in a connected vehicle environment. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/
426.

Fernandez:2023:OCR

[2510] Arturo J. Fernández. Optimal confidence regions for Weibull parameters and quantiles under progressive censoring. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/9/427.

Bowman:2023:PIN

[2511] Brett Bowman, Chad Oian, Jason Kurz, Taufiquar Khan, Eddie Gil, and Nick Gamez. Physics-informed neural networks for the heat equation with source term under various boundary conditions. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/9/428.

Chetkin:2023:BOB

[2512] Egor I. Chetkin, Sergei L. Shishkin, and Bogdan L. Kozyrskiy. Bayesian opportunities for brain-computer interfaces: Enhancement of the existing classification algorithms and out-of-domain detection. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/429.

Surendran:2023:ISR

[2513] Ranjini Surendran, Ines Chihi, J. Anitha, and D. Jude Hemanth. Indoor scene recognition: an attention-based approach using feature selectionbased transfer learning and deep liquid state machine. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/430.

Zahedi:2023:RCM

[2514] Esmaeil Zahedi, Mohamad Saraee, Fatemeh Sadat Masoumi, and Mohsen Yazdinejad. Regularized contrastive masked autoencoder model for machinery anomaly detection using diffusion-based data augmentation. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
9/431.

Lin:2023:SAG

[2515] Guo Lin and Yongfeng Zhang. Sparks of Artificial General Recommender (AGR): Experiments with ChatGPT. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/16/9/432.

Nayak:2023:RMA

[2516] Suraj Kumar Nayak, Bikash Pradhan, Biswaranjan Mohanty, Jayaraman Sivaraman, Sirsendu Sekhar Ray, Jolanta Wawrzyniak, Maciej Jarzebski, and Kunal Pal. A review of methods and applications for a heart rate variability analysis. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/433.

Ibanez:2023:CML

[2517] Javier Ibáñez, José M. Alonso, Emilio Defez, Pedro Alonso-Jordá, and Jorge Sastre. Computing the matrix logarithm with the Romberg integration method. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/434.

Apostolopoulos:2023:EAI

[2518] Dimitris J. Apostolopoulos, Ioannis D. Apostolopoulos, Nikolaos D. Papathanasiou, Trifon Spyridonidis, and George S. Panayiotakis. Explainable artificial intelligence method (ParaNet+) localises abnormal parathyroid glands in scintigraphic scans of patients with primary hyperparathyroidism. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/435.

Halawi:2023:MOD

[2519] Ola N. Halawi, Faisal N. Abu-Khzam, and Sergio Thoumi. A multiobjective degree-based network anonymization method. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/436.

Fawwaz:2023:DPC

[2520] Muhammad Amir Fawwaz, Kishore Bingi, Rosdiazli Ibrahim, P. Arun Mozhi Devan, and B. Rajanarayan Prusty. Design of PIDD^{α} controller for robust performance of process plants. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/9/437.

Wang:2023:QQC

[2521] Yu Wang, Alister Machado, and Alexandru Telea. Quantitative and qualitative comparison of decision-map techniques for explaining classification models. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/438.

Rodriguez-Sevillano:2023:AWW

[2522] Ángel Antonio Rodríguez-Sevillano, Rafael Bardera-Mora, Alejandra López-Cuervo-Alcaraz, Daniel Anguita-Mazón, Juan Carlos Matías-García, Estela Barroso-Barderas, and Jaime Fernández-Antón. Automation of winglet wings geometry generation for its application in TOR-NADO. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/439.

Turov:2023:AFD

[2523] Artem T. Turov, Fedor L. Barkov, Yuri A. Konstantinov, Dmitry A. Korobko, Cesar A. Lopez-Mercado, and Andrei A. Fotiadi. Activation function dynamic averaging as a technique for nonlinear 2D data denoising in distributed acoustic sensors. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/9/440.

Canali:2023:PIM

[2524] Claudia Canali, Caterina Gazzotti, Riccardo Lancellotti, and Felice Schena. Placement of IoT microservices in fog computing systems: a comparison of heuristics. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/9/441.

Wang:2023:PDC

[2525] Huan Wang, Shi Qiu, Huping Ye, and Xiaohan Liao. A plant disease classification algorithm based on attention MobileNet V2. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/442.

Favier:2023:TBA

[2526] Gérard Favier and Alain Kibangou. Tensor-based approaches for nonlinear and multilinear systems modeling and identification. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/9/443.

Karmitsa:2023:NOB

[2527] Napsu Karmitsa, Sona Taheri, Kaisa Joki, Pauliina Paasivirta, Adil M. Bagirov, and Marko M. Mäkelä. Nonsmooth optimization-based hyperparameter-free neural networks for large-scale regression. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/444.

Oghalai:2023:ASO

[2528] Thomas P. Oghalai, Ryan Long, Wihan Kim, Brian E. Applegate, and John S. Oghalai. Automated segmentation of optical coherence tomography images of the human tympanic membrane using deep learning. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/9/
445.

Robitzsch:2023:IAR

[2529] Alexander Robitzsch. Implementation aspects in regularized structural equation models. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/9/446.

Wang:2023:CGA

[2530] Jiaqi Wang, Bruce Golden, and Carmine Cerrone. Carousel greedy algorithms for feature selection in linear regression. Algorithms (Basel), 16 (9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/9/447.

Zotov:2023:NNB

[2531] Mikhail Zotov, Dmitry Anzhiganov, Aleksandr Kryazhenkov, Dario Barghini, Matteo Battisti, Alexander Belov, Mario Bertaina, Marta Bianciotto, Francesca Bisconti, Carl Blaksley, Sylvie Blin, Giorgio Cambiè, Francesca Capel, Marco Casolino, Toshikazu Ebisuzaki, Johannes Eser, Francesco Fenu, Massimo Alberto Franceschi, Alessio Golzio, Philippe Gorodetzky, Fumiyoshi Kajino, Hiroshi Kasuga, Pavel Klimov, Massimiliano Manfrin, Laura Marcelli, Hiroko Miyamoto, Alexey Murashov, Tommaso Napolitano, Hiroshi Ohmori, Angela Olinto, Etienne Parizot, Piergiorgio Picozza, Lech Wiktor Piotrowski, Zbigniew Plebaniak, Guillaume Prévôt, Enzo Reali, Marco Ricci, Giulia Romoli, Naoto Sakaki, Kenji Shinozaki, Christophe De La Taille, Yoshiyuki Takizawa, Michal Vrábel, and Lawrence Wiencke. Neural network based approach to recognition of meteor tracks in the mini-EUSO telescope data. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/448.

Salcido:2023:DLL

[2532] E. Anyuli Alvarez Salcido, Daniel Centeno-Barreda, Yukio Rosales, Ricardo Lopéz-Gutiérrez, Sergio Salazar, and Rogelio Lozano. Design of a lower limb exoskeleton: Robust control, simulation and experimental results. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/9/449.

Kumar:2023:ITA

[2533] Mohit Kumar, Bernhard A. Moser, Lukas Fischer, and Bernhard Freudenthaler. An information theoretic approach to privacy-preserving interpretable and transferable learning. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/9/450.

Manescu:2023:HAH

[2534] Alexandru-Razvan Manescu and Bogdan Dumitrescu. HyperDE: an adaptive hyper-heuristic for global optimization. Algorithms (Basel), 16 (9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/9/451.

Sami:2023:IYB

[2535] Abdullah As Sami, Saadman Sakib, Kaushik Deb, and Iqbal H. Sarker. Improved YOLOv5-based real-time road pavement damage detection in road infrastructure management. Algorithms (Basel), 16(9), September 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/9/452.

AlAmin:2023:EBW

[2536] Ahmed Al Amin, Junho Hong, Van-Hai Bui, and Wencong Su. Emerging 6G/B6G wireless communication for the power infrastructure in smart cities: Innovations, challenges, and future perspectives. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/10/474.

Alviano:2023:HTD

[2537] Mario Alviano. Hashcash tree, a data structure to mitigate denial-of-service attacks. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/462.

Bergomi:2023:EGD

[2538] Mattia G. Bergomi and Massimo Ferri. Exploring graph and digraph persistence. Algorithms (Basel), 16(10):??, October 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/465.

Cagliero:2023:DDA

[2539] Luca Cagliero, Lorenzo Canale, and Laura Farinetti. Data-driven analysis of student engagement in time-limited computer laboratories. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
10/464.

Carballo:2023:CDT

[2540] Jose Antonio Carballo, Javier Bonilla, Jesús Fernández-Reche, Bijan Nouri, Antonio Avila-Marin, Yann Fabel, and Diego-César Alarcón-Padilla. Cloud detection and tracking based on object detection with convolutional neural networks. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/10/487.

Chou:2023:PDS

[2541] Xiaochen Chou and Enza Messina. Problem-driven scenario generation for stochastic programming problems: a survey. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/479.

Dai:2023:MFS

[2542] Manli Dai and Zhongyi Jiang. Multiprocessor fair scheduling based on an improved slime mold algorithm. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/10/473.

Ezeife:2023:SSP

[2543] Christie I. Ezeife and Hemni Karlapalepu. A survey of sequential pattern based e-commerce recommendation systems. Algorithms (Basel), 16(10): ??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/10/467.

Faiyaz:2023:CDR

[2544] Chowdhury Abrar Faiyaz, Pabel Shahrear, Rakibul Alam Shamim, Thilo Strauss, and Taufiquar Khan. Comparison of different radial basis function networks for the electrical impedance tomography (EIT) inverse problem. Algorithms (Basel), 16(10):??, October 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/461.

Garcia-Olmedo:2023:MFF

[2545] Francisco Miguel García-Olmedo, Jesús García-Miranda, and Pedro González-Rodelas. Mathematical foundation of a functional implementation of the CNF algorithm. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/10/459.

Gholami:2023:ETC

[2546] Fatemeh Gholami, Zahed Rahmati, Alireza Mofidi, and Mostafa Abbaszadeh. On enhancement of text classification and analysis of text emotions using graph machine learning and ensemble learning methods on non-English datasets. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/10/470.

Grignaffini:2023:ADS

[2547] Flavia Grignaffini, Maurizio Troiano, Francesco Barbuto, Patrizio Simeoni, Fabio Mangini, Gabriele D’Andrea, Lorenzo Piazzo, Carmen Cantisani, Noah Musolff, Costantino Ricciuti, and Fabrizio Frezza. Anomaly detection for skin lesion images using convolutional neural network and injection of handcrafted features: a method that bypasses the preprocessing of dermoscopic images. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/10/466.

Hashemi:2023:BPP

[2548] Seyed Mohammad Hashemi, Ruxandra Mihaela Botez, and Georges Ghazi. Blockchain PoS and PoW consensus algorithms for airspace management application to the UAS-S4 Ehécatl. Algorithms (Basel), 16(10): ??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/10/472.

Hou:2023:SRM

[2549] Xinxiang Hou, Andong Qiu, Lu Yang, and Zhouwang Yang. Shelvedretrieved method for weakly balanced constrained clustering problems. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
10/492.

Kiany:2023:EUB

[2550] Katayoon Kiany, Abolfazl Baghbani, Hossam Abuel-Naga, Hasan Baghbani, Mahyar Arabani, and Mohammad Mahdi Shalchian. Enhancing ultimate bearing capacity prediction of cohesionless soils beneath shallow foundations with grey box and hybrid AI models. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/456.

Lange:2023:CHD

[2551] Kenneth Lange. Computation of the Hausdorff distance between two compact convex sets. Algorithms (Basel), 16(10):??, October 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/10/471.

Lin:2023:FEF

[2552] Yun-Wei Lin, Yuh-Hwan Liu, Yi-Bing Lin, and Jian-Chang Hong. FenceTalk: Exploring false negatives in moving object detection. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/10/481.

Lozi:2023:SRA

[2553] René Lozi. Survey of recent applications of the chaotic Lozi map. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
10/491.

Lu:2023:TPC

[2554] Shiyu Lu, Huamin Yang, and Cheng Han. TransPCGC: Point cloud geometry compression based on transformers. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/484.

Mao:2023:ERF

[2555] Ye-Jiao Mao, Andy Yiu-Chau Tam, Queenie Tsung-Kwan Shea, Yong-Ping Zheng, and James Chung-Wai Cheung. eNightTrack: Restraint-free depth-camera-based surveillance and alarm system for fall prevention using deep learning tracking. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/10/477.

Margerit:2023:IIG

[2556] William Margerit, Antoine Charpentier, Cathy Maugis-Rabusseau, Johann Christian Schön, Nathalie Tarrat, and Juan Cortés. IGLOO: an iterative global exploration and local optimization algorithm to find diverse low-energy conformations of flexible molecules. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/476.

Mathesul:2023:CDC

[2557] Shubham Mathesul, Debabrata Swain, Santosh Kumar Satapathy, Ayush Rambhad, Biswaranjan Acharya, Vassilis C. Gerogiannis, and Andreas Kanavos. COVID-19 detection from chest X-ray images based on deep learning techniques. Algorithms (Basel), 16(10):??, October 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/10/494.

Mercuri:2023:EAA

[2558] Valeria Mercuri, Martina Saletta, and Claudio Ferretti. Evolutionary approaches for adversarial attacks on neural source code classifiers. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
10/478.

Mosquera-Rojas:2023:AMS

[2559] Gonzalo Mosquera-Rojas, Cylia Ouadah, Azadeh Hadadi, Alain Lalande, and Sarah Leclerc. Automatic myocardium segmentation in delayedenhancement MRI with pathology-specific data augmentation and deep learning architectures. Algorithms (Basel), 16(10):??, October 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/10/488.

Ochoa-Munoz:2023:MFA

[2560] Andrés F. Ochoa-Muñoz and Javier E. Contreras-Reyes. Multiple factor analysis based on NIPALS algorithm to solve missing data problems. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
10/457.

ONeill:2023:AOA

[2561] Vyas O’Neill and Ben Soh. An aspect-oriented approach to timeconstrained strategies in smart city IoT applications. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/10/454.

Pachniak:2023:RSS

[2562] Elliot Pachniak, Wei Li, Tomonori Tanikawa, Charles Gatebe, and Knut Stamnes. Remote sensing of snow parameters: a sensitivity study of retrieval performance based on hyperspectral versus multispectral data. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
10/493.

Passacantando:2023:LCR

[2563] Mauro Passacantando and Fabio Raciti. Lipschitz continuity results for a class of parametric variational inequalities and applications to network games. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/10/458.

Patel:2023:MES

[2564] Tulsi Patel, Mark W. Jones, and Thomas Redfern. Manifold explorer: Satellite image labelling and clustering tool with using deep convolutional autoencoders. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/469.

Pokushko:2023:AAB

[2565] Mariia Pokushko, Alena Stupina, Inmaculada Medina-Bulo, Svetlana Ezhemanskaya, Roman Kuzmich, and Roman Pokushko. Algorithm for application of a basic model for the data envelopment analysis method in technical systems. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/460.

Robbins:2023:SMM

[2566] Edmund Robbins, Thu Thu Hlaing, Jonathan Webb, and Nezamoddin N. Kachouie. Supervised methods for modeling spatiotemporal glacier variations by quantification of the area and terminus of mountain glaciers using remote sensing. Algorithms (Basel), 16(10):??, October 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/10/486.

Sakaida:2023:DMC

[2567] Miu Sakaida, Takaaki Yoshimura, Minghui Tang, Shota Ichikawa, and Hiroyuki Sugimori. Development of a mammography calcification detection algorithm using deep learning with resolution-preserved image patch division. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/483.

Siddique:2023:SRR

[2568] Muhammad Talha Siddique, Paraskevas Koukaras, Dimosthenis Ioannidis, and Christos Tjortjis. SmartBuild RecSys: a recommendation system based on the smart readiness indicator for energy efficiency in buildings. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/10/482.

Soper:2023:UOM

[2569] Daniel S. Soper. Using an opportunity matrix to select centers for RBF neural networks. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/455.

Tanti:2023:DMM

[2570] Lili Tanti, Syahril Efendi, Maya Silvi Lydia, and Herman Mawengkang. A decision-making model to determine dynamic facility locations for a disaster logistic planning problem using deep learning. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/10/468.

Ukaegbu:2023:MOO

[2571] Uchechi Ukaegbu, Lagouge Tartibu, and C. W. Lim. Multi-objective optimization of a solar-assisted combined cooling, heating and power generation system using the greywolf optimizer. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/463.

Xing:2023:SGN

[2572] Siyuan Xing and Jian-Qiao Sun. Separable Gaussian neural networks: Structure, analysis, and function approximations. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/453.

Xu:2023:IEC

[2573] Wei Xu, Xunhong Zhu, Liping Zhang, and Jun Gao. The iterative exclusion of compatible samples workflow for multi-SNP analysis in complex diseases. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/480.

Yang:2023:DAM

[2574] Qing Yang, Bingyu Song, Yingguo Chen, Lei He, and Pei Wang. A distributed autonomous mission planning method for the low-orbit imaging constellation. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/475.

Yousefi:2023:DNN

[2575] Mahsa Yousefi and Ángeles Martínez. Deep neural networks training by stochastic quasi-Newton trust-region methods. Algorithms (Basel), 16 (10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/10/490.

Zhang:2023:NNE

[2576] Yifan Zhang and Quanmin Zhu. Neural network-enhanced fault diagnosis of robot joints. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/10/489.

Zhang:2023:RIM

[2577] Hairuilong Zhang, Yangsong Gu, and Lee D. Han. Representing and inferring massive network traffic condition: a case study in Nashville, Tennessee. Algorithms (Basel), 16(10):??, October 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/10/485.

Adekoya:2023:DMP

[2578] Adekunle Rotimi Adekoya and Mardé Helbig. Decision-maker’s preference-driven dynamic multi-objective optimization. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/11/504.

Adel:2023:GMS

[2579] Tameem Adel and Mark Levene. A general model for side information in neural networks. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/526.

Alarfaj:2023:DDF

[2580] Fawaz Khaled Alarfaj and Jawad Abbas Khan. Deep dive into fake news detection: Feature-centric classification with ensemble and deep learning methods. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/507.

Ali:2023:WBC

[2581] Mohamad Abou Ali, Fadi Dornaika, and Ignacio Arganda-Carreras. White blood cell classification: Convolutional neural network (CNN) and vision transformer (ViT) under medical microscope. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/11/525.

Appakaya:2023:PDC

[2582] Sai Bharadwaj Appakaya, Ruchira Pratihar, and Ravi Sankar. Parkinson’s disease classification framework using vocal dynamics in connected speech. Algorithms (Basel), 16(11):??, November 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/509.

Aviles:2023:AAR

[2583] Roberto Avilés, Diego Brito de Souza, José Pino-Ortega, and Julen Castellano. Agreement, accuracy, and reliability of a new algorithm for the detection of change of direction angle based on integrating inertial data from inertial sensors. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/11/496.

Boche:2023:TDR

[2584] Holger Boche, Yannik N. Böck, Ullrich J. Mönich, and Frank H. P. Fitzek. Trustworthy digital representations of analog information — an application-guided analysis of a fundamental theoretical problem in digital twinning. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/514.

Carpitella:2023:RSS

[2585] Silvia Carpitella, Bruno Brentan, Antonella Certa, and Joaquín Izquierdo. A recommendation system supporting the implementation of sustainable risk management measures in airport operations. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
11/511.

Chen:2023:MOO

[2586] Sirui Chen, Yuming Tian, and Lingling An. Multi-objective order scheduling via reinforcement learning. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/16/11/495.

Chen:2023:SNL

[2587] Ren-Raw Chen, Cameron D. Miller, and Puay Khoon Toh. Search on an NK landscape with swarm intelligence: Limitations and future research opportunities. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/527.

Dalal:2023:TWL

[2588] Ketan Dalal, Luc Devroye, and Ebrahim Malalla. Two-way linear probing revisited. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/500.

DellOglio:2023:SSR

[2589] Pietro Dell’Oglio, Alessandro Bondielli, and Francesco Marcelloni. A system to support readers in automatically acquiring complete summarized information on an event from different sources. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/11/513.

Huang:2023:RCM

[2590] Xinbo Huang, Zhiwei Song, Chao Ji, Ye Zhang, and Luya Yang. Research on a classification method for strip steel surface defects based on knowledge distillation and a self-adaptive residual shrinkage network. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
11/516.

Kadane:2023:TKA

[2591] Joseph B. Kadane. Two Kadane algorithms for the maximum sum subarray problem. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/519.

Korkhov:2023:FBM

[2592] Vladimir Korkhov, Ivan Gankevich, Anton Gavrikov, Maria Mingazova, Ivan Petriakov, Dmitrii Tereshchenko, Artem Shatalin, and Vitaly Slobodskoy. Finding bottlenecks in message passing interface programs by scalable critical path analysis. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/11/505.

Koutoulakis:2023:AMS

[2593] Emmanouil Koutoulakis, Louis Marage, Emmanouil Markodimitrakis, Leone Aubignac, Catherine Jenny, Igor Bessieres, and Alain Lalande. Automatic multiorgan segmentation in pelvic region with convolutional neural networks on 0.35 T MR-Linac images. Algorithms (Basel), 16(11): ??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/11/521.

Kozin:2023:ASF

[2594] Alexey Kozin, Anton Gerasimov, Maxim Bakaev, Anton Pashkov, and Olga Razumnikova. Automating stimulation frequency selection for SSVEP-based brain–computer interfaces. Algorithms (Basel), 16(11): ??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/11/502.

Latif:2023:ICG

[2595] Ehsan Latif and Ramviyas Parasuraman. On the intersection of computational geometry algorithms with mobile robot path planning. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/11/498.

Liu:2023:OSI

[2596] Jia-Bao Liu, Muhammad Faisal Nadeem, and Yilun Shang. Overview of the special issue on ‘Deep Neural Networks and Optimization Algorithms’. Algorithms (Basel), 16(11):??, November 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/497.

Nickkar:2023:DDR

[2597] Amirreza Nickkar and Young-Jae Lee. Dynamic demand-responsive feeder bus network design for a short headway trunk line. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/11/506.

Nickkar:2023:OIS

[2598] Amirreza Nickkar and Young-Jae Lee. Optimal integrated singleframework algorithm for the multi-level school bus network problem. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
11/528.

Nizharadze:2023:PGD

[2599] Nika Nizharadze, Arash Farokhi Soofi, and Saeed Manshadi. Predicting the gap in the day-ahead and real-time market prices leveraging exogenous weather data. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/11/508.

Nunez-Valdez:2023:SIA

[2600] Edward Rolando Núñez-Valdez. Special issue on algorithms in decision support systems vol. 2. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/11/512.

Pacella:2023:UMR

[2601] Massimo Pacella, Matteo Mangini, and Gabriele Papadia. Utilizing mixture regression models for clustering time-series energy consumption of a plastic injection molding process. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/16/11/524.

Rahman:2023:MLB

[2602] Md. Ashikur Rahman, Lway Faisal Abdulrazak, Md. Mamun Ali, Imran Mahmud, Kawsar Ahmed, and Francis M. Bui. Machine learning-based approach for predicting diabetes employing socio-demographic characteristics. Algorithms (Basel), 16(11):??, November 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/503.

Risser:2023:DPU

[2603] Laurent Risser, Agustin Martin Picard, Lucas Hervier, and Jean-Michel Loubes. Detecting and processing unsuspected sensitive variables for robust machine learning. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/11/510.

Rixen:2023:CML

[2604] Jöran Rixen, Nico Blass, Simon Lyra, and Steffen Leonhardt. Comparison of machine learning classifiers for the detection of breast cancer in an electrical impedance tomography setup. Algorithms (Basel), 16(11): ??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/11/517.

Rovito:2023:DNL

[2605] Luigi Rovito, Andrea De Lorenzo, and Luca Manzoni. Discovering nonlinear Boolean functions by evolving Walsh transforms with genetic programming. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/499.

Scribano:2023:DDM

[2606] Carmelo Scribano, Danilo Pezzi, Giorgia Franchini, and Marco Prato. Denoising diffusion models on model-based latent space. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/11/501.

Tallapally:2023:UGN

[2607] Dharahas Tallapally, John Wang, Katerina Potika, and Magdalini Eirinaki. Using graph neural networks for social recommendations. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
11/515.

Tzelepis:2023:IIR

[2608] Theodoros Tzelepis, George Matlis, Nikos Dimokas, Petros Karvelis, Paraskevi Malliou, and Anastasia Beneka. An intelligent injury rehabilitation guidance system for recreational runners using data mining algorithms. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/523.

Umair:2023:NAB

[2609] Sahibzada Muhammad Ahmad Umair and Tughrul Arslan. A narrowdown approach based on machine learning for indoor localization. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
11/529.

Vidakovic:2023:PAP

[2610] Marin Vidaković and Kruno Milicević. Performance and applicability of post-quantum digital signature algorithms in resource-constrained environments. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/11/518.

Wang:2023:RFA

[2611] Li-Na Wang, Guoqiang Zhong, Yaxin Shi, and Mohamed Cheriet. Relational Fisher analysis: Dimensionality reduction in relational data with global convergence. Algorithms (Basel), 16(11):??, November 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/11/522.

Ye:2023:PCO

[2612] Daoyu Ye and Shengxiang Deng. Period cycle optimization of integrated energy systems with long-term scheduling consideration. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/11/530.

Zhang:2023:IOD

[2613] Linhua Zhang, Ning Xiong, Xinghao Pan, Xiaodong Yue, Peng Wu, and Caiping Guo. Improved object detection method utilizing YOLOv7-tiny for unmanned aerial vehicle photographic imagery. Algorithms (Basel), 16(11):??, November 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/11/520.

Abiola:2023:NDR

[2614] Abiodun Abiola, Francisca Segura Manzano, and José Manuel Andújar. A novel deep reinforcement learning (DRL) algorithm to apply artificial intelligence-based maintenance in electrolysers. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/12/541.

Ali:2023:BCR

[2615] Mohamad Abou Ali, Fadi Dornaika, and Ignacio Arganda-Carreras. Blood cell revolution: Unveiling 11 distinct types with ‘naturalize’ augmentation. Algorithms (Basel), 16(12):??, December 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/562.

Alidaee:2023:ECF

[2616] Bahram Alidaee, Haibo Wang, and Lutfu S. Sua. An efficient closed-form formula for evaluating r-flip moves in quadratic unconstrained binary optimization. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/557.

Alzahrani:2023:NDL

[2617] A. Khuzaim Alzahrani, Ahmed A. Alsheikhy, Tawfeeq Shawly, Ahmed Azzahrani, and Yahia Said. A novel deep learning segmentation and classification framework for leukemia diagnosis. Algorithms (Basel), 16(12): ??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/12/556.

Amiri:2023:ILF

[2618] Farhad Amiri, Mohsen Eskandari, and Mohammad Hassan Moradi. Improved load frequency control in power systems hosting wind turbines by an augmented fractional order PID controller optimized by the powerful owl search algorithm. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/12/539.

Antonelli:2023:SISb

[2619] Laura Antonelli and Lucia Maddalena. Special issue on ‘Algorithms for Biomedical Image Analysis and Processing’. Algorithms (Basel), 16(12): ??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/12/544.

Ayub:2023:EOD

[2620] Nasir Ayub, Tayyaba, Saddam Hussain, Syed Sajid Ullah, and Jawaid Iqbal. An efficient optimized DenseNet model for aspect-based multilabel classification. Algorithms (Basel), 16(12):??, December 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/12/548.

Belcastro:2023:ECP

[2621] Loris Belcastro, Domenico Carbone, Cristian Cosentino, Fabrizio Marozzo, and Paolo Trunfio. Enhancing cryptocurrency price forecasting by integrating machine learning with social media and market data. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/542.

Bendebane:2023:MCD

[2622] Lamia Bendebane, Zakaria Laboudi, Asma Saighi, Hassan Al-Tarawneh, Adel Ouannas, and Giuseppe Grassi. A multi-class deep learning approach for early detection of depressive and anxiety disorders using Twitter data. Algorithms (Basel), 16(12):??, December 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/543.

Butt:2023:PID

[2623] Usman Javed Butt, Osama Hussien, Krison Hasanaj, Khaled Shaalan, Bilal Hassan, and Haider al Khateeb. Predicting the impact of data poisoning attacks in blockchain-enabled supply chain networks. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/12/549.

Chekroun:2023:DLB

[2624] Mohammed Chekroun, Youssef Mourchid, Igor Bessières, and Alain Lalande. Deep learning based on EfficientNet for multiorgan segmentation of thoracic structures on a 0.35 T MR-Linac radiation therapy system. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
16/12/564.

Cisneros:2023:ASH

[2625] Juan Cisneros, Alain Lalande, Binnaz Yalcin, Fabrice Meriaudeau, and Stephan Collins. Automatic segmentation of histological images of mouse brains. Algorithms (Basel), 16(12):??, December 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/553.

Daniilidou:2023:GFI

[2626] Athina Daniilidou, Avrilia Konguetsof, Georgios Souliotis, and Basil Papadopoulos. Generator of fuzzy implications. Algorithms (Basel), 16(12): ??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/12/569.

Espada:2023:FOD

[2627] Joaquim Espada, Alexandre P. Francisco, Tatiana Rocher, Luís M. S. Russo, and Cátia Vaz. On finding optimal (dynamic) arborescences. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/559.

Feng:2023:LGN

[2628] Miao Feng and Jean Meunier. A lightweight graph neural network algorithm for action recognition based on self-distillation. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/12/552.

Freijo:2023:RSD

[2629] Clara Freijo, Joaquin L. Herraiz, Fernando Arias-Valcayo, Paula Ibáñez, Gabriela Moreno, Amaia Villa-Abaunza, and José Manuel Udías. Robustness of single- and dual-energy deep-learning-based scatter correction models on simulated and real chest X-rays. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/12/565.

Gomez:2023:LAC

[2630] Juan F. Gomez, Antonio R. Uguina, Javier Panadero, and Angel A. Juan. A learnheuristic algorithm for the capacitated dispersion problem under dynamic conditions. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/12/532.

Han:2023:NDA

[2631] Xiaoyu Han, Chenyu Li, Zifan Wang, and Guohua Liu. NDARTS: A differentiable architecture search based on the Neumann series. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/12/536.

Hermes:2023:NPM

[2632] Jacques Hermes, Marcus Rosenblatt, Christian Tönsing, and Jens Timmer. Non-parametric model-based estimation of the effective reproduction number for SARS-CoV-2. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/16/12/533.

Juan:2023:SNH

[2633] Angel A. Juan, Markus Rabe, Majsa Ammouriova, Javier Panadero, David Peidro, and Daniel Riera. Solving NP-hard challenges in logistics and transportation under general uncertainty scenarios using fuzzy simheuristics. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/570.

Kalogiratou:2023:CTD

[2634] Zacharoula Kalogiratou and Theodoros Monovasilis. Construction of two-derivative Runge–Kutta methods of order six. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/12/558.

Khan:2023:FLL

[2635] Sarwar Shah Khan, Muzammil Khan, and Yasser Alharbi. Fast local Laplacian filter based on modified Laplacian through bilateral filter for coronary angiography medical imaging enhancement. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/12/531.

Khan:2023:VBC

[2636] Md. Al-Masrur Khan, Seong-Hoon Kee, and Abdullah-Al Nahid. Visionbased concrete-crack detection on railway sleepers using dense U-net model. Algorithms (Basel), 16(12):??, December 2023. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/568.

Lao:2023:PPT

[2637] Liming Lao, Dangkui Du, and Pengzhan Chen. Predicting pedestrian trajectories with deep adversarial networks considering motion and spatial information. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/566.

Majumder:2023:HDP

[2638] Annwesha Banerjee Majumder, Somsubhra Gupta, Dharmpal Singh, Biswaranjan Acharya, Vassilis C. Gerogiannis, Andreas Kanavos, and Panagiotis Pintelas. Heart disease prediction using concatenated hybrid ensemble classifiers. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/12/538.

Makaram:2023:DLB

[2639] Navaneethakrishna Makaram, Sarvagya Gupta, Matthew Pesce, Jeffrey Bolton, Scellig Stone, Daniel Haehn, Marc Pomplun, Christos Papadelis, Phillip Pearl, Alexander Rotenberg, Patricia Ellen Grant, and Eleonora Tamilia. Deep learning-based visual complexity analysis of electroencephalography time-frequency images: Can it localize the epileptogenic zone in the brain? Algorithms (Basel), 16(12):??, December 2023. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/16/12/567.

Mavrovouniotis:2023:MPA

[2640] Michalis Mavrovouniotis, Maria N. Anastasiadou, and Diofantos Hadjimitsis. Measuring the performance of ant colony optimization algorithms for the dynamic traveling salesman problem. Algorithms (Basel), 16(12): ??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/12/545.

Nigro:2023:ICA

[2641] Libero Nigro and Franco Cicirelli. Improving clustering accuracy of Kmeans and random swap by an evolutionary technique based on careful seeding. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/572.

Nikitas:2023:AAU

[2642] Panos Nikitas and Efthymia Nikita. Assessing algorithms used for constructing confidence ellipses in multidimensional scaling solutions. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/535.

Oluwasakin:2023:OPI

[2643] Ebenezer O. Oluwasakin and Abdul Q. M. Khaliq. Optimizing physicsinformed neural network in dynamic system simulation and learning of parameters. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/547.

Pantale:2023:CAF

[2644] Olivier Pantalé. Comparing activation functions in machine learning for finite element simulations in thermomechanical forming. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/16/12/537.

Porras:2023:CSC

[2645] José Manuel Porras, Juan Alfonso Lara, Cristóbal Romero, and Sebastián Ventura. A case-study comparison of machine learning approaches for predicting student’s dropout from multiple online educational entities. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/554.

Rosyida:2023:ACP

[2646] Isnaini Rosyida and Christiana Rini Indrati. An algorithm for coloring of picture fuzzy graphs based on strong and weak adjacencies, and its application. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/551.

Scrucca:2023:IDI

[2647] Luca Scrucca. On the influence of data imbalance on supervised Gaussian mixture models. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/563.

Sinitsyn:2023:EFM

[2648] Dmitry O. Sinitsyn, Alexandra G. Poydasheva, Ilya S. Bakulin, Alfiia H. Zabirova, Dmitry Yu. Lagoda, Natalia A. Suponeva, and Michael A. Piradov. Estimating the frequencies of maximal theta-gamma coupling in EEG during the N-back task: Sensitivity to methodology and temporal instability. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/540.

Trinh:2023:DDB

[2649] Cindy Trinh, Silvia Lasala, Olivier Herbinet, and Dimitrios Meimaroglou. On the development of descriptor-based machine learning models for thermodynamic properties: Part 2 — applicability domain and outliers. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/573.

Villaescusa:2023:EAB

[2650] Sara Lledó Villaescusa and Rafael Lahoz-Beltra. Evolutionary algorithms in a bacterial consortium of synthetic bacteria. Algorithms (Basel), 16 (12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/16/12/571.

Vrtal:2023:TDU

[2651] Matej Vrtal, Radek Fujdiak, Jan Benedikt, Pavel Praks, Radim Bris, Michal Ptacek, and Petr Toman. Time-dependent unavailability exploration of interconnected urban power grid and communication network. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/561.

Wang:2023:DEC

[2652] Li-Na Wang, Hongxu Wei, Yuchen Zheng, Junyu Dong, and Guoqiang Zhong. Deep error-correcting output codes. Algorithms (Basel), 16(12): ??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/16/12/555.

Wang:2023:IDB

[2653] Yi Wang, Yating Xu, Tianjian Li, Tao Zhang, and Jian Zou. Image deblurring based on convex non-convex sparse regularization and plug-and-play algorithm. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/16/12/574.

Wang:2023:SDA

[2654] Langyu Wang, Yan Zhang, Yahong Lin, Shuai Yan, Yuanyuan Xu, and Bo Sun. Ship detection algorithm based on YOLOv5 network improved with lightweight convolution and attention mechanism. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/534.

Zhang:2023:OST

[2655] Xiaobo Zhang, Huashun Li, Jingzhao Li, and Xuehai Zhou. OrthoDETR: A streamlined transformer-based approach for precision detection of orthopedic medical devices. Algorithms (Basel), 16(12):??, December 2023.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/12/550.

Zhang:2023:WTP

[2656] Shuo Zhang, Emma Robinson, and Malabika Basu. Wind turbine predictive fault diagnostics based on a novel long short-term memory model. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/16/
12/546.

Zhao:2023:SOD

[2657] Kexin Zhao, Rui Jiang, and Jun He. Stereo 3D object detection using a feature attention module. Algorithms (Basel), 16(12):??, December 2023. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/16/12/560.

Huang:2024:UMR

[2658] Jih-Jeng Huang and Chin-Yi Chen. Using Markov random field and analytic hierarchy process to account for interdependent criteria. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/1/1.

Angrisano:2024:VEU

[2659] Antonio Angrisano, Giovanni Cappello, Salvatore Gaglione, and Ciro Gioia. Velocity estimation using time-differenced carrier phase and Doppler shift with different grades of devices: From smartphones to professional receivers. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/2.

Muhsen:2024:SSR

[2660] Dena Kadhim Muhsen, Ahmed T. Sadiq, and Firas Abdulrazzaq Raheem. A survey on swarm robotics for area coverage problem. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/1/3.

Gutenschwager:2024:CDD

[2661] Kai Gutenschwager, Markus Rabe, and Jorge Chicaiza-Vaca. Comparing direct deliveries and automated parcel locker systems with respect to overall CO_{2} emissions for the last mile. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/4.

Tsoulos:2024:SIS

[2662] Ioannis G. Tsoulos. Special issue “Algorithms in Data Classification”. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/1/5.

Aikodon:2024:PDR

[2663] Nosa Aikodon, Sandra Ortega-Martorell, and Ivan Olier. Predicting decompensation risk in intensive care unit patients using machine learning. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/1/
6.

Gaipov:2024:HGG

[2664] Konstantin Gaipov, Daniil Tausnev, Sergey Khodenkov, Natalya Shepeta, Dmitry Malyshev, Aleksey Popov, and Lev Kazakovtsev. Heuristic greedy-gradient route search method for finding an optimal traffic distribution in telecommunication networks. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/7.

Prezja:2024:EEB

[2665] Fabi Prezja, Leevi Annala, Sampsa Kiiskinen, and Timo Ojala. Exploring the efficacy of base data augmentation methods in deep learning-based radiograph classification of knee joint osteoarthritis. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/8.

Hsieh:2024:SAM

[2666] Fu-Shiung Hsieh. A self-adaptive meta-heuristic algorithm based on success rate and differential evolution for improving the performance of ridesharing systems with a discount guarantee. Algorithms (Basel), 17 (1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/9.

Legrand:2024:EDA

[2667] François Legrand, Richard Macwan, Alain Lalande, Lisa Métairie, and Thomas Decourselle. Effect of data augmentation on deep-learning-based segmentation of long-axis cine-MRI. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/1/10.

Manca:2024:AOM

[2668] Vincenzo Manca. The Archimedean origin of modern positional number systems. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/11.

DellAmico:2024:CMS

[2669] Mauro Dell’Amico, Jafar Jamal, and Roberto Montemanni. Compact models to solve the precedence-constrained minimum-cost arborescence problem with waiting times. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/1/12.

Alghanim:2024:MLM

[2670] Firas Alghanim, Ibrahim Al-Hurani, Hazem Qattous, Abdullah Al-Refai, Osamah Batiha, Abedalrhman Alkhateeb, and Salama Ikki. Machine learning model for multiomics biomarkers identification for menopause status in breast cancer. Algorithms (Basel), 17(1), January 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/1/13.

Gao:2024:ICM

[2671] Depeng Gao, Shuai Wang, Yuwei Yang, Haifei Zhang, Hao Chen, Xiangxiang Mei, Shuxi Chen, and Jianlin Qiu. An intelligent control method for servo motor based on reinforcement learning. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/14.

Branco:2024:CRT

[2672] António Pedro Branco, Cátia Vaz, and Alexandre P. Francisco. Computing RF tree distance over succinct representations. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/15.

Kamiwaki:2024:ELE

[2673] Yuto Kamiwaki and Shinji Fukuda. Effect of the light environment on image-based SPAD value prediction of radish leaves. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/16.

Pellicer:2024:AEF

[2674] David S. Pellicer and Emilio Larrodé. Analysis of the effectiveness of a freight transport vehicle at high speed in a vacuum tube (hyperloop transport system). Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/17.

Rezvanian:2024:CCA

[2675] Alireza Rezvanian, S. Mehdi Vahidipour, and Ali Mohammad Saghiri. CaAIS: Cellular automata-based artificial immune system for dynamic environments. Algorithms (Basel), 17(1), January 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/18.

Corceiro:2024:LML

[2676] Ana Corceiro, Nuno Pereira, Khadijeh Alibabaei, and Pedro D. Gaspar. Leveraging machine learning for weed management and crop enhancement: Vineyard flora classification. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/1/19.

Sabatier:2024:AFD

[2677] Jocelyn Sabatier and Christophe Farges. Algorithms for fractional dynamical behaviors modelling using non-singular rational kernels. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/1/20.

Ionescu:2024:EUA

[2678] Stefan Ionescu, Camelia Delcea, Nora Chirita, and Ionut Nica. Exploring the use of artificial intelligence in agent-based modeling applications: a bibliometric study. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/21.

Pau:2024:TFF

[2679] Danilo Pau, Andrea Pisani, and Antonio Candelieri. Towards full forward on-tiny-device learning: a guided search for a randomly initialized neural network. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/22.

Santana:2024:NAI

[2680] Dthenifer Cordeiro Santana, Gustavo de Faria Theodoro, Ricardo Gava, João Lucas Gouveia de Oliveira, Larissa Pereira Ribeiro Teodoro, Izabela Cristina de Oliveira, Fábio Henrique Rojo Baio, Carlos Antonio da Silva Junior, Job Teixeira de Oliveira, and Paulo Eduardo Teodoro.

A new approach to identifying sorghum hybrids using UAV imagery using multispectral signature and machine learning. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/23.

Scutari:2024:EKL

[2681] Marco Scutari. Entropy and the Kullback–Leibler divergence for Bayesian networks: Computational complexity and efficient implementation. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/24.

Panagiotakis:2024:PSO

[2682] Costas Panagiotakis. Particle swarm optimization-based unconstrained polygonal fitting of 2D shapes. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/1/25.

El-Mageed:2024:HSS

[2683] Amr A. Abd El-Mageed, Ayoub Al-Hamadi, Samy Bakheet, and Asmaa H. Abd El-Rahiem. Hybrid sparrow search-exponential distribution optimization with differential evolution for parameter prediction of solar photovoltaic models. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/26.

Sakalauskas:2024:PAC

[2684] Virgilijus Sakalauskas and Dalia Kriksciuniene. Personalized advertising in e-commerce: Using clickstream data to target high-value customers. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/1/27.

Fan:2024:SMB

[2685] Yiming Fan and Meng Wang. Specification mining based on the ordering points to identify the clustering structure clustering algorithm and model checking. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/28.

Tatit:2024:NME

[2686] Pornrawee Tatit, Kiki Adhinugraha, and David Taniar. Navigating the maps: Euclidean vs. road network distances in spatial queries. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/1/29.

Maksymov:2024:QIN

[2687] Ivan S. Maksymov. Quantum-inspired neural network model of optical illusions. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/30.

Haque:2024:GAS

[2688] Sardar Anisul Haque, Mohammad Tanvir Parvez, and Shahadat Hossain. GPU algorithms for structured sparse matrix multiplication with diagonal storage schemes. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/31.

Begum:2024:IWU

[2689] Mahbuba Begum, Sumaita Binte Shorif, Mohammad Shorif Uddin, Jannatul Ferdush, Tony Jan, Alistair Barros, and Md Whaiduzzaman. Image watermarking using discrete wavelet transform and singular value decomposition for enhanced imperceptibility and robustness. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/1/32.

Khan:2024:QIP

[2690] Azal Ahmad Khan, Salman Hussain, and Rohitash Chandra. A quantuminspired predator-prey algorithm for real-parameter optimization. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/1/33.

Shokouhifar:2024:EHM

[2691] Mohammad Shokouhifar, Mohamad Hasanvand, Elaheh Moharamkhani, and Frank Werner. Ensemble heuristic-metaheuristic feature fusion learning for heart disease diagnosis using tabular data. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/34.

Cellupica:2024:IDT

[2692] Alessio Cellupica, Marco Cirelli, Giovanni Saggio, Emanuele Gruppioni, and Pier Paolo Valentini. An interactive digital-twin model for virtual reality environments to train in the use of a sensorized upper-limb prosthesis. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/35.

Li:2024:RVE

[2693] Zheng Li, Xinkai Chen, Jiaqing Fu, Ning Xie, and Tingting Zhao. Reducing Q-value estimation bias via mutual estimation and softmax operation in MADRL. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/36.

Li:2024:ORL

[2694] Jiaming Li, Ning Xie, and Tingting Zhao. Optimizing reinforcement learning using a generative action-translator transformer. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/1/37.

Amyot:2024:AFA

[2695] Romain Amyot, Noriyuki Kodera, and Holger Flechsig. Atom filtering algorithm and GPU-accelerated calculation of simulation atomic force microscopy images. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/38.

Su:2024:PDB

[2696] Jiao Su, Yi An, Jialin Wu, and Kai Zhang. Pedestrian detection based on feature enhancement in complex scenes. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/39.

Kirpichenko:2024:BBE

[2697] Stanislav Kirpichenko, Lev Utkin, Andrei Konstantinov, and Vladimir Muliukha. BENK: The Beran Estimator with Neural Kernels for estimating the heterogeneous treatment effect. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/40.

Ho:2024:EMO

[2698] Ken Jom Ho, Ender Özcan, and Peer-Olaf Siebers. Efficient multiobjective simulation metamodeling for researchers. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/1/41.

Zhong:2024:DEB

[2699] Deyuan Zhong, Liangda Fang, and Quanlong Guan. Dictionary encoding based on tagged sentential decision diagrams. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/42.

Heberger:2024:FEM

[2700] Károly Héberger. Frequent errors in modeling by machine learning: a prototype case of predicting the timely evolution of COVID-19 pandemic. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/1/43.

Tsoulos:2024:NAI

[2701] Ioannis G. Tsoulos and V. N. Stavrou. Numerical algorithms in III– V semiconductor heterostructures. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/1/44.

Zhang:2024:DDD

[2702] Yuzhu Zhang and Hao Xu. Distributed data-driven learning-based optimal dynamic resource allocation for multi-RIS-assisted multi-user ad-hoc network. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/1/45.

Neroni:2024:BRD

[2703] Mattia Neroni, Massimo Bertolini, and Angel A. Juan. A biasedrandomized discrete event algorithm to improve the productivity of automated storage and retrieval systems in the steel industry. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/1/46.

Hojas-Mazo:2024:FBS

[2704] Wenny Hojas-Mazo, Francisco Maciá-Pérez, José Vicente Berná Martínez, Mailyn Moreno-Espino, Iren Lorenzo Fonseca, and Juan Pavón. Framework based on simulation of real-world message streams to evaluate classification solutions. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/1/47.

Ganguli:2024:ABP

[2705] Tushar Ganguli and Edwin K. P. Chong. Activation-based pruning of neural networks. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/1/48.

Nayyeri:2024:CBC

[2706] Mojtaba Nayyeri, Modjtaba Rouhani, Hadi Sadoghi Yazdi, Marko M. Mäkelä, Alaleh Maskooki, and Yury Nikulin. Correntropy-based constructive one hidden layer neural network. Algorithms (Basel), 17(1), January 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/1/49.

Volkov:2024:VNA

[2707] Konstantin Volkov. Vectorized numerical algorithms to solve internal problems of computational fluid dynamics. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/2/50.

deCampos:2024:IRM

[2708] Luis M. de Campos, Juan M. Fernández-Luna, Juan F. Huete, Francisco J. Ribadas-Pena, and Néstor Bolaños. Information retrieval and machine learning methods for academic expert finding. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/2/51.

Juan:2024:ECE

[2709] Pin-Hung Juan and Ja-Ling Wu. Enhancing communication efficiency and training time uniformity in federated learning through multi-branch networks and the Oort algorithm. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/2/52.

Sulaiman:2024:PSS

[2710] Abdullahi T. Sulaiman, Habeeb Bello-Salau, Adeiza J. Onumanyi, Muhammed B. Mu’azu, Emmanuel A. Adedokun, Ahmed T. Salawudeen, and Abdulfatai D. Adekale. A particle swarm and smell agent-based hybrid algorithm for enhanced optimization. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/2/53.

Du:2024:SCE

[2711] Yunkang Du and Zuoliang Xu. Simultaneous calibration of European option volatility and fractional order under the time fractional Vasicek model. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/54.

Hajnal:2024:BNS

[2712] Péter Hajnal. Binary numeration system with alternating signed digits and its graph theoretical relationship. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/2/55.

Marum:2024:FWP

[2713] João Paulo Oliveira Marum, H. Conrad Cunningham, J. Adam Jones, and Yi Liu. Following the writer’s path to the dynamically coalescing reactive chains design pattern. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/2/56.

Supajaidee:2024:AMW

[2714] Nattakan Supajaidee, Nawinda Chutsagulprom, and Sompop Moonchai. An adaptive moving window kriging based on K-means clustering for spatial interpolation. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/57.

Mao:2024:FOF

[2715] Wei-Lung Mao, Sung-Hua Chen, and Chun-Yu Kao. Fractional-order fuzzy PID controller with evolutionary computation for an effective synchronized gantry system. Algorithms (Basel), 17(2), February 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/2/58.

Vadla:2024:EPD

[2716] Mahammad Khalid Shaik Vadla, Mahima Agumbe Suresh, and Vimal K. Viswanathan. Enhancing product design through AI-driven sentiment analysis of Amazon reviews using BERT. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/2/59.

Wang:2024:LSS

[2717] Ziyi Wang, Xinran Li, Luoyang Sun, Haifeng Zhang, Hualin Liu, and Jun Wang. Learning state-specific action masks for reinforcement learning. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
2/60.

Fotakis:2024:ETS

[2718] Dimitris Fotakis, Panagiotis Patsilinakos, Eleni Psaroudaki, and Michalis Xefteris. Efficient time-series clustering through sparse Gaussian modeling. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/61.

Idrisov:2024:PCG

[2719] Baskhad Idrisov and Tim Schlippe. Program code generation with generative AIs. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/62.

Cavallaro:2024:GDM

[2720] Claudia Cavallaro, Carolina Crespi, Vincenzo Cutello, Mario Pavone, and Francesco Zito. Group dynamics in memory-enhanced ant colonies: The influence of colony division on a maze navigation problem. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/2/63.

More:2024:EID

[2721] Shweta More, Moad Idrissi, Haitham Mahmoud, and A. Taufiq Asyhari. Enhanced intrusion detection systems performance with UNSW-NB15 data analysis. Algorithms (Basel), 17(2), February 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/64.

Werner:2024:SIS

[2722] Frank Werner. Special issue: “2022 and 2023 Selected Papers from Algorithms’ Editorial Board Members”. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/2/65.

Orti:2024:GSA

[2723] José Antonio López Ortí, Francisco José Marco Castillo, and María José Martínez Usó. A geometrical study about the biparametric family of anomalies in the elliptic two-body problem with extensions to other families. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/66.

Durasevic:2024:AAG

[2724] Marko Durasević, Domagoj Jakobović, Stjepan Picek, and Luca Mariot. Assessing the ability of genetic programming for feature selection in constructing dispatching rules for unrelated machine environments. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/2/67.

Zhu:2024:HAC

[2725] Kefei Zhu, Xu Yang, Yanbo Zhang, Mengkun Liang, and Jun Wu. A heterogeneity-aware car-following model: Based on the XGBoost method. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/68.

Li:2024:FAD

[2726] Wenjun Li, Xueying Yang, Chao Xu, and Yongjie Yang. An FPT algorithm for directed co-graph edge deletion. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/2/69.

Sandu:2024:NDL

[2727] Andra Sandu, Ioana Ioanas, Camelia Delcea, Margareta-Stela Florescu, and Liviu-Adrian Cotfas. Numbers do not lie: a bibliometric examination of machine learning techniques in fake news research. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/2/70.

Tassiopoulou:2024:ATR

[2728] Styliani Tassiopoulou, Georgia Koukiou, and Vassilis Anastassopoulos. Algorithms in tomography and related inverse problems — a review. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
2/71.

Lazzarinetti:2024:ABM

[2729] Giorgio Lazzarinetti, Riccardo Dondi, Sara Manzoni, and Italo Zoppis. An attention-based method for the minimum vertex cover problem on complex networks. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/72.

Moutsopoulou:2024:ASC

[2730] Amalia Moutsopoulou, Markos Petousis, Georgios E. Stavroulakis, Anastasios Pouliezos, and Nectarios Vidakis. μ-analysis and μ-synthesis control methods in smart structure disturbance suppression with reduced order control. Algorithms (Basel), 17(2), February 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/73.

Milanic:2024:GAD

[2731] Matija Milanic and Rok Hren. GPU adding-doubling algorithm for analysis of optical spectral images. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/2/74.

Alderighi:2024:AIP

[2732] Stefano Alderighi, Paolo Landa, Elena Tànfani, and Angela Testi. Assessing the impact of patient characteristics on genetic clinical pathways: a regression approach. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/75.

Munoz-Zavala:2024:LRS

[2733] Angel E. Muñoz-Zavala, Jorge E. Macías-Díaz, Daniel Alba-Cuéllar, and José A. Guerrero-Díaz de León. A literature review on some trends in artificial neural networks for modeling and simulation with time series. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
2/76.

Puchalski:2024:AUC

[2734] Jacek G. Puchalski, Janusz D. Fidelus, and Paweł Fotowicz. Algorithms utilized for creep analysis in torque transducers for wind turbines. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/2/77.

Naser:2024:RML

[2735] Marwah Abdulrazzaq Naser, Aso Ahmed Majeed, Muntadher Alsabah, Taha Raad Al-Shaikhli, and Kawa M. Kaky. A review of machine learning’s role in cardiovascular disease prediction: Recent advances and future challenges. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/78.

Wang:2024:RGD

[2736] Haojie Wang, Pingqing Fan, Xipei Ma, and Yansong Wang. Research on gangue detection algorithm based on cross-scale feature fusion and dynamic pruning. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/79.

Nardin:2024:QIA

[2737] Antonella Nardin and Fabio D’Andreagiovanni. A quantum-inspired ant colony optimization algorithm for parking lot rental to shared e-scooter services. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/80.

Wnuk:2024:AAA

[2738] Marian Wnuk. Adaptive antenna array control algorithm in radiocommunication systems. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/81.

Jeng:2024:OMP

[2739] Mingyoung Jeng, Alvir Nobel, Vinayak Jha, David Levy, Dylan Kneidel, Manu Chaudhary, Ishraq Islam, Evan Baumgartner, Eade Vanderhoof, Audrey Facer, Manish Singh, Abina Arshad, and Esam El-Araby. Optimizing multidimensional pooling for variational quantum algorithms. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
2/82.

Buscher:2024:CSI

[2740] Keno Jann Büscher, Jan Philipp Degel, and Jan Oellerich. A comprehensive survey of isocontouring methods: Applications, limitations and perspectives. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/2/83.

Htun:2024:MDH

[2741] Nyo Me Htun, Toshiaki Owari, Satoshi Tsuyuki, and Takuya Hiroshima. Mapping the distribution of high-value broadleaf tree crowns through unmanned aerial vehicle image analysis using deep learning. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/2/84.

Maphosa:2024:IAA

[2742] Mfowabo Maphosa, Wesley Doorsamy, and Babu Paul. Improving academic advising in engineering education with machine learning using a real-world dataset. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/85.

Zafar:2024:NHO

[2743] Fiza Zafar, Alicia Cordero, Husna Maryam, and Juan R. Torregrosa. A novel higher-order numerical scheme for system of nonlinear load flow equations. Algorithms (Basel), 17(2), February 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/86.

Souza:2024:TRL

[2744] Gleice Kelly Barbosa Souza, Samara Oliveira Silva Santos, André Luiz Carvalho Ottoni, Marcos Santos Oliveira, Daniela Carine Ramires Oliveira, and Erivelton Geraldo Nepomuceno. Transfer reinforcement learning for combinatorial optimization problems. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/2/87.

Guo:2024:ALP

[2745] Lin Guo, Anand Balu Nellippallil, Warren F. Smith, Janet K. Allen, and Farrokh Mistree. An adaptive linear programming algorithm with parameter learning. Algorithms (Basel), 17(2), February 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/2/88.

Pascoe:2024:UDE

[2746] Sean Pascoe. On the use of data envelopment analysis for multi-criteria decision analysis. Algorithms (Basel), 17(3), March 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/89.

Tyagi:2024:OSE

[2747] Suryakant Tyagi and Sándor Szénási. Optimizing speech emotion recognition with deep learning and grey wolf optimization: a multi-dataset approach. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/3/90.

Wang:2024:MAB

[2748] Jie Wang, Jie Yang, Jiafan He, and Dongliang Peng. Multiaugmentation-based contrastive learning for semi-supervised learning. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/91.

Lu:2024:IIS

[2749] Hexin Lu, Xiaodong Zhu, Jingwei Cui, and Haifeng Jiang. An iris image super-resolution model based on Swin transformer and generative adversarial network. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/92.

Dawid:2024:WCG

[2750] Philip Dawid. What is a causal graph? Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/3/93.

Hamza:2024:IDF

[2751] Mukhtar Fatihu Hamza. Improved decentralized fractional-order control of higher-order systems using modified flower pollination optimization. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/94.

Lange:2024:CFW

[2752] Kenneth Lange. Closest farthest widest. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/95.

Rybczak:2024:DML

[2753] Monika Rybczak and Krystian Kozakiewicz. Deep machine learning of MobileNet, efficient, and inception models. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/96.

Pedraza:2024:DNN

[2754] Anibal Pedraza, Lucia Gonzalez, Oscar Deniz, and Gloria Bueno. Deep neural networks for HER2 grading of whole slide images with subclasses levels. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/

Paraschos:2024:RLB

[2755] Panagiotis D. Paraschos, Georgios K. Koulinas, and Dimitrios E. Koulouriotis. Reinforcement learning-based optimization for sustainable and lean production within the context of Industry 4.0. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/3/98.

Das:2024:ESU

[2756] Saikat Das, Mohammad Ashrafuzzaman, Frederick T. Sheldon, and Sajjan Shiva. Ensembling supervised and unsupervised machine learning algorithms for detecting distributed denial of service attacks. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/3/99.

Adriani:2024:CDA

[2757] Andrea Adriani, Stefano Serra-Capizzano, and Cristina Tablino-Possio. Clustering/distribution analysis and preconditioned Krylov solvers for the approximated Helmholtz equation and fractional Laplacian in the case of complex-valued, unbounded variable coefficient wave number μ. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/100.

Zhassuzak:2024:AGA

[2758] Mukhtar Zhassuzak, Marat Akhmet, Yedilkhan Amirgaliyev, and Zholdas Buribayev. Application of genetic algorithms for periodicity recognition and finite sequences sorting. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/101.

Jing:2024:RCT

[2759] Fengwei Jing, Fenghe Li, Yong Song, Jie Li, Zhanbiao Feng, and Jin Guo. Root cause tracing using equipment process accuracy evaluation for looper in hot rolling. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/102.

UlAinTahir:2024:ODA

[2760] Noor Ul Ain Tahir, Zuping Zhang, Muhammad Asim, Junhong Chen, and Mohammed ELAffendi. Object detection in autonomous vehicles under adverse weather: a review of traditional and deep learning approaches. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/3/103.

Lalapura:2024:SER

[2761] Varsha S. Lalapura, Veerender Reddy Bhimavarapu, J. Amudha, and Hariram Selvamurugan Satheesh. A systematic evaluation of recurrent neural network models for edge intelligence and human activity recognition applications. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/104.

Chumachenko:2024:AIA

[2762] Dmytro Chumachenko and Sergiy Yakovlev. Artificial intelligence algorithms for healthcare. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/105.

Shahrabadi:2024:AOD

[2763] Somayeh Shahrabadi, Telmo Adão, Emanuel Peres, Raul Morais, Luís G. Magalhães, and Victor Alves. Automatic optimization of deep learning training through feature-aware-based dataset splitting. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/3/106.

Kolosnitsyn:2024:APM

[2764] Anton Kolosnitsyn, Oleg Khamisov, Eugene Semenkin, and Vladimir Nelyub. Application of the parabola method in nonconvex optimization. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/107.

Caballero:2024:DMT

[2765] Pablo Caballero, Luis Gonzalez-Abril, Juan A. Ortega, and Áurea Simon-Soro. Data mining techniques for endometriosis detection in a datascarce medical dataset. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/108.

Wu:2024:ASC

[2766] Yuhuan Wu and Yonghong Wu. Application of split coordinate channel attention embedding U2Net in salient object detection. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/3/109.

Che:2024:EVO

[2767] Shuang Che, Yan Chen, Longda Wang, and Chuanfang Xu. Electric vehicle ordered charging planning based on improved dual-population genetic moth-flame optimization. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/3/110.

Pendharkar:2024:MCG

[2768] Parag C. Pendharkar. A Markov chain genetic algorithm approach for non-parametric posterior distribution sampling of regression parameters. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/111.

Hubai:2024:EDA

[2769] András Hubai, Sándor Szabó, and Bogdán Zaválnij. Exploratory data analysis and searching cliques in graphs. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/112.

Zhao:2024:ENN

[2770] Anni Zhao, Arash Toudeshki, Reza Ehsani, Joshua H. Viers, and Jian-Qiao Sun. Evaluation of neural network effectiveness on sliding mode control of delta robot for trajectory tracking. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/113.

Reshadi:2024:DSM

[2771] MohammadHossein Reshadi, Wen Li, Wenjie Xu, Precious Omashor, Albert Dinh, Scott Dick, Yuntong She, and Michael Lipsett. Deep-shallow metaclassifier with synthetic minority oversampling for anomaly detection in a time series. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/ 1999-4893/17/3/114. See correction [3049].

Saleem:2024:IIH

[2772] Sharoon Saleem, Fawad Hussain, and Naveed Khan Baloch. IWO–IGA — a hybrid whale optimization algorithm featuring improved genetic characteristics for mapping real-time applications onto 2D network on chip. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/3/
115.

Laffitte:2024:PMA

[2773] Marcos E. González Laffitte and Peter F. Stadler. Progressive multiple alignment of graphs. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/116.

Cui:2024:FPG

[2774] Bo Cui, Lingyun Wang, Guangxi Li, and Xian Ren. Field programmable gate array-based acceleration algorithm design for dynamic star map parallel computing. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/117.

Parr:2024:ADS

[2775] Thomas Parr, Karl Friston, and Peter Zeidman. Active data selection and information seeking. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/118.

Li:2024:PMC

[2776] Yanjun Li, Takaaki Yoshimura, Yuto Horima, and Hiroyuki Sugimori. A preprocessing method for coronary artery stenosis detection based on deep learning. Algorithms (Basel), 17(3), March 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/119.

Vo:2024:EEG

[2777] Minh-Quan Vo, Thu Nguyen, Michael A. Riegler, and Hugo L. Hammer. Efficient estimation of generative models using Tukey depth. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/3/120.

Kyurkchiev:2024:MSC

[2778] Nikolay Kyurkchiev, Tsvetelin Zaevski, Anton Iliev, Vesselin Kyurkchiev, and Asen Rahnev. Modeling of some classes of extended oscillators: Simulations, algorithms, generating chaos, and open problems. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/3/121.

Si:2024:HIC

[2779] Xiaonan Si, Lei Wang, Wenchang Xu, Biao Wang, and Wenbo Cheng. Highly imbalanced classification of gout using data resampling and ensemble method. Algorithms (Basel), 17(3), March 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/122.

Abdul-Hassan:2024:ETO

[2780] Noori Y. Abdul-Hassan, Zainab J. Kadum, and Ali Hasan Ali. An efficient third-order scheme based on Runge–Kutta and Taylor series expansion for solving initial value problems. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/123.

Belaroussi:2024:EVE

[2781] Rachid Belaroussi, Elie Issa, Leonardo Cameli, Claudio Lantieri, and Sonia Adelé. Exploring virtual environments to assess the quality of public spaces. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/3/124.

Xie:2024:GGD

[2782] Xuanyuan Xie and Jieyu Zhao. GDUI: Guided diffusion model for unlabeled images. Algorithms (Basel), 17(3), March 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/3/125.

Krivovichev:2024:ATS

[2783] Gerasim V. Krivovichev and Valentina Yu. Sergeeva. Analysis of a twostep gradient method with two momentum parameters for strongly convex unconstrained optimization. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/3/126.

Levner:2024:FAH

[2784] Eugene Levner, Vladimir Kats, Pengyu Yan, and Ada Che. Fast algorithm for high-throughput screening scheduling based on the PERT/ CPM project management technique. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/3/127.

Franz:2024:NAB

[2785] Karly S. Franz, Grace Reszetnik, and Tom Chau. On the need for accurate brushstroke segmentation of tablet-acquired kinematic and pressure data: The case of unconstrained tracing. Algorithms (Basel), 17 (3), March 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/3/128.

Tian:2024:PFI

[2786] Xin Tian and Yuan Meng. PDEC: a framework for improving knowledge graph reasoning performance through predicate decomposition. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/3/129.

Dizaji:2024:CBM

[2787] Navid Khalili Dizaji and Mustafa Dogan. A comprehensive brain MRI image segmentation system based on contourlet transform and deep neural networks. Algorithms (Basel), 17(3), March 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/3/130.

Moldovan:2024:CPA

[2788] Ionut Dragos Moldovan, Abdalla Almukashfi, and António Gomes Correia. A computational platform for automatic signal processing for bender element sensors. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/131.

Wagner:2024:IDP

[2789] Torrey Wagner, Dennis Guhl, and Brent Langhals. The impact of data preparation and model complexity on the natural language classification of Chinese news headlines. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/4/132.

Romero-Gonzalez:2024:BSD

[2790] Julio-Alejandro Romero-González, Diana-Margarita Córdova-Esparza, Juan Terven, Ana-Marcela Herrera-Navarro, and Hugo Jiménez-Hernández. Background subtraction for dynamic scenes using Gabor filter bank and statistical moments. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/4/133.

Gabriel:2024:IMM

[2791] Ana Teresa Gabriel, Cláudia Quaresma, and Pedro Vieira. An innovative mathematical model of the spine: Predicting Cobb and intervertebral angles using the 3D position of the spinous processes measured by vertebral metrics. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/

Davoodi:2024:TCL

[2792] Mansoor Davoodi and Justin M. Calabrese. Test center location problem: a bi-objective model and algorithms. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/4/135.

Combs:2024:UVG

[2793] Kara Combs, Adam Moyer, and Trevor J. Bihl. Uncertainty in visual generative AI. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/4/136.

Markus:2024:TEA

[2794] Nenad Markus and Mirko Suznjević. Theoretical and empirical analysis of a fast algorithm for extracting polygons from signed distance bounds. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/137.

Ganopoulou:2024:DCD

[2795] Maria Ganopoulou, Efstratios Kontopoulos, Konstantinos Fokianos, Dimitris Koparanis, Lefteris Angelis, Ioannis Kotsianidis, and Theodoros Moysiadis. Delving into causal discovery in health-related quality of life questionnaires. Algorithms (Basel), 17(4), April 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/138.

Lee:2024:TVB

[2796] Alvin Lee, Suet-Peng Yong, Witold Pedrycz, and Junzo Watada. Testing a vision-based autonomous drone navigation model in a forest environment. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/
139.

Zolfaghari:2024:MVR

[2797] Hussein Zolfaghari, Hossein Karimi, Amin Ramezani, and Mohammadreza Davoodi. Minimizing voltage ripple of a DC microgrid via a particle-swarm-optimization-based fuzzy controller. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/4/140.

Soltan:2024:CRB

[2798] Armin Soltan and Peter Washington. Challenges in reducing bias using post-processing fairness for breast cancer stage classification with deep learning. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/4/141.

Fuladi:2024:DEF

[2799] Shubhendu Kshitij Fuladi and Chang-Soo Kim. Dynamic events in the flexible job-shop scheduling problem: Rescheduling with a hybrid metaheuristic algorithm. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/142.

Klawonn:2024:OFB

[2800] Frank Klawonn and Georg Hoffmann. An objective function-based clustering algorithm with a closed-form solution and application to reference interval estimation in laboratory medicine. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/4/143.

Silva:2024:AIE

[2801] Hugo Silva, Vítor Duque, Mário Macedo, and Mateus Mendes. Aiding ICD-10 encoding of clinical health records using improved text cosine similarity and PLM-ICD. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/144.

Xue:2024:PSB

[2802] Sen Xue, Chengyu Wu, Jing Han, and Ao Zhan. PDASTSGAT: an STSGAT-based multipath data scheduling algorithm. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/4/145.

Dogan:2024:MOB

[2803] Vedat Dogan and Steven Prestwich. Multi-objective BiLevel optimization by Bayesian optimization. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/4/146.

Moreno-Carbonell:2024:PLR

[2804] Santiago Moreno-Carbonell and Eugenio F. Sánchez-Úbeda. A piecewise linear regression model ensemble for large-scale curve fitting. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/4/147.

Gheibi:2024:PAC

[2805] Sanaz Gheibi, Tania Banerjee, Sanjay Ranka, and Sartaj Sahni. Path algorithms for contact sequence temporal graphs. Algorithms (Basel), 17 (4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/4/148.

Juarez-Lucero:2024:NAD

[2806] Jorge Juarez-Lucero, Maria Guevara-Villa, Anabel Sanchez-Sanchez, Raquel Diaz-Hernandez, and Leopoldo Altamirano-Robles. A new algorithm for detecting GPN protein expression and overexpression of IDC and ILC Her2+ subtypes on polyacrylamide gels associated with breast cancer. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/
149.

Matrenin:2024:SIF

[2807] Pavel V. Matrenin, Valeriy V. Gamaley, Alexandra I. Khalyasmaa, and Alina I. Stepanova. Solar irradiance forecasting with natural language processing of cloud observations and interpretation of results with modified Shapley additive explanations. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/4/150.

Li:2024:REF

[2808] Ruoyang Li, Shuping Xiong, Yinchao Che, Lei Shi, Xinming Ma, and Lei Xi. Research on efficient feature generation and spatial aggregation for remote sensing semantic segmentation. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/4/151.

Huang:2024:RAC

[2809] Jih-Jeng Huang and Chin-Yi Chen. Resource allocation of cooperative alternatives using the analytic hierarchy process and analytic network process with Shapley values. Algorithms (Basel), 17(4), April 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/4/152.

Robitzsch:2024:SIC

[2810] Alexander Robitzsch. Smooth information criterion for regularized estimation of item response models. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/4/153.

Argyros:2024:HNL

[2811] Ioannis K. Argyros, Santhosh George, Samundra Regmi, and Christopher I. Argyros. Hybrid Newton-like inverse free algorithms for solving nonlinear equations. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/154.

Alahmed:2024:IRD

[2812] Shahad Alahmed, Qutaiba Alasad, Jiann-Shiun Yuan, and Mohammed Alawad. Impacting robustness in deep learning-based NIDS through poisoning attacks. Algorithms (Basel), 17(4), April 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/155.

McMillan:2024:SWS

[2813] Kyle McMillan, Rosa Qiyue So, Camilo Libedinsky, Kai Keng Ang, and Brian Premchand. Spike-weighted spiking neural network with spiking long short-term memory: a biomimetic approach to decoding brain signals. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/
156.

Ghidarcea:2024:PNS

[2814] Mircea Ghidarcea and Decebal Popescu. Prime number sieving — a systematic review with performance analysis. Algorithms (Basel), 17 (4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/4/157.

Guo:2024:PSL

[2815] Jiachen Guo and Wenjie Luo. Point-Sim: a lightweight network for 3D point cloud classification. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/158.

Piazza:2024:EAP

[2816] Carla Piazza, Sabina Rossi, and Daria Smuseva. Efficient algorithm for proportional lumpability and its application to selfish mining in public blockchains. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/4/159.

Al-Betar:2024:CEF

[2817] Mohammed Azmi Al-Betar, Ammar Kamal Abasi, Zaid Abdi Alkareem Alyasseri, Salam Fraihat, and Raghad Falih Mohammed. A communication-efficient federated learning framework for sustainable development using lemurs optimizer. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/4/160.

Safta:2024:APN

[2818] Wiem Safta and Ahmed Shaffie. Advancing pulmonary nodule diagnosis by integrating engineered and deep features extracted from CT scans. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/161.

Garaev:2024:RAA

[2819] Roman Garaev, Bader Rasheed, and Adil Mehmood Khan. Not so robust after all: Evaluating the robustness of deep neural networks to unseen adversarial attacks. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/162.

Chen:2024:QRN

[2820] Yuan Chen and Abdul Khaliq. Quantum recurrent neural networks: Predicting the dynamics of oscillatory and chaotic systems. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/4/163.

Bian:2024:DRL

[2821] Ye Bian, Chengyong Si, and Lei Wang. Diabetic retinopathy lesion segmentation method based on multi-scale attention and lesion perception. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/164.

Yin:2024:RFI

[2822] Chenglong Yin, Fei Zhang, Bin Hao, Zijian Fu, and Xiaoyu Pang. Research on a fast image-matching algorithm based on nonlinear filtering. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/4/165.

Sanchez-Rada:2024:PAM

[2823] J. Fernando Sánchez-Rada, Raquel Vila-Rodríguez, Jesús Montes, and Pedro J. Zufiria. Predicting the aggregate mobility of a vehicle fleet within a city graph. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/4/166.

Torenvliet:2024:EDM

[2824] Nick Torenvliet and John Zelek. Evaluating diffusion models for the automation of ultrasonic nondestructive evaluation data analysis. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/4/167.

Chen:2024:CCC

[2825] Jialu Chen and Baohua Yuan. CCFNet: Collaborative cross-fusion network for medical image segmentation. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/4/168.

Pingos:2024:SOU

[2826] Michalis Pingos, Panayiotis Christodoulou, and Andreas S. Andreou. Security and ownership in user-defined data meshes. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/4/169.

Guijarro:2024:MSM

[2827] María Guijarro, Juan Bayon, Daniel Martín-Carabias, and Joaquín Recas. A multi-stage method for logo detection in scanned official documents based on image processing. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/4/170.

Wessel:2024:PIS

[2828] Niels Wessel, Mariana Sprincean, Ludmila Sidorenko, Ninel Revenco, and Svetlana Hadjiu. Pediatric ischemic stroke: Clinical and paraclinical manifestations — algorithms for diagnosis and treatment. Algorithms (Basel), 17(4), April 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/4/171.

Zhao:2024:IBS

[2829] Yanchi Zhao, Jianhua Cheng, and Jing Cai. Improved brain storm optimization algorithm based on flock decision mutation strategy. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/172.

Jiang:2024:ODA

[2830] Bo Jiang, Guofu Ding, Jianlin Fu, Jian Zhang, and Yong Zhang. An overview of demand analysis and forecasting algorithms for the flow of checked baggage among departing passengers. Algorithms (Basel), 17 (5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/173.

Zhen:2024:OBI

[2831] Qianqian Zhen, Liang Wu, and Guoying Liu. An oracle bone inscriptions detection algorithm based on improved YOLOv8. Algorithms (Basel), 17 (5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/174.

Javed:2024:CPD

[2832] Khadija Javed, Ren Shengbing, Muhammad Asim, and Mudasir Ahmad Wani. Cross-project defect prediction based on domain adaptation and LSTM optimization. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/175.

Xiong:2024:SAT

[2833] Shufeng Xiong, Wenjie Tian, Haiping Si, Guipei Zhang, and Lei Shi. A survey of the applications of text mining for the food domain. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/176.

Chen:2024:MPU

[2834] Xiao Chen, Yu Wu, and Shuting Xu. Mission planning of UAVs and UGV for building inspection in rural area. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/5/177.

Tompra:2024:SML

[2835] Konstantina-Vasiliki Tompra, George Papageorgiou, and Christos Tjortjis. Strategic machine learning optimization for cardiovascular disease prediction and high-risk patient identification. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/178.

Warner:2024:CID

[2836] Charles Warner, Fanyou Wu, Rado Gazo, Bedrich Benes, Nicole Kong, and Songlin Fei. CentralBark image dataset and tree species classification using deep learning. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/179.

Phuntsho:2024:MNP

[2837] Tshewang Phuntsho and Tad Gonsalves. Maximizing net present value for resource constraint project scheduling problems with payments at event occurrences using approximate dynamic programming. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/5/180.

Vasileiadis:2024:AEC

[2838] Vasileios Vasileiadis, Christos Kyriklidis, Vayos Karayannis, and Constantinos Tsanaktsidis. Application of evolutionary computation to the optimization of biodiesel mixtures using a nature-inspired adaptive genetic algorithm. Algorithms (Basel), 17(5), May 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/181.

Islam:2024:SBB

[2839] Kazi Aminul Islam, Hongyi Wu, Chunsheng Xin, Rui Ning, Liuwan Zhu, and Jiang Li. Sub-band backdoor attack in remote sensing imagery. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/182.

Buchmeister:2024:ADA

[2840] Borut Buchmeister and Natasa Vujica Herzog. Advancements in data analysis for the work-sampling method. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/5/183.

Perez-Ramos:2024:ABM

[2841] Jorge L. Perez-Ramos, Selene Ramirez-Rosales, Daniel Canton-Enriquez, Luis A. Diaz-Jimenez, Gabriela Xicotencatl-Ramirez, Ana M. Herrera-Navarro, and Hugo Jimenez-Hernandez. Algorithm based on morphological operators for shortness path planning. Algorithms (Basel), 17 (5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/184.

Sun:2024:LIC

[2842] Haijing Sun, Wen Zhou, Yichuan Shao, Jiaqi Cui, Lei Xing, Qian Zhao, and Le Zhang. A linear interpolation and curvature-controlled gradient optimization strategy based on Adam. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/5/185.

Delabre:2024:SCS

[2843] Mattéo Delabre and Nadia El-Mabrouk. Synesth: Comprehensive syntenic reconciliation with unsampled lineages. Algorithms (Basel), 17 (5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/186.

Markova:2024:MMM

[2844] Valentina Markova, Todor Ganchev, Silvia Filkova, and Miroslav Markov. MMD-MSD: a multimodal multisensory dataset in support of research and technology development for musculoskeletal disorders. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/187.

Hernandez-Resendiz:2024:DDA

[2845] Jaciel David Hernandez-Resendiz, Edgar Tello-Leal, and Marcos Sepúlveda. A data-driven approach to discovering process choreography. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/
188.

Mohammed:2024:IIU

[2846] Sarfaraz Ahmed Mohammed and Anca L. Ralescu. Insights into image understanding: Segmentation methods for object recognition and scene classification. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/5/189.

Zhang:2024:DQN

[2847] Changfan Zhang, Shuo Zhou, Jing He, and Lin Jia. Deep Q-network algorithm-based cyclic air braking strategy for heavy-haul trains. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/190.

Theodorakatos:2024:MMS

[2848] Nikolaos P. Theodorakatos, Rohit Babu, Christos A. Theodoridis, and Angelos P. Moschoudis. Mathematical models for the single-channel and multi-channel PMU allocation problem and their solution algorithms. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/191.

Alves:2024:EEI

[2849] Victor Alves, Florentino Fdez-Riverola, Jorge Ribeiro, José Neves, and Henrique Vicente. Encouraging eco-innovative urban development. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/192.

Forbes:2024:AGE

[2850] Alistair Forbes. The algorithm of Gu and Eisenstat and D-optimal design of experiments. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/193.

Galka:2024:WLS

[2851] Andreas Galka. The weighted least-squares approach to state estimation in linear state space models: The case of correlated noise terms. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/194.

Kannan:2024:EPS

[2852] Shiva Kumar Kannan and Urmila Diwekar. An enhanced particle swarm optimization (PSO) algorithm employing quasi-random numbers. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/195.

Tang:2024:PCF

[2853] Peng Tang, Chuangxin Zhao, Shizhe Liang, and Yuehong Dai. Performance-constraint fault tolerant control to aircraft in presence of actuator deviation. Algorithms (Basel), 17(5), May 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/196.

Ying:2024:ISS

[2854] Wen Ying, Zhaohui Wang, Hui Li, Sheng Du, and Man Zhao. Intelligent ship scheduling and path planning method for maritime emergency rescue. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/
197.

Adamko:2024:TCP

[2855] Peter Adamko. Three cube packing for all dimensions. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/5/198.

Liu:2024:MMV

[2856] Xin Liu, Mingjiang Pang, Qiang Li, Jiehan Zhou, Haiwen Wang, and Dawei Yang. MVACLNet: a multimodal virtual augmentation contrastive learning network for rumor detection. Algorithms (Basel), 17 (5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/199.

Peyman:2024:SLT

[2857] Mohammad Peyman, Xabier A. Martin, Javier Panadero, and Angel A. Juan. A sim-learnheuristic for the team orienteering problem: Applications to unmanned aerial vehicles. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/5/200.

Cholevas:2024:ADB

[2858] Christos Cholevas, Eftychia Angeli, Zacharoula Sereti, Emmanouil Mavrikos, and George E. Tsekouras. Anomaly detection in blockchain networks using unsupervised learning: a survey. Algorithms (Basel), 17 (5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/201.

Shoman:2024:ETS

[2859] Maged Shoman, Tarek Ghoul, Gabriel Lanzaro, Tala Alsharif, Suliman Gargoum, and Tarek Sayed. Enforcing traffic safety: a deep learning approach for detecting motorcyclists’ helmet violations using YOLOv8 and deep convolutional generative adversarial network-generated images. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/202.

Zhang:2024:STB

[2860] Huanlong Zhang, Bin Zhou, Yangyang Tian, and Zhe Li. Segmentation and tracking based on equalized memory matching network and its application in electric substation inspection. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/5/203.

Zarrouk:2024:TDF

[2861] Tarik Zarrouk, Mohammed Nouari, Jamal-Eddine Salhi, Mohammed Abbadi, and Ahmed Abbadi. Three-dimensional finite element modeling of ultrasonic vibration-assisted milling of the nomex honeycomb structure. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/204.

Rad:2024:TWA

[2862] Mahbubeh Askari Rad, Alibek Kruglikov, and Xuhua Xia. Three-way alignment improves multiple sequence alignment of highly diverged sequences. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/5/
205.

Kesireddy:2024:EMC

[2863] Adarsh Kesireddy and F. Antonio Medrano. Elite multi-criteria decision making-Pareto front optimization in multi-objective optimization. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/206.

Ali:2024:AIM

[2864] Esraa H. Ali, Sawsan Sadek, Georges Zakka El Nashef, and Zaid F. Makki. Advanced integration of machine learning techniques for accurate segmentation and detection of Alzheimer’s disease. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/207.

Holasova:2024:CAC

[2865] Eva Holasova, Radek Fujdiak, and Jiri Misurec. Comparative analysis of classification methods and suitable datasets for protocol recognition in operational technologies. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/208.

Muncino:2024:MHA

[2866] David M. Munciño, Emily A. Damian-Ramírez, Mayra Cruz-Fernández, Luis A. Montoya-Santiyanes, and Juvenal Rodríguez-Reséndiz. Metaheuristic and heuristic algorithms-based identification parameters of a direct current motor. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/209.

Liu:2024:PSO

[2867] Yang Liu, Huadong Wang, and Yan Ma. Particle swarm optimizationbased model abstraction and explanation generation for a recurrent neural network. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/5/210.

Liu:2024:SLS

[2868] Chein-Shan Liu, Chung-Lun Kuo, and Chih-Wen Chang. Solving leastsquares problems via a double-optimal algorithm and a variant of the Karush–Kuhn–Tucker equation for over-determined systems. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/5/211.

Koehl:2024:GSP

[2869] Patrice Koehl and Henri Orland. A general statistical physics framework for assignment problems. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/5/212.

Chuu:2024:EHG

[2870] Eric Chuu, Yabo Niu, Anirban Bhattacharya, and Debdeep Pati. EPSOM-Hyb: a general purpose estimator of log-marginal likelihoods with applications in probabilistic graphical models. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/5/213.

Woodrum:2024:IQQ

[2871] Charles Woodrum, Torrey Wagner, and David Weeks. Improving 2-5 qubit quantum phase estimation circuits using machine learning. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/214.

Zhang:2024:MCB

[2872] Lei Zhang, Xiaoke Wang, Michael Rawson, Radu Balan, Edward H. Herskovits, Elias R. Melhem, Linda Chang, Ze Wang, and Thomas Ernst. Motion correction for brain MRI using deep learning and a novel hybrid loss function. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/5/215.

Violante:2024:IMP

[2873] João Pedro Costa Violante, Marcela A. G. Machado, Amanda dos Santos Mendes, and Túlio S. Almeida. An interface to monitor process variability using the binomial ATTRIVAR SS control chart. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/5/216.

Wu:2024:FLM

[2874] Sheng Wu and Jihong Guan. Fault location method based on dynamic operation and maintenance map and common alarm points analysis. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/5/217.

Kim:2024:BSR

[2875] Ryan Kim and Paul M. Torrens. Boundary SPH for robust particlemesh interaction in three dimensions. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/5/218.

Punnen:2024:KPC

[2876] Abraham P. Punnen and Jasdeep Dhahan. The knapsack problem with conflict pair constraints on bipartite graphs and extensions. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/5/219.

Zhang:2024:IFG

[2877] Yiqun Zhang, Honglei Xu, Yang Li, Gang Lin, Liyuan Zhang, Chaoyang Tao, and Yonghong Wu. An integer-fractional gradient algorithm for back propagation neural networks. Algorithms (Basel), 17(5), May 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/5/220.

Dixon:2024:HLA

[2878] Jose Dixon, Oluwatunmise Akinniyi, Abeer Abdelhamid, Gehad A. Saleh, Md Mahmudur Rahman, and Fahmi Khalifa. A hybrid learningarchitecture for improved brain tumor recognition. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/6/221.

Craveirinha:2024:MPP

[2879] José Craveirinha, João Clímaco, Rita Girão-Silva, and Marta Pascoal. Multiobjective path problems and algorithms in telecommunication network design-overview and trends. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/6/222.

Martins:2024:ISB

[2880] Rodrigo Martins, Fátima Rodrigues, Susana Costa, and Nelson Costa. Inertial sensors-based assessment of human breathing pattern: a systematic literature review. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/223.

Hachem:2024:BES

[2881] Hassan Hachem and Candy Abboud. Bayesian estimation of simultaneous regression quantiles using Hamiltonian Monte Carlo. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/224.

Odeh:2024:CPP

[2882] John Owoicho Odeh, Xiaolong Yang, Cosmas Ifeanyi Nwakanma, and Sahraoui Dhelim. Context privacy preservation for user validation by wireless sensors in the industrial metaverse access system. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/225.

Ferdowsi:2024:GPG

[2883] Arman Ferdowsi and Maryam Dehghan Chenary. Gain and pain in graph partitioning: Finding accurate communities in complex networks. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/6/226.

Sharma:2024:EAF

[2884] Neeraj Anand Sharma, Rishal Ravikesh Chand, Zain Buksh, A. B. M. Shawkat Ali, Ambreen Hanif, and Amin Beheshti. Explainable AI frameworks: Navigating the present challenges and unveiling innovative applications. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/6/227.

Morosanu:2024:APL

[2885] Bogdan Morosanu, Marian Negru, and Constantin Paleologu. Automated personalized loudness control for multi-track recordings. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/6/228.

Mohammed:2024:ECN

[2886] Maadh Rajaa Mohammed and Ali Makki Sagheer. Employing a convolutional neural network to classify sleep stages from EEG signals using feature reduction techniques. Algorithms (Basel), 17(6), June 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/6/229.

Ninpan:2024:MCA

[2887] Khwansiri Ninpan, Shuzhang Huang, Francesco Vitillo, Mohamad Ali Assaad, Lies Benmiloud Bechet, and Robert Plana. Mitigating coactivity conflicts and resource overallocation in construction projects: a modular heuristic scheduling approach with Primavera P6 EPPM integration. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/6/230.

Chang:2024:PCC

[2888] Victor Chang, Karl Hall, Qianwen Ariel Xu, Folakemi Ololade Amao, Meghana Ashok Ganatra, and Vladlena Benson. Prediction of customer churn behavior in the telecommunication industry using machine learning models. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/6/
231.

Mehat:2024:MPS

[2889] Yann Méhat, Sylvain Sagot, Egon Ostrosi, and Dominique Deuff. A multi-process system for investigating inclusive design in user interfaces for low-income countries. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/232.

Tian:2024:DDL

[2890] Guangqiang Tian and Fuzhong Wang. Data-driven load frequency control for multi-area power system based on switching method under cyber attacks. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/6/
233.

Zolfagharinia:2024:UPT

[2891] Hossein Zolfagharinia, Mehdi Najafi, Shamir Rizvi, and Aida Haghighi. Unleashing the power of tweets and news in stock-price prediction using machine-learning techniques. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/6/234.

Iniguez-Lomeli:2024:CEU

[2892] Francisco Javier Iñiguez-Lomeli, Edgar Eliseo Franco-Ortiz, Ana Maria Silvia Gonzalez-Acosta, Andres Amador Garcia-Granada, and Horacio Rostro-Gonzalez. A comprehensive exploration of unsupervised classification in spike sorting: a case study on macaque monkey and human pancreatic signals. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/235.

Morakis:2024:HML

[2893] Dimitrios Morakis and Adam Adamopoulos. Hybrid machine learning algorithms to evaluate prostate cancer. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/6/236.

Barba:2024:CAA

[2894] Alessandro Barba, Luca Bertazzi, and Bruce L. Golden. Competitive analysis of algorithms for an online distribution problem. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/237.

Chowdhury:2024:SHE

[2895] Jaher Hassan Chowdhury, Qian Liu, and Sheela Ramanna. Simple histogram equalization technique improves performance of VGG models on facial emotion recognition datasets. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/6/238.

Bai:2024:LSI

[2896] Zhongwei Bai, Peng Yu, Yan Liu, and Jin Guo. Linear system identification-oriented optimal tampering attack strategy and implementation based on information entropy with multiple binary observations. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/6/239.

Samkunta:2024:FEB

[2897] Jirayu Samkunta, Patinya Ketthong, Nghia Thi Mai, Md Abdus Samad Kamal, Iwanori Murakami, and Kou Yamada. Feature extraction based on sparse coding approach for hand grasp type classification. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/240.

Engelbrecht:2024:FLA

[2898] Andries Engelbrecht and Robert Gouldie. Fitness landscape analysis of product unit neural networks. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/6/241.

Azimov:2024:CSM

[2899] Rustam Azimov and Efthimios Providas. A comparative study of machine learning methods and text features for text authorship recognition in the example of Azerbaijani language texts. Algorithms (Basel), 17 (6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/6/242.

Gaytan:2024:TCN

[2900] Armando Gaytan, Ofelia Begovich-Mendoza, and Nancy Arana-Daniel. Training of convolutional neural networks for image classification with fully decoupled extended Kalman filter. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/6/243.

Sharaf:2024:ARA

[2901] Mohamed A. Sharaf, Heba Helal, Nazar Zaki, Wadha Alketbi, Latifa Alkaabi, Sara Alshamsi, and Fatmah Alhefeiti. Automated recommendation of aggregate visualizations for crowdfunding data. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/244.

Pascoe:2024:MAH

[2902] Sean Pascoe, Anna Farmery, Rachel Nichols, Sarah Lothian, and Kamal Azmi. A modified analytic hierarchy process suitable for online survey preference elicitation. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/245.

Evans:2024:MQF

[2903] William Evans and David Kirkpatrick. Minimizing query frequency to bound congestion potential for moving entities at a fixed target time. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/6/246.

Ferdous:2024:NMV

[2904] Sayeda Muntaha Ferdous, Shafayat Bin Shabbir Mugdha, and Iman Dehzangi. New multi-view feature learning method for accurate antifungal peptide detection. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/247.

Volkau:2024:NGN

[2905] Ihar Volkau, Sergei Krasovskii, Abdul Mujeeb, and Helen Balinsky. A non-gradient and non-iterative method for mapping 3D mesh objects based on a summation of dependent random values. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/6/248.

Gussenbauer:2024:SCC

[2906] Johannes Gussenbauer, Matthias Templ, Siro Fritzmann, and Alexander Kowarik. Simulation of calibrated complex synthetic population data with XGBoost. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/249.

Chen:2024:RDF

[2907] Chengming Chen, Xuejun Ren, and Guoqing Cheng. Research on distributed fault diagnosis model of elevator based on PCA-LSTM. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/6/250.

Pan:2024:DCH

[2908] Tingzhe Pan, Jue Hou, Xin Jin, Zhenfan Yu, Wei Zhou, and Zhijun Wang. Distributed control of hydrogen-based microgrids for the demand side: a multiagent self-triggered MPC-based strategy. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/251.

Lebedeva:2024:PHS

[2909] Albina V. Lebedeva, Margarita I. Samburova, Vyacheslav V. Razin, Nikolay V. Gromov, Svetlana A. Gerasimova, Tatiana A. Levanova, Lev A. Smirnov, and Alexander N. Pisarchik. Prediction of hippocampal signals in mice using a deep learning approach for neurohybrid technology applications. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/6/252.

Cabodi:2024:HMC

[2910] Gianpiero Cabodi, Paolo Enrico Camurati, Marco Palena, and Paolo Pasini. Hardware model checking algorithms and techniques. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/6/253.

Faa:2024:QAA

[2911] Gavino Faa, Massimo Castagnola, Luca Didaci, Fernando Coghe, Mario Scartozzi, Luca Saba, and Matteo Fraschini. The quest for the application of artificial intelligence to whole slide imaging: Unique prospective from new advanced tools. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/254.

Nunez-Valdez:2024:GES

[2912] Edward Rolando Núñez-Valdez and Vicente García-Díaz. Guest editorial for the special issue “New Trends in Algorithms for Intelligent Recommendation Systems”. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/255.

Mittal:2024:SCA

[2913] Vikas Mittal, Kanta Prasad Sharma, Narmadha Thangarasu, Udandarao Sarat, Ahmad O. Hourani, and Rohit Salgotra. Synthesis of circular antenna arrays for achieving lower side lobe level and higher directivity using hybrid optimization algorithm. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/6/256.

Sun:2024:NNN

[2914] Yue Sun, Zhaohong Huang, Guorong Cai, Jinhe Su, and Zheng Gong. NSBR-Net: a novel noise suppression and boundary refinement network for breast tumor segmentation in ultrasound images. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/6/257.

Inza:2024:AMD

[2915] Ernesto Parra Inza, Nodari Vakhania, José María Sigarreta Almira, and José Alberto Hernández-Aguilar. Approximating a minimum dominating set by purification. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/258.

Elhairach:2024:UOD

[2916] Mohamed Limam El hairach, Amal Tmiri, and Insaf Bellamine. Univariate outlier detection: Precision-driven algorithm for single-cluster scenarios. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/6/259.

Montesinos-Lopez:2024:EDA

[2917] Osval A. Montesinos-López, Arvinth Sivakumar, Gloria Isabel Huerta Prado, Josafhat Salinas-Ruiz, Afolabi Agbona, Axel Efraín Ortiz Reyes, Khalid Alnowibet, Rodomiro Ortiz, Abelardo Montesinos-López, and José Crossa. Exploring data augmentation algorithm to improve genomic prediction of top-ranking cultivars. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/6/260.

Rahman:2024:AIB

[2918] Md. Shahidur Rahman, Sowrav Chowdhury, Mirza Rasheduzzaman, and A. B. M. S. U. Doulah. Artificial intelligence-based algorithms and healthcare applications of respiratory inductance plethysmography: a systematic review. Algorithms (Basel), 17(6), June 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/261.

Chen:2024:ELE

[2919] Shijuan Chen, Zongmei Li, Xiaofeng Du, and Qin Nie. EAND-LPRM: Enhanced attention network and decoding for efficient license plate recognition under complex conditions. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/6/262.

Li:2024:RBI

[2920] Saiya Li, Jinhe Su, Guoqing Jiang, Ziyu Huang, and Xiaorong Zhang. 3D reconstruction based on iterative optimization of moving least-squares function. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/6/263.

Cheng:2024:OCP

[2921] Rei-Heng Cheng, Chang-Wu Yu, and Zuo-Li Zhang. Optimizing charging pad deployment by applying a quad-tree scheme. Algorithms (Basel), 17 (6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/6/264.

Fachada:2024:AIM

[2922] Nuno Fachada and Nuno David. Artificial intelligence in modeling and simulation. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/6/265.

Liu:2024:ROA

[2923] Chein-Shan Liu, Chih-Wen Chang, and Chung-Lun Kuo. Reorthogonalized/affine GMRES and orthogonalized maximal projection algorithm for solving linear systems. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/6/266.

Ghanbari:2024:SSS

[2924] Alireza Ghanbari, Gholam Hassan Shirdel, and Farhad Maleki. Semi-self-supervised domain adaptation: Developing deep learning models with limited annotated data for wheat head segmentation. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/6/267.

Berglund:2024:PUT

[2925] Martin Berglund, Henrik Björklund, and Johanna Björklund. Parsing unranked tree languages, folded once. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/6/268.

Aydin:2024:WRL

[2926] Mehmet Emin Aydin, Rafet Durgut, and Abdur Rakib. Why reinforcement learning? Algorithms (Basel), 17(6), June 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/6/269.

Begovich:2024:SFA

[2927] Ofelia Begovich, Adrián Lizárraga, and Antonio Ramírez-Treviño. Sensor fusion architecture for fault diagnosis with a predefined-time observer. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/6/270.

Santos:2024:TSI

[2928] João A. M. Santos, Miguel S. E. Martins, Rui M. Pinto, and Susana M. Vieira. Towards sustainable inventory management: a many-objective approach to stock optimization in multi-storage supply chains. Algorithms (Basel), 17(6), June 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/6/271.

Sun:2024:IAA

[2929] Haijing Sun, Hao Yu, Yichuan Shao, Jiantao Wang, Lei Xing, Le Zhang, and Qian Zhao. An improved Adam’s algorithm for stomach image classification. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/
272.

Sadolewski:2024:VCS

[2930] Jan Sadolewski and Bartosz Trybus. Verification of control system runtime using an executable semantic model. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/7/273.

Zanlorenzi:2024:MCD

[2931] Guilherme Zanlorenzi, Anderson Luis Szejka, and Osiris Canciglieri Junior. Multi-criteria decision support system for automatically selecting photovoltaic sets to maximise micro solar generation. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/274.

Pais:2024:SDC

[2932] Namitha Pais, Nalini Ravishanker, and Sanguthevar Rajasekaran. Supervised dynamic correlated topic model for classifying categorical time series. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/275.

Wang:2024:AYN

[2933] Shijie Wang, Zekun Zhang, Qingqing Chao, and Teng Yu. AFE-YOLOv8: a novel object detection model for unmanned aerial vehicle scenes with adaptive feature enhancement. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/276.

Kim:2024:AIV

[2934] Wangdo Kim, Duarte Araujo, MooYoung Choi, Albert Vette, and Eunice Ortiz. Algorithmic implementation of visually guided interceptive actions: Harmonic ratios and stimulation invariants. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/277.

Mo:2024:DSF

[2935] Tianjin Mo, Qiao Xiao, Hongyi Zhang, Ren Li, and Yunsong Wu. Domain-specific few-shot table prompt question answering via contrastive exemplar selection. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/278.

Yonekura:2024:SNP

[2936] Kazuo Yonekura. A short note on physics-guided GAN to learn physical models without gradients. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/279.

Aldamani:2024:LXR

[2937] Raghad Aldamani, Diaa Addeen Abuhani, and Tamer Shanableh. LungVision: X-ray imagery classification for on-edge diagnosis applications. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/280.

Carbone:2024:OAB

[2938] Patrick Carbone, Celina Alba, Alexis Bennett, Kseniia Kriukova, and Dominique Duncan. Optimizing automated brain extraction for moderate to severe traumatic brain injury patients: The role of intensity normalization and bias-field correction. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/7/281.

Ramos-Valdovinos:2024:QPA

[2939] Miguel Angel Ramos-Valdovinos, Prisciluis Caheri Salas-Navarrete, Gerardo R. Amores, Ana Lilia Hernández-Orihuela, and Agustino Martínez-Antonio. Qualitative perturbation analysis and machine learning: Elucidating bacterial optimization of tryptophan production. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/7/282.

Kats:2024:MAE

[2940] Vladimir Kats and Eugene Levner. Maximizing the average environmental benefit of a fleet of drones under a periodic schedule of tasks. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/7/283.

Jaradat:2024:ELP

[2941] Shadi Jaradat, Richi Nayak, Alexander Paz, and Mohammed Elhenawy. Ensemble learning with pre-trained transformers for crash severity classification: a deep NLP approach. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/284.

Khomidov:2024:NEA

[2942] Mavlonbek Khomidov and Jong-Ha Lee. The novel EfficientNet architecture-based system and algorithm to predict complex human emotions. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/285.

Habeb:2024:EVA

[2943] Mohamed H. Habeb, May Salama, and Lamiaa A. Elrefaei. Enhancing video anomaly detection using a transformer spatiotemporal attention unsupervised framework for large datasets. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/7/286.

Tao:2024:EPS

[2944] Ning Tao, Anthony Ventresque, Vivek Nallur, and Takfarinas Saber. Enhancing program synthesis with large language models using manyobjective grammar-guided genetic programming. Algorithms (Basel), 17 (7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/17/7/287.

Shiryayeva:2024:ODP

[2945] Olga Shiryayeva, Batyrbek Suleimenov, and Yelena Kulakova. Optimal design of I-PD and PI-D industrial controllers based on artificial intelligence algorithm. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/7/288.

Urumov:2024:FFB

[2946] Georgy Urumov, Panagiotis Chountas, and Thierry Chaussalet. Fuzzy fractional Brownian motion: Review and extension. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/289.

Alkhamisi:2024:FLB

[2947] Abrar Alkhamisi, Iyad Katib, and Seyed M. Buhari. Federated learningbased security attack detection for multi-controller software-defined networks. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/290.

Ghidarcea:2024:PTT

[2948] Mircea Ghidarcea and Decebal Popescu. Prime time tactics sieve tweaks and boosters. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/
291.

Ahmad:2024:CKT

[2949] Hawraz A. Ahmad and Tarik A. Rashid. Central Kurdish text-to-speech synthesis with novel end-to-end transformer training. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/292.

Chen:2024:HPM

[2950] Jianzhang Chen, Shuo Zhou, Jie Qiu, Yixin Xu, Bozhe Zeng, Wanchuan Fang, Xiangying Chen, Yipeng Huang, Zhengquan Xu, and Youqin Chen. A histogram publishing method under differential privacy that involves balancing small-bin availability first. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/7/293.

Mosqueda-Arvizu:2024:LET

[2951] Carlos-Antonio Mosqueda-Arvizu, Julio-Alejandro Romero-González, Diana-Margarita Córdova-Esparza, Juan Terven, Ricardo Chaparro-Sánchez, and Juvenal Rodríguez-Reséndiz. Logical execution time and time-division multiple access in multicore embedded systems: a case study. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/294.

Rjeib:2024:VEV

[2952] Hasanein D. Rjeib and Gabor Kecskemeti. VMP-ER: an efficient virtual machine placement algorithm for energy and resources optimization in cloud data center. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/7/295.

Bruno:2024:HAD

[2953] Roberto Bruno. Hardness and approximability of dimension reduction on the probability simplex. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/296.

Vais:2024:CSI

[2954] Ram Ishwar Vais, Kuldeep Sahay, Tirumalasetty Chiranjeevi, Ramesh Devarapalli, and ukasz Knypi ski. Crystal symmetry-inspired algorithm for optimal design of contemporary mono passivated emitter and rear cell solar photovoltaic modules. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/297.

Wang:2024:SIM

[2955] Shuling Wang, Fengze Jiang, and Xiaojin Gong. A sparsity-invariant model via unifying depth prediction and completion. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/298.

Mihova:2024:MGC

[2956] Evangelina I. Mihova and Yuri N. Sotskov. Mixed graph colouring as scheduling a partially ordered set of interruptible multi-processor tasks with integer due dates. Algorithms (Basel), 17(7), 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/299.

Chen:2024:CRT

[2957] Ivane Delos Santos Chen, Chieh-Ming Yang, Shang-Shu Wu, Chih-Kang Yang, Mei-Juan Chen, Chia-Hung Yeh, and Yuan-Hong Lin. Continuous recognition of teachers’ hand signals for students with attention deficits. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/7/300.

Lyons:2024:CC

[2958] Steven Lyons and Raju Rangaswami. To cache or not to cache. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/7/301.

Prabhakar:2024:SSC

[2959] Sunil Kumar Prabhakar and Dong-Ok Won. SCMs: Systematic conglomerated models for audio cough signal classification. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/302.

Hafshejani:2024:ITS

[2960] Sajad Fathi Hafshejani, Daya Gaur, and Robert Benkoczi. On implementing a two-step interior point method for solving linear programs. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/7/303.

Li:2024:SCP

[2961] Yong Li, Qidan Zhu, and Ahsan Elahi. Sequential convex programming for nonlinear optimal control in UAV trajectory planning. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/7/304.

Arezoumand:2024:ETA

[2962] Sara Arezoumand and Omar Smadi. Equity in transportation asset management: a proposed framework. Algorithms (Basel), 17(7), 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/7/305.

Pazmino:2024:PEF

[2963] Raúl Pazmiño, Wilson Pavon, Matthew Armstrong, and Silvio Simani. Performance evaluation of fractional proportional–integral–derivative controllers tuned by heuristic algorithms for nonlinear interconnected tanks. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/306.

Schaa:2024:EER

[2964] Hans Schaa and Nicolas A. Barriga. Evaluating the expressive range of Super Mario Bros level generators. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/7/307.

Chen:2024:AVP

[2965] Yan Chen, Gang Liu, Longda Wang, and Bing Xia. Automatic vertical parking reference trajectory based on improved immune shark smell optimization. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/
308.

Alcaraz-Chavez:2024:RTT

[2966] Jesus Eduardo Alcaraz-Chavez, Adriana del Carmen Téllez-Anguiano, Juan Carlos Olivares-Rojas, and Ricardo Martínez-Parrales. Real-time tracking and detection of cervical cancer precursor cells: Leveraging SIFT descriptors in mobile video sequences for enhanced early diagnosis. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/7/309.

Adibi:2024:MBG

[2967] Aria Adibi and Hovhannes A. Harutyunyan. Messy broadcasting in grid. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/7/310.

Bouyuklieva:2024:GAG

[2968] Stefka Bouyuklieva, Iliya Bouyukliev, Valentin Bakoev, and Maria Pashinska-Gadzheva. Generating m-ary Gray codes and related algorithms. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
7/311.

Jia:2024:NWS

[2969] Zijie Jia, Zhijian Fang, and Huaxiong Zhang. Normalization of Web of Science institution names based on deep learning. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/312.

Dey:2024:NHC

[2970] Bishwajit Dey, Gulshan Sharma, and Pitshou N. Bokoro. A novel hybrid crow search arithmetic optimization algorithm for solving weighted combined economic emission dispatch with load-shifting practice. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/7/313.

Yoshioka:2024:RQM

[2971] Hitoshi Yoshioka and Hirotada Hashimoto. A reliability quantification method for deep reinforcement learning-based control. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/7/314.

Orlova:2024:AIB

[2972] Ekaterina V. Orlova. Artificial intelligence-based system for retinal disease diagnosis. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/7/315.

Abundez:2024:TAL

[2973] Itzel M. Abundez, Roberto Alejo, Francisco Primero Primero, Everardo E. Granda-Gutiérrez, Otniel Portillo-Rodríguez, and Juan Alberto Antonio Velázquez. Threshold active learning approach for physical violence detection on images obtained from video (frame-level) using pretrained deep learning neural network models. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/7/316.

Ponce-Silva:2024:AAS

[2974] Mario Ponce-Silva, Óscar Sánchez-Vargas, Claudia Cortés-García, Jesús Aguayo-Alquicira, and Susana Estefany De León-Aldaco. Algorithm for assessment of the switching angles in the unipolar SPWM technique for single-phase inverters. Algorithms (Basel), 17(7), 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/7/317.

Bodory:2024:EDM

[2975] Hugo Bodory, Federica Mascolo, and Michael Lechner. Enabling decision making with the modified causal forest: Policy trees for treatment assignment. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/7/
318.

Tobon:2024:ASM

[2976] Andrés Tobón, Carlos Andrés Ramos-Paja, Martha Lucía Orozco-Gutíerrez, Andrés Julián Saavedra-Montes, and Sergio Ignacio Serna-Garcés. Adaptive sliding-mode controller for a zeta converter to provide high-frequency transients in battery applications. Algorithms (Basel), 17(7), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/7/319.

Femminella:2024:CRL

[2977] Mauro Femminella and Gianluca Reali. Comparison of reinforcement learning algorithms for edge computing applications deployed by serverless technologies. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/320.

Liu:2024:MHS

[2978] Zhaofeng Liu, Xiaoqing Zheng, Anke Xue, Ming Ge, and Aipeng Jiang. Multi-head self-attention-based fully convolutional network for RUL prediction of turbofan engines. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/321.

Solis-Villarreal:2024:ECO

[2979] José-Alberto Solís-Villarreal, Valeria Soto-Mendoza, Jesús Alejandro Navarro-Acosta, and Efraín Ruiz y Ruiz. Energy consumption outlier detection with AI models in modern cities: a case study from North-Eastern Mexico. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/322.

Thorjussen:2024:CTC

[2980] Christian B. H. Thorjussen, Kristian Hovde Liland, Ingrid Måge, and Lars Erik Solberg. Computational test for conditional independence. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/323.

Wang:2024:TCR

[2981] Jianping Wang, Youchao Wang, Boyan Chen, Xiaoyue Jia, and Dexi Pu. Trajectory classification and recognition of planar mechanisms based on ResNet18 network. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/324.

Udhayasekar:2024:LSA

[2982] Sethu Vinayagam Udhayasekar, Karthik K. Srinivasan, Pramesh Kumar, and Bhargava Rama Chilukuri. Label-setting algorithm for multidestination k simple shortest paths problem and application. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/325.

Scavino:2024:EIP

[2983] Edgar Scavino, Mohd Amiruddin Abd Rahman, Zahid Farid, Sadique Ahmad, and Muhammad Asim. Enhancing indoor positioning accuracy with WLAN and WSN: a QPSO hybrid algorithm with surface tessellation. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/8/326.

Daribayev:2024:QAE

[2984] Beimbet Daribayev, Aksultan Mukhanbet, Nurtugan Azatbekuly, and Timur Imankulov. A quantum approach for exploring the numerical results of the heat equation. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/327.

Rabie:2024:HFS

[2985] Tamer Rabie, Mohammed Baziyad, and Ibrahim Kamel. High-fidelity steganography: a covert parity bit model-based approach. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/328.

Narigina:2024:CNN

[2986] Marta Narigina, Andrejs Romanovs, and Yuri Merkuryev. Convolutional neural network-based digital diagnostic tool for the identification of psychosomatic illnesses. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/329.

Tous:2024:LRA

[2987] Ruben Tous. Lester: Rotoscope animation through video object segmentation and tracking. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/330.

Alhujaylan:2024:SIS

[2988] Abeer I. Alhujaylan and Manar I. Hosny. A swarm intelligence solution for the multi-vehicle profitable pickup and delivery problem. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/331.

Usmani:2024:PAT

[2989] Mehak Usmani, Zulfiqar Ali Memon, Adil Zulfiqar, and Rizwan Qureshi. Preptimize: Automation of time series data preprocessing and forecasting. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/8/332.

Awasthi:2024:DLB

[2990] Amruta Awasthi, Lenka Krpalkova, and Joseph Walsh. Deep learningbased Boolean, time series, error detection, and predictive analysis in container crane operations. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/333.

Zhang:2024:EOM

[2991] Lijun Zhang, Han Zou, and Yungang Zhu. An efficient optimization of the Monte Carlo tree search algorithm for Amazons. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/8/334.

Kwiek:2024:CSC

[2992] Patrycja Kwiek and Ma gorzata Jakubowska. Color standardization of chemical solution images using template-based histogram matching in deep learning regression. Algorithms (Basel), 17(8), 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/335.

Briones-Baez:2024:PMS

[2993] Mónica Fabiola Briones-Báez, Luciano Aguilera-Vázquez, Nelson Rangel-Valdez, Cristal Zuñiga, Ana Lidia Martínez-Salazar, and Claudia Gomez-Santillan. Pitfalls in metaheuristics solving stoichiometric-based optimization models for metabolic networks. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/8/336.

Bakker:2024:HPH

[2994] Wim Bakker, Frank van Ruitenbeek, Harald van der Werff, Christoph Hecker, Arjan Dijkstra, and Freek van der Meer. Hyperspectral Python: HypPy. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/8/337.

Madanayake:2024:DMS

[2995] Arjuna Madanayake, Keththura Lawrance, Bopage Umesha Kumarasiri, Sivakumar Sivasankar, Thushara Gunaratne, Chamira U. S. Edussooriya, and Renato J. Cintra. Design of multichannel spectrum intelligence systems using approximate Discrete Fourier Transform algorithm for antenna array-based spectrum perception applications. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/8/338.

AlSerhali:2024:CSP

[2996] Antonio Al Serhali and Joachim Niehren. Complete subhedge projection for stepwise hedge automata. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/339.

Lin:2024:HOF

[2997] Yu Lin, Yanhua Qiu, Hao Chen, Jun Zhou, Jiayi He, Penghua Du, and Dafan Liu. Hierarchical optimization framework for layout design of star–tree gas-gathering pipeline network in discrete spaces. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/340.

Li:2024:FME

[2998] Qiang Li, Xiao Liao, Wei Cui, Ying Wang, Hui Cao, and Qingshu Guan. Fast minimum error entropy for linear regression. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/8/341.

Gafar:2024:OGS

[2999] Mona G. Gafar, Amr A. Abohany, Ahmed E. Elkhouli, and Amr A. Abd El-Mageed. Optimization of gene selection for cancer classification in high-dimensional data using an improved African vultures algorithm. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/342.

Modrak:2024:RRL

[3000] Vladimir Modrak, Ranjitharamasamy Sudhakarapandian, Arunmozhi Balamurugan, and Zuzana Soltysova. A review on reinforcement learning in production scheduling: an inferential perspective. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/8/343.

Zhuang:2024:IBD

[3001] Kai Zhuang, Daniel Trad, and Amr Ibrahim. Inversion-based deblending in common midpoint domain using time domain high-resolution radon. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/344.

Lizcano:2024:PTC

[3002] Leonardo Lizcano, Eduardo Angulo, and José Márquez. Precedence table construction algorithm for CFGs regardless of being OPGs. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/345.

Al-Betar:2024:ECS

[3003] Mohammed Azmi Al-Betar, Zaid Abdi Alkareem Alyasseri, Noor Kamal Al-Qazzaz, Sharif Naser Makhadmeh, Nabeel Salih Ali, and Christoph Guger. EEG channel selection for stroke patient rehabilitation using BAT optimizer. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/346.

Shen:2024:CRP

[3004] Yufei Shen and Wenxing Zhou. Classification and regression of pinhole corrosions on pipelines based on magnetic flux leakage signals using convolutional neural networks. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/347.

Tarazona-Torres:2024:PMS

[3005] Luis Tarazona-Torres, Ciro Amaya, Alvaro Paipilla, Camilo Gomez, and David Alvarez-Martinez. The parallel machine scheduling problem with different speeds and release times in the ore hauling operation. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/8/348.

Grad:2024:ESI

[3006] Sorin-Mihai Grad. Editorial for the special issue on recent advances in nonsmooth optimization and analysis. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/8/349.

Piazza:2024:NSN

[3007] Mattia Piazza, Enrico Bertolazzi, and Marco Frego. A non-smooth numerical optimization approach to the three-point Dubins problem (3PDP). Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
8/350.

Lazari:2024:MOR

[3008] Vasiliki Lazari, Athanasios Chassiakos, and Stylianos Karatzas. Multiobjective resource-constrained scheduling in large and repetitive construction projects. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/351.

Boujou:2024:DAG

[3009] Moncef Boujou, Rabah Iguernaissi, Lionel Nicod, Djamal Merad, and Séverine Dubuisson. In-depth analysis of GAF-Net: Comparative fusion approaches in video-based person re-identification. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/8/352.

Gamil:2024:EAA

[3010] Seham Gamil, Feng Zeng, Moath Alrifaey, Muhammad Asim, and Naveed Ahmad. An efficient AdaBoost algorithm for enhancing skin cancer detection and classification. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/353.

Matkerim:2024:MLA

[3011] Bazargul Matkerim, Aksultan Mukhanbet, Nurislam Kassymbek, Beimbet Daribayev, Maksat Mustafin, and Timur Imankulov. Machine learning analysis using the black oil model and parallel algorithms in oil recovery forecasting. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/354.

Al-Gburi:2024:MOU

[3012] Abbas Fadhil Jasim Al-Gburi, Mohd Zakree Ahmad Nazri, Mohd Ridzwan Bin Yaakub, and Zaid Abdi Alkareem Alyasseri. Multi-objective unsupervised feature selection and cluster based on symbiotic organism search. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/8/355.

Bosse:2024:VMP

[3013] Stefan Bosse. A virtual machine platform providing machine learning as a programmable and distributed service for IoT and edge on-device computing: Architecture, transformation, and evaluation of integer discretization. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/8/
356.

Hajipour:2024:SDP

[3014] Vahid Hajipour, Sajjad Jalali, Francisco Javier Santos-Arteaga, Samira Vazifeh Noshafagh, and Debora Di Caprio. A system design perspective for business growth in a crowdsourced data labeling practice. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/357.

Zhang:2024:DCM

[3015] Yingjuan Zhang and Jochen Einbeck. Directed clustering of multivariate data based on linear or quadratic latent variable models. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/358.

Lee:2024:ECT

[3016] Chuan-Min Lee. Exploring clique transversal variants on distancehereditary graphs: Computational insights and algorithmic approaches. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/359.

Noitz:2024:DSE

[3017] Matthias Noitz, Christoph Mörtl, Carl Böck, Christoph Mahringer, Ulrich Bodenhofer, Martin W. Dünser, and Jens Meier. Detection of subtle ECG changes despite superimposed artifacts by different machine learning algorithms. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/360.

Aizenberg:2024:FDS

[3018] Igor Aizenberg and Alexander Vasko. Frequency-domain and spatialdomain MLMVN-based convolutional neural networks. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/361.

Jaiteh:2024:UML

[3019] Musa Jaiteh, Edith Phalane, Yegnanew A. Shiferaw, Karen Alida Voet, and Refilwe Nancy Phaswana-Mafuya. Utilization of machine learning algorithms for the strengthening of HIV testing: a systematic review. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/362.

Greco:2024:CVA

[3020] Danilo Greco, Majid Fasihiany, Ali Varasteh Ranjbar, Francesco Masulli, Stefano Rovetta, and Alberto Cabri. Computer vision algorithms on a Raspberry Pi 4 for automated depalletizing. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/8/363.

Moqurrab:2024:HHR

[3021] Syed Atif Moqurrab, Hari Mohan Rai, and Joon Yoo. HRIDM: Hybrid residual/inception-based deeper model for arrhythmia detection from large sets of 12-lead ECG recordings. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/8/364.

Konstantinov:2024:IPB

[3022] Yuri Konstantinov, Anton Krivosheev, and Fedor Barkov. An image processing-based correlation method for improving the characteristics of Brillouin frequency shift extraction in distributed fiber optic sensors. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/365.

Khanov:2024:DTO

[3023] Alexander Khanov, Anastasija Shulzhenko, Anzhelika Voroshilova, Alexander Zubarev, Timur Karimov, and Shakeeb Fahmi. Determining thresholds for optimal adaptive discrete cosine transformation. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/366.

Ding:2024:ADV

[3024] Kelin Ding, Jiayi Luo, Haohang Huang, John M. Hart, Issam I. A. Qamhia, and Erol Tutumluer. Augmented dataset for vision-based analysis of railroad ballast via multi-dimensional data synthesis. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/367.

Abubakar:2024:ICD

[3025] Abdulrazaq Nafiu Abubakar, Mustapha Kamel Khaldi, Mujahed Aldhaifallah, Rohit Patwardhan, and Hussain Salloum. Identification of crude distillation unit: a comparison between neural network and Koopman operator. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/8/368.

Alsinet:2024:CBP

[3026] Teresa Alsinet, Josep Argelich, Ramón Béjar, and Santi Martínez. On the complexity of the bipartite polarization problem: From neutral to highly polarized discussions. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/369.

He:2024:JOS

[3027] Zhenli He, Liheng Li, Ziqi Lin, Yunyun Dong, Jianglong Qin, and Keqin Li. Joint optimization of service migration and resource allocation in mobile edge–cloud computing. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/8/370.

Abdelmaguid:2024:BOD

[3028] Tamer F. Abdelmaguid. Bi-objective, dynamic, multiprocessor openshop scheduling: a hybrid scatter search–tabu search approach. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/8/371.

Mohammadzadeh:2024:CAR

[3029] Saeed Mohammadzadeh, Hamidreza Heydari, Mahdi Karimi, and Araliya Mosleh. Correlation analysis of railway track alignment and ballast stiffness: Comparing frequency-based and machine learning algorithms. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/372.

Kowalenko:2024:AVT

[3030] Victor Kowalenko. Algorithms for various trigonometric power sums. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/8/373.

Chiang:2024:EGM

[3031] Ching-Shoei Chiang. Extended general Malfatti’s problem. Algorithms (Basel), 17(8), 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/8/374.

Iliopoulou:2024:SFD

[3032] Sofia Iliopoulou, Panagiotis Tsinganos, Dimitris Ampeliotis, and Athanassios Skodras. Synthetic face discrimination via learned image compression. Algorithms (Basel), 17(9), September 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/375.

Sandulescu:2024:IIA

[3033] Virginia Sandulescu, Marilena Ianculescu, Liudmila Valeanu, and Adriana Alexandru. Integrating IoMT and AI for proactive healthcare: Predictive models and emotion detection in neurodegenerative diseases. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
9/376.

Gaur:2024:SBB

[3034] Daya Gaur, Shahadat Hossain, and Rishi Ranjan Singh. Star bicolouring of bipartite graphs. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/377.

Capel:2024:PPE

[3035] Manuel I. Capel, Alberto Salguero-Hidalgo, and Juan A. Holgado-Terriza. Parallel PSO for efficient neural network training using GPGPU and Apache Spark in edge computing sets. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/378.

Shingade:2024:HPS

[3036] Sandip Shingade, Rajdeep Niyogi, and Mayuri Pichare. Hybrid particle swarm optimization-Jaya algorithm for team formation. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/379.

Caramia:2024:INB

[3037] Massimiliano Caramia and Giuseppe Stecca. An improved negotiationbased approach for collecting and sorting operations in waste management and recycling. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/380.

Al-sudani:2024:MBA

[3038] Ahlam Hanoon Al-sudani, Basheera M. Mahmmod, Firas A. Sabir, Sadiq H. Abdulhussain, Muntadher Alsabah, and Wameedh Nazar Flayyih. Multithreading-based algorithm for high-performance Tchebichef polynomials with higher orders. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/9/381.

Naranjo:2024:RQC

[3039] Miguel A. Naranjo and Luis A. Fletscher. Review quantum circuit synthesis for Grover’s algorithm oracle. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/17/9/382.

Dayyani:2024:CBS

[3040] Faranak Dayyani, Charlene H. Chu, Ali Abedi, and Shehroz S. Khan. Correlations between social isolation and functional decline in older adults after lower limb fractures using multimodal sensors: a pilot study. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/
383.

Chen:2024:MSD

[3041] Juan Chen and Rui Huang. Multi-source data-driven local-global dynamic multi-graph convolutional network for bike-sharing demands prediction. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/9/384.

Ouali:2024:ERA

[3042] Soufiyan Ouali and Said El Garouani. Efficient and robust Arabic automotive speech command recognition system. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/385.

Kabashkin:2024:ACD

[3043] Igor Kabashkin. Algorithms of cross-domain redundancy management for resilient of dual-priority critical communication systems. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/386.

Kandilarov:2024:NAI

[3044] Juri D. Kandilarov and Lubin G. Vulkov. Numerical algorithms for identification of convection coefficient and source in a magnetohydrodynamics flow. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/9/387.

Cai:2024:DRB

[3045] Xueyi Cai and Youhua Qian. Detection and reconstruction of bursting oscillations in complex systems using the HAVOK analysis framework. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/
388.

Zhang:2024:IMC

[3046] Chunfu Zhang, Jing Bai, Yanchun Liang, Adriano Tavares, Lidong Wang, Tiago Gomes, and Sandro Pinto. An improved multi-chaotic public key algorithm based on Chebyshev polynomials. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/389.

Papadimitriou:2024:RBF

[3047] Vasso E. Papadimitriou, Georgios N. Aretoulis, and Jason Papathanasiou. Radial basis function (RBF) and multilayer perceptron (MLP) comparative analysis on building renovation cost estimation: The case of Greece. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/9/390.

Reyes:2024:FMA

[3048] Christian Reyes, Julio C. Ramos-Fernández, Eduardo S. Espinoza, and Rogelio Lozano. Fuzzy modelling algorithms and parallel distributed compensation for coupled electromechanical systems. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/391.

Reshadi:2024:CRA

[3049] MohammadHossein Reshadi, Wen Li, Wenjie Xu, Precious Omashor, Albert Dinh, Jun Xiao, Scott Dick, Yuntong She, and Michael Lipsett. Correction: Reshadi et al. deep-shallow metaclassifier with synthetic minority oversampling for anomaly detection in a time series. algorithms 2024, 17, 114. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/392. See [2771].

Viveros:2024:MDS

[3050] Andres Viveros, Pablo Adasme, Ali Dehghan Firoozabadi, and Enrique San Juan. Minimizing the density of switch-controller latencies over total latency for software-defined networks. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/393.

Charilogis:2024:CPS

[3051] Vasileios Charilogis, Ioannis G. Tsoulos, and Anna Maria Gianni. Combining parallel stochastic methods and mixed termination rules in optimization. Algorithms (Basel), 17(9), September 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/394.

Golyandina:2024:ICS

[3052] Nina Golyandina and Nikita Zvonarev. Information criteria for signal extraction using singular spectrum analysis: White and red noise. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/395.

Torkzaban:2024:APA

[3053] Nariman Torkzaban, Anousheh Gholami, John S. Baras, and Bruce L. Golden. AASA: a priori adaptive splitting algorithm for the split delivery vehicle routing problem. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/9/396.

Alhamd:2024:AFD

[3054] Qusay Alhamd, Mohsen Saniei, Seyyed Ghodratollah Seifossadat, and Elaheh Mashhour. Advanced fault detection in power transformers using improved wavelet analysis and LSTM networks considering current transformer saturation and uncertainties. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/17/9/397.

He:2024:DIA

[3055] Yuxin He, Ping Huang, Weihang Hong, Qin Luo, Lishuai Li, and Kwok-Leung Tsui. In-depth insights into the application of recurrent neural networks (RNNs) in traffic prediction: a comprehensive review. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/398.

Radhika:2024:IFA

[3056] M. N. Radhika, Mahendra Mallick, and Xiaoqing Tian. IMM filtering algorithms for a highly maneuvering fighter aircraft: an overview. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/399.

Kor:2024:RLG

[3057] Peyman Kor, Reidar B. Bratvold, and Aojie Hong. Is reinforcement learning good at American option valuation? Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/400.

Kulandai:2024:PLS

[3058] Arockia David Roy Kulandai and Thomas Schwarz. Performance of linear and spiral hashing algorithms. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/9/401.

Chen:2024:IEM

[3059] Jong-Chen Chen and Yin-Zhen Chen. Integrating eye movement, finger pressure, and foot pressure information to build an intelligent driving fatigue detection system. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/9/402.

Zou:2024:LFO

[3060] Kuansheng Zou, Yue Wang, Baowei Liu, and Zhaojun Zhang. Load frequency optimal active disturbance rejection control of hybrid power system. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/9/403.

Carapina:2024:CSM

[3061] Mia Carapina, Ognjen Stanicić, Ivica Dodig, and Davor Cafuta. A comparative study of maze generation algorithms in a game-based mobile learning application for learning basic programming concepts. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/404.

Manawongsakul:2024:SBA

[3062] Dulyawat Manawongsakul and Karn Patanukhom. A segmentation-based automated corneal ulcer grading system for ocular staining images using deep learning and Hough circle transform. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/405.

Binsawad:2024:ADA

[3063] Muhammad Binsawad and Bilal Khan. Advanced detection of abnormal ECG patterns using an optimized LADTree model with enhanced predictive feature: Potential application in CKD. Algorithms (Basel), 17 (9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/9/406.

Wang:2024:IVD

[3064] Dongming Wang, Xing Xu, Xuewen Xia, and Heming Jia. Interactive 3D vase design based on gradient boosting decision trees. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/407.

Walschots:2024:UWT

[3065] Quinty Walschots, Milan Zarchev, Maurits Unkel, and Astrid Kamperman. Using wearable technology to detect, monitor, and predict major depressive disorder — a scoping review and introductory text for clinical professionals. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/408.

Azim:2024:EML

[3066] Sayed Mehedi Azim, Aravind Balasubramanyam, Sheikh Rabiul Islam, Jinglin Fu, and Iman Dehzangi. Explainable machine learning model to accurately predict protein-binding peptides. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/409.

Gyarmati:2024:AON

[3067] László Gyarmati, Csaba Mihálykó, and Éva Orbán-Mihálykó. Algorithm for option number selection in stochastic paired comparison models. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
9/410.

Maione:2024:MLF

[3068] Francesco Maione, Paolo Lino, Guido Maione, and Giuseppe Giannino. A machine learning framework for condition-based maintenance of marine diesel engines: a case study. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/9/411.

Quintana-Rojas:2024:CMB

[3069] Juliana Quintana-Rojas, Rafael Amaya-Gómez, and Nicolas Ratkovich. Compatibility model between encapsulant compounds and antioxidants by the implementation of machine learning. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/412.

Tsai:2024:RAD

[3070] Yu-Ting Tsai, Jin-Shyan Lee, and Chien-Yu Huang. Research on applying deep learning to visual-motor integration assessment systems in pediatric rehabilitation medicine. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/9/413.

Wu:2024:VAM

[3071] Tongyu Wu, Minjie Li, and Shuangjiao Lin. Vulnerability analysis of a multilayer logistics network against cascading failure. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/414.

Berrone:2024:MVP

[3072] Stefano Berrone and Moreno Pintore. Meshfree variational-physics-informed neural networks (MF-VPINN): an adaptive training strategy. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/9/

Rashed:2024:AOC

[3073] Noor A. Rashed, Yossra H. Ali, and Tarik A. Rashid. Advancements in optimization: Critical analysis of evolutionary, swarm, and behaviorbased algorithms. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/416.

Fakhouri:2024:HFV

[3074] Hussam N. Fakhouri, Ahmad Sami Al-Shamayleh, Abdelraouf Ishtaiwi, Sharif Naser Makhadmeh, Sandi N. Fakhouri, and Faten Hamad. Hybrid four vector intelligent metaheuristic with differential evolution for structural single-objective engineering optimization. Algorithms (Basel), 17 (9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/9/417.

Delcea:2024:MFB

[3075] Camelia Delcea, Ionut Nica, Stefan Ionescu, Bianca Cibu, and Horatiu Tibrea. Mapping the frontier: a bibliometric analysis of artificial intelligence applications in local and regional studies. Algorithms (Basel), 17 (9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/9/418.

Haljasmaa:2024:WCC

[3076] Kristina I. Haljasmaa, Andrey M. Bramm, Pavel V. Matrenin, and Stanislav A. Eroshenko. Weather condition clustering for improvement of photovoltaic power plant generation forecasting accuracy. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/419.

Labonte:2024:AFS

[3077] Gilles Labonté. An algorithm to find the shortest path through obstacles of arbitrary shapes and positions in 2D. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/9/420.

Gafarov:2024:CAV

[3078] Evgeny R. Gafarov and Frank Werner. Connected and autonomous vehicle scheduling problems: Some models and algorithms. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/421.

Tamin:2024:CQB

[3079] Owen Tamin and Samsul Ariffin Abdul Karim. Cubic q-Bézier triangular patch for scattered data interpolation and its algorithm. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/9/422.

Abdulla:2024:MBE

[3080] Hemin Sardar Abdulla, Azad A. Ameen, Sarwar Ibrahim Saeed, Ismail Asaad Mohammed, and Tarik A. Rashid. MRSO: Balancing exploration and exploitation through modified rat swarm optimization for global optimization. Algorithms (Basel), 17(9), September 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/9/423.

Amiri:2024:OOS

[3081] Behzad Amiri, Ashkan Jahanbani Ghahfarokhi, Vera Rocca, and Cuthbert Shang Wui Ng. Optimization of offshore saline aquifer CO_{2} storage in Smeaheia using surrogate reservoir models. Algorithms (Basel), 17 (10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/10/452.

Arif:2024:EML

[3082] Muhammad Shoaib Arif, Ateeq Ur Rehman, and Daniyal Asif. Explainable machine learning model for chronic kidney disease prediction. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/443.

Azami:2024:DFA

[3083] Pegah Azami and Kalpdrum Passi. Detecting fake accounts on Instagram using machine learning and hybrid optimization algorithms. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/10/425.

BenJelloun:2024:OCO

[3084] Rim Ben Jelloun, Khalid Jebari, and Abdelaziz El Moujahid. Open competency optimization: a human-inspired optimizer for the dynamic vehicle-routing problem. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/449.

Cardenas-Haro:2024:ESS

[3085] José Antonio Cárdenas-Haro, Mohamed Salem, Abraham N. Aldaco-Gastélum, Roberto López-Avitia, and Maurice Dawson. Enhancing security in social networks through machine learning: Detecting and mitigating Sybil attacks with SybilSocNet. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/442.

Chen:2024:EDR

[3086] Liyuan Chen, Avanka Lowe, and Jing Wang. Early detection of residual/recurrent lung malignancies on post-radiation FDG PET/CT. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/435.

Cui:2024:IIG

[3087] Xining Cui, Kaidong Yang, Xiaoqing Wang, and Peng Duan. An improved iterated greedy algorithm for solving collaborative helicopter rescue routing problem with time window and limited survival time. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/431.

Dev:2024:HRH

[3088] Deepali Goyal Dev and Vishal Bhatnagar. Hybrid RFSVM: Hybridization of SVM and random forest models for detection of fake news. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/459.

Dihmani:2024:NCA

[3089] Hanane Dihmani, Abdelmajid Bousselham, and Omar Bouattane. A new computer-aided diagnosis system for breast cancer detection from thermograms using metaheuristic algorithms and explainable AI. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/10/462.

Dormido-Canto:2024:EPP

[3090] Sebastián Dormido-Canto, Joaquín Rohland, Matías López, Gonzalo Garcia, Ernesto Fabregas, and Gonzalo Farias. Enhancing photovoltaic power predictions with deep physical chain model. Algorithms (Basel), 17 (10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/10/445.

Eisenbach:2024:IRI

[3091] Markus Eisenbach, Andreas Gebhardt, Dustin Aganian, and Horst-Michael Gross. Improving re-identification by estimating and utilizing diverse uncertainty types for embeddings. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/430.

Falcao:2024:FDI

[3092] Dinis Falcão, Francisco Reis, José Farinha, Nuno Lavado, and Mateus Mendes. Fault detection in industrial equipment through analysis of time series stationarity. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/455.

He:2024:API

[3093] Yumin He, Alexandre Dolgui, and Milton Smith. An algorithm for part input sequencing of flexible manufacturing systems with machine disruption. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/10/470.

Hernandez-Barragan:2024:IKR

[3094] Jesus Hernandez-Barragan, Josue Plascencia-Lopez, Michel Lopez-Franco, Nancy Arana-Daniel, and Carlos Lopez-Franco. Inverse kinematics of robotic manipulators based on hybrid differential evolution and Jacobian pseudoinverse approach. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/454.

Huang:2024:SRA

[3095] Jianyun Huang, Qiuwei Yang, Hongfei Cao, and Jiwei Ma. Structural reliability analysis using stochastic finite element method based on Krylov subspace. Algorithms (Basel), 17(10):??, October 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/10/424.

Huynh:2024:RMF

[3096] Harry Hoa Huynh and Cheol-Hong Min. Rotating machinery fault detection using support vector machine via feature ranking. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/10/441.

Irshad:2024:OAR

[3097] Muhammad Nauman Irshad, Imran Ali Khoso, Muhammad Muzamil Aslam, and Rardchawadee Silapunt. Optimized accelerated overrelaxation method for robust signal detection: a metaheuristic approach. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/463.

Jain:2024:PTA

[3098] Anuj Jain and Sartaj Sahni. Polynomial time algorithm for shortest paths in interval temporal graphs. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/468.

Jiao:2024:ESM

[3099] Yanli Jiao, Xiaoyue Zhu, Xinkai Ji, Long Zhang, and Zhibin Li. Empirical studies on merging characteristics in temporary highway work zones with different traffic states and layouts. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/450.

Kshirasagar:2024:ISW

[3100] Shreya Kshirasagar, Andre Guntoro, and Christian Mayr. Impact of sliding window variation and neuronal time constants on acoustic anomaly detection using recurrent spiking neural networks in automotive environment. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/10/440.

Lastrucci:2024:AHB

[3101] Andrea Lastrucci, Antonia Pirrera, Graziano Lepri, and Daniele Giansanti. Algorethics in healthcare: Balancing innovation and integrity in AI development. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/10/432.

Mahmood:2024:MSE

[3102] Nasir Mahmood, Sohail Masood Bhatti, Hussain Dawood, Manas Ranjan Pradhan, and Haseeb Ahmad. Measuring student engagement through behavioral and emotional features using deep-learning models. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/458.

Margenov:2024:PAP

[3103] Svetozar Margenov and Dimitar Slavchev. Performance analysis and parallel scalability of numerical methods for fractional-in-space diffusion problems with adaptive time stepping. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/453.

Maria:2024:SOF

[3104] Gheorghe Maria and Daniela Gheorghe. In silico optimization of a fedbatch bioreactor for tryptophan production using a structured hybrid model and several algorithms including a Pareto-optimal front. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/428.

Matanga:2024:NGO

[3105] Yves Matanga, Pius Owolawi, Chunling Du, and Etienne van Wyk. Niching global optimisation: Systematic literature review. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/10/448.

Mayouf:2024:SOP

[3106] Chiva Mayouf, Ahmed Salhi, Fanta Haidara, Fatima Zahra Aroua, Ragab A. El-Sehiemy, Djemai Naimi, Chouaib Aya, and Cheikh Sidi Ethmane Kane. Solving optimal power flow using new efficient hybrid jellyfish search and moth flame optimization algorithms. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/10/438.

Molnar:2024:DCM

[3107] Miklos Molnar. Degree-constrained minimum spanning hierarchies in graphs. Algorithms (Basel), 17(10):??, October 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/10/467.

Moscato:2024:NAP

[3108] Pablo Moscato and Andrew Ciezak. A new approximation for the perimeter of an ellipse. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/10/464.

Nie:2024:RSD

[3109] Xiaoying Nie, Chunqing Zhou, Zhaoxi Yu, and Gang Yang. Realistic simulation of dissolution process on rock surface. Algorithms (Basel), 17 (10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/10/466.

Oleh:2024:RRT

[3110] Ugonna Oleh, Roman Obermaisser, and Abu Shad Ahammed. A review of recent techniques for human activity recognition: Multimodality, reinforcement learning, and language models. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/434.

Pellicer:2024:TSW

[3111] David S. Pellicer and Emilio Larrodé. Theoretical study of the wear of a reduced-diameter wheel for freight wagons, based on its diameter. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/437.

Pham:2024:ACR

[3112] Tan-Hanh Pham, Godwyll Aikins, Tri Truong, and Kim-Doang Nguyen. Adaptive compensation for robotic joint failures using partially observable reinforcement learning. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/10/436.

Qin:2024:PRI

[3113] Changming Qin, Zhiwen Wang, Linghui Zhang, Qichang Peng, Guixing Lin, and Guanlin Lu. Pedestrian re-identification based on weakly supervised multi-feature fusion. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/426.

Renz:2024:MMT

[3114] Robert Renz and Albert Albers. Multi-material topology optimization on separate tetrahedral meshes with explicit design resolution by means of remeshing. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/10/460.

Rezayiye:2024:MCD

[3115] Reza Khoshkbary Rezayiye, Clemente Ibarra-Castanedo, and Xavier Maldague. Methods for corrosion detection in pipes using thermography: a case study on synthetic datasets. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/10/439.

Saad:2024:ACP

[3116] Marwa Saad, Wesam Almasri, Tanvirul Hye, Monzurul Roni, and Changiz Mohiyeddini. Analysis of ChatGPT-3.5’s potential in generating NBME-standard pharmacology questions: What can be improved? Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/469.

Salgotra:2024:NAP

[3117] Rohit Salgotra, Harmanjeet Singh, Gurpreet Kaur, Supreet Singh, Pratap Singh, and Szymon Lukasik. A novel approach to predict the Asian Exchange Stock Market Index using artificial intelligence. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/17/10/457.

Santos:2024:ASP

[3118] Vitor Santos. Analytical solution for the problem of point location in arbitrary planar domains. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/444.

Seck-Tuoh-Mora:2024:MOM

[3119] Juan Carlos Seck-Tuoh-Mora, Ulises Hernandez-Hurtado, Joselito Medina-Marín, Norberto Hernández-Romero, and Liliana Lizárraga-Mendiola. Multi-objective majority-minority cellular automata algorithm for global and engineering design optimization. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/10/433.

Shafeek:2024:APS

[3120] Yamama A. Shafeek and Hazem I. Ali. Application of particle swarm optimization to a hybrid H_{∞}/sliding mode controller design for the triple inverted pendulum system. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/427.

Stepanova:2024:ASM

[3121] Alina I. Stepanova, Alexandra I. Khalyasmaa, Pavel V. Matrenin, and Stanislav A. Eroshenko. Application of SHAP and multi-agent approach for short-term forecast of power consumption of gas industry enterprises. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/447.

Tsoulos:2024:IGA

[3122] Ioannis G. Tsoulos, Vasileios Charilogis, Dimitrios Tsalikakis, and Alexandros Tzallas. Improving the generalization abilities of constructed neural networks with the addition of local optimization techniques. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
10/446.

Tsoulos:2024:LCN

[3123] Ioannis G. Tsoulos, Vasileios Charilogis, and Dimitrios Tsalikakis. Local crossover: a new genetic operator for grammatical evolution. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/10/461.

Usvakangas:2024:CSF

[3124] Isla Usvakangas, Ronja Tuovinen, and Pekka Neittaanmäki. The current state and future of the urban cold chain: a review of algorithms for environmental optimization. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/465.

Zagubnaya:2024:ACS

[3125] Olga A. Zagubnaya and Yaroslav R. Nartsissov. An algorithm for creating a synaptic cleft digital phantom suitable for further numerical modeling. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/10/451.

Zheng:2024:RAS

[3126] Duo Zheng, Keren Li, and Jie Yang. Response-aided score-matching representative approaches for big data analysis and model selection under generalized linear models. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/10/456.

Zhou:2024:IUB

[3127] Jianmin Zhou, Jian Huang, Jikang Liu, and Jingbo Liu. Improved U2Netbased surface defect detection method for blister tablets. Algorithms (Basel), 17(10):??, October 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/10/429.

Galantai:2024:NMS

[3128] Aurél Galántai. The Nelder–Mead simplex algorithm is sixty years old: New convergence results and open questions. Algorithms (Basel), 17(11): 1–21, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/11/523.

Almaqtari:2024:EAS

[3129] Hani Almaqtari, Feng Zeng, and Ammar Mohammed. Enhancing Arabic sentiment analysis of consumer reviews: Machine learning and deep learning methods based on NLP. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/17/11/495.

Ansori:2024:ELM

[3130] Moch. Fandi Ansori, Kuntjoro Adji Sidarto, Novriana Sumarti, and Iman Gunadi. On the estimation of logistic models with banking data using particle swarm optimization. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/11/507.

Bairooz:2024:IEJ

[3131] Jalal Jabbar Bairooz and Farhad Mardukhi. An innovative enhanced JAYA algorithm for the optimization of continuous and discrete problems. Algorithms (Basel), 17(11):??, November 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/472.

Beltran-Escobar:2024:RRC

[3132] Miguel Beltrán-Escobar, Teresa E. Alarcón, Jesse Y. Rumbo-Morales, Sonia López, Gerardo Ortiz-Torres, and Felipe D. J. Sorcia-Vázquez. A review on resource-constrained embedded vision systems-based tiny machine learning for robotic applications. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/476.

Bo:2024:RCA

[3133] Shi Bo and Minheng Xiao. Root cause attribution of delivery risks via causal discovery with reinforcement learning. Algorithms (Basel), 17(11): ??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/11/498.

Cacao:2024:UIS

[3134] Gabriel Farias Cacao, Dongping Du, and Nandini Nair. Unsupervised image segmentation on 2D echocardiogram. Algorithms (Basel), 17(11): ??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/11/515.

Chakrabarti:2024:SME

[3135] Subir. K. Chakrabarti, Jianan Chen, and Qin Hu. Stationary Markov equilibrium strategies in asynchronous stochastic games: Existence and computation. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/490.

Chen:2024:GES

[3136] Kun Chen, Gang Liu, Genshen Chen, Zhengping Weng, and Qiyu Chen. GLMI: an efficient spatiotemporal index leveraging geohash and piecewise linear models for optimized query performance. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/11/474.

daCruz:2024:LSH

[3137] José Pedro Gomes da Cruz, Matthias Winkenbach, and Hugo Tsugunobu Yoshida Yoshizaki. Local search heuristic for the two-echelon capacitated vehicle routing problem in educational decision support systems. Algorithms (Basel), 17(11):??, November 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/509.

Du:2024:PDD

[3138] Xin Du, Jun Qi, Jiyi Kang, Zezhong Sun, Chunxin Wang, and Jun Xie. Partial discharge data augmentation and pattern recognition method based on DAE-GAN. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/11/487.

Dudin:2024:FAQ

[3139] Alexander Dudin, Sergei Dudin, Agassi Melikov, and Olga Dudina. Framework for analysis of queueing systems with correlated arrival processes and simultaneous service of a restricted number of customers in scenarios with an infinite buffer and retrials. Algorithms (Basel), 17(11): ??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/11/493.

Erdogan:2024:ECB

[3140] Mehmet Erdogan and Mustafa Dogan. Enhanced curvature-based fabric defect detection: a experimental study with Gabor transform and deep learning. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/506.

Favero:2024:PNP

[3141] Luiz Paulo Fávero, Helder Prado Santos, Patrícia Belfiore, Alexandre Duarte, Igor Pinheiro de Araújo Costa, Adilson Vilarinho Terra, Miguel Ângelo Lellis Moreira, Wilson Tarantin Junior, and Marcos dos Santos. A proposal for a new Python library implementing stepwise procedure. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/502.

Gerlein:2024:CPP

[3142] Eduardo A. Gerlein, Francisco Calderón, Martha Zequera-Díaz, and Roozbeh Naemi. Can the plantar pressure and temperature data trend show the presence of diabetes? A comparative study of a variety of machine learning techniques. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/11/519.

Gouton:2024:ESR

[3143] Pierre Gouton, Kacoutchy Jean Ayikpa, and Diarra Mamadou. Estimating the spectral response of eight-band MSFA one-shot cameras using deep learning. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/473.

Grosset:2024:FMA

[3144] Juliette Grosset, Ouzna Oukacha, Alain-Jérôme Fougères, Moïse Djoko-Kouam, and Jean-Marie Bonnin. Fuzzy multi-agent simulation for collective energy management of autonomous industrial vehicle fleets. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/484.

He:2024:AEM

[3145] Mengjie He, Yujie Zhang, Yi Liu, Yang Shen, and Chao Mi. Automated evaluation method for risk behaviors of quay crane operators at ports using virtual reality. Algorithms (Basel), 17(11):??, November 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/11/508.

Hui:2024:QNN

[3146] Peng Hui, Jiamuyang Zhao, Changxin Li, and Qingzhen Zhu. Quotient network — a network similar to ResNet but learning quotients. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/521.

Ikponmwoba:2024:DRL

[3147] Eloghosa Ikponmwoba and Opeoluwa Owoyele. Deephive: a reinforcement learning approach for automated discovery of swarm-based optimization policies. Algorithms (Basel), 17(11):??, November 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/11/500.

Kabashkin:2024:NBF

[3148] Igor Kabashkin. NFT-based framework for digital twin management in aviation component lifecycle tracking. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/494.

Kadyrzhan:2024:AAE

[3149] Aruzhan Kadyrzhan, Akhat Bakirov, Dina Shaltykova, and Ibragim Suleimenov. Application of the algebraic extension method to the construction of orthogonal bases for partial digital convolutions. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/11/496.

Kalitzin:2024:TRA

[3150] Stiliyan Kalitzin. Topological reinforcement adaptive algorithm (TORE-ADA) application to the alerting of convulsive seizures and validation with Monte Carlo numerical simulations. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/516.

Karampinis:2024:APF

[3151] Ioannis Karampinis, Georgios Souliotis, Basil Papadopoulos, and Athanasios Karabinis. An algorithm for producing fuzzy implications via conical sections. Algorithms (Basel), 17(11):??, November 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/11/499.

Khalil:2024:ASE

[3152] Rayan Khalil, Guilherme Vieira Hollweg, Akhtar Hussain, Wencong Su, and Van-Hai Bui. Assessment of solar energy generation toward netzero energy buildings. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/11/528.

Knypinski:2024:MAO

[3153] Lukasz Knypiński, Ramesh Devarapalli, and Marcin Kamiński. Metaheuristic algorithms in optimal design of engineering problems. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/522.

Kojima:2024:OPE

[3154] Takeshi Kojima, Yusaku Yamaguchi, Omar M. Abou Al-Ola, and Tetsuya Yoshinaga. Optimizing parameters for enhanced iterative image reconstruction using extended power divergence. Algorithms (Basel), 17 (11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/11/512.

Krishnamurthy:2024:ISB

[3155] Bhargavi Krishnamurthy and Sajjan G. Shiva. Indetermsoft-set-based D∗ extra lite framework for resource provisioning in cloud computing. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/479.

Levner:2024:FAC

[3156] Eugene Levner and Dmitry Tsadikovich. Fast algorithm for cyber-attack estimation and attack path extraction using attack graphs with AND/ OR nodes. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/504.

Lionnie:2024:NAR

[3157] Regina Lionnie, Julpri Andika, and Mudrik Alaydrus. A new approach to recognize faces amidst challenges: Fusion between the opposite frequencies of the multi-resolution features. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/529.

Liu:2024:SAI

[3158] Jiaqi Liu, Ziyang Wang, HaiYing Wang, and Nalini Ravishanker. Subsampling algorithms for irregularly spaced autoregressive models. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/524.

Lu:2024:TRC

[3159] Aiguo Lu, Pandi Liu, Jie Yang, Zhe Li, and Ke Wang. Tailhook recognition for carrier-based aircraft based on YOLO with bi-level routing attention. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/11/488.

Ludwig:2024:PBA

[3160] Simone A. Ludwig, Jamil Al-Sawwa, and Aaron Mackenzie Misquith. Parallelization of the bison algorithm applied to data classification. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/501.

Machado:2024:MLA

[3161] Rhea Machado, Reshen Soorinarain Dodhy, Atharve Sehgal, Kate Rattigan, Aparna Lalwani, and David Waynforth. A machine learning approach to identifying risk factors for long COVID-19. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/11/485.

Mandravickaite:2024:ASL

[3162] Justina Mandravickaite, Egle Rimkiene, Danguole Kotryna Kapkan, Danguole Kalinauskaite, and Tomas Krilavicius. Automatic simplification of Lithuanian administrative texts. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/533.

Martinez:2024:LCS

[3163] Eulalia Martínez and Arleen Ledesma. Local convergence study for an iterative scheme with a high order of convergence. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/11/481.

Meza:2024:DLA

[3164] Gerardo Meza, Deepak Ganta, and Sergio Gonzalez Torres. Deep learning approach for arm fracture detection based on an improved YOLOv8 algorithm. Algorithms (Basel), 17(11):??, November 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/471.

Moutsopoulou:2024:NIC

[3165] Amalia Moutsopoulou, Markos Petousis, Georgios E. Stavroulakis, Anastasios Pouliezos, and Nectarios Vidakis. Novelty in intelligent controlled oscillations in smart structures. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/11/505.

Neirz:2024:ARS

[3166] Pablo Neirz, Hector Allende, and Carolina Saavedra. Attribute relevance score: a novel measure for identifying attribute importance. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/11/518.

Olanrewaju:2024:MES

[3167] Oludolapo Akanni Olanrewaju, Fabio Luiz Peres Krykhtine, and Felix Mora-Camino. Minimum-energy scheduling of flexible job-shop through optimization and comprehensive heuristic. Algorithms (Basel), 17(11): ??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/11/520.

Padhy:2024:AML

[3168] Manjog Padhy, Umar Muhammad Modibbo, Rasmita Rautray, Subhranshu Sekhar Tripathy, and Sujit Bebortta. Application of machine learning techniques to classify Twitter sentiments using vectorization techniques. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/486.

Parlangeli:2024:DDF

[3169] Gianfranco Parlangeli. Data-driven formation control for multi-vehicle systems induced by leader motion. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/514.

Pinter:2024:NOA

[3170] János D. Pintér, Ignacio Castillo, and Frank J. Kampas. Nonlinear optimization and adaptive heuristics for solving irregular object packing problems. Algorithms (Basel), 17(11):??, November 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/480.

Poghosyan:2024:ECQ

[3171] Arnak Poghosyan, Lusine Poghosyan, and Rafayel Barkhudaryan. On an exact convergence of quasi-periodic interpolations for the polyharmonic-Neumann eigenfunctions. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/11/497.

Ramos-Martinez:2024:SAC

[3172] Moises Ramos-Martinez, Felipe D. J. Sorcia-Vázquez, Gerardo Ortiz-Torres, Mario Martínez García, Mayra G. Mena-Enriquez, Estela Sarmiento-Bustos, Juan Carlos Mixteco-Sánchez, Erasmo Misael Rentería-Vargas, Jesús E. Valdez-Resendiz, and Jesse Yoe Rumbo-Morales. Sleep apnea classification using the mean Euler–Poincaré characteristic and AI techniques. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/11/527.

Renza:2024:AVI

[3173] Diego Renza, Ernesto Moya-Albor, and Adrian Chavarro. Adversarial validation in image classification datasets by means of cumulative spectral gradient. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/531.

Samarathunga:2024:ABT

[3174] Prabhath Samarathunga, Hossein Rezaei, Maheshi Lokumarambage, Thushan Sivalingam, Nandana Rajatheva, and Anil Fernando. An autoencoder-based task-oriented semantic communication system for M2M communication. Algorithms (Basel), 17(11):??, November 2024.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/11/492.

Sanz-Sanchez:2024:PPG

[3175] Alejandro Sanz-Sanchez, Francisco B. García, Pablo Mesas-Lafarga, Joan Prats-Climent, and María José Rodríguez-Álvarez. PIPET: a pipeline to generate PET phantom datasets for reconstruction based on convolutional neural network training. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/11/511.

Shili:2024:YOL

[3176] Mohamed Shili, Osama Sohaib, and Salah Hammedi. You only look once version 5 and deep simple online and real-time tracking algorithms for real-time customer behavior tracking and retail optimization. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/11/525.

Sverko:2024:FED

[3177] Zoran Sverko, Sasa Vlahinić, and Peter Rogelj. A framework for evaluating dynamic directed brain connectivity estimation methods using synthetic EEG signal generation. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/11/517.

Teng:2024:LRM

[3178] Matthew Kit Khinn Teng, Haibo Zhang, and Takeshi Saitoh. LGNMNet-RF: Micro-expression detection using motion history images. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/11/491.

Tsai:2024:EFP

[3179] Ming-Chang Tsai, Scott Thomas, and Marc Klimstra. Exponential functions permit estimation of anaerobic work capacity and critical power from less than 2 min all-out test. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/17/11/532.

Upadhyay:2024:ROD

[3180] Balendu Bhooshan Upadhyay, Shubham Kumar Singh, I. M. Stancu-Minasian, and Andreea Madalina Rusu-Stancu. Robust optimality and duality for nonsmooth multiobjective programming problems with vanishing constraints under data uncertainty. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/11/482.

Vasei:2024:IBR

[3181] Tahereh Vasei, Harshil Gediya, Maryam Ravan, Anand Santhanakrishnan, David Mayor, and Tony Steffert. Investigating brain responses to transcutaneous electroacupuncture stimulation: a deep learning approach. Algorithms (Basel), 17(11):??, November 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/477.

Wei:2024:BNA

[3182] Qingbo Wei, Nanfeng Zhang, Yuan Gao, Cheng Chen, Li Wang, and Jingfeng Yang. Bus network adjustment pre-evaluation based on biometric recognition and travel spatio-temporal deduction. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/11/513.

Wu:2024:LBT

[3183] Liangcheng Wu, Kai Lin, Xiaoyu Lin, and Juan Lin. List-based threshold accepting algorithm with improved neighbor operator for 0–1 knapsack problem. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/478.

Xiao:2024:EER

[3184] Minheng Xiao and Shi Bo. Electroencephalogram emotion recognition via AUC maximization. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/11/489.

Xu:2024:CRA

[3185] Cong Xu and Ravi Sankar. A comprehensive review of autonomous driving algorithms: Tackling adverse weather conditions, unpredictable traffic violations, blind spot monitoring, and emergency maneuvers. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/526.

Xu:2024:VBS

[3186] Wentao Xu, Jin Qian, Yueyang Wu, Shaowei Yan, Yongting Ni, and Guangjin Yang. A VIKOR-based sequential three-way classification ranking method. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/530.

Yuvaraj:2024:MLF

[3187] Rajamanickam Yuvaraj, Shivam Chadha, A. Amalin Prince, M. Murugappan, Md. Sakib Bin Islam, Md. Shaheenur Islam Sumon, and Muhammad E. H. Chowdhury. A machine learning framework for classroom EEG recording classification: Unveiling learning-style patterns. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/11/503.

Zaini:2024:IBF

[3188] Farah Anishah Zaini, Mohamad Fani Sulaima, Intan Azmira Wan Abdul Razak, Mohammad Lutfi Othman, and Hazlie Mokhlis. Improved bacterial foraging optimization algorithm with machine learning-driven shortterm electricity load forecasting: a case study in Peninsular Malaysia. Algorithms (Basel), 17(11):??, November 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
11/510.

Zhou:2024:ASM

[3189] Shengpei Zhou, Nanfeng Zhang, Qin Duan, Jinchao Xiao, and Jingfeng Yang. Automatic scheduling method for customs inspection vehicle relocation based on automotive electronic identification and biometric recognition. Algorithms (Basel), 17(11):??, November 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/11/483.

Zulfqar:2024:AFP

[3190] Sukana Zulfqar, Zenab Elgamal, Muhammad Azam Zia, Abdul Razzaq, Sami Ullah, and Hussain Dawood. ACT-FRCNN: Progress towards transformer-based object detection. Algorithms (Basel), 17(11): ??, November 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/11/475.

Abarzua:2024:ASP

[3191] H. Abarzúa, C. Melo, S. E. Restrepo, S. Vergara, D. Sbarbaro, G. Cañas, G. Lima, G. Saavedra, and J. Cariñe. Adaptive-step perturb-and-observe algorithm for multidimensional phase noise stabilization in fiber-based multi-arm Mach–Zehnder interferometers. Algorithms (Basel), 17(12): ??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/12/534.

Abdel-Hakim:2024:EMA

[3192] Alaa E. Abdel-Hakim, Abdel-Monem M. Ibrahim, Kheir Eddine Bouazza, Wael Deabes, and Abdel-Rahman Hedar. Ellipsoidal K-means: an automatic clustering approach for non-uniform data distributions. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/551.

Ahmed:2024:SNU

[3193] Muhammad Waqas Ahmed, Muhammad Adnan, Muhammad Ahmed, Davy Janssens, Geert Wets, Afzal Ahmed, and Wim Ectors. From stationary to nonstationary UAVs: Deep-learning-based method for vehicle speed estimation. Algorithms (Basel), 17(12):??, December 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/12/558.

Alanis:2024:CNE

[3194] Alma Y. Alanis, Gustavo Munoz-Gomez, Nancy F. Ramirez, Oscar D. Sanchez, and Jesus G. Alvarez. A complex network epidemiological approach for infectious disease spread control with time-varying connections. Algorithms (Basel), 17(12):??, December 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/577.

Alsaffar:2024:EID

[3195] Ali Mohammed Alsaffar, Mostafa Nouri-Baygi, and Hamed Zolbanin. Enhancing intrusion detection systems with dimensionality reduction and multi-stacking ensemble techniques. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/550.

Anes:2024:HFR

[3196] Vitor Anes and António Abreu. A hybrid FMEA-ROC-CoCoSo approach for improved risk assessment and reduced complexity in failure mode prioritization. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/585.

Arteaga:2024:SAE

[3197] Marco A. Arteaga, Evert J. Guajardo-Benavides, and Pablo Sánchez-Sánchez. Stability analysis and experimental validation of standard proportional-integral-derivative control in bilateral teleoperators with time-varying delays. Algorithms (Basel), 17(12):??, December 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/12/580.

Bonny:2024:PMA

[3198] Talal Bonny, Wafaa Al Nassan, Khaled Obaideen, Tamer Rabie, Maryam Nooman AlMallahi, and Swati Gupta. Primary methods and algorithms in artificial-intelligence-based dental image analysis: a systematic review. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/567.

Castellanos-Cardenas:2024:RDD

[3199] Duby Castellanos-Cárdenas, Norha L. Posada, Andrés Orozco-Duque, Lina M. Sepúlveda-Cano, Fabio Castrillón, Oscar E. Camacho, and Rafael E. Vásquez. A review on data-driven model-free sliding mode control. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/12/543.

Castillo:2024:EDL

[3200] Sebastian-Josue Castillo and Ferenc Izsák. Efficient discretization of the Laplacian: Application to moving boundary problems. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/12/541.

Dai:2024:STF

[3201] Zhuangzhuang Dai, Vincent Gbouna Zakka, Luis J. Manso, Martin Rudorfer, Ulysses Bernardet, Johanna Zumer, and Manolya Kavakli-Thorne. Sensors, techniques, and future trends of human-engagement-enabled applications: a review. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/560.

deOliveira:2024:DIN

[3202] Izabela Cristina de Oliveira, Ricardo Gava, Dthenifer Cordeiro Santana, Ana Carina da Silva Cândido Seron, Larissa Pereira Ribeiro Teodoro, Mayara Favero Cotrim, Regimar Garcia dos Santos, Rita de Cássia Félix Alvarez, Carlos Antonio da Silva Junior, Fábio Henrique Rojo Baio, and Paulo Eduardo Teodoro. Detection of irrigated and non-irrigated soybeans using hyperspectral data in machine-learning models. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/542.

Dong:2024:SSE

[3203] Chenyang Dong, Shiyu Xi, Yinchao Che, Shufeng Xiong, Xinming Ma, Lei Xi, and Shuping Xiong. Single-stage entity-relation joint extraction of pesticide registration information based on HT-BES multi-dimensional labeling strategy. Algorithms (Basel), 17(12):??, December 2024. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/17/12/559.

Du:2024:RPD

[3204] Sheng Du, Zixin Huang, Li Jin, and Xiongbo Wan. Recent progress in data-driven intelligent modeling and optimization algorithms for industrial processes. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/569.

Espiga-Fernandez:2024:SAP

[3205] Francisco Espiga-Fernández, Álvaro García-Sánchez, and Joaquín Ordieres-Meré. A systematic approach to portfolio optimization: a comparative study of reinforcement learning agents, market signals, and investment horizons. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/12/570.

Fakhouri:2024:HAP

[3206] Hussam N. Fakhouri, Mohannad S. Alkhalaileh, Faten Hamad, Najem N. Sirhan, and Sandi N. Fakhouri. Hybrid Arctic puffin algorithm for solving design optimization problems. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/589.

Fasano:2024:IDQ

[3207] Giovanni Fasano, Christian Piermarini, and Massimo Roma. Issues on a 2-dimensional quadratic sub-problem and its applications in nonlinear programming: Trust-region methods (TRMs) and linesearch based methods (LBMs). Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/563.

Fraihat:2024:VAB

[3208] Salam Fraihat, Qusai Shambour, Mohammed Azmi Al-Betar, and Sharif Naser Makhadmeh. Variational autoencoders-based algorithm for multi-criteria recommendation systems. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/561.

Ghiani:2024:ILL

[3209] Gianpaolo Ghiani, Gianluca Solazzo, and Gianluca Elia. Integrating large language models and optimization in semi- structured decision making: Methodology and a case study. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/582.

Han:2024:MSI

[3210] Tao Han, Tingting Li, Quanzeng Liu, Yourui Huang, and Hongping Song. A multi-strategy improved honey badger algorithm for engineering design problems. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/573.

Hong:2024:ADV

[3211] Chih-Chiang Hong. Advanced dynamic vibration control algorithms of materials Terfenol-D Si_{3}N_{4} and SUS304 plates/cylindrical shells with velocity feedback control law. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/539.

Hossen:2024:TIF

[3212] Md Helal Hossen, Daniel Gibney, and Sharma V. Thankachan. Text indexing for faster gapped pattern matching. Algorithms (Basel), 17(12): ??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/12/537.

Kanda:2024:ABM

[3213] Akichy Adon Jean Rodrigue Kanda, Amanvon Ferdinand Atta, Zacrada Françoise Odile Trey, Michel Babri, and Ahmed Dooguy Kora. Anchorbased method for inter-domain mobility management in software-defined networking. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/566.

Kavouras:2024:DBL

[3214] Ioannis Kavouras, Ioannis Rallis, Dimitra Zouli, Emmanuel Sardis, Eftychios Protopapadakis, Anastasios Doulamis, and Nikolaos Doulamis. A dilemma-based learning-to-rank approach for generative design in urban architectural regeneration. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/538.

Kawano:2024:LMI

[3215] Yuta Kawano, Koichi Kobayashi, and Yuh Yamashita. Linear matrix inequality-based design of structured sparse feedback controllers for sensor and actuator networks. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/590.

Khodamoradi:2024:AIL

[3216] Farshad Khodamoradi, Javad Rezazadeh, and John Ayoade. Accurate indoor localization with IoT devices and advanced fingerprinting methods. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/12/544.

Kim:2024:UPR

[3217] Yong Han Kim, Wei Ye, Ritbik Kumar, Finn Bail, Julia Dvorak, Yanchao Tan, Marvin Carl May, Qing Chang, Ragu Athinarayanan, Gisela Lanza, John W. Sutherland, Xingyu Li, and Chandra Nath. Unlocking the potential of remanufacturing through machine learning and datadriven models — a survey. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/562.

Kitova:2024:SDM

[3218] Kalina Kitova, Ivan Ivanov, and Vincent Hooper. Stroke dataset modeling: Comparative study of machine learning classification methods. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/571.

Konstantopoulou:2024:GEI

[3219] Despoina Konstantopoulou, Paraskevi Zacharia, Michail Papoutsidakis, Helen C. Leligou, and Charalampos Patrikakis. GAGAN: Enhancing image generation through hybrid optimization of genetic algorithms and deep convolutional generative adversarial networks. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/12/584.

Kouzari:2024:TSP

[3220] Elia Kouzari and Ioannis Stamelos. A tertiary study for process mining. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/548.

Krivovichev:2024:SOE

[3221] Gerasim V. Krivovichev. Stability optimization of explicit Runge–Kutta methods with higher-order derivatives. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/535.

Kucukmehmetoglu:2024:HEG

[3222] Mehmet Küçükmehmetoglu, Yasin Fahjan, and Muhammed Ziya Paköz. Handling exponentially growing strategies in spatial cooperative games: The case of the European Union. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/554.

Lagana:2024:OAN

[3223] Filippo Laganà, Salvatore A. Pullano, Giovanni Angiulli, and Mario Versaci. Optimized analytical-numerical procedure for ultrasonic sludge treatment for agricultural use. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/592.

Lai:2024:BSC

[3224] Fei Lai and Chaoqun Huang. Balancing safety and comfort: a novel automatic braking control method using seventh-degree polynomials. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/545.

Lazaro:2024:IAU

[3225] Carla Lázaro and Cecilio Angulo. Iterative application of UMAP-based algorithms for fully synthetic healthcare tabular data generation. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/591.

Lin:2024:DAV

[3226] Yao-San Lin, Hung-Yu Chen, Mei-Ling Huang, and Tsung-Yu Hsieh. Data augmentation for voiceprint recognition using generative adversarial networks. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/583.

Mann:2024:PRD

[3227] Kevin Mann and Henning Fernau. Perfect Roman domination: Aspects of enumeration and parameterization. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/576.

McKendall:2024:SAA

[3228] Alan McKendall and Yugesh Dhungel. A simulated annealing algorithm for the generalized quadratic assignment problem. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/12/540.

Nagy:2024:TWB

[3229] Balázs Nagy, Ahmad Nader Fasseeh, Jonathan D. Campbell, Zoltán Kaló, Kareem Ahmed El-Fass, Rok Hren, and Bertalan Németh. Tradeoffs when building and running cohort and patient-level Markov simulation models. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/546.

Naimipour:2024:UAD

[3230] Naveed Naimipour, Shahin Khobahi, Mojtaba Soltanalian, Haleh Safavi, and Harry C. Shaw. Unfolded algorithms for deep phase retrieval. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/587.

Nasir:2024:IEC

[3231] Inzamam Mashood Nasir, Sara Tehsin, Robertas Damasevicius, and Rytis Maskeliunas. Integrating explanations into CNNs by adopting spiking attention block for skin cancer detection. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/557.

Nassih:2024:RPB

[3232] Rym Nassih and Abdelaziz Berrado. A random PRIM based algorithm for interpretable classification and advanced subgroup discovery. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/565.

Nigro:2024:VME

[3233] Libero Nigro. Verifying mutual exclusion algorithms with non-atomic registers. Algorithms (Basel), 17(12):??, December 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/536.

Pan:2024:IFD

[3234] Yuxin Pan, Yinsheng Chen, Xihong Fei, Kang Wang, Tian Fang, and Jing Wang. Intelligent fault diagnosis for rotating mechanical systems: an improved multiscale fuzzy entropy and support vector machine algorithm. Algorithms (Basel), 17(12):??, December 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/588.

Papadakis:2024:ARS

[3235] Harris Papadakis, Paraskevi Fragopoulou, and Costas Panagiotakis. Applying recommender systems to predict personalized film age ratings for parents. Algorithms (Basel), 17(12):??, December 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/578.

Parasyris:2024:DAA

[3236] Antonios Parasyris, Vassiliki Metheniti, George Alexandrakis, Georgios V. Kozyrakis, and Nikolaos A. Kampanis. Data assimilated atmospheric forecasts for digital twin of the ocean applications: a case study in the South Aegean, Greece. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/586.

Raftopoulos:2024:FTS

[3237] George Raftopoulos, Gregory Davrazos, and Sotiris Kotsiantis. Fair and transparent student admission prediction using machine learning models. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/572.

Rahman:2024:UUA

[3238] Md. Mahfuzur Rahman, Sunzida Siddique, Marufa Kamal, Rakib Hossain Rifat, and Kishor Datta Gupta. UAV (Unmanned Aerial Vehicle):

Diverse applications of UAV datasets in segmentation, classification, detection, and tracking. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/17/12/594.

Rehman:2024:HAR

[3239] Saeed Ur Rehman, Anwar Ali, Adil Mehmood Khan, and Cynthia Okpala. Human activity recognition: a comparative study of validation methods and impact of feature extraction in wearable sensors. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/17/12/556.

Saldana-Ulloa:2024:TGN

[3240] Diego Saldaña-Ulloa, Guillermo De Ita Luna, and J. Raymundo Marcial-Romero. A temporal graph network algorithm for detecting fraudulent transactions on online payment platforms. Algorithms (Basel), 17(12): ??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/12/552.

Shams:2024:CAP

[3241] Mudassir Shams and Bruno Carpentieri. Computational analysis of parallel techniques for nonlinear biomedical engineering problems. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/575.

Sheng:2024:MAC

[3242] Lei Sheng, Honghui Chen, and Xiliang Chen. A multi-agent centralized strategy gradient reinforcement learning algorithm based on state transition. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
17/12/579.

Solano-Lopez:2024:ESG

[3243] Pablo Solano-López, Jorge Saavedra, and Raúl Molina. Exploratory study of a green function based solver for nonlinear partial differential equations. Algorithms (Basel), 17(12):??, December 2024. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/564.

Solarte-Sanchez:2024:QRB

[3244] Miguel Solarte-Sanchez, David Marquez-Viloria, Andrés E. Castro-Ospina, Erick Reyes-Vera, Neil Guerrero-Gonzalez, and Juan Botero-Valencia. m-QAM receiver based on data stream spectral clustering for optical channels dominated by nonlinear phase noise. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/17/12/553.

Syzonov:2024:NIF

[3245] Oleksandr Syzonov, Stefania Tomasiello, and Nicola Capuano. New insights into fuzzy genetic algorithms for optimization problems. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/17/
12/549.

Tian:2024:UTA

[3246] Haitao Tian and Pierre Payeur. Unsupervised temporal adaptation in skeleton-based human action recognition. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/581.

Wang:2024:DDT

[3247] Li Wang, Feng Chen, Yujia Hu, Zhiyao Zheng, and Kexin Zhang. Disassembly of distribution transformers based on multimodal data recognition and collaborative processing. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/17/12/595.

Xiao:2024:ODN

[3248] Chun Xiao, Qiong Cao, Yulu Ren, and Xiaoqing Han. Optimization of distribution network current protection for inverter-based distributed power access. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/555.

Zhang:2024:MMS

[3249] Jianqiang Zhang, Renyao Chen, Shengwen Li, Tailong Li, and Hong Yao. MGKGR: Multimodal semantic fusion for geographic knowledge graph representation. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/17/12/593.

Zhou:2024:MAD

[3250] Shengpei Zhou, Nanfeng Zhang, Qin Duan, Xiaosong Liu, Jinchao Xiao, Li Wang, and Jingfeng Yang. Monitoring and analyzing driver physiological states based on automotive electronic identification and multimodal biometric recognition methods. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/547.

Zhou:2024:RII

[3251] Ruihong Zhou, Baokang Xiang, Long Wu, Yanli Hu, Litong Dou, and Kaifeng Huang. Research on intelligent identification method for pantograph positioning and skateboard structural anomalies based on improved YOLO v8 algorithm. Algorithms (Basel), 17(12):??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/17/12/574.

Zupan:2024:RHO

[3252] Hugo Zupan, Niko Herakovic, and Janez Zerovnik. A robust heuristics for the online job shop scheduling problem. Algorithms (Basel), 17(12): ??, December 2024. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/17/12/568.

Kim:2025:EAA

[3253] Hyundong Kim, Soobin Kwak, Moumni Mohammed, Seungyoon Kang, Seokjun Ham, and Junseok Kim. An efficient and accurate adaptive timestepping method for the Landau–Lifshitz equation. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/1.

Abd-Elhameed:2025:NGC

[3254] Waleed Mohamed Abd-Elhameed, Ramy M. Hafez, Anna Napoli, and Ahmed Gamal Atta. A new generalized Chebyshev matrix algorithm for solving second-order and telegraph partial differential equations. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/2.

Grieco:2025:AAD

[3255] Antonio Grieco, Pierpaolo Caricato, and Paolo Margiotta. Algorithmic advances for 1.5-dimensional two-stage cutting stock problem. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/3.

Li:2025:MSA

[3256] Hongyu Li, Lei Chen, Jian Zhang, and Muxi Li. A multi-surrogate assisted multi-tasking optimization algorithm for high-dimensional expensive problems. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/4.

Qiao:2025:ODO

[3257] Yulong Qiao, Zixing Wu, Qian Zhang, Qin Xu, and Ge Jin. Optimal design of one-sided exponential adaptive EWMA scheme based on median run length. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/5.

Zhang:2025:MIV

[3258] Xu Zhang, Xiaojun Kang, Hong Yao, and Lijun Dong. Measuring the inferential values of relations in knowledge graphs. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/6.

Belacel:2025:CRC

[3259] Nabil Belacel. A closest resemblance classifier with feature interval learning and outranking measures for improved performance. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/1/7.

Ding:2025:FVR

[3260] Jian Ding, Yun Yi, Tinghua Wang, and Tao Tian. Fire video recognition based on local and global adaptive enhancement. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/8.

Babaei:2025:MOO

[3261] Salar Babaei, Mehran Khalaj, Mehdi Keramatpour, and Ramin Enayati. Multi-objective optimization for green BTS site selection in telecommunication networks using NSGA-II and MOPSO. Algorithms (Basel), 18 (1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/9.

Barba:2025:OMP

[3262] Paolo Di Barba, Arash Ghafoorinejad, Maria Evelina Mognaschi, Fabrizio Dughiero, Michele Forzan, and Elisabetta Sieni. Optimal multiphysics synthesis of a dual-frequency power inductor using deep neural networks and Gaussian process regression. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/1/10.

Guo:2025:ADR

[3263] Qiang Guo, Fenghe Li, Hengwen Liu, and Jin Guo. Anomaly detection and root cause analysis for energy consumption of medium and heavy plate: a novel method based on Bayesian neural network with Adam variational inference. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/11.

Ghorbani:2025:AMC

[3264] Mohammad Kazem Ghorbani, Nasser Talebbeydokhti, Hossein Hamidifar, Mehrshad Samadi, Michael Nones, Fatemeh Rezaeitavabe, and Shabnam Heidarifar. Application of multi-criteria decision-making models for assessment of education quality in water resources engineering. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/12.

Mo:2025:MOO

[3265] Dong Mo, Qiuwen Li, Yan Sun, Yixin Zhuo, and Fangming Deng. Multiobjective optimization scheduling of a wind-solar energy storage microgrid based on an improved OGGWO algorithm. Algorithms (Basel), 18 (1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/13.

Drousiotis:2025:MPS

[3266] Efthyvoulos Drousiotis, Alessandro Varsi, Alexander M. Phillips, Simon Maskell, and Paul G. Spirakis. A massively parallel SMC sampler for decision trees. Algorithms (Basel), 18(1), January 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/14.

Dolgui:2025:SIS

[3267] Alexandre Dolgui, David Lemoine, María I. Restrepo, and Frank Werner. Special issue on scheduling theory and algorithms for sustainable manufacturing. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/15.

Khan:2025:SIA

[3268] Muhammad Adnan Khan. Special issue “Algorithms for Feature Selection (2nd Edition)”. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/16.

Mao:2025:FIP

[3269] Dongfang Mao, Haojie Lin, and Xuyang Lou. Fingerprinting indoor positioning based on improved sequential deep learning. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/17.

Zhang:2025:ODC

[3270] Jing Zhang, Xiaokai Zhu, Te Chen, and Guowei Dou. Optimal dynamics control in trajectory tracking of industrial robots based on adaptive Gaussian pseudo-spectral algorithm. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/1/18.

Aguilar-Gonzalez:2025:EME

[3271] Abiel Aguilar-González and Alejandro Medina Santiago. Ego-motion estimation for autonomous vehicles based on genetic algorithms and CUDA parallel processing. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/19.

Mim:2025:EEL

[3272] Mithila Akter Mim, M. R. Khatun, Muhammad Minoar Hossain, Wahidur Rahman, and Arslan Munir. Exploring early learning challenges in children utilizing statistical and explainable machine learning. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/20.

Su:2025:DRC

[3273] Weixian Su and Donglin Jing. DDL R-CNN: Dynamic direction learning R-CNN for rotated object detection. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/1/21.

Pacella:2025:SFS

[3274] Massimo Pacella, Antonio Papa, Gabriele Papadia, and Emiliano Fedeli. A scalable framework for sensor data ingestion and real-time processing in cloud manufacturing. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/1/22.

Calzada-Garcia:2025:RIK

[3275] Ana Calzada-Garcia, Juan G. Victores, Francisco J. Naranjo-Campos, and Carlos Balaguer. A review on inverse kinematics, control and planning for robotic manipulators with and without obstacles via deep neural networks. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/23.

Liu:2025:MST

[3276] Tzu-Hsin Liu, He-Yao Hsu, and Fu-Min Chang. Multi-server two-way communication retrial queue subject to disaster and synchronous working vacation. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/24.

Esquivel-Cruz:2025:HEV

[3277] Eduardo Esquivel-Cruz, Francisco Beltran-Carbajal, Ivan Rivas-Cambero, José Humberto Arroyo-Núñez, Ruben Tapia-Olvera, and Daniel Guillen. Hybrid empirical and variational mode decomposition of vibratory signals. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/25.

Kshemkalyani:2025:IRB

[3278] Ajay D. Kshemkalyani and Anshuman Misra. Impossibility results for Byzantine-tolerant state observation, synchronization, and graph computation problems. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/26.

Ding:2025:YTO

[3279] Dong Ding, Zhengrong Deng, and Rui Yang. YOLO-TC: an optimized detection model for monitoring safety-critical small objects in tower crane operations. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/27.

Shili:2025:SIC

[3280] Mohamed Shili, Salah Hammedi, and Mahmoud Elkhodr. Spatial intelligence in e-commerce: Integrating mobile agents with GISs for a dynamic recommendation system. Algorithms (Basel), 18(1), January 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/1/28.

Martinenghi:2025:PCG

[3281] Davide Martinenghi. Parallelizing the computation of grid resistance to measure the strength of skyline tuples. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/1/29.

Liu:2025:RMS

[3282] Binglin Liu, Qian Li, Zhihua Zheng, Yanjia Huang, Shuguang Deng, Qiongxiu Huang, and Weijiang Liu. A review of multi-source data fusion and analysis algorithms in smart city construction: Facilitating real estate management and urban optimization. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/1/30.

Feng:2025:CPD

[3283] Wang Feng, Sihai Tang, Shengze Wang, Ying He, Donger Chen, Qing Yang, and Song Fu. Characterizing perception deep learning algorithms and applications for vehicular edge computing. Algorithms (Basel), 18 (1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/31.

Pan:2025:DPC

[3284] Tingzhe Pan, Jue Hou, Xin Jin, Chao Li, Xinlei Cai, and Xiaodong Zhou. Differentially private clustered federated load prediction based on the Louvain algorithm. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/32.

deLucena:2025:MCT

[3285] Mateus Martinez de Lucena, Josafat Leal Ribeiro, Matheus Wagner, and Antônio Augusto Fröhlich. Multi-component temporal-correlation seismic data compression algorithm based on the PCA and DWT. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/1/33.

Bouchouras:2025:IAI

[3286] Georgios Bouchouras and Konstantinos Kotis. Integrating artificial intelligence, Internet of Things, and sensor-based technologies: a systematic review of methodologies in autism spectrum disorder detection. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/34.

Barragan:2025:EHE

[3287] Guillermo Barragan, Sebastian Atarihuana, Edgar Cando, and Victor Hidalgo. Enhancing hydraulic efficiency of Pelton turbines through computational fluid dynamics and metaheuristic optimization. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/1/35.

Mo:2025:DNE

[3288] Dong Mo, Qiuwen Li, Yan Sun, Yixin Zhuo, and Fangming Deng. Design of a new energy microgrid optimization scheduling algorithm based on improved grey relational theory. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/1/36.

Husain:2025:SVS

[3289] Gazi Husain, Daniel Nasef, Rejath Jose, Jonathan Mayer, Molly Bekbolatova, Timothy Devine, and Milan Toma. SMOTE vs. SMOTEENN: a study on the performance of resampling algorithms for addressing class imbalance in regression models. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/1/37.

deArmas:2025:SVN

[3290] Jesica de Armas and José A. Moreno-Pérez. A survey on variable neighborhood search for sustainable logistics. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/1/38.

Zungu:2025:SID

[3291] Ntandoyenkosi Zungu, Peter Olukanmi, and Pitshou Bokoro. Synth-SecureNet: an improved deep learning architecture with application to intelligent violence detection. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/1/39.

Garcke:2025:FDS

[3292] Jochen Garcke and Steffen Ruttscheidt. Finite differences on sparse grids for continuous-time heterogeneous agent models. Algorithms (Basel), 18 (1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/40.

Wang:2025:MOP

[3293] Zhaohui Wang, Meng Zhang, Shanqing Liang, Shuang Yu, Chengchun Zhang, and Sheng Du. A multi-objective path-planning approach for multi-scenario urban mobility needs. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/1/41.

Liu:2025:MMS

[3294] Botao Liu, Changqi Shi, and Ming Zhao. MSEANet: Multi-scale selective edge aware network for polyp segmentation. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/1/42.

Kavakli-Thorne:2025:AAP

[3295] Manolya Kavakli-Thorne and Zhuangzhuang Dai. AI algorithms for positive change in digital futures. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/1/43.

Hamed:2025:LBF

[3296] Alaa Hamed, Arabi Keshk, and Anas Youssef. A lexicon-based framework for mining and analysis of Arabic comparative sentences. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/44.

Hernandez-Venegas:2025:VWC

[3297] Ulises Hernandez-Venegas, Jesus Hernandez-Barragan, Irene Gomez Jimenez, Gabriel Martinez-Soltero, and Alma Y. Alanis. Vertex-weighted consensus-based formation control with area constraints and collision avoidance. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/45.

Muhammad:2025:APL

[3298] Iftikhar Muhammad and Marco Rospocher. On assessing the performance of LLMs for target-level sentiment analysis in financial news headlines. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/
46.

Liu:2025:AED

[3299] Binglin Liu, Weijia Zeng, Weijiang Liu, Yi Peng, and Nini Yao. AIenhanced design and application of high school geography field studies in China: a case study of the Yellow (Bohai) Sea migratory bird habitat curriculum. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/1/47.

Zhu:2025:GGE

[3300] Ke Zhu, Donghui Luo, Zhengzheng Fu, Zhihang Xue, and Xianghang Bu. Grounding grid electrical impedance imaging method based on an improved conditional generative adversarial network. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/48.

Ilyushin:2025:AMA

[3301] Pavel Ilyushin, Boris Papkov, Aleksandr Kulikov, and Konstantin Suslov. Algorithm and methods for analyzing power consumption behavior of industrial enterprises considering process characteristics. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/1/49.

Mendoza-Andrade:2025:SAC

[3302] Nidia Mendoza-Andrade, Efrain Ruiz y Ruiz, and Suemi Rodriguez-Romo. Solution algorithms for the capacitated location tree problem with interconnections. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/1/50.

Leal:2025:CML

[3303] Hugo M. Leal, Ramiro S. Barbosa, and Isabel S. Jesus. Control of a mobile line-following robot using neural networks. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/1/51.

Fang:2025:ALP

[3304] Zhijian Fang, Xiang Gao, Huaxiong Zhang, Jingpeng Tang, and Qiang Gao. Application layer protocol identification method based on ResNet. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/1/52.

Zeng:2025:ASP

[3305] Xinhong Zeng and Yongxiang Wang. Analysis and simulation of polishing robot operation trajectory planning. Algorithms (Basel), 18(1), January 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/1/53.

Raitsis:2025:COC

[3306] Tomer Raitsis, Yossi Elgazari, Guy E. Toibin, Yotam Lurie, Shlomo Mark, and Oded Margalit. Code obfuscation: a comprehensive approach to detection, classification, and ethical challenges. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/54.

Khan:2025:MDA

[3307] Shehroz S. Khan, Tianyu Shi, Simon Donato-Woodger, and Charlene H. Chu. Mitigating digital ageism in skin lesion detection with adversarial learning. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/55.

Wang:2025:ANE

[3308] Chi Wang, Lin Shi, and Junru Luo. Adaptive noise exploration for neural contextual multi-armed bandits. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/56.

Yoon:2025:QAU

[3309] Hyun Jin Yoon, Daye Yoon, Sungmin Jun, Young Jin Jeong, and Do-Young Kang. Quantitative analysis using PMOD and FreeSurfer for three types of radiopharmaceuticals for Alzheimer’s disease diagnosis. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
2/57.

Mohammad:2025:CEM

[3310] Mohammad Mohammad, Olga Ibryaeva, Vladimir Sinitsin, and Victoria Eremeeva. A computationally efficient method for the diagnosis of defects in rolling bearings based on linear predictive coding. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/58.

Mariettou:2025:AIA

[3311] Savina Mariettou, Constantinos Koutsojannis, and Vassilios Triantafillou. Artificial intelligence and algorithmic approaches of health security systems: a review. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/59.

Dronyuk:2025:ACG

[3312] Ivanna Dronyuk. Algorithms for calculating generalized trigonometric functions. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/60.

Hsieh:2025:IHG

[3313] Yu-Heng Hsieh, Chen-Chun Kao, and Shyan-Ming Yuan. Imitating human go players via vision transformer. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/18/2/61.

Wang:2025:ADH

[3314] Yunge Wang, Lingling Zhang, Tong Si, Graham Bishop, and Haijun Gong. Anomaly detection in high-dimensional time series data with scaled Bregman divergence. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/62.

Gameiro:2025:EPB

[3315] Tiago Gameiro, Tiago Pereira, Hamid Moghadaspoura, Francesco Di Giorgio, Carlos Viegas, Nuno Ferreira, João Ferreira, Salviano Soares, and António Valente. Evaluation of PID-based algorithms for UGVs. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
2/63.

Abdelfattah:2025:CSD

[3316] Waleed Mohammed Abdelfattah. A comparative study of differential quadrature methods for METE nanobeam vibrations. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/64.

Shao:2025:SNL

[3317] Yuxiang Shao, Longyang Zhou, Xiang Li, Chunsheng Feng, and Xinyu Jin. SA-Net: Leveraging spatial correlations spatial-aware net for multiperspective robust estimation algorithm. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/2/65.

Ji:2025:AFO

[3318] Ruipu Ji, Shokrullah Sorosh, Eric Lo, Tanner J. Norton, John W. Driscoll, Falko Kuester, Andre R. Barbosa, Barbara G. Simpson, and Tara C. Hutchinson. Application framework and optimal features for UAV-based earthquake-induced structural displacement monitoring. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/2/66.

Wang:2025:CMS

[3319] Xiao Wang, Wenshuai Suo, and Rong Wang. CSpredR: a multi-site mRNA subcellular localization prediction method based on fusion encoding and hybrid neural networks. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/67.

Nguyen:2025:CRP

[3320] Linh Thi Hoai Nguyen, Tomoyuki Shirai, and Takeshi Tsuji. Characterization of rock pore geometry and mineralization process via a random walk-based clogging scheme. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/68.

Kamal:2025:EHD

[3321] Hesham Kamal and Maggie Mashaly. Enhanced hybrid deep learning models-based anomaly detection method for two-stage binary and multiclass classification of attacks in intrusion detection systems. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/69.

Masot-Llima:2025:ADT

[3322] Sergi Masot-Llima and Artur Garcia-Saez. Advantages of density in tensor network geometries for gradient-based training. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/70.

Mu:2025:ENE

[3323] Pengyu Mu, Hongwei Zhao, and Ke Ma. EGSDK-Net: Edge-guided stepwise dual kernel update network for panoptic segmentation. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/71.

Jha:2025:DWY

[3324] Aditya Jha and Peter A. Gloor. Do what you say-computing personal values associated with professions based on the words they use. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/72.

Li:2025:IMF

[3325] Jiajin Li, Steve King, and Ian Jennions. Intelligent multi-fault diagnosis for a simplified aircraft fuel system. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/73.

Theodorakopoulos:2025:OAS

[3326] Leonidas Theodorakopoulos, Aristeidis Karras, and George A. Krimpas. Optimizing Apache spark MLlib: Predictive performance of large-scale models for big data analytics. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/74.

Liu:2025:NMN

[3327] Lei Liu, Ziyi Wang, Xiaohan Zhang, Yan Zhuang, and Yongbo Liang. A novel model for noninvasive haemoglobin detection based on visibility network and clustering network for multi-wavelength PPG signals. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/2/75.

Ambashankar:2025:BSC

[3328] Akash Ambashankar and Hovhannes A. Harutyunyan. Broadcasting in stars of cliques and path-connected cliques. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/2/76.

Torkey:2025:SDM

[3329] Hanaa Torkey, Sonia Hashish, Samia Souissi, Ezz El-Din Hemdan, and Amged Sayed. Seizure detection in medical IoT: Hybrid CNN-LSTM-GRU model with data balancing and XAI integration. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/77.

Mohammed:2025:TDO

[3330] Aqeel Abdulazeez Mohammed, Ahlam Hanoon Al-sudani, Alaa M. Abdul-Hadi, Almuntadher Alwhelat, Basheera M. Mahmmod, Sadiq H. Abdulhussain, Muntadher Alsabah, and Abir Hussain. Threedimensional object recognition using orthogonal polynomials: an embedded kernel approach. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/78.

Marusak:2025:AMA

[3331] Piotr M. Marusak. Analytical MPC algorithm using steady-state process model. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/79.

Li:2025:ESH

[3332] Haitao Li, Chenyu Wang, Shufu Yuan, Hui Zhu, Bo Li, Yuexin Liu, and Li Sun. Energy scheduling of hydrogen hybrid UAV based on model predictive control and deep deterministic policy gradient algorithm. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/2/80.

Bucura:2025:MLM

[3333] Cosmina Elena Bucura and Paolo Giudici. Machine learning models to predict Google stock prices. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/81.

Usman:2025:PDD

[3334] Cathryn Usman, Saeed Ur Rehman, Anwar Ali, Adil Mehmood Khan, and Baseer Ahmad. Pneumonia disease detection using chest X-rays and machine learning. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/82.

Kovacs:2025:PDE

[3335] László Kovács and Ali Jlidi. Process discovery for event logs with multioccurrence event types. Algorithms (Basel), 18(2), February 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/2/83.

Colucci:2025:APM

[3336] Giovanni Paolo Colucci, Paola Battilani, Marco Camardo Leggieri, and Daniele Trinchero. Algorithms for plant monitoring applications: a comprehensive review. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/84.

Moiseev:2025:OEB

[3337] Ilia Moiseev, Ksenia Balabaeva, and Sergey Kovalchuk. Open and extensible benchmark for explainable artificial intelligence methods. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/2/85.

Dhesi:2025:DEV

[3338] Simran Singh Dhesi, Pratik Adusumilli, Nishant Ravikumar, Mohammed A. Waduud, Russell Frood, Alejandro F. Frangi, Garry Mc-Dermott, James H. F. Rudd, Yuan Huang, Jonathan R. Boyle, Maysoon Elkhawad, David E. Newby, Nikhil Joshi, Jing Yi Kwan, Patrick Coughlin, Marc A. Bailey, and Andrew F. Scarsbrook. Development and external validation of [18F]FDG PET-CT-derived radiomic models for prediction of abdominal aortic aneurysm growth rate. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/86.

Moore:2025:BBO

[3339] Alex Moore and Jie Zhang. Batch-to-batch optimization control of fedbatch fermentation process based on recursively updated extreme learning machine models. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/87.

Aghamiri:2025:NPD

[3340] Hamid Reza Aghamiri, S. Abolfazl Hosseini, James R. Green, and B. John Oommen. Nonparametric probability density function estimation using the Padé approximation. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/88.

Tehsin:2025:GGA

[3341] Sara Tehsin, Inzamam Mashood Nasir, and Robertas Damasevicius. GATransformer: a graph attention network-based transformer model to generate explainable attentions for brain tumor detection. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/89.

Cui:2025:ESC

[3342] Hongzhen Cui, Shenhui Ning, Shichao Wang, Wei Zhang, and Yunfeng Peng. ECG signal classification using interpretable KAN: Towards predictive diagnosis of arrhythmias. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/90.

Hong:2025:ADV

[3343] Chih-Chiang Hong. Advanced dynamic vibration of Terfenol-D control law on functionally graded material plates/cylindrical shells in unsteady supersonic flow. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/91.

Liu:2025:OMT

[3344] Zhiwei Liu, Sihui Zhou, and Song Liu. Optimization of multimodal transport paths considering a low-carbon economy under uncertain demand. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
2/92.

Viktoratos:2025:ARE

[3345] Iosif Viktoratos and Athanasios Tsadiras. Advancing real-estate forecasting: a novel approach using Kolmogorov–Arnold networks. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/2/93.

Awad:2025:EBM

[3346] Sohaib R. Awad and Faris S. Alghareb. Encoding-based machine learning approach for health status classification and remote monitoring of cardiac patients. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/94.

Alghlayini:2025:ENI

[3347] Saifeddin Alghlayini, Mohammed Azmi Al-Betar, and Mohamed Atef. Enhancing non-invasive blood glucose prediction from photoplethysmography signals via heart rate variability-based features selection using metaheuristic algorithms. Algorithms (Basel), 18(2), February 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/2/95.

Matsuzaka:2025:DCP

[3348] Yasunari Matsuzaka and Ryu Yashiro. The diagnostic classification of the pathological image using computer vision. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/2/96.

Manzoor:2025:PBS

[3349] Shahid Manzoor, Noor Shamsiah Othman, and Mohammed W. Muhieldeen. Preamble-based signal-to-noise ratio estimation for adaptive modulation in space-time block coding-assisted multiple-input multiple-output orthogonal frequency division multiplexing system. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/97.

Radocaj:2025:PPR

[3350] Petra Radocaj, Dorijan Radocaj, and Goran Martinović. Pediatric pneumonia recognition using an improved DenseNet201 model with multiscale convolutions and mish activation function. Algorithms (Basel), 18 (2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/98.

Bano:2025:LMW

[3351] Farheen Bano, Faisal Ahmad, Mohammad Shahid, Mahfooz Alam, Faraz Hasan, and Mohammad Sajid. A levelized multiple workflow heterogeneous earliest finish time allocation model for Infrastructure as a Service (IaaS) cloud environment. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/99.

Samaroo:2025:PIA

[3352] Kyle Samaroo, Abdul Waheed Awan, Siva Marimuthu, Muhammad Naveed Iqbal, Kamran Daniel, and Noman Shabbir. Performance investigation of active, semi-active and passive suspension using quarter car model. Algorithms (Basel), 18(2), February 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/100.

Wang:2025:LDL

[3353] Yizhao Wang, Ziye Guo, Haitao Luo, Jing Liu, and Ruohua Zhou. A lightweight deep learning approach for detecting external intrusion signals from optical fiber sensing system based on temporal efficient residual network. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/101.

Pardini:2025:IAD

[3354] Gabriel R. Pardini, Paulo M. Tasinaffo, Elcio H. Shiguemori, Tahisa N. Kuck, Marcos R. O. A. Maximo, and William R. Gyotoku. Improved algorithm to detect clandestine airstrips in Amazon RainForest. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/2/102.

Lopera:2025:BCC

[3355] Juan Camilo Londoño Lopera, Freddy Bolaños Martinez, and Luis Alejandro Fletscher Bocanegra. Building a custom crime detection dataset and implementing a 3D convolutional neural network for video analysis. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/2/
103.

Carcangiu:2025:TAL

[3356] Sara Carcangiu and Augusto Montisci. A training algorithm for locally recurrent neural networks based on the explicit gradient of the loss function. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/104.

Nanni:2025:ATM

[3357] Loris Nanni, Matteo De Gobbi, Roger DE Almeida Matos Junior, and Daniel Fusaro. Advancing taxonomy with machine learning: a hybrid ensemble for species and genus classification. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/2/105.

Penelas:2025:MLD

[3358] Gonçalo Penelas, Luís Barbosa, Arsénio Reis, João Barroso, and Tiago Pinto. Machine learning for decision support and automation in games: a study on vehicle optimal path. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/106.

Deng:2025:MSI

[3359] Le Deng, Ting Shu, and Jinsong Xia. Multi-strategy improved artificial rabbit algorithm for QoS-aware service composition in cloud manufacturing. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/2/107.

Perianez-Pascual:2025:BSR

[3360] Jorge Perianez-Pascual, Juan D. Gutiérrez, Laura Escobar-Encinas, Álvaro Rubio-Largo, and Roberto Rodriguez-Echeverria. Beyond spectrograms: Rethinking audio classification from EnCodec’s latent space.

Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
2/108.

Zarzycki:2025:OIP

[3361] Hubert Zarzycki. Optimizing investment portfolios with bacterial foraging and robust risk management. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/2/109.

Jewpanya:2025:OTI

[3362] Parida Jewpanya, Pinit Nuangpirom, Siwasit Pitjamit, and Warisa Nakkiew. Optimized travel itineraries: Combining mandatory visits and personalized activities. Algorithms (Basel), 18(2), February 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/2/110.

Castellanos-Buitrago:2025:EEM

[3363] Sebastián F. Castellanos-Buitrago, Pablo Maya-Duque, Walter M. Villa-Acevedo, Nicolás Muñoz-Galeano, and Jesús M. López-Lezama. Enhancing energy microgrid sizing: a multiyear optimization approach with uncertainty considerations for optimal design. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/2/111.

Kobayashi:2025:EMV

[3364] Yasuaki Kobayashi, Kazuhiro Kurita, Kevin Mann, Yasuko Matsui, and Hirotaka Ono. Enumerating minimal vertex covers and dominating sets with capacity and/or connectivity constraints. Algorithms (Basel), 18 (2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/112.

Lazzarinetti:2025:DDL

[3365] Giorgio Lazzarinetti, Riccardo Dondi, Sara Manzoni, and Italo Zoppis. DLMinTC+: a deep learning based algorithm for minimum timeline cover on temporal graphs. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/113.

Awdaa:2025:EMH

[3366] Mohammed Ajel Awdaa, Elaheh Mashhour, Hossein Farzin, and Mahmood Joorabian. Energy management and hosting capacity evaluation of a hybrid AC-DC micro grid including photovoltaic units and battery energy storage systems. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/114.

Dosa:2025:SPI

[3367] György Dósa, András Éles, Angshuman Robin Goswami, István Szalkai, and Zsolt Tuza. Solution of bin packing instances in Falkenauer T class: Not so hard. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/2/115.

Wang:2025:KDP

[3368] Xiao-Song Wang, Anoop Kumar Maurya, Muhammad Ishtiaq, Sung-Gyu Kang, and Nagireddy Gari Subba Reddy. Knowledge discovery in predicting martensite start temperature of medium-carbon steels by artificial neural networks. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/2/116.

Gad:2025:TIT

[3369] Ibrahim Gad. TOCA-IoT: Threshold optimization and causal analysis for IoT network anomaly detection based on explainable random forest. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
2/117.

Sittijuk:2025:RCS

[3370] Pramote Sittijuk, Narin Petrot, and Kreangsak Tamee. Robust client selection strategy using an improved federated random high local performance algorithm to address high non-IID challenges. Algorithms (Basel), 18(2), February 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/2/118.

Eid:2025:UOA

[3371] Mahmoud M. Eid, Kamal ElDahshan, Abdelatif H. Abouali, and Alaa Tharwat. Using optimization algorithms for effective missing-data imputation: a case study of tabular data derived from video surveillance. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/119.

Bologna:2025:FFL

[3372] Guido Bologna, Jean-Marc Boutay, Damian Boquete, Quentin Leblanc, Deniz Köprülü, and Ludovic Pfeiffer. Fidex and FidexGlo: From local explanations to global explanations of deep models. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/120.

Jiao:2025:DDS

[3373] Hao Jiao, Chen Wu, Lei Wei, Jinming Chen, Yang Xu, and Manyun Huang. A data-driven state estimation based on sample migration for low-observable distribution networks. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/3/121.

Herendi:2025:AAC

[3374] Tamás Herendi and Viktória Padányi. Automata and arithmetics in canonical number systems. Algorithms (Basel), 18(3), March 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/3/122.

Martinenghi:2025:SIC

[3375] Davide Martinenghi. Simplified integrity checking for an expressive class of denial constraints. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/123.

Tankovic:2025:TMD

[3376] Nikola Tanković, Robert Sajina, and Ivan Lorencin. Transforming medical data access: The role and challenges of recent language models in SQL query automation. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/124.

Xu:2025:SSA

[3377] Shaowen Xu and Der-Horng Lee. Spatial shape-aware network for elongated target detection. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/125.

Feesantia:2025:IDS

[3378] Thipagon Feesantia, Tipsuda Arunrat, and Nimit Nimana. Incremental delayed subgradient method for decentralized nonsmooth convex-concave minimax optimization. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/126.

Aguilar-Gonzalez:2025:RED

[3379] Abiel Aguilar-González and Alejandro Medina Santiago. Road event detection and classification algorithm using vibration and acceleration data. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/
127.

Toronov:2025:UCH

[3380] Vladislav Toronov, Nima Soltani, Leeanne Leung, Rohit Mohindra, and Steve Lin. Using coherent hemodynamic spectroscopy model to investigate cardiac arrest. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/128.

Konishi:2025:MIA

[3381] Tomokazu Konishi. Means and issues for adjusting principal component analysis results. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/129.

Ding:2025:IPD

[3382] Yucheng Ding, Yingfeng Zhang, Jianfeng Huang, and Shitong Peng. Incremental pyraformer-deep canonical correlation analysis: a novel framework for effective fault detection in dynamic nonlinear processes. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/130.

Wang:2025:FGM

[3383] Yahui Wang and Yaqing Liu. Fine-grained mapping between daily activity features in smart homes. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/3/131.

Moghimi:2025:AML

[3384] Seyed Morteza Moghimi, Thomas Aaron Gulliver, Ilamparithi Thirumarai Chelvan, and Hossen Teimoorinia. Adaptive machine learning for automatic load optimization in connected smart green townhouses. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/132.

Zhang:2025:LED

[3385] Biao Zhang, Chishe Wang, and Jie Wang. A lightweight and efficient detection transformer for highway abandoned objects. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/133.

Marcos:2025:EDA

[3386] Luella Marcos, Paul Babyn, and Javad Alirezaie. Edge detection attention module in pure vision transformer for low-dose X-ray computed tomography image denoising. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/3/134.

Carpentieri:2025:DCT

[3387] Bruno Carpentieri. Data compression with a time limit. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/3/135.

Nassih:2025:BCC

[3388] Rym Nassih and Abdelaziz Berrado. Breast cancer classification using an adapted bump-hunting algorithm. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/3/136.

Gomez:2025:CIT

[3389] Pablo García Gómez, Inés González-Rodríguez, and Camino R. Vela. Computing idle times in fuzzy flexible job shop scheduling. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/3/137.

Lei:2025:LLA

[3390] Yi Lei, Xin Wen, Yanrong Hao, Ruochen Cao, Chengxin Gao, Peng Wang, Yuanyuan Guo, and Rui Cao. LSTGINet: Local attention spatiotemporal graph inference network for age prediction. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/138.

Sutton:2025:AMG

[3391] Nate M. Sutton, Blanca E. Gutiérrez-Guzmán, Holger Dannenberg, and Giorgio A. Ascoli. Automated measurement of grid cell firing characteristics. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/
139.

Schulhof:2025:FMO

[3392] Noah Schulhof, Pattara Sukprasert, Eytan Ruppin, Samir Khuller, and Alejandro A. Schäffer. Finding multiple optimal solutions to an integer linear program by random perturbations of its objective function. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/140.

DeLorenzis:2025:PSP

[3393] Emilio De Lorenzis and Davide Martinenghi. Partitioning strategies for parallel computation of flexible skylines. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/3/141.

Dong:2025:UFR

[3394] Hao Dong, Peng Lu, Xiang Jiang, Xu Chen, Puyan Wang, and Junpeng Zhu. Uncertainty feasible region analysis for microgrids in the coordination with distribution system considering interactive power deviation. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/142.

Zhang:2025:HDE

[3395] Yue Zhang, Nobuo Funabiki, Erita Cicilia Febrianti, Amang Sudarsono, and Chenchien Hsu. A hair drawing evaluation algorithm for exactness assessment method in portrait drawing learning assistant system. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/143.

Nikolovski:2025:AAH

[3396] Vlatko Nikolovski, Dimitar Trajanov, and Ivan Chorbev. Advancing AI in higher education: a comparative study of large language modelbased agents for exam question generation, improvement, and evaluation. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/144.

Almutairi:2025:PDU

[3397] Entissar Almutairi, Maysam Abbod, and Ziad Hunaiti. Prediction of diabetes using statistical and machine learning modelling techniques. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/145.

Hua:2025:DPP

[3398] Xia Hua, Tengteng Zhang, and Jun Cao. Dynamic path planning for vehicles based on causal state-masking deep reinforcement learning. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/146.

Zuo:2025:PCB

[3399] Jiwei Zuo, Junfu Fan, Kuan Li, Qingyun Liu, Yuke Zhou, and Yi Zhang. Parallel CUDA-based optimization of the intersection calculation process in the Greiner–Hormann algorithm. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/3/147.

Wang:2025:CIS

[3400] Ruixue Wang and Ning Zhao. CEEMDAN-IHO-SVM: a machine learning research model for valve leak diagnosis. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/3/148.

Bialecki:2025:SRD

[3401] Mariusz Białecki. Simple rules of a discrete stochastic process leading to Catalan-like recurrences. Algorithms (Basel), 18(3), March 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/3/149.

Rozinek:2025:RTF

[3402] Ondrej Rozinek, Jaroslav Marek, Jan Panus, and Jan Mares. Real-time fuzzy record-matching similarity metric and optimal Q-gram filter. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/150.

Siyamayambo:2025:DMS

[3403] Claris Siyamayambo, Edith Phalane, and Refilwe Nancy Phaswana-Mafuya. Diagnosis and management of sexually transmitted infections using artificial intelligence applications among key and general populations in Sub-Saharan Africa: a systematic review and meta-analysis. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/151.

Shen:2025:NES

[3404] Kedi Shen, Jieyu Zhao, and Min Xie. A novel equivariant self-supervised vector network for three-dimensional point clouds. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/152.

Tasinaffo:2025:ETU

[3405] Paulo M. Tasinaffo, Gildárcio S. Gonçalves, Johnny C. Marques, Luiz A. V. Dias, and Adilson M. da Cunha. The Euler-type universal numerical integrator (E-TUNI) with backward integration. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/153.

Pham:2025:AFB

[3406] Tin H. Pham and Bijan Raahemi. An adaptive feature-based quantum genetic algorithm for dimension reduction with applications in outlier detection. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/3/154.

Peter:2025:TGS

[3407] Ojonugwa Oluwafemi Ejiga Peter, Opeyemi Taiwo Adeniran, Adetokunbo Mac-Gregor John-Otumu, Fahmi Khalifa, and Md Mahmudur Rahman. Textguided synthesis in medical multimedia retrieval: a framework for enhanced colonoscopy image classification and segmentation. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/3/155.

Chow:2025:QCM

[3408] James C. L. Chow. Quantum computing and machine learning in medical decision-making: a comprehensive review. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/3/156.

Brohi:2025:AUA

[3409] Sarfraz Brohi and Qurat ul-ain Mastoi. AI under attack: Metric-driven analysis of cybersecurity threats in deep learning models for healthcare applications. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/3/157.

Han:2025:GOU

[3410] Xu Han and Dian Wang. Gradual optimization of university course scheduling problem using genetic algorithm and dynamic programming. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/158.

Bosch:2025:TOI

[3411] Paul Bosch, Ernesto Parra Inza, Ismael Rios Villamar, and José Luis Sánchez-Santiesteban. Total outer-independent domination number: Bounds and algorithms. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/159.

AlHwaitat:2025:HOA

[3412] Ahmad K. Al Hwaitat, Hussam N. Fakhouri, Jamal Zraqou, and Najem Sirhan. Hybrid optimization algorithm for solving attack-response optimization and engineering design problems. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/3/160.

Elbayed:2025:DLF

[3413] Zakaria Elbayed and Abdelmjid Qadi EI Idrissi. Deep learning in financial modeling: Predicting European put option prices with neural networks. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/3/161.

Radac:2025:CHP

[3414] Mircea-Bogdan Radac, Valentin-Dan Muller, and Samuel Ciucurita. Control of high-power slip ring induction generator wind turbines at variable wind speeds in optimal and reliable modes. Algorithms (Basel), 18 (3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/3/162.

Adeniran:2025:EMB

[3415] Opeyemi Taiwo Adeniran, Blessing Ojeme, Temitope Ezekiel Ajibola, Ojonugwa Oluwafemi Ejiga Peter, Abiola Olayinka Ajala, Md Mahmudur Rahman, and Fahmi Khalifa. Explainable MRI-based ensemble learnable architecture for Alzheimer’s disease detection. Algorithms (Basel), 18 (3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/3/163.

Liu:2025:TSM

[3416] Jiaqi Liu and Tianyu Liu. A two-stage multi-objective optimization algorithm for solving large-scale optimization problems. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/164.

Hvattum:2025:WSH

[3417] Lars Magnus Hvattum. Where to split in hybrid genetic search for the capacitated vehicle routing problem. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/3/165.

Qiu:2025:GBF

[3418] Junhong Qiu, Xinxin Zhang, Tao Wang, Huiying Hou, Siyuan Wang, and Tiejun Yang. A GNN-based false data detection scheme for smart grids. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/3/
166.

Chen:2025:DPP

[3419] Minru Chen, Binglin Liu, Mingzhi Liang, and Nini Yao. Decoding PM_{2.5} prediction in Nanning Urban Area, China: Unraveling model superiorities and drawbacks through SARIMA, Prophet, and LightGBM. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/167.

Lin:2025:CSG

[3420] Jhe-Wei Lin, Tang-Wei Su, and Che-Cheng Chang. Chinese story generation based on style control of transformer model and content evaluation method. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/3/168.

Farghali:2025:AEP

[3421] Rasha A. Farghali, Adewale F. Lukman, Zakariya Algamal, Murat Genc, and Hend Attia. An alternative estimator for Poisson-inverse-Gaussian regression: The modified Kibria-Lukman estimator. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/169.

Tudoroiu:2025:IDI

[3422] Roxana-Elena Tudoroiu, Mohammed Zaheeruddin, Daniel-Ioan Curiac, Mihai Sorin Radu, and Nicolae Tudoroiu. Investigations into the design and implementation of reinforcement learning using deep learning neural networks. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/3/170.

Dang:2025:GSA

[3423] Tuan Phuong Dang, Manh Tuan Nguyen, Adnan F. Alhaj Hasan, and Talgat R. Gazizov. Generation of sparse antennas and scatterers based on optimal current grid approximation. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/3/171.

Shi:2025:MMM

[3424] Peicheng Shi, Wenchao Wu, and Aixi Yang. MPVF: Multi-modal 3D object detection algorithm with pointwise and voxelwise fusion. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/3/172.

Faria:2025:DSM

[3425] Pedro Faria, Victor Alves, José Neves, and Henrique Vicente. Data science in the management of healthcare organizations. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/3/173.

Liu:2025:AFO

[3426] Binglin Liu, Weijia Zeng, Weijiang Liu, Yi Peng, and Nini Yao. Algorithms facilitating the observation of urban residential vacancy rates: Technologies, challenges and breakthroughs. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/3/174.

Arabi:2025:PBA

[3427] Marzieh Sadat Arabi and Anjali Awasthi. A PSO-based approach for the optimal allocation of electric vehicle parking lots to the electricity distribution network. Algorithms (Basel), 18(3), March 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/3/175.

Chaudhary:2025:SMP

[3428] Manu Chaudhary, Kareem El-Araby, Alvir Nobel, Vinayak Jha, Dylan Kneidel, Ishraq Islam, Manish Singh, Sunday Ogundele, Ben Phillips, Kieran Egan, Sneha Thomas, Devon Bontrager, Serom Kim, and Esam El-Araby. Solving multidimensional partial differential equations using efficient quantum circuits. Algorithms (Basel), 18(3), March 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/3/176.

Abdelazeem:2025:ULC

[3429] Mohamed Abdelazeem, Amgad Abazeed, Hussain A. Kamal, and Mudathir O. A. Mohamed. Ultra-low-cost real-time precise point positioning using different streams for precise positioning and precipitable water vapor retrieval estimates. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/198.

Alrasheed:2025:PIS

[3430] Sadeem Alrasheed, Suliman Aladhadh, and Abdulatif Alabdulatif. Protecting intellectual security through hate speech detection using an artificial intelligence approach. Algorithms (Basel), 18(4), April 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/4/179.

Alviano:2025:UAE

[3431] Mario Alviano and Pierpaolo Sestito. User armor: An extension for AppArmor. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/4/185.

Amannah:2025:DDI

[3432] Constance Amannah, Kingsley Friday Attai, and Faith-Michael Uzoka. A data-driven intelligent methodology for developing explainable diagnostic model for febrile diseases. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/190.

Angelina:2025:SAA

[3433] Q Angelina, Khadija Begum, Hee-Cheol Kim, Sushanta Tripathy, Deepak Singhal, and Saranjit Singh. A structural analysis of AI implementation challenges in healthcare. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/189.

Bai:2025:OSM

[3434] Wei Bai, Chengxin Fu, Bin Zhao, Gen Li, and Zhihong Yao. An optimal scheduling model for connected automated vehicles at an unsignalized intersection. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/4/194.

Banko:2025:AML

[3435] Márton Bendegúz Bankó, Szymon Dyszewski, Michaela Králová, Márton Bertalan Limpek, Maria Papaioannou, Gaurav Choudhary, and Nicola Dragoni. Advancements in machine learning-based intrusion detection in IoT: Research trends and challenges. Algorithms (Basel), 18 (4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/209.

Barbre:2025:EWE

[3436] Zachary Barbre and Gang Li. Enhanced wind energy forecasting using an extended long short-term memory model. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/206.

Bhattacharya:2025:IDR

[3437] Sayantan Bhattacharya, Dimitris M. Christodoulou, and Silas G. T. Laycock. Image deconvolution to resolve astronomical X-ray sources in close proximity: The NuSTAR images of SXP 15.3 and SXP 305. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/4/191.

Cafaro:2025:NGB

[3438] Rosa Cafaro, Barbara Cardone, Valeria D’Ambrosio, Ferdinando Di Martino, and Vittorio Miraglia. A new GIS-based detection technique for urban heat islands using the fuzzy C-means clustering algorithm: a case study of Naples, (Italy). Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/228.

Chen:2025:PPA

[3439] Wei Chen, Qiuyue Liu, Hancong Li, and Jian Zou. The prediction performance analysis of the lasso model with convex non-convex sparse regularization. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/4/195.

Gallardo:2025:PSM

[3440] Claudia Sifuentes Gallardo, Misael Zambrano de la Torre, Daniel Alaniz Lumbreras, Efren Gonzalez-Ramirez, José Ismael De la Rosa Vargas, Carlos Olvera-Olvera, José Ortega Sigala, Omar Alejandro Guirette-Barbosa, Oscar Cruz Domínguez, and Héctor Durán Muñoz. A preliminary study on machine learning techniques to classify cardiovascular diseases in Mexico. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/202.

Garcia-Ruiz:2025:MDA

[3441] Álvaro García-Ruiz, Pablo Fernández-Arias, and Diego Vergara. Mathematics-driven analysis of offshore green hydrogen stations. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/237.

Gonzalez-Prida:2025:DTA

[3442] Vicente González-Prida, Carlos Parra Márquez, Pablo Viveros Gunckel, Fredy Kristjanpoller Rodríguez, and Adolfo Crespo Márquez. Digital transformation in aftersales and warranty management: a review of advanced technologies in I4.0. Algorithms (Basel), 18(4), April 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/4/231.

Guo:2025:DLB

[3443] Lei Guo, Xin Guo, and Feiya Lv. Distributed load-balancing method for CCA parallel component applications. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/4/211.

Gutierrez:2025:POS

[3444] Juan D. Gutiérrez, Emilio Delgado, Carlos Breuer, José M. Conejero, and Roberto Rodriguez-Echeverria. Prompt once, segment everything: Leveraging SAM 2 potential for infinite medical image segmentation with a single prompt. Algorithms (Basel), 18(4), April 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/227.

Gutman:2025:HDR

[3445] David Gutman, Nir Perel, Oana Barbulescu, and Oded Koren. A hybrid dimensionality reduction procedure integrating clustering with KNNbased feature selection for unsupervised data. Algorithms (Basel), 18 (4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/188.

Han:2025:PSH

[3446] Gawon Han, Keith Wachowicz, Nawaid Usmani, Don Yee, Jordan Wong, Arun Elangovan, Jihyun Yun, and B. Gino Fallone. Patient-specific hyperparameter optimization of a deep learning-based tumor autocontouring algorithm on 2D liver, prostate, and lung cine MR images: a pilot study. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/
233.

Huang:2025:UCT

[3447] Yi-Sheng Huang, Yi-Shun Weng, and Chun-Yu Shih. Utilizing cell transmission models to alleviate accident-induced traffic congestion in twoway grid networks. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/193.

Ilyushin:2025:AIP

[3448] Pavel Ilyushin, Bulat Gaisin, Ildar Shahmaev, and Konstantin Suslov. An algorithm for identifying the possibilities of cascading failure processes and their development trajectories in electric power systems. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/183.

Islam:2025:QCS

[3449] Ishraq Islam, Vinayak Jha, Sneha Thomas, Kieran F. Egan, Alvir Nobel, Serom Kim, Manu Chaudhary, Sunday Ogundele, Dylan Kneidel, Ben Phillips, Manish Singh, Kareem El-Araby, Devon Bontrager, and Esam El-Araby. Quantum circuit synthesis using fuzzy-logic-assisted genetic algorithms. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/4/178.

Jansson:2025:ESA

[3450] Jesper Jansson. Editorial: Surveys in algorithm analysis and complexity theory, part II (special issue). Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/4/212.

Khan:2025:TEP

[3451] Umar Khan, Qaiser Riaz, Mehdi Hussain, Muhammad Zeeshan, and Björn Krüger. Towards effective Parkinson’s monitoring: Movement disorder detection and symptom identification using wearable inertial sensors. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/

Khiat:2025:EED

[3452] Salim Khiat, Sidi Ahmed Mahmoudi, Sédrick Stassin, Lillia Boukerroui, Besma Senaï, and Saïd Mahmoudi. An efficient explainability of deep models on medical images. Algorithms (Basel), 18(4), April 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/4/210.

Kisel:2025:DLB

[3453] Ivan Kisel, Robin Lakos, and Gianna Zischka. Deep-learning-based optimization of the signal/background ratio for Λ particles in the CBM experiment at FAIR. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/229.

Krishnamurthy:2025:SRP

[3454] Bhargavi Krishnamurthy and Sajjan G. Shiva. Scalable resource provisioning framework for fog computing using LLM-guided Q-learning approach. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/
230.

Li:2025:CAM

[3455] Haitao Li, Chenyu Wang, Shufu Yuan, Hui Zhu, and Li Sun. A capacity allocation method for long-endurance hydrogen-powered hybrid UAVs based on two-stage optimization. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/4/196.

Ludwig:2025:GPS

[3456] Simone A. Ludwig. Guided particle swarm optimization for feature selection: Application to cancer genome data. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/220.

Lv:2025:MFF

[3457] Zekai Lv, Yufeng Guo, Shangbin Yang, Linlin Du, Rui Gao, Jinti Sun, Jiaqi Han, Hui Zhang, and Qiang Wang. Multiscale feature fusion with self-attention for efficient 6D pose estimation. Algorithms (Basel), 18 (4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/207.

Marcheva:2025:QCD

[3458] Plamena I. Marcheva, Ivan K. Ivanov, and Stoil I. Ivanov. On the Qconvergence and dynamics of a modified Weierstrass method for the simultaneous extraction of polynomial zeros. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/205.

Martinez:2025:IER

[3459] Mariano Romero Martínez, Pedro Carmona Ibáñez, and José Pozuelo Campillo. The impact of environmental risk on business failure: a fuzzy-set qualitative comparative analysis approach with extreme gradient boosting feature selection. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/4/225.

Moulik:2025:OEC

[3460] Bedatri Moulik, Sanmukh Kaur, and Muhammad Ijaz. Optimized energy consumption of electric vehicles with driving pattern recognition for real driving scenarios. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/204.

Muller:2025:CAI

[3461] Felipe Martins Müller, Vanessa Andréia Schneider, Olinto Cesar Bassi de Araujo, Claudio Roberto Scheer Júnior, and Guilherme Lopes Weis. Comparing an artificial intelligence planner with traditional optimization methods: a case study in the dairy industry. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/219.

Munaye:2025:HDL

[3462] Yirga Yayeh Munaye, Metadel Addis, Yenework Belayneh, Atinkut Molla, and Wasyihun Admass. Hybrid deep learning methods for human activity recognition and localization in outdoor environments. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/235.

Myllis:2025:HPC

[3463] Georgios Myllis, Alkiviadis Tsimpiris, Stamatios Aggelopoulos, and Vasiliki G. Vrana. High-performance computing and parallel algorithms for urban water demand forecasting. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/4/182.

Nigro:2025:PPD

[3464] Libero Nigro and Franco Cicirelli. Proving properties of Dekker’s algorithm for mutual exclusion of N processes. Algorithms (Basel), 18 (4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/226.

Nurmaini:2025:IEE

[3465] Siti Nurmaini, Wisnu Jatmiko, Satria Mandala, Bambang Tutuko, Erwin Erwin, Alexander Edo Tondas, Annisa Darmawahyuni, Firdaus Firdaus, Muhammad Naufal Rachmatullah, Ade Iriani Sapitri, Anggun Islami, Akhiar Wista Arum, and Muhammad Ikhwan Perwira. iCOR: End-to-end electrocardiography morphology classification combining multi-layer filter and BiLSTM. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/236.

Panagiotopoulos:2025:CMR

[3466] Fotios Panagiotopoulos and Vassilios Chatzis. A continuous monitoring and reallocation method for successful decisions in change management. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/184.

Pandey:2025:IVF

[3467] Abhishek Kumar Pandey, Kedarnath Senapati, Ioannis K. Argyros, and G. P. Pateel. Improving vertebral fracture detection in C-spine CT images using Bayesian probability-based ensemble learning. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/4/181.

Perez-Quezadas:2025:PMB

[3468] Nora Isabel Pérez-Quezadas, Héctor Benítez-Pérez, and Adrián Durán-Chavesti. Proposal of a methodology based on using a wavelet transform as a convolution operation in a convolutional neural network for feature extraction purposes. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/221.

Pliakos:2025:AFS

[3469] Chris Pliakos, Giorgos Efrem, Dimitrios Terzis, and Pericles Panagiotou. An automated framework for streamlined CFD-based design and optimization of fixed-wing UAV wings. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/4/186.

Poturic:2025:SAE

[3470] Vanja Cotić Poturić, Sanja Candrlić, and Ivan Drazić. A scoring algorithm for the early prediction of academic risk in STEM courses. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/177.

Rahmani:2025:ICS

[3471] Mohammed Rahmani, Abderrahmane Nitaj, and Mhammed Ziane. Improved cryptanalysis of some RSA variants. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/223.

Ramos-Martinez:2025:IIC

[3472] Moises Ramos-Martinez, Jorge A. Brizuela-Mendoza, Carlos A. Torres-Cantero, Gerardo Ortiz-Torres, Felipe D. J. Sorcia-Vázquez, Mario A. Juarez, Jair de Jesús Cambrón Navarrete, Juan Carlos Mixteco-Sánchez, Mayra G. Mena-Enriquez, Rafael Murrieta Yescas, and Jesse Y. Rumbo-Morales. Implementation of an intelligent controller based on neural networks for the simulation of pressure swing adsorption systems. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/215.

Robitzsch:2025:CAL

[3473] Alexander Robitzsch. Computational aspects of L_{0} linking in the Rasch model. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/
213.

Roman:2025:IFS

[3474] Ali Roman, Md Mostafizer Rahman, Sajjad Ali Haider, Tallha Akram, and Syed Rameez Naqvi. Integrating feature selection and deep learning: a hybrid approach for smart agriculture applications. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/222.

Sakovich:2025:WBO

[3475] Nikita Sakovich, Dmitry Aksenov, Ekaterina Pleshakova, and Sergey Gataullin. Wavelet-based optimization and numerical computing for fault detection method-signal fault localization and classification algorithm. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/

Saleem:2025:ARL

[3476] Omer Saleem, Aliha Tanveer, and Jamshed Iqbal. Adaptive reconfigurable learning algorithm for robust optimal longitudinal motion control of unmanned aerial vehicles. Algorithms (Basel), 18(4), April 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/4/180.

Sanchez:2025:GSM

[3477] Oscar D. Sanchez, Luz M. Reyes, Arturo Valdivia-González, Alma Y. Alanis, and Eduardo Rangel-Heras. Gaslike social motility: Optimization algorithm with application in image thresholding segmentation. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/199.

Shahabikargar:2025:CGA

[3478] Maryam Shahabikargar, Amin Beheshti, Wathiq Mansoor, Xuyun Zhang, Eu Jin Foo, Alireza Jolfaei, Ambreen Hanif, and Nasrin Shabani. ChurnKB: a generative AI-enriched knowledge base for customer churn feature engineering. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/238.

Song:2025:EVR

[3479] Bowen Song and Rui Xu. Electric vehicle routing problem with heterogeneous energy replenishment infrastructures under capacity constraints. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/216.

Srivastava:2025:KTT

[3480] Hari Mohan Srivastava, Bidu Bhusan Jena, Susanta Kumar Paikray, and Umakanta Misra. Korovkin-type theorems for positive linear operators based on the statistical derivative of deferred Cesàro summability. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/218.

Tsoulos:2025:UBP

[3481] Ioannis G. Tsoulos, Vasileios Charilogis, and Dimitrios Tsalikakis. Utilizing a bounding procedure based on simulated annealing to effectively locate the bounds for the parameters of radial basis function networks. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/234.

Wang:2025:ESD

[3482] Bo Wang, Xinyu Sun, and Ying Song. Edge server deployment strategy based on queueing search meta-heuristic algorithm. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/200.

Wang:2025:PSR

[3483] Pengzhao Wang, Xueqin Tian, Kai Sun, Yi Huang, Zhidong Wang, and Li Sun. Power system reliability assessment considering coal-fired unit peaking characteristics. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/197.

Wang:2025:RGC

[3484] Yujie Wang, Yang Cao, Zhong Qian, Jianping Xia, Xuhong Kang, Yixian Zhu, Yanan Yang, Wendong Zhang, Shaohua Chen, and Guoqing Wu. Research on grid-connected speed control of hydraulic wind turbine based on enhanced chaotic particle swarm optimization fuzzy PID. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/4/187.

Wani:2025:CSM

[3485] Shahid Ahmad Wani, SMK Quadri, Mohammad Shuaib Mir, and Yonis Gulzar. A comparative study of machine learning techniques for cell annotation of scRNA-Seq data. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/4/232.

Yin:2025:NOL

[3486] Xijian Yin and Yanjun Liu. A new orthogonal least squares identification method for a class of fractional Hammerstein models. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/4/201.

Yu:2025:IRA

[3487] Yifan Yu, Shuxi Chen, Depeng Gao, and Jianlin Qiu. An intelligent reliability assessment and prognosis of rolling bearings using adaptive cyclostationary blind deconvolution and AdaBoost-mixed kernel relevance vector machine. Algorithms (Basel), 18(4), April 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/4/192.

Yuan:2025:OVE

[3488] Meng Yuan and Hongjun Li. Optimal view estimation algorithm and evaluation with deviation angle analysis. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/4/224.

Zattoni:2025:GMC

[3489] Elena Zattoni, Anna Maria Perdon, and Giuseppe Conte. Geometric methods and computational tools for switching structured linear systems. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/4/208.

Zhao:2025:PCP

[3490] Peng Zhao, Xiaoyu Liu, Xuqi Su, Di Wu, Zi Li, Kai Kang, Keqin Li, and Armando Zhu. Probabilistic contingent planning based on hierarchical task network for high-quality plans. Algorithms (Basel), 18(4), April 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/4/214.

Abdikerimova:2025:DMS

[3491] Gulzira Abdikerimova, Dana Khamitova, Akmaral Kassymova, Assyl Bissengaliyeva, Gulsara Nurova, Murat Aitimov, Yerlan Alimzhanovich Shynbergenov, Moldir Yessenova, and Roza Bekbayeva. Development of a model for soil salinity segmentation based on remote sensing data and climate parameters. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/285.

Abolhassanifar:2025:EML

[3492] Mohammad Saeid Abolhassanifar, Reza Saadati, Mohammad Bagher Ghaemi, and Donal O’Regan. Existence and Mittag-Leffler stability for the solution of a fuzzy fractional system with application of Laplace transforms to solve fractional differential systems. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/264.

Amat:2025:ABR

[3493] Sergio Amat, Sonia Busquier, Carlos D. Gómez-Carmona, Manuel Gómez-López, and José Pino-Ortega. Algorithm-based real-time analysis of training phases in competitive canoeing: an automated approach for performance monitoring. Algorithms (Basel), 18(5), May 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/5/242.

Bravo-Lopez:2025:CCF

[3494] Esteban Bravo-López, Tomás Fernández, Chester Sellers, and Jorge Delgado-García. Combination of conditioning factors for generation of landslide susceptibility maps by extreme gradient boosting in cuenca, Ecuador. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/5/258.

Cuchipe:2025:ASR

[3495] Wilmer Clemente Cunuhay Cuchipe, Johnny Bajaña Zajia, Byron Oviedo, and Cristian Zambrano-Vega. Advanced sales route optimization through enhanced genetic algorithms and real-time navigation systems. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/260.

Cuzzocrea:2025:DAB

[3496] Alfredo Cuzzocrea, Abderraouf Hafsaoui, and Carmine Gallo. Detecting and analyzing botnet nodes via advanced graph representation learning tools. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/
253.

Dabek:2025:AND

[3497] Aleksander Dabek, Lorenzo Mantovani, Susanna Mirabella, Michele Vignati, and Simone Cinquemani. Advancements in non-destructive detection of biochemical traits in plants through spectral imaging-based algorithms: a systematic review. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/255.

Dimov:2025:CES

[3498] Ivan Dimov and Rayna Georgieva. Cutting-edge stochastic approach: Efficient Monte Carlo algorithms with applications to sensitivity analysis. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/252.

Fernandez-Echeverria:2025:EAE

[3499] Eduardo Fernández-Echeverría, Gregorio Fernández-Lambert, Luis Enrique García-Santamaría, Loecelia Ruvalcaba-Sánchez, Roberto Angel Melendez-Armenta, Juan Manuel Carrión-Delgado, and José Aparicio-Urbano. An evolutionary algorithm for efficient flexible workstation layout in multi-product handcrafted manufacturing. Algorithms (Basel), 18 (5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/270.

Galuppini:2025:TMN

[3500] Giacomo Galuppini and Paolo Venini. A transfer matrix norm-based framework for multimaterial topology optimization: Methodology and MATLAB implementation. Algorithms (Basel), 18(5), May 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/5/282.

Han:2025:QII

[3501] Gawon Han, Arun Elangovan, Jordan Wong, Asmara Waheed, Keith Wachowicz, Nawaid Usmani, Zsolt Gabos, Jihyun Yun, and B. Gino Fallone. Quantifying intra- and inter-observer variabilities in manual contours for radiotherapy: Evaluation of an MR tumor autocontouring algorithm for liver, prostate, and lung cancer patients. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/5/290.

He:2025:AEI

[3502] Ying He, Donger Chen, Isabella Xu, Wang Feng, Qing Yang, Ted Tsao, and Song Fu. Assessing environmental influences on flash storage for vehicle computing: a quantitative and analytical investigation. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/269.

Hu:2025:NAP

[3503] Hengrui Hu, Anai N. Kothari, and Anjishnu Banerjee. A novel algorithm for personalized federated learning: Knowledge distillation with weighted combination loss. Algorithms (Basel), 18(5), May 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/5/274.

Huang:2025:FLF

[3504] Wanwei Huang, Wenqiang Gui, Yingying Li, Qingsong Lv, Jia Zhang, and Xi He. Faulty links’ fast recovery method based on deep reinforcement learning. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/5/241.

Huang:2025:IDB

[3505] Zicheng Huang, Zuopeng Liang, Shibo Zhou, and Shuntao Zhang. An improved density-based spatial clustering of applications with noise algorithm with an adaptive parameter based on the sparrow search algorithm. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/
273.

Kansizoglou:2025:AIH

[3506] Ioannis Kansizoglou, Christos Kokkotis, Theodoros Stampoulis, Erasmia Giannakou, Panagiotis Siaperas, Stavros Kallidis, Maria Koutra, Paraskevi Malliou, Maria Michalopoulou, and Antonios Gasteratos. Artificial intelligence and the human-computer interaction in occupational therapy: a scoping review. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/276.

Karousos:2025:ACT

[3507] Nikos Karousos, Despoina Pantazi, George Vorvilas, and Vassilios S. Verykios. Assigning candidate tutors to modules: a Preference Adjustment Matching Algorithm (PAMA). Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/5/250.

Kaviani:2025:FCI

[3508] Ehsan Kaviani and Kalpdrum Passi. Forecasting cancer incidence in Canada by age, sex, and region until 2026 using machine learning techniques. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/
265.

Korn:2025:SIE

[3509] Ralf Korn and Laurena Ramadani. Simulating intraday electricity consumption with ForGAN. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/256.

Krutikov:2025:AAS

[3510] Vladimir Krutikov, Elena Tovbis, and Lev Kazakovtsev. Adam algorithm with step adaptation. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/268.

Kumar:2025:ICT

[3511] N. Manoj Kumar, Yousef Farhaoui, R. Vimala, M. Anandan, M. Aiswarya, and A. Radhika. Innovative control techniques for enhancing signal quality in power applications: Mitigating electromagnetic interference. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/5/288.

Li:2025:DPD

[3512] Shuai Li, Peicheng Shi, Aixi Yang, Heng Qi, and Xinlong Dong. Dualpriority delayed deep double Q-network (DPD3QN): a dueling double deep Q-network with dual-priority experience replay for autonomous driving behavior decision-making. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/291.

Li:2025:MSU

[3513] Ziheng Li, Benyuan Chen, Wanting Wang, Hui Lv, Qinghua Lv, Jie Chen, Yan Wang, Juan Li, and Cheng Zhang. A method for synthesizing ultra-large-scale clock trees. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/249.

Li:2025:VPP

[3514] Yonglin Li, Zhao Liu, Changtao Kan, Rongfei Qiao, Yue Yu, and Changgang Li. A virtual power plant-integrated proactive voltage regulation framework for urban distribution networks: Enhanced termite life cycle optimization algorithm and dynamic coordination. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/251.

Liang:2025:DSE

[3515] Haijun Liang, Yimin He, Hanwen Chang, and Jianguo Kong. DeCGAN: Speech enhancement algorithm for air traffic control. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/245.

Loudiki:2025:NAC

[3516] Laila Loudiki and Mustapha Kchikech. A new algorithm for computing the distance and the diameter in circulant graphs. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/261.

Ma:2025:ITA

[3517] Junxia Ma, Qilin Liu, Zixu Yang, and Bo Wang. Improved trimming ant colony optimization algorithm for mobile robot path planning. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/240.

Malakov:2025:DOS

[3518] Ivo Malakov, Velizar Zaharinov, and Hasan Hasansabri. Development of optimal size range of modules for driving of automatic sliding doors. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/248.

Mercado:2025:LAN

[3519] Jhon Mercado, Carlos Mendoza-Cardenas, Luis Fletscher, and Natalia Gaviria-Gomez. Learning analytics in a non-linear virtual course. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/284.

Nikolopoulos:2025:GNN

[3520] Anastasios Nikolopoulos and Vangelis D. Karalis. Generative neural networks for addressing the bioequivalence of highly variable drugs. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/266.

Panagiotakis:2025:FEC

[3521] Costas Panagiotakis. Fast equipartition of complex 2D shapes with minimal boundaries. Algorithms (Basel), 18(5), May 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/277.

Parlangeli:2025:LCO

[3522] Gianfranco Parlangeli. Laplacian controllability and observability of multi-agent systems: Recent advances in tree graphs. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/239.

Penate-Rodriguez:2025:SLH

[3523] Helen C. Peñate-Rodríguez, Gilberto Rivera, J. Patricia Sánchez-Solís, and Rogelio Florencia. The scientific landscape of hyper-heuristics: a bibliometric analysis based on Scopus. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/5/294.

Peng:2025:EED

[3524] Hao Peng, Juan Bayon, Joaquin Recas, and Maria Guijarro. Efficient expiration date recognition in food packages for mobile applications. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/286.

Pratibha:2025:RFH

[3525] Pratibha and Naveed Mahmud. A reconfigurable framework for hybrid quantum-classical computing. Algorithms (Basel), 18(5), May 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/5/271.

Qin:2025:SNN

[3526] Panke Qin, Yongjie Ding, Ya Li, Bo Ye, Zhenlun Gao, Yaxing Liu, Zhongqi Cai, and Haoran Qi. Spiking neural networks optimized by improved cuckoo search algorithm: a model for financial time series forecasting. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/
262.

Sarsenbay:2025:GJR

[3527] Shakhmar Sarsenbay, Asset Kabdiyev, Iraklis Varlamis, Christos Sardianos, Cemil Turan, Bobir Razhametov, and Yermek Kazym. Generating job recommendations based on user personality and Gallup tests. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/275.

Suhardiman:2025:SCG

[3528] Basuki Suhardiman, Kuntjoro Adji Sidarto, and Novriana Sumarti. A simulation of contact graph routing for Mars–Earth data communication. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/293.

Sun:2025:DDE

[3529] Yange Sun, Siyu Geng, Chengyi Zheng, Chenglong Xu, Huaping Guo, and Yan Feng. DEPANet: a differentiable edge-guided pyramid aggregation network for strip steel surface defect segmentation. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/5/279.

Sun:2025:SOM

[3530] Jiandong Sun, Shixun You, Di Hua, Zhiwei Xu, Peiyao Wang, and Zihang Yang. Simulation and optimization of multi-phase terminal trajectory for three-dimensional anti-ship missiles based on hybrid MOPSO. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/278.

Szenasi:2025:DLB

[3531] Sándor Szénási, Gábor Légrádi, and Gábor Kovács. Deep learningbased step size determination for hill climbing metaheuristics. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/298.

Tao:2025:IMS

[3532] Wenzheng Tao, Sarang Joshi, and Ross Whitaker. Integrated model selection and scalability in functional data analysis through Bayesian learning. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/5/254.

Tavafi:2025:ERP

[3533] Mehdi Tavafi, Kalpdrum Passi, and Robert Ohle. Early risk prediction in acute aortic syndrome on clinical data using machine learning. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/257.

Trinh:2025:UAV

[3534] Mien L. Trinh, Dung T. Nguyen, Long Q. Dinh, Mui D. Nguyen, De Rosal Ignatius Moses Setiadi, and Minh T. Nguyen. Unmanned aerial vehicles (UAV) networking algorithms: Communication, control, and AI-based approaches. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/5/244.

Tussupov:2025:AST

[3535] Jamalbek Tussupov, Akmaral Kassymova, Ayagoz Mukhanova, Assyl Bissengaliyeva, Zhanar Azhibekova, Moldir Yessenova, and Zhanargul Abuova. Analysis of short texts using intelligent clustering methods. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/289.

Urbina-Salas:2025:MGI

[3536] Ismael Urbina-Salas, David Granados-Lieberman, Martín Valtierra-Rodríguez, Claudia Adriana Ramírez-Valdespino, and David Aarón Rodríguez-Alejandro. Modeling of global and individual kinetic parameters in wheat straw torrefaction: Particle swarm optimization and its impact on elemental composition prediction. Algorithms (Basel), 18 (5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/5/283.

Valova:2025:GRS

[3537] Irena Valova, Katerina G. Gabrovska-Evstatieva, Tsvetelina Kaneva, and Boris I. Evstatiev. Generation of realistic synthetic load profile based on the Markov chains theory: Methodology and case studies. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/5/287.

Wang:2025:RED

[3538] Chia-Nan Wang, Nhat-Luong Nhieu, and Yu-Cin Ye. A regret-enhanced DEA approach to mapping renewable energy efficiency in Asia’s growth economies. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/5/297.

Wawrzyniak:2025:LAD

[3539] Wojciech Wawrzyniak. A local 6-approximation distributed algorithm for minimum dominating set problem in planar triangle-free graphs. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/280.

Xiong:2025:PSO

[3540] Zhongqin Xiong, Shichang Huang, Shen Ren, Yutong Lin, Zewen Li, Dongyu Li, and Fangming Deng. Particle swarm optimization support vector machine-based grounding fault detection method in distribution network. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/
259.

Yang:2025:AAF

[3541] Xianyang Yang, Wei Shu, Huoping Wang, Haifeng Li, Yi Wang, Di Zhang, Jiayu Liu, Deyang Wang, and Wangsui Xiao. Adjustment algorithm for free station control network of ultra-large deepwater jacket. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/5/292.

Ying:2025:HHP

[3542] Yuhong Ying, Haoyuan Li, Yiwen Zhong, and Min Lin. HPANet: Hierarchical path aggregation network with pyramid vision transformers for colorectal polyp segmentation. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/281.

Young:2025:BML

[3543] Richard J. Young, Alice M. Matthews, and Brach Poston. Benchmarking multiple large language models for automated clinical trial data extraction in aging research. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/296.

Yousif:2025:SEA

[3544] Sally Taha Yousif and Basheera M. Mahmmod. Speech enhancement algorithms: a systematic literature review. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/5/272.

Yuan:2025:ADO

[3545] Fujiang Yuan, Zihao Zuo, Yang Jiang, Wenzhou Shu, Zhen Tian, Chenxi Ye, Junye Yang, Zebing Mao, Xia Huang, Shaojie Gu, and Yanhong Peng. AI-driven optimization of blockchain scalability, security, and privacy protection. Algorithms (Basel), 18(5), May 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/5/263.

Zhang:2025:DDD

[3546] Bin Zhang, Dan Wang, and Fuzhong Wang. Data-driven distributed model-free adaptive predictive control for multiple high-speed trains under false data injection attacks. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/267.

Zhang:2025:PPT

[3547] Kai Zhang, Wangqing Niu, and Kailun Zhang. Phase plane trajectory planning for double pendulum crane anti-sway control. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/5/246.

Zheng:2025:BII

[3548] Danqi Zheng, Jiyun Qin, Zhen Liu, Qinglei Zhang, Jianguo Duan, and Ying Zhou. BWO-ICEEMDAN-iTransformer: a short-term load forecasting model for power systems with parameter optimization. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/5/243.

Zheng:2025:HGS

[3549] Fuyuan Zheng, Kanglong Cheng, Kai Yang, Ning Li, Yu Lin, and Yiwen Zhong. A heuristics-guided simplified discrete harmony search algorithm for solving 0–1 knapsack problem. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/295.

Zhou:2025:MFF

[3550] Meng Zhou, Xiaoyi Zhou, Zhijian Li, Xinyue Liu, and Chengming Chen. A multi-feature fusion algorithm for fatigue driving detection considering individual driver differences. Algorithms (Basel), 18(5), May 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/5/247.

Chen:2025:IDM

[3551] Yan Chen, Yi Chen, Yang Guo, Longda Wang, and Gang Liu. Improved directional mutation moth-flame optimization algorithm via gene modification for automatic reverse parking trajectory optimization. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/299.

Tan:2025:ROD

[3552] Pengliu Tan, Zhi Wang, and Xin Chang. Road obstacle detection method based on improved YOLOv5. Algorithms (Basel), 18(6), June 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/6/300.

Nodal:2025:EOH

[3553] Yarelis Valdivia Nodal, Luis Angel Iturralde Carrera, Araceli Zapatero-Gutiérrez, Mario Antonio Álvarez Guerra Plasencia, Royd Reyes Calvo, José M. Álvarez-Alvarado, and Juvenal Rodríguez-Reséndiz. Energy optimization in hotels: Strategies for efficiency in hot water systems. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/301.

Gunasekaran:2025:EVG

[3554] Hemalatha Gunasekaran, Nesaian Reginal Wilfred Blessing, Umar Sathic, and Mohammad Shahid Husain. Enhanced viral genome classification using large language models. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/302.

Onyenagubo:2025:FRU

[3555] Chisom Onyenagubo, Yasser Ismail, Radian Belu, and Fred Lacy. Forecasting the remaining useful life of lithium-ion batteries using machine learning models — a Web-based application. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/303.

Li:2025:RCMa

[3556] Xiaoxu Li, Jiahao Wang, Jianqiang Wang, Jixuan Wang, Jiaming Chen, and Xuelian Yu. Research on CNC machine tool spindle fault diagnosis method based on DRSN-GCE model. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/304.

Kabir:2025:SSL

[3557] Sami Kabir, Mohammad Shahadat Hossain, and Karl Andersson. A semi-supervised-learning-aided explainable belief rule-based approach to predict the energy consumption of buildings. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/305.

Hakizimana:2025:NDR

[3558] Gedeon Hakizimana and Agapito Ledezma Espino. Nomological deductive reasoning for trustworthy, human-readable, and actionable AI outputs. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/
306.

Zhang:2025:SAB

[3559] Xuan Zhang and Chaojie Wang. Structure approximation-based preconditioning for solving tempered fractional diffusion equations. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/307.

Scarpa:2025:PPR

[3560] Marco Scarpa, Marco Garofalo, Francesco Longo, and Salvatore Serrano. Polarization in political rallies: a Markovian agent-based model for opinion dynamics. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/6/308.

Freixes:2025:CAN

[3561] Alfons Freixes, Javier Panadero, Angel A. Juan, and Carles Serrat. Combining the A∗ algorithm with neural networks to solve the team orienteering problem with obstacles and environmental factors. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/309.

Zhabinets:2025:SDB

[3562] Maxim Zhabinets, Benjamin Tyler, Martin Lukac, Shinobu Nagayama, Ferdinand Molnár, and Michitaka Kameyama. Synthetic data-based algorithm selection for medical image classification under limited data availability. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/6/310.

Randhawa:2025:DPT

[3563] Shan Randhawa, Abbas Shojaee, Elisa Sorrentino, Yifan Li, Azza Abouzied, and Dennis Shasha. DataToCare: Predicting treatments for intensive care unit patients based on similarity of abnormalities. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/311.

Cirstea:2025:BAM

[3564] Ilie Cirstea, Andrei-Flavius Radu, Delia Mirela Tit, Ada Radu, Gabriela Bungau, and Paul Andrei Negru. Bibliometric analysis of medical waste research using Python-driven algorithm. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/312.

Li:2025:IAF

[3565] Aiyou Li, Huoyou Li, Yanfang Liu, and Guoli Ji. Improved asynchronous federated learning for data injection pollution. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/313.

Aghileh:2025:MPS

[3566] Marzieh Aghileh, Anabela Tereso, Filipe Alvelos, and Maria Odete Monteiro Lopes. Multi-project scheduling with uncertainty and resource flexibility: a narrative review and exploration of future landscapes. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/314.

Dumitrescu:2025:GPS

[3567] Adrian Dumitrescu. General position subset selection in line arrangements. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/
315.

Xue:2025:AGH

[3568] Xinru Xue, Ting Shu, and Jinsong Xia. Automated generation of hybrid metaheuristics using learning-to-rank. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/316.

Jaradat:2025:ADQ

[3569] Shadi Jaradat, Nirmal Acharya, Smitha Shivshankar, Taqwa I. Alhadidi, and Mohammad Elhenawy. AI for data quality auditing: Detecting mislabeled work zone crashes using large language models. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/317.

Zhou:2025:RKS

[3570] Yixin Zhou, Zepeng Ma, and Lei Fu. A review of key signal processing techniques for structural health monitoring: Highlighting non-parametric time-frequency analysis, adaptive decomposition, and deconvolution. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/318.

Wang:2025:WAE

[3571] Yuhonghao Wang, Wenxin Li, Xingmin Qi, and Yinzhang Yu. A weight assignment-enhanced convolutional neural network (WACNN) for freight volume prediction of sea-rail intermodal container systems. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/319.

daSilva:2025:BDS

[3572] Edvan Gomes da Silva, Fernando Rocha Moreira, Marcus Aurélio Carvalho Georg, Rildo Ribeiro dos Santos, Luiz Antônio Ribeiro Júnior, and Rafael Rabelo Nunes. Binary decision support using AHP: a model for alternative analysis. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/6/320.

Mahmood:2025:CRA

[3573] Yasir Mahmood, Nof Yasir, Nita Yodo, Ying Huang, Di Wu, and Roy A. McCann. Comprehensive risk assessment of power grids using fuzzy Bayesian networks through expert elicitation: a technical analysis. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/321.

Zhang:2025:MBE

[3574] Keshu Zhang, Jiandong Jia, Xiaodan Shangguan, and Jing Dong. Modelica-based energy management of PEMFC hybrid power system of vehicle. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/
322.

Gurski:2025:CWM

[3575] Frank Gurski and Ruzayn Quaddoura. The clique-width of minimal series-parallel digraphs. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/6/323.

Decker:2025:QEH

[3576] Jonathan Decker, Vincent Florens Hasse, and Julian Kunkel. Q8S: Emulation of heterogeneous Kubernetes clusters using QEMU. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/324.

Rafique:2025:MRM

[3577] Wajid Rafique. ML-RASPF: a machine learning-based rate-adaptive framework for dynamic resource allocation in smart healthcare IoT. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/325.

Lappas:2025:MCW

[3578] Demetris Lappas, Vasileios Argyriou, and Dimitrios Makris. Masked convolutions within skip connections for video anomaly detection. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/326.

Bedalli:2025:FCA

[3579] Erind Bedalli, Shkelqim Hajrulla, Rexhep Rada, and Robert Kosova. Fuzzy clustering approaches based on numerical optimizations of modified objective functions. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/6/327.

Rescigno:2025:SAP

[3580] Adele Anna Rescigno. Selected algorithmic papers from the international workshop on combinatorial algorithms 2024. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/328.

Lu:2025:HMG

[3581] Aiguo Lu, Zican Li, Yanwei Liu, Pandi Liu, and Ke Wang. Hybrid multigranularity approach for few-shot image retrieval with weak features. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/329.

Bashkirtseva:2025:SDS

[3582] Irina A. Bashkirtseva, Lev B. Ryashko, Ivan N. Tsvetkov, and Alexander N. Pisarchik. Stochastic disruption of synchronization patterns in coupled non-identical neurons. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/330.

Naseeb:2025:MLD

[3583] Amna Naseeb, Muhammad Zain, Nisar Hussain, Amna Qasim, Fiaz Ahmad, Grigori Sidorov, and Alexander Gelbukh. Machine learning- and deep learning-based multi-model system for hate speech detection on Facebook. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/6/331.

Wang:2025:KBV

[3584] Yadong Wang, Xinyao Duan, Jiang Wang, Jin Guo, and Minglei Han. Kriging-based variable screening method for aircraft optimization problems with expensive functions. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/332.

Waqar:2025:EHA

[3585] Muhammad Waqar, Muhammad Bilal Shahnawaz, Sajid Saleem, Hassan Dawood, Usman Muhammad, and Hussain Dawood. Enhancing heart attack prediction: Feature identification from multiparametric cardiac data using explainable AI. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/6/333.

Turar:2025:AGG

[3586] Olzhas Turar, Maksat Mustafin, and Darkhan Akhmed-Zaki. Adaptive grid generation by solving one-dimensional diffusion equation using physics-informed neural networks. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/334.

Levin:2025:ABS

[3587] David Levin and Nira Gruberger. Adapted B-spline quasi-interpolation for approximating piecewise smooth functions. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/335.

Chavarria:2025:NSP

[3588] Aleida Chavarria, Ramon Palau, and Raúl Santiago. Navigating stakeholders perspectives on artificial intelligence in higher education. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/336.

Prinzi:2025:MLA

[3589] Francesco Prinzi, Ines Prata Machado, and Carmelo Militello. Machine learning algorithms for biomedical image analysis and their applications. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/337.

Habib:2025:MOB

[3590] Maria Habib, Victor Vicente-Palacios, and Pablo García-Sánchez. A multi-objective bio-inspired optimization for voice disorders detection: a comparative study. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/6/338.

Tsamis:2025:CED

[3591] George Tsamis, Georgios Evangelos, Aris Papakostas, Giannis Vassiliou, Michael Grafanakis, Alexandros Garefalakis, Michalis Vassalos, Anastasia Mylona, and Nikos Papadakis. Cost-effective design, content management system implementation and artificial intelligence support of Greek government AADE, myDATA Web service for generic government infrastructure, a complete analysis. Algorithms (Basel), 18(6), June 2025.

CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/339.

Grieco:2025:BKP

[3592] Antonio Grieco, Pierpaolo Caricato, and Paolo Margiotta. Beyond Kanban: POLCA-constrained scheduling for job shops. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/340.

Guilbault:2025:ESE

[3593] Raynald Guilbault. S-EPSO: a socio-emotional particle swarm optimization algorithm for multimodal search in low-dimensional engineering applications. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/6/341.

Komatina:2025:OPT

[3594] Nikola Komatina and Dragan Marinković. Optimization of PFMEA team composition in the automotive industry using the IPF-RADAR approach. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/6/342.

Nasef:2025:SVF

[3595] Daniel Nasef, Demarcus Nasef, Michael Sher, and Milan Toma. A standardized validation framework for clinically actionable healthcare machine learning with knee osteoarthritis grading as a case study. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/343.

Boiangiu:2025:NCC

[3596] Costin-Anton Boiangiu, Giorgiana-Violeta Vlasceanu, Constantin-Eduard Staniloiu, Nicolae Tarba, and Mihai-Lucian Voncila. A novel connected-components algorithm for 2D binarized images. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/344.

Majeed:2025:AGS

[3597] Bilal Majeed, Rajkumar Sarma, Ayman Youssef, Douglas Mota Dias, and Conor Ryan. Automatic generation of synthesisable hardware description language code of multi-sequence detector using grammatical evolution. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/345.

Zhang:2025:ITA

[3598] Hongbin Zhang, McKaylee Robertson, Sarah L. Braunstein, David B. Hanna, Uriel R. Felsen, Levi Waldron, and Denis Nash. Inferring the timing of antiretroviral therapy by zero-inflated random change point models using longitudinal data subject to left-censoring. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/346.

Chen:2025:NMA

[3599] Xi Chen, Feng Wang, Li Xu, Taiwu Xia, Minhao Wang, Gangping Chen, Longyu Chen, and Jun Zhou. A novel model for accurate daily urban gas load prediction using genetic algorithms. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/347.

Rasheed:2025:NTT

[3600] Saburi Rasheed, Olaniyi S. Iyiola, Segun I. Oke, and Bruce A. Wade. Numerical and theoretical treatments of the optimal control model for the interaction between diabetes and tuberculosis. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/348.

Barsanti:2025:TEV

[3601] Sara Gonizzi Barsanti and Ernesto Nappi. Testing the effectiveness of voxels for structural analysis. Algorithms (Basel), 18(6), June 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/6/349.

Sinthong:2025:MEB

[3602] Phanwadee Sinthong, Nam Nguyen, Vijay Ekambaram, Arindam Jati, Jayant Kalagnanam, and Peeravit Koad. Memory-efficient batching for time series transformer training: a systematic evaluation. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/350.

Yang:2025:RIL

[3603] Jingjing Yang. An RMSprop-incorporated latent factorization of tensor model for random missing data imputation in structural health monitoring. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/

Munaye:2025:CES

[3604] Yirga Yayeh Munaye, Wasyihun Admass, Yenework Belayneh, Atinkut Molla, and Mekete Asmare. ChatGPT in education: a systematic review on opportunities, challenges, and future directions. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/352.

Arauzo-Azofra:2025:EOC

[3605] Antonio Arauzo-Azofra, Jose Molina-Baena, and Maria Luque-Rodriguez. Evolutionary optimization for the classification of small molecules regulating the circadian rhythm period: a reliable assessment. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/353.

Pan:2025:RHV

[3606] Yuede Pan, Kaifeng Zhong, Yubao Xie, Mingzhang Pan, Wei Guan, Li Li, Changye Liu, Xingjia Man, Zhiqing Zhang, and Mantian Li. A review of hybrid vehicles classification and their energy management strategies: an exploration of the advantages of genetic algorithms. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/354.

Munaye:2025:EDM

[3607] Yirga Yayeh Munaye, Aneas Bekele Workneh, Yenework Belayneh Chekol, and Atinkut Molla Mekonen. Effective detection of malicious uniform resource locator (URLs) using deep-learning techniques. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/355.

Tepsan:2025:IHT

[3608] Worawit Tepsan, Aniwat Phaphuangwittayakul, Saronsad Sokantika, and Napat Harnpornchai. Identifying hubs through influential nodes in transportation network by using a gravity centrality approach. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/356.

Tang:2025:COH

[3609] Yongli Tang, Zhenlun Gao, Ya Li, Zhongqi Cai, Jinxia Yu, and Panke Qin. Crude oil and hot-rolled coil futures price prediction based on multidimensional fusion feature enhancement. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/357.

Cheng:2025:DYR

[3610] Siqiang Cheng, Lingshan Chen, and Kun Yang. DGSS-YOLOv8s: a real-time model for small and complex object detection in autonomous vehicles. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/
358.

Li:2025:AAD

[3611] Danlei Li, Nirmal-Kumar C. Nair, and Kevin I-Kai Wang. ADDAEIL: Anomaly detection with drift-aware ensemble-based incremental learning. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/
359.

Caridi:2025:BBN

[3612] Yuri Caridi, Andrea Cucuzzella, Fabio Vicini, and Stefano Berrone. A box-bounded non-linear least square minimization algorithm with application to the JWL parameter determination in the isentropic expansion for highly energetic material simulation. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/360.

Boche:2025:CCR

[3613] Holger Boche and Christian Deppe. The computability of the channel reliability function and related bounds. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/6/361.

Akritidis:2025:WWI

[3614] Leonidas Akritidis and Panayiotis Bozanis. WIRE: a weighted item removal method for unsupervised rank aggregation. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/362.

Li:2025:AGA

[3615] Wenxin Li and Haifeng Zhang. Algorithms for game AI. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/363.

Yang:2025:TSH

[3616] Yong Yang, Caixia Chen, Shiming Yuan, and Yonghua Yan. A two-step high-order compact corrected WENO scheme. Algorithms (Basel), 18 (6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/6/364.

Li:2025:ECA

[3617] Zhengang Li, Xinpei Wu, and Junseok Kim. An efficient computational algorithm for the nonlocal Cahn–Hilliard equation with a spacedependent parameter. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/6/365.

Yuan:2025:CBD

[3618] Qingyang Yuan and Bo Zhang. A clustering-based dimensionality reduction method guided by POD structures and its application to convective flow problems. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/6/366.

Hadad:2025:TSP

[3619] Yossi Hadad and Baruch Keren. A two-stage bin packing algorithm for minimizing machines and operators in cyclic production systems. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/367.

Veledar:2025:SEA

[3620] Emir Veledar, Lili Zhou, Omar Veledar, Hannah Gardener, Carolina M. Gutierrez, Jose G. Romano, and Tatjana Rundek. Synthesizing explainability across multiple ML models for structured data. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/6/368.

Zhang:2025:FFE

[3621] Shizheng Zhang, Kunpeng Wang, Zhihao Liu, Min Huang, and Sheng Huang. The fine feature extraction and attention re-embedding model based on the swin transformer for pavement damage classification. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/6/369.

Luo:2025:DPD

[3622] Yirong Luo, Yanping Du, Zhaohua Wang, Jingtian Mo, Wenxuan Yu, and Shuihai Dou. DScanNet: Packaging defect detection algorithm based on selective state space models. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/370.

Brank:2025:SCI

[3623] Janez Brank and Tomaz Hocevar. Substring counting with insertions. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/6/371.

Ma:2025:LRP

[3624] Junxia Ma, Yongxuan Sang, Yaoli Xu, and Bo Wang. A linear regression prediction-based dynamic multi-objective evolutionary algorithm with correlations of Pareto front points. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/372.

Liu:2025:SSS

[3625] Yaoxian Liu, Kaixin Zhang, Yue Sun, Jingwen Chen, and Junshuo Chen. A small-sample scenario optimization scheduling method based on multidimensional data expansion. Algorithms (Basel), 18(6), June 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/6/373.

Shadimetov:2025:WOQ

[3626] Kholmat Shadimetov and Khojiakbar Usmanov. Weighted optimal quadrature formulas in Sobolev space and their applications. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/374.

Gaspar:2025:PFS

[3627] Joana S. D. Gaspar, Maria A. R. Loja, and Joaquim I. Barbosa. Plane frame structures: Optimization and design solutions clustering. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/375.

Zheng:2025:PSM

[3628] Hong Zheng, Zhenhui Xu, Qihong Pan, Zhenzhen Zhao, and Xiangjie Kong. Plugging small models in large language models for POI recommendation in smart tourism. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/7/376.

Ahmadi:2025:ADR

[3629] Nur Ahmadi, Hansel Valentino Tanoto, and Rinaldi Munir. Automated digit recognition and measurement-type classification from blood pressure monitor images using deep learning. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/377.

Santos:2025:ASA

[3630] Gonçalo Santos, Joana Cardoso-Fernandes, and Ana C. Teodoro. Aether-Geo: a spectral analysis interface for geologic mapping. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/378.

Sub-r-pa:2025:SSK

[3631] Chayanon Sub r pa and Rung-Ching Chen. Self-supervised keypoint learning for the geometric analysis of road-marking templates. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/379.

Zhang:2025:DTL

[3632] Nan Zhang, Qijing Lin, and Zhuangde Jiang. Decision-time learning and planning integrated control for the mild hyperbaric chamber. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/380.

Pandey:2025:HST

[3633] Rupesh Krishna Pandey, Balendu Bhooshan Upadhyay, Subham Poddar, and Ioan Stancu-Minasian. Hestenes–Stiefel-type conjugate direction algorithm for interval-valued multiobjective optimization problems. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/381.

Rusu:2025:TMW

[3634] Andrei Rusu, Petru Dobra, Mihai Hulea, and Radu Miron. Time management in wireless sensor networks for industrial process control. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/382.

Yang:2025:SCS

[3635] Wenhao Yang, Liqun Kuang, Song Wang, Xie Han, Rong Guo, Yongpeng Wang, Haifeng Yue, and Tao Wei. Spatiotemporal contextual 3D semantic segmentation for intelligent outdoor mining. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/383.

Hamdi:2025:CAA

[3636] Abdelouahed Hamdi, Arezou Karimi, Farshid Mehrdoust, and Samir Brahim Belhaouari. Comparative analysis of artificial neural networks and evolutionary algorithms in DEA-β-MSV portfolio optimization. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/384.

Chab:2025:ATG

[3637] Robert Chab, Fei Li, and Sanjeev Setia. Algorithmic techniques for GPU scheduling: a comprehensive survey. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/385.

Gurski:2025:BTW

[3638] Frank Gurski and Robin Weishaupt. The behavior of tree-width and path-width under graph operations and graph transformations. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/386.

Chen:2025:EEA

[3639] Gang Chen, Shan Hua, Changhao Fan, Chun Wang, Shuchong Wang, and Li Sun. An enhanced error-adaptive extended-state Kalman filter model predictive controller for supercritical power plants. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/387.

Torrens:2025:SSP

[3640] Paul M. Torrens, Ryan Kim, and Kaishuu Shinozaki-Conefrey. Situationally sensitive path planning. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/7/388.

Shams:2025:CEF

[3641] Mudassir Shams and Bruno Carpentieri. Chaos-enhanced fractionalorder iterative methods for the stable and efficient solution of nonlinear engineering problems. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/389.

Abdelazeem:2025:AAR

[3642] Mohamed Abdelazeem, Amgad Abazeed, Abdulmajeed Alsultan, and Amr M. Wahaballa. Assessment of affordable real-time PPP solutions for transportation applications. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/7/390.

Barrera-Fernandez:2025:PDA

[3643] Juan M. Barrera-Fernández, Juan Pablo Manzo Hernández, Kevin Miramontes Escobedo, Alberto Vázquez-Cervantes, and Julio-César Solano-Vargas. PI-D^{Æ}: An adaptive PID controller utilizing a new adaptive exponent (Æ) algorithm to solve derivative term issues. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/391.

Stepanov:2025:BDC

[3644] Arthur Stepanov, Hristo Zhivomirov, Ivaylo Nedelchev, Todor Ganchev, and Penka Stateva. Bottlenose dolphins’ clicks comply with three laws of efficient communication. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/392.

Yang:2025:UFR

[3645] Jiayi Yang, Jiao Liang, Huimin Pan, Yuting Cai, Quanli Gao, and Xihan Wang. A unified framework for recognizing dynamic hand actions and estimating hand pose from first-person RGB videos. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/393.

Le:2025:SVS

[3646] Van-Hung Le and Thi-Ha-Phuong Nguyen. A survey of visual SLAM based on RGB-D images using deep learning and comparative study for VOE. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/
394.

Elgindy:2025:SEA

[3647] Kareem T. Elgindy. Super-exponential approximation of the Riemann– Liouville fractional integral via Gegenbauer-based fractional approximation methods. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/395.

Zain:2025:ROC

[3648] Muhammad Zain, Nisar Hussain, Amna Qasim, Gull Mehak, Fiaz Ahmad, Grigori Sidorov, and Alexander Gelbukh. RU-OLD: a comprehensive analysis of offensive language detection in Roman Urdu using hybrid machine learning, deep learning, and transformer models. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/396.

Ispas-Gil:2025:ASS

[3649] Decebal Aitor Ispas-Gil, Ekaitz Zulueta, Javier Olarte, and Jose Manuel Lopez-Guede. Algorithms for simulation of shunt currents in a vanadium redox flow battery. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/397.

Alfadli:2025:ATM

[3650] Nawaf Mijbel Alfadli, Eman Mostafa Oun, and Ali Wagdy Mohamed. Auto-tuning memory-based adaptive local search gaining-sharing knowledge-based algorithm for solving optimization problems. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/398.

Alevizos:2025:IGG

[3651] Vasileios Alevizos, Emmanouil V. Gkouvrikos, Ilias Georgousis, Sotiria Karipidou, and George A. Papakostas. Inspiring from galaxies to green AI in Earth: Benchmarking energy-efficient models for galaxy morphology classification. Algorithms (Basel), 18(7), July 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/399.

Zhang:2025:TSP

[3652] Yijie Zhang and Jizhou Chen. The true shortest path of obstacle grid graph is solved by SGP vertex extraction and filtering algorithm. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/400.

Hemdan:2025:SSH

[3653] Ezz El-Din Hemdan and Amged Sayed. Smart and secure healthcare with digital twins: a deep dive into blockchain, federated learning, and future innovations. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/401.

Nikolova:2025:ESE

[3654] Elena V. Nikolova. On the extended simple equations method (SEsM) for obtaining numerous exact solutions to fractional partial differential equations: a generalized algorithm and several applications. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/402.

Wang:2025:RAD

[3655] Xizheng Wang, Gang Li, and Zijian Bian. Research on APF-Dijkstra path planning fusion algorithm based on steering model and volume constraints. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/
403.

Brandao:2025:USF

[3656] Pedro Brandão and Carla Silva. Unveiling the shadows-a framework for APT’s defense AI and game theory strategy. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/404.

Kopitsa:2025:PME

[3657] Constantina Kopitsa, Ioannis G. Tsoulos, and Vasileios Charilogis. Predicting the magnitude of earthquakes using grammatical evolution. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/405.

Alrekaby:2025:SIT

[3658] Suhad Naji Alrekaby, Maisa’a Abid Ali Khodher, Layth Kamil Adday, and Reem Aljuaidi. Secure image transmission using multilevel chaotic encryption and video steganography. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/406.

Rigopouli:2025:WOA

[3659] Kalliopi Rigopouli, Dimitrios Kotsifakos, and Yannis Psaromiligkos. The whale optimization algorithm and Markov chain Monte Carlo-based approach for optimizing teacher professional development in creative learning design with technology. Algorithms (Basel), 18(7), July 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/7/407.

Li:2025:ATA

[3660] Chenhao Li, Jiyin Zhang, Weilin Chen, and Xiaogang Ma. An approach to trustworthy article ranking by NLP and multi-layered analysis and optimization. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/408.

Wang:2025:LTF

[3661] Xiaoyu Wang, Te Chen, and Jiankang Lu. Longitudinal tire force estimation method for 4WIDEV based on data-driven modified recursive subspace identification algorithm. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/7/409.

Perez-Portabella:2025:IML

[3662] Antonio Pérez-Portabella, Jorge de Andrés-Sánchez, Mario Arias-Oliva, and Mar Souto-Romero. Integrating machine learning techniques and the theory of planned behavior to assess the drivers of and barriers to the use of generative artificial intelligence: Evidence in Spain. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/410.

Kabashkin:2025:DDT

[3663] Igor Kabashkin. Development of digital training twins in the aircraft maintenance ecosystem. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/411.

Levner:2025:ARE

[3664] Eugene Levner and Boris Kriheli. Analysis of reliability and efficiency of information extraction using AI-based chatbot: The more-for-less paradox. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/
412.

Amuda:2025:DSC

[3665] Kamorudeen Amuda, Almustapha Wakili, Tomilade Amoo, Lukman Agbetu, Qianlong Wang, and Jinjuan Feng. Detecting SARS-CoV-2 in CT scans using vision transformer and graph neural network. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/413.

Rupanetti:2025:IAS

[3666] Dulana Rupanetti, Corissa Uberecken, Adam King, Hassan Salamy, Cheol-Hong Min, and Samantha Schmidgall. An industry application of secure augmentation and gen-AI for transforming engineering design and manufacturing. Algorithms (Basel), 18(7), July 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/414.

Tadros:2025:EOM

[3667] Mina Tadros. Engine optimization model for accurate prediction of friction model in marine dual-fuel engine. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/415.

Brkic:2025:EES

[3668] Dejan Brkić. European efficiency schemes for domestic gas boilers: Estimation of savings in heating of settlements. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/416.

Warner:2025:ESA

[3669] Charles Warner, Fanyou Wu, Rado Gazo, Bedrich Benes, and Songlin Fei. Environmental sensitivity in AI tree bark detection: Identifying key factors for improving classification accuracy. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/417.

Yu:2025:RIS

[3670] Guanfei Yu and Kunlun Zhang. Research on the integrated scheduling of imaging and data transmission for Earth observation satellites. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/418.

Kapogianni:2025:USS

[3671] Nikoletta-Anna Kapogianni, Angeliki Sideraki, and Christos-Nikolaos Anagnostopoulos. Using smartwatches in stress management, mental health, and well-being: a systematic review. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/419.

Masri:2025:CAB

[3672] Yahya Masri, Zifu Wang, Anusha Srirenganathan Malarvizhi, Samir Ahmed, Tayven Stover, David W. S. Wong, Yongyao Jiang, Yun Li, Qian Liu, Mathieu Bere, Daniel Rothbart, Dieter Pfoser, and Chaowei Yang. Comparative analysis of BERT and GPT for classifying crisis news with Sudan conflict as an example. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/7/420.

Deng:2025:SSA

[3673] Jianzhi Deng, Shuilian Lv, Yun Li, Liping Luo, Yishan Su, Xiaolin Wang, and Xinzhi Liu. Ship scheduling algorithm based on Markov-modulated fluid priority queues. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/421.

Zhang:2025:FTB

[3674] Xubo Zhang and Yang Luo. Feature transformation-based few-shot classincremental learning. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/422.

Xia:2025:DMV

[3675] Dawei Xia, Guizhou Cao, Jiayao Pan, Xinghai Wang, Kai Meng, Yuancheng Sun, and Zhenlong Wu. Dynamic modeling and validation of peak ability of biomass units. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/7/423.

Mou:2025:FLB

[3676] Yi Mou, Long Zhou, Weizhen Chen, Jianguo Liu, and Teng Li. Filter learning-based partial least squares regression and its application in infrared spectral analysis. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/424.

Rosenthal:2025:RAC

[3677] Robert Rosenthal, Nils Albersmann, and Mohieddine Jelali. Robust algorithm for calculating the alignment of guide rolls in slab continuous casting machines. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/425.

He:2025:CMF

[3678] Ping He, Hanxue Zhang, Shufu Cao, and Yali Wu. Cross-modal fake news detection method based on multi-level fusion without evidence. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/426.

Zhao:2025:IRL

[3679] Jun Zhao, Jintian Ji, Robail Yasrab, Shuxin Wang, Liang Yu, and Lingzhen Zhao. Interpretable reinforcement learning for sequential strategy prediction in language-based games. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/427.

Fang:2025:SYL

[3680] Yue Fang, Chenbo Yang, Jie Li, and Jingmin Tu. SPL-YOLOv8: a lightweight method for rape flower cluster detection and counting based on YOLOv8n. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/428.

Arevalo:2025:SMM

[3681] Paul Arévalo, Dario Benavides, Danny Ochoa-Correa, Alberto Ríos, David Torres, and Carlos W. Villanueva-Machado. Smart microgrid management and optimization: a systematic review towards the proposal of smart management models. Algorithms (Basel), 18(7), July 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/7/429.

Veysi:2025:EVS

[3682] Shima Veysi, Mohammad Moshfeghi, Amir Sadrfaridpour, and Peiman Emamy. Electric vehicle sales forecast for the UK: Integrating machine learning, time series models, and global trends. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/430.

Li:2025:MLR

[3683] Dandan Li, Shaojun Zheng, Chenyu Zheng, Lingling Xie, and Li Yan. Machine learning reconstruction of Wyrtki jet seasonal variability in the Equatorial Indian Ocean. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/431.

Rajabzadeh-Oghaz:2025:CVC

[3684] Hamidreza Rajabzadeh-Oghaz, Josie Elwell, Bradley Schoch, William Aibinder, Bruno Gobbato, Daniel Wessell, Vikas Kumar, and Christopher P. Roche. Clinical validation of a computed tomography imagebased machine learning model for segmentation and quantification of shoulder muscles. Algorithms (Basel), 18(7), July 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/432.

Yang:2025:LEP

[3685] Xiong Yang, Hao Wang, Qi Zhou, Lei Lu, Lijuan Zhang, Changming Sun, and Guilu Wu. A lightweight and efficient plant disease detection method integrating knowledge distillation and dual-scale weighted convolutions. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/433.

Alam:2025:EDA

[3686] Md Minul Alam and Shahram Latifi. Early detection of Alzheimer’s disease using generative models: a review of GANs and diffusion models in medical imaging. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/434.

Raftopoulos:2025:CRB

[3687] George Raftopoulos, Nikos Fazakis, Gregory Davrazos, and Sotiris Kotsiantis. A comprehensive review and benchmarking of fairness-aware variants of machine learning models. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/435.

Jayasinghe:2025:TRD

[3688] Udara Jayasinghe, Prabhath Samarathunga, Thanuj Fernando, and Anil Fernando. Transmit and receive diversity in MIMO quantum communication for high-fidelity video transmission. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/436.

Dabees:2025:PIR

[3689] Aisha Dabees and Mehmet Hakan Karaata. Possibility and the impossibility of reliable broadcast: a 1-safe and reliable broadcast algorithm in the presence of arbitrary initialization. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/7/437.

Lin:2025:HPU

[3690] Wen-Piao Lin and Yi-Shun Lu. A high-precision UWB-based indoor positioning system using time-of-arrival and intersection midpoint algorithm. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/

Navascues:2025:IMT

[3691] María A. Navascués. Iterative matrix techniques based on averages. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/439.

Liu:2025:TET

[3692] Chein-Shan Liu and Botong Li. Two extrapolation techniques on splitting iterative schemes to accelerate the convergence speed for solving linear systems. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/440.

Dong:2025:FRP

[3693] Xuwei Dong, Jiashuo Yuan, and Jinpeng Dai. Frost resistance prediction of concrete based on dynamic multi-stage optimisation algorithm. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/441.

Lee:2025:ANP

[3694] Joo-Yeon Lee, Gang-Gyoo Jin, and Gun-Baek So. Adaptive nonlinear proportional-integral-derivative control of a continuous stirred tank reactor process using a radial basis function neural network. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/7/442.

Pelosi:2025:EIC

[3695] Daniele Pelosi, Diletta Cacciagrano, and Marco Piangerelli. Explainability and interpretability in concept and data drift: a systematic literature review. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/
443.

Hammad:2025:FDC

[3696] Ma’mon Abu Hammad, Imane Zouak, Adel Ouannas, and Giuseppe Grassi. Fractional discrete computer virus system: Chaos and complexity algorithms. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/444.

Boukham:2025:DSL

[3697] Houda Boukham, Kawtar Younsi Dahbi, and Dalila Chiadmi. Domainspecific languages for algorithmic graph processing: a systematic literature review. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/445.

Du:2025:OEC

[3698] Zhuoqun Du, Yisheng Liu, Yuyan Xue, and Boyang Liu. Optimizing economy with comfort in climate control system scheduling for indoor ice sports venues’ spectator zones considering demand response. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/7/446.

Amigo:2025:IIS

[3699] Salvador Amigó, Antonio Caselles, Joan C. Micó, and Pantaleón D. Romero. Integrating inferential statistics and systems dynamics: a study of short-term happiness evolution in response to a dose of alcohol and caffeine. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/
447.

Liu:2025:HPC

[3700] Ruofei Liu and Junjiang Zhu. High-precision chip detection using YOLObased methods. Algorithms (Basel), 18(7), July 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/7/448.

Ma:2025:DMC

[3701] Zheng Ma, Rui Zhang, and Lang Gao. Detection model for 5G core PFCP DDoS attacks based on Sin-Cos-bIAVOA. Algorithms (Basel), 18 (7), July 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/7/449.

Kascelan:2025:ENM

[3702] Jovan Kascelan, Ruoxi Yang, and Dennis Shasha. EpiInfer: a non-Markovian method and system to forecast infection rates in epidemics. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/7/450.

Galli:2025:SPG

[3703] Carlo Galli, Maria Teresa Colangelo, Marco Meleti, and Elena Calciolari. The specialist’s paradox: Generalist AI may better organize medical knowledge. Algorithms (Basel), 18(7), July 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/7/451.

Singh:2025:BBS

[3704] Vikash Singh, Matthew Khanzadeh, Vincent Davis, Harrison Rush, Emanuele Rossi, Jesse Shrader, and Pietro Lio’. Bayesian binary search. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/452.

Baibulova:2025:HCG

[3705] Makbal Baibulova, Murat Aitimov, Roza Burganova, Lazzat Abdykerimova, Umida Sabirova, Zhanat Seitakhmetova, Gulsiya Uvaliyeva, Maksym Orynbassar, Aislu Kassekeyeva, and Murizah Kassim. A hybrid CNN-GRU-LSTM algorithm with SHAP-based interpretability for EEG-based ADHD diagnosis. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/8/453.

Li:2025:DLH

[3706] Ze-Long Li, Bai Jiang, Liang Xu, Zhe Lu, Zi-Teng Wang, Bin Liu, Si-Ye Jia, Hong-Dan Liu, and Bing Li. DRFAN: a lightweight hybrid attention network for high-fidelity image super-resolution in visual inspection applications. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/454.

Khan:2025:ERF

[3707] Koffka Khan. Entropy-regularized federated optimization for non-IID data. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/
455.

Onwuchekwa:2025:SEC

[3708] Daniel Onwuchekwa, Omar Hekal, and Roman Obermaisser. Scheduling the exchange of context information for time-triggered adaptive systems. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/456.

Enisoglu:2025:NWT

[3709] Ramazan Enisoglu and Veselin Rakocevic. A novel wavelet transform and deep learning-based algorithm for low-latency Internet traffic classification. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/457.

Yang:2025:PMT

[3710] Yujun Yang, Yimei Yang, and Huijuan Liao. A predictive method for temperature based on ensemble EMD with linear regression. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/458.

Jantschi:2025:APO

[3711] Lorentz Jäntschi. Algorithm providing ordered integer sequences for sampling with replacement confidence intervals. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/459.

Grubits:2025:NCF

[3712] Péter Grubits, Tamás Balogh, and Majid Movahedi Rad. A nonlinear computational framework for optimizing steel end-plate connections using the finite element method and genetic algorithms. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/460.

Aizenberg:2025:IFD

[3713] Igor Aizenberg and Yurii Tovt. Intelligent frequency domain image filtering based on a multilayer neural network with multi-valued neurons. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/461.

Ai:2025:LPM

[3714] Wenji Ai, Jiaxuan Zou, Zongchao Liu, Shaodi Wang, and Shuai Teng. Light propagation and multi-scale enhanced DeepLabV3+ for underwater crack detection. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/462.

Wang:2025:IKA

[3715] Binghao Wang, Lei Jing, and Xiang Li. Inverse kinematics-augmented sign language: a simulation-based framework for scalable deep gesture recognition. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/463.

Sigut:2025:USV

[3716] Jose Sigut, Francisco Fumero, and Tinguaro Díaz-Alemán. Using Shapley values to explain the decisions of convolutional neural networks in glaucoma diagnosis. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/464.

Zhou:2025:EFR

[3717] Jie Zhou and Chunhua Wang. Enhancing ferroptosis-related protein prediction through multimodal feature integration and pre-trained language model embeddings. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/465.

Averina:2025:ACO

[3718] Tatyana Averina and Konstantin Rybakov. Analysis and conditional optimization of projection estimates for distribution of random variable using Legendre polynomials. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/8/466.

Tahernezhad-Javazm:2025:EBG

[3719] Farajollah Tahernezhad-Javazm, Andrew Colligan, Imelda Friel, Simon J. Hickinbotham, Paul Goodall, Edgar Buchanan, Mark Price, Trevor Robinson, and Andy M. Tyrrell. EvoDevo: Bioinspired generative design via evolutionary graph-based development. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/467.

Gao:2025:DIS

[3720] Depeng Gao, Wenjie Liu, Shuxi Chen, Jianlin Qiu, Xiangxiang Mei, and Bingshu Wang. Document image shadow removal based on illumination correction method. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/468.

Pankoon:2025:DSS

[3721] Ontima Pankoon and Nimit Nimana. Delayed star subgradient methods for constrained nondifferentiable quasi-convex optimization. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/469.

Iacobescu:2025:HAA

[3722] Paul Iacobescu and Ioan Susnea. Hybrid ARIMA-ANN for crime risk forecasting: Enhancing interpretability and predictive accuracy through socioeconomic and environmental indicators. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/470.

Tuesta:2025:PMM

[3723] Sebastian Tuesta, Nahum Flores, and David Mauricio. Prediction of the maximum and minimum prices of stocks in the stock market using a hybrid model based on stacking. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/8/471.

Xu:2025:AOP

[3724] Biao Xu, Fan Ouyang, Yangyang Li, Kun Yu, Fei Ao, Hui Li, and Liming Tan. An adaptive overcurrent protection method for distribution networks based on dynamic multi-objective optimization algorithm. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/472.

Yang:2025:NLF

[3725] Qiaohe Yang, Yueshun He, Hongmao Chen, Youyong Wu, and Zhihua Rao. A novel lightweight facial expression recognition network based on deep shallow network fusion and attention mechanism. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/473.

Liu:2025:ECS

[3726] Binglin Liu, Liwen Li, Hang Ren, Jianwan Qin, and Weijiang Liu. Evaluation of competitiveness and sustainable development prospects of French-speaking African countries based on TOPSIS and adaptive LASSO algorithms. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/474.

Fasseeh:2025:GRS

[3727] Ahmad Nader Fasseeh, Rasha Ashmawy, Rok Hren, Kareem ElFass, Attila Imre, Bertalan Németh, Dávid Nagy, Balázs Nagy, and Zoltán Vokó. Generating realistic synthetic patient cohorts: Enforcing statistical distributions, correlations, and logical constraints. Algorithms (Basel), 18 (8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/475.

Xia:2025:BSM

[3728] Shucan Xia and Lipu Zhang. Bee swarm Metropolis–Hastings sampling for Bayesian inference in the Ginzburg–Landau equation. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/476.

Bajana-Zajia:2025:CBA

[3729] Johnny Bajaña-Zajía, José Ramón Trillo, Francisco Javier Cabrerizo, and Juan Antonio Morente-Molinera. Consensus-based automatic group decision-making method with reliability and subjectivity measures based on sentiment analysis. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/477.

Chen:2025:HGS

[3730] Juan Chen and Qiao Li. Heterogeneous graph structure learning for next point-of-interest recommendation. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/18/8/478.

Ma:2025:CSS

[3731] Mingzhao Ma, Feng Wang, Lirong Xiong, Yuhonghao Wang, and Wenxin Li. Charging station siting and capacity determination based on a generalized least-cost model of traffic distribution. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/479.

Hamza:2025:QRF

[3732] Rafik Hamza. A quantum-resistant FHE framework for privacypreserving image processing in the cloud. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/480.

Su:2025:MSU

[3733] Na Su, Jingtao Wang, and Yun Pan. Multi-scale universal style-transfer network based on diffusion model. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/8/481.

Akmeemana:2025:MLS

[3734] Anuradha Akmeemana and Michael E. Sigman. Machine learning subjective opinions: an application in forensic chemistry. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/482.

George:2025:ICA

[3735] Santhosh George, Manjusree Gopal, Samhitha Bhide, and Ioannis K. Argyros. An improved convergence analysis of a multi-step method with high-efficiency indices. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/483.

Fekete:2025:TPC

[3736] Imre Fekete, Ferenc Izsák, Vendel P. Kupás, and Gustaf Söderlind. Tolerance proportionality and computational stability in adaptive parallelin-time Runge–Kutta methods. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/8/484.

Zhang:2025:NDS

[3737] Daokun Zhang, Jian Zhang, Yanhe Jia, and Mengjie Liao. A novel dualstrategy approach for constructing knowledge graphs in the home appliance fault domain. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/485.

Wei:2025:DBC

[3738] Lan Wei and Catherine Mooney. DenseNet-based classification of EEG abnormalities using spectrograms. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/8/486.

Zhou:2025:RUN

[3739] Peishan Zhou, Stephen So, and Belinda Schwerin. Robust U-nets for fetal R-peak identification in electrocardiography. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/487.

Hajighasemi:2025:FFC

[3740] Saeide Hajighasemi and Michael Breuß. Fuzzy Frankot–Chellappa algorithm for surface normal integration. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/8/488.

Akallouch:2025:CLL

[3741] Oussama Akallouch and Khalid Fardousse. In-context learning for lowresource machine translation: a study on tarifit with large language models. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/
489.

Vassiliou:2025:EDM

[3742] Giannis Vassiliou, George Tsamis, Stavroula Chatzinikolaou, Thomas Nipurakis, and Nikos Papadakis. Enhancing discoverability: a metadata framework for empirical research in theses. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/490.

Zheng:2025:EIH

[3743] Muqing Zheng, Chenxu Liu, Samuel Stein, Xiangyu Li, Johannes Mülmenstädt, Yousu Chen, and Ang Li. An early investigation of the HHL quantum linear solver for scientific applications. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/491.

Khansari:2025:ESP

[3744] Hadi Sadeghi Khansari, Mostafa Abbaszadeh, Gholamreza Heidary Joonaghany, Hamidreza Mohagerani, and Fardin Faraji. Epileptic seizure prediction using a combination of deep learning, time-frequency fusion methods, and discrete wavelet analysis. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https:

//www.mdpi.com/1999-4893/18/8/492.

Sousa:2025:CCA

[3745] João T. Sousa and Ramiro S. Barbosa. Comparison of classical and artificial intelligence algorithms to the optimization of photovoltaic panels using MPPT. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/493.

Lokman:2025:WQE

[3746] Amar Lokman, Wan Zakiah Wan Ismail, and Nor Azlina Ab Aziz. Water quality evaluation and analysis by integrating statistical and machine learning approaches. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/494.

Li:2025:HYL

[3747] Liang Li, Yangfei He, Yingying Wei, Hucheng Pu, Xiangge He, Chunlei Li, and Weiliang Zhang. HSS-YOLO lightweight object detection model for intelligent inspection robots in power distribution rooms. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/495.

Mauladdawilah:2025:DFS

[3748] Husein Mauladdawilah, Mohammed Balfaqih, Zain Balfagih, María del Carmen Pegalajar, and Eulalia Jadraque Gago. Deep feature selection of meteorological variables for LSTM-based PV power forecasting in highdimensional time-series data. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/8/496.

Chu:2025:SAF

[3749] Ruobo Chu, Schweitzer Patrick, and Kai Yang. Series arc fault detection method based on time domain imaging and long short-term memory network for residential applications. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/8/497.

Han:2025:QSC

[3750] Longzhe Han, Tian Zeng, Jia Zhao, Xuecai Bao, Guangming Liu, and Yan Liu. Queue stability-constrained deep reinforcement learning algorithms for adaptive transmission control in multi-access edge computing systems. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/498.

Brohi:2025:RLA

[3751] Sarfraz Brohi, Qurat ul-ain Mastoi, N. Z. Jhanjhi, and Thulasyammal Ramiah Pillai. A research landscape of agentic AI and large language models: Applications, challenges and future directions. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/499.

Zeng:2025:ERV

[3752] Weijia Zeng, Binglin Liu, Yi Hu, Weijiang Liu, Yuhe Fu, Yiyue Zhang, and Weiran Zhang. Estimation of residential vacancy rate in underdeveloped areas of China based on Baidu Street View residential exterior images: a case study of Nanning, Guangxi. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/500.

Jayasinghe:2025:HFI

[3753] Udara Jayasinghe, Thanuj Fernando, and Anil Fernando. High-fidelity image transmission in quantum communication with frequency domain multi-qubit techniques. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/501.

Ortiz-Bayliss:2025:HMT

[3754] José Carlos Ortiz-Bayliss, Alonso Vela Morales, and Ivan Amaya. How much is too much? Facing practical limitations in hyper-heuristic design for packing problems. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/502.

Levashova:2025:PFD

[3755] Natalia Levashova, Pavel Levashov, Dmitry Erofeev, and Alla Sidorova. The problem of formation destruction in carbon dioxide storage: a microscopic model. Algorithms (Basel), 18(8), August 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/503.

Botta:2025:SPE

[3756] Marco Botta, Davide Cavagnino, Alessandro Druetto, Maurizio Lucenteforte, and Annunziata Marra. A survey of printable encodings. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/8/504.

Fedkin:2025:AHP

[3757] Yevgeniy Fedkin, Natalya Denissova, Gulzhan Daumova, Ruslan Chettykbayev, and Saule Rakhmetullina. Avalanche hazard prediction in East Kazakhstan using ensemble machine learning algorithms. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/505.

Ma:2025:AMG

[3758] Xinyi Ma, Xu Qin, and Li Lv. Algorithmic modeling of Generation Z’s therapeutic toys consumption behavior in an emotional economy context. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/506.

Anemeta:2025:CCS

[3759] Don Divin Anemeta and Rafal Rzepka. Cross-cultural safety judgments in child environments: a semantic comparison of vision-language models and humans. Algorithms (Basel), 18(8), August 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/507.

Li:2025:HCC

[3760] Junmin Li, Fei Wang, Wenguang Guo, Zhengyong Zhou, Shuaike Miao, and Te Chen. Hierarchical coordination control of distributed drive intelligent vehicle based on TSMPC and tire force optimization allocation. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/8/508.

Gogia:2025:ASH

[3761] Eshaa Gogia, Arash Dehzangi, and Iman Dehzangi. Automated subregional hippocampus segmentation using 3D CNNs: a computational framework for brain aging biomarker analysis. Algorithms (Basel), 18 (8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/509.

Okano:2025:EAI

[3762] Marcelo T. Okano, William Aparecido Celestino Lopes, Sergio Miele Ruggero, Oduvaldo Vendrametto, and João Carlos Lopes Fernandes. Edge AI for industrial visual inspection: YOLOv8-based visual conformity detection using Raspberry Pi. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/8/510.

Li:2025:CCA

[3763] Hui Li, Yunshi Tao, Huan Wang, Zhe Wang, and Qingzheng Liu. CASF: Correlation-alignment and significance-aware fusion for multimodal named entity recognition. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/8/511.

Ruiz-Torrubiano:2025:MLS

[3764] Rubén Ruiz-Torrubiano, Deepak Dhungana, Sarita Paudel, and Himanshu Buckchash. Modeling local search metaheuristics using Markov decision processes. Algorithms (Basel), 18(8), August 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/512.

Yuan:2025:MAI

[3765] Yanhong Yuan, Shuangsheng Duo, Xuming Tong, and Yapeng Wang. A multimodal affective interaction architecture integrating BERT-based semantic understanding and VITS-based emotional speech synthesis. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/8/513.

Lahraoui:2025:NEB

[3766] Younes Lahraoui, Saiida Lazaar, Youssef Amal, and Abderrahmane Nitaj. A novel ECC-based method for secure image encryption. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/514.

Zhang:2025:FBA

[3767] Ao Zhang, Kelei Li, and Haihua Wang. A fusion-based approach with Bayes and DeBERTa for efficient and robust spam detection. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/515.

Cakebread:2025:SST

[3768] Jack Cakebread, Warren G. Jackson, Daniel Karapetyan, Andrew J. Parkes, and Ender Özcan. Student surpasses the teacher: Apprenticeship learning for quadratic unconstrained binary optimisation. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/516.

Qin:2025:HBA

[3769] Panke Qin, Bo Ye, Ya Li, Zhongqi Cai, Zhenlun Gao, Haoran Qi, and Yongjie Ding. Hybrid BiLSTM-ARIMA architecture with whale-driven optimization for financial time series forecasting. Algorithms (Basel), 18 (8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/517.

Wang:2025:AEA

[3770] Chia-Nan Wang and Giovanni Cahilig. Algorithmic efficiency analysis in innovation-driven labor markets: a super-SBM and Malmquist productivity index approach. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/518.

Zayani:2025:MPM

[3771] Marii Zayani, Abdelmalek Toumi, and Ali Khalfallah. MHD-protonet: Margin-aware hard example mining for SAR few-shot learning via dualloss optimization. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/519.

Saad:2025:FLA

[3772] Amani Saad, Andries P. Engelbrecht, and Salman A. Khan. Fitness landscape analysis for the differential evolution algorithm. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/520.

Costa:2025:ASF

[3773] João Costa, Rui Silva, Gonçalo Martins, Jorge Barreiros, and Mateus Mendes. Analysis of the state and fault detection of a plastic injection machine — a machine learning-based approach. Algorithms (Basel), 18 (8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/521.

Devanna:2025:EFD

[3774] Rosa Pia Devanna, Miguel Torres-Torriti, Kamil Sacilik, Necati Cetin, and Fernando Auat Cheein. Exploring the frequency domain point cloud processing for localisation purposes in arboreal environments. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/8/522.

Vladov:2025:IMG

[3775] Serhii Vladov, Lyubomyr Chyrun, Eduard Muzychuk, Victoria Vysotska, Vasyl Lytvyn, Tetiana Rekunenko, and Andriy Basko. Intelligent method for generating criminal community influence risk parameters using neural networks and regional economic analysis. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/523.

Wei:2025:FMF

[3776] Shuting Wei, Mian Jia, and Junyi Duan. FE-MCFN: Fuzzy-enhanced multi-scale cross-modal fusion network for hyperspectral and LiDAR joint data classification. Algorithms (Basel), 18(8), August 2025. CO-DEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.

com/1999-4893/18/8/524.

Duraisingam:2025:ENA

[3777] Aruna Duraisingam, Daniele Soria, and Ramaswamy Palaniappan. Examining the neurophysiology of attentional habituation to repeated presentations of food and non-food visual stimuli. Algorithms (Basel), 18 (8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/8/525.

Montemanni:2025:SMS

[3778] Roberto Montemanni and Derek H. Smith. On solving the minimum spanning tree problem with conflicting edge pairs. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/526.

Saggio:2025:NGA

[3779] Giovanni Saggio, Paolo Roselli, Luca Pietrosanti, Alessandro Romano, Nicola Arangino, Martina Patera, and Antonio Suppa. A new geometric algebra-based classification of hand bradykinesia in Parkinson’s disease measured using a sensory glove. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/8/527.

Chen:2025:DNL

[3780] Qiang Chen, Yuanfeng Luo, Wu Yuan, Ruiliang Zhang, and Yunshou Mao. DGNCA-net: a lightweight and efficient insulator defect detection model. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/528.

Chang:2025:OAM

[3781] Xueli Chang, Yue Wang, Heping Zhang, Bogdan Adamyk, and Lingyu Yan. Optimized adaptive multi-scale architecture for surface defect recognition. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/8/529.

Khuwuthyakorn:2025:BES

[3782] Pattaraporn Khuwuthyakorn, Abdullah Lakhan, Arnab Majumdar, and Orawit Thinnukool. Blockchain-enabled self-autonomous intelligent transport system for drone task workflow in edge cloud networks. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/8/530.

Islam:2025:MPA

[3783] Md Saiful Islam, Tushar Kanti Roy, and Israt Jahan Bushra. Marine predators algorithm-based robust composite controller for enhanced power sharing and real-time voltage stability in DC-AC microgrids. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/8/531.

Pant:2025:IKM

[3784] Yuv Raj Pant, Larry Leigh, and Juliana Fajardo Rueda. Improving Kmeans clustering: a comparative study of parallelized version of modified K-means algorithm for clustering of satellite images. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/8/532.

Lin:2025:HRH

[3785] Yuhuai Lin, Kelei Li, and Haihua Wang. High-resolution human keypoint detection: a unified framework for single and multi-person settings. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/8/533.

Tan:2025:RCM

[3786] Jinhui Tan, Shuai Teng, Zongchao Liu, Wei Mao, and Minghui Chen. Research on commuting mode split model based on dominant transportation distance. Algorithms (Basel), 18(8), August 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/8/534.

Huang:2025:SRS

[3787] Pengcheng Huang, Canyu Chen, and Xinmeng Wan. A strongly robust secret image sharing algorithm based on QR codes. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/535.

Wang:2025:RLH

[3788] Xiaochuan Wang, Na Li, and Xingchen Jin. A reinforcement learning hyper-heuristic with cumulative rewards for dual-peak time-varying network optimization in heterogeneous multi-trip vehicle routing. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/536.

Shi:2025:OOD

[3789] Hongyan Shi, Xiaofeng Bai, and Chenshuai Bai. Optimization of object detection network architecture for high-resolution remote sensing. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/
537.

Huang:2025:HHF

[3790] Shaojie Huang, Tianzhong Wang, Weiquan Liu, Yingchao Piao, Jinhe Su, Guorong Cai, and Huilin Xu. HFGAD: Hierarchical fine-grained attention decoder for gaze estimation. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/538.

Wasana:2025:OCS

[3791] Sanduni Wasana and Upeksha Perera. Optimal choice of the shape parameter for the radial basis functions method in one-dimensional parabolic inverse problems. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/9/539.

Feng:2025:ADS

[3792] Yuming Feng and Violeta Leoreanu-Fotea. Algorithm for determining the strong fuzzy grade of a hypergroup. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/540.

Chai:2025:ASM

[3793] Yanfeng Chai, Jiashu Li, and Qiang Zhang. Accelerating subgraph matching through advanced compression and label filtering. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/541.

Kwakye-Boateng:2025:POS

[3794] Patricia Kwakye-Boateng, Lagouge Tartibu, and Jen Tien-Chien. Performance optimization of a silica gel-water adsorption chiller using grey wolf-based multi-objective algorithms and regression analysis. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/542.

Gaopale:2025:HML

[3795] Kesalopa Gaopale, Takashi Sasaoka, Akihiro Hamanaka, and Hideki Shimada. Hybrid machine learning model for blast-induced peak particle velocity estimation in surface mining: Application of sparrow search algorithm in ANN optimization. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/543.

Chen-Charpentier:2025:MHD

[3796] Benito Chen-Charpentier. Modeling how the different parts of the immune system fight viruses. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/9/544.

Arowolo:2025:ARD

[3797] Micheal Olaolu Arowolo, Sulaiman Olaniyi Abdulsalam, Rafiu Mope Isiaka, Kingsley Theophilus Igulu, Bukola Fatimah Balogun, Mihail Popescu, and Dong Xu. Agentic RAG-driven multi-omics analysis for PI3K/AKT pathway deregulation in precision medicine. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/545.

Puentes-Atencio:2025:MSA

[3798] Andres F. Puentes-Atencio, Daniel Mejia-Parra, Ander Arbelaiz, Carlos Cadavid, and Oscar Ruiz-Salguero. Mixed 1D/2D simplicial approximation of volumetric medial axis by direct palpation of shape diameter function. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/9/546.

Hashemian:2025:LSS

[3799] Saya Hashemian, Zak Khan, Pulkit Kalhan, and Yang Liu. Learning SMILES semantics: Word2Vec and transformer embeddings for molecular property prediction. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/9/547.

Yang:2025:CLM

[3800] Wanqun Yang and Jun Chen. Comparison of linear MPC and explicit MPC for battery cell balancing control. Algorithms (Basel), 18 (9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/548.

Redouane:2025:NHI

[3801] Kheireddine Redouane and Ashkan Jahanbani Ghahfarokhi. A new hybrid intelligent system for predicting bottom-hole pressure in vertical oil wells: a case study. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/9/549.

Sakan:2025:LHF

[3802] Kairat Sakan, Kunbolat Algazy, Nursulu Kapalova, and Andrey Varennikov. Lightweight hash function design for the Internet of Things: Structure and SAT-based cryptanalysis. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/550.

Ordonez:2025:SSA

[3803] Pablo Ordóñez, Ying Xie, Xinyue Zhang, Chloe Yixin Xie, Santiago Acosta, and Issac Guitierrez. SPADE: Superpixel adjacency driven embedding for three-class melanoma segmentation. Algorithms (Basel), 18 (9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/551.

Bommanapally:2025:MTS

[3804] Vidya Bommanapally, Amir Akhavanrezayat, Parvathi Chundi, Quan Dong Nguyen, and Mahadevan Subramaniam. Multi-task semi-supervised approach for counting cones in adaptive optics images. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/552.

Kelmendi:2025:ADO

[3805] Ardian Kelmendi and George Pappas. AI-driven optimization of functional feature placement in automotive CAD. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/553.

Kavargyris:2025:FSG

[3806] Dimitrios Christos Kavargyris, Konstantinos Georgiou, Eleanna Papaioannou, Theodoros Moysiadis, Nikolaos Mittas, and Lefteris Angelis. Future skills in the GenAI era: a labor market classification system using Kolmogorov–Arnold networks and explainable AI. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/554.

Simko:2025:OVC

[3807] Juraj Simko, Michal Prazenica, Roman Konarik, Slavomír Kascák, and Peter Klco. Output voltage control of a synchronous Buck DC/DC converter using artificial neural networks. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/555.

Kabir:2025:REA

[3808] Sami Kabir, Mohammad Shahadat Hossain, and Karl Andersson. A review of explainable artificial intelligence from the perspectives of challenges and opportunities. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/9/556.

Almonte:2025:DDP

[3809] Carol Santos Almonte, Oscar Sanchez Jimenez, Eduardo Souza de Cursi, and Emmanuel Pagnacco. Data-driven prior construction in Hilbert spaces for Bayesian optimization. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/557.

Shalabayeva:2025:MLM

[3810] Laura Shalabayeva, Pilar Bahílo Mateu, Marc Romeu Ferras, Javier Díaz-Carnicero, Alberto Budía, and David Vivas-Consuelo. Machine learning models for predicting postoperative complications and hospitalization after percutaneous nephrolithotomy. Algorithms (Basel), 18 (9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/558.

Wang:2025:RDL

[3811] Longda Wang, Zhang Wu, Yang Liu, and Yan Chen. Research on dualloop ADRC for PMSM based on opposition-based learning hybrid optimization algorithm. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/9/559.

Lancia:2025:SFW

[3812] Giuseppe Lancia and Marcello Dalpasso. Speeding up Floyd–Warshall’s algorithm to compute all-pairs shortest paths and the transitive closure of a graph. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/9/560.

Tentsidou:2025:SGG

[3813] Evmorfia Tentsidou and Haridimos Kondylakis. A SHAP-guided grouped L1 regularization method for CRISPR-Cas9 off-target predictions. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/561.

Tasinaffo:2025:OET

[3814] Paulo M. Tasinaffo, Gildárcio S. Gonçalves, Johnny C. Marques, Luiz A. V. Dias, and Adilson M. da Cunha. An overview of the Euler-type universal numerical integrator (E-TUNI): Applications in non-linear dynamics and predictive control. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/562.

Pascoal:2025:DIS

[3815] David Pascoal, Telmo Adão, Agnieszka Chojka, Nuno Silva, Sandra Rodrigues, Emanuel Peres, and Raul Morais. Democratizing IoT for smart irrigation: a cost-effective DIY solution proposal evaluated in an actinidia orchard. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/9/563.

Escobar-Anel:2025:GAA

[3816] Marcos Escobar-Anel and Yiyao Jiao. The generative adversarial approach: a cautionary tale of finite samples. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/564.

Bargellini:2025:EBE

[3817] Antonio E. Bargellini, Daniele Ritelli, and Giulia Spaletta. Extended von Bertalanffy equation in Solow growth modelling. Algorithms (Basel), 18 (9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/565.

Xu:2025:IAD

[3818] Chuanfang Xu, Chengyu Zhang, Mingxia Xu, Jiaqing Chen, Longda Wang, and Zhaoyu Han. Improved active disturbance rejection speed tracking control for high-speed trains based on SBWO algorithm. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/566.

Marasovic:2025:FSB

[3819] Tea Marasović and Vladan Papić. Few-shot breast cancer diagnosis using a Siamese neural network framework and triplet-based loss. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/567.

Chicharro:2025:STI

[3820] Francisco I. Chicharro, Neus Garrido-Saez, and Julissa H. Jerezano. Secant-type iterative classes for nonlinear equations with multiple roots. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/
568.

Li:2025:RCMb

[3821] Xiaoxu Li, Jixuan Wang, Jianqiang Wang, Jiahao Wang, Jiamin Liu, Jiaming Chen, and Xuelian Yu. Research on CNC machine tool spindle fault diagnosis method based on deep residual shrinkage network with dynamic convolution and selective kernel attention model. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/569.

Anumula:2025:UPR

[3822] Chaitanya Vardhini Anumula, Tanvi Banerjee, and William Lee Romine. Understanding physiological responses for intelligent posture and autonomic response detection using wearable technology. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/570.

Mohammed:2025:EWS

[3823] Hussein S. Mohammed, Hayam K. Mustafa, and Omar A. Abdulkareem. Enhancing wireless sensor networks with Bluetooth low-energy mesh and ant colony optimization algorithm. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/571.

Marevac:2025:MVS

[3824] Elmin Marevac, Esad Kadusić, Natasa Zivić, Nevzudin Buzadija, Edin Tabak, and Safet Velić. Multimodal video summarization using machine learning: a comprehensive benchmark of feature selection and classifier performance. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/9/572.

Skapinyecz:2025:RTO

[3825] Róbert Skapinyecz. Recent trends in the optimization of logistics systems through discrete-event simulation and deep learning. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/573.

Jayasinghe:2025:QFD

[3826] Udara Jayasinghe, Thanuj Fernando, and Anil Fernando. A quantum frequency-domain framework for image transmission with three-qubit error correction. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/9/574.

Wang:2025:SDP

[3827] Dongwei Wang, Sally McClean, Lingkai Yang, Ian McChesney, and Zeeshan Tariq. A smooth-delayed phase-type mixture model for humandriven process duration modeling. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/575.

Liu:2025:OGS

[3828] Zongchao Liu, Shuai Teng, and Shaodi Wang. Optimized GRU with self-attention for bearing fault diagnosis using Bayesian hyperparameter tuning. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/9/576.

Burks:2025:MAA

[3829] Jamison H. Burks, Wendy Hartogensis, Stephan Dilchert, Ashley E. Mason, and Benjamin L. Smarr. Multiscale average absolute difference (MSAAD): a computationally efficient and nonparametric adaptation of line length for noisy, uncontrolled wearables time series. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/577.

Knights:2025:MMN

[3830] Vesna Antoska Knights, Tatjana Atanasova-Pacemska, and Jasenka Gajdos Kljusurić. Mathematical modelling and numerical analysis of turbulence models (in a two-stage laboratory turbine). Algorithms (Basel), 18 (9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/578.

Kovtunenko:2025:PMN

[3831] Victor A. Kovtunenko and Olena M. Atlasiuk. Poroelastic medium with non-penetrating crack driven by hydraulic fracture: FEM approximation using HHT-alpha and semi-smooth Newton methods. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/579.

Grieger:2025:DES

[3832] Niklas Grieger, Siamak Mehrkanoon, and Stephan Bialonski. Dataefficient sleep staging with synthetic time series pretraining. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/580.

Vici:2025:DVW

[3833] Alessandro Vici, Ann Zeuner, and Alessandro Giuliani. Detecting very weak signals: a mixed strategy to deal with biologically relevant information. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/9/581.

Zafar:2025:HSG

[3834] Fiza Zafar, Alicia Cordero, Sadia Mujtaba, and Juan R. Torregrosa. A hybrid Steffensen-genetic algorithm for finding multi-roots of nonlinear equations and applications to biomedical engineering. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/582.

Yelemessov:2025:AOC

[3835] Kassym Yelemessov, Dinara Baskanbayeva, Leyla Sabirova, Nikita V. Martyushev, Boris V. Malozyomov, Tatayeva Zhanar, and Vladimir I. Golik. Algorithmic optimal control of screw compressors for energyefficient operation in smart power systems. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/583.

Liu:2025:PPI

[3836] Juan Liu and Guofeng Mei. PISI: Physical information based solverinteractive network structure reconstruction. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/584.

Morales:2025:PFC

[3837] Marco A. Morales, Dania A. Pérez-Muñoz, J. Alejandro Hernández-González, Miguel Alvarado-Flores, and Sinuhé Ruiz-Salgado. 2D/3D pattern formation comparison using spectral methods to solve nonlinear partial differential equations of condensed and soft matter. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/9/585.

Daovisan:2025:RXI

[3838] Hanvedes Daovisan, Charin Suwanwong, Pitchada Prasittichok, Narulmon Prayai, and Phaktada Choowan. Reclaiming XAI as an innovation in healthcare: Bridging rule-based systems. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/586.

Alanis:2025:EKF

[3839] Alma Y. Alanis. Exploring Kalman filtering applications for enhancing artificial neural network learning. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/587.

Gkologkinas:2025:CAU

[3840] Georgios Dimitrios Gkologkinas, Konstantinos Ntouros, Eftychios Protopapadakis, and Ioannis Rallis. A comparative analysis of U-Net architectures with dimensionality reduction for agricultural crop classification using hyperspectral data. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/9/588.

Wu:2025:IAD

[3841] Jie Wu, Zhengwei Chang, Wei Zhang, Haina Rong, and Tingting Zeng. Importance assessment of distribution network nodes based on an improved MBCC-HITS algorithm. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/9/589.

Zhou:2025:SSM

[3842] Ming Zhou, Hancheng Yu, Zhongchen Li, and Yupu Zhang. Selfsupervised monocular depth estimation based on differential attention. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/9/
590.

Lokman:2025:IFL

[3843] Amar Lokman, Wan Zakiah Wan Ismail, Nor Azlina Ab Aziz, and Anith Khairunnisa Ghazali. Interactive fuzzy logic interface for enhanced real-time water quality index monitoring. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/9/591.

Liu:2025:UMT

[3844] Zhuofu Liu, Lichen Niu, Zhixin Di, and Meimei Liu. UCINet: a multitask network for umbilical coiling index measurement in obstetric ultrasound. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/9/592.

Luque-Hernandez:2025:CEC

[3845] Francisco Javier Luque-Hernández, Sergio Aquino-Britez, Josefa Díaz-Álvarez, and Pablo García-Sánchez. A comparison of energy consumption and quality of solutions in evolutionary algorithms. Algorithms (Basel), 18(9), September 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/9/593.

Al-dabbagh:2025:ALD

[3846] Bashar Sami Nayyef Al-dabbagh, Agapito Ledezma Espino, and Araceli Sanchis de Miguel. An active learning and deep attention framework for robust driver emotion recognition. Algorithms (Basel), 18(10): ??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/669.

Al-Qazzaz:2025:IMO

[3847] Noor Kamal Al-Qazzaz, Sawal Hamid Bin Mohd Ali, and Siti Anom Ahmad. Investigation of the MobileNetV2 optimal feature extraction layer for EEG-based dementia severity classification: a comparative study. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/620.

Almasri:2025:EGG

[3848] Wesam Almasri, Marwa Saad, and Changiz Mohiyeddini. Evaluating Google Gemini’s capability to generate NBME-standard pharmacology questions using a 16-criterion NBME rubric. Algorithms (Basel), 18(10): ??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/612.

AlZubi:2025:NNA

[3849] Shadi AlZu’bi, Fatima Quiam, Ala’ M. Al-Zoubi, and Muder Almiani. Neural network architectures for secure and sustainable data processing in e-government systems. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/601.

Angelov:2025:HAB

[3850] Anton Angelov and Milena Lazarova. Hybrid artificial bee colony algorithm for test case generation and optimization. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/668.

Arbab:2025:IEI

[3851] Erfan Arbab, Kalpdrum Passi, and Robert Ohle. Impact of an educational intervention on BPPV management: an interrupted time series analysis. Algorithms (Basel), 18(10):??, October 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/606.

Bettahi:2025:MEM

[3852] Abdelkarim Bettahi, Fatima-Zahra Belouadha, and Hamid Harroud. A modular and explainable machine learning pipeline for student dropout prediction in higher education. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/662.

Bouyukliev:2025:SPA

[3853] Iliya Bouyukliev, Dushan Bikov, and Maria Pashinska-Gadzheva. Strategies for parallelization of algorithms for integer partition. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/655.

Brand:2025:COS

[3854] Jonathan Brand, Ryan Kochis, Vasav Shah, and Wentai Liu. Characterizing and optimizing spatial selectivity of peripheral nerve stimulation montages and electrode configurations in silico. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/635.

Bungau:2025:BMS

[3855] Gabriela S. Bungau, Andrei-Flavius Radu, Ada Radu, Delia Mirela Tit, and Paul Andrei Negru. Bibliometric mapping of soil chemicalization and fertilizer research: Environmental and computational insights. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/660.

Byun:2025:VFD

[3856] Subin Byun and Jeongil Seo. Visual feature domain audio coding for anomaly sound detection application. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/646.

Cabrera-Ibarra:2025:GBA

[3857] Hugo Cabrera-Ibarra, David Hernández-Granados, and Lina Riego-Ruiz. A graph-based algorithm for detecting long non-coding RNAs through RNA secondary structure analysis. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/652.

Calude:2025:SSI

[3858] Cristian S. Calude, Patrick Gladding, Alec Henderson, and Nikola Kasabov. SAIN: Search-And-INfer, a mathematical and computational framework for personalised multimodal data modelling with applications in healthcare. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/605.

Cedeno-Munoz:2025:TMT

[3859] Joel Cedeño-Muñoz, Cristian Zambrano-Vega, and Antonio J. Nebro. TMP-M2Align: a topology-aware multiobjective approach to the multiple sequence alignment of transmembrane proteins. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (elec-

tronic). URL https://www.mdpi.com/1999-4893/18/10/640.

Chauhan:2025:EAR

[3860] Sumika Chauhan, Govind Vashishtha, and Prabhkiran Kaur. An effective approach to rotatory fault diagnosis combining CEEMDAN and featurelevel integration. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/644.

Cifuentes:2025:DMD

[3861] Lina María Yara Cifuentes, Ernesto Cadena Muñoz, and Rafael Cubillos Sánchez. Development of a model for detecting spectrum sensing data falsification attack in mobile cognitive radio networks integrating artificial intelligence techniques. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/596.

Coquelet:2025:EPI

[3862] Boris Coquelet, Gilles Dejaegere, and Yves De Smet. Evaluating the Promethee II ranking quality. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/597.

daSilva:2025:MLR

[3863] Abner Fernandes Souza da Silva, Virginia Aparecida da Silva Moris, João Eduardo Azevedo Ramos da Silva, Murilo Aparecido Voltarelli, and Tiago F. A. C. Sigahi. Machine learning in reverse logistics: a systematic literature review. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/650.

Del-Valle-Soto:2025:REE

[3864] Carolina Del-Valle-Soto, José A. Del-Puerto-Flores, Leonardo J. Valdivia, Aimé Lay-Ekuakille, and Paolo Visconti. A resilient energyefficient framework for jamming mitigation in cluster-based wireless sensor networks. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/614.

deOliveira:2025:ECR

[3865] Nazário Augusto de Oliveira and Leonardo Fernando Cruz Basso. Explaining corporate ratings transitions and defaults through machine learning. Algorithms (Basel), 18(10):??, October 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/608.

Feng:2025:IGE

[3866] Xu Feng, Hongzhou Chai, Jinkai Feng, and Yunlong Wu. An information granulation-enhanced kernel principal component analysis method for detecting anomalies in time series. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/658.

Filgueira:2025:BDE

[3867] Jade Filgueira, Thiago Antonini Alves, Clodomir Santana, Attilio Converti, Carmelo J. A. Bastos-Filho, and Hugo Siqueira. Binary differential evolution with a limited maximum number of dimension changes. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-

4893 (electronic). URL https://www.mdpi.com/1999-4893/18/10/621.

Garagnani:2025:DMT

[3868] Filippo Garagnani and Vittorio Maniezzo. Deconstructing a minimalist transformer architecture for univariate time series forecasting. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/645.

Guo:2025:CPB

[3869] Lei Guo, Hongye Li, Min Pang, Kaowei Liu, Xie Han, and Fengguang Xiong. Cosine prompt-based class incremental semantic segmentation for point clouds. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/648.

Gutierrez-Sanchez:2025:BBS

[3870] Gresky Gutiérrez-Sánchez, María Purificación Vicente-Galindo, and Purificación Galindo-Villardón. Bootstrap-based stabilization of sparse solutions in tensor models: Theory, assessment, and application. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/602.

Hlupic:2025:TMM

[3871] Nikica Hlupić, Danko Basch, Edouard Ivanjko, and Martin Gregurić. Two modifications of MinSum algorithm for efficient system-optimal traffic assignment. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/609.

Huang:2025:RSG

[3872] Xiaoyao Huang, Shuo Quan, and Jie Wu. A recursive solution to the global maximum minimum cut problem with a fixed sink. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/665.

Ifrim:2025:HNS

[3873] Ana-Maria Ifrim and Ionica Oncioiu. A hybrid numerical-semantic clustering algorithm based on scalarized optimization. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/607.

Jiang:2025:BSP

[3874] Kun Jiang and Zhihong Fu. Backward signal propagation: a symmetrybased training method for neural networks. Algorithms (Basel), 18(10): ??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/594.

Jiang:2025:MOA

[3875] Fan Jiang, Zhiwei Ma, Zheng Gong, and Jumei Zhou. MemristiveAdamW: an optimization algorithm for spiking neural networks incorporating memristive effects. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/618.

Jing:2025:FLI

[3876] Yurong Jing, Zhiyong Tao, and Sen Lin. FocusNet: a lightweight insulator defect detection network via first-order Taylor importance assessment and knowledge distillation. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/649.

Kapusta:2025:COG

[3877] Jan Kapusta, Waldemar Bauer, and Jerzy Baranowski. Contextual object grouping (COG): a specialized framework for dynamic symbol interpretation in technical security diagrams. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/642.

Karras:2025:MMT

[3878] Christos Karras, Leonidas Theodorakopoulos, Aristeidis Karras, George A. Krimpas, Charalampos-Panagiotis Bakalis, and Alexandra Theodoropoulou. MCMC methods: From theory to distributed Hamiltonian Monte Carlo over PySpark. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/661.

Koltunov:2025:MCW

[3879] Nikita Koltunov, Egor Guguchkin, Oleg Samovarov, Liudmila Mikhailova, and Evgeny Karpulevich. A method for calculating whole-genome sequencing outcomes from trio data. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/610.

Konidaris:2025:IBR

[3880] Dimitrios Konidaris, Polyzois Soumplis, Andreas Varvarigos, and Panagiotis Kokkinos. Intent-based resource allocation in edge and cloud computing using reinforcement learning. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/627.

Kostopoulos:2025:ESS

[3881] Georgios Kostopoulos, Nikos Fazakis, Sotiris Kotsiantis, and Yiannis Dimakopoulos. Enhancing semi-supervised learning in educational data mining through synthetic data generation using tabular variational autoencoder. Algorithms (Basel), 18(10):??, October 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/663.

Kozelkov:2025:AFV

[3882] Andrey Kozelkov, Andrey Kurkin, Andrey Puzan, Vadim Kurulin, Natalya Tarasova, and Vitaliy Gerasimov. Application features of a VOF method for simulating boiling and condensation processes. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/604.

Laurent:2025:ESA

[3883] Matthew Larnet Laurent, George Edward Marquis, Maria Gonzalez, Ibrahim Tansel, and Sabri Tosunoglu. Embedded sensing in additive manufacturing metal and polymer parts: a comparative study of integration techniques and structural health monitoring performance. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/613.

Li:2025:GGM

[3884] Guangpeng Li, Li Li, and Guoyong Cai. A GIN-guided multiobjective evolutionary algorithm for robustness optimization of complex networks. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/666.

Li:2025:GMG

[3885] Xueping Li, Yanbo Li, Yuhang Li, and Yuan Yang. GMHCA-MCBILSTM: a gated multi-head cross-modal attention-based network for emotion recognition using multi-physiological signals. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/664.

Li:2025:QSH

[3886] Rex Li and Chunyu Liu. Quantifying statistical heterogeneity and reproducibility in cooperative multi-agent reinforcement learning: a metaanalysis of the SMAC benchmark. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/653.

Li:2025:RWS

[3887] Zhen Li, Luhong Wang, Lingzhong Meng, and Guang Yang. Research on warship system resilience based on intelligent recovery with improved ant colony optimization. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/626.

Li:2025:TBF

[3888] Yi Li, Xingzhou Deng, Ang Yang, and Jing Gao. A transformer-based framework for DDoS attack detection via temporal dependency and behavioral pattern modeling. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/628.

Lizarraga:2025:PHA

[3889] Jorge A. Lizarraga, Alberto J. Pita, Javier Ruiz-Leon, Alma Y. Alanis, Luis F. Luque-Vega, Rocío Carrasco-Navarro, Carlos Lara-Álvarez, Yehoshua Aguilar-Molina, and Héctor A. Guerrero-Osuna. Phymastichus-hypothenemus algorithm for minimizing and determining the number of pinned nodes in pinning control of complex networks. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/637.

Lo:2025:OPE

[3890] Wai Lun Lo, Henry Shu Hung Chung, Richard Tai Chiu Hsung, Hong Fu, Tony Yulin Zhu, Tak Wai Shen, and Harris Sik Ho Tsang. Optimal parameter estimation for solar PV panel based on ANN and adaptive particle swarm optimization. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/598.

Luo:2025:OHD

[3891] Shucheng Luo, Xiangbin Meng, Xinfu Pang, Haibo Li, and Zedong Zheng. Optimized hybrid deep learning framework for short-term power load interval forecasting via improved crowned crested porcupine optimization and feature mode decomposition. Algorithms (Basel), 18(10): ??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/659.

Marino:2025:AEP

[3892] Estefanía Avilés Mariño and Antonio Sarasa Cabezuelo. AI-enhanced PBL and experiential learning for communication and career readiness: an engineering pilot course. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/634.

Matei:2025:DTR

[3893] Ioana Matei and Victor-Valeriu Patriciu. A decision tree regression algorithm for real-time trust evaluation of battlefield IoT devices. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/641.

Mireles-Preciado:2025:HSC

[3894] Carlos Gabriel Mireles-Preciado, Diana Carolina Toledo-Pérez, Roberto Augusto Gómez-Loenzo, Marcos Aviles, and Juvenal Rodríguez-Reséndiz. Hardware-software co-design architecture for real-time EMG feature processing in FPGA-based prosthetic systems. Algorithms (Basel), 18(10): ??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/617.

Mitsos:2025:AAP

[3895] Dimitrios Mitsos and Vassilis Poulopoulos. Addressing air pollution challenges: an integrated algorithmic approach towards safeguarding built heritage. Algorithms (Basel), 18(10):??, October 2025. CODEN AL-GOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/619.

Monovasilis:2025:TDR

[3896] Theodoros Monovasilis and Zacharoula Kalogiratou. Two-derivative Runge–Kutta methods with frequency dependent coefficients for solving orbital or oscillatory problems. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/603.

Moutassem:2025:PEC

[3897] Zaki Moutassem, Doha Bounaim, and Gang Li. A physics-enhanced CNN-LSTM predictive condition monitoring method for underground power cable infrastructure. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/600.

Musazade:2025:EIA

[3898] Nurlan Musazade, József Mezei, and Xiaolu Wang. Exploring the impact of AI tools on cognitive skills: a comparative analysis. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/631.

Nanni:2025:DEM

[3899] Loris Nanni and Sheryl Brahnam. Deep ensembles and multisensor data for global LCZ mapping: Insights from So2Sat LCZ42. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/657.

Naskinova:2025:HSM

[3900] Irina Naskinova, Mikhail Kolev, Dilyana Karova, and Mariyan Milev. Hybrid stochastic-machine learning framework for postprandial glucose prediction in type 1 diabetes. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/623.

Panteleev:2025:DLP

[3901] Ivan Panteleev, Mikhail Semin, Evgenii Grishin, Denis Kormshchikov, Anastasiya Iziumova, Mikhail Verezhak, Lev Levin, and Oleg Plekhov. Deep learning prediction of exhaust mass flow and CO emissions for underground mining application. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/630.

Pasupuleti:2025:FSM

[3902] Satya Dev Pasupuleti and Simone A. Ludwig. Feature selection method based on simultaneous perturbation stochastic approximation technique evaluated on cancer genome data classification. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/622.

Pereira:2025:ABD

[3903] Sílvia de Castro Pereira, Eduardo J. Solteiro Pires, and Paulo B. de Moura Oliveira. The ACO-BmTSP to distribute meals among the elderly. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/10/667.

Polo:2025:THM

[3904] Aura Polo, Nelson Cárdenas-Bolaño, Lácides Antonio Ripoll Solano, Lely A. Luengas-Contreras, and Carlos Robles-Algarín. TWISS: a hybrid multi-criteria and wrapper-based feature selection method for EMG pattern recognition in prosthetic applications. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/633.

Ribeiro:2025:ORM

[3905] Bruno Ribeiro, José Baptista, and Adelaide Cerveira. Optimizing renewable microgrid performance through hydrogen storage integration. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN

1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/18/
10/656.

Santos:2025:AID

[3906] Luís F. F. M. Santos, Miguel Ángel Sánchez-Tena, Cristina Alvarez-Peregrina, and Clara Martinez-Perez. Artificial intelligence-driven diagnostics in eye care: a random forest approach for data classification and predictive modeling. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/647.

Shafique:2025:EMB

[3907] Shehzaib Shafique, Gian Luca Bailo, Monica Gori, Giulio Sciortino, and Alessio Del Bue. Enhancing mobility for the blind: an AI-powered bus route recognition system. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/616.

Shussett:2025:AAD

[3908] Daniel Shussett and Veljko Dubljević. Applying the agent-deed-consequence (ADC) model to smart city ethics. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/625.

Sisic:2025:DLB

[3909] Nedim Sisić and Peter Rogelj. Deep learning for brain MRI tissue and structure segmentation: a comprehensive review. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/636.

Tang:2025:IST

[3910] Liqing Tang, Shuxin Wang, Jintian Ji, Siyuan Yin, Robail Yasrab, and Chao Zhou. Integrating structured time-series modeling and ensemble learning for strategic performance forecasting. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/611.

Tarapiah:2025:EEL

[3911] Saed Tarapiah, Linda Abbas, Oula Mardawi, Shadi Atalla, Yassine Himeur, and Wathiq Mansoor. Evaluating the effectiveness of large language models (LLMs) versus machine learning (ML) in identifying and detecting phishing email attempts. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL

https://www.mdpi.com/1999-4893/18/10/599.

Vaidya:2025:EML

[3912] Soham Sanjay Vaidya, Raja Hashim Ali, Shan Faiz, Iftikhar Ahmed, and Talha Ali Khan. Evaluating machine learning techniques for brain tumor detection with emphasis on few-shot learning using MAML. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/624.

Wang:2025:PWF

[3913] Jue Wang, Xinyi Xu, Changxin Ye, and Wei Huangfu. Point-wise fullfield physics neural mapping framework via boundary geometry constrained for large thermoplastic deformation. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/651.

Wosiak:2025:ITW

[3914] Agnieszka Wosiak, Michał Sumiński, and Katarzyna Zykwińska. Impact of temporal window shift on EEG-based machine learning models for cognitive fatigue detection. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/629.

Wu:2025:TAF

[3915] Weitao Wu, Zengwen Zhang, Zhong Xiang, and Miao Qian. Textureadaptive fabric defect detection via dynamic subspace feature extraction and luminance reconstruction. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://

www.mdpi.com/1999-4893/18/10/638.

Xing:2025:EOE

[3916] Ruiming Xing, Baihua Li, Shirin Dora, Michael Whittaker, and Janette Mathie. Engineering-oriented explainable machine learning and digital twin framework for sustainable dairy production and environmental impact optimisation. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/670.

Xu:2025:TGR

[3917] Menghui Xu, Yang Yan, Qiuyan Wang, Hanning Chen, and Zhao Zhang. TRed-GNN: a robust graph neural network with task-relevant edge disentanglement and reverse process mechanism. Algorithms (Basel), 18 (10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic).

URL https://www.mdpi.com/1999-4893/18/10/632.

Yang:2025:OUS

[3918] Xiaodong Yang, Yue Chang, Lun Cheng, Yujing Su, and Tao Wang. Optimal unit scheduling considering multi-scenario source-load uncertainty and frequency security. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/595.

Yedilkhan:2025:OHA

[3919] Didar Yedilkhan, Arman Zhalgasbayev, Sabina Saleshova, and Nursultan Khaimuldin. OIKAN: a hybrid AI framework combining symbolic inference and deep learning for interpretable information retrieval models. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH.

ISSN 1999-4893 (electronic). URL https://www.mdpi.com/1999-4893/
18/10/639.

Yuan:2025:GBS

[3920] Jiheng Yuan and Jian-Yu Li. A Gumbel-based selection data-driven evolutionary algorithm and its application to Chinese text-based cheating official accounts mining. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.

mdpi.com/1999-4893/18/10/643.

Zhao:2025:ARE

[3921] Duo Zhao, Qichao Tang, Lei Ma, Yongkui Sun, and Jieyu Lei. An adaptive rapidly-exploring random trees algorithm based on cross-entropy optimization. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893 (electronic). URL https://www.mdpi.com/

1999-4893/18/10/615.

Zhao:2025:HRI

[3922] Yiying Zhao and Baohua Jin. HIPACO: an RSSI indoor positioning algorithm based on improved ant colony optimization algorithm. Algorithms (Basel), 18(10):??, October 2025. CODEN ALGOCH. ISSN 1999-4893

(electronic). URL https://www.mdpi.com/1999-4893/18/10/654.